A Bibliography of Publications of the USENIX Association: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

03 February 2024 Version 3.20

Title word cross-reference

#1 [Dar93b]. #15 [Nys99]. #ifdef [SC92].

24 × 7 [Sco97]. 3 [Ell97]. < [Sun97]. > [Sun97]. $HOME [Uhl91, ZD93b].

*BSD [Den99a].

-or- [Bor92, Det91, Lib96a, MH92, RL99, VZ91]. -request [Cha92b].

.rhosts [Cra94b].

/tmp [OT90].

1 [ASWW91, BCF^{+}91, BM91b, HHW92, LPLF91, Mit91, ZRB^{+}93]. 1-5 [Ano93a]. 1.2 [For98, GMPS97]. 1/2 [Pik91a]. 10 [Far96e]. 10base [dL95]. 10base-T [dL95]. 12 [Shi98]. 13th [Ano99b]. 1990’s [Hae91a]. 1998 [You97b]. 1999 [Den99c]. 1st [Ano98e, Ano98f, Ano98g, Ano99x].

2 [PAO93]. 2.0 [Rob99]. 2.X [Pom94f]. 2000 [Cum99a, Lie99]. 20th [Cla92]. 21st [Sch96i]. 2d [Han97]. 2K [Ano99w]. 2nd [Ano98h, Ano98i, Ano98j, Ano98k, Ano98z, Ano99f, Ano99u, Ano99g, Jen98, Kno97, Pot97, Quo91, USE98a, USE99a, USE99b, USE99c, Uma97b, Ant95].

3 [PJS^{+}94, Vix92]. 3-ring [GHN91]. 3.0 [BDMVL93, FL94, FGB91, JCS^{+}91, LHF^{+}93, MSS98, NYM92, NN94b, Pat93, RMGB91, WS96, Wie92]. 3.5 [Lei98].

30-Amp [Eva96d]. 32 [Car98b, CH97]. 32-Bit [CH97]. 386 [Kat94]. 3D [FEBM97]. 3DFS [Roo92]. 3rd [Ann97, Ano97a, Ano98a, Ano98l, Ano98m, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99-31, Bar94d, USE98b, USE99d]. 3wish [Ull97].

4 [DG94, RvIG98]. 4.0 [CBB^{+}91, Gee90]. 4.3BSD [BL90c, Mac91]. 4.4 [AMP^{+}96]. 4.4BSD [Bos93, McK95, PM95a, SP95]. 4.4BSD-Lite [PM95a]. 4GL [Sch97b]. 4GL-database [Sch97b]. 4th [Ano98-27, Ano99m, Ano99n, Uma96c].

5 [DW98b, LIN^{+}93]. 5.0 [Bru97, Bru98, Cox97d, Cum99b, WT99]. 5.005 [Hal98d]. 5000 [KP93a]. 5th [Ano98n, Ano99s, Ano99t, Ano99-29].

6.0 [Gra99a]. 64 [YBA93]. 64-bit [MWO^{+}93, YBA93]. 6bone [Fin97a]. 6th [Ano97b, Ano98o, Ano99-28].

7th [Ano97c, Ano98-28, Ano99v, Ano99o, Rub98].

8-1 [Pik91a]. 8-1/2 [Pik91a]. 80/20 [Dar98a, Dar98d]. 8th [Ano99-32].

9 [Pik91a, PPD^{+}95, Pre92, PW93, Wel94]. 9-12 [USE99l]. ’90s [Coo92, Fle96]. ’91 [Cha92a]. ’92 [And93]. ’93 [Num94a, Ros93, SCP^{+}93, Jac93]. ’94 [Ano94f, CPK94, Num94b, Old94b, WP94, Rot95]. ’95 [Ben95, Sch95e, LBSL97]. 95/NT [LBSL97]. ’96 [Ano96d, Var96, Var97, DSS96, Har96b, USE96f]. ’97 [Ano98y, Ano97g, Ano97p]. ’98 [Ano99-35, vW98, CRH99, WWDF98, Ano98c, Ano98d, Ano98r]. ’99 [Ano98u, Ano98i, Ano98j, Ano99r, Ano99s, Ano99t, Ano99b, Ano99-31, Par99b, USE99a, USE99e, USE99g, USE99f, USE99h, USE99c, USE99j, Mil99h, USE99l, Ano98t]. 9th [Ano99p, Ano99q].

= [DeW91].

A&M [SSH96]. ABCs [Gun98a]. Abilities [Zwi91a]. Abstract [Cre97, FO96, Ken91, Mur92, WAKS97, ATS94]. Abstraction [EL94, ZJ90]. Abstractions [CMS93, Cou96a, Kel99, SM92]. Abstractor [GC90]. Abuses [Vix98]. Academic [DBC^{+}99, SSH93a]. Academics [PBEDT97]. Access [ALO90, AGF90, AR94, Byr93, CLS92, DMF97, EW90, GK99, GM99, HJ92, Hed98, Hic99, Hil96, IL90, IL93, KS99, LeF92, Lea98, Lee98, MK95, Mol95, NB97, PB97, Pol93, Roy93, RW93, Spi99, STM^{+}99, Str90, Sug92, Sum94, TA92, TJM^{+}99, Ven92, WCO^{+}90]. Accesses [BD96, Ede93, Kel99]. Accommodate [MA90]. Accommodating [NWEF92]. Account [CKDS90, Cur91, HG96, RDM95, Spe96, Spe97b]. Accounting [Sim91c]. Accounts [Arn98, Cox98c]. Accountworks [Arn98]. Accurate [Bla90]. ACE [Sch99]. Achievement [Ano95b, Ano99a]. Achieving [FC96]. ACID [Win94b]. ACL [Cox97b]. ACLs [BS98, Str90]. ACM97 [Sal97b]. ACMAINT [CKDS90, Cur91]. Acme [Pik94]. Acquisitions [Lea99]. Across [Dar98a, Kah95, MWCR90, MWCR91, ACO92]. Act [Mot91]. Action [Gla96, WH97]. Actions [Ano95d, KL96, Ros95b]. Activation [WWW95]. Activations [BDMVL93]. Active [CGLR92, FSM^{+}91, Jon95, Leh98, SUZ98, Ten96, VDAA99, Cox97a]. ActiveX [WN98]. Activities [Hae90e, Mil95, NKM96, Sto93e, Sto93c, Sto93d, Sto93f, Sto94b, Sto94a, Sto94c, Wal92d, Wal92e, Wal92f, Wal92g, Wal93c, Wal93d]. Activity [Hae90a, Hae90b, Hae90c, Hae90d, Hae91b, Hae91c, Man91, SW91a, Wal91c, Wal91d, Wal91e, Wal92h, Wal92i]. Actor [IG99]. Ad [KCVP95, Fuc97]. Ad-Hoc [KCVP95]. Ada [Gar90b]. ADAPT [RH98a]. Adaptable [GLW95, PM99]. Adaptation [NPS95a, NPS95b, ZS99]. Adaptations [SB98b]. Adapter [RM97, TC95]. Adaptive [AS93, BL97, BS97b, CL98, CK96b, CPM^{+}98, DBCF95, DKB95a, DKB95b, MP90]. Added [Rei92b]. Addiction [Kol93a]. Adding [BDMVL93, BD92, DSS90, Fly99f, Maf95, Sel97a, Shi91, Swa97a, Swa97b]. Additional [Shi91, Sto98e]. Address [Bla93, Che98b, FLB98, KW99, KC96, Leo91, Mor93, MWO^{+}93, Rit90, SM98a, YBA93, YHK96]. Addressed [CMS90]. Addressing [CJ95]. ADL [Vis96]. ADL/C [Vis96]. Admin [dL95]. Administering [Byr93, Car99b, Chr97a, MFM95]. Administration [Abb94, All96, Ano97g, Ano98q, Ano98-28, Ano98y, Ano98-29, Ano99j, Ano99k, Ano99u, Ano99b, Ano99x, Ano99-35, Arc93, Ben93, BGMR91, Bla93, CWT^{+}97, Cas95, Var96, CJ92, Cou96a, Cuc90, Dar93b, DSS96, Eva95a, EG97, Fin96, Fis96c, GSS99b, Gré93, Grg93, Har94, HCH95, Hug96, HW93, Kap94, Kub93, Lib90b, Lib91b, McD92, McN93, MM96a, Nat96, Par92, Pie96, Pla97, Pla99, Pom94g, Sat90, Sat93e, Sat94, Sca97a, Sch92b, Sel97e, SO99, Sim92c, Sim93b, Spe96, Spe97b, Sto93b, TSO^{+}96, Tho98a, Tom96, Tro96, USE91f, USE92e, USE93f, USE94h, USE94m, USE95h, USE96j, USE97a, USE97c, USE98e, USE98g, USE99g, USE99b, USE99i, Ver97, YL98, ZSD90, Zwi93e, Zwi93c, Zwi94e, Zwi94g, Zwi94f, Zwi94h, Ano98e, Ano99-33, Jen98]. Administrator [Bar94b, Gir99b, GC98, Hal99d, Kni96, Kol94j, Kolxx, Mar97, MSN93, Mor92, ML92, RL99, USE90f, Zie93, ZK94, Zwi97]. Administrators [Ano99v, Dar93a, Fis96b, JS93, Sno96, USE93e]. Admins [May95, Sha92b]. Adrift [DG95, DGT96]. Adult [Kle98a, Kle98b]. Advanced [Cal93, Col92a, DG96, DSS96, Mol95, Sch95e]. Advancement [Rob99]. Advances [BH95]. Advancing [Lam91b]. Advantage [Lea95a]. Advantages [KSU92]. Adventure [GBK^{+}95]. Adventures [SCD94]. Adverse [RL99]. Advocate [Lim97]. Aerospace [RA98]. Affinity [DM92]. Affordable [ACTW96]. AFRAID [SW96]. AFS [DW98a, GBK^{+}95, HHS92, HH93, Lam90, Lam91a, Sto92, Sto93a]. AFS-client [DW98a]. After [BS97a, Pet93, Sat93a]. Again [MRK97, Cuc96, Too98]. Against [Car91a, Car93a, Jon95, Kle99b, SP99b, Wal93a]. Age [Bar94a]. Aged [Dar99d, Kol93a]. Agenda [Ana99]. Agent [Ano99r, AD93, Ben94a, CSL^{+}97, EFJ96, GNMR97, Gra96, NB97, Gra96]. Agent-based [EFJ96]. Agent-Mediated [AD93]. Agents [MLC98, Var95a, ZS99, Ano99r]. Ago [Sal97i, Sal98h, Sal98i, Sal98j, Sal98k, Sal98l, Sal98m, Sal99a, Sal99b, Sal99c, Sal99d, Sal99e, She93a, She93b, She93c, She94a, She94b]. Agora [GS96a]. Agreement [Ano94a]. agrep [WM92]. AGUS [RDM95]. Ahead [PTB^{+}97, Win94a]. AHS [AG97]. Aid [Muf99]. Aided [CSYG94]. ain’t [MH92]. Air [Rei97]. AirJava [Mil99k]. AIX [CGHK95, Tal95]. AKA [SSH96]. Alabama [IEE90]. Alamos [CJC^{+}92]. Alex [Cat92]. Algebraic [Ken96]. Algorithm [BK91, GS99b, Owe97, Win94a]. Algorithms [BHS95, BS93b, CI97, CS90, KTP^{+}96]. Alias [Dam93]. Aliases [Det91]. Alleviating [LB97]. Alliance [MS95]. Allocation [BH99, BL90b, MS93a, SS96b]. Allocator [Bon94, SGG^{+}99]. Allocators [Loe96]. Almost [Rob92]. Along [Mil99f]. Alpha [Rey92, CH97, CGLR97, GL97]. Alpha/NT [CGLR97, GL97]. AltaVista [Gri97]. Alternate [Cho98]. Alternative [AMP98, Den92, Wal97b, Wal98, Pat93]. Alternatives [DMST95, DCK^{+}94]. Always [Kol96i, Zwi92c, Wal95a]. Am [Zwi97]. Amanda [dSG93]. AMD [Ste93b]. Amoeba [DOKT91, van92]. Among [AMML96, CBE93]. Amortized [JY96]. Amortized-Signature [JY96]. Amp [Eva96d]. Anagrams [Var95b]. Anaheim [USE90g, USE97e, Ano97n, Far97g]. Analogies [Cha94, Zwi93b]. Analysis [Ber99b, Car94, ES95, Eva97, GMNR99, Hal99c, Jae95, JMM^{+}99, KP93a, KPR94, LKHA94, MS96b, MDS99, MK93, MSS98, MZA96, MM92, SO92b, SBRH96, Ste93a, WS96, WVT94, WVS^{+}99, kc99a]. Analyst [Ken90]. Analytical [LD97]. Analyze [VC92]. Analyzed [RW91]. Analyzing [Sin98d]. Anarchistic [SSH96]. Anatomy [Bis92b, BR98, Lor95, SM98b]. Andrew [Wat96, Kol96d, Wat96]. Anecdote [Koe94]. Animated [McC98g]. Animation [BK91, Ell97, Rao99b]. Ann [USE92h, USE95i]. Annotated [Spe93b]. Annotation [CLL97, GL99]. Announcement [Ano97b, Ano97a, Ano97c, Ano98o, Ano98b, Ano98h, Ano98c, Ano98e, Ano98n, Ano98d, Ano98a, Ano98p, Ano98u, Ano98f, Ano98i, Ano98r, Ano98q, Ano98s, Ano98t, Ano98l, Ano98v, Ano98j, Ano98k, Ano98m, Ano98g, Ano98w, Ano99h, Ano99i, Ano99r, Ano99s, Ano99t, Ano99f, Ano99j, Ano99k, Ano99u, Ano99v, Ano99b, Ano99m, Ano99w, Ano99c, Ano99d, Ano99e, Ano99x, Ano99g, Ano99p, Ano99q, Ano99o, Ano99l, Ano99n, Ano99y, Ano99z]. Announcements [KG93, Sal94d, Sat93c]. Annual [Ano97b, Ano98o, Ano98u, Ano98v, Ano98-28, Ano98-32, Ano98-38, Ano99h, Ano99i, Ano99v, Ano99c, Ano99d, Ano99e, Ano99-28, Ano99-52, Ano99-55, CWT^{+}97, Den99a, Fis97, Ros96i, USE94m, USE99m, USE96e, USE96i, USE96k, USE97d, USE97e, USE98d, USE98f, USE98h]. Anomaly [BBK99, GS99a]. Anonymizing [Mar98].

Anonymous

[CHTY96, CdR99, Pom99, Rot90, YBA93]. ANSI [Laj96, Sim94a]. Answer [Chr90a, Chr91, NN92a, Sch92b]. Answering [Cha92b]. Anti [CV93, Sze99, WCO^{+}90]. Anti-Spam [Sze99]. Anti-Tampering [CV93]. Anti-Viral [WCO^{+}90]. Antidotes [Lea97]. Antigone [MPH99]. Antonio [USE92g, USE98i, Wil98]. Any [MB96]. Anything [Hae92, Rob92]. Apache [Don99, Koh99, Mee98]. Apertos [iIYL95]. API [PBS98]. Apparent [Zwi91a]. Appelman [Dar97c, Kol96e, Mil97c]. Applet [Sch96e]. Appletalk [Ran97]. Applets [BCE98, Lea97, Nag96, Rao99b]. Appliance [DBD^{+}99, HLM94]. Application [AT92, AAL92, Bar98, Bel96a, BG97, Bre93, BCLC99, CFL94a, CFL94b, Dea90, DFG^{+}96, Eng97, Fin97c, FB96, GDFR90, Gue92, HO94, HK99, IMC93, JD94, KGMP^{+}96, KSKD99, KCCT94, LR94, Lar94, Lin90a, MO99, NN93a, NPS95a, NPS95b, PG92, Pol97, RMGB91, RBJ91, Ros96h, Sat95, SC99, SW90, TW93, WH92, Wal97b, Wal98, ZD93a, ZPMD97, Bir94]. Application- [FB96]. Application-Aware [NPS95a, NPS95b]. Application-Controlled [CFL94a, CFL94b]. Applications [ACFM99, Ano98i, Ano99r, Ano99y, Arn95b, Bah90, BBG^{+}99, CB90, CH97, Coo97b, DDD92, DBCF95, DZL97, Den95, Fuc97, GS94a, Gol90a, GWTB96, HdJB97, HS99c, Kam93, Kle94, Laz98, Lea98, LBBA99, LWS^{+}94, MJS97, MDK96, Nee90, Noe98, PK93, Pom94b, PTB^{+}97, SPK^{+}96, SL96, STW93a, Sch94b, Sch97b, SD98b, SL95, SW95, SW90, USE94g, Vck98, Vin96, Wal94b, WU97, WMB99, YT95, USE96i, Seo93a]. Applications/Mobile [Ano99r]. Applied [Dav94b]. Apply [DF95]. Applying [Chr98a, Lit92, POS^{+}99, Sze99, CG99b]. Approach [AW99b, BTC97, Bis92a, BL90c, CD94a, CHK^{+}98, DT97, Duc97, GB98, GG91, GA94, KS94c, KMR91, KCVP95, Kur90, LaP94, LD97, Lyc94, MB96, Mud99, MM92, NMMS97b, New95, Par90, Par93a, PE92, RM97, SB93, Sel91, Sin97d, Sto93b, TSE90, WS99, WCO^{+}90, ZTM96, ZSBC98]. Approaches [LS98a, SEKN92]. Approaching [Mil99b]. Approximate [WM92]. Approximating [Pol93]. Apps [Wal97a]. Apricots [Dij97]. April [USE90j, USE91k, USE92f, USE92i, USE93h, USE94e, USE94g, USE94m, USE95i, USE98c, USE99g, USE99l, Ano99-58]. Arbor [USE92h, USE95i]. Archaeology [Gra92b]. archie [ED92]. Architectural [Hil92, KLA^{+}90, Rey92, Ros92, Sch99]. Architecturally [Rus99a]. Architecture [BB^{+}91, Bas98, Bed90, BVAD98, BSKH93, BFH^{+}92, BCVC91, BJM^{+}96, CMS90, CDK^{+}99, CBB97a, DBGM^{+}98, Den95, DB96b, FB96, Gol92, GSR93, GMPS97, Gut99, IB93, IM93, LPV^{+}91, MB98, MO99, MJ93, MR97a, Mol95, NK98, OSS^{+}92, PKB^{+}91, Raw93, Roh96, SB94, Seb91, Ske92, Tor96, TC95, Vau90, VBD^{+}92, VMMD98, WL98, YTJ91, ZBCS99, Bra99, BCLC99, RMS^{+}99, SSL^{+}99]. Architectures [GV91, Lan92, MPS98, MR97a, Sch95b, TB99, USE92i, USE93k, Wel92, Wel94]. Archive [LT97, Zwi91b]. Archiving [Mil99a, SM99, Sin98c]. Area [BVAD98, FHLW99, IP94, LWS^{+}94, MET98a, RAC97, Sel97a, SS94b, VDAA99, WB93, YGM95]. Areas [Bel96b, Tay99]. Aren’t [Ous90a]. Arjuna [Par90, Par93a, PSWL95]. Ark [Ins96f]. Arla [DW98a]. ARPANET [Ste95]. Array [SW96]. Arresting [Mad92]. Art [Bar94d, Bar92d, Eng93, Rus99b, She92b, Uma96c, kc99a]. articles [Far96i]. ASA [Ano99r]. ASA/MA [Ano99r]. ASCII [HF91]. Aspect [Sat93e]. Aspects [Eva95b, GB92]. Assembly [Kap92a]. Assertion [Gau92]. Assertions [Coo97a]. Assignment [Win94a]. Assistant [Pet94].

ASTLOG [Cre97]. ASX [Sch94b].

Asynchronous [AMML96, BC91, NN94a]. AT&T [Joh98, Mog90]. Athena [BR98]. Atlanta [Ano99h, Den99c, USE91j]. ATM [BE98, ATS94, BZT99, BC94, Jac94, MBH94, NN94b, NPW93, SMSM98, WVT94, Zea97]. ATOM [ES95]. Atomic [BN92, CHTY96, MKC92]. Atomicity [ST96]. ATP [CV93]. Attached [Puc91a]. Attack [Jon95, KI99, LA98]. Attack-Resistant [LA98]. Attacks [AM98b, CPM^{+}98, GMNR99, Jas96, Kle99b, Lev98, MDS99, Rod93, SP99b, Van98]. Attitude [Zwi94a]. Attri [SKCB97]. Attribute [SO92b]. Attributed [Lam91b]. AU’99 [Ano99v]. Auction [MW98b]. Auctions [HTK98, KF98]. Audience [SBB96]. Audio [AOG92, AHL^{+}91, GLP^{+}93, Haw90a, KH97b, RBB^{+}91, SW91b, TP91]. Audio/Video [AOG92]. AudioFile [GLP^{+}93]. Audit [Cra94b, Sat92]. Auditals [Lanxx]. Auditing [Cox98a, Muf95]. Auditor [RL91]. August [USE90i, USE92k, USE92n, USE92o, USE93i, USE94f, USE97a, USE97f, USE98e, USE98a, USE98b, USE99e]. Aurora [GMR95]. Austin [USE99a]. Australia [Ano99v]. Authenticated [RH93]. Authenticating [KS99]. Authentication [ADKN90, Bak99, Bec99, BM91a, Bis91, Bru98, DVPL92, FL96, Gol90b, Hil97, Ins95b, IH98, KCCT94, MAM95, Mee98, SSH93b, SA95a, SKAM99, TAP90]. Authoring [DG91]. Authorities [Ell96]. Authority [RG95]. Authorization [FL96, SSH96]. Authors [Bin99]. AutoCacher [Min93a]. Autochangers [TSE90]. Autofs [CS93, Lab99]. AutoHack [Muf95]. AutoInstall [FL98]. AUTOLOAD [PS92]. Autoloading [Chr92]. Automate [Lib90b, Lib91b, Par99a]. Automated [AK94, Car98b, GMS96, HA93, HM92, HH99, Kap94, KSKD99, LPS99, Mar96b, NLG^{+}99, Rid94, VW99a]. Automatic [CG99a, CCK98, CPM^{+}98, Fin95a, FD98, HS95b, HS99b, KW96, Kin90, KC99b, ML91a, MDK96, RDM95, SBRH96, Spi98]. Automatically [MSS99, Pul98, SHM93]. Automating [Fin94a, Fis96c, Lib96a, Sco97, TB99]. Automation [AKS98, Cla92, Fin97b, SSZ95]. Automounter [Lab99, Ste93b, CS93]. Autonomous [DLO92]. Autonomy [Wan91]. Autos [Tab94]. Availability [BS92a, Bar97, FP93, VF95]. Available [Atk96, BEM91, BS92b, HSOS99, KS94b, McK98]. Avoid [Lib96a]. Avoiding [Fer95, LFR^{+}98, Lim99c]. Award [Ano95b, Ano99a, Chu99]. Awards [Ano95c, Ano99-55, Pet96]. Aware [CI97, CE99, CRS99, FK99, NPS95a, NPS95b, RASS97]. Away [Den99a, Dij96b]. Awk [Yig97, Ker91, Spe91, Kno97]. AWT [O’N97b, Nag96].

B3 [E^{+}91]. Back [Dar96d, Dar97b, Mit96, Mit97, Phi95]. Backbone [Hei90]. Backend [RC97]. Background [Lib94]. Backtracking [CA96]. Backup [BMLH97, BL96, Eng91, Han99b, HMF^{+}99, Kol91, Lam90, Lam91a, Leb97, MK92, PK91, Pre97, Pre98, Rom90, Rom91a, Ros94, Shu91b, dSGM92, dSG93, Wei91, WWDF98, Zwi91b, Mel99b]. Backup/Restore [BMLH97]. Backups [AK94, Sel98a]. Bad [Kor97b, Ous96]. Badge [SUZ98]. Bags [Kol98k]. Bakeoff [Wil93]. Bal [Chr97b]. Balancing [GG91, Sch95a]. Balloons [Cra96]. Baltic [Kol99h]. Baltimore [USE92m]. Bandit [NN92c]. Bandits [Bia96b]. Bandwidth [Cha99, GK97, LB97, MK95, NN92c, SW98, SO94, SCD94]. Bank [Lod93]. Banking [Sch95g]. Banks [Gee94]. Barb [Kol99a]. Barbara [USE97b]. Barn [HS93a]. Base [HW93, Lar94, Law94, Mey91, Mil93, NB94, SV90b, SV91]. Based [Alf94, Bab90, BGP^{+}94, BK97, BHJV92, BMLH97, Ber92, BP93, BKM97, Bir95, BAKW92, Boy91, BBG^{+}99, BCLC99, IG99, Car93c, CLL97, CN90, CCK^{+}91, CK96a, Cho98, Chr90b, Cla97, Cou96a, DT97, Ess90, FKSV94, Gau92, Gol99, GR91, GR93, GLOR91, HS93b, ISS97, IS99, KNM95, KS94c, KC99b, Kra99, LBLM90, Lon95, MB98, MSS99, Mas91, MKB90, Mot95, MW96, Neu92b, NN93a, NLG^{+}99, Ous91b, Pul98, RS91, Ros96h, SM91, Sel95a, Skl91, SBRH96, SC93, SKP94, Spi99, TSM95, WYOT93, WWWK95, WAP99, WS99, WVS^{+}99, YKB94, ZSBC98, ZBCS99, ZIL96, Ama98, AB96, ABB^{+}97, BBK99, EFJ96, GM99, GWP99, HK99, How95, Kan91, KLB98, Kle99b, KCVP95, LB98, MJS97, McN93, NUO^{+}97, PBS97, SD98a, Sav99, SH98a, SK99, SKS92, TAP90, TJM^{+}99, YL99, Min95]. Bases [NB94]. Bash [Ram94]. Basic [CS90, Pol93]. Basis [Man98]. Batch [CS92, CS94]. Battle [Chexx]. BayLISA [McD96]. BBFS [HR92]. bbn [Sch92a]. bbn-public [Sch92a]. BC [CA96]. BDL [BA97]. Be [BF99, Bel92, Tab94, Zwi99a, Zwi99b, PCMI91, Ros98b, Wal95a, Zwi98]. Beach [USE92e, USE92l]. Beam [Eir94]. Beans [NK98, Rie97]. Beast [CDGT98]. Beating [AEK96]. Bedfellows [Dar97d]. Beeper [Zwi94f]. Behaved [Mur93]. Behavior [BA97, Che94a, GSS99a, SBS99, Tra90, VK98]. Behaviors [Kin90]. Behaviour [Lu97]. Behind [Muf95, Rat98]. Being [Dar96f, Men94]. Bell [LRNL97, Bah90, PPD^{+}95]. Bell-LaPadula [Bah90]. Belongs [Wel91a]. Benchmark [WN92, SM98b]. Benchmarking [KW93, SA95b, WK93, Rob99]. Benchmarks [Fin90]. Beneath [Poo94]. Berferd [Che92]. Berkeley [Ano94a, Skl91, Ste93b, VS90]. Best [Joh93a]. Beta [Fla97]. Better [BF99, Dar99a, Hol97a, Lyc94, McC97a, McC97b, McC97c, McC97d, McC98e, McC98f, Mog94a, NC92, Pol96, Sch96d]. Between [Chr97b, GC96, KA94, RK97, vBC^{+}92]. Beyond [Bah90, EKB^{+}92, For99b, GK99, GMPS97, Zwi92d, Ste95]. Bias [WRJ92]. Bible [She92a]. Bibliographic [GLW94]. Bids [HTK98]. Big [Ahr93, Mel97a, Sch96c, Tra90, Car99a]. BigDog [FL96]. Biggest [Chr98b]. Bill [Kol98b]. Binaries [And91]. Binary [CWS90, Gar90b, GS94b, GLW95, HB99]. BIND [Col93, Dar97f, Vix95, Vix99]. Binding [PB96, YCO94]. Bindings [HS95b]. Bingo [Kle99c]. Biology [BG97]. Biometrics [Cal99, For99a]. BirLix [SHK92]. Birthday [Sto98a]. Bit [CH97, Dra91, MWO^{+}93, LZ97, YBA93]. Black [Add96]. Block [AS93, DR99]. blocking [GC96]. Blocks [ST96]. BNR [Rat94]. Board [Ano95d, Ano97k, Dar98a, Eva94a, Kre97, Mil95, Mil97d, Par93b, Par93c, Sch93a, Wil95c, You93c, You93d, You94a, You94b, You94c, You94d, You95b, You95c, You96a, You96b, You97c, You97d, You93a, You93b, You97a, You98b, You98a, You98c, You99b, You99c, You99d, You99a]. Body [Hal99d]. BOF [Sal92f, Sma92a]. Booby [Ven92]. Book [Ann98, Ano99-27, Bar92a, Bar92b, Bar92c, Bea97, Bea98, Bir90, Bir95, Bra92b, Chi97, Chr97a, Chr98a, Chr98c, Chr98d, Chr98b, Chr98e, Chr98f, Chr98g, Cof97, Col91a, Col91b, Col92a, Cuc90, Eng93, Far92, Far97a, Fly98b, Fly99a, Gor90, Gwy90, Hae91a, Han97, Han99b, Haw90b, Hum97a, Hum98a, Jen97, Jen98, Joh97a, Kno97, Kol95a, Kol97a, Kot98, Laz97, Laz98, Laz99, Lea90b, Lea92b, Lea95a, Lea97, Lea98, NN92a, O’N97a, O’N97b, O’N98, Pax91, Pom97a, Pot97, Quo91, Roo97, Roo98a, Roo98b, Roo99, Sal93g, Sal96c, Sal97c, Sal97d, Sal98b, Sal98c, Sal98d, Sal98e, Sal98f, Sal98g, Sal99g, Sal99h, Sal99i, Sal99j, Sal99k, Sal99l, Seo93a, Seo93b, She92a, She92b, Sie97b, Sie97a, Sie98a, Sim92b, Sim92a, Sim93a, Spe95b, Spe95a, Spe97a, Sub98, Uma97a, Uma98a, Uma98b]. Book [Van99, Wal92a, Zwi92b, NN92a]. Bookworm [Sal92c, Sal92d, Sal92e, Sal92a, Sal92b, Sal93a, Sal93b, Sal93c, Sal93d, Sal93h, Sal93e, Sal94a, Sal94b, Sal94c, Sal94e, Sal94f, Sal94g, Sal95b, Sal95c, Sal95d, Sal95a, Sal96c, Sal96d, Sal96e, Sal96f, Sal96g, Sal96h, Sal97c, Sal97d, Sal97e, Sal97a, Sal97f, Sal97g, Sal98b, Sal98c, Sal98d, Sal98e, Sal98f, Sal98g, Sal99g, Sal99h, Sal99i, Sal99j, Sal99k, Sal99l]. Boot [Bra92a, Rom91b]. Bootstrapping [TH98]. Boss [Ins96c]. Boston [USE94j, USE97d, USE98b]. Bottleneck [ACG^{+}98]. Bottlenecks [AEK96, ACO92]. Bottom [Bor92, Sim96a]. Boulder [USE99c]. Bouncer [Hol97a]. Boundaries [MWCR90, MWCR91]. Bounds [KL97]. Box [Add96, BS92a]. Branch [VR99]. Branchstart [Tom98]. Brave [Ano97d, Swa96]. Braveheart [Cuc95]. Brazilian [dSM96]. Brazos [SB97b]. Breaches [BKS90]. Breadth [Sto96a]. Break [Bel95, BS97a, Han92]. Break-in [BS97a]. Break-ins [Bel95]. Breaking [FM99, Gee94, Haa90, SS99]. Bricks [Dam95]. Bridge [Lim99d, TLK99]. Bridging [Lar94]. Brief [NWRP^{+}91]. Briefs [Ano99-53]. Briggs [JS93]. Bring [Ahl99]. Bringing [HS99a, Jon91, Mot95]. Bro [Pax98]. Broadband [HR92, LWS^{+}94, NN94a]. Broadcast [AKS98, Kan91, SSZ95]. Broken [Mil99c]. Broker [MB97]. Browser [AM98b, Bal96, Ins96e, KL93]. Browsers [HS95a]. Browsing [LB97, MSG97]. Bruce [Muf99]. Brute [CD98, KI99]. Bryan [Kol95j]. BSD [NN92a, AMP^{+}96, Kat94, LKY93, MJ93, McK98, SSB92, YLK93]. BSD/386 [Kat94]. BSDI [Sma92a].

Budgeting [MH99]. Buffer [CPM^{+}98, MM93]. Buffer-Overflow [CPM^{+}98]. Buffering [BS96, PDZ99, Spi99]. Buffers [BKW94, Koo93]. Build [CV96, Dea93, Pol94, TM97, YCE^{+}97]. Builder [Dam93, Uhl95]. Building [ALO90, Add96, Arn95b, BKC94, Cal96, Cha92a, CN90, ES95, Fre91, GZ94, HdJB97, Hir97, Hop96, Ins96d, JPR96, Kap92b, MAM95, Mee97, Mur93, NKM93, PBR91, RS91, SPK^{+}96, Shi97, ST96, SW90, SUZ98, Tan96, Uhl91, Vin96, Woo98, WMB99, Yor99, YM99, Dep94]. Built [Win94b]. Bureau [GH98]. Burlington [USE90e]. Burt [Mel99b]. Bus [OW94, PS93, Ske92]. bused [Dar97a]. Business [Avo98, Hae91a, JABA98, Lea95a, Mue94a, Dep94, Kol94g]. Business-to-Business [JABA98]. bute [SKCB97]. Butter [Lea99]. buzzerd [HM92]. Buzzword [Kle99c]. Bytecode [Lew96, MMBC97, PW97]. Bytecodes [BM98a, LZ97]. Bytes [MP93]. Byzantine [CL99].

C [Fly99a, Kol95a, Laj96, Quo91, SS94a, Tie93b, USE90j, USE92k, USE92o, USE94e, Ada94, Add96, AT92, AB91, AGG^{+}94, BKC94, BR94, Bea96, Ben94b, Bis92a, Bli90, Box96, BPR92, CFLM92, Car90a, Car91a, CGMV91, Car93a, Car91b, CSYG94, Cog90, CHS91b, Cop94, DeW91, Dea90, Det92, Dil95, DSS90, DMR92, EP91, ED94, FN91, FWF^{+}90, Gau91, GG98, GS94b, GS90, GR91, GR93, GC90, GKS90, Gra92a, Gue92, Hae90a, Hae90b, Hae90c, Hae90d, HS91a, HJ92, HS95b, HG98, IL90, IL93, JJ90, Joh98, Jon91, JSS90, KA94, Ker91, Klu90, KS90, Koe92, KS93, KS94c, Lae95, LM94, LN95, Lau95, Lea90a, Lea93, LUM91, LMU93, Lin92, Loe96, Mar91b, Mar93, Mar96b, Mar96c, McC97a, McC97b, McC97c, McC97d, McC98e, McC98f, ML91a, Mey91, MMR92]. C [Mog90, Mue96a, Mur92, Mur93, Nil94, Pal92, PR94, Par90, Par93a, Par95, PBF^{+}96, RM90, RW91, RC95, RMC90, Sak92a, Sak92b, Sel90, She95, Sim94a, Sma97, SC92, Spe93b, Spe97a, SL92, Str93a, Str93b, Tie90, Tie93a, USE90j, USE91k, USE92n, Vin96, Vis96, Wag92, Wal91a, Wal93a, Wal93b, Wal94a, Wat96, Wil97a, Wyb90, ZJ90]. C/C [Fly99a]. C2 [Eps92, Ste92]. C3 [FPSW91]. CA [USE92l, USE93f, USE93m, USE94h, USE95h, USE97c]. Cable [Pom94b]. Cache [CDN^{+}96, CMS90, DM92, Dou93, GS96b, KW97, Mac94, Min93a, MS94, RLBC94, SKS90, YADL99, MH92]. Cache-Affinity [DM92]. Cached [VBD^{+}92]. Caches [BNMK99, BMH98, DMF97, KW97, OW94]. Caching [BKW94, BP99, Bel96a, BKM97, BA91, Bon94, CFL94a, CFL94b, CI97, CD94b, DWAP94, DHR97, Hau94, IC97, KTP^{+}96, KAD96a, KAD96b, KL97, KPR94, LCC94, LC99, LPJI98, MH92, NH93, PDZ99, Pom94b, Roy93, SH96, Sch95b, SO92b, SAYZ99, VA98, Wel91b, Wel91c, WVS^{+}99]. CAD [RAF92, TLK99]. Caldera [Fia95]. Calendar [Sal98a]. California [USE90g, USE90j, USE91f, USE91h, USE92e, USE92j, USE93k, USE93j, USE93l, USE94i, USE94f, USE94l, USE95k, USE96e, USE96i, USE96h, USE96k, USE97b, USE97e, USE97h, USE98f, USE99g, USE99f, USE99l, USE99m]. Call [Ano95c, Ano98b, Ano98c, Ano98d, Ano98a, Bin99, HG93, Kle97, Kle99a, San90, Sel90, Wil95c, Wil95b]. Caller [JKS^{+}93]. Calliope [HSE96]. Calls [Pat93, PBF^{+}96, PA90]. Calvinistic [Poo94]. Cambridge [USE93i, USE94e, USE99k]. Came [VZ91]. Camel [Pom95c]. Camels [Hop92]. CAMP [Ber99b, Ano98-43]. Campus [Cam94, How90, How91, Sel91]. Can [BF99, DF95, GGI^{+}91, Hae92, Lim99d, Mue96b, PCMI91, Ros95c, Sie96, CG05, Man91]. Canada [USE94g, USE95j, USE96g]. Can’t [WT99, Man91]. Capability [Bar97, MD95]. capable [Tal95]. Capacity [BD97, Chr98d]. Capital [Lip95]. Capture [MJ93]. car [Zwi93b]. Card [MK99, OK99, SKAM99]. Cards [Gut98a, KCCT94, AG99, Bak99, BF99, Gir99a, GSTY96, LW99, NFQ99, SS99]. Career [Dar98d]. Cartoon [Spe95b]. CASA [CF94]. Case [Car91a, Car93a, CMB^{+}98a, Chr98b, CJJ^{+}98, DH99, Eng92, IMC93, KOO95, Mad92, MP99, Mil98a, Mog99, Pol97, RMC90, Sch93c, SPR96a, SF97b, SUZ98, Wal91a, Wal93a, WIK92, WD95, TM97]. Casper [VBD^{+}92]. Casting [Ste95]. Catalog [Bar92c]. Cataloging [Cal93]. Catalogs [Kel95]. Caubview [BMH98]. Caubweb [LM97]. Causal [Boy91, NSM99]. Causes [Oka98a, Oka98b]. Cautionary [AK96, Kol94f]. CC [Pom94a, Bas98]. cc-NUMA [Bas98]. CCEL [DMR92]. ccNUMA [BCC^{+}99]. Cdiff [Gra92a]. CDMA [Kar93]. Cdt [Vo97]. Celebrant [Ree97]. Celebrating [Jon94]. Cellular [Kar93]. Center [HN97, Car99a]. Centers [dSM96]. Central [Har94, SCD94, Ano97o]. Centralized [HA92, MM96a, Sch92b]. Centrally [Mot91]. Century [Kol94e, Cla92, Sch96i]. CERT [Car99a]. Certificate [NN98, TJM^{+}99]. Certificate-based [TJM^{+}99]. Certificates [DS92, Sim96d, TAP90]. Certification [Ano99-53, Ber99d, CDGT98, Dij98, Dij99b, Ell96, GD96a, Kol99a, LA98, McD99, SH94, Yor99]. Certify [OK93]. Cfengine [BS98, Bur99a, Bur99b, Bur99c]. Cget [Che97]. CGI [Lib96b, Mah96, Tay95b, Vck98]. Cha [CHHL99]. Cha-Cha [CHHL99]. Chaining [Jas96]. Challenges [CSYG94, Nes99, Rei99, SO99, Sno96]. Championships [Kol98a]. Change [Ben94a, DFKM97, HN97, Mil97f, NPW93, Pla99]. Changer [Bis92b]. Changes [DB96a, Hum99a, Mil97d, Rom91b]. Changing [Bis90, McK93b, Mil93, Pal92]. Channel [USE92f]. Channels [DFGG97].

Chaos [Cou97, Wil99]. Char [Kol97b]. Character [Ped90]. Characteristics [DMF97, WCS93]. Characterization [Hei90, Kei90]. Characters [Gsc95a]. Charity [Gri98]. Chase [Ano97e]. Cheap [Gra98c, Kol94j, MJ92, War99b]. Cheating [ACG^{+}98]. Checker [FS90, Kin90, Sat92, ZL96]. Checkers [KS94a]. Checking [BD96, HTWW96, HT92, HT94, KW96, NL96, PKA98]. Checkpoint [Rot99b, Rot99c, SSBK98]. Checkpoint/Restart [Rot99b]. Checkpointing [ASSB99, LS92, PBKL95]. Checks [Sze99]. Chest [Fly99a]. Cheswick [Kol98b]. Chicago [USE96j, USE99j]. Chime [SD98a]. Chocolate [Lea99]. Choices [CIM92, Spa92]. Choosing [Leb97, Spi97]. CHORUS [AAO92, KSAL96, BG93, BGH^{+}92, BCC^{+}91, Roz92, WIK92]. CHORUS/MiX [KSAL96]. Chrg [TL96]. Chrg-http [TL96]. CIA [Gra92b]. CIFS [Swa97a, Swa97b]. Cincinnati [USE93g]. Ciphers [DR99]. Circuits [Eva96d]. Circulation [FKT^{+}99]. Circus [GHN90, GHN91]. Cisco [Chr98c, Chr98e, RFR99]. City [USE95f, Car99a]. CitySpace [CV96, Vel95]. Clair [Kol98d]. Clara [USE93m, USE99g, USE99l]. Clarification [Cox98b, Rei98a]. Clarity [Pom94e]. Clark [Pol93]. Clark-Wilson [Pol93]. Class [BKC94, Den92, FN91, GS94b, Lea92a, Lu97, Mar96b, Mil96b, NB94, Van91]. Classes [HG98, Mey91]. ‘Classical [BE98]. Classifier [BGP^{+}94]. Classroom [Cat90, CKM95]. Cleaning [BHS95]. Cleanup [Gre91]. Clerk [Lim97]. Click [RAM98]. Click-Through [RAM98]. Client [All95, BGG^{+}98, Car98b, DWAP94, DMF97, GS94a, Gol99, GB97, Gue92, KSKD99, Sie97a, WWDF98, ZPMD97, DW98a]. Client-Relationships [BGG^{+}98]. Client-Server [Gue92]. Client-side [KSKD99]. Client/Server [GS94a]. Clients [LCL98, YCE^{+}97]. Climbing [JSS90]. Clinical [AB96]. Clock [MT93]. Clocks [DG95, DGT96, Yah94]. Cloned [JR91, Zwi92d]. Cloning [JR91, Soq97]. Close [Pul98]. Closet [SF97a]. Closing [Man98]. Clouds [DCM^{+}90, DAM^{+}91a, DAM^{+}91b]. Clue [YCO94]. Cluster [BCLC99, DMST95, FM99, FLA94, KCVP95, MGM^{+}98, VDA^{+}98, ZSBC98, ZBCS99]. Cluster-Based [BCLC99, ZSBC98, ZBCS99, KCVP95]. Clustering [SSB^{+}95]. Clusters [ASSB99, AP97, BC97, YL99]. CM [LIN^{+}93]. CM-5 [LIN^{+}93]. CMD [CMB^{+}98a]. CMU [Wat96]. CO [USE91e]. COBEA [MB98]. Cockpit [MSN93]. Coconut [SKS92]. Code [Ano97l, CL97, Dar97c, DSS90, Fly97, FWF^{+}90, Gra98a, Gra98b, Gra98e, Gra98d, Gra98c, Gra99c, Gra99d, Gra99e, Gra99b, KR99, Lan94a, Lan94b, MT93, Mil97b, MMBC97, RF95, RW91, SG97b, HG98]. Codebook [GS99b]. Coding [Fly97]. Coexistence [Vas91]. Cohabitation [BG93]. Coherence [AAL92, CLD^{+}97, MS94]. Coherency [FCNL94, PK96]. Coign [HS99b]. Coir [Sun97]. COLA [KK92b]. Collaboration [Gon96, LM99b, TSM95]. Collaborative [Eva93b, KLB98, PK93, SPK^{+}96]. Collaboratively [Lib97a]. Collected [DK95]. Collection [ACTW96, Det92, ED94, FS94, Nil94, Sel90, Van90]. Collections [Chr97b]. Collector [EP91]. Collide [Sca97b]. Colorado [USE90f, USE91e, USE99c]. Columbia [BKS90]. Column [Mog90, Rob90]. COM [CHW98, HS99c, WL98]. Combinators [CD97]. Combining [MSG97, SA95a]. Come [CV96]. COMERA [WL98]. Comes [Bas98]. Coming [Dar98b]. Command [Ous90b]. Commandments [Spe93b]. Commands [HK95b, Pol93]. Comment [Col92b]. Commerce [Avo98, CST95, GS96a, dPG98, GD96a, HS96, HTWW96, Hon95a, Hon95b, JABA98, KS98, LHB96, Man95, MMS97, MW98b, New95, Sch95g, SBV95, Sim96d, ST96, TW96, USE95g, USE96h, USE98b, Wri96, YT95, Ann97, Ano98a, Ano98m, Ano98-31, Far97a]. Commercial [BBD^{+}96, HN97, PK91, Pre97, Sch96a]. Committee [Nem99a, WKR98, You97b]. Commodity [GPRA98, Ism98, TM99]. Common [Sha97]. Communicating [BS90]. Communication [AOW96b, AOW96a, BZM91, BVG^{+}97, CD94a, CLMS91, CDK^{+}99, CK96a, CG92a, DDD92, Eva93b, GBB93, HS94, IGP95, Lie92, Maf95, MB91, MPH99, Pol96, RAC97, RAF92, Sal96b, Sat93e, ZJ90]. Communications [dC93]. Community [Cou96a, Eva95c, Gee98, HS91c, Sch92a, Sto95b, Woo94]. Community-Based [Cou96a]. Companies [Joh94d]. Companion [Bar92b]. Compaq [BGT98, FM99]. Comparative [Ste92]. Comparison [CL98, CRH99, CS90, DOKT91, GES98, KTP^{+}96, LB96, SEKN92, SSB^{+}95, SS96a, SS96b, Wel92, Wel94, ZY98]. Compartmented [Cho99]. Compatibility [GLOR91]. Compilation [Eng97, FST91]. Compile [GG98]. Compiled [MMBC97]. Compiler [DG94, Lew96, MDK96, PBF^{+}96, PTB^{+}97, RC95, RC97, SCD90]. Compiler-Inserted [MDK96]. Compilers [FST91, LM99a, SG97b]. Compiling [Bot98, FST91, JM94]. Complete [FL98, McK98, Ver97]. Completed [Ano98-43]. Complex [RL98]. Complexity [Ros96h, Sho97, SW95]. Compliance [Dav96]. Compliant [KAD96a, KAD96b]. Complicated [Kol94c]. Component [BBG^{+}99, HHLS90, KC99b, NK98, SUZ98, WCN93]. Component-Based [BBG^{+}99, KC99b]. Components [Add96, Box96, CBE93, Cow99, MPZD98, Pol94, SHAS95]. Compose [GG97]. Composing [Loe96]. Composition [JD94, MM96b, NB94, Sah95, SG97b]. Comprehensive [LV99]. Compressed [Tau91, VcC97]. Compression [CE99, CDL99, Dou93, Haw90a]. Compromise [BBP99]. Computation [IG99, KMR91]. Computational [Mel97b, RA98]. Compute [BGT98, NLZ99]. Computer [Bas98, Bur98, Cha96, Chexx, CSYG94, Dij96b, Eng93, Haw90b, Hic99, HL93, Hop92, KI99, KBM^{+}96, Kol94g, LaP94, Lev98, Lin94, Pie94a, Pie95, She92b, TMJY91, Van98, Ven96, Kol94f, Kol95i]. Computer-Participative [Lin94]. Computer-Related [Kol95i]. Computerized [Var95a]. Computers [ASG^{+}93, AP97, ASB^{+}99, BF99, Chr97b, Cox98a, DCK^{+}94, DKB95a, DKB95b, Har90, Har91, LKHA94, O’D90b, PBEDT97, Tab94, Tro96, Whixx]. Computing [Abb94, Ano98u, Ano98i, Ant95, Bab90, BLO^{+}91, BDD^{+}93, BL97, BPR92, CD97, Car93c, CMB^{+}98a, CCW97, Dan94, Dil95, Eva94c, EG97, How90, How91, Hum96a, IGP95, Jon94, Kol93c, KPR94, MPS98, MGM^{+}98, MQ97, NAU93, NPS95a, NPS95b, Pie99b, RAV95, Sat95, Sel97c, TSO^{+}96, TC95, USE90a, USE90b, USE90c, USE90d, USE91a, USE91b, USE91c, USE91d, USE92a, USE92b, USE92c, USE92d, USE93a, USE93b, USE93c, USE93d, USE93i, USE94a, USE94b, USE94c, USE94d, USE95a, USE95b, USE95c, USE95d, USE95i, USE96a, USE96b, USE96c, USE96d, WYOT93, WW97, WB93, Wei95b, WS93, WCM93, Win96, ZD93a, Kol94f, LH90, Pax93]. concatenation [MdR99]. Concept [Bir90]. Concern [MR97a]. Concert [AGG^{+}94]. Concert/C [AGG^{+}94]. Concurrency [AL93, BOM95, BD92, CJJ^{+}98, GKAS99, Sel90]. Concurrent [AB91, BA97, CW96, LH97, LCH92, PE92, Puc99, Sch96g, SHM93, STV90]. Conditional [GK99, VZ91]. Conditions [BD96]. Conduct [Lan94a, Lan94b]. Conduit [ZJ90]. Conference [Ano93b, Ano94f, Ano96d, Ano97i, Ano98h, Ano98e, Ano98n, Ano98u, Ano98-28, Ano98x, Ano98y, Ano98-29, Ano98-30, Ano98-31, Ano98-33, Ano98-27, Ano98-32, Ano98z, Ano99s, Ano99t, Ano99f, Ano99v, Ano99b, Ano99x, Ano99-28, Ano99-35, Ano99-31, Ano99-30, Ano99-33, Ano99-34, Ano99-29, Ano99-32, Aut95b, Aut95a, Bai97, CWT^{+}97, CPK94, Var96, Dyk91, mH96, Isl95, Jon98, Kar96, McD92, Met95, Num94b, Old94a, Old94b, PS97, Red98, Num94a, Ros93, Rot95, SCP^{+}93, Sal97h, Sal92f, SS94a, SD92, USE90e, USE90f, USE90g, USE90h, USE91e, USE91f, USE91g, USE91i, USE91l, USE92e, USE92g, USE92k, USE92j, USE92n, USE92o, USE93f, USE93g, USE93l, USE94h, USE94j, USE94e, USE94l, USE94m, USE95h, USE95k, USE95e, USE95l, USE96g, USE96j, USE97b, USE97c, USE97e, USE97g, USE98c, USE98e, USE98f]. Conference [USE98g, USE99a, USE99g, USE99f, USE99b, USE99i, USE99m, Wag92, Wal94a, WSR93, WUS^{+}92, WP94, Zwi93d, vW98, USE90j, USE91k, USE96k, USE98d, USE98h, USE98i, All92, Ano98x, Den99a, Ano97b, Ano98o, Ano98f, Ano98q, Ano98v, Ano98-29, Ano98-32, Ano99j, Ano99k, Ano99u, Ano99w, Ano99c, Ano99d, Ano99e, Ano99o, Ano99-28]. Conferences [Ano96b, Sim97]. Conferencing [Mot95, NWRP^{+}91, TP91]. Confessions [Fis96a]. Config [RM94a]. Configurable [DG94, Puc91a, VZ91]. Configuration [And94, AR91, Bur95, Chr98d, EL92, Eva97, Fin94a, Fin97b, GZ94, JS97, JV95b, JV95a, Kin91, KC99b, Oka92, Pla99, Rat94, SL93, Sim91b, ZL96, dSdCF^{+}98, dSdS98]. Configurations [Cal96, RM94a, Wal95d]. Configurator [JS97]. Configure [deL99]. Configuring [Pag98, PBS97]. Confining [WSB^{+}96]. Conflict [RLBC94]. Conflicts [Kol94g, KS95, RHR^{+}94]. Conformance [Lib92]. Connected [Duc97, HH95a, HH95b]. Connecting [AT92, Mos94a]. Connection [CFHB99, KMS^{+}93]. Connection-Oriented [KMS^{+}93]. Connections [Uhl90, Ylo96, KC93]. Connectivity [BTS^{+}99, CS97, ME99, MS94]. Conquering [Dan94]. Conservative [BL90c]. Consider [McD99]. Consideration [War94]. Considerations [Chr90b, Hel92, IC92, SW91b]. Considered [LT91, Sim99, SC92]. Consistency [AAO92, CD92, FS94, Gol92, GS96b, KFJ94, Mac94, PKA98, TD92, YADL99, ZIL96]. Consistent [CGS^{+}96, MdR99]. Console [FR90, Fin91, Mat98]. Consortium [Wat96]. Constrained [AD93]. Constraint [BK97, Ken96]. Constraints [CA96, HT92, HT94, JR99]. Constructing [Sma97]. Construction [Mas91]. Consultant [Cha96, Dar96f]. Consulting [Sim96a, Sto96a]. Consumption [MQ97]. Container [SBV95, Vo97, WN98]. Containers [BDM99, Wei95a]. Content [GM99, JPR96, Leh98, MAH98, SD98b, SAYZ99, YL99]. Content-based [GM99, YL99]. Content-Derived [MAH98]. Contest [Ano96c, Ano97o]. Context [SP91b, Str90, Tay95b]. Context-Sensitive [Tay95b]. Continuations [Dea93, Lie92]. Continuous [AOG92, TN96]. Contract [Cla96]. Contributions [Sch92a]. Contributor [Ros96b]. Contributors [Ano99-54]. Control [ALO90, AGF90, AAL92, AR91, BHJV92, BOM95, BA97, BGMR91, CLS92, CD92, Cla97, CMSvE99, DL93, Eng97, FL96, Fut95, GSPW98, Hel92, HK99, HN97, KR99, Lee98, Lin94, Min95, Mit91, MQ97, NB97, Pre99, Rei97, Rob92, RAF92, Rud97, Sin99b, STM^{+}99, TJM^{+}99, Ven92, WCO^{+}90].

Controlled [CFL94a, CFL94b, FKT^{+}99]. Controller [ES92]. Controllers [Str99]. Controlling [Kin90, Lib91a]. Controls [Gam93, Str90]. Controversy [Car91a, FG90, Wal91a]. Conundrum [Kol96k]. Conventional [DV90, DF95]. Convergence [NPW93]. Convergent [CG99b]. Conversation [RK97, Spi97]. Cookies [Tay96e]. Cool [Lea95b]. Cooperation [BG93]. Cooperative [CBE93, DWAP94, SH96]. Coordinated [BC97, GMNR99]. Coordination [Car99a]. COOTS [Ano98-27, Ano99s, Ano99t, Ano99-29, Har96b, PS97, USE95k, USE96g, USE97g, USE98c, USE99f]. COOTS’98 [Ano98b]. Coping [AL93, Har99a]. Coprocessors [SA98, YT95]. Cops [FS90]. Copy [Chu96, MdR99]. Copying [EP91]. CORBA [BPT98, DFGG97, HdJB97, KAD96a, KAD96b, MB98, Maf95, Mar96a, Maz96, MK98, MCU97, NMMS97a, NMMS97b, RM97, SUZ98, Vin96]. CORBA-Based [MB98]. Core [MDK96]. Corner [Lim97]. Corporate [CF95, Dan94, HS93a, Lea99, Lod93]. Correction [Riz98]. Corrections [Ano99-50]. Correspondences [WCN93]. Corrigendum [Sak92a]. Cortical [WW91]. COSNIX [Kap92a]. Cost [Ber99b, CI97, Pom94b, Yah94]. Cost-Aware [CI97]. Costs [Sim96d]. Could [LO90, Tab94, War99a]. Counterpoint [Kol92, SP94]. Country [McR95]. Couple [SP99a]. Coupled [FBAE^{+}92, PS93, RS91]. CPANic [Hal99a]. CPANorama [Hal99b]. CPU [CS90, FP93, FS96, GGV96, Hel92, JR99, MT93, WWDS94]. Cput [Che97]. cql [Fow94]. Crack [De 90]. Cracker [Che92, Kle90a, Mad92]. Cracker-Case [Mad92]. Crackers [De 90, Mad92]. Craft [Bir95]. Crash [Hid94]. Create [Arn98]. Creating [DBK98, DG94, Han92, Hun97, Lea95b, LRNL97, Man97, SD98b]. Creation [CKDS90, Cur91]. Cricket [Hax96, All99]. CRISIS [BVAD98]. Criteria [Cog90]. Critical [Men93, Rou97a, Spexx]. Critique [Sak92a, Sak92b]. Cro [Tap99]. Cro-Magnon [Tap99]. Cross [JS95]. Crossing [Bar92a]. Crossword [Ano99-43, Ano99-44, Ano99-58, Ano99-56, Ano99-57]. Crushing [VB92]. Cryptographic [CF95, GK99, Gut99, LHB96, NK98, Nys99, SK98]. Cryptographic-Token [Nys99]. Cryptography [Avo98, Dav96, Lac93, Ros96e, Sch96i, dRHG^{+}99, USE96i, Dav94b]. CryptoLib [Lac93]. Cryptosystems [FY98]. CSRG [McK98]. Cultural [Lea96b]. Curing [Lib90a]. Currency [dPG96, Mur99a, Tan96]. Current [JCS^{+}91, Lip95]. Curve [Mil98c]. Curves [Men98]. Custom [EL92]. Customer [Kub92, Nat92, Rus99b, Sma92b, Zwi92c]. Customers [Saf92]. Customizable [AM98a, AF94, BP99]. Customization [Lea90a, RP92, Wel95, WCM93]. Customized [JR91, KK92b]. Customizing [JR91, STW93a, Won93]. Cut [BCC^{+}94, BM96]. Cut-and-Paste [BM96]. Cute [Wan91]. CVS [Ber90, CdR99]. Cybermed [Sal96a]. Cyclic [Fri97]. Cygwin32 [Noe98].

D [Ell97]. Daemons [LeF92, Sum94]. Dagger [Che94b]. Dalek [OCH90]. Dallas [USE91i, USE91l]. Damon [MW98a]. Dan [Dar97c, Kol96e, Kol98c, Mil97c]. dancing [TM97]. Dangerous [Zwi98, Zwi99a, Zwi99b]. Dangers [Rei99]. Dare [Dar95d]. Dark [Hum99b, Kle98a, Kle98b, Mel96b, Ran99]. Dashboard [Cla97]. Data [AS92, AHB^{+}98, AOG92, AUH99, ACO92, BFRS99, CL91b, CL91a, CGCS91, CJC^{+}92, CK96b, DA92b, EBG99, FP93, FST99, FPSW91, GM93, Gra99e, Gut96, Haw90a, HF91, HF96, HN97, How98b, IGP95, IK95, IC97, KPR94, LS98a, LS93, Lu97, NN95b, PGH^{+}91, PCMI91, PK96, Pol97, Puc91a, Puc91b, SW98, Sch96c, SB97a, Sri93, Ste93a, SV90b, SV91, Vo97, War94, WWDF98, kc99a]. Data-Intensive [FST99]. Data-Propagation [IK95]. Data-Race [PK96]. Data-Rates [CL91b]. Database [CS97, CMSvE99, Eng91, Fin94b, Fin95a, Fow94, GLW94, HF91, JV95b, JV95a, Mot91, Pag98, RM97, SO99, SD98b, Shi91, Urb90, Urb91, WS99, Sch97b, Lar91]. Databases [AT92, KA94, SH98b]. Datacenters [YM99]. DataMesh [Wil92a]. Date [NP99]. Day [Kor97a, Roc97]. DB [Pom95a]. DBM [BN92]. DBMS [DeW91, Ska94]. Dbx [Lin90b]. DC [USE90h, USE91k, USE94m, USE99e]. DCE [Seo93a, GS94a, Wiz96, Seo93b]. DCE/OSF [GS94a]. DCOM [BCLC99, HdJB97]. DCS [NWRP^{+}91]. Dead [DeV93]. Deadline [Ano97f]. Deadlines [Ano98-40]. Dealing [Bec99, Bee92, Mar96c, SESS96, WCN93, Ins96f]. Dear [Ano96a, CG99b]. Death [Kol97a]. Debate [Kol99a, OK93]. Debian [Den99a]. Debugger [CB90, HK97, LM94, May90, May92, OCH90, PBS97, Vck98, Win94b]. Debugging [Cop94, DK95, GH93, GL93, Hov91, Kap97, PVL91, RRB99, RBJ91, Kot98]. Deceit [SBM90]. December [USE97h]. Decentralising [HCH95]. Decision [Ske92]. Decision-Support [Ske92]. Decisions [GADG96]. Declarative [FST99]. Decompilation [PW97]. Deconstructing [Lim99b]. DEcorum [KLA^{+}90]. DECstation [KP93a]. Deducing [BM98a]. Deductions [GB97]. DeeJay [MK92]. Deep [Kol95a, SF97b, Vix99]. Defamation [Nas96]. Default [GD91]. Defects [Dav96]. Defending [Kle99b, SP99b]. Defenses [Kle99b]. Definitive [Fla97, Lea96d]. Delay [OT90]. Delayline [IP94]. Delays [AS92]. Delegation [Gré93, Grg93, HSK99, Law94, NL98, RG95]. Delegations [Bee92]. Deletion [Gut96]. Delivering [ZD93b]. Delivery [Zea97].

Delta [OS92]. Deltas [BDR97]. Demand [CIC94]. Demise [Hum96a]. Demos [Uhl91]. Demultiplexing [MD92, YBMM94]. Denial [SP99b]. Denmark [JABA98, Mad92]. Dependencies [Cow99]. Dependent [AHS95, CG92a]. Deploying [CF95, Hed98]. Deployment [Cum99a, Fin97a, SCW^{+}99]. Depot [MWCR90, MWCR91, Won93, CW92b, RM94b]. Depot-Lite [RM94b]. Derivation [GG98, NB94]. Derived [MAH98, PV95]. Description [MR97a, WAKS97]. Descriptions [Dar93a, Dar93c, Fis96b, USE93e]. Design [AEK96, Ano97a, Ano98l, Ano99m, Ano99-31, BCK94a, BMHR91, BCK94b, CGHK95, Che90, CDL99, CSYG94, Cog90, CG90, Var97, Cuc91, DG96, DCM^{+}90, DMST95, FLB98, FN91, FSV94, FK98, GG97, GGI^{+}91, Gra92b, GB97, GWP99, Gut99, HG96, HLM94, HPS97, IM93, IC92, JWS98, JMM^{+}99, Joh95a, KP93b, KK99, Kot98, LKY93, Lea92a, LH97, Lea96f, LC93, Lyc94, MSL94, Maf96, Mar96c, MO99, Mit95, MQ97, Mul97, MWO^{+}93, Nee90, Nes99, NZ99, NN92a, OB99b, Ols93, OW94, PSWL95, PJS^{+}94, Pul99, PHS96a, POS^{+}99, PHS96b, Sab90, SCW^{+}99, ShCR93, SL94, Sch96a, Sch96g, STW93b, Sim94b, Smi91, Sos95, SSCK93, SW91b, TLK99, TMC97, USE94i, USE96f, USE99h, Var95a, Vis96, Wal93b, WIK92, YL98, YLK93]. Designer [Tay95a]. Designers [Hae92]. Designing [Dea90, Gor90, Hai97, Lam95, O’N98, SUK92, WWDF98, CG05]. Desks [Tru98]. Desktop [CR94, Fia95, LBBA99, LFR^{+}98, Lim99c, Uhl93, Zea97, dILM98]. Desktops [KLB98]. Detaching [LM97]. Details [Pom96f]. Detecting [GMNR99, HJ92, Mad92, Pax98, RAM98]. Detection [Ano98g, Ano99x, Ano99-34, BZT99, BBK99, Cas96b, Cas96a, CPM^{+}98, DH99, EM99, For99b, GSS99a, GS99a, HL93, KS94a, LS98a, LPS99, Mel99a, ML91a, PK96, Roe99, RFR99, SU99, SST92, USE99l, ZSB94]. Determination [PR94]. Detours [HB99]. Deutsch [Hac98a]. Develop [Bar98]. Developed [Phi91]. Developers [Ano97d, Cha99, Zwi93d]. Developing [BCR91a, BCR91b, Cae97, CF95, Fin90, Laz98, LFB92, Muf99, Sch94b]. Development [Bea97, BSKH93, BG97, Ber90, Bir94, BCGS91, CSL^{+}97, DBCF95, Fly98a, GMPS97, Har95, HO94, Hen90, Kha98, Kle94, Owe97, Rei98b, SSL96, SW90, Tho96, USE94g, Wal94b, Wyb90]. Device [Cof97, GD96b, GSR93, Hun97, Raw93, RN93, TMC97, TB94]. Devices [Eng92, HS99a]. Devil [Pom96f]. DHCP [LR98a, VW99a]. Diagrams [MM96b]. Dial [BK93, ME99]. Dial-In [BK93, ME99]. Dictionary [Jas96, Joh97a, NV93]. Did [MP93]. DIDDLY [Lar91]. DIDS [SST92]. Diego [Pea93, USE91f, USE93k, USE93j, USE93l, USE94h, USE96k, USE97c, USE98f, USE99f]. Diff [Gra92a]. Difference [Chr93b, Sie96]. Different [Dij96a]. Digest [Hon95a, Hon95b, USE97a]. DigiBox [SBV95]. Digital [AOG92, FN98, FKT^{+}99, Kar93, Pet94, RBB^{+}91, CH97, dPG96, Lar91, MW96]. Digital-Ticket-Controlled [FKT^{+}99]. Digitally [Sim96d]. Digitized [SF97a]. Digs [Gal93]. Dijker [Kol99a]. Dilemmas [Kol94f]. Direct [Rub91]. Directed [Gra92a, Kot94]. Directing [Sos95]. Directions [Sch95e, Sch96b, Str93b]. Directories [CIC94, Sin97d]. Directors [Ano95d, Wil95c, You97a, You98a, You99a]. Directory [AM98a, ED92, Gre99, Hed98, How90, How91, Lea98, MC99, RL91]. Directory-Enabled [Lea98]. Disaster [Dar96e, SESS96]. Discardable [Sub91]. Discipline [Moe92]. Disclosure [Rau99]. Disconnected [HH93, MBM95, SKM^{+}93]. Discontinued [Ros98b]. Discourse [ASB^{+}99]. Discovering [GS94c, WCS93]. Discovery [HS93b, Kre94, SEKN92, ZD93a, SSK97]. Discussing [Kol97b]. Discussion [Mem93, PBEDT97, Sca98, AGF90, ADKN90, BCR90]. Disguise [vR95]. Disk [AOG92, CL91a, DKM94, DKB95a, DKB95b, ES92, Fre97, GK97, Hid94, Kot94, LKHA94, Ric97, RW93, SCO90, Sin97c, SS96b, Swa93, Vanxx, WAP99, Won93]. Disk-directed [Kot94]. Disks [And92, SW96, Van97]. Display [Fly99f, SD98b, Spr96b]. Disposable [ASB^{+}99]. Disputes [AVS98]. Dissemination [Hil96, YGM95]. Dissimilar [Har90, Har91]. DiSTiL [SB97a]. Distributed [AMML96, AAO92, ACFM99, AEK96, AAL92, AHL^{+}91, Arn95b, AGG^{+}94, BL90a, BB^{+}91, Bel99, Ber98, BZM91, Bir94, BSSD96, BA91, BA92, BH91, BD93, Box96, Boy91, BCC^{+}91, IG99, CL91a, Car95, CW92a, CNTS93, CD92, CW96, CMS93, CGCS91, CCGS92, CG90, CG92a, CJ92, CGS^{+}96, Cou95, Cra90, CKDS90, Cur91, DBGL^{+}93, DDD92, DLO92, DCM^{+}90, DAM^{+}91a, DAM^{+}91b, DCWB92, DL93, Den92, Den95, Dil95, DOKT91, DG91, EW90, FCNL94, FHG93, Fle93, FHLW99, FWF^{+}90, FPSW91, FLA94, FK98, FK99, Fuc97, GS96a, GG97, GGI^{+}91, GM93, Gol92, GLW94, GG91, GMS96, Gré93, Grg93, GV91, GWP99, HK95b, HdJB97, HSE96, HHLS90, HCH95, Hop96, HS99b, IEE90, IMC93, ISS97, KCDZ94, KSKD99, KSAL96, KFJ94, KMR91, KC99b, Lar91, LCH92, LMC94, Lu97, Mem93, MSL94, MPS98]. Distributed [Maf93, MB91, MD98, MM96a, MK98, MU98, MQ97, MR95, MLM94, MH92, NL98, NMMS97b, Nes99, NN92c, NWRP^{+}91, PT92, Par90, Par93a, PCMI91, PK93, PS93, PG92, Pre92, PBR91, RA98, Rei92a, Rey92, RAF92, Roz92, SL94, Sch94b, Sch96g, Sco93, Seb91, SM92, SO92b, Shi97, Shu91a, SBM90, SM98a, SST92, SS97, SS94b, STV90, SW95, STM^{+}99, SLC94, Tan95, TSO^{+}96, TJM^{+}99, TJH^{+}94, USE91j, USE92l, USE93j, Ull97, VB92, Vau90, VMMD98, Vin96, Wag90, WWWK95, WW97, Wei95a, Wel94, Win94a, WWW95, WRW96a, WRW96b, YKB94, YCO94, YL98, YTJ91, ZTM96, ZSB94, PT93b]. Distributing [DFG^{+}96]. Distribution [BCK94a, CK96b, Cou97, Gal96, LCL98, MJS97, MGZ93, MR97b, NC94, OG95, SM91, SK99, SSZ^{+}93, VCV92, dSdS98]. Distributions [BCK94b]. Distributor [BH99]. Diverse [SSL^{+}99]. Dixie [Gai94]. DMA [Tho98b]. DMZ [Mel97c]. DNS [AM97, CB96, Col93, Gal96, Kra99, Lar99, LeF99, RH98a, Rei99, Vix95, deL99]. Do [CMB^{+}98b, Eva94b, Fri97, Ins97b, Lim99d, Nat96, Pre97, Rue96, SS99, Tru98, hvESV^{+}97, Mor92]. Doc [San90]. Document [BNMK99, Chr97b, DHR97, Hei96, HK99, KL93]. Document-based [HK99]. Documentation [Mue96d, She92b, VZ91, WS99]. Documenting [CI93, Sin97d]. Documents [Bvv91, CSM97, HS91b]. Does [Car90a, PW97]. Dogfight [Chexx]. Dogzilla [Chexx]. Doing [AM97, Pom93d]. doit [Fle92a]. Domain [Ano97i, Ano98h, Ano98x, Ano99f, ABB^{+}97, BSS^{+}95, BSSW96, DVPL92, FNP97, Fuc97, Gil96, Har90, Har91, Hud97, KS97, LM99a, Maz96, NAOP99, Rou97a, SM91, SV94, Str99, TMC97, USE97b, USE99a, WSB^{+}96, Bel95, RH98b]. Domain-Specific [Ano97i, Ano98h, Ano98x, Ano99f, Hud97, KS97, LM99a, USE97b, USE99a]. Domains [Bru98, BGÖS98, Car99b, HCC^{+}98, Lei98, MR97a, Pom96a]. Don’t [BTS^{+}99, Hal99a, Mel96b, Zwi93b, WD95]. Door [Zwi94g]. Doors [HS93a]. Dossier [MCLY94]. Double [Spi99]. Double-Buffering [Spi99]. down [DKB95a, DKB95b]. Download [JPR96]. Dozen [Har96a]. DP [Arn95b, LSR95, PSJL97]. DPMI [GMR93].

Dr. [Kol98d]. Draft [BC91, Sto98e]. Drag [YD94]. Drag-and-Drop [YD94]. Dragons [Bel92]. Drawing [KH97a, RL98]. Dreams [Kol95b, Lan91, Som97]. Drinking [CHK^{+}98, LOM94]. Drive [LKHA94, Vanxx, Zwi93b]. Driven [KTP^{+}96, KFJ94, MCLY94, Sat92, SO92b, Sjö99, MR96, SGG^{+}99, Den95]. Driver [Cof97, CRH99, DG94, GD96b, LR98b, TB94, vL98]. Drivers [Bel90, GSR93, Hun97, Raw93, TMC97, TIF^{+}97]. Driving [All99, Gee98]. Drop [YD94]. Dropping [VB92]. DRUMS [BH91]. DSA [CMS93]. DSL [SeaBP99, USE99a]. DSM [FS94, PBS98, SH98a, SB97b]. DSMs [IS99]. Dual [Jas96, Mar96c]. Dual-workfactor [Jas96]. Duckling [Cox98a]. DUEL [GH93]. Dues [Ano98-36, Ber99c, You95a, You96c, You97e]. Dumb [Bor92]. Dummynet [Riz98]. Dump [MK92]. DUNE [BDKM93, PA90]. DUNE-iX [BDKM93]. DVSM [SWS92, WSS92]. Dynamic [Bar98, BL90b, BP99, BPT98, Bot98, Bvv91, Cur94, DHR97, EM99, EN96, GSPW98, HG98, ISS97, IC97, JS97, Joh95a, KYBR99, MDC95, NH93, PB96, PS96, RLBC94, SAYZ99, TM99, Uhl90, WL95, WN98, YCO94, ZSG92]. Dynamic-Binding [YCO94]. Dynamically [HCL95]. Dynamically-Linked [HCL95]. Dynascope [Sos95].

E-Commerce [MMS97]. E-mail [Bar94c, SBRH96, Hol97a]. Ear [Lanxx]. Early [MPZD98, NN95b, PBR91, Sal96i]. Earth [Sal96a]. Ease [Tay97a]. Easier [Hid94]. Easy [Bea96, MK95, Opp96, Kol94h]. Eater [Ins96a]. Eats [Dam95]. Eclipse [BGÖS98]. Ecology [MS95, MET98a]. Economic [Var95a]. Economical [Mof92]. Economics [Mue95d, Roo97]. Ed [Jen98, Joh97a, Kno97, Pot97, Han97]. ed. [Uma96c]. EDI [JABA98, SPR96a].

EDistribution [LR98a]. Editable [Phi91]. Editing [IS91, KTBL93, RL98]. Edition [Spe93b, Bar94d, Bea98, Fla97, Uma97b, Quo91]. Editor [Ano98-35, Ano98-37, Bar96, Car98a, Dam98, Dar95a, Dar95b, Dar95c, Dar96a, How95, Kol94b, Kol94d, Kol94c, Kol95b, Kol95c, Kol95d, Kol95f, Kol95e, Kol96a, Kol96b, MW98a, McD93a, McD93b, McD94, Pee96, Per98b, Tho90b, Ano99-50, Ano99-45, Ano99-46, Ano99-47, Ano99-48, Ano99-49, Cal98, Che98a, Cox98b, Gar98, Rei98a, Sal98a, Van98, Vix98]. Editorial [Avo96, Dar96f, Dar96h, Hon94, Pre95, Ree95, Sch96f]. Editors [HS91b, KP93b]. Edits [Sin97a]. Education [Sto96b]. Effect [BS93a]. Effective [And92, Cou96a, Hal98a, Hal98b, Hal98c, Hal99c, Hal99b, MR97b, SW98, Yig97]. Effects [BS96, Van97]. Efficient [ASSB99, Alf94, BNMK99, Bed90, BS92c, CG92a, Cra94a, DR99, ED94, FLA94, IGP95, Koe92, KWB^{+}98, MD92, MS93a, MMBC97, OS94, SH96, SKS90, SUZ98, TD92, Vo97, YL99, YBMM94]. Efficiently [Jas96]. Effort [MS96a, Sto96e]. Egg [Joh93d]. Eighth [Ano98p, USE94h, USE99e]. ELAN [Bra99]. Election [Ano97k, Ano98-34, Ber99a, Dar96b, Mil97d, Wil95c, Wil95b, You97a, You98a, You99a]. Electra [Maf93]. Electronic [Ann97, Ano98a, Ano98m, Ano98-31, Avo98, Bah96, BGH^{+}95, BGJY98, Bis92c, CST95, CBB97b, Cuc91, Cur99, ED92, GS96a, dPG98, GD96a, HS96, HTK98, HTWW96, Hon95a, Hon95b, HS99a, JABA98, KS98, Lee98, LO95, LHB96, Man95, MW98b, PHS96a, PHS96b, SV94, Sch95g, Sim96d, SP96, ST96, Tan96, TW96, USE95g, USE96h, USE98b, Wri96, YT95]. Elementary [Cas96b, Cas96a, CG99b]. Elements [Haw90b]. Eleventh [USE97c, Ano98y]. ELF [Arn90]. Eliminating [MR96]. Elimination [SBRH96]. Elk [LB94]. Elliptic [Men98].

Emacs [Chi97, Bra92b]. Email [Cam94, Ins97a, Mil99a, Pom97c, Sin98c, ZD93b]. Embeddable [Ous90b]. Embedded [Ano99y, Ano99-30, BBG^{+}99, Ell97, GK97, JB99, KR99, LS99, LM99a, PBS97, SO99, USE99k, Wil97a]. Embedding [ATS94, Gra99c]. Embraces [Ts’97]. EMERALD [NP99]. Emergence [Woo94]. Emergency [Mel96b]. Emerging [Hua98, Mar91a]. Empire [Dew97]. Employment [All96, Cla96]. Emulation [CWS90, Dra91, GD96b, IP94, JCS^{+}91, PBS98, RHG98]. Emulator [DBCF95, Pat93]. EMV [VW99b]. Enabled [Lea98]. Enabling [Gon96, Leh98, WU97]. Encapsulating [Lin92]. Enclaves [Gon96]. Encrypt [WT99]. Encrypted [Jas96]. Encrypting [Bla94, LW99]. Encryption [BRW99, BB95b, LT91, LW99]. End [FSV94, Hub98, Mil98b, NLZ99, vR95]. End-User [FSV94]. Ended [Fer95]. Endpoints [YBMM94]. Ends [Vo96]. Endured [Che92]. Enemy [HS91c]. Energy [IGP95, NUO^{+}97, WWDS94]. Enforcement [BSS^{+}95, BSSW96, WSB^{+}96]. Enforcing [Pov99]. Engine [Bar98, Bur95, CLL97, FD98, Puc91b, VcC97]. Engineer [Lea96f, Lie99]. Engineering [ACFM99, AG97, Bir90, BGT98, Ker95, MBM93, NAOP99, PK91, Rou97a, SCW^{+}99, SG97a, WD95, Eva94c]. Enhance [Tay97e]. Enhanced [Bis92c, LB98, PM95b]. Enhancements [Lab99]. Enhancing [McC98g, Zwi91a]. Enough [Zwi98, Zwi99a, Zwi99b]. Ensemble [BVG^{+}97]. Enterprise [AW99a, BD93, GZ94, Gru96, Hei99, Her97, Lar99, Laz99, Lie99, LBSL97, LCL98, Pre97, Woo98, WWDF98]. Enterprise-specific [GZ94]. Enterprise-Wide [Gru96]. Entire [MW94]. Entity [MM96b]. Entity-Relationship [MM96b]. Entrepreneurship [RK97]. Environment [Abb94, AHL^{+}91, BS92a, Bis92a, Bla91, Bog95, BGT98, BCGS91, Cas95, CMB^{+}98a, CCW97, CBE93, CG92a, CV96, CJ92, Cra90, Cuc91, DBC^{+}99, DZL97, Dil95, DG91, EBG99, EL94, Eva94c, Fur91, GF95, GMR95, GWTB96, GMR93, GMS96, HM92, Har94, Hau94, How90, How91, JKS^{+}93, Kan91, KDF^{+}93, Kap94, KTBL93, KOO95, Leb97, Lin94, Mah91, Men93, ML90, MK93, MMBC97, MZA96, NR94, Pal92, PG92, RAV95, Rid94, Rou97a, Roy93, RH93, SSH93a, SCW^{+}99, SM91, Sat92, Sat93b, SKM^{+}93, Sel91, Sma91, SW90, TLK99, TSO^{+}96, TM97, VCV92, WYOT93, Wag90, Wal97b, Wal98, Wil99, WCM93, Won93, Zwi99c, dILM98, SW91b, Col92a]. Environments [AD93, BSKH93, BSSD96, CW92b, DK95, EL92, GSTY96, NL98, RAF92, Ske92, Sri93, SKAM99, Bir94]. Episode [CAK^{+}92]. Equations [Van91]. Equivalence [SAYZ99]. Erratum [Ano98-35, Per98b]. Error [Moe92, Riz98]. Errors [DT97, HJ92, ML91a]. Escalator [Mel98]. Escape [Dew97]. Essence [HS93b]. ESSENSE [BHJV92]. Establishing [AMP98, Byr93, Ell96]. Estimation [MT93]. Etch [RVL^{+}97]. Eternal [AMP98]. Ethernet [Bec99, cC99, McR95, Pre98]. Ethical [Kol94f, Lan94a, Lan94b, Kol94g]. Ethics [Ano97l, DK99, Dar97c, Kol94f, Mil97b]. ETI [BH99]. European [Ano97o, Mur99a]. EurOpen [Dun91a, USE92f, Wal92c, Wil91a, Wil91b, Wil92b]. EurOpen.SE [Sae98]. EurOpen/USENIX [USE92f]. Evaluate [EWCS96]. Evaluating [KS94b, WR98, lWS99]. Evaluation [BS93b, BCVC91, DMST95, Jae95, Kev93, MSL94, NLG^{+}99, SL94, SSZ^{+}93, SS94b, TN91, WT99, ZIL96]. Even [Zwi92c]. Evening [Che92]. Event [CCK^{+}91, DFGG97, Ess90, MB98]. Event-Based [CCK^{+}91, Ess90]. Events [KL96, Sal98a, WWWK95]. Ever [Eva94b, Hil97, Mil93, Wal95c].

Ever-Changing [Mil93]. Ever-Growing [Eva94b]. Everyone [Cam94, Cox97c, Gwy90]. Everything [Wal95c, MAM95]. Evidence [PV95]. Evil [Dew97]. Evolution [BZM91, Bra99, Cat90, CMB^{+}98a, FHG93, HS93a, Lin90b, MO99, SCD90, Str93a, SCD94, TB99, Wal93b]. Evolutionary [IFT^{+}91]. Evolvable [GS94b]. Evolving [FL94, Kolxx, Ros90]. Examining [Cre97]. Example [LH97]. Exception [CFLM92, FK92, Kin91, KS90, KS93, Mue96a, Sel90, Tie90, Tie93a]. Exception-Safe [Mue96a]. Exceptions [Gau92]. Exchange [Cat90, Jas96]. Exchanges [DBGM^{+}98]. Exclusion [MJ92]. Exec [IJ97]. Executable [JPR96]. Executables [CGLR97, GL97, RVL^{+}97, Tau91]. Executing [GSR93, Raw93]. Execution [CG99a, DLVW98, FHL^{+}99, Gam93, Mas91, PS96]. Exercise [Tau91]. Exhibition [Rot95]. existing [LBSL97]. Exiting [CFSW99]. exmh [Wel95]. Expanding [Car99a, MS98]. Expansion [Mil98a]. expect [Lib90a, Lib90b, Lib91a, Lib91b, Lib95, Lib96a]. Expected [KC96]. Expenditure [Kol94a]. Experience [ABB^{+}97, BCF^{+}93, BGH^{+}92, BC94, BCC^{+}91, BCR91b, CMB^{+}98b, CLD^{+}97, CFH^{+}93, Duc91, Fre91, GKS90, Hun95, JD94, JR99, Kam93, Ker95, MRK97, NUO^{+}97, NKM93, NP99, Pin92, PBR91, Ros96c, SKM^{+}93, Ska94, SC92, WRJ92, ZWM^{+}99]. Experiences [ACO92, BFM92, BCR91a, CDK^{+}99, DCWB92, FN91, FWF^{+}90, FPSW91, FK92, GS94c, HHW92, JCS^{+}91, KDF^{+}93, KGL^{+}91, KS94a, LPS99, Lun90, MGZ93, Mot95, NWEF92, PST^{+}92, Rei92a, TP91, Tho96, Tor96, USE91j, USE92l, USE93j, UKGS94, Wyb90, MAM95]. Experimental [BC91, CDM90, Fle93, IEE90, Mem93]. Experimentation [MS93b]. Experiments [NKT93]. Expert [Kol95a, Kur90]. Expires [CMB^{+}98b]. Explicit [GK97, MP96]. Exploiting [CL91b, FP93, GK97, KSU92, NMMS97a, SAYZ99]. Explorer [LN95]. Exploring [Far96f, KL97, Lib95]. Exporting [CIC94, ST93]. Expressions [Hum97a, Hum99b, Uma97a]. Extend [Dou93, Ts’97]. Extended [Ben93, MM96b, Str90]. Extending [AISS97, BMK96, Gau91, MA90, MS98, Mil99d, SKCB97, SH98a, Sel90, Sun97, Tho98a, WL95]. Extensibility [GPRA98, Lea94, Mil96b]. Extensible [AM98a, AP97, BKC94, Bis90, CCK^{+}91, CV96, FB96, HK97, Kor94, May92, PB96, Rhe98, SB98a, SMSM98, SSL96, Sma97, Tho90b, WL98]. Extension [Car92b, DBK98, EN96, KTN93, Lib97b, Par99a, RS91, SS96a, LB94]. Extensions [CMSvE99, Den92, HS95a, NYM91, NYM92, NL96, SESS96, SL95, SC96, Ull97, Web93, Sto98e, Chi97]. Extent [MK91]. Extent-like [MK91]. External [AR94, Boy91, KN93b, LCC94, MA90, Sub91]. Exu [RG95]. Eye [Gir99b]. Eyes [Kol95e].

Facility [BDM99, Bla90, Cou97, DDD92, MB91, MCLY94, MKC92, NKM96, SS97, SSBK98, Byr93]. Fact [Sat93a]. Factoids [Wil94]. Facts [MS97]. Fads [Ous91a]. Fail [MD95]. Failed [Pom96c]. Failover [Bar97]. Failsafe [Pul99]. Failure [Lea96f]. Fair [Ess90]. FALCON [GV91]. Fall [Joh93d, USE90a, USE91a, USE92a, USE93a, USE94a, USE95a, USE96a]. False [BS93a]. Families [Pol94]. Family [BGH^{+}95, BC94]. Fancy [Dar96g]. Fangled [Sno96]. Fantasy [MS97]. Faq [How98a]. Faq-O-Matic [How98a]. Farewell [Joh96a]. Farm [BGT98]. Farms [NLZ99]. Farrow [Car98a]. Fast [BS92a, Bar91, FST91, HF91, HS91d, KW96, KV91, Lie92, MD95, O’D90b, OT90, OOM^{+}93, OLB93, Ous90a, PKA98, RAC97, Sch96c, SM99, SU99, SCB93, WM92]. Faster [Car94, Ous90a, Spe92, Sto92, Sto93a]. Fault [Bab90, BMHR91, CL99, CN90, CBL90, Ede93, HCK^{+}98, KMBT92, KSAL96, Maf95, MCU97, NMMS97a, SL96, SC93]. Fault-Tolerance [KMBT92, Maf95]. Fault-Tolerant [Bab90, CN90]. Favor [NN92b]. FAXes [Pin92]. FDDI [NN95a]. Fdist [SM91]. Fe [USE93h, USE94k, USE98c]. Fear [Ran99]. Feasibility [NFQ99]. Feature [FKSV94]. Feature-Based [FKSV94]. Features [Ken91, Laj96, MS98, Shi91, Sin98a]. February [USE99h, Ano99-58]. Federation [BPT98]. Feedback [GSPW98, MP90, SGG^{+}99]. Feedback-driven [SGG^{+}99]. Ferret [KL93]. FFS [SS96b]. Ficus [GHM^{+}90, PGH^{+}91, RHR^{+}94]. FidoNet [Cro98]. Field [Zie93, RM90]. FIFS [Alm99]. Fifth [Ano98u, USE91f, USE95f, USE97d, USE99f, Bai97, Far95d, Pie94b]. Fifty [Mel96a]. Fifty-one [Mel96a]. Filaments [FLA94]. File [AISS97, All97, AHB^{+}98, Alm99, AEK96, AOG92, Arn90, BL90b, Bar94c, Bel96a, BBS95, BEM91, BD96, BHS95, BA91, BA92, Bla94, Bor98, BM96, CFL94a, CFL94b, CS92, CS94, Cat90, Che97, CAK^{+}92, CR94, Cou97, DWAP94, DMST95, EP92, EP93, FG91, FC96, FHKR93, Fow94, GP94, GM99, Gre91, GA94, GWP99, GHM^{+}90, HS93b, HO92, HHLS90, HLM94, HMF^{+}99, KNM95, KLA^{+}90, Kei90, KW93, KP97, KSS93, KV91, Kot92, KL96, KPR94, KS95, LD97, LIN^{+}93, Man94, MW94, MS92, MKB90, MK93, MK95, Min93a, ML93, MH92, Nad92, NKM93, Neu92b, Neu92a, OT90, Oka92, Ols93, OS92, PT92, PT93a, PGH^{+}91, Pea92, PKA98, RHR^{+}94, RL91, Roo92, Roy93, SM91, Sel95a, SBMS93, SSB^{+}95, SO94, SBM90, SS94b, SKS90, Swa97a, Swa97b, SDH^{+}96, TD92, TSM95, USE92h]. File [Vanxx, WAP99, Wel91a, Wel91b, Wel91c, Wel92, Wel94, WK93, ZD93a, MK91, SKS92, Per92]. File-Based [TSM95]. File-System [KL96]. File-Systems [BM96]. Files [BS92c, Cra94b, GK97, Kar90, KSU92, Lam91b, Man94, Mof92, PV95, Rom91b, SSB92, VC92]. Filesharing [Ran97]. Filesystem [BS98, Cat92, Cou96a, FM94, Hen90, HR92, HAB^{+}98, IHR99, LBLM90, LPLF91, McK95, Sum94, Web93, Fri97]. Filesystems [CFSW99, SGM91]. filetsf [Sel95a]. Filter [CB96, MJ93, Pla93]. Filterfresh [BCH^{+}98]. Filtering [Cha92c, CHS91c, IGP95, LP94, Mol95, CB96]. Filters [NZ99]. Final [Pie99b]. Financial [Ano95a, Ber99c, Gee94, Gee98, SPR96a]. Finding [All95, Man94]. Fine [CLS92, Ede93, FLA94, IS99, JLI98, MP90, TM99]. Fine-Grain [Ede93, FLA94, IS99, MP90]. Fine-Grained [CLS92, JLI98, TM99]. Fingerprinting [Hei96]. Finite [MSS98]. Finite-State [MSS98]. Fire [CHK^{+}98, RL99]. Firehose [LOM94]. Firewall [BK93, Cal96, For99b, Kah95, Lim99d, Mat99, Muf95, Sch97a, TW93, Zwi99c]. Firewall-Router [Cal96]. Firewalls [Bel99, Coo99, HSOS99, Ins96d, MM96a, RA94, Far94]. Firmware [Bra92a]. First [Ano99-33, Hon95a, Ila99, Kre97, Men94, Mil98h, Mit95, Sub98, USE94i, USE95g, USE99g, New95, Sub98]. Fish [Kol97h]. Fishing [CK96b]. Fit [Her97, Ism98]. Fits [Her97, Lib90a]. Five [BKS90, Wil98]. Flash [KNM95]. Flash-Memory [KNM95]. Flask [SSL^{+}99]. Flat [Fow94]. Flattening [Mil98c]. Flexibility [Bar97, MB96, Wel95]. Flexible [ACTW96, CHS91b, Eir94, ES95, Fur91, Gra96, Gre91, HS91a, HMS94, Hil96, JY96, KN93b, KS95, MPH99, MMBC97, Mul97, OLB93, RH98a, SBM90, VDAA99, WW94]. Flexibly [JPR96]. Flight [Lal98]. FLIPC [BSSD96]. Flooding [Ins96f]. Floors [Lip95]. Flow [RFR99]. Fluke [FHL^{+}99]. fly [Lew96]. Focus [Dar96f]. focusing [USE96i]. FOG [GS90]. FOG/C [GS90]. Foiling [Kle90a]. Follow [Moh99b]. Follow-Up [Moh99b]. Food [Ins96a]. Force [CD98, KI99, Kle98a, Kle98b]. Forecasting [Swa93]. Form [ABB^{+}97, Eva95c, Sch95e]. Form-based [ABB^{+}97]. Formal [ALLO92, Kev93, LaP94, SA98]. Format [Nys99, Sha97, PT93b]. Formats [GM93, Pom95b]. Formed [Pov99]. Forming [Hec96]. Formula [IK95]. Fornax [Hal94]. FORTRAN [Spe93a]. Fortress [Cra94a]. Forum [Gri97]. Forward [BGT98, Riz98]. Forwarding [Pom96a, WYOT93, WGC^{+}95]. Foundation [HBvdRA98]. Four [AW99b, Kol96j, Pie99b]. Four-Star [AW99b]. Fourth [Kar96, USE90f, USE96e, USE98c, Ros96i]. Fragmentation [McC99d]. Fragmented [GS90]. Frame [MM93, Tay97a]. Frames [Tay97e]. Framework [AM98a, Alm99, AVS98, AF94, Bah96, BN96, BKC94, CI93, Cho98, Dil95, FN98, HK99, Hol97b, JHC94, Lae95, Lau95, MB96, Mah91, MWCR90, MWCR91, MPH99, PHS96a, PHS96b, RA98, Sch94b, Sch99, Shi97, SB98b, Tan96, Tho98b, ZPMD97, DBK98]. Frameworks [CIM92, CSYG94, Dea90, MSS99]. Franca [MR95]. Francisco [All92, USE90j, USE92j, USE94l]. Free [AK94, Ano99-45, BBD^{+}96, Noe98, Sch95d]. FreeBSD [Lei99, LR98b, Mat98, SS97]. FreeNIX [USE98d]. Freeware [MGM^{+}98, ZWM^{+}99]. Fremont [WCS93]. Frequency [Zwi92a]. Frequently [BS97b, LC99, SW96]. Freudian [Poo94]. Friend [Gra98c]. Friendly [Cox97d, CJ92]. Friends [Joh93a]. Frigate [KP97]. Front [Sch95c]. Frontend [NN93a, Sch97b]. Frozen [Kar90]. FTP [Pom99, Rot90, SSH93b]. Fulfilled [Kol95b]. Full [BNMK99, Kol96i, MLL95, Rau99]. Fun [Joh90, Pol96, Pom95b, Wil94].

Function [PBF^{+}96, Sch90]. Functional [Bah96, Dar96g, MLL95]. Functionality [CD94b]. Functions [HB99, Lae95, Lau95, Spe92, TIF^{+}97, HR93]. Fundamentals [Hof99]. Funding [Den99b]. Future [BG99, Bla99, Dar96d, Dar97b, Fle93, Gee98, Ins96b, Jon91, Kha98, Mem93, Mit96, Mit97, Mue95c, PM99, Qua98, Rau99, Sch95e, Tor96, VZ91]. Future-Adaptable [PM99]. Futures [AC97, Sca98, Sun97]. FX [CH97].

GA [USE91j]. Game [Hax96]. Ganymede [AM98a]. Gap [Lar94]. Garbage [Det92, DK95, ED94, FS94, Nil94, Sel90, Van90]. Garbage- [DK95]. GASH [Abb94, Abb95]. Gatekeeper [SW92a]. Gateway [Car90b, Che90, JJ90, AT92]. Gateways [Bec99, CB96]. Gatherer [Tap99]. GAUD [Urb90, Urb91]. Gbit [MBH94]. Gbit/sec [MBH94]. Geant4 [NUO^{+}97]. General [FN98, GS98, Hal94, HO94, OOM^{+}93, Vo97, WUB^{+}90]. General-Purpose [Hal94, FN98]. Generalized [ALO90, GGI^{+}91, JCS^{+}91, RLS^{+}97, YD94]. Generate [Che94b]. Generated [MSS99, SHM93]. Generating [Sin98b, Sin99a]. Generation [CG99a, Fin95a, GMR95, GLOR91, Gut98b, HS95b, KW93, Lar94, NN96a, OSS^{+}92, Par95, RDM95, Sim96b, SB97b, Spi98, WK93]. Generator [CH95, GS90, Sch90]. Generators [GKS90]. Generic [Bah96, HS95a, Lin90a, MBM95, SJ95, Tom98, Wei95a]. GENESIS [WUB^{+}90, AKS98, SSZ95]. GeNMSim [GNMR97]. Genome [SMD^{+}94]. GeNUAdmin [Har94]. Geographic [MLL95]. Geometric [Mil96b]. Gesture [Rub91]. Get [Gra98c, Gru96]. Getting [Fri97, Ins95a, Kol94c, Ous90a, Zwi94k]. Gig [Mue95e]. Gigabit [CF94, HJK^{+}92, Pac93, Pre98]. Gigabit-Networked [HJK^{+}92]. Given [Den99a]. Giving [Saf92]. Glass [Dam95]. Glimpse [BG99, MW94]. Glish [PS93]. Global [AISS97, Cat92, GMR95, Neu92b, Pik91b]. Globally [Mot91]. Glory [Men94]. GNOME [dILM98]. GNU [Bra92b, Chi97, BS98, Jen97, OCH90, Ram94, Tie93b, VWM98, Sec95]. GNU/Cfengine [BS98]. Go [Dij99a, MP93, Zwi94i]. Goal [Mil99d]. God [Ros94]. Going [Chr90b, GMPS97]. Gold [Ano97o, Den98, Kol98k, Kol98l, BCC^{+}97]. Golden [Joh93d]. Goldsmith [Kol98d]. Good [Cam93, Den98, Hal98c, Kol95c, Kor97b, Mue95e, Mue96c]. Goose [Joh93d]. Got [Hed98, Jac94, Lea99]. Gothic [PBR91]. Gotta [Han92]. Governance [Hec96]. GPERF [Sch90]. GPS [Den95]. GPS-Driven [Den95]. Grace [Ree97]. Graftals [Lanxx]. Grain [Ede93, FLA94, IS99, MP90]. Grained [CLS92, JLI98, TM99]. Grammars [SB99]. Grant [Cow99]. Grants [Ber99f, NN92b]. Granularity [CHS91a, MS94]. Graph [NV93, Sch94a, Sch94c]. Grapher [Oet98b]. Graphic [Spr96b, SW90, Tho90b]. Graphical [HK97, Hir97, JMM^{+}99, PN99, SMSM98, SSL96, Uhl95]. Graphically [She92b]. Graphics [BH95, Jef94, FEBM97, Ker82]. Graphs [Che94b, EN96, Pla93]. GraphTab [Hor92]. Grasshopper [DVdB^{+}94]. Grateful [LFB92]. Great [Kol99a]. Greatest [Mad92]. Greetings [O’D90a, O’D94a, O’D94b]. grep [Tay97b]. Group [Abb94, Dun90c, Dun91b, Dun91c, HW93, KP93b, Maf95, Maf96, MPH99, MCU97, Urb90, Urb91]. GroupDesign [KTBL93]. Grouping [GK97]. GroupLens [MRK97]. Groups [Sie96, SMDP97, Var92a, Var92b]. Groupware [AL93, Ros96h, SMDP97]. Growing [Eva94b, Kol97c]. Growth [CMB^{+}98a]. GTrace [PN99]. Guaraná [OB99b]. Guaranteed [BH99]. Guaranteeing [For99b]. Guarantees [PK96]. Guard [McK93b]. Guarding [Cra94a]. Guerrilla [HW93]. Guest [Avo96, Hon94, Hov91, Ree95, Sch96f]. GUI [Hop96, Mul97, NUO^{+}97, Sch97b]. Guidance [Lon95]. Guide [BB^{+}91, BCF^{+}93, Bar94c, Bar92c, Cof97, Fla97, Gsc95b, Gsc95a, Gwy90, Har97b, Lea96d, O’B99a, Seo93a, Sim94a, Spe95b, Uma96a, Zie93, CFH^{+}93, Col91a]. Guidelines [HS93a, Sma90, Sma91]. Guild [Ano99v]. Guile [Lor95]. GUIs [Wal97a].

H.323 [AW99a]. HA-NFS [BS92b]. HACC [ZBCS99]. Hack [Bow98, Wan91]. Hacker [Mud99, Rad97, Rod93, Far94]. Hackers [Bru99b, Joh97a]. hacking [Muf95]. Hamlyn [BJM^{+}96]. Hancock [BFRS99]. Hand [BF99]. Hand-Held [BF99]. Handbook [Uma96a, Van99, NN95a, Cuc90]. Handing [Mue96a]. Handle [Oka92]. Handler [PB97]. Handling [AVS98, BM91b, CFLM92, DCWB92, FK92, KS90, KS93, Koo93, Lib94, Moe92, Sel90, Tie90, Tie93a]. Handoff [BB95a]. Hands [Har99a, Lea96a]. Hands-On [Lea96a]. Handwriting [Kem93]. Happened [Cam93]. Happy [Kam98a, Roc97, Sto98a]. Hard [CJ99, SS99]. Hardening [Rot90]. Hardware [Gra98b, Gsc95a, McR93, Ous90a, RLBC94, SH98a, Sel97c, SA98, Wal95d, WB93]. Harissa [MMBC97]. Harmful [Sim99, SC92]. Harmony [FSM^{+}91]. Harness [Ama98]. Harnessing [MPS98]. Hash [Sch90]. Hashing [SY91]. Hater [Dam95]. Haystack [All95]. HDTV [IS91]. Heels [Gal93]. Heimdal [DW98b]. Held [Ano97m, BF99]. Hello [PT93b]. Help [EL92, Eng93, Gra98c, Tru98]. Helper [GWTB96]. Helping [Cha99, EL92]. Helps [Car99a, Den99b]. Here [Bas98, Gru96, San93]. Hermes [BL90a]. Heterogeneity [NWEF92]. Heterogeneou [Cou97]. Heterogeneous [And91, CCW97, CBE93, DVPL92, EW90, EBG99, Fis96c, Gré93, Grg93, Har94, KTBL93, MK92, MU98, PE92, SM91, Sat92, Sat93b, Sri93, TSO^{+}96, Vas91, Wag90, Wal95d, dSdS98]. Heuristic [BHS95]. Hide [Det91, Man91]. Hierarchical [CDN^{+}96, FL96, GM99, GGV96, YADL99]. Hierarchies [Mar96c]. Hierarchy [Ken91]. High [ATS94, AAL92, And94, Ano92, Ano97m, ACTW96, BS92a, BPR92, CL91b, CL91a, CK96a, CJC^{+}92, ES95, FHLW99, FSV94, Gil90, GH93, Hae95b, HPS97, Kel99, Kor94, Laz92, LCC94, Lee94, LPJI98, MPS98, MK95, MFM95, MZA96, NUO^{+}97, NH93, NLZ99, Par94, Ped90, Pol94, PHS96a, PHS96b, RAC97, SHAS95, SU99, SD98b, SO94, Spr96b, SCD94, TCI98, TJH^{+}94, TB94, USE94f, USE94k, Zea97, ZS99, vR95, Pax93]. High-Bandwidth [MK95, SCD94]. High-End [NLZ99, vR95]. High-Level [And94, FSV94, GH93, Kel99, SU99]. High-Performance [BPR92, CK96a, CJC^{+}92, FHLW99, LPJI98, PHS96a, RAC97, TCI98, TB94]. High-Speed [MPS98, MZA96, PHS96b, USE94f, SHAS95]. Higher [Lae95, Lau95, Puc99, HR93]. Higher-Order [Lae95, Lau95, Puc99, HR93]. HighLight [KSS93]. Highlights [Ano96b, Eva94a, Kar96, Met95, Par93b, Par93c, Rub98, WCR99]. Highly [BEM91, BS92b, CBB97b, CW96, HK95a, Kap94, LBLM90, NKM93]. Highly-Parallelized [LBLM90]. Highway [Sch95g]. Hijacking [HHS92]. Hildebrand [Kol98c]. Hint [CG99a]. Hints [PA90, SH96]. HiPEC [LCC94]. HIPPI [CF94]. Hire [Kol94j]. Historical [Sch97a]. History [BNMK99, DGT96, Eva93a, Hir97, Joh98, Kap92a, Laz97, NN95b, SC93]. Hit [LPJI98, RAM98]. Hit-Server [LPJI98]. HLFSD [ZD93b]. hobgoblin [RL91]. Hoc [KCVP95, Fuc97]. Holes [Lea97]. Home [CIC94, Dar98c, Mot95, ZIL96]. Home-Based [ZIL96]. Homed [Swa92]. Hook [Ros96d]. Hooked [Lea96c]. Hopper [Ree97]. Horror [Ano96c]. Horse [Dar99c]. Horses [HS93a, TW96]. Hose [CHK^{+}98]. Host [BBP99, Det91, Rod93, Spe96, Spe97b, TSO^{+}96, ZL96, SV90a]. Hostile [GSTY96, Lea97, SKAM99]. Hostname [Gol99]. Hosts [JR91, KMS^{+}93, Zwi92d]. Hot [BCH^{+}98, Kle99c, ZD93a]. HotWire [LM94]. House [Eps92]. Houses [Dam95]. How-To [Bin99, Don99, Mil99e, Pom99, deL99]. HP [SMPR98, WIK92]. HP-UX [SMPR98]. HPP [DHR97]. HSM [CFSW99]. HTML [CSM97, DHR97, Fin95a, Lea95b, Lea96d, WU96]. htmlview.tcl [Fly99f]. HTTP [BCE98, TL96]. Huge [Cou96b]. Human [Lea96f]. Humanistic [Man98]. Humans [Bug96]. Hume [Kol96d]. Humor [NZ93a]. Hungry [DKM94]. Hunter [Tap99]. Hunter-Gatherer [Tap99]. Huntsville [IEE90]. Hybrid [AG97, GG91]. Hydrology [CS97]. Hydrology-Meteorology [CS97]. Hype [Kol96c]. Hyperlinks [Duc99]. Hypermedia [Mar91a]. Hypertext [Fly99f, San93]. Hypertools [Bos96]. HyTime [Mar91a].

I-CAMP [Ber99b]. i.e [USE98f]. I/O [ACG^{+}98, BS96, BC91, CL91b, CL91a, CG99a, FP93, FGB91, KV91, Kot94, KSU92, McC98a, MDK96, PDZ99, RvIG98]. IAW [Ken90]. IBDL [VK98]. IC [KCG98]. Iclient [Ber98]. Iclient/Iserver [Ber98]. Icon [Jef94]. ID [BG99, JKS^{+}93, Mud99, USE99l]. Idea [Ous96]. Ideas [Kol95c, Wal93b]. Identification [LUM91, LMU93, SL92].

Identity [Ell96]. IDL [MR95]. Idle [Arn95a, DMY99]. Idleness [GBS^{+}95]. IEC [Dun90c, Dun90b]. IEEE [Ano98i, Wal91b, IEE90, Sat95]. If [Arn91, dPG96, LO90, Lim99d]. Ifmail [Cro98]. Igor [Pie96]. II [Ber99b, Ber90, BGMR91, Chr93b, Gun98a, Kol94d, Pom96e, SPR96a, USE90i, USE91j, Wal92a]. III [Gun98b, USE92l, USE92m, USE93h, USE94m]. IISME [Sto96b]. IK [SSZ^{+}93]. IKE [Sim99]. IKE/ISAKMP [Sim99]. iKP [BGH^{+}95]. Illinois [USE99j]. Illuminated [She92a]. Illustrated [Mos96, NN96b, Pax94]. Illustrating [She92b]. Image [Bos96, CE99, Fly99f, Whixx]. Images [SF97a]. ImageTcl [Owe97]. Imaging [Laz98, PHS96a, PHS96b]. Imake [Dub94, Hae94a]. IMHO [Dam97a, Dam99, Hub98]. Immunology [Bur98]. Impact [Car93c, Laz97, Mue95a]. Implement [Ska94]. Implementation [ALLO92, Alf94, Ano99m, Ano99-31, AMP^{+}96, Bal92, BB^{+}91, Bar91, Bla90, BC94, BC91, BJM^{+}96, CFLM92, CFL94b, CD94a, CD92, CBE93, CGHK95, CDL99, CSYG94, CHS91b, CDM90, Var97, Cro98, Cuc91, DCM^{+}90, DGT96, DM92, FNP97, FLB98, GB92, Gar90a, GM93, GHM^{+}90, HS91a, HG96, Heu90, IB93, IM93, JHC94, Joh95a, JR99, KPI98, KN93a, KWB^{+}98, LKY93, LS97, LHF^{+}93, LC93, LP94, Mac91, MB91, MD96, Mar91b, Mar93, MSN93, Mit95, Mue93, MWO^{+}93, NN92a, NN96b, OB99b, Ols93, OOM^{+}93, PSWL95, PJS^{+}94, Pul99, Ros96c, SB94, ShCR93, STW93b, SLM^{+}90, SBMS93, SSB92, SeaBP99, Sos95, Spi97, SG97b, Sto92, SMPR98, TMC97, Tie90, Tie93a, USE94i, USE96f, USE99h, VcC97, WIK92, YL98, YLK93, Ano97a, Ano98l]. Implementations [CL94]. Implemented [HCK^{+}98]. Implementing [ALLO92, Alm99, BR94, CGLR92, CL98, CB90, CCW97, DMZ98, Gam93, HCC^{+}98, HHW92, Hol97b, KOO95, Lu97, MKC92, NSM99, RLS^{+}97, SS92, SWS92, TJH^{+}94, TIF^{+}97, VS90, WGC^{+}95, WSS92, YD94, ZWM^{+}99]. Implications [Bas98, MBMS95]. Implicitly [JD94]. Importance [Bre93, Mue96d, WR98]. Important [Sat93e]. Impressions [Jac93, Kre97]. Improve [CSM97, DWAP94, Den99b, FP93]. Improved [CBL90, OCH90]. Improvement [KP93a]. Improvements [BCF^{+}91, Kle90a, Kle90b, Pee96]. Improving [CL97, IC97, Jus94, PV95, SKP94]. In-Kernel [FP93, LHF^{+}93]. ina [Sma91]. Incident [Car99a, Ber99b]. Incidents [BKS90]. Incl [VC92]. Include [VC92]. Including [Bir95, CFSW99]. Incognito [HT96]. Incoming [MD92]. Incorporating [Eng97]. incr [Ber98, FEBM97, McL95, Ulf95, Ull97]. Increased [Bar97]. Increasing [Ber92, SW98, Sat90]. Incremental [FST91, OG95]. Independence [ESS96, Cow99]. Independent [Add96, Ant95, AHS95, ML90, NAU93, PT92, Rub95, Rub96b, SW96, SS95, Tho98b, USE93i, USE95i]. Index [Ano93a]. Indexing [DLO92, HS93b]. Indicator [JS93]. Indicators [BBP99]. Indirect [BB95a]. Individual [MR97b]. Industrial [Tho96]. Industry [LAG^{+}96, Sto96b]. inetd [RW97]. Inexpensive [Coo99]. Inference [Koe94]. Inferring [Cow99]. Informal [HS91c]. Informatics [Pie94b]. Information [AB96, Bar94a, Bar94b, Bia96b, DF95, EFJ96, Gol92, GC90, Hae95a, IL90, IL93, Isa95, Kol94g, Lea92a, MLL95, Mue95d, New95, Nys99, Opp96, She92b, SBV95, Ske92, Sum94, Var95b, VF95, WD95, YGM95, Kol94h]. Infrared [ASG^{+}93]. Infrastructure [EFJ96, Gru96, Hol97b, Isa95, Mee97, NWEF92, Per98a, Rou97b, TH98, HMT^{+}98]. Infusion [Ben94a]. ing [Dam97b].

Inheritance [Car90a, Car91a, CGMV91, Car93a, Car91b, FS96, Mar96c, Sak92a, Sak92b, Wal91a, Wal93a]. Initial [Hau94, ML91a]. Initializers [Rei92b]. Initiatives [Sto96b]. Injury [Oka98a, Oka98b]. INN [Fri97]. Innovations [Phe95]. Innovative [SPK^{+}96]. Inode [BL90b]. Inodes [GK97]. Input [Rao99c]. ins [Bel95]. Insecure [RG95, Wal95a]. Insecurity [Sal96i]. Inserted [MDK96]. Insertion [MDC95]. Inspection [Kur90]. Install [Pom99]. Installability [Sim94b]. Installation [Ano98q, Ano98-29, Ano99j, Ano99k, Ano99u, FL98, GES98, HH99, LL98, MWCR90, MWCR91, Mue96c, OS94, Oet98a, Sjö99, Soq97, Tom98, USE90f, USE91f, USE98e, USE99b, Wan91]. Installed [HW93]. Installing [CM94, RM94a]. Institute [Fin96]. Instruction [CL97]. Instructional [CPA93]. Instructions [Ros96a]. Instrumentation [RVL^{+}97, TM99]. Instrumenting [HS99c, LZ97]. Insulator [NFQ99]. Insuring [CBL90]. Integrated [EW90, GZ94, Lar91, NAU93, RAV95, RRB99, iW91]. Integrating [AHL^{+}91, Bea96, BM96, BCGS91, CD92, FCNL94, GM99, How90, How91, Kem93, LBSL97, MMR92, Ove91, Rub91]. Integration [BOM95, Cox97d, IH99, KSKD99, TSE90, WCN93]. Integrity [KS94a]. Intel [BMLH97, RVL^{+}97]. Intel-based [BMLH97]. Intelligence [Man98, Ken90]. Intelligent [Eng92, FC96, LS99, LWS^{+}94, WWDF98]. IntelliJuke [BKM97]. Intensional [EP92]. Intensive [FST99]. Intentionally [Cou97]. Inter [Maz96, NMMS97a, WCN93]. Inter-Domain [Maz96]. Inter-ORB [NMMS97a]. Inter-relational [WCN93]. Interaction [EFJ96, ESSB93, Lib90a, Mey91, Shu91a]. Interactions [MK98]. Interactive [Ell97, EW90, EWCS96, Lib91a, Lib92, Lib96a, MCU97]. Interactive-Group [MCU97]. Interactively [Sin97a]. Intercepting [HS99c]. Interception [HB99, NSM99, NMMS97b]. Interconnect [Hua98]. Interdependent [Win94a]. Interface [BN96, BHK^{+}90, Ben93, BCLC99, BJM^{+}96, Dra90, ES95, FHG93, FHL^{+}99, GS94b, Har95, HO94, Hou96, HLR97, KGMP^{+}96, KN93b, Lin90a, LP94, Lor95, Mar91b, Mar93, Mas91, McK95, MA90, MBH94, NN94b, NPS95a, NPS95b, Ped90, Pik91b, Pik94, Ros90, SMSM98, SKS90, SW90, Uhl95, VK98, Wel95, WBSL94, NMMS97a, Rao99e]. Interfaces [AHS95, ACO92, Hov91, KR99, Kot92, McC98g, Mue95c, SSL96, Ull97, WIK92, WN98]. Interfacing [KS94c]. Interim [Cae97]. Interlanguage [Ham96]. Intermediate [SCD90, Sha97]. Intermittent [MS94]. Internal [Cum99a, Mot95]. International [Ano98u, Ano98f, Ano98-44, Dun90a, Kar96, Kol96j, Mad92, Pie94b, San93, vW98, Pie94a]. Internationalization [Dun90c, Gsc95b, Gsc95a]. Internet [Ano98j, Ano98-33, Ano99g, Bar94d, Bar92a, Bar92b, Bar92c, Chr98b, Far94, Far95e, PB97, Roo97, Ste95, USE97h, USE99c, Uma96c, Ahl99, BM98b, Bar94b, Bar94c, Bar92d, BFM92, BP93, BTS^{+}99, Byr93, Car99a, CDN^{+}96, Che90, Che97, Dam97b, DLO92, ED92, Gon96, GB97, HS93a, HSOS99, VW99b, HK99, HS99a, Ins96d, JKS^{+}93, Kap92b, KLB98, Kle98a, Kle98b, Kre94, KF98, KRJM98, Lam97, Lar99, Lea96b, LS98b, LFB92, MC99, MJS97, Mar98, Mog99, MFM95, Mos94a, NB97, NMMS97a, Nkc99, NN96a, New95, PPC97, Qua98, RA94, Rod93, Ros96h, SSH93a, SBB96, Sal92g, SW92a, Sch96d, SEKN92, SA95b, SPR96a, Sim96b, Sug92, TAP90, WYOT93, Ylo96, YM99, kc99a, Hum98a, Laz97]. Internet-based [KLB98]. InterNetNews [Sal92g]. Internetworked [HS91c]. Internetworking [NN94c, IM93]. Interop [Cha92a]. Interoperability [CG92b]. Interoperable [Gre99, Sri93]. Interpersonal [JS93]. Interposition [Cur94]. Interpretation [Ede93]. Interpreted [Cou95, She95]. Interpreter [Gai94, Mat97]. Interpretive [FEBM97]. Interrogation [Sam95]. Interrupt [Col96, MR96, SCB93]. Interrupt-driven [MR96]. Interview [Dar97c, Hac98a, Kol95g, Kol96d, Kol96e, Kol96f, Kol96g, Kol96h, Kol97b, Kol98b, Kol98d, Kol98c, Kol98e, Mue94a]. Interviews [Kol99a, LC93]. Intra [DVPL92]. Intra-Domain [DVPL92]. Intranet [Ann98, CHHL99, Muf99, Roo99]. Intrinsics [HO94]. Introducing [MS92]. Introduction [Ful96, Hau92, Kre94, Mar96a, Mil97b, Rie97]. Intruder [Gir99b, Rod93]. Intruder-Administrator [Gir99b]. Intruders [Bru99a, Pax98, SSH93a]. Intrusion [Ano98g, Ano99x, Ano99-34, BZT99, Cas96b, Cas96a, DH99, EM99, For99b, GSS99a, KS94a, LS98a, LPS99, Mel99a, Roe99, RFR99, SU99, SST92, USE99l, WMB99]. Intrusion-Detection [For99b]. Intrusion-Tolerant [WMB99]. Intrusions [SBS99]. Intuitive [HF96]. Invasion [Sal96a, Chr98b]. invasive [Car91b]. InvenTcl [FEBM97]. Inventor [FEBM97]. Inversion [Mof92, Ols93]. Investigations [MDS99]. Invisible [Bru99a]. Invited [HMT^{+}98, Kre94, Pla97]. Invocation [VR99, WW97, KWB^{+}98, Rao99d]. IO [PDZ99]. IO-Lite [PDZ99]. IOI [Kol98k, Pie99a]. ION [Puc91b]. IP [BE98, BP93, Chr98c, Chr98e, Jen98, Laz99, ATS94, BB95a, Bel96b, BCC^{+}94, Cha92c, CGHK95, CF94, CHS91c, CJ95, CG92b, Hof99, Ins96b, Jac94, KP93a, Lim99a, Mos96, NN94c, NN96b, Pax94, Sca98, Swa92, TP91, Wil93, YHK96]. ipasswd [Hie92]. IPC [Bar91, Ber92, Dra90, KTN93, Koo93, OOM^{+}93, RN93]. IPI [NC92]. IPng [NN96a, Sim96b]. IPsec [Atk96, KPI98, Met98b]. IPv4 [FLB98]. IPv6 [Met98b, AMP^{+}96, Atk96, DMZ98, Fin97a, FLB98]. IPv6/IPv4 [FLB98]. IRC [Bru99b]. IRCAM [LPV^{+}91]. Iron [Ahr93]. Irrelevance [Ber92]. ISAKMP [Sim99]. ISC [LR98a]. ISDN [NN94a]. Iserver [Ber98]. ish [Sal97h]. Islands [USE92f]. Isn’t [Rus99a, Zwi92c]. ISO [Dun90c, Dun90b, Dun91b, Wal92b]. ISO/IEC [Dun90c, Dun90b]. Isolation [MD95]. ISP [Ila99]. ISP’s [CF98]. Issue [Ano99-37, Ano99-38, Ano99-39, Ano99-40, Ano99-41, Ano99-42]. Issues [BA92, CHS91b, DM92, FBAE^{+}92, Fin99b, Gai94, GF95, GGI^{+}91, GB97, HS91a, HR92, KP93b, NYKT94, Nes99, Ros96e, Sab90, SLM^{+}90, Shu91b, TJH^{+}94, TB94, Urn97, Vix95]. Italy [BFM92]. Itch [Dar97h]. IV [USE91e, USE93j, USE93m]. Iway [Tay96d]. Iway-80 [Tay96d]. IX [Far95c, USE95h, BDKM93].

Jack [Ber96]. Jacks [Bec99]. Jacl [LS97]. Jam [Tay96d, Sei94]. January [USE90h, USE91i, USE92j, USE93l, USE94l, USE95e, USE95l, USE96k, USE97e, USE98i]. Japanese [Hae95a]. Jargons [NAOP99]. Java [Hum98a, Lea97, O’N97a, O’N97b, Roo98a, BM98a, BCH^{+}98, BCE98, Bot98, IG99, BCC^{+}97, CJJ^{+}98, CMSvE99, DG96, DBK98, Far96f, Far96i, Far96g, Far96h, Far97i, Far97j, Far97k, Far98g, GMPS97, Har97b, HCC^{+}98, HHSW97, Kap97, KWB^{+}98, LS97, LH97, Lea96c, Lea96e, LZ97, LV99, LS98b, McC98b, McC98c, McC98d, McC98a, McC99d, McC99a, McC99b, McC99c, MS98, Mit96, Mit97, MK99, MMBC97, Nag96, NK98, OK99, Par99a, PW97, PTB^{+}97, Rao98d, Rao98a, Rao98b, Rao98c, Rao99a, Rao99b, Rao99c, Rao99e, Rao99d, Rie97, RWWB96a, RWWB96b, Sie98b, Sla97, TW96, VR99, WW97, WRW96a, WRW96b]. Java-Based [IG99]. Java-Object [BCC^{+}97]. JavaScript [AKM99, Fla97, O’N98]. JDK [For98]. Jersey [USE92f, RF95]. Jgraph [Pla93]. Jihad [Dam97a]. Jim [Kol96f]. JMAS [IG99]. Job [Dar93a, Dar93c, Fis96b, USE93e]. Jobs [MM92, RH93]. Jockey [MK92]. John [Chu99, Kol95g, Kol98e, Sal99m]. Johnny [WT99]. Johnson [Kol96h]. JOIE [CCK98]. Joining [ARH95]. Joint [Man98]. Jonah [ZWM^{+}99]. Jose [USE96i]. Journalism [Kol96b]. Journey [Kor97a]. JPEG [CE99, SW91b]. JTC1 [Dun90c, Dun90b]. JTC1/SC22/WG15 [Dun90c, Dun90b]. Jukebox [BKM97, GGWB99]. Jukebox-Based [BKM97]. July [USE95g, USE95j, USE96e, USE96i, USE97d, USE99b, USE99d, Kap92a, Wal91b]. Jump [Cla96]. JumpStart [Hei99]. June [USE90g, USE91g, USE92g, USE93g, USE94j, USE95f, USE95k, USE96g, USE97g, USE98d, USE98h, USE99m, Ano99-43, Eva94a]. Just [BTS^{+}99, Bow98, CL97, Dar96g, HSOS99, Kol94c, Pol96]. Just-in-Time [CL97]. Justice [Mil99j]. JVM [Rao98d].

Kawa [Bot98]. Keep [Car99a, Pom93d, SL93]. Keeping [Bra90, Bra91, Sim92b]. Keeps [Kol94c]. Kerberos [ARH95, BM91a, Bru98, DW98b, Dav95, DG95, DGT96, HRT96, IH99, Lun90, SA95a, Ste92, Ts’97]. Kernel [Alf94, BMHR91, Bon94, CLMS91, CWS90, CD94b, CRRS93, CGCS91, EKB^{+}92, FP93, FHL^{+}99, FK92, GD91, GLOR91, HO90b, HR91, HHLS90, LKY93, Lan92, LeF94, LHF^{+}93, LS92, MP92, MS93a, MR96, NL96, NKM96, OSS^{+}92, RFW93, Rey92, RN93, Rua90, Sch95b, SESS96, SM97, SKP94, TIF^{+}97, USE92i, USE93k, UKGS94, Wel91a, Wie92, YLK93, YR93, BM98b, CNTS93, JHC94]. Kernel-Based [Alf94]. Kernel-Kernel [CLMS91]. Kernel-Level [CGCS91]. Kernelized [DA92b]. Kernels [CPW93, Lan92, McR93, PVL91, SCB93, TM99, USE92i, vBC^{+}92]. Key [Bla94, CGHK95, DS92, Dav96, Gal96, HMT^{+}98, Jas96, LA98, MS95, Ros96a, Ros96f, Ros96g, Ros97, Ros98a, Ros98b, SSH96, Sto96a, TAP90]. Key-signing [Ros96a, Ros96f, Ros96g]. Keyed [LW99]. KeyKOS [BFH^{+}92, FL93]. Keynote [Bla93, Che98b, Leo91, Mor93, Rit90, Tho90b]. Keyspace [CD98]. KHEPERA [FNP97]. Killing [Joh93d]. Kind [Dij96a]. Kinzler [Sin98a]. Kit [GMPS97, LB94]. Knit [Hal98d]. Know [Hil97, Ins95a, Wal95c, Wil94, Zie93, Zwi91b]. Knowledge [BHJV92, CW96, Cla97, Hal99d]. Knowledge-Based [BHJV92, Cla97]. Knows [Poo94]. Kolstad [Kol99a, RK97]. Korn [Bor93, Mos94b]. KornShell [Kor96]. Krakatoa [PW97]. ksh [Kor94].

L. [Hac98a]. Lab [Kra98a, Rid94]. Labeled [NKM96]. Laboratory [Dav91, Smi91, SMD^{+}94, Lar91]. Labs [LRNL97, PPD^{+}95]. LACHESIS [SA95b]. LADDIS [KW93, WN92, WK93]. Laid [Joh93d]. Lake [USE95f]. Lame [Bee92]. Laminator [Zwi93c]. LAN [Han92, dC93]. Land [Ano97d]. Lanes [McR95]. Language [AG97, ABB^{+}97, AGG^{+}94, BL90a, BOM95, BA97, Bli90, BFRS99, BD92, Car95, Car92b, CLD^{+}97, Coo97a, Coo97b, Cou95, Cre97, DAM^{+}91a, Den92, Ell97, Fow94, Ful96, GH93, GL99, Hae94b, Hal94, Hop92, JD94, KH97a, Kap92a, KS97, Kor94, LB94, Lea96e, LM99b, MQ97, NZ99, Ous90b, Ous91b, Par99a, Pol94, Sal96b, SCD90, Sch96d, Sch96e, Spe91, SF97a, SG97b, SS95, TMC97, Tho90b, VZ91, Vis96, VK98, Wal93b, WAKS97, Win94b, iW91, Ker82].

Languages [Ano97i, AM98b, Bea96, BRH^{+}93, Bot98, FNP97, Fuc97, Gai94, Hae95b, Hud97, LR94, Lan90, MR97a, Mue96a, NN93a, NZ99, SB94, SG97a, USE94k, USE97b, USE99a, Ano98x, Ano98h, Ano99f, Van99]. LANs [Fis96c]. LaPadula [Bah90]. Large [Ano97g, Ano98q, Ano98-29, Ano99j, Ano99k, Ano99u, BA91, BFRS99, Cam93, CHK^{+}98, CCW97, Chr98b, CMS93, CFSW99, Cra90, Dar91, DH99, Fut95, GES98, GMS96, GB97, Hea97, How90, How91, Kol97i, Koo93, Man94, MO99, Mof92, MW98b, MS94, NWEF92, Pol97, RFD98, SV94, SA95a, SSB92, SW91a, Ske92, SH98b, TC92, USE90f, USE91f, USE97a, USE98e, USE99b, Wan91, Wil99, YBMM94, Sca97a]. Large-Scale [BFRS99, CMS93, GES98, GMS96, MW98b, Ske92]. Larger [HW93]. Last [Ano99-44, Hax96]. Latency [BDR97, BSSD96, Cha99, CRH99, EWCS96, GA94, KL97, LBBA99, LB97, MBH94, VF95, WVT94, YBA93]. Lava [LPJI98]. Law [Mue98, Ven96]. Laws [Sim92c]. Layer [Bal92, BB95b, IB93, KOO95, Noe98]. Layout [CL97, How95]. Lazy [CGS^{+}96, DB96b, ZIL96]. lbnamed [Sch95a]. LDAP [Lea98, Hed98, Mar97]. Lead [Cra96]. Lead-weighted [Cra96]. Leads [Ran99]. Leaks [HJ92]. Learn [Hae92]. Learned [ALLO92, CKM95, Gee94, Mac91]. Learning [GSS99a, Jac94, JSS90, LS99, Lib96a, Pot97, Bor93, Gil93, Kol94i, Mos94b]. Lease [CMB^{+}98b]. Leashes [Dar97a]. Lee [MW98a]. Legacy [ACFM99, MO99]. Legal [GADG96, Ros95b, Wri96]. Lehman [Rou97b]. Less [Fri97]. Lessons [ALLO92, CGLR92, Che98b, CKM95, Gee94, GLOR91, Mac91, Phe95, SPK^{+}96]. Letter [Ano97h, Ano98-35, Ano99-45, Bar96, Cal98, Car98a, Che98a, Cox98b, Dam98, Eva96a, Eva96b, Eva96c, Gar98, Hum96b, Hum98b, Hum98c, Joh92, Joh93c, Joh93a, Joh93b, Joh94d, Joh94b, Joh94c, Joh95b, Joh95c, Joh95d, Joh96b, MW98a, Mil97e, Mil98e, Mil98f, Per98b, Rei98a, Sal98a, Van98, Vix98, Zwi94b]. Letters [Ano96e, Ano97j, Ano98-42, Ano99-50, Ano99-46, Ano99-47, Ano99-48, Ano99-49, Aut94, Pee96]. Level [ASSB99, AISS97, And94, CDK^{+}99, CGCS91, Dea93, FK92, FHKR93, FSV94, Fre97, GH93, GSR93, Hae95b, ISS97, Kel99, Kor94, MPS98, MA90, MP99, Min93a, OW94, PS93, Ped90, Pol94, Raw93, RN93, SP91a, SU99, SKS90, Sze99, USE94k, MJ93, MH92, RMS^{+}99]. Leveraging [CK96a]. lex [Pax91]. LFS [RO90]. Liaison [KGL^{+}91]. Libckpt [PBKL95]. Libraries [BKC94, BTC97, Cog90, CHS91b, Den92, GS94b, GL99, HS91a, HS95b, Jon91, NV93, OLB93, Sab90, See90]. Library [All95, Arn95b, Ben94b, BN92, Cur94, DBGL^{+}93, Dea93, Det92, FN91, Kor96, Lam95, Lin92, Lu97, Mar96b, Mue93, MW96, ShCR93, SST90, SB97a, Sun97, Van91, Vo97]. libscheme [Ben94b]. LIBTP [SO92a]. License [AG99, Fin97c]. Life [Cam94, Gwy90, Kol96i, Pet93, Sim90, Sim91a, Ros95a]. Lifetime [Ano95b, KC96]. Lifetimes [OK99]. Lights [O’B99a]. Lightweight [CGS^{+}96, Gau91, Hai97, HHLS90, Lea98, MMS97, PHS98, Roe99, ShCR93, SM92, SG97a, SS92, HG98, Hed98]. Like [Tab94, MK91, HK95a]. Likeable [DW98a]. Limit [Cha94]. Limitations [BM91a, Cha99, ES97]. Linda [Mee97]. Line [CK96b, CFSW99, Ins96a, Sim96a, Woo94, Dou93, Shu91b, Uma98b]. Lingua [MR95]. Link [Mot91, SW98, Sim92b]. Linkage [Fen96]. Linked [HCL95]. Linking [GSB^{+}93, NH93]. Links [Det91, Fly99f]. lint [ML91a]. Linux [Ano99h, Ano99i, Den99c, Gra99a, Bla99, Den99a, GD96b, HH99, Mat99, RHG98, Tor96, Tro96, Yor99]. Lions [Chu99, Sal99m]. LISA [Ano98c, Ano98r, Ano98y, Ano99u, Ano99b, Ano99-35, Abb95, And93, AP99, Ano97i, Aut95b, Ben95, Cha98, Chr93b, Var96, DSS96, Dyk91, Far95c, Far96e, Kol94a, Old94a, Par99b, Red98, Num94a, Sch95e, SD92, USE91e, USE92e, USE93f, USE94h, USE95h, USE96j, USE97c, USE98g, USE99i]. LISA-NT [Ano99u, WCR99]. LISA/NT [Red98]. List [Eva94b, LT97, SK99, VWM98, WCO^{+}90]. Lists [CHK^{+}98, Cha92b, Kol97i, Rue96, Uma98a]. Lite [Wil97b, PDZ99, PM95a, RM94b]. Little [Har92, Lan90, Swa96, War99a]. Live [DeV93, DFKM97, Zwi93b]. Livelock [MR96]. Living [Pom94a, Uma95]. lmbench [MS96b]. Load [GG91, MGZ93, Sch95a]. Loadable [CJ98]. Loading [ASWW91, Gra98a, SW92b]. loads [BM98b]. Local [And92, LWS^{+}94, RAC97, Sie96, WB93, Won93]. Locality [CL97, GKAS99]. Locally [Mot91]. Locating [DT97]. Location [Ant95, Gol99, NAU93, PGH^{+}91, USE93i, USE95i, VDAA99]. Location-Independent [Ant95, NAU93, USE93i, USE95i]. Locator [AMP98]. Lock [BS92b, CHS91a, CPW93]. Locking [Bor98, Ska94, UKGS94]. Locks [BS97b]. Log [BHS95, CS92, CS94, KSS93, OS94, SBMS93, WAP99]. Log-Structured [BHS95, CS92, CS94, SBMS93, KSS93]. Logbook [Sim93b]. Loge [ES92]. Logging [CGS^{+}96, HS97, NSM99, SSB^{+}95, SO94, Tay96c, VGT95]. Logic [Bar98, CG99b, iW91, iW91]. Logical [HMF^{+}99]. Login [ARH95, Hau94, Hil97, Rat94, USE93n, Ylo96, Sal98h, Sal98i, Sal98j, Sal98k, Sal98l, Aut94, She93a, She93b, She93c, She94a, She94b]. Logs [Fin97c, GB98, RFR99, SK98]. Lonely [May95]. Long [BA91, DeV93, Kor97a, RH93, USE92e]. Long-Term [BA91]. Longer [FG90]. Longest [PP99]. Look [Har99a, Sch97a, Win94a, Hec94].

Look-Ahead [Win94a]. Lookaside [BKW94]. Looks [Joh97b]. Lookup [Duc94]. Lookups [LeF99]. Loops [SBRH96]. Loosely [PS93, RS91]. Loosely-Coupled [PS93]. Los-Alamos [CJC^{+}92]. Losing [HN97]. Lottery [WW94]. Louisiana [USE95e, USE99h]. Lousy [Mue96c]. Love [Kol95h]. Low [Bla90, BSSD96, Fre97, Gil90, Kap94, Sel94, YBA93]. Low-Latency [YBA93]. Low-Level [Fre97]. Low-Overhead [Bla90]. lpd [Sel95a, Fle96]. lpr [Sel95a]. lpr/lpd [Sel95a]. LPRng [PM95b]. LRP [DB96b]. Lucent [LRNL97]. LUDE [DBGL^{+}93]. Luke [Too98]. Lured [Che92].

MA [Ano99r, USE94e, Har99a]. Mac [IJ97]. Mach [USE90e, USE91h, USE93h, Bab90, Bacxx, Bar91, BDMVL93, BB93, Bla90, Bla91, BCC^{+}91, BL90c, BCCR91, CNTS93, CB90, CMS90, CN90, Dra91, Dra92, Duc91, ES90, Fin90, FL94, FGB91, GBB93, GD96b, Gol90a, GD91, Hov91, JCS^{+}91, KTN93, Koo93, Kup93, LBLM90, LHF^{+}93, MA90, MGZ93, MZDG93, Min95, Mit91, Mit93, NYM91, NYM92, NKT93, OOM^{+}93, Pat93, PAO93, RMGB91, ST93, Seb91, SP91a, SJ95, TN96, TNR90, TN91, Vau90, Wie92, ZTM96]. Mach-Based [Min95, LBLM90]. Mach-US [SJ95]. Mach/4.3BSD [BL90c]. Machine [And94, CGLR92, Pom96a, RF95, Rot99a, Tho98b, d’A92a, LV99]. Machine-Code [RF95]. Machine-Independent [Tho98b]. Machines [Bal92, FHL^{+}96, Oka92, SL94, SK98]. MacOS [BG93]. Macro [DHR97, HO90a]. Macro-Preprocessing [DHR97]. Made [Kol94h, Opp96]. MAEstro [DG91]. Magic [Rus99b]. Magnetic [Gut96]. Magnon [Tap99]. Mail [Bis92c, Bor92, Cha92b, CBB97b, Cuc91, Gru96, Har90, Har91, Hol97a, Pom96a, SV94, Wel95, Bar94c, Pom94a, SBRH96]. Mailer [Pom96a]. Mailers [Bor92].

Mailing [CHK^{+}98, Cha92b, Kol97i, LT97, Uma98a, VWM98]. Mailman [VWM98]. Mainframe [HJK^{+}92, VB92]. Maintained [Mot91]. Maintaining [Dav94a, Sat93b]. Maintenance [Cou96a, CKDS90, Cur91, Fly97, Sel91, Sma90, Sma91]. Major [Ros96b, Spe91]. MajorCool [Hou96]. Majordomo [Cha92b, Hou96, SK99]. Make [HSOS99, Sat94, Sie96, Sei94, Sin98b, Sin99a]. Making [BDD^{+}93, Cam94, Cla96, Hid94, Ism98, Kam98a, Maf93, MRK97, MP96, RM97, Uma95]. Malloc [Kam98b]. Man [Ada94, Chr90a, Chr91, KS94c, Shi91]. Manage [BTS^{+}99, Cha92b, Cra94b, HN97, SP99a, Tro96]. Managed [BJM^{+}96]. Management [AM98a, AW99b, Ano96a, AR91, AF94, AG99, BCK94a, BCK94b, BDM99, BHK^{+}90, BPT98, Bla93, Bla94, BM91b, BS93b, CGHK95, Cho98, Coo95, CG99b, DK95, Fin97b, Fle92a, FPSW91, FY98, GHN90, GHN91, GZ94, GS99b, Ham97a, HG96, Hea97, Hic99, Hog95, IFT^{+}91, Joh97b, JV95b, JV95a, JP99, KM92, KSS93, Lar94, LCH92, LO90, LKHA94, Lie92, MKPF99, Maz96, MAH98, MM92, PGH^{+}91, Pla99, Pom97a, PS92, PM98, RN93, Rot99a, Sch95d, SP91a, Shi97, SL95, Sie97b, SKS90, SCB93, VCV92, WH92, WW94, War99b, WWDF98, ZTM96, dSdCF^{+}98, SK99, Bir90]. Manager [BS92b, CCK^{+}91, GD91, IK95, Ken96, LaS90, Rat94, RO90, dSGM92, dSG93, MW98b, VWM98, dL95]. Managing [All99, And91, BTS^{+}99, BS98, Bur99a, Bur99b, Bur99c, Cha92a, Chr98e, Chr98f, CM94, CW92b, Dar98d, DFG^{+}96, Eva94b, Har99b, KYBR99, Lea99, Lib97a, Mel96b, Men93, Mil93, PBEDT97, RBB^{+}91, RH98a, Ros96h, RM94b, Rud97, Rue96, SV94, Sha92b, Sim92a, Sri93, Sub91, SH98b, Uma98a, Wil99, HSOS99]. Manipulation [Puc91a, Rub91]. MANs [Han92]. Manual [Bra90, Bra91, CD99, Zwi93b]. Many [Har92, LO90, Pom96a, Poo94]. Map [Sha92a]. Mapped [KSU92]. Mapping [RLBC94, SMD^{+}94]. Maps [Ada93]. March [USE91j, USE92l, USE99k, Kol97a]. Marches [Kol94b, Kol97j]. Marching [Kol97e]. Marcus [RK97]. Mark [Ano99-36, Kol97b]. Market [EFJ96, MW96, MW98b]. Market-Based [MW96]. Marketeers [Joh93a]. Marketplace [Mar91a, Wal93b]. Markets [Lip95]. MarketSpace [EFJ96]. Marks [KCDZ94]. Maryland [Ano97m]. MasPar [HHW92]. Mass [BLO^{+}91, GBK^{+}95, USE98b]. Massachusetts [USE93i, USE94j, USE97d, USE99k]. Massive [BA92]. Massively [MK95, Nes99, ZRB^{+}93]. Master [Tay95b, Tay96a, Tay97b, Tay97a, Tay97c, Tay97e, Tay97d]. Mastering [Hum97a, Uma97a]. Matching [WM92]. Math [SST90, Sto96b]. Matic [How98a]. MATLAB [MP99, JM94]. Maturity [Kub93]. Maven [Sal93g]. Max [Gun97]. Maxim [Eva95a]. Maximizing [GKAS99]. May [App96, USE92h, USE99f, USE99j]. Maybe [Pol96]. MBONE [Mot95]. MC [KBM^{+}96]. MCE [RAV95]. MD [USE92m]. Mean [App96, Fin95b]. Means [AMP98, NC94]. Measure [Fin90, MS96a]. Measured [DMF97, Wel91b, Wel91c]. Measurement [EM99, KP93a, Kub92, Sav99, Swa97a, Swa97b, WRJ92, kc99a]. Measurements [HO90b, Nkc99]. Measures [Car90b]. Measuring [BD97, BTS^{+}99, CPW93, ES97]. Mechanism [AOW96b, AOW96a, Bis91, ESS96, EL94, Fut95, Gau92, RM94a, RM94b, SLS92, Sum94, TA92, Var95a, YCO94, HG98]. Mechanisms [BMHR91, CHS91a, CBE93, Gil90, GM99]. Med [LFB92]. Medals [Pie95]. Media [Alt93, AOG92, AHS95, BLO^{+}91, Dav91, DBK98, FSM^{+}91, HK95b, NN92c, TN96, Kan91]. Media-Dependent [AHS95]. Media-Independent [AHS95]. Mediated [AD93, GD96a]. MediaView [Phi91]. Medical [PHS96a, PHS96b]. MedJava [JWS98]. MEDLINE [LFB92]. Meet [Ano99-51, FHL^{+}96]. Meeting [Ano98-38, Ano99-52, Dun90c, Dun90b, Fis97, Gra91, LO95, Par93b, Par93c, Wal91b, You93a, You93b, You93c, You93d, You94a, You94b, You94c, You94d, You95b, You95c, You96a, You96b, You97c, You97d, You98b, You99b, You99c, You99d]. Meetings [Ano95d, You98c]. Meets [Fin99a, Hop92, Lam97, Sin97b, Swa96]. Mega [Jae95, Lam95, McL95, Uhl96, Ulf95]. Mega-Widget [Ulf95, Uhl96]. Mega-Widgets [McL95, Jae95]. Megabyte [OW94]. Member [Ano98-36, Ber99c, Kre97, Mil98d, You95a, You96c, You97e]. Members [Mey91, Sto94d, Sto94e, Sto94f, Sto95c, Sto95d, Sto95e, Sto95a, Sto96f, Sto96g, Sto96d, Sto97a, Sto97b, Sto97c, Sto97d, Sto97e, Sto97f, Sto98b, Sto98c, Sto98d, Sto98e]. Memorex [Kol97d]. Memoriam [Ano99-36, Sal99m]. Memories [GM93, Tay96e]. Memory [AAL92, AF94, Bal92, BB93, BCF^{+}91, BS93a, Bon94, Boy91, BCCR91, CLMS91, CD92, Che94a, CK96a, CRRS93, CG90, CG91, CG92a, CGS^{+}96, DWAP94, Dou93, Fle93, Fra92, GKAS99, GD91, GL93, Gut96, HJ92, IC92, ISS97, KNM95, KCDZ94, KS99, KFJ94, LCC94, Loe96, MD96, McC99d, MS93a, MKB90, Min93b, MZA96, OW94, SB98a, SL94, SH98a, SP91a, SS97, SUK92, ZTM96, ZSB94, ZIL96]. Memory-Based [CK96a]. Menace [LKCT96]. Mergers [Lea99, Sch93c]. Merging [HAB^{+}98]. Mess [Sho97]. Message [AD93, Dan94, IC92, IMC93, KFJ94]. Message-Driven [KFJ94]. Message-Oriented [Dan94]. Messages [AOW96a, Cam93, Koo93, SKAM99, YBMM94, AOW96b]. Messaging [BSSD96, CK96a]. Meta [iIYL95]. Meta-Objects [iIYL95]. Metadata [GP94, VGT95]. Metalanguage [DMR92]. Metaphor [Ben93]. Meteorology [CS97]. Metering [SK96]. Mether [Min93b]. Mether-NFS [Min93b]. Method [BN96, BMLH97, KWB^{+}98, Mar99, Rao99d, VR99, WW97]. Methodology [RRB99]. Methods [CL98, Cra94a, GES98, LPS99, RA94, Sat93b]. Metis [ZPMD97]. Metric [CE99]. Metrics [DFKM97, Hog95, Kub92, LA98]. Mexico [USE93h, USE94k]. MH [GHN90, GHN91]. MHEG [Mar91a]. mhz [SM98b]. MI [USE92h]. Michigan [USE95i]. Micro [BMHR91, FK92, GLOR91, HO90a, Lan92, OSS^{+}92, USE92i, vBC^{+}92, CNTS93, SM98b]. micro-benchmark [SM98b]. Micro-Kernel [BMHR91, FK92, GLOR91, CNTS93]. Micro-Kernels [Lan92, USE92i, vBC^{+}92]. Microcomputer [BCGS91]. Microcomputer-Oriented [BCGS91]. Microeconomics [DF95]. Microkernel [BCF^{+}93, Ber92, FHG93, FL93, GMR93, HK93, MGZ93, MZDG93, MS93b, MWO^{+}93, WIK92, ZTM96]. Microkernel-Based [Ber92]. Microkernels [FHL^{+}96, Pet93, USE93k]. Micropayments [PHS98, Tan95, TL96, JY96]. Microprocessor [HHLS90]. Microsecond [MBH94]. Microsoft [Cum99a, Cum99b, Ham97a, SCW^{+}99, Sar99, Ts’97, VDA^{+}98]. Microsystems [MMR92]. Middle [Dar95e, Kol93a]. Middleware [BCC^{+}97, Dan94, GB97, MW98b]. MIDI [Sch92c]. Migrating [FL94, PG92, TLK99]. Migration [ASSB99, CCK^{+}95, Cum99b, Fre91, ISS97, LS92, MK93, MZDG93]. Migratory [Coo97b]. MIKE [CNTS93, Kol96g]. Millenium [Gas98]. Millennium [BCLC99]. Millicent [Man95]. Millions [SP99a]. Millipage [IS99]. Millipede [ISS97]. MIMD [GV91, Kot94].

MIME [GF95]. Mind [Ben94a, HN97]. Mini [Gra99a]. Mini-Review [Gra99a]. Minimal [GS96a]. Minimalist [Pik91b]. Minimizing [OW94]. Mining [LS98a, PP99]. Mintabs [HHW92]. Mirror [All99, Pul98]. Misbehaved [SESS96]. Mischievous [Arn90]. Misoneism [Arn90]. Miss [Ano96a]. Mission [Men93, Rou97a]. Mistake [Hae94a]. Misuse [GS99a]. Mitigate [Arn90]. MITRE [Gol90b]. MiX [KSAL96]. Mixing [MMBC97]. MJDLM [SK99]. mkpkg [Sta98]. Mkserv [RP92]. ML [Koe94]. MMU [DMY99]. MOA [MLC98]. Moat [DBD^{+}99]. MobiCom [Ano98u]. Mobile [ASG^{+}93, Ano98u, Ano98i, Ano99r, Ant95, BDD^{+}93, BP93, BCC^{+}94, IG99, BCC^{+}97, Den95, DCK^{+}94, DKB95a, DKB95b, Gra96, IGP95, IM93, KR99, KMS^{+}93, KPR94, MC99, MLC98, NPS95a, NPS95b, RAV95, RASS97, RASA98, SKM^{+}93, Sat95, STW93a, TC95, USE93i, USE95i, WYOT93, WB93, WS93, ZD93a, ZS99, RMS^{+}99]. Mobile-Agent [Gra96]. Mobility [BZCS96, TIF^{+}97]. Mode [Alm99, Hun97, NN94a]. Model [ALLO92, AKM99, Bah90, CD94b, CG90, CRH99, DV90, FG91, FL94, For98, HTWW96, LDOW98, Lim99b, MMR92, Neu92b, RHKP95, Sch93b, SA98, Wie92, WRW96a, WRW96b, dSdCF^{+}98, Kub93]. Modeling [AG97, Ber99b, DG96, Ell97, LaP94, Mil96b, SS99, Sim96d]. Models [DF95, FHL^{+}99, HL93, MZA96, Par92, Zwi94e]. Modern [Sch95b, Sch96b]. Modes [TB99]. Modified [Min93b]. Modula [Vix92]. Modula-3 [Vix92]. Modular [CD94a, CGHK95, NKM93, Rhe98, SO92a, WH97, YTJ91]. Modularity [BTC97, WIK92]. Module [Mee98]. Modules [Fur91, FO96, IH98, Koh99]. Monadic [PH99, WH97]. Monads [SeaBP99]. Money [Ano98-41, CST95, dPG96, MR98, Sim96a].

Monitor [Cra94a, Kin91, Wal92b]. Monitoring [AP97, Arn95a, BHJV92, BTS^{+}99, Bla92, Cox98d, Ede93, Fin94b, Fin97c, GB98, HA92, HA93, HM92, HMS94, HS97, Kis90, Kis91, Kra98a, Pre99, RLS^{+}97, SB93, SBS99, SW91a, SC99, USE99l, Ven92, Ano98g, Ano99x, Ano99-34]. Monolithic [JHC94]. Monotonic [HG93]. Monsters [SF97a]. Montage [Haa90, WN98]. Montana [Ham97b]. Monterey [USE91h, USE93f, USE94i, USE95h, USE95k, USE96e, USE97h, USE99m]. Month [Ano99-44]. Months [VB92]. Moose [Ros96c]. Morgan [GMR95]. MosquitoNet [BZCS96]. most [Mor92, Ous96]. Mostly [WB93]. motd [Kol97f, Kol97g, Kol98f, Kol98g, Kol98h, Kol98i, Kol98j, Kol99b, Kol99c, Kol99d, Kol99e, Kol99f, Kol99g]. Motel [ML93]. Motif [Dav94a]. Mount [OT90]. Mounter [ML93]. Mounts [PM95a]. Move [dPG98]. Movement [DA92b]. Moves [Sch93c]. MOVIE [Uhl91]. Movies [Cuc95, Cuc96, SP91b]. Moving [BVG^{+}97, BGT98, CMB^{+}98b, Cla96, CFSW99, GD91, Hun95, Mil99f, Sch96c, Uhl93]. Mow [SMH95]. MP [CHS91a, HHW92, Pea92, Tal95]. MP-1 [HHW92]. MP-capable [Tal95]. MPVM [CCK^{+}95]. MrMean [Rad97]. MRTG [Oet98b]. MS [Bru97, FM94, Pla97]. MS-DOS [FM94]. MS/Security [Bru97]. MS/Zero [Pla97]. MSN [Sar99]. MTEX [TLK99]. MTX [Uhl90]. Mucking [LeF94]. MUD [Pol96]. MUDs [Eva93b]. Multi [BGJY98, BSKH93, CJ98, GF95, Gar90a, GBK^{+}95, Jon91, Kan91, KBM^{+}96, MS92, MH92, NN92c, PKB^{+}91, Rei98b, RDM95, Sam95, STW93b, SV90a, Spe96, Spe97b, Swa92, Tom98, Van97, WN92, ZL96, Oet98b]. Multi-Homed [Swa92]. Multi-Host [Spe96, Spe97b, ZL96, SV90a]. Multi-level [MH92]. Multi-Media [NN92c, Kan91]. Multi-OS [Tom98]. Multi-Platform [Rei98b, RDM95, Sam95]. Multi-Policy [CJ98]. Multi-Processor [Gar90a]. Multi-purpose [BGJY98]. Multi-resident [GBK^{+}95]. Multi-structured [MS92]. Multi-thread [PKB^{+}91]. Multi-Threaded [Jon91, STW93b]. Multi-User [BSKH93, GF95]. Multi-Vendor [WN92]. Multi-Zone [Van97]. Multiapplication [Gir99a]. Multicast [CHW98, CK96b, GWP99, LC99, LOM94, MBM95, Mot95, vBC^{+}92]. Multicast-Based [Mot95, GWP99]. Multicomputer [Roy93, SWS92, WSS92, Pea93]. Multicomputers [ZRB^{+}93]. Multidatabase [BS93b, CBE93]. Multihost [Sch91]. Multihub [Cuc91]. MultiLane [McR95]. Multilevel [d’A92a, dC93]. Multilingual [Spe97c]. Multimedia [BK97, BH99, Bor92, Bvv91, DCWB92, DBK98, DG91, Ell97, FSM^{+}91, GGV96, HS94, HK95b, HSE96, HS91b, KGL^{+}91, LKY93, Lin94, LR98b, MS95, MLM94, NYM91, NYM92, Owe97, PCMI91, Phi91, Sal93g, SW95, YLK93]. Multimedia/Realtime [NYM91, NYM92]. MultiPersonality [Hop96]. Multiple [BGG^{+}98, CL91b, Car90a, Car91a, CGMV91, Car93a, CCGS92, FP90, GS98, GMR93, HCC^{+}98, KP93b, MBMS95, NB94, Oka92, RC97, Sah95, SMPR98, Wal91a, Wal93a, YBMM94]. Multiple-Backend [RC97]. Multiplexing [YHK96]. Multiprocessing [Ano92, EKB^{+}92, Laz92, Sch95b]. Multiprocessor [Bal92, BCVC91, CPW93, CLMS91, CS90, FK92, GKAS99, HO90b, IC92, Kot92, MS93a, PVL91, Puc91a, RS91, Rus91, SLM^{+}90, SUK92, USE91j, USE92l, USE93j, UKGS94, ZY98, ZSG92]. Multiprocessors [Bre93, CMS93, HJ97, Kot94, LS93, OW94, Vas91]. Multiprotocol [Bor98]. Multithreaded [Ama98, CB90, PG92, TCI98]. Multithreading [EKB^{+}92, PST^{+}92, Pea92, SPY^{+}93]. MultiView [IS99]. Murphy [Ven96]. Music [AB91, Lan90, Tho90b, LH90, Haw90b]. Musical [Lan91, LPV^{+}91]. Musings [Far95c, Far95a, Far95b, Far96e, Far96a, Far96b, Far96c, Far96d, Far97g, Far97b, Far97c, Far97d, Far97e, Far97f, Far98b, Far98c, Far98d, Far98e, Far98f, Far98a, Far99a, Far99b, Far99c, Far99d, Far99e, Far99f, Ins95b]. Must [Wal95a]. Mutual [Bak99, MJ92]. MVHS [Shi98]. MVM [GMR93]. My [Dew97, Ins96c, Lea99, Zwi94j, Pin92]. Myers [JS93]. Myers-Briggs [JS93]. Myrinet [ACG^{+}98]. Mythology [Leo91]. Myths [Par94].

Nachos [CPA93]. Name

[Bel95, Det91, Gil96, LSR95, NR94, RH98b, Sch95a, Sch96b, SO92b, Zwi92a]. Names [MAH98, VDAA99]. Namespaces [Lib97a, McL95]. Naming [CDM90, MC99, MS92, SV94, YCO94]. Nanokernel [BFH^{+}92]. Nashville [USE91g]. National [Gsc95a, Hei90, dSM96, Kap92b]. Native [Rao99e]. Navigation [Tay97a, YKB94]. NBC [AKS98, SSZ95]. Necessary [PCMI91]. NED [May92]. Need [Car90a, Pre97, hvESV^{+}97, Zwi93b]. Needle [All95]. Needles [Hop92]. Negotiation [BN96, MW96]. Neighborhood [SPK^{+}96]. Nemesis [Han99a]. NeoWebScript [Leh98]. Net [Bac95, CK96b, Hec96, Nas96, Ste95]. NETA [USE99g, WCR99]. NETA/LISA [WCR99]. NETA/LISA-NT/Security [WCR99]. NetBench [Swa97a, Swa97b]. NetBill [CTS95]. NetBIOS [Her97, Ran97]. NetBSD [HK95a, Tho98b, Tho98c]. NetCents [PHS98]. Netherlands [Pie95]. Netizens [Laz97]. NetKuang [ZL96]. NetMapper [Gol99]. Netnews [Bar94c, Uma96a]. NetPebbles [MPZD98]. NetReg [VW99a]. Netscape [Lea96a, Lev96]. NetScript [PM98]. NetWare [ARH95, MMP94]. Network [Abb94, AW99a, ASG^{+}93, All99, AW99b, AHB^{+}98, And91, And92, AK94, ACG^{+}98, Ano98e, Ano98g, Ano99x, Ano99-33, Ano99-34, AUH99, AR94, Bar91, Bel90, BEM91, Bla92, Box96, BCR90, Bur99a, Bur99b, Bur99c, Cha92c, Cha99, CH95, Cox98d, Dar91, DBCF95, DS92, DBD^{+}99, DB96b, DH99, Fia95, FLB98, Fle92a, Fle92b, For99b, FL98, GLP^{+}93, Gir99b, Gol99, GS99b, GS94c, HJ97, HM92, Har97b, HG96, HMS94, HO92, HK92, HR91, Hug96, HHSW97, IP94, IB93, IFT^{+}91, Jen98, Kam93, KGL^{+}91, KMS^{+}93, KOO95, Kis90, Kis91, KCCT94, Lar94, LeF92, Lea99, LCH92, LO90, LRNL97, LWS^{+}94, LFB92, LCL98, Mar99, May90, May92, MET98a, MK92, ML91b, MPZD98, MD95, NYKT94, NN94b, O’N97a, OOM^{+}93, Pax98, Pom94b, Pom96d, Pom96e, Pom97b, Pre99, PS92]. Network [Pul98, RS91, RG95, RASS97, RLS^{+}97, Roo98a, SMSM98, Sar99, Sav99, Sch95d, SHAS95, SL93, SSK97, She93d, Shi98, SW91a, SL95, Sie97b, dSGM92, dSG93, SW92b, SH98b, Sum94, SCD94, TSM95, TA92, USE99g, USE99l, Ven92, Wal95d, War99b, WS99, Wel91b, Wel91c, WUB^{+}90, WVT94, WBSL94, WCS93, Kap92b, Lal98, LBSL97, dL95]. Network-aware [RASS97]. Network-Transparent [GLP^{+}93]. Networked [CCM^{+}90, Eva93b, HJK^{+}92, Rot99a, VCV92]. Networking [Ano98u, Ano98-28, Ano99x, BP93, CWT^{+}97, Cas95, Dij96a, FB96, Lee94, MPS98, Mil99k, MZA96, Pac93, Pea99, SKP94, TB94, USE94f, USE94m, Laz99]. Networks [Bel90, BZT99, BL97, CCW97, cC99, Cho99, Chr98e, Chr98g, Cou97, CJ95, Das97, FHLW99, dPG96, GS99a, GSPT95, Gre99, KYBR99, KCVP95, Lim99a, MPS98, ME99, PW93, Roe99, SL96, Sel97d, Ten96, War94, dSdS98, BE98, NN95b]. Neural [GS99a, WUB^{+}90, WW91]. Neurophysiology [AB96]. Neutral [WN92]. Never [Nem99b, Pom94g, Ros95c, Sim93b, Zwi93c, Zwi94g, Zwi94f, Zwi94h]. Newport [USE92l]. News [Chr97a, CBB97a, Dar96h, Eva95c, Fri97, LOM94, Ros98a, Sae98, Sto95b, SC92, Sal97i]. Newspace [BLO^{+}91]. Newsreader [Her96]. newu [Sch91, SV90a]. Next [GMR95, KW93, NN96a, OSS^{+}92, Sim96b, WK93, She95, TMJY91]. NeXTWORLD [Dyk92]. NFR [LPS99]. NFS [And92, BS92b, Bla92, Duc94, EBG99, HLM94, Jus94, Kei90, KW93, Mac91, Mac94, Min93a, Min93b, Mog92, Mog94b, NC92, PJS^{+}94, Rob99, SB93, Sim92a, SW92b, VGT95, WN92, WK93, dSdCF^{+}98]. NIC [War99a]. Night [Kor97a]. Nightly [Oka92]. Nine [Lim99b, VB92]. Ninja [GGWB99]. Ninth [USE95h]. NIS [CAY99, Sim92a]. NIX [Sal99b]. NLnet [the98, HBvdRA98]. NLOG [Fin99a]. NLOG/NMAP [Fin99a]. nlp [Fle92b]. NM [USE98c]. NMAP [Fin99a]. No [CBL90, FG90, Har99a]. No-Fault [CBL90]. Node [FT99]. Nomads [BDD^{+}93]. Nominating [Nem99a, WKR98, You97b]. Nominations [Ano95c, Wil95c, Wil95b]. Non [Car91b, Car99b, GC96, JHC94, LW99, Win96, Yah94]. Non-blocking [GC96]. Non-Encrypting [LW99]. Non-invasive [Car91b]. Non-Monolithic [JHC94]. Non-Synchronized [Yah94]. Non-Technical [Win96]. Non-Windows [Car99b]. NonStop [Nor91]. NordU99 [Ano98s]. Nostalgia [Mel97b, Sha92a]. Notable [Bar94a]. Notational [CST95]. Note [AK96, Mil99g]. Notebook [Man97]. Notes [DSS96, Tay95a, Arn98]. Nothing [AM97, Cox97c, Phi95]. Notice [Ano98-38, Ano99-52]. Notification [HA93, HM92, Mee97]. Notifications [AMML96]. November [USE91h, USE93f, USE94i, USE96h, USE99i]. Nowhere [Dar95e]. NR [Her96]. NRL [Met98b]. NSFNET [Hei90]. Nsync [BK97]. NT [Ano97g, Ano97p, Ano98q, Ano98k, Ano98w, Ano98-29, Ano98z, Ano99j, Ano99k, Ano99u, Bru97, Bru98, Car99b, Cox97d, Han99b, Sca97a, Sie98a, lWS99, ASSB99, Ahl99, AHB^{+}98, ACFM99, Alm99, Ano97d, Arn98, AC97, BVG^{+}97, Bir93, BGT98, BCC^{+}99, BC97, BCLC99, CRVZ97, Car99b, CCW97, CDK^{+}99, CJ99, CH97, Cof97, CGLR97, CRH99, Cox97b, Cox98c, CD99, Cum99b, Cut92, DBC^{+}99, Das97, DMZ98, ES97, EG97, FL98, GES98, GSS99b, GL97, Gun97, Gun98a, Gun98b, HJ97, Hea97, Hil97, HAB^{+}98, HPS97, HCK^{+}98, IH98, ISS97, JR99, Kor97c, Kor97a, Kra99, Kra98a, LeB99, Lei98, LFR^{+}98, Lim99c, hLhCN98, LBSL97, Lyc96, ME99, MD98, NLZ99, O’B99a, PS96, Pla99, Rat98, Red98, RW97, Rus99a, SCW^{+}99, SD98a, Sat93d, SH98a, Sjö99, SSBK98, Str99, TLK99, TCI98]. NT [USE97a, USE97f, USE98e, USE98a, USE99b, USE99d, Wal95b, Wal97a, Wal98, WWDF98, ZY98, Zea97, Zwi99c]. NT-Based [ISS97, Kra99]. NT-SwiFT [HCK^{+}98]. NT/2000 [Ano99k]. NT/95/98 [WWDF98]. NT/Security [WCR99]. NTFS [Fre97, vL98]. NTnix [Jon98]. NT^{T M} [RvIG98, Wal97b]. Nucleus [HK93]. NUMA [Bas98, Bre93, LS93, UKGS94]. NUMACROS [LS93]. Numbers [Gut98b, Kol95h]. Numerical [Spe93a]. nuthin [MH92]. Nutshell [Som97, Sie98a].

O [How98a, ML93, AT92, ACG^{+}98, BS96, BC91, CL91b, CL91a, CG99a, FP93, FGB91, KV91, Kot94, KSU92, McC98a, MDK96, PDZ99, RvIG98, Rus99a]. O-R [AT92]. O.S. [Hov91]. O/S [Rus99a]. Oakland [USE94f, USE96h]. OAT [SKCB97]. OATH [Ken91]. Object [Ano98n, Ano98-27, Ano99s, Ano99t, Ano99-29, AOW96b, AOW96a, Arn90, BCK94a, BB^{+}91, BCF^{+}93, BCK94b, BMK96, Bir95, Bog95, Bon94, BCC^{+}91, BCC^{+}97, CLS92, CI93, CDN^{+}96, CFH^{+}93, Cop94, DDD92, DLVW98, DeW91, DV90, FHG93, FL93, FK98, FK99, GF95, GS94b, Gor90, GS90, GKS90, Gra92b, Hae94b, HO90a, Hal98c, Ham96, Isl95, JV95b, JV95a, Ken91, KN93a, KP97, KMR91, KAD96a, KAD96b, Lea95a, MSS99, MSL94, MPS98, Maf93, Maf96, MO99, ML90, Min95, MCU97, Mog90, MDC95, MK99, MR95, Mue95b, MWO^{+}93, NUO^{+}97, NL98, NZ99, OK99, PT92, Phi91, PHS96a, PS97, PHS96b, RHKP95, RA98, RM97, Rus91, SKCB97, SL94, Sch94b, SHAS95, She95, Shi97, SMDP97, STM^{+}99, SJ95, Tho96, TB99, USE95k, USE96g, USE97g, USE98c, USE99f]. Object [Ulf95, Wei95a, WL95, WRW96a, WRW96b, WVS^{+}99, iW91, Wyb90, dSdCF^{+}98, CNTS93, Ros95d]. Object-Based [Bir95]. Object-Caching [Bon94]. Object-Groups [SMDP97]. Object-Oriented [Ano98n, Ano98-27, Ano99s, Ano99t, Ano99-29, AOW96b, AOW96a, BB^{+}91, BCF^{+}93, BMK96, Bog95, BCC^{+}91, CLS92, CI93, Cop94, DDD92, DLVW98, DeW91, FHG93, FL93, Gor90, GKS90, Gra92b, HO90a, Isl95, JV95b, JV95a, KP97, KMR91, MSS99, Maf93, MO99, MWO^{+}93, NUO^{+}97, NZ99, Phi91, PHS96a, PS97, PHS96b, RA98, Rus91, Sch94b, SHAS95, Shi97, TB99, USE95k, USE96g, USE97g, USE98c, USE99f, Ulf95, WL95, iW91, Wyb90, dSdCF^{+}98, Hae94b, Ken91, KN93a, CNTS93]. Object-Replication [MCU97]. Objecting [Joh94a]. Objective [Bog95]. Objective-Tcl [Bog95]. Objects [AMML96, BCH^{+}98, Ber98, BA97, CGLR92, CW96, CJJ^{+}98, DAM^{+}91b, FHLW99, GS94b, GL93, HK93, How98b, iIYL95, Joh94a, KS94c, Lam91b, LC99, McL95, MCLY94, MLC98, MKC92, NMMS97b, RM97, See90, WWW95]. Observations [Spa92]. Observing [Van97]. Obtuse [Coo97b]. OC3MON [ACTW96]. October [Dar93a, IEE90, USE90e, USE90f, USE91e, USE91f, USE92e, USE93e, USE93m, USE94k, USE96f, USE96j, USE97b, USE97c, USE99a, USE99c]. O’Dell [Kol96g]. Off [Dar99b, Ros96d]. Office [AFJ^{+}99, Pol96]. Offline [HSK99]. Oh [Pin92]. Ohio [USE93g, Rom90, Rom91a, RFR99]. Ohio-State [Rom91a]. Oil [Cur99]. Old [Ada93, Ada94, Kol97c, Sha92b]. OLE [Box96]. OLTP [YR93]. Olympiad [Pie94b, Kol93c, Kol96j, Kol99h, Pie94a, Pie99b]. Olympics [Pie95]. OMNICONF [Hid94]. On-Line [CFSW99, Woo94, Dou93, Shu91b]. On-the-fly [Lew96]. ONC [ESS96, Mat97]. One [Arn91, Cha98, CF98, Hal98d, Her97, HHLS90, Hub98, Oka92, Pom94b, Pom96a, Rub95, Rub96b, Sat93b, Mel96a, Mor92, MAM95]. One-Time [Rub95, Rub96b, MAM95]. Ones [Sha92b]. Ongoing [SSH93a]. Online [Gra98c, PK96, Sal96a, Tay97c, She92b]. Only [cC99, Lib97b, ZD93a, ATS94]. Ontario [USE94g, USE95j]. onto [ACFM99, Uhl93]. OO [Sch96g]. OODCE [Dil95]. Open [Bra92a, Cas95, CCK^{+}91, CBB97b, Col92b, DBC^{+}99, DCWB92, DZL97, EFJ96, Fer95, GSPT95, Hae91a, Har99b, Kap92b, Kev93, Kol98a, Lyc96, Rot95, Sto99b, Tho90a, Wal95b, dSM96, FEBM97, Wal95c]. Open-Ended [Fer95]. OpenBSD [KPI98, dR98, dRHG^{+}99]. OpenDist [OG95]. OpenDoc [Ing96]. Opened [Kol95e]. OpenGL [Esp97]. Opening [CdR99, KG93, Pet96, RC93, Sat93c, Sch96h]. OPENNT [Wal98, RW97]. OPENNT^{T M} [Wal97b]. OpenSim [MM92]. Operates [Min93a]. Operating [AISS97, Ano97a, Ano98l, Ano99m, Ano99-31, BDKM93, BBS95, Ber92, BGÖS98, BBG^{+}99, BCR91a, BCR91b, CPW93, CD94b, Cho99, Chr90b, CPA93, CCGS92, Var97, DCM^{+}90, DAM^{+}91a, DVdB^{+}94, DV90, Fin90, GKAS99, GS98, GGI^{+}91, GPRA98, Gil93, GGV96, GC96, Han99a, iIYL95, JLI98, Kap92a, KCDZ94, KN93a, KC96, LB96, Lam97, Lea94, LR98b, MMP94, Mit95, MP96, MLM94, NYM91, NKM93, Nor91, Ous90a, Ous91a, PVL91, PBR91, Qui94, RBJ91, Rei92a, Ros92, Roz92, Rud97, RMC90, Rus91, SB98a, SLS92, SHK92, SLM^{+}90, SM98a, SCB93, SUK92, Tal95, TM99, Tho98c, USE94i, USE96f, USE99h, UKGS94, Vas91, Vau90, Web93, Won90, d’A92a]. Operation [Duc94, Duc97, HH93, HH95a, HH95b, MBM95, O’B99a, SKM^{+}93]. Operations [DBGM^{+}98, Hea97, Min95]. Operator [NC94, Shu91a]. OPIE [MAM95]. Opinion [Bac95, Eva95d, IS93, Sat93d, Sec95]. Opinions [GADG96]. Opportunistic [BRW99, OS94]. Optical [TSE90]. OptiMach [Bacxx]. Optimal [CS92, CS94, OW94, Spi99, Swa92]. Optimisation [iIYL95]. Optimistic [Bacxx, BDR97, Duc94, Rei92a]. Optimization [ISS97, Koo93, PBF^{+}96, POS^{+}99, RVL^{+}97, SKP94]. Optimized [CK96a, Lu97, WWDF98]. Optimizer [CGLR97, GL97]. Optimizing [DMY99, ESSB93, GL99, HCL95, Mue94b, RC97]. Option [OT90]. Optional [HO94]. Options [Ros95c]. Oranges [Dij97]. ORB [NMMS97a, NLG^{+}99]. OrbixWeb [NSM99]. ORBs [POS^{+}99]. Orca [KMBT92]. Orchestra [Haw90a]. Orchestration [WW91]. Order [CG92a, Cou97, Lae95, Lau95, Puc99, HR93]. Order-Dependent [CG92a]. Ordinary [KP97]. Oregon [USE90i, USE92o, USE97g]. O’Reilly [Far95e]. Organisational [Dun90d]. Organization [PW93, SV94, WVS^{+}99]. Organization-Based [WVS^{+}99].

Organizational [MWCR90, MWCR91]. Organizing [CHHL99, ES92, Mah91, SP96, Wil99]. ORGCON [ALLO92]. Oriented [Ano98n, Ano98-27, Ano99s, Ano99t, Ano99-29, AOW96b, AOW96a, BCK94a, BB^{+}91, BCF^{+}93, BCK94b, BMK96, Bir95, Bog95, BCC^{+}91, BCGS91, CLS92, CI93, Cat90, Cop94, DDD92, Dan94, DLVW98, DeW91, FHG93, FL93, Gor90, GKS90, Gra92b, HO90a, Isl95, JV95b, JV95a, KMS^{+}93, KP97, KMR91, MSS99, Maf93, MO99, May90, ML90, Mog90, Mue95b, MWO^{+}93, NUO^{+}97, NZ99, PT92, Phi91, PHS96a, PS97, PHS96b, RA98, Roh96, Roo92, Rus91, Sch94b, SHAS95, She95, Shi97, SW90, SUZ98, TB99, USE95k, USE96g, USE97g, USE98c, USE99f, Ulf95, WL95, iW91, Wyb90, dSdCF^{+}98, CNTS93, Hae94b, Ken91, KN93a]. Origin [RH98b]. Orleans [USE95e, USE95l, USE98d, USE98h, USE99h]. Orthogonal [Ros92]. Orthogonally [DVdB^{+}94]. OS/2 [PAO93]. OSDI [Ano98t, Ano99m, Ano99-31, Var97, USE94i, USE96f, USE99h]. OSes [Den99a]. OSF [ASWW91, BCF^{+}91, BM91b, Dil95, GS94a, LPLF91, Mit91, ZRB^{+}93]. OSF/1 [ASWW91, BCF^{+}91, BM91b, LPLF91, Mit91, ZRB^{+}93]. OSI [CG92b, Dun91c, Maz96]. Other [DBGM^{+}98, DMY99, HSOS99, Lan92, SPY^{+}93, Sch96e, TW93, USE92i, USE93k, DFKM97]. OTSO [KMR91]. Ought [Zwi91b]. Our [Ano99-54, Mil99d, You98d, You99e, dLRT93, Wal95a]. Ousterhout [Kol95g]. Out-Of-Core [MDK96]. Output [Rao99c]. Outreach [Shi98]. Outstanding [Ano99a]. Overcoming [ACO92, KH97b]. Overflow [CPM^{+}98]. Overhauling [Coo92]. Overhead [Bla90, Sim96a]. Overlaying [KK92b]. Overview [Chr93a, Cox97a, DLM90, Dos91, GMPS97, Hil92, JCS^{+}91, KLA^{+}90, MMP94, NWRP^{+}91, Red98, Rey92, Roz92, Sca97a, Sch99, SSZ^{+}93, Sto98e, dRHG^{+}99]. Own [Ins96c, Tay97b].

PA [WIK92]. PA-Risc [WIK92]. Pace [Car99a]. Package [DL93, GHN90, MAH98, Qui94, SSH93a, SKCB97, SY91]. packaging [Sta98]. Packet [BGP^{+}94, Cha92c, CB96, CHS91c, LO90, MJ93, Mol95, Skl91, WYOT93, WGC^{+}95, YBMM94, AUH99]. Packet-filtering [CB96]. Packets [Cam93, LO90, MD92, Swa92]. Pad [BH95]. Page [Dra91, Ede93, GS98, IS99, MA90, RLBC94, SC93]. Page-Based [IS99]. Pageable [MKB90]. Pager [MD96, MA90, Sub91]. Pagers [Boy91]. Pages [Fin96, Lea95b, Sin98b, Sub91, Tay95b, Tay96c, SP99a, BD93, Sin99a]. Pagesize [SMPR98]. Paging [Han99a, KN93b]. PAI [KGMP^{+}96]. Painkiller [Pom94g]. Panda [BRH^{+}93]. Panel [AGF90, ADKN90, BCR90, Cla97, CKM95, Fin99b, Lea94, Mem93, Par92, PBEDT97, Sca98]. Paper [Har96a, She92b]. Paradigm [CCGS92, Mil99c]. Parallel [AOW96b, BRH^{+}93, BL97, Bre93, IG99, CGLR92, Das97, FBAE^{+}92, Gar90a, HO90a, IMC93, JD94, KMBT92, KTP^{+}96, LS93, LIN^{+}93, MSS99, Mas91, MK95, SL96, Sch96g, Sco93, dSGM92, SLC94, TJH^{+}94, WRJ92, ZRB^{+}93, SD98a]. Parallelism [FLA94, Sun97]. Parallelization [BL90c, CBB^{+}91]. Parallelized [LBLM90, NYKT94]. Parallelizing [Ber90, BM91b]. Paralysis [Hal99c]. Parameterization [Bli90, TC92]. Parser [GKS90]. Parsing [Car94, Par99a, SP91b]. Part [Bur99a, Bur99b, Bur99c, Cas96b, Cas96a, Gsc95b, Gsc95a, Ing96, Nat92, Pom96d, Pom96e, Sin99a, Ste93b, Zwi98, Zwi99a, Zwi99b]. Partially [Duc97, HH95a, HH95b, Pov99]. Participative [Lin94]. Particle [FWF^{+}90]. Partitioned [Vas91]. Partitioning [HS99b]. Party [Kam93, Pom94b]. Passing [AD93, IMC93]. Passive [Bla92, SSK97]. Password [Bis90, Bis92b, Coo95, Cra90, Hau92, Hie92, Jas96, Kle90a, Kle90b, LT91, PM99, Spa92, Tay99]. Password-Protecting [Tay99]. Passwords [Bug96, De 90, JMM^{+}99, KI99, Lib94, MAM95, Rub95, Rub96b]. Paste [BM96]. Patch [Fut95, Tap99, Car98b]. Patches [LL98]. PatchReport [Sin97b]. Patent [Den99b]. Path [HG93, IFT^{+}91]. pathalias [Sal92h]. PATHFINDER [BGP^{+}94]. Paths [Duc94, FP93, MP96, Sin98d]. Patients [Sal96a]. Pattern [BGP^{+}94, GG97, JS97, MSS99, Maf96, Mue96a, Roh96, WM92]. Pattern-Based [BGP^{+}94, MSS99]. Pattern-Matching [WM92]. Pattern-Oriented [Roh96]. Patterns [DLVW98, LH97, Mar96c, NZ99, NLG^{+}99, POS^{+}99, RW93, Sch96g]. Payment [AVS98, Bah96, BN96, BGH^{+}95, BGJY98, GSPT95, KGMP^{+}96, RAM98]. Payments [DBGM^{+}98, VW99b]. PayTree [JY96]. PC [BMLH97, Cho98, Gra98a, Gra98b, Gra98d, LCL98, Rud97, Tro96]. PC-UNIX [Cho98]. PCs [CIC94, Gra98e, Hae92, Ism98]. pdb [May90]. PDC [Car99b]. Peanut [Lea99]. Pebble [BBG^{+}99]. Peek [Dar97f]. Peer [SK96, SMDP97]. Peer-to-Peer [SK96]. Peers [Zwi94j]. PEM [Fra93, GB92, Rat93]. Pen [Car93c]. Pen-Based [Car93c]. Pentagon [Map96]. Pentium [LB96]. People [RMS^{+}99, CG05]. Perfect [Hec96, Sch90, Uhl96]. Performance [AAL92, Ano92, ACTW96, Ber92, BTS^{+}99, BS93a, BS93b, BS96, BPR92, BCVC91, CFL94b, CPW93, Cha99, CBL90, CDK^{+}99, CK96a, CF94, CRS99, CJC^{+}92, CRH99, DWAP94, EWCS96, ES95, Fin90, FHLW99, GP94, GLOR91, HO90b, HCL95, HL93, HPS97, Hum98a, IMC93, IC97, JWS98, Jus94, Kei90, Kol95d, LB96, Laz92, LCC94, LHF^{+}93, LPJI98, MB91, McC98b, McC98c, McC98d, McC98a, McC99d, McC99a, McC99b, McC99c, MK91, MS96b, MBMS95, Mog99, Mue94c, NYKT94, NH93, NLG^{+}99, Par94, Pax93, PS96, PHS96a, PHS96b, RvIG98, RAC97, RMC90, Sch92c, SHK92, SSB^{+}95, SSK97, SD98b, dSGM92, Sim93a, SHM93, SKP94, Spr96b, SW92b, Swa97a, Swa97b, TCI98, TB94, Urn97, Wel91b, Wel91c, WRJ92, lWS99, YR93, ZY98, ZIL96, ZS99, kc99a, BM98b]. Performer [KS98]. Perils [BS90]. Perimeter [AR94]. Peripheral [Spi99]. Peripherals [Hua98]. Perl [Uma97b, And97, Bow98, Chr92, Chr93a, Hal98a, Hal98b, Hal98c, Hal98d, Hal99c, Hal99b, Hop92, Kol94i, Pat97, Pom93a, Pom93b, Pom93c, Pom94e, Pom94c, Pom94d, Pom95c, Pom95a, Pom95b, Pom96b, Pom96f, Pom96c, Pom96d, Pom96e, Pom97c, Pom97b, Sch95a, SW91a, Spi97, Pot97, Sie97a]. Perl5 [Bea94]. Permanent [Cla96]. Permissions [HAB^{+}98]. Permits [Uhl90]. Persistence [DeW91, Ros92, SSZ^{+}93, VBD^{+}92]. Persistent [CRRS93, DVdB^{+}94, KW99, MCLY94, RM97, Rod93, Vau90]. Person [RMS^{+}99]. Person-level [RMS^{+}99]. Personal [BLO^{+}91, BMH98, Haw90a, Joh98, MC99, Man97, Tro96, Whixx, FT99, Pet94]. Personality [PAO93]. Personnel [Kap94]. Perspective [Hic99, Pom97a, Sal92h, Tho96]. Perspectives [Kei90, Sat90]. Pesky [Ros95c]. Petabyte [Mil99b]. Peter [Hac98a]. PGP [MR98, Ros95a, Ros95b, Ros96a, Ros96c, Ros96f, Ros96g, Ros97, Ros98a, Ros98b, SA95a, WT99]. Phantom [Cou95]. Phase [Wil92a]. Phil [Ros95a, Ros96d]. Phone [Sno96]. PhoneStation [Uhl93]. Physical [Dou93, HMF^{+}99, SP91a, Ull97]. Physical-World [Ull97]. Physicians [Sal96a]. Physics [NUO^{+}97]. PIC [Ker82]. Pickling [RWWB96a, RWWB96b]. Picks [Pie99b]. Picture [KH97a, Mel97a, Tra90].

Picture-Drawing [KH97a]. Pieces [Haa90]. Piggyback [KW97]. Pinpointing [Urn97]. Pints [Wil97b]. Pipe [Spi99]. Pipe-Based [Spi99]. Pipes [Sin97a]. Pitfalls [BS90, SPY^{+}93]. Pits [TM97, KM92]. PKCS [Nys99]. PKIX [ZWM^{+}99]. Placement [Bre93]. Places [LO95]. Placing [AS92]. Plan [Lon95, Pik91a, PPD^{+}95, Pre92, PW93, Wel94]. Plan-Based [Lon95]. Planning [Chr98d, Sie97b]. Plans [Kha98]. PLASMA [Kra98b]. Plastic [HS91b]. Platform [BRH^{+}93, Cro98, DBD^{+}99, HO94, JS95, MWCR90, MWCR91, Rei98b, RDM95, SSZ^{+}93, CNTS93, Sam95]. Platinum [FK92]. PLATO [Woo94]. Play [CV96, OH95]. PLC [cC99]. Plea [Pom94e]. Please [Lim98]. PLOD [Pom93d]. Plotting [Pla93]. Plug [OH95, Lon95]. Plug-And-Play [OH95]. PLUG-IN [Lon95]. Pluggable [IH98]. Plugin [Lev96]. Plus [Dav95, SW90]. PN [FT99]. Poet [Tay97d]. Poetry [Hop92]. Point [All96, Cox97b, Cox98c, LL98, Lyc96, SP94]. Point/Counterpoint [SP94]. Pointer [MBMS95]. Pointers [Ede92, Ham97b, Mey91]. Policies [BP99, CFL94a, CL98, CBL90, Cra96, DKB95a, DKB95b, Kol94h, MA90, MU98, Opp96, RLBC94, SS96b, SSL^{+}99, WGC^{+}95]. Policy [ALLO92, BL90b, Bla91, CJ98, Gir99a, HS91c, Kin90, Min95, Mog94a, Mol95, Sat92, War94, ZSD90]. Politics [Nat96]. Polymorphism [GR91, GR93]. Pong [HMS94]. Poor [KS94c]. Port [Joh98, Koo93, Mil98h, YHK96, Bea94]. Port-Address [YHK96]. Portability [FG90, FKSV94, Sel94, SC92, Wal97b, Wal98]. Portable [Arn95b, BL90a, BRH^{+}93, CFLM92, Dea90, HK95a, LKHA94, MS96b, SO92a, SC96, Web93, vL98]. Portals [SP95]. Porting [ACFM99, CDK^{+}99, Dav94a, Gau91, Kor97c, Noe98, TC92]. Portland [USE90i, USE92k, USE92n, USE92o, USE97g]. Ports [SF97b]. Pose [Sno96]. POSIX [Col91a, Sto99a, Alf94, Arc93, Bla99, BC91, Dun90d, Dun91b, Dun91c, Lyc96, Mue93, Sto98a, Wal91b, Wal95b]. POSIX.1 [Col91b]. POSIX.1d [Sto98e]. POSIX.1h [Rot99c, Rot99b]. POSIX.1m [Rot99c, Rot99b]. Possibilities [ES97]. Possible [Mue95c, Str93b]. Postloading [Joh90]. PostScript [Pla93]. Potential [Vix98]. Potentials [VW99b]. Potpourri [Ins95d]. Power [Bea98, Dam95, DKM94, Kol93b, LKHA94, MDS99, Rat98]. Power-Hungry [DKM94]. PPP [Kot98]. Practical [Avo98, CL99, Far92, Gsc95b, Gsc95a, Lea92b, SB93, Sel94]. Practically [Gut98b]. Practice [Bru99a, Lip95, Pin92]. Practices [HS91c]. Practicum [Pom93a, Pom93b, Pom93c, Pom94e, Pom94c, Pom94d, Pom95c, Pom95a, Pom95b, Pom96b, Pom96f, Pom96c, Pom96d, Pom96e, Pom97c, Pom97b]. Practitioner [Lyc94]. Pragmatic [Bis92a, CHS91b, HS91a, MB96]. Pragmatica [Dam93]. Predict [PP99]. Predictable [FC96, TNR90]. Predicting [KL96]. Predictive [GA94, KPR94]. Predictors [VR99]. Preemptive [MD98]. Prefetching [BKW94, Duc99, KTP^{+}96, KL97, LD97, MDK96, SC93]. Prefix [Car94]. Preliminary [Vis96]. Premier [Lie99]. Preparation [Dar96e]. Prepares [Mar91a]. Preparing [Gas98, McD99]. Preprocessing [DHR97]. Preprocessor [GS99b]. Present [Jac94]. Presentation [FSM^{+}91, Pet96]. Presentations [HMT^{+}98]. Presenting [She92b]. President [Zwi94b, Ano97h, Eva96a, Eva96b, Eva96c, Hum96b, Hum98b, Hum98c, Joh92, Joh93c, Joh93a, Joh93b, Joh94d, Joh94b, Joh94c, Joh95b, Joh95c, Joh95d, Joh96b, Mil97a, Mil97e, Mil98e, Mil98f, Mil99g]. Presses [Ano97n, Mau98]. Pretoria [Kol97e]. Preventing [Jas96, SBS99]. Prevention [CPM^{+}98, DH99, Oka98a, Oka98b, SU99]. Prevention/Detection [SU99]. Prime [Men98, BS97a]. Primer [Quo91]. Primitives [BCLC99, GBB93, MMS97]. Principle [POS^{+}99]. Principles [KK99, Sak92a, Sak92b]. Print [All97, Pul99, RFD98]. Printer [PM95b, Sim91b]. Printing [Fin94a, Fle92b, Woo98]. Prior [KL96]. Priorities [Col96]. Priority [Bla91, SCB93]. Priv [Hil96]. Privacy [Ano99-59, Avo98, Bis92c, Gee99, Uma98b]. Private [Cho99, DS92, DBD^{+}99, Gre99, HTK98]. Private-Key [DS92]. Privatization [RP92]. Privilege [NC94, SLS92]. Privileged [Hic99, Hil96]. Privileges [Gil90, WH92]. Proactive [Bis92b, CH95, Hie92]. Probably [Zwi91b]. Probes [GMNR99]. Probing [CL94]. Problem [Bel96b, CF98, FG90, Fin96, HL93, Mog99, Rue96, Tie93b]. Problems [Car91b, Chr90a, Chr91, DK95, KH97b, LB97, Pea99, WCS93, Zwi94a]. Procedure [PA90, Sel90]. Procedures [RL99]. Proceedings [AP99, IEE90, USE90e, USE90f, USE90g, USE90h, USE91e, USE91f, USE91g, USE91h, USE91l, USE92f, USE92e, USE92g, USE92k, USE92h, USE92i, USE92j, USE92n, USE93f, USE93g, USE93h, USE93i, USE93k, USE93j, USE93l, USE94h, USE94i, USE94j, USE94e, USE94f, USE94k, USE94g, USE94l, USE95f, USE95g, USE95h, USE95i, USE95j, USE95k, USE95e, USE95l, USE96e, USE96i, USE96f, USE96h, USE96g, USE96j, USE96k, USE97b, USE97c, USE97d, USE97e, USE97f, USE97h, USE98d, USE98c, USE98e, USE98f, USE98g, USE98h, USE98a, USE98b, USE99a, USE99e, USE99g, USE99f, USE99b, USE99i, USE99h, USE99c, USE99j, USE99d, USE99k, USE99l, USE90j, USE90i, USE91k, USE92o, USE93m, USE98i, USE91i]. Process [BBS95, BM91b, Den99b, FG91, Fre91, Kev93, Kin91, Lea95a, LS92, MZA96, NKM96, RS91, RAF92, Rua90, Rus91, SBS99, Shi97, Uhl90]. Process-Specific [BBS95]. Processes [CG99b, Gau91, Gil90, Lib91a, Lib94, ShCR93, SPY^{+}93]. Processing [Bar98, BZM91, BFRS99, Das97, DB96b, FBAE^{+}92, FL93, Hae95a, Ism98, LWS^{+}94, MB91, SLC94, YTJ91, SD98a]. Processor [Gar90a, OSS^{+}92, Nad92]. Processors [Bla91, KK99]. Procurement [Sie97b]. Product [Gra99c, Kol96k, Sch96a]. Production [AKS98, GMR95, IS91, LR94, SB99]. Productional [Sma91]. Professional [Chr98c, Eva95c, Zwi97]. Professionalism [Eva95b, Eva95d, Lim94]. Profile [Kni96, NKM96, SS94b, SKP94, ZK94]. Profiles [GSS99a, Moh99a, Moh99b, MR97b]. Profiling [Cur94, HG93, LV99, Mar99, MT93, McR93, NKM96, WR98]. Profit [Joh90, Pol96]. Program [ASWW91, And91, Bis90, Che94b, CCK98, CV93, DSS90, ES95, Fly98a, GSS99a, GDFR90, HG93, IS91, KA94, Kin90, LN95, Lin90a, MLL95, PK91, SM99, Yor99, dSM96, HG98, Ano97m]. Programmable [OCH90, SA98, WAP99]. Programmer [AKM99, Col91a, Sto98e]. Programmers [Cof97, Pik94, Sch95b, Spe93b]. Programming [Ano97o, AGG^{+}94, BL90a, BN96, BRH^{+}93, Bis92a, Box96, BD92, CCGS92, Col92a, CG99b, Cou95, Cur97, DAM^{+}91a, DAM^{+}91b, Den92, Ful96, Gar98, Hae94b, Hal94, Hal98a, Hal98b, Hal98c, Hal99c, Hal99b, Har97b, Hop92, HHSW97, Jef94, Jen97, Kol95a, Kol98a, Lan96, LH97, Lea93, Lea96e, Lea98, LS93, Lin94, MSS99, ML91a, ML90, Mog90, MPZD98, NN93a, NPS95a, NPS95b, Par90, Par93a, Puc99, Roo98b, SBB96, Sch94a, SHAS95, Sch94c, She93d, Sie97a, Spe91, STV90, SF97a, TCI98, Van99, WL95, Wil97a, WBSL94, iW91, Yig97, CNTS93, Uma97b, O’N97a].

Programs [AT92, Ano98-44, Cop94, GMR93, GLW95, Gra92a, HJ92, HCL95, JLI98, KMBT92, Lib92, Lib96a, Maf93, Mur92, NN93a, RASS97, RBJ91, Rei92b, Sco93, WSB^{+}96, Zwi91b]. Progress [SW93, WUS^{+}92]. Project [Ber99b, Cow99, HK95a, SM98a, Vel95, WH92, Wil92a]. Projects [Den98]. Promise [Hud97]. Promotions [KRJM98]. Proof [Wri96]. Propagation [IK95]. Proportion [SGG^{+}99]. Proportional [Che98a, WW94]. Proportional-Share [WW94]. Proposal [Ano98-40]. Proposals [Ano97f, Kle97, Kle99a]. Proposed [Sch93b]. Propulsion [RA98]. Prospects [Sco93]. Prospero [Neu92b, Neu92a, NAU93]. Protect [Saf92]. Protected [BTC97]. Protecting [Chr98g, SBV95, Tay99]. Protection [BBS95, HCC^{+}98, JLI98, YBA93]. Protocol [BP93, CGHK95, CH95, CK96b, CTS95, FB96, FLB98, GS96a, GS94c, HR91, Kah95, KOO95, LKY93, Mac91, Mog92, NMMS97a, NN96a, PHS98, Sim96b, SEYG96, SLC94, WS96, YLK93, Hed98, KS98, Lea98]. Protocols [BGH^{+}95, Bel96b, CLD^{+}97, DGT96, GG97, HTWW96, JHC94, KW96, LHB96, Man95, MMR92, NYKT94, NN96b, NK98, Par94, Tan95, ZIL96]. Prototype [ALLO92, AKS98, BSS^{+}95, BSSW96, E^{+}91, SST92, Wie92]. Prototyping [NUO^{+}97, SSZ95]. Provide [BS92a, CL91b, CL91a, NMMS97a, SA95a]. Providers [SA95b]. Providing [BP99, BGÖS98, Fur91, KMS^{+}93, LFR^{+}98, Lim99c, Min95, SKAM99, SB98b]. Proxies [MKPF99]. Proxy [CI97, CDL99, DMF97, Hun97, KW97]. Pseudo [Bel90, CH95]. Pseudo-Network [Bel90]. Pseudo-Random [CH95]. Pseudonymous [BGG^{+}98]. Psyche [SLM^{+}90]. Psychic [Zwi91a]. Pthreads [HC97]. PtTcl [HC97]. Public [Bec99, Byr93, Dav96, Gal96, dPG96, Kap92b, LA98, PB97, TAP90, TA92, Sch92a, HMT^{+}98]. Public-Key-based [TAP90]. Publicly [HSOS99, KS94b]. Publishing [Phi91]. Pulling [Hub98]. Purify [HJ92]. Purpose [GS98, Hal94, HO94, KH97a, WUB^{+}90, BGJY98, FN98]. purposes [Ous96]. Push [MJS97]. Push-based [MJS97]. Put [Den99c]. Putting [O’D90b]. Puzzle [Ano99-43, Ano99-44, Ano99-58, Ano99-56, Ano99-57]. PVM [CCK^{+}95]. Python [Ful96, Gar98, Lut99, Roo98b].

Q&A [Dij98]. QNX [Hil92]. Qualcomm [Kar93]. Quality [BGÖS98, CE99, FK98, FK99, Pol94, Zea97, Mor92]. Quality-Aware [CE99]. Quantitative [LKHA94]. Quantum [MQ97]. Quarter [Kol94e]. QuaSR [GMS96]. Queries [HF91]. Query [Fow94]. Querying [LFB92, WCN93]. Question [Dar97e]. Questions [Fer95, Wil95a]. Queue [GHN90, GHN91]. Queueing [Cho98, Col96]. Quit [Lim98]. Quite [Mac94].

R [BFH^{+}92, AT92]. RACC [ZSBC98]. Race [BD96, PK96, Ste93a]. Radical [Lea94]. Radio [Far95e, ME99]. Radix [MBM93]. Raid [BZM91, LMC94, MB91]. Rainless [Hum98c]. RainMan [PPC97]. Rains [Kaw96]. RAMA [MK95]. RAND [Urb90, Urb91]. Random [CH95, Gut98b]. Randomization [CH95]. Randomized [MT93]. Ranum [RK97]. Rapid [Bea97, CMB^{+}98a, FNP97, Fre97, LeF99, Sch94a, Sch94c]. Rapidly [Bar98, Car99a]. Rapporteur [Dun90c]. rat [SM99]. Rate [DFKM97, SGG^{+}99]. Rates [CL91b, CL91a, PCMI91]. Rather [Zwi91b]. Raven [RN93]. RBK [Kol95e]. RC [Rom91b]. RCS [Chr98a]. Rdist [Coo92]. Re [Eva94c]. Re-engineering [Eva94c]. Reach [Ano94a]. Read [ZD93a]. Read-Only [ZD93a]. Readings [Lea90b].

Reads [OS94]. Ready [Men98]. Real [Ahl99, AL93, AB91, AOG92, And97, Bah90, BDKM93, BB93, BZT99, BSSD96, Bug96, Cam94, CRVZ97, CJ99, cC99, Cla97, CV96, DZL97, Det91, EBG99, FPSW91, Gee99, HS94, Heu90, KTN93, Kol97d, KH97b, hLhCN98, Mel99a, NKT93, Nil94, Ove91, Pac93, Pat97, Pax98, POS^{+}99, Rey92, ST93, SGG^{+}99, TN96, TNR90, TN91, Tra90, WGC^{+}95, ZSG92]. Real-Rate [SGG^{+}99]. Real-Time [AB91, AOG92, BDKM93, CRVZ97, cC99, Cla97, DZL97, HS94, Heu90, KTN93, KH97b, NKT93, Nil94, Ove91, Pax98, POS^{+}99, Rey92, TN96, TNR90, TN91, ZSG92, BZT99, CJ99, CV96, FPSW91, hLhCN98]. Real-World [Pac93]. realistic [BM98b]. Realities [Par94]. Reality [Lan91]. Really [Car90a, DW98a]. Realms [ARH95]. Realtime [KSZ92, NYM91, NYM92, Sch92c, Sto98e]. Rear [All99]. Rear-View [All99]. Rearrangement [AS93, Uhl90]. Reasons [Har96a]. Receive [MR96]. Receiver [DB96b]. Recipes [Spe93a]. Reclaim [BL90b]. Recognition [Eva95c, Kem93, Rub91]. Reconciliation [FP90]. Reconciling [ALLO92]. Reconfigurable [MS93b]. Reconfiguration [GSPW98]. Reconfigured [Man98]. Record [Cro95, FM99]. Record-Breaking [FM99]. Recordable [KLB98]. Recorder [Lal98]. Recoverable [CRRS93]. Recoverable-Persistent [CRRS93]. Recovery [Bacxx, BS92a, FCNL94, Gol90a, Hid94, Mel99b, Mog92, Mog94b, Pre97, Sel98a, Sel98b]. Recursive [FHL^{+}96]. Red [Gra99a]. RedAlert [SC99]. Redesign [GKS90]. Redesigning [PSJL97]. Redirecting [Pat93]. Redirection [CRS99]. Reduce [AS92]. Reduced [WWDS94]. Reducing [Fly97, GA94, LBBA99, Rei92b]. Reduction [BDR97, KL97]. Redundancy [Sri93]. Redundant [SW96, Sim91b]. Redux [Sei94]. Reengineering [Lea95a]. refdbms [GLW94, Gol95]. Reference [BNMK99, Dra91, Pom96c, ZWM^{+}99, dSM96, Chr98c, O’N97b]. References [Gol95, ISS97, SM92]. Refinement [CIM92]. Reflection [Mil96a, Par99a]. Reflections [Car92a, Dar97g]. Refunds [DBGM^{+}98]. Registration [VW99a]. Regression [Lib92]. Regular [Hum97a, Hum99b, KS97, Uma97a]. Reimplementing [MP92]. Rejecting [Har97a]. Related [Hae90e, Hae91b, Hae91c, Sto93e, Sto93c, Sto93d, Sto93f, Sto94b, Sto94c, Wal91c, Wal91d, Wal91e, Wal92h, Wal92i, Wal92d, Wal92e, Wal92f, Wal92g, Wal93c, Wal93d, Kol95i, Sto94a]. Relational [AT92, Fin94b, HF96, WS99, WCN93]. Relations [WCN93]. Relationship [MM96b]. Relationships [BGG^{+}98, Mil96b, Mur93]. Relatively [HF91]. RelaX [MKC92]. Relaxed [KFJ94]. Relays [TW93]. Release [Bos93, CBB^{+}91, CGS^{+}96, DG94, PST^{+}92, ZIL96, EKB^{+}92, Pea92, SLS92, VS90]. Releases [Den99a, McK98, Wat96]. Relevant [Sto94d, Sto94e, Sto94f, Sto95c, Sto95d, Sto95e, Sto95a, Sto96f, Sto96g, Sto96c, Sto96d, Sto97a, Sto97b, Sto97c, Sto97d, Sto97e, Sto97f, Sto98b, Sto98c, Sto98d, Sto98e]. Reliability [DFGG97, Lib94, Sel97b, Sel97c, Sel97d, Sel97e, Sel98a, Sel98b, Sel98c, Sel99a, Sel99b, Sel99c, Wan91]. Reliable [Arn95b, BMHR91, BL97, BS92c, GG97, LC99, LFR^{+}98, Lim99c, MD96, NMMS97b, Nil94, OS94, Par90, Par93a, SBRH96, vBC^{+}92]. Relief [Sim96c]. Rely [WD95]. Remarks [KG93, Pet96, RC93, Sat93c, Sch96h]. Remedies [AM98b]. Remember [Ros95c]. Reminder [Ana99]. Remote [CM94, CG90, DWAP94, GSS99b, KWB^{+}98, MD96, PA90, SMSM98, Sch97b, Sel90, WW97, Rao99d].

Remotely [LW99]. Remoting [WL98]. Removable [Alt93]. Renaissance [Rus91]. Reno [Mac91]. Renumbering [LKCT96]. Repeat [Dar97e]. Repeating [Dar98f, PP99]. Repelling [Far94]. Repetitive [Oka98a, Oka98b]. Replacement [BNMK99, Dra91, Ins95c, MA90, ZD93a]. Replay [Cro95]. Replicated [AS92, FHKR93, GHM^{+}90, PGH^{+}91, SW98, TD92, ZD93a]. Replication [BCH^{+}98, BCC^{+}97, MCU97]. Report [And93, Ano98-28, Ano98x, Ano98y, Ano98-29, Ano98-30, Ano98-31, Ano98-33, Ano98-27, Ano98-32, Ano98z, Ano99-28, Ano99-35, Ano99-31, Ano99-30, Ano99-33, Ano99-34, Ano99-29, Ano99-32, Ant95, Bai97, Ben95, Ber99e, Bla99, Cow99, DF93, Den98, Dun90c, Dun90b, Dun91a, Dun91b, Dun91c, Dyk91, Dyk92, Far95d, Gra91, Hae95b, Har96b, Isl95, Jon98, Lan92, Lee94, McD92, McK93a, Mil98b, Mit93, Mit95, MPZD98, Mur99b, PT93a, Pie99a, Ree97, Ros93, Rot95, Rot99c, Sal97h, Sal97b, Sal92f, Sat95, Sch95e, SS94a, SST90, SD92, Ska94, SW93, Sma92a, Sto97a, Sto97b, Sto98a, Sto98b, Sto98c, Sto98d, Sto98e, Sto99b, Sto99a, Wag92, Wal91b, Wal92b, Wal92c, Wal94b, Wil91a, Wil91b, Wil92b, WUS^{+}92, WKR98, Zwi93d, Per92]. Reporting [Car99a, Sam95]. Reports [Ano94f, Ano96d, Ano99-35, Ano99-31, Ano99-33, Ano99-34, Ano99-32, Aut95b, Aut95a, Bai97, CWT^{+}97, CPK94, Var96, Var97, Num94b, Old94a, Old94b, PS97, Num94a, Ros96i, Rub96a, Rue96, SCP^{+}93, WS93, WP94, WSR93]. Repository [CdR99]. Representation [Che98a, Mur92]. Representing [HF96, RW91]. Reprise [RW91]. Req [Sel97a]. Request [KM92, Wil99, Cha92b, Rhe98, Sha92b]. Requests [Lim99b, Sco97]. Requirements [Swa93]. Research [BG97, Ber99f, LAG^{+}96, LRNL97, Mem93, Ous91a, Tho96, Wil92a].

Reservation [BGÖS98]. Reservations [JR99]. resident [GBK^{+}95]. Resilient [Lee98]. Resistance [AK96]. Resistant [KK99, LA98]. Resolution [Gol99, KS95, RLBC94, SS95]. Resolving [RHR^{+}94]. Resource [AMP98, BH99, BDM99, BGMR91, CMSvE99, ESSB93, HS93b, Kin91, Mit91, MQ97, NB97, SEKN92, Sim91c, Swa93, WW94, NN93b]. Resources [Bar94b, HG93, Mue96a, TJM^{+}99, VDAA99]. Response [Abb95, CF98, Dam98, Dub94, FC96, MW98a, Mil97c, PT93a, SSH93a, SH94, SB93, Sco97, Swa97a, Swa97b, Wiz96]. Responsibility [BS97a]. Restart [Rot99b, Rot99c]. Restore [BMLH97, Kol91, ML93, Han99b]. Restore-o-Mounter [ML93]. Restores [Sel98a, Sel98b]. Restoring [BS97a]. Restricting [LeF92]. Result [SAYZ99, VA98]. Results [And95, Ano97k, Ano98-34, Ber99a, BC94, CHHL99, Dar96b, Kol98m, Mil98d, Pie94a]. Resume [Cla99, Cla99]. RETHER [cC99]. Retires [Wal94a]. Retrieval [AOG92, CSM97, Kre94, Qui94, SM99]. Retriever [LFB92]. Retrofitting [Kam93]. Retrospective [AP99, Dav94a, Kol90]. Reusability [FSV94]. Reusable [Pol94, Spa92]. Reveal [PW97]. Revelation [Kol96a, Leo91]. Reverse [LeF99]. Review [Ann98, Ano99-27, Bar94c, Bar94d, Bar92a, Bar92b, Bar92c, Bea97, Bea98, Bia96b, Bir90, Bir93, Bir94, Bir95, Bor93, Bra92b, Cha96, Chi97, Chr97a, Chr98a, Chr98c, Chr98d, Chr98b, Chr98e, Chr98f, Chr98g, Cof97, Col91a, Col91b, Col92a, Col93, Cuc90, Cur97, Dav94b, Dep94, Eng93, Far94, Far95e, Far96f, Far97a, Far97h, Fia94, Fla97, Fly98b, Fly99a, Gil93, Gor90, Gra99a, Gwy90, Hae91a, Hae95a, Han97, Han99b, Haw90b, Hec94, Hum97a, Hum98a, Ins96d, Jen97, Jen98, Joh97a, Kat94, KC93, Kno97, Kol93b, Kol94f, Kol94g, Kol94h, Kol94i, Kol94e, Kol95a, Kol95i, Kol97a, Kot98, Lal98, Laz97, Laz98, Laz99, Lea90b, Lea92b, Lea93, Lea95a, Lea95b, Lea96b, Lea96f, Lea96d, Lea96c, Lea96e, Lea97, Lea98, Lib95, Mon93, Mos94a, Mos94b, Mos96, NN92a, NN93b]. Review [NN94a, NN94b, NN94c, NN95b, NN95a, NN96a, NN96b, O’N97a, O’N97b, O’N98, Pax91, Pax93, Pax94, Pom97a, Pot97, Quo91, Roo97, Roo98a, Roo98b, Roo99, Sal93g, Sal93f, Sch95b, Sch96d, Seo93a, Seo93b, She92a, She92b, She93d, Sie97b, Sie97a, Sie98a, Sim92a, Sim93a, Sim94a, Sim96b, Spe93a, Spe95b, Spe95a, Spe97a, Ste95, Ste92, Sub98, Uma96a, Uma96b, Uma96c, Uma97a, Uma97b, Uma98a, Uma98b, Van99, Wal92a, Yig97, Zwi93a]. Reviewers [Ano99-27, Bin99]. Reviews [Far92, Sal96c, Sal97c, Sal97d, Sal98b, Sal98c, Sal98d, Sal98e, Sal98f, Sal98g, Sal99g, Sal99h, Sal99i, Sal99j, Sal99k, Sal99l, Sim92b, Zwi92b]. Revised [Dra90, KS90]. Revision [Sin99b, Rud97]. Revisitation [SM97]. Revisited [Kam98b, ML93, Nem99b, SCO90, Sin99b, Wil98]. Revocable [BL96]. Revocation [NN98]. Revolution [Bas98, DF95]. Rewritable [TSE90]. Rewrite [Sch94a, Sch94c]. Rhymes [Hae94a]. Rich [Far99g]. Riding [Ano98-39, Mel98]. Right [Jac94, Zwi92c]. Rik [Car98a]. Ring [GHN90, GHN91]. RIPE [ML90]. Risc [WIK92]. Risk [FY98]. Risks [VW99b, Sim96d, Kol95i]. River [BCLC99]. RIVL [SS95]. rlogin [KOO95]. RMI [BCH^{+}98, KWB^{+}98]. Road [Kap92b, Tho96]. Roadmap [JP99]. Rob [Kol99a, RK97]. Robot [ML90]. Robotic [SKS92]. Robotics [PH99]. Role [Fle93, Kolxx, Lea96f, McN93]. Role-based [McN93]. Roles [Gil90]. Roll [Tay97b]. Rollouts [Hei99]. Romanian [Rot95]. Room [Bia96a]. Root [Sim90, Sim91a, Tho98a, WSB^{+}96, dLRT93]. Rootkits [Bru99a]. Rope [Hub98]. ROSE [Rot95]. Round [Dij99a]. Route [LO90, SB98b]. Route-Specific [SB98b]. Router [Cal96, CHS91c, LO90, Oet98b]. Routers [Cho98, Chr98e]. Routing [BCC^{+}94, Chr98c, KCVP95, Lar99, RMS^{+}99, Skl91, Swa92, YL99, kc99a]. RP3 [BCR91a, BCR91b]. RPC [Duc91, ESS96, Mat97, MMR92, SSH93b, Tie93b, WWWK95, YBA93]. RSA [Dav95, FY98]. Rscan [Sam95]. Rsh [Ins95c]. RSI [Oka98b, Har99a]. RT [KTN93]. RT-IPC [KTN93]. RTX [CRVZ97, CJ99]. Rule [Dar98d, LaP94]. Rule-Set [LaP94]. Rules [Gee98, Han92, Sch94a, Sch94c]. Run [BL90a, Det92, FBAE^{+}92, ISS97, Lea92a, NL96, SL92, Man91]. Run-Time [BL90a, Det92, FBAE^{+}92, ISS97, NL96]. Running [CH97, DSS90, GMR93, RH93, HG98]. Runtime [IL90, IL93]. Runtimes [BS97b]. Rush [BCC^{+}97]. RUST [Rue96].

s [Lea99, Rus99a]. Safe [Bor98, Che97, ED94, Kah95, KV91, KS95, LO95, Mue96a, NL96, Sma97, MdR99, LB98, LDOW98]. Safe-Tcl [LB98, LDOW98]. Safe-Tcl/Tk-based [LB98]. Safe/Fast [KV91]. Safety [Cra96, For99b]. SAGE [Ano99v, Ano93b, Ano95d, Ano97k, Ano97l, Ano99-53, Ber99d, Ber99e, Dar93c, Dar95b, Dar95c, Dar96f, Dar96b, Dar96h, Dar97c, Dij98, Dij99b, Eng93, Eva93a, Eva94a, Eva96c, Far92, Fis96b, Gas96, Jac93, Kni96, Kol92, Kre97, Lan94a, Lan94b, McD92, McD93a, McD93b, Mil95, Mil97b, Mil97d, Mil97a, Mil97e, Mil99h, Mil99i, Mor92, NZ93a, Nat92, NZ93b, Nat93, Par93b, Par93c, PEZ93, Sat93e, Sch93a, Sch93b, Sim92b, Sim92a, SD92, Sim93a, Sim93b, Sma92b, Var92a, Var92b, Wil95c, Wil95b, Wil98, Zie93, Zwi92b, Zwi92c, Zwi93a, Zwi93d, Zwi93b, Zwi93c, Zwi94b, ZK94, Zwi94c, Zwi94d]. SAGE-AU [Jac93, Lan94a, Lan94b]. SAGE-AU’99 [Ano99v]. Salamander [MJS97]. Salary [Kni96, ZK94]. Sales [KRJM98, Uma95]. Salt [USE95f]. Samba [All97, Kha98]. Same [AM97]. SAMM [Kub93]. Sample [Kol97b]. Sampling [MT93]. San [USE90j, USE91f, USE92g, USE92j, USE93k, USE93j, USE93l, USE94h, USE94l, USE96i, USE96k, USE97c, USE98f, USE98i, USE99f, All92, Pea93, Wil98]. Sandbox [GMPS97]. SANE [Ano98d, Ano98f, vW98]. SANS [Ano98-28, And95, Ano96d, CWT^{+}97, Chr93b, Dar98e, Met95, USE94m]. Santa [USE93h, USE93m, USE94k, USE97b, USE98c, USE99g, USE99l]. SASE [VcC97]. SATAN [Ins95a, Chr98g]. Satisfaction [Kub92]. satool [MSN93]. Saturated [HJ97]. Saturday [SMH95]. Saved [FST91]. Sawmill [SO94]. SC22 [Dun90c, Dun90b]. Scalability [Gun98a, Gun98b, SDH^{+}96, VDA^{+}98]. Scalable [AP97, Bal92, BM98b, Bar98, CBB97b, CBB97a, Cou97, GSS99b, Hei96, HSE96, LC99, SC99, STM^{+}99, SUK92, SUZ98, Ver97, YCE^{+}97]. Scale [Ano97g, BA92, BFRS99, CMS93, Fut95, GES98, GMS96, Hea97, MW98b, NWEF92, Pol97, RFD98, Sca97a, SA95a, Ske92, TC92, USE97a]. Scaling [Gru96, HW93, SA95a]. SCCS [Chr98a]. Scenario [NLG^{+}99]. Scenario-Based [NLG^{+}99]. SCFS [IHR99]. Scheduled [BS97b]. Scheduler [BDMVL93, CJ98, Ess90, GGV96, GV91, MR91]. Schedules [Bea97]. Scheduling [BH99, Bla91, BC97, CJ98, CFHB99, CS90, DM92, ESSB93, FS96, HJ97, Hel92, KSZ92, KH97b, hLhCN98, MP90, MD98, Rus91, SGG^{+}99, WW94, WWDS94, SCO90]. Schema [HF96, WCN93]. Schemas [Fin95a]. Scheme [Dar91, HJ97, Lee98, PM99, Ben94b, JB99]. Schemes [RAM98, SM97]. Schizophrenic [SSCK93]. School [Ano97m]. SCI [SH98a]. SCI-based [SH98a]. Science [Gra93, Kol94g, Rus99b, Sto96b]. Scientific [BPR92, EG97, Ker95]. SCO [RHG98]. Scope [Car95]. Scoped [Lu97]. Scores [Kol98l]. Scout [MP96, SP99b]. SCRAPE [Kin91]. Script [Lan96, MK98]. Scripting [AM98b, Bea96, Coo97b, Ing96, NZ99, SB94]. Scripts [Kis90, Kis91, Lib91a, Mue94b, Oka92, Lib96b]. SCSI [NC92, Pom94b, Rob92]. SE [Sal99b]. Search [CHHL99, CD98, MW94, MB97, Sch96d, Tay97b, Uhl96, VcC97, Ver97, Vo96]. Searching [Fre97, HS91d, MSG97]. Seattle [USE92i, USE96f, USE97a, USE97f, USE98e, USE98a, USE99b, USE99i, USE99d, Hum98c]. sec [MBH94]. Second [Bea98, Var97, the98, GLOR91, HBvdRA98, IEE90, OW94, USE96f, USE96g, USE95i, USE96h, Hec94]. Second-Generation [GLOR91]. Second-Level [OW94]. Secondary [MKPF99]. Secretary [Gas96]. Secrets [Kol95a]. Secure [AHB^{+}98, AUH99, AR94, BGH^{+}95, BCE98, BCC^{+}94, BGG^{+}98, CGHK95, Che90, Gal96, dPG96, Gil96, GWTB96, Gon96, Gra96, Gut96, Hil96, HAC^{+}98b, Ins95c, KS99, KOO95, Lee98, MPH99, MK99, NL98, Nee90, PHS98, Pre99, RG95, SSH93b, SK98, SM99, SA98, SEYG96, Sug92, TA92, Tho98a, Won90, WBSL94, Yah94, YT95, Ylo96, d’A92a, dC93, CG05, PB97, VTB95]. SecureBase [Kam93]. Securely [Cox97b]. SecureShare [Bor98]. Securing [BE98, Kra99, LeB99]. Security [ARH95, AHB^{+}98, Ann98, Ano97c, Ano98p, Ano98-28, Ano98-30, Ano99p, Ano99q, Ano99-32, AM98b, AKM99, Bah90, BHJV92, BK93, BKS90, BVAD98, Bel96b, BFM92, BCR90, Bur99a, Bur99b, Bur99c, Car90b, Car93b, CL98, CWT^{+}97, Cas95, Cha92c, Che98b, Chr90b, CTS95, Cra90, Cra94a, Cra96, CG05, DBC^{+}99, DS92, Dav95, DG95, DF93, ESS96, FS90, Far92, Far94, Far95d, Far97a, For98, GB92, GF95, Gee94, Gir99a, GMPS97, Gra99d, Gre99, Gut99, HK92, HT92, HT94, Hie92, Ins95d, Ins96b, IB93, Kam93, KS94b, KW96, Kle90a, Kle90b, Kol94h, Kol97b, KYBR99, Kra98b, Kur90, Lam97, Lea97, LDOW98, Lib94, Lin90a, Mah96, MMS97, McK93a, MS98, MU98, Mon93, Muf95, Muf99, Opp96, Per98a, Pom97a, PV95, Ran99, Rao98d, Rub96a, Rub98, SSH93a, Sat90, Sat92, SS99, Sel99b]. Security [SP96, SSL^{+}99, TIF^{+}97, USE90i, USE92m, USE93m, USE94m, USE95f, USE96i, USE98i, USE99e, Ven96, Vix95, WCO^{+}90, WCR99, WD95, Bru97, McD92, Lea92b, Roo98a, Roo99]. sed [Kno97]. SEDMS [USE91j, USE92l, USE93j]. See [Dar99b]. Seek [AS92]. Seen [Arn91, Kle98a, Kle98b]. Segments [GSB^{+}93]. Selecting [Pul98]. Selectively [Har97a]. Self [ES92, Han99a, Man98, PBS97, SBV95]. Self-configuring [PBS97]. Self-Organizing [ES92]. Self-Paging [Han99a]. Self-Protecting [SBV95]. Selling [BBD^{+}96, Ran99]. Semantic [HS93b]. Semantically [RW91]. Semantics [BS96, Eng97, MQ97, WH97]. Semi [RS91]. Semi-Loosely [RS91]. Seminar [Ano97m, Sel95b]. Sender [BJM^{+}96]. Sender-Managed [BJM^{+}96]. Sendmail [Car93d, Kol97i, Han97, Fia94, Har97a, BRW99, Sze99]. sendmail.cf [Dar91]. SENDS [SV94]. Sensitive [Tay95b]. SENSS [Muf99]. Separation [Mar91b, Mar93]. SEPP [Oet98a]. September [USE91f, USE92m, USE93k, USE93j, USE94h, USE95h, USE96j, USE98f, USE98b]. Sequential [Mas91, RvIG98, SHM93]. Series [Ano98-37, HL93, Far96i]. Serpent [BHK^{+}90]. Server [AAO92, All97, BD97, BDM99, BCH^{+}98, Bar97, BEM91, BKM97, BH91, CDL99, Det91, Don99, DB96b, ES90, FR90, Fin91, Fin97c, GLP^{+}93, GS94a, Gue92, HSE96, HLM94, IC97, Jus94, Kei90, KW93, Kra99, LPJI98, hLhCN98, LSR95, Man91, Mat98, Mil99e, Mog99, OS94, PT92, Roo92, SBB96, Sch95a, Sos95, Swa93, Swa97a, Swa97b, TN96, Tom98, VGT95, WK93, YL98, YL99, LFR^{+}98, Lim99c, lWS99]. Server-Side [SBB96]. Servers [BMHR91, BGG^{+}98, CFHB99, GS94c, HR91, HHLS90, HPS97, KSKD99, LeF99, LHF^{+}93, NKT93, SP99a, SHM93, SJ95, Swa92, SCD94, VF95, ZSBC98, ZBCS99, BM98b, Chr97a, Chr98d]. Service [BN96, BD93, BGÖS98, CD97, Cat90, CBB97b, ED92, Fow93, FK98, FK99, HK92, How90, How91, JS97, Lin90a, MC99, MBM95, NR94, RFD98, Ros96f, Ros96g, Ros97, Ros98a, Ros98b, SB98a, Sch96b, SA95b, SP99b, TD92, VDA^{+}98]. Services [ABB^{+}97, Bah96, DLO92, DBD^{+}99, EW90, GF95, Gee94, Gre99, GB97, HMS94, HJK^{+}92, JS97, JCS^{+}91, KMS^{+}93, LeB99, LFR^{+}98, Lim99c, Mar96a, MDC95, MFM95, MW96, Sel97d, Ska94, SP96, YCE^{+}97, d’A92a, dC93]. ServiceTrak [Fin99a]. Serving [Bel96a]. Session [BB95b, FL96, Man98]. Session-Layer [BB95b]. Set [Ben94a, Don99, LaP94, McK98, Mil99d, Mil99e, Tan95, Ulf95]. Sets [KS97]. Setting [Col96, Cox97b, Mat99]. Setup [Sch91, SV90a, EL92]. Seven [Dar97h]. Seven-Year [Dar97h]. Seventh [Ano98-30, Ano99w, USE93f, USE98i]. SFIO [KV91]. sfs [LIN^{+}93, Rob99]. sh [Hae94b, Sin98b, Sin99a]. Shadow [Hau92]. Share [Ess90, MR91, WW94, BGMR91]. Shared [AAL92, Bal92, BTC97, BS93a, Boy91, BCCR91, CLMS91, CD92, CFH^{+}93, CMS93, CGCS91, CGS^{+}96, Fle93, GKAS99, GSB^{+}93, GM93, GS94b, GL93, IC92, ISS97, KTBL93, KCDZ94, Kel99, KFJ94, MSL94, MS93a, Mil91, Min93b, OLB93, OW94, Sab90, SL94, See90, SS97, SUK92, ZSB94, ZIL96, SSK97]. Shared-Bus [OW94]. Shared-Memory [CLMS91, MS93a, OW94]. Sharing [AHB^{+}98, BTC97, BS93a, Bor98, EBG99, Ham96, IS99, KA94, Lu97, MWCR90, MWCR91, MK99, Oet98a, WVS^{+}99]. Shaving [RAM98]. Shell [Abb94, Fow93, LaS90, Lan96, Mue94b, NC94, Spi97, Uhl90, HR93, Ram94, Bor93, Mos94b]. SHIFT [AG97]. Shootouts [Gir99b]. Shopping [HT96]. Short [Ano98-37, BCC^{+}94, Dar93b]. Short-Cut [BCC^{+}94]. Showcase [Den99c, Ano99h, Ano99i]. Shownet [Cha92a]. Shuse [Spe96, Spe97b]. SIC [Klu90]. Side [BBD^{+}96, Hum99b, Kle98a, Kle98b, Ran99, SBB96, KSKD99]. SIFT [YGM95]. Sign [GC98]. Sign-On [GC98]. Signal [BM91b, Ism98]. Signature [GR91, GR93, JY96]. Signature-Based [GR91, GR93]. Signatures [BR94, Wri96]. Signed [Sim96d, SKAM99]. Signing [Ros97, Ros98a, Ros98b, Ros96a, Ros96f, Ros96g]. Silicon [NFQ99]. Silicon-On-Insulator [NFQ99]. SIMD [KI99]. SimICS [MLM^{+}98]. SimICS/sun4m [MLM^{+}98]. Similar [Man94]. Similarities [BM98a]. Simmons [Sim92c]. Simple [AK94, Bel96a, Cou96a, Cou97, DBGM^{+}98, HK97, Jon95, Pom94b, Spe97c, TN96, WWW95]. Simplicity [Hal98c]. Simplifying [Sto93b]. Simulation [AG97, Bed90, FWF^{+}90, Klu90, MM92, NUO^{+}97, RAV95, RA98, ShCR93, WUB^{+}90, WW91]. Simulator [GNMR97, SMSM98]. Simulators [BM96]. Simultaneous [Van91]. Since [Sal99f]. Sing [Bor92]. Single [ARH95, CBB97a, Duc94, GC98, KC96, LL98, MI93, Mur99a, MWO^{+}93, Rei98b, SM98a, Wal95c]. Single-Address-Space [KC96]. Single-Platform [Rei98b]. Site [Bur95, Fin97b, FP90, Tay96e, Tay97e, Tay99]. Sites [CM94, FST99, Sal92g]. Sitting [MI93]. Six [Pre98]. Sixth [CWT^{+}97, Rub96a, USE92e, USE96i, USE98f]. Size [CS92, CS94, Her97]. Sizes [GS98, MBMS95]. Sketch [MM93]. Skills [Mue94a]. Slab [Bon94]. SLAC [CCW97]. Slate [RL98]. SLIC [GPRA98].

Slicing [DT97, RRB99]. Slicing-Based [DT97]. Slick [Tay97a]. slides [DW98a, DW98b]. SLINK [Cou96a]. Sloth [GBS^{+}95]. Slow [KS99]. Small [GK97, Haa90, HW93, Koo93, SP99a, Spi98, Tau91, Wal92a, Wan91]. Smart [AFJ^{+}99, AOW96a, Ede92, Gut98a, Ham97b, Kel95, KCCT94, MM93, Mil99k, SGM91, YCE^{+}97, AG97, AOW96b, AG99, Bak99, BF99, Gir99a, GSTY96, LW99, NFQ99, SS99, SKAM99]. SMART-AHS [AG97]. Smartcard [Ano99z, USE99j, IH99, KK99]. Smartcards [DR99, IHR99, MDS99]. Smashing [Ins97b]. SMB [EBG99]. SMTP [Har90, Har91]. SNA [Laz99]. Snake [Cur99]. Sniff [Bis92a]. SNMP [Maz96, Zwi98, Zwi99a, Zwi99b]. Snort [Roe99]. SNP [WBSL94]. Social [WD95]. Socket [BC94]. Sockets [RAC97, VS90]. SOCKS [KK92a]. Soft [hLhCN98, Mac94, EL94]. Software [Ano99-45, Arc93, Atk96, AG99, BBD^{+}96, BCK94a, BKW94, BK93, BCK94b, Bea97, BSKH93, Ber90, Bir90, Bir95, BAKW92, cC99, Col96, CW92b, DBGL^{+}93, DFG^{+}96, Den95, Den99b, Dep94, Eir94, EM99, EL94, Far97h, Fle92a, Fut95, GES98, Gra99a, Gut98b, GL99, Hai97, Har99b, HK92, Hen90, HCK^{+}98, Jen97, JV95a, KFJ94, Lac93, LR94, Lal98, Lam91b, Lea90b, Lod93, LCL98, MWCR90, MWCR91, MO99, MR97a, Mue96c, MM96b, MR97b, Oet98a, OG95, PBS98, PS93, Rei98b, Roh96, RM94b, SK96, Sel91, Sel94, Sim94b, SB99, Sjö99, Sma90, Sma91, Smi91, Spe97a, SG97a, TMJY91, TB99, VCV92, WH92, WB93, Won93, Wyb90, ZSB94, Zwi91a, Sta98, Gor90]. Software-Only [cC99]. SOI [NFQ99]. Solaris [Chr98d, Alt93, Chu96, Dam95, KBM^{+}96, MD95, PKA98, Pom94f, Sec95, ZY98, Zwi93d]. Solid [Gut96, Mil96b]. Solid-State [Gut96]. Solution [Ano93c, Ano99-43, Ano99-44, Ano99-58, NN94a, RH98a, RH98b, WWDF98].

Solutions [Sat90, Spe97a]. Solved [FG90]. Solving [Tie93b, Van91]. SOM [Ham96]. Sombrero [SM98a]. Some [Bed90, CBL90, Hua98, Joh93a, Met95, Mue95d, Rom91b]. Something [Phi95]. Sometimes [Han92]. Somewhere [Dar95e, Kol97e]. Sort [FM99, MBM93, SP99a, BCLC99]. Sorting [Sin98c]. Source [CdR99, FP90, Gra98a, Gra98b, Gra98e, Gra98d, Gra98c, Gra99c, Gra99d, Gra99e, Gra99b, Har99b, MP99, PW97, Sat93b, Sto99b, Too98, hvESV^{+}97]. Source-Level [MP99]. Sourcebook [Far97a]. Sources [BM98a]. South [LRNL97]. Space [EP93, Koe92, KC96, MWO^{+}93, SM98a, Vix99, YBA93]. Space-Efficient [Koe92]. Spaces [AFJ^{+}99, Mil99k]. Spacio [IS91]. Spacio-Temporal [IS91]. Spam [CF98, Cox97c, Ins97a, Mue98, Sze99, Har97a]. SPAND [SSK97]. Spanner [Cox98d]. Sparcstation [Uhl91]. Speakers [GH98]. SPEC [WN92]. Special [KH97a, Loe96]. Special-Purpose [KH97a]. Specialization [ZS99]. Species [Pom97c]. Specific [AAL92, Ano97i, Ano98h, Ano98x, Ano99f, ABB^{+}97, BBS95, FNP97, FB96, Fuc97, Hud97, KS97, LM99a, SB98b, TMC97, USE97b, USE99a, WR98, GZ94]. Specification [Bah96, FST99, FK98, LM99b, Mas91, NN94b, Vis96, VK98, Wal95c]. Specifications [SU99]. Specifying [HT92, HT94]. Speculative [CG99a]. Speeches [Bar94a]. Speed [ATS94, Dar98b, Lee94, MPS98, MZA96, PHS96b, TJH^{+}94, USE94f, SHAS95]. Speedier [Mog99]. Speeding [Hau94]. Speedup [WRJ92]. Spent [Dew97]. Spike [CGLR97, GL97]. Spin [DKB95a, DKB95b]. Spin-down [DKB95a, DKB95b]. SPINACH [PB97]. Spins [Pom95c]. Splicing [CRS99, SMD^{+}94]. Splitting [Lim99a]. SPM [Coo95]. Spool [CBB97a, RFD98]. Spooler [PM95b].

Spooling [Wag90]. Spot

[Fly99f, Fly99b, Fly99c, Fly99d, Fly99e].

SPP [Sel94]. Spread [JABA98].

Spreadsheet [RRB99]. Spreadsheets [RRB99]. Spring [Kol98m, USE90b, USE91b, USE92b, USE92f, USE93b, USE94b, USE95b, USE96b, HK93, NR94, RHKP95]. Springs [USE90f, USE91e]. Sprite [DOKT91, HO90b, Kup93, Wel91b, Wel91c, Wel92, Wel94]. Spritely [Mog92, Mog94b]. SQL [Arn98]. SQL 2 HTML [Fin95a]. Squid [Mog99]. SRA [SSH93b]. SRASS [Rot99c, Rot99b]. SSH [Far97h, Tho98a, Ylo96]. SSL [MSS98, WS96]. ssmail [BRW99]. SSU [Tho98a]. Stable [HF91, Lei99]. Stack [Ins97b]. Stacked [Sto92]. StackGuard [CPM^{+}98]. Stacking [SW93]. Staff [Ano99-51, SP99a]. Stage [Che97]. Stage-Safe [Che97]. Staging [SeaBP99]. Standard [Arc93, Bla99, Col91b, GM93, KCDZ94, Mur99b, Nys99, Rat94, RLBC94, Rot99c, Sto98a, Sto98e, Sto99b, Sto99a, Sun97, BCE98]. Standardization [Dun90a]. Standards [Dun90d, Dun91a, Gar90b, Hae90a, Hae90b, Hae90c, Hae90d, Hae90e, Hae91b, Hae91c, Mar91a, Sto93e, Sto93c, Sto93d, Sto93f, Sto94d, Sto94e, Sto94f, Sto94b, Sto94c, Sto95c, Sto95d, Sto95e, Sto95a, Sto96e, Sto96f, Sto96g, Sto96c, Sto96d, Sto97a, Sto97b, Sto97c, Sto97d, Sto97e, Sto97f, Sto98b, Sto98c, Sto98d, Sto98e, Sto99b, Wal91c, Wal91d, Wal91e, Wal92h, Wal92i, Wal92d, Wal92e, Wal92f, Wal92g, Wal93c, Wal93d]. Stanley [GMR95]. Star [AW99b]. STARS [BH91]. Start [the98, HBvdRA98]. Starting [Cox97b, Cox98c]. Startup [LBBA99]. State [FST91, Gut96, Joh93b, MSS98, RWWB96a, RWWB96b, RFR99, Sjö99, kc99a, Rom90, Rom91a]. State-Driven [Sjö99]. Statement [Ano99-59]. Statements [Ano95a, Ber99c]. Static [PR94, Rei92b, Ste93a]. Statically [Mur92]. Station [Pre99, vR95]. Statistical [BH91, Mar99, SBRH96]. Statistics [ACTW96, kc99a, Spe95b]. Status [JCS^{+}91, Ros95b, SST90]. Staying [Wan91]. STEL [VTB95]. Step [Dij96b, Kol91, Lim99b]. Steve [Kol96h, Sin98a]. Stevens [Far99g]. Stewart [Kol98e]. still [SMH95]. Sting [Sav99]. STL [Gla96]. Stochastic [Klu90]. Stock [WB93]. Stop [Ano97n, Mau98]. Storage [AOG92, ACG^{+}98, AUH99, BKM97, BS92c, DK95, DCK^{+}94, Eng92, GBK^{+}95, HK92, IFT^{+}91, KSS93, MKPF99, OS94, RBB^{+}91, RO90, SKS92, TN96, TJH^{+}94, dC93]. Store [Wei95a, KW99]. Stories [Ann98]. Story [Ano96c, Bob99, Sto99b]. Strain [Oka98a, Oka98b]. Strange [Chr98b, Dar97d, San90]. Strata [Poo94]. Strategic [MS95]. Strategies [BA91, CJ95, Cum99b, KW99]. Strategy [GG97, Hae91a, Pom94b]. Stream [KSU92, MMR92]. Streams [CL91b, RASA98, Rao99c, WGC^{+}95, Gar90a, SPY^{+}93]. Street [HRT96, Kap94, KS98]. Stress [Sim96c, Pea93]. Stretching [Her97]. Strikes [Cuc96]. String [HS91d, KS94c, KV91, Spe92, MdR99]. String-Based [KS94c]. String/File [KV91]. Strings [KS97]. Striped [HO92]. Striping [CL91a]. strlcat [MdR99]. strlcpy [MdR99]. Strong [Gut98b]. Strongboxes [HS96]. Stronger [MAM95]. Structural [Leo91, MR95]. Structure [Bvv91, CSM97, Dun90d, GC96, HF96]. Structured [BHS95, BCCR91, CS92, CS94, SBMS93, TCI98, KSS93, MS92]. Structures [Fre97, Lea94, SB97a]. Stub [Par95]. Stubs [HO90a]. Student [All96, Ano97f, Ano98-40, Ber99f, Cow99, NN92b, Shi98]. Students [Ano96e, Ano98-42]. Studied [Che92]. Studies [PS96, SF97b]. Study [CMB^{+}98a, CJJ^{+}98, DFKM97, DH99, EBG99, GGI^{+}91, GS99a, IMC93, KOO95, KW97, Mog99, Pol97, RvIG98, RMC90, Sch93c, SPR96a, SUZ98, WIK92, WD95]. Stuff [BBD^{+}96, Sch95f]. Style [Ano98-39, Lea93]. Subclassing [BOM95]. Subcommittee [Ano99-53]. Subsequences [PP99]. Substrate [Hai97, MJS97]. Subsystem [Bir95, CRVZ97, CD94a, DB96b, Fre91, Wal97b, Wal98]. Subtyping [BOM95, MR95]. Success [MS95, Sto99b]. Successful [Dep94, Lea96f, Sto96a]. Succumbing [Kle98a, Kle98b]. Suite [Hau92, LKY93, Ver97, YLK93, Pea93]. Suited [BCF^{+}93]. Summaries [Blo92]. Summary [Abb95, Ano95d, Mil95, You93a, You93b, You93c, You93d, You94a, You94b, You94c, You94d, You95b, You95c, You96a, You96b, You97c, You97d, You98b, You98c, You99b, You99c, You99d]. Summer [Dew97, Kol90, USE90c, USE90g, USE91c, USE91g, USE92c, USE92g, USE93c, USE93g, USE94c, USE94j, USE95c, USE96c, WUS^{+}92, CPK94, Old94b, Ros93, SCP^{+}93, Sal92f, WP94]. Sun [Hum98a, ACO92, MMR92, Ove91, SKS90]. SUN-Transputer [ACO92]. sun4m [MLM^{+}98]. Sunday [SMH95]. SunOS [KSZ92, LeF92, PKB^{+}91, Ste92]. Super [Zwi92b]. Super-Users [Zwi92b]. Supercomputer [Chr90b]. Supercomputing [MK93]. Superhighway [Bia96b, Var95b]. Superhighways [McR95]. SuperInformation [Sch95g]. Superuser [CCM^{+}90, Car93d, Sug92]. Support [Ano92, BB95a, Bal98, BCF^{+}93, BDD^{+}93, BRH^{+}93, CJ98, CFH^{+}93, CRRS93, CGCS91, CSL^{+}97, DAM^{+}91a, DBCF95, Den98, DHR97, FBAE^{+}92, GS98, Gau91, HS94, HR91, Hov91, iIYL95, JS95, KCG98, Laz92, LV99, Lie99, LR98b, MBM95, MU98, MLM94, MZA96, NAU93, NZ99, RFW93, RM90, Ros92, Rus99b, SK98, Sco97, Sel90, Ske92, Spe97c, SSL^{+}99, STV90, SF97a, SMPR98, Uma95, VR99, Web93, YL99, YR93, Eng93].

Supported [Gri98, PCMI91, Vau90]. Supporting [AW99a, BZCS96, BCC^{+}91, BCCR91, Cop94, FK92, KS94b, KP93b, KC99b, LS92, PE92, SMDP97, VBD^{+}92]. Supports [Min93b]. Suppressing [SW98]. Suppression [BZT99]. Surfer [Kle99b]. Surfing [PP99]. SurfIt [Bal96]. Surprises [Kol96i, Mel97c]. Survey [And95, GS94c, HS91c, Kle90a, Kle90b, Kni96, Kol94a, Mil98d, Won90, ZK94]. Surveys [Tay97c]. Survival [Mue94a]. Survive [Mue96b]. Surviving [Pom94f, SESS96]. Sustaining [Kol95d]. SUUG [Gra91]. SVR4 [BGH^{+}92, CHS91a, Cur97, SPY^{+}93]. SVR4.2 [Nad92]. SVR4/MP [CHS91a]. SWAPEROO [DBGM^{+}98]. Swatch [HA92, HA93]. Sweet [Dar98c]. SWiFT [GSPW98, HCK^{+}98, CL91a, LMC94]. Swift/RAID [LMC94]. SWIG [Bea96]. Switch [CB96]. Switched [CJ95]. Switches [Bec99, GSPT95]. swm [LaS90]. Sylvan [BCVC91]. Symbolic [Det91, GG98, May90]. Symbolism [Fen96]. Symlinks [Mot91]. Symmetric [Sch95b, ZY98]. Symphony [WW91]. Symposium [Ano97a, Ano98l, Ano98j, Ano98-33, Ano99m, Ano99g, Ano99-31, Var97, DF93, Far95d, Hae95b, Lee94, McK93a, Mit93, Mit95, Rub96a, Rub98, USE91h, USE91j, USE92l, USE92m, USE93h, USE93i, USE93k, USE93j, USE93m, USE94i, USE94f, USE94k, USE94g, USE95f, USE95i, USE96i, USE96f, USE97h, USE98a, USE98i, USE99e, USE99h, USE99c, USE99d, Wal94b, WS93, Ant95, Ano97c, Ano98p, Ano98k, Ano98-30, Ano98z, Ano99p, Ano99q, Ano99l, Ano99n, Ano99-32]. SYN [Ins96f]. Synchronization [GC96, Heu90, Rei92a, Rua90, Rus91, SM97, TN91, ZSG92]. Synchronize [Chr97b]. Synchronized [SW91b, Yah94]. Synchronous [CD94a, PK93, SMDP97]. Synctree [LL98]. Synergy [GC96]. Syntax [Cre97, Gra92a, WAKS97]. Synthesis [MP92, Whixx]. Synthesizing [SU99]. Sys [May95, Sha92b, dL95]. Sysadmin [BS97a, Mil99j, Nem99b, Cuc95, Cuc96]. SysAdmins [MH99, Wil95a]. Sysctl [DL93]. syslogd [Wal95d, SW91a]. syssumm [Moh99a, Moh99b]. System [AAO92, AT92, AISS97, All96, AHB^{+}98, AOG92, AK94, And94, AB96, AKS98, Ano92, Ano97g, Ano98q, Ano98-28, Ano98-29, Ano99j, Ano99k, Ano99u, Ano99v, AG97, BL90a, BB^{+}91, BDKM93, BL90b, Bar94b, BHK^{+}90, Bel96a, BGJY98, BM91a, Bel95, Ben93, BK91, BBS95, BGMR91, BP93, BZM91, BTS^{+}99, BVG^{+}97, BSSD96, Bla94, BL96, BCC^{+}91, Bra90, Bra91, BCC^{+}99, BGÖS98, BBG^{+}99, BCR91a, IG99, BCR91b, CLS92, Cal93, CBB^{+}91, CI93, CPW93, CWT^{+}97, Car98b, CN90, CBE93, CHHL99, CD94b, CK96a, CFH^{+}93, Chr90b, CPA93, CAK^{+}92, CR94, CJC^{+}92, CDM90, Coo95, CG99b, Cow99, Cuc90, CKDS90, Cur91, DWAP94, Dar93a, Dar93b, DSS96, DCM^{+}90, DAM^{+}91a, DAM^{+}91b, DL93, DVdB^{+}94, DMST95, EP92, EL92, EWCS96, Eps92, Eva95a, Eva97, FNP97, FS90, FG91, FC96, Fin96]. System [Fin90, Fis96b, FGB91, FHLW99, FHKR93, FSM^{+}91, GHN91, GKAS99, GS98, Gau91, GPRA98, Gil90, Gil96, GMR95, Gol90b, GLW94, GGV96, Gra96, Gra99d, Gre91, GC96, GB97, GA94, GWP99, GC98, GL93, Hal99d, Han99a, HK95a, HA92, HA93, HS93b, Har94, HG96, Har90, Har91, HMS94, HO92, HLM94, HL93, Hug96, HS99b, HW93, HMF^{+}99, HLR97, iIYL95, ISS97, JS93, JLI98, JKS^{+}93, Kap92a, Kap94, Kar93, KNM95, KLA^{+}90, KCDZ94, KN93a, KP97, Kin91, KSAL96, Klu90, Kni96, KM92, KSS93, Kol94j, Kolxx, KAD96a, KAD96b, Kot92, KL96, LaP94, Lam90, Lam91a, Lam97, Laz92, LeF92, Lea90b, Leo91, Lib90b, Lib91b, Lim97, LIN^{+}93, LMC94, LR98b, MSL94, MB91, Man94, Mar97, MO99, McD92, MKB90, McN93, MK92, MSN93, MK95, MK98, Min95, MMP94, Moh99a]. System [Moh99b, MD95, MLL95, Mor92, ML92, MP96, MLM94, Mul97, Mur99b, Nad92, NYM91, Nat96, Nee90, NKM93, Neu92b, Neu92a, NWRP^{+}91, NKM96, Nor91, Oet98a, OT90, Ols93, OS92, Owe97, PVL91, PT92, PT93a, PDZ99, PB96, Par95, Par92, PBS97, Pat93, PPC97, Pax98, PST^{+}92, Pea92, PKA98, Phi91, Pie96, Pik91a, Pom94g, Pom97a, PM95b, Pre92, PBR91, PS92, Pul99, PM98, Qui94, RS91, RG95, RBJ91, RHR^{+}94, RDM95, RWWB96a, RWWB96b, RL99, Ros92, Rot99a, RM94a, RC97, Roz92, Rus91, SB98a, SSZ95, Sat90, SM91, Sat93e, Sat94, SL94, Sca97a, Sch92b, SK96, SHK92, SLM^{+}90, Sel95a, Sel97e, SBMS93, SSB^{+}95, SO99, She93d, Shi91, SO92b, SO94, Shu91a, SBM90, SST90, Sim92c, Sim93a, Sim93b, Sim91c, SM98a, SST92, Sno96, SC99, SS94b]. System [SB97b, SSL^{+}99, SCB93, SV90b, SV91, Sto93b, Str90, SUK92, SH98b, SSCK93, SDH^{+}96, SUZ98, TSM95, Tal95, TM99, Tan96, TCI98, Tho98c, TJH^{+}94, TNR90, Tom96, Tra90, USE90f, USE93e, USE94m, USE97a, USE98e, USE99b, UKGS94, Urn97, VW99a, Vau90, VMMD98, WWWK95, Wat96, Web93, Wel91a, Wel91b, Wel91c, Wel94, WRW96a, WRW96b, WCS93, Woo98, YL98, Zie93, ZSD90, Zwi93e, Zwi93c, Zwi94e, Zwi94g, Zwi94f, Zwi94h, ZK94, Zwi97, dSdCF^{+}98, dSdS98, MK91, SD98a, SKS92, Bel95, Che94a, CG92b, DG94, ES90, E^{+}91, EKB^{+}92, Gil93, GHM^{+}90, HHW92, Kah95, Kub93, Neu92b, PGH^{+}91, Pea92, Ped90, RH98b, Rud97, SLS92, TP91, VS90, Wel92]. Systems [ACFM99, Alm99, AEK96, Ano97a, Ano98i, Ano98l, Ano98j, Ano98y, Ano98-27, Ano99r, Ano99b, Ano99m, Ano99l, Ano99n, Ano99-35, Ano99-31, Ano99-29, AOW96b, AVS98, BH99, BB95a, BCF^{+}93, BDM99, Ber92, Bla93, BHS95, BA91, BA92, BM96, Cae97, Cal93, CS92, CS94, CW92a, Cho99, CBB97b, CF95, CCGS92, Col92b, CG90, Var96, CJ92, Cra94a, CKDS90, Cur91, Cur97, DG96, DCWB92, DA92b, DV90, DOKT91, DB96b, EP93, Eva93b, EG97, FCNL94, Fle93, For99b, FPSW91, FSV94, FK98, FK99, GP94, GGI^{+}91, GPRA98, Gil90, Gol92, GM99, GG91, Hae91a, Ham97a, HM92, Hel92, Hic99, HCH95, HH99, Hum96a, HW93, IEE90, IC92, JPR96, KR99, Kev93, KC99b, KC96, Kur90, LB96, LH90, Lea94, LS99, Lyc96, Mem93, MW94, MU98, Mit95, MR95, MH92, Nes99]. Systems [Ous90a, Ous91a, Pag98, Par94, PE92, PS93, Per92, PS97, RA98, Rei92a, Rot95, Rud97, RMC90, SLS92, Sat95, Sch95b, Sch96g, Sco93, SU99, Sho97, Sma97, Sno96, Spexx, Spe91, STV90, STM^{+}99, Tan95, Tho90a, TJH^{+}94, USE90a, USE90b, USE90c, USE90d, USE91a, USE91b, USE91c, USE91d, USE91f, USE91j, USE92a, USE92b, USE92c, USE92d, USE92e, USE92h, USE92l, USE93a, USE93b, USE93c, USE93d, USE93f, USE93j, USE94a, USE94b, USE94c, USE94d, USE94h, USE94i, USE95a, USE95b, USE95c, USE95d, USE95h, USE96a, USE96b, USE96c, USE96d, USE96f, USE96g, USE96j, USE97c, USE97g, USE97h, USE98c, USE98g, USE99f, USE99i, USE99h, USE99c, USE99k, Vas91, Ver97, Wal95a, Wal95b, WAP99, Win96, Won90, YKB94, ZD93a, ZSG92, ZIL96, Zwi94k, dSM96, CG05, Ano98-33, Ano98n, Ano99s, Ano99t]. Systems [Ano99g, Ano99y, Ano99-30]. SysView [CJ92].

T [dL95]. Table [How95, Skl91].

Table-based [How95]. Tables [YCO94]. TAE [SW90]. Take [Rom90, Rom91a]. Taken [Ano95d]. Taking [Cha94]. Tale [Joh94d]. Tales [Kol94f]. Talk [ASB^{+}99, BTS^{+}99, Far95e, Pla97, Wal92a]. Tamed [Uma96b]. Taming [Bea97, CDGT98, SW95]. Tamper [AK96, KK99]. Tamper-Resistant [KK99]. Tampering [CV93]. TAMU [SSH93a]. Tangled [Sho97]. Tape [Cal93, SKS92]. Tape-based [SKS92]. Tapes [Wei91, Far95e]. Tapeworm [Kel99]. Taps [Kol96j]. Tar [TM97]. Task [DMY99, Gau91, MZDG93, Rhe98]. Tasks [Bacxx, BS97b, GSR93, Lib90b, Lib91b, Raw93, Sto93b, Win94a]. Tax [Rei92b]. Taxonomies [SP96]. Tcl [Ano97b, Ano98o, Ano99w, Ano99o, Ano99-28, Bai97, Ber98, FEBM97, Fly98b, McL95, Ros96i, Spe95a, USE95j, USE96e, USE97d, USE98f, FD98, Ama98, AB96, Bal98, Bar98, BG97, Bob99, Bog95, CLL97, CS97, CA96, Coo97a, CSL^{+}97, CKM95, DBK98, EN96, FEBM97, Fly98a, Gra96, Gri97, Gri98, HS95b, HK95b, HC97, Hir97, Hop96, Ing96, IJ97, IK95, Jae95, KCG98, Ken96, Ker95, KLB98, Kor96, LS97, Leh98, LO95, LDOW98, Lew96, Lib96b, Lib97a, Lib97b, Lin94, LSR95, LM97, Lon95, Lor95, Mah96, Mar96b, Mee97, MAH98, MK98, MLL95, ML92, NUO^{+}97, Ous90b, Ous91b, PBS97, PSJL97, Pet94, RL98, Ros96h, RC95, RC97, SPK^{+}96, SKCB97, SBB96, Sal97h, SSZ95, Sch95d, Sch95c, Sch96a, Sch96e, SD98b, She95, SL95, SSL96, Spe97c, Spr96b, TM97, Ull97]. Tcl [Wal97a, WU96, WL95]. Tcl-Based [Ros96h, Ama98, PBS97]. Tcl-BC [CA96]. Tcl-DP [LSR95, PSJL97]. Tcl-ish [Sal97h]. Tcl/2K [Ano99w]. Tcl/Tk [Ano97b, Ano98o, Ano99w, Ano99o, Ano99-28, Bai97, Fly98b, Ros96i, USE95j, USE96e, USE97d, USE98f, FD98, AB96, BG97, CLL97, CS97, Fly98a, Gri98, Hir97, Ing96, Jae95, KCG98, Ker95, KLB98, LB98, Lon95, Lor95, MLL95, NUO^{+}97, Pet94, RL98, SKCB97, SSZ95, SSL96, Spr96b, TM97, Wal97a, WU96]. Tcl/Tk-Based [CLL97, NUO^{+}97]. Tcl/Tk/ [FEBM97]. TclDG [EN96]. TclJava [SC96]. TclOSAScript [IJ97]. TclProp [IK95]. Tclsh [Fly99f, Fly99b, Fly99c, Fly99d, Fly99e].

TclSolver [Ken96]. TCP

[Chr98c, Jen98, Laz99, BB95a, CF94, Chu96, CRS99, CL94, CJ95, CG92b, Hof99, Jon95, MD92, Mon93, Mos96, NN94c, NN96b, Pax94, Sav99, Sca98, TP91, Ven92, Wil93, WVT94]. TCP-based [Sav99]. TCP/IP [Chr98c, Jen98, Laz99, BB95a, CF94, CJ95, CG92b, Hof99, Mos96, NN94c, NN96b, Pax94, Sca98, TP91, Wil93]. TCPDUMP [Cox98d]. TDBM [BN92]. Teaching [And97, Bor92, Pat97, Tom96]. Team [Kol98l]. TeamRooms [Ros96h]. Teams [Den99c]. Tech [Sel94]. Technical [Ano98v, Ano98-32, Ano99c, Ano99d, Ano99e, Aut95a, Eng93, USE92k, USE92o, USE95e, USE97e, USE99m, Uma95, Win96, USE96k, USE98d, USE98h]. Technique [CI93]. Techniques [Bed90, BMK96, Fly97, GMNR99, Mar98, MO99]. Technologies [Ano98n, Ano98j, Ano98-33, Ano98-27, Ano99s, Ano99t, Ano99g, Ano99-29, Hua98, Ins96e, Isl95, PS97, Sch97a, SS96a, Spexx, USE95k, USE96g, USE97g, USE97h, USE98c, USE99f, USE99c]. Technologist [Joh97b]. Technology [BCF^{+}93, Ben94a, CSYG94, Cro98, Han92, Kol91, Kol94b, Kol94g, Lar94, NFQ99, TMJY91, USE99j, WD95, Ano99z, Lea95a]. Teicher [Kol97b]. Telecom [KCG98]. Telecommunication [CJJ^{+}98, Sim92b]. Telecommunications [NWEF92]. Telephone [Sco97, Uhl93]. Telephony [AHL^{+}91]. TeleRead [Ano93c]. Telethon [Gri98]. Television [Kan91]. Tell [Wil94]. TELNET [SSH93b, VTB95]. telnetd [RW97]. Template [Law94, Sun97]. Templates [GG98, SD98b]. Temporal [FSM^{+}91, IS91]. Ten [Var95b, Spe93b]. Tennessee [USE91g]. Tenth [Var96, USE96j]. Tenwen [Eva94c]. Terabyte [FM99]. Terabytes [Pre98]. Term [BA91]. Terminal [Bia96a, lWS99]. Terminals [Ped90]. Termination [RL99]. Tertiary [KSS93]. Test [Ama98, FD98, Pea93]. Testbed [CF94, Kap92b, RAV95, RBB^{+}91]. Testing [Bir95, GMS96, Lib92, MS96a, RRB99, SB99, VK98, Zwi91b, Bir95]. Texas [USE92g, USE98i, USE99a, KW99, SSH96]. Text [Mof92, Qui94, Spi98, VcC97, PT93b]. Texts [She92a]. Thanks [Ano99-27, Ano99-54, Kol95j, Kol97h, Mil98g, You98d, You99e]. Them [Arn91]. Themselves [EL92, Saf92]. There [Bel92, CD99, Den98, Eps92, Gru96, Kel93, Kel92a, Kel92b, Pet93, Zwi94i, Jon98]. They’re [Ede92, Ham97b]. Thin [ZPMD97]. Thin-Client [ZPMD97]. Things [Jac94, Sub98, Wil94, Zwi91b]. Thinking [Tau91]. Thinnet [dL95]. Third [Ano98-31, Joh97a, Kam93, Mil99i, Pom94b, PS97, SB97b, USE94m, USE97g, USE99h]. Third-Party [Kam93, Pom94b]. Thirteenth [USE99i]. Those [Kol95h, Lib90a, Ros95c]. Thoughts [Bac95, ML91a, Mue95d]. Thousand [Wil97b]. Thousands [SP99a]. Thread [ASSB99, BC97, FL94, Heu90, ISS97, Lie92, ZY98, PKB^{+}91]. Threaded [Jon91, STW93b]. Threads [Alf94, DAM^{+}91b, Dea93, Duc91, GL93, Hai97, Lit92, Mue93, Ous96, SS92, ZSG92]. Threat [LKCT96, NPW93, Win96]. Threats [SS99]. Three [TB99, Wel94]. Threshold [FY98, Bar92a]. Throughput [FP93, KP93a]. Thrown [Dam95]. Thwarting [DKM94]. Tickerplants [BGH^{+}90]. Ticket [FN98, FKT^{+}99]. Tickets [Wil99]. Tickle [Sch95f]. Tightly [FBAE^{+}92]. Tightly-Coupled [FBAE^{+}92]. Tilde [CDM90]. Tiled [Fra92]. Time [AL93, AB91, AOG92, Arn95a, BL90a, BDKM93, BSSD96, CRVZ97, CL97, cC99, Cla97, DZL97, Det92, FBAE^{+}92, FC96, GG98, GS94c, HS94, Har92, Heu90, HL93, ISS97, JR99, KTN93, Kol94a, KH97b, Kre97, Lea92a, Men98, Mil99e, NKT93, NL96, Nil94, Ove91, Pax98, POS^{+}99, Rey92, Roo92, Rub95, Rub96b, ST93, SB93, SL92, Swa97a, Swa97b, TN96, TNR90, TN91, Tra90, WGC^{+}95, ZSG92, BZT99, CJ99, CV96, FPSW91, hLhCN98, SMH95, MAM95, PTB^{+}97, RBJ91]. Time-Oriented [Roo92]. Timeliness [Yah94]. Timers [ST93]. Timing [Bla90]. Tiny [Cam93]. Tip [Ros95c]. TITAN [FPA98]. Tix [Lam95]. Tk [Ano97b, Ano98o, Ano99w, Ano99o, Ano99-28, Bai97, FEBM97, Fly98b, Ros96i, Spe95a, USE95j, USE96e, USE97d, USE98f, FD98, AB96, Bea94, BG97, CS97, Cro95, CKM95, Esp97, Fly98a, Gri98, Hir97, Ing96, IK95, Jae95, JS95, KCG98, Ker95, KLB98, Lev96, Lon95, Lor95, MLL95, ML92, Pet94, RL98, SPK^{+}96, SKCB97, SSZ95, Sch97b, SSL96, Spr96b, TM97, Uhl95, Wal97a, WU96]. Tk-Based [CLL97, LB98, NUO^{+}97]. TkMan [Phe95]. TkPerl [Bea94]. TkReplay [Cro95]. Tksh [Kor96]. tmp [OT90]. To-Do [Rue96]. Toaster [Swa96]. Toba [PTB^{+}97]. Token [CST95, GD96a, Nys99]. Token-Mediated [GD96a]. Tokens [KS99]. Told [Nem99b, Pom94g, Sim93b, Zwi93c, Zwi94g, Zwi94f, Zwi94h]. Tolerance [CL99, HCK^{+}98, KMBT92, KSAL96, Maf95, MCU97, NMMS97a, SL96]. Tolerant [Bab90, BMHR91, CN90, WMB99]. Tomatoes [ASB^{+}99]. Tool [Ana99, Bea96, Cal96, Che94b, CAY99, Chr97b, CW92b, Eir94, Fis96b, Fle92a, Fle92b, GS94c, HJ92, IP94, JS93, KCG98, KS94b, KH97b, LZ97, LCH92, MW94, Mon93, Muf99, PN99, Pie96, RLS^{+}97, Rhe98, Ric97, Sat94, Sav99, SV94, SA95b, Sha92b, Sma97, Ste93a, SW90, TL96, VC92, WUB^{+}90, WM92, YGM95, ZSD90, Sta98, Mel99b]. Toolbox [Koh99]. Toolkit [BK97, CS97, Duc97, GSPW98, HO94, HK95b, Hop96, KP93b, LB98, LO95, ML91b, NUO^{+}97, NN93a, Ous91b, PK93, Ped90, Phi91, Quo91, RF95, RA94, Bea94, Spe95a]. ToolMan [Sin97d, Sin97a, Sin97b, Sin97c, Sin98d, Sin98a, Sin98b, Sin98c, Sin99a, Sin99b]. Tools [BAKW92, Che97, ES95, Eva93b, Fly98a, Har95, Kol93b, Kre94, LT97, LO90, Mah91, May95, MS96b, ML92, Mot95, MM92, Nem99b, Pom94g, Saf92, Sim93b, Sin98d, Sma90, Sma91, Spi94, SG97a, Spi98, Tro96, Uhl91, WU96, Zwi93e, Zwi93c, Zwi94g, Zwi94f, Zwi94h, Bea98, Fly98b]. Top [May95, MGZ93, Var95b, CNTS93, MZDG93, LeF94]. Topics [Ano98-37, Dar93b, DSS96, Kle99c, Sch95e]. Topics/Buzzword [Kle99c]. Topology [Pul98, kc99a]. TOPSY [SWS92, WSS92]. Torn [Ano98-41, MR98]. Tornado [GKAS99]. Toronto [USE94g, USE95j, USE96g]. Torture [Zwi91b]. Torture-testing [Zwi91b]. Tournament [Kol98m]. Toy [Zwi94h]. TPCs [Gun98a]. Trace [GB97, KTP^{+}96, KPR94, PV95, SKCB97, Sah95, SO92b]. Trace-Driven [KTP^{+}96, SO92b]. Tracefiles [Cou96b]. Traceroute [PN99]. Traces [PS96, SKCB97]. Tracing [Bla92, Cur94, Mad92]. Track [Pom93d, SL93, USE98d]. Tracking [Bru99b, DB96a, Lei99, Rhe98, RM94a, Wal95d, Zwi94k]. Trader [BPT98]. Trades [Ber96]. Trading [Lip95]. Traditional [BMK96, FST91]. Traffic [Cho98, Hei90, Mar99, Oet98b, OW94, Rei97, SH98b, Tay96d]. Training [Sha92b]. Transaction [BZM91, BS93b, BBK99, BCC^{+}97, CCK^{+}91, CTS95, FL93, MB91, MKC92, Spexx, YTJ91]. Transaction-based [BBK99]. Transactional [CLS92, SB98a]. Transactions [BCE98, BN92, CHTY96, JABA98, Lee98, LS98b, Pov99, SO92a, SEYG96]. Transcoding [CE99]. Transfer [ACO92, Bar94c, NN94a, Sel95a]. Transformation [CCK98, SB97a]. Transformations [MP99]. Transformer [SB98b]. Transition [Rei98b]. Translation [BKW94, KW99, KA94, LP94]. Translator [FLB98, Ker91, YHK96]. Transparent [CCK^{+}95, GLP^{+}93, Gol90a, KMBT92, PBKL95, SL96, SSBK98, TSE90, VA98, YHK96]. Transport [Che97, CHW98, MBM95, Sal92g, SLC94, VDAA99]. Transportable [Bvv91, SW90]. Transputer [ACO92]. Trapeze [ACG^{+}98]. Trapeze/Myrinet [ACG^{+}98]. Traps [Ven92]. trash [MH92]. Travails [CD92]. Tread [KCDZ94]. Treasure [Fly99a]. Treasures [Lea96b]. Treatment [Oka98a, Oka98b]. Tree [JSS90, Sat93b, Skl91]. Tree-Based [Skl91]. Trees [Cre97, HF96, KS97, Koe92]. Trenches [CF98, Ann98]. Trends [Lev98]. Trials [Har90, Har91]. Tribeca [SH98b]. Tribulations [Har90, Har91]. Tribute [Far99g]. Trick [Tay97a]. Tricks [DMY99, Lim99d]. Trip [Dar98e, Dyk91, Dyk92]. Triple [NPW93]. Triples [Pol93]. Tripwire [KS94a, KS94b]. Triveni [CJJ^{+}98]. Trojan [TW96]. TRON [BBS95]. Trouble [Cal99, GHN90, GHN91]. Trouble-MH [GHN90, GHN91]. Troubleshooting [Hof99]. True [Gra93, Ous91a]. Truly [Cop94]. Trust [AMP98, BS97a, Fin95b, Gil90, LA98, MR98, SA95a, YKB94]. Trust-Based [YKB94]. Trusted [CW92a, Cho99, ES90, E^{+}91, LaP94, Seb91, SA98]. Trusting [Cra94b, SA98]. Trustworthiness [PV95]. Truth [For99a, Mil99j]. Tuning [CHS91a, Hum98a, Kol97i, Mac91, VR99, Sim93a]. Tunnel [CGHK95, RFW93]. Tunnels [SB98b]. Tuple [HF96]. Turkey [Pie99a]. Turnin [Cat90]. Turning [Lim97, Tal95]. Tutorial [Kle97, Kle99a, Lea96a]. Tutorials [Ano98w]. Twelfth [Ano99-35, USE98g]. Twelve [AP99]. Twenty [Sal97i, Sal98h, Sal98i, Sal98j, Sal98k, Sal98l, Sal98m]. Twist [Tom96]. Twisting [McR95]. Two [Ano95c, BS93b, DOKT91, Joh94d, LR94, LO90, Mad92, Nat92, Phe95, Spe97b, ZIL96].

TX [USE91i, USE91l]. TxRx [Mat97]. Tycho [HLR97, RL98]. Type [BSS^{+}95, BSSW96, Bli90, DT97, HSOS99, IL90, IL93, JS93, Ken91, Koe94, Lea92a, LUM91, LMU93, Mur93, PR94, SL92, Vo97, WSB^{+}96]. Typecast [Zwi92d]. Typed [Mur92, Sha97]. typesetting [Ker82]. Typing [Det92, RC97].

U [KGMP^{+}96, Sal99b]. U-PAI [KGMP^{+}96]. UAMS [Sto93b]. Ubiquitous [KR99, Wei95b]. UC [Ano94a]. Ucode [SCD90]. UDB [SV90b, SV91]. UDP [KP93a]. UDP/IP [KP93a]. Ufo [AISS97]. UFS [LPLF91]. Ugly [Cox98a, Kor97b]. UI [Wat96]. UIDs [Cal98]. UIMS [HO90a, Kle94]. Ultrix [Gee90]. Unattended [Soq97]. Uncompressed [Zea97]. Uncontrollable [Lib90a]. Underlying [dSdCF^{+}98]. Understand [Cha99, Pom96c]. Understanding [Mue94c, Seo93b]. Undesirable [Kin90]. Unified [MU98, PDZ99]. Uniform [Mah91, NR94]. Uniformity [Gré93, Grg93]. Unifying [Sum94]. Union [PM95a]. UNIQNAME [Dos91, DLM90]. Unique [Cal98]. Uniserver [Gee90]. United [Joh93d]. Units [KA94]. Universal [CCGS92, KGMP^{+}96]. Universe [Ros95a]. University [Fin99b, BKS90, Bra99, RFR99, SSH96]. UNIX [Bea98, Col92a, Cuc90, Gil93, Gwy90, Lea90b, Lea92b, NN92a, Sal97i, Sim93a, ALO90, AS93, ATS94, And91, Ano97d, Arn91, Arn98, BSS^{+}95, BS92a, BDD^{+}93, BGH^{+}90, BBS95, BGMR91, BAKW92, Bor98, BCC^{+}91, BBP99, BCGS91, Cal93, CBB^{+}91, Car93b, CCM^{+}90, CW92a, CWS90, Cho98, Chr90b, CJ92, Cro98, CS90, CKDS90, Cur91, Cur97, DF93, DeV93, Eng91, Eps92, ESSB93, Eva97, Far92, Far95d, FG91, FM94, Fra92, Fre91, FST91, GB92, Gar90b, GDFR90, GES98, Gra98a, Gra98b, Gra98e, Gra98d, Gra98c, Gra99c, Gra99d, Gra99e, Gra99b, GLOR91, Hae90a, Hae90b, Hae90c, Hae90d, Hae90e, Hae91b, Hae91c, Hae92, HG93, HK95a, Hau94, Hel92, HT92, HT94, HHLS90, HAB^{+}98, IB93, IHR99, Kap92a, KI99, Kem93, KN93a, Kol93b, Kol94e, Kor97c, Kor97d, Kor97e, Kor97a, Kur90, LB96, LKY93]. UNIX [Lei98, LT91, LS92, LPLF91, LBSL97, MWCR90, MWCR91, MS92, McK93a, MK91, Mil91, MK93, Mue93, Mue94c, MZA96, Nad92, Nee90, NR94, Nor91, O’D90b, PBS98, PK93, Pin92, PBKL95, Pol93, Pov99, Rei97, Ros95c, Roy93, RW93, Sal99a, Sat90, Sat93e, Sch95b, Sel94, SY91, SO92a, SBMS93, SSB92, She93d, SW91a, SST90, Sim91c, Sin97d, Skl91, SKP94, Spi94, SMD^{+}94, SJ95, Sto93e, Sto93c, Sto93d, Sto93f, Sto94b, Sto94a, Sto94c, Str90, SW91b, TLK99, TP91, TSO^{+}96, USE90i, USE92m, USE93m, USE94g, USE95f, Uma96b, VB92, Wal91c, Wal91d, Wal91e, Wal92h, Wal92i, Wal92d, Wal92e, Wal92f, Wal92g, Wal93c, Wal93d, Wal94b, Wal95c, Wal97b, Wal97a, WCM93, Won90, YLK93, YR93, ZRB^{+}93, Zie93, Zwi92b, d’A92a, dC93, AHB^{+}98, Gam93, GMR95, Har94, Joh98, Mil98h]. Unix [NC94, Qui94, Tho90a]. UNIX-like [HK95a]. UNIX-related [Sto94a, Hae90e, Hae91b, Hae91c, Sto93e, Sto93c, Sto93d, Sto93f, Sto94b, Sto94c, Wal91c, Wal91d, Wal91e, Wal92h, Wal92i, Wal92d, Wal92e, Wal92f, Wal92g, Wal93c, Wal93d]. UNIX/Windows [Bor98]. UNIX(R) [Noe98, Wal98]. Unpredictable [BS97b]. Unsecure [LT91]. Untangling [Sho97]. Until [Mel96b]. Untrusted [GWTB96, SK98]. Untruths [For99a]. Upcoming [Sin98d]. Update [Ber99d, Car93b, Dij99b, Eir94, GP94, Hae90a, Hae90b, Hae90c, Hae90d, Hae90e, Hae91b, Hae91c, Lim99a, Mog94a, NN98, Ros95b, Ros96e, Ros97, Sto93e, Sto93c, Sto93d, Sto93f, Sto94d, Sto94e, Sto94f, Sto94b, Sto94a, Sto94c, Sto95c, Sto95d, Sto95e, Sto95a, Sto96f, Sto96g, Sto96c, Sto96d, Sto97a, Sto97b, Sto97c, Sto97d, Sto97e, Sto97f, Sto98b, Sto98c, Sto98d, Sto98e, Wal91c, Wal91d, Wal91e, Wal92h, Wal92i, Wal92d, Wal92e, Wal92f, Wal92g, Wal93c, Wal93d, HG98, Sto99a]. Updated [LC99]. Updater [FC96]. Updates [Car98b]. Upgrade [SMH95]. Upgrades [Hid94, LL98, Rid94]. Upgrading [Lar94, Lim97, MI93]. UPS [Mel96b]. URL [CRS99]. URL-Aware [CRS99]. USA [Ano97o, Kol93c, Kol98k, Kol98l, Pie95, Pie99b, USE90f, USE90g, USE90h, USE91f, USE91g, USE91h, USE91i, USE92e, USE92g, USE92h, USE92i, USE93f, USE93g, USE93h, USE93i, USE93k, USE93j, USE93l, USE93m, USE94h, USE94i, USE94j, USE94f, USE94k, USE94l, USE95f, USE95g, USE95h, USE95i, USE95k, USE95e, USE96j, USE96k, USE97a, USE97c, USE97e, USE99a, USE99e, USE99f, USE99b, USE99i, USE99c, USE99j, USE99d, USE99k, USE99m]. Usability [WT99, CG05]. USACO [Ano96e, Ano98-43, Ano98-42, Kol96j, Kol97j, Kol98a, Kol98m, Kol99h]. Usage [Bla90, Cur94, Fin94b, HS93a, Ric97, Sin97c, SS94b]. Use [And92, Bea96, CK96b, Fin97c, Kah95, Mot95, MR97b, Too98, CG05]. Useful [CGMV91, MRK97, Rom91b]. USELINUX [mH96, Ano97n]. Usenet [Chr97a, Laz97, Ada93, Bis91, LOM94, Sha92a, Uma96a, Ins97a, MRK97, Sal92g, Swa93, Swa96, Zwi92a, Chr98f]. USENIX [Ann97, Ano97c, Ano98a, Ano98p, Ano98v, Ano98j, Ano98k, Ano98m, Ano98g, Ano98-30, Ano98-31, Ano98-27, Ano98-32, Ano98z, Ano99w, Ano99c, Ano99d, Ano99e, Ano99x, Ano99g, Ano99p, Ano99q, Ano99o, Ano99l, Ano99n, Ano99-29, Ano99-32, Far95d, USE91i, You97b, USE99m, Ano96b, Ano97j, Ano97p, Ano98-33, Ano98-44, Ano99-51, Ano99-55, Ano99-58, Ano99-56, Ano99-57, Ano99-59, Aut95a, Cow99, Den98, Den99a, Den99b, Den99c, Fis96a, Fis97, Gol95, Hon95a, Hon95b, Hum97b, Isl95, Kol90, LH90, Mau98, Nem99a, Ros96a, Ros96f, Ros96g, Ros97, Ros98a, Ros98b, Rub96a, Rub98, Sal98a, Sal98m, Sal99c, Sal99d, Sal99e, Sal92f, SS94a, Sim97, Sto94d, Sto94e, Sto94f, Sto95c, Sto95d, Sto95e, Sto95a, Sto96f, Sto96g, Sto96c, Sto96d, Sto97a, Sto97b, Sto97c, Sto97d, Sto97e, Sto97f]. USENIX [Sto98b, Sto98c, Sto98d, Sto98e, Vel95, Wal94a, WUS^{+}92, You94d, You95b, You95c, You98c, Ano97n]. Usenix/Uselinux [Ano97n]. User [ASSB99, AISS97, Alm99, AKM99, Arn95a, Bar94c, BHK^{+}90, BSKH93, Ben93, BGMR91, CDK^{+}99, CV96, CJ92, Cra94b, Dea93, EP93, ESSB93, FK92, FHKR93, FSV94, Fur91, GF95, Gol90b, GSR93, Har95, HO94, Hun97, HLR97, ISS97, Kol96k, KCCT94, Lim99b, Lon95, MPS98, Mas91, MJ93, MA90, Mil93, Mil96b, Mue95c, MR97b, NB97, NN94b, PS93, Ped90, Pik91b, Pik94, Raw93, RN93, ST93, Sch91, Sch92a, SP91a, SV90a, Sho97, Sim94a, SSL96, SKS90, SV90b, SV91, SW90, Sze99, Uhl95, Uma96a, Urb90, Urb91, VZ91, WR98, Wel95, EL92]. User-Extensible [CV96, SSL96]. User-friendly [CJ92]. User-Level [ASSB99, AISS97, CDK^{+}99, Dea93, FK92, FHKR93, GSR93, ISS97, MPS98, MA90, PS93, Raw93, SP91a, SKS90, MJ93]. User-Mode [Alm99, Hun97]. User-Network [NN94b]. user-setup [EL92]. User-Specific [WR98]. User-Visible [Sho97]. User/Product [Kol96k]. Users [And97, Arn98, EL92, Gil90, HO94, Kam98a, KP97, Lea96a, MC99, Pat97, SP99a, Sel99c, VB92, Zwi92b, dLRT93]. Uses [Sah95]. Using [BNMK99, BS92a, Bar98, Bel95, BCE98, BG97, BMK96, Box96, BPR92, BCLC99, CL91a, Car99b, CD94a, CWS90, CBB97b, CGCS91, CSM97, DWAP94, Dar98d, DBK98, Dea93, Dou93, EWCS96, Far96i, Far96g, Far96h, Far97i, Far97j, Far97k, Far98g, FD98, For99b, FY98, GG97, GS99a, GBB93, GKS90, Gsc95a, HSOS99, Har97a, Hed98, VW99b, HC97, Hir97, HHLS90, HH99, HL93, Hor92, Hug96, iIYL95, KR99, KCG98, KS99, KS94a, KOO95, KSS93, Kra98b, LPS99, Leh98, LC99, Lib90b, Lib91b, Lin94, Lon95, Lu97, LW99, Lut99, Mat98, McC97a, McC97b, McC97c, McC97d, McC98e, McC98f, MAM95, Mee97, MAH98, MK98, MPZD98, MLL95, NSM99, NAU93, NLZ99, Par99a, PS96, Pol94, Pre98, PBR91, PA90, Rao98d, Rao98a, Rao98b, Rao98c, Rao99a, Rao99b, Rao99c, Rao99e, Rao99d, Rob92, RVL^{+}97]. Using [ShCR93, SSZ95, SH96, Sch95d, Sch97b, SeaBP99, Sie98b, SB99, Ska94, Sla97, SKP94, Spi99, TAP90, Tay97e, TP91, Tro96, TM97, WGC^{+}95, Wie92, YCE^{+}97, BMHR91, Ber98, CS97, FEBM97, GA94, Han92, MP90, SD98b, SUZ98]. USITS [Ano98j, USE99c]. USIX [BSSW96, Mue96b]. USL [Ano94a]. Utah [USE95f]. Utilities [HSOS99]. Utility [Gre91, Pre97]. Utilization [Kra98a, MT93]. Utilizing [Fre97, SKS92]. Utopia [Tho96]. UTS [Rua90]. UUCP [Bar94c]. UWIN [Kor97d, Kor97e]. UX [SMPR98].

V [CBB^{+}91, DG94, EKB^{+}92, Pea92, SLS92, VS90, CWS90, Dyk91, FG91, FHG93, PST^{+}92, She93d, Str90]. v3 [Rhe98]. V5 [IH99, Ts’97]. Vacation [Dew97]. Valentine [Roc97]. Validation [KW97]. Value [Ous91a, Rei92b, Mor92]. Vanderbilt [Bra99]. Vanguard [FHG93]. Variable [TD92]. Variable-Consistency [TD92]. VarietyCash [BGJY98]. Various [NN93a]. Vassal [CJ98]. Vault [AUH99]. Vaulted [Cho99]. Vehicle [Spi97]. Vendor [Nem99b, Pom94g, Sim93b, WN92, Zwi93c, Zwi94g, Zwi94f, Zwi94h]. Vendor-Neutral [WN92]. Verification [Nee90]. Verifying [LHB96]. Verilog [JB99]. Verischemelog [JB99]. Version [CCK^{+}95, Dar97f, Jef94, LFB92, NN94b, PJS^{+}94, Wat96]. Versus [Cha99, SSB^{+}95]. Very [BA91, BFRS99, CHK^{+}98, GH93, Hae95b, MFM95, O’D90b, Pol94, USE94k]. Vests [Cra96]. VHLL [USE94k]. VI [SD92, USE92e]. Via [Cur94, DS92, Sug92, BGÖS98, Car94, Col96, PK96, Roc97, Wal97b, Wal98]. Video [AW99a, AOG92, CLL97, LWS^{+}94, Mil91, Ove91, RBB^{+}91, SS95, SW91b, TMC97, Zea97]. Videoconferencing [HAC^{+}98b]. View [All96, All99, AKM99, Bia96a, Bla93, Dav94a, Dun90d, Lyc96, Mar97]. Viewer [SPK^{+}96]. Viewing [DB96a]. Views [Kol92, Mor92, Nat92, NZ93b, Nat93, PEZ93, Sat93e, Sma92b, Zie93, Zwi92c, Zwi93b, Zwi93c, Zwi94c, Zwi94d]. ViewStation [LWS^{+}94]. VII [USE93f]. VIII [Aut95b, Old94a, USE94h]. Viral [WCO^{+}90]. Virtual [All95, Ano97m, Bal92, Bel90, BCF^{+}91, BCCR91, CRRS93, Cho99, CG91, CG92a, CV96, DBD^{+}99, FHL^{+}96, Fra92, Gre99, Kel95, LCC94, LV99, MLM^{+}98, McK95, Mey91, Min93b, NB94, Neu92b, PE92, PBF^{+}96, SH98a, Tay97d, Uhl93, VMMD98, VR99, WAP99, ZIL96, BCLC99]. Virtually [CMS90]. Virtual(R) [New95]. Visible [Sho97]. Vision [Sto96a]. Visual [GB98, LM94, Hop96]. Visualization [Bos96, DLVW98, Hug96, Ker95]. Visualized [Sin97c]. Visualizer [LN95]. Visualizing [BMH98, Cou96b, Tra90]. VLAN [Bra99]. VLAN/ELAN [Bra99]. VM [Bot98, CBL90]. Vmalloc [Vo96]. VMS [Wie92]. Vnode [SKS90, LBLM90, Ros90, Sto92, Wel92, Wel94]. Vnodes [Joh95a, SW93]. VNS [LFB92]. Voice [AW99a, Ben94c]. Voices [Pin92]. Vol.I [NN96b]. Vol.II [NN96b]. Vols [Wal92a]. Volume [Bos96, MFM95, Mos96, PGH^{+}91]. Volumes [Ano93a, Koo93]. Volunteer [Bia96a, Fis96a]. Volunteers [You98d, You99e]. Voodoo [Lar99]. VPN [Cho99]. VPNs [Gre99]. vs [HMF^{+}99]. VT [USE90e]. Vulnerabilities [AM98b, BS90]. Vulnerability [Rau99, ZL96].

W [Kol98d]. WA [USE92i, USE99i]. Wafe [NN93a]. Wait [Mel96b]. Waldo [Kol96f]. Wall [HRT96, Kap94]. Wallet [DBGM^{+}98]. walls [CHK^{+}98]. WAN [Muf95, YADL99]. WAN-hacking [Muf95]. Want [Kol94j]. Wanted [Hil97, Wal95c]. wants [LAG^{+}96]. Warehouse [Pol97]. WARLOCK [Ste93a]. Warp [RBJ91]. Washington [USE90h, USE91k, USE94m, USE96f, USE97a, USE97f, USE98e, USE98a, USE99e, USE99b, USE99d]. WAT [PTB^{+}97]. Watching [Ins96e]. Watson [CG99b, MR91]. Way [GGI^{+}91, Mil99j, Mue96a, PTB^{+}97]. Weak [Gol92]. Weak-Consistency [Gol92]. Weather [BTS^{+}99]. Web [Far97a, MR98, SA95a, Sie97a, Sin99a, AMP98, AM98b, BNMK99, BD97, BMLH97, BMH98, CD97, CS97, CFHB99, Dav95, Don99, DB96a, DFKM97, DMF97, FST99, FL96, GS96b, Hir97, Hou96, HPS97, IC97, Kle99b, Kra98b, KW97, KL97, LB98, LeF99, Lea95b, LBBA99, LC99, LM97, Man97, MS97, MKPF99, PP99, Pom95c, Roc97, SP99a, SD98b, Sin98b, SAYZ99, TL96, Tay95a, Tay95b, Tay96c, Tay96a, Tay96e, Tay97b, Tay97a, Tay97c, Tay97e, Tay97d, Tay99, VF95, WS99, WU97, WVS^{+}99, YL98, YL99, ZSBC98, ZBCS99]. Web-Based [BMLH97, WS99, Kle99b]. Web-Object [WVS^{+}99]. Web-to-Database [CS97]. WebGlimpse [MSG97]. Weblint [Bow98]. Webmaster [Pee96, Tay96b, Tay96e, Koh99, Tay96c, Tay96d, Tay98a, Tay98b, Tay99]. Webpages [Leh98]. Website [Kle98a, Kle98b]. WebWiseTclTk [LB98]. Weeks [Mel96a]. weighted [Cra96]. Weightless [SPY^{+}93]. Weiser [Ano99-36]. Welcome [Car99a]. Well [BCF^{+}93, Mur93, Pov99]. Well-Behaved [Mur93]. Well-Formed [Pov99]. WG15 [Dun90c, Dun90b]. Whatever [Cam93]. Where [Dar99c, Ins96f, McN93, MP93]. Wherein [Kol95e]. Wherever [Zwi94i]. Which [Che92, Gir99a, Min93a, Min93b]. White [BD93, Fin96]. White-pages [BD93]. Whither [Hum97b, Mur99b, Nat92, Sma92b, Ano97j]. Who [Cra94b, McN93, Poo94, Zwi94j]. Whois [All95]. Whole [Bar92c, Duc94]. Whom [Cra94b]. Wide [AMP98, BVAD98, CS97, Dav95, DFKM97, DMF97, Gru96, GS96b, IP94, KW97, LB98, MET98a, Pea99, PP99, SS94b, TL96, VDAA99, VF95, YGM95]. Wide-Area [IP94, SS94b, VDAA99, YGM95]. Widely [TJM^{+}99]. Widget [BH95, BAKW92, Esp97, Lyc94, Ulf95, Uhl96]. Widgets [Lam95, McC98g, McL95, Ulf95, Jae95]. Wild [Bea97]. Wiles [Pom96d, Pom96e, Pom97b]. Will [ASB^{+}99, NPW93]. Wilson [Pol93]. Wily [Far94, Kle99b]. Win32 [HB99, Kor97b, Noe98, PBS98, Puc99, RVL^{+}97, Wal97a]. Win32/Intel [RVL^{+}97]. WinACIF [KCG98]. Wind [RFW93]. Window [Che94a, CG92b, ES90, E^{+}91, Kah95, LaS90, Mur99b, PE92, Ped90, Pik91a, TP91]. Windowing [PE92]. Windows [Ano97p, Ano98q, Ano98k, Ano98w, Ano98-29, Ano98z, Ano99j, Ano99k, Ano99u, Ano99l, Ano99n, Bir93, Haa90, SCW^{+}99, Sca97a, Sho97, Sie98a, USE97a, USE97f, USE98e, USE98a, USE99d, ASSB99, Ahl99, Alm99, Ano97g, AC97, Bor98, BGT98, BCC^{+}99, Bru97, Bru98, BC97, BCLC99, CRVZ97, Car99b, CJ99, Cof97, CRH99, Cox97d, Cum99a, Das97, DMZ98, ES97, Far97h, GMR93, GSS99b, HPS97, HCK^{+}98, IH98, JR99, Kor97c, Kor97d, Kor97e, Kor97a, LeB99, Lea96a, Lie99, hLhCN98, LBSL97, Lyc96, MD98, NLZ99, O’B99a, PS96, Pla97, Pla99, Rat98, RvIG98, Rob90, RW97, Rus99a, SD98a, Sjö99, Spi94, TCI98, USE99b, Uhl90, Wal97b, Wal97a, Wal98, lWS99, WWDF98, ZY98, Zea97]. Windows(R) [JP99]. Winners [Ano96c]. Wins [Ano97o, Kol99h, Pie95, Rou97b]. Winter [Ano94f, Num94b, USE90d, USE90h, USE91d, USE91i, USE91l, USE92d, USE92j, USE93d, USE93l, USE94d, USE94l, USE95d, USE96d, WSR93]. WIP [Blo92]. Wireless [IGP95, KDF^{+}93, TC95]. Wires [OH95]. Wiring [Sim92b]. Wisconsin [RFW93]. Wisdom [Dar99d, Dar99e]. Within [BCGS91, EBG99]. Without [CCM^{+}90, Car93d, Cha92b, Ell96, Hal99c, HN97, Lie92, NL96, Sim90, Sim91a, Str99, VB92, Wei91, Zwi93b]. Wizards [Dar99b]. WMCSA [Ano98i]. Wolfpack [BVG^{+}97]. Women [Jon94]. Woodstock [Kol94d]. Words [Dar99e, Kol95f, Sal93f, Sal93f]. Work [Cam94, GHN90, GHN91, Zwi94k]. Work-Queue [GHN90, GHN91]. Workbench [Ken90]. Workbook [Cha96]. Workbook-Review [Cha96]. workfactor [Jas96]. Workflow [PPC97]. Working [Dun91b, Dun91c, Kol96l, Kor97b, Var92a, Var92b]. Workload [Win94a, kc99a]. Works [Den98, WUS^{+}92]. Works-in-Progress [WUS^{+}92]. Workshop [Ann97, Ano97g, Ano97p, Ano98a, Ano98i, Ano98m, Ano98g, Ano98-31, Ano99x, Ano99y, Ano99z, Ano99-30, Ano99-34, Hon95a, Hon95b, IEE90, Lan92, PT93a, Per92, Sch95e, USE90e, USE92f, USE92h, USE92i, USE95g, USE95j, USE96e, USE96h, USE97a, USE97d, USE97f, USE98b, USE99j, USE99k, USE99l, USE90i, Bai97, DSS96, Ros96i, Sat95, Sca97a]. Workstation [And92, AHL^{+}91, BC97, HS94, Kan91, KGL^{+}91, KH97b, Kra98a, LPV^{+}91, MLM^{+}98, Mue94c, NLZ99, Ove91, RP92, Sch92c, SSCK93, vR95, SW91b, BR98]. Workstation-based [Kan91]. Workstations [BL97, Fin94b, FLA94, Har92, HJK^{+}92, KCDZ94, MI93, SL96, SMH95]. World [And97, Ano97d, AHS95, Chr98b, DW98b, Dav95, EBG99, Gee99, Kol94c, Mel99a, NPW93, Pac93, Pat97, Pea99, Ull97, VF95, PT93b, AMP98, CS97, DFKM97, DMF97, GS96b, KW97, LB98, McD92, PP99, TL96]. World-Wide [Pea99]. World-Wide-Web [DMF97]. Worlds [Sca97b, VMMD98]. Worth [Dar98f, Gra99e, Sto96e]. Worthwhile [Den98]. Would [Dar97e, LO90, Zwi91b]. WP [Wat96]. WRAPPER [Ven92]. Wrappers [Mon93]. Wrapping [GS94a, Mar96b]. Wrapup [All92]. Wrench [Cox98d]. Write [CS92, CS94, Har96a, Jus94, VZ91, ZSB94]. Writing [CLD^{+}97, Chi97, Cla99, FWF^{+}90, Gue92, JJ90, KS94b, Lib96b, Lib97b, Seo93a]. Wrong [Jac94]. wshdbg [Vck98]. Wux [Spi94]. WWW [Bal96, BDR97, BCE98, CI97, HS95a, Kar96, LB97, TW96]. WWWhither [Dam97b].

X [E^{+}91, Mur99b, NN93b, BAKW92, Cox97a, CG92b, Dav94a, ES90, JHC94, Kah95, KDF^{+}93, LaS90, NN93a, Ped90, TP91, TW93, USE96j, Wal95c, YD94, vR95]. X-kernel [JHC94]. X.500 [How90, How91]. X/Open [Wal95c]. X11 [Che94a, GBB93, Ous91b]. x86 [CH97]. XDR [MMR92]. XDR/RPC [MMR92]. XFS [SDH^{+}96]. Xhier [Sel91]. XI [Ano97i, USE97c]. XIELib [Laz98]. XII [USE98g]. XIII [USE99i]. Xinu [CG91]. Xlib [STW93b]. xmb [TM97]. XML [Bal98, SD98b]. XODUS [WUB^{+}90]. Xscal [Cou96b]. Xt [HO94]. XTP [SLC94].

Y2K [Dam99]. yacc [Pax91]. Yard [SMH95]. YaST [HH99]. Year [Cha98, Dar97h, Joh93b, Mil97f, Mil98b]. Years [AP99, Lib97b, Phe95, Sal97i, Sal98h, Sal98i, Sal98j, Sal98k, Sal98l, Sal98m, Sal99a, Sal99b, Sal99c, Sal99d, Sal99e, Sal99f, She93a, She93b, She93c, She94a, She94b, Wil98].

York [USE95g]. You’Goo [Dar96c]. You’re [Pom93d]. Yourself [CV96, FO96, Lim97, Sel98c]. You’ve [Arn91, Wal95c]. Ypmake [CAY99].

Zebra [HO92]. Zen [Bar94d, Bar92d, Uma96c]. Zephyr [Mee97, WAKS97]. Zero [Cha98, Chu96, Pla97]. Zero-Copy [Chu96]. Zimmerman [Ros95a]. Zimmermann [Ros96d]. Zone [Van97, Van90]. Zoo [Pea93]. Zoomable [BH95].

References

Ananthanarayanan:1992:ASC

[AAL92] R. Ananthanarayanan, M. Ahamad, and R. J. LeBlanc. Application specific coherence control for high performance distributed shared memory. In USENIX [USE92l], pages 109– 128. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Abrosimov:1992:DCS

[AAO92] Vadim Abrosimov, François Armand, and Maria Ines Ortega. A distributed consistency server for the CHORUS system. In USENIX [USE92l], pages 129– 148. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Anderson:1991:CRT

[AB91] David Anderson and Jeff Bilmes. Concurrent real-time music in C++. In USE-NIX [USE91k], pages 147–162. LCCN QA76.73.C153 U27 1991.

Andrews:1996:CNI

[AB96] Martin B. Andrews and Richard C.

Burgess. A clinical neurophysiology information system based on Tcl/Tk. In USE-NIX [USE96e], pages 231–235. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/andrews.html.

Abbey:1994:GAS

[Abb94] Jonathan Abbey. The group administration shell and the GASH network computing environment. In USENIX [USE94h], pages 191–203. ISBN 1-880446-64-2. LCCN QA 76.76 O63

L37 1994. URL http://www.
usenix.org/publications/library/
proceedings/lisa94/abbey.
html.

Abbey:1995:RLS

[Abb95] Jonathan Abbey. Response to LISA summary on GASH. ;login: the USENIX Association newsletter, 20(1):21–??, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Atkins:1997:EDS

[ABB^{+}97] David Atkins, Thomas Ball, Michael Benedikt, Glenn Bruns, Kenneth Cox, Peter Mataga, and Kenneth Rehor. Experience with a domain specific language for form-based services. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/atkins.html.

Artale:1997:WNF

[AC97] Frank Artale and Felipe Cabrera. Windows NT futures, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/technical.html.

Unpublished invited talk at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Allsopp:1999:PLE

[ACFM99] N. K. Allsopp, T. P. Cooper, P. Ftakas, and P. C. Macey. Porting legacy engineering applications onto distributed NT systems. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/allsopp.html.

Anderson:1998:CBN

[ACG^{+}98] Darrell C. Anderson, Jeffrey S.

Chase, Syam Gadde, Andrew J. Gallatin, Kenneth G. Yocum, and Michael J. Feeley. Cheating the I/O bottleneck: Network storage with Trapeze/ Myrinet. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/anderson.html.

Atkins:1992:EOD

[ACO92] M. Stella Atkins, Yan Chen, and Florina Olariu. Experiences: Overcoming data transfer bottlenecks across SUN-Transputer interfaces. In USE-NIX [USE92b], pages 159–192.

Apisdorf:1996:OFA

[ACTW96] Joel Apisdorf, K. Claffy, Kevin Thompson, and Rick Wilder. OC3MON: Flexible, affordable, high performance statistics collection. In USE-NIX [USE96j], pages 97–

112. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/kc.html.

Athan:1993:AMM

[AD93] Andrew Athan and Daniel Duchamp. Agent-mediated message passing for constrained environments. In USENIX [USE93i], pages 103–107. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993. URL http://

www.usenix.org/publications/
library/proceedings/mobile93/
athan.html.

Adams:1993:MOU

[Ada93] Rick Adams. More old USENET maps. ;login: the USENIX Association newsletter, 18(3):21– 22, May/June 1993. CODEN LOGNEM. ISSN 1044-6397.

Adams:1994:OMC

[Ada94]  Evan Adams.  The old

man and the C. In USE-NIX [USE94j], pages 15–26. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/adams.html.

Addesso:1996:BIB

[Add96] Mark Addesso. Building independent black box components in C++. In USENIX [USE96g], pages 25–34. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/addesso.html.

Alagappan:1990:PDA

[ADKN90] K. Alagappan, A. M. De Alvare, D. Klein, and C. Neuman. Panel and discussion on authentication. In USENIX [USE90i], page ?? LCCN QA 76.9 A25 U55 1990.

Amer:1996:BBD

[AEK96] Ahmed Amer and Amr El-Kadi.

Beating bottlenecks in the design of distributed file systems. ;login: the USENIX Association newsletter, 21(6):14–23, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Attardi:1994:CMM

[AF94] Giuseppe Attardi and Tito Flagella. A customizable memory management framework. In USENIX [USE94e], pages 123–142. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/attardi.html.

Akyol:1999:SOS

[AFJ^{+}99] Bora A. Akyol, Matt Fredette, Alden W. Jackson, Rajesh Krishnan, David Mankins, Craig Partridge, Nicholas Shectman, and Gregory D. Troxel. Smart office spaces. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999. URL

http://www.usenix.org/publications/
library/proceedings/es99/
akyol.html.

Antoniotti:1997:SSA

[AG97] Marco Antoniotti and Aleks Göllü. SHIFT and SMART-AHS: a language for hybrid system engineering modeling and simulation. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/antoniotti.
html.

Aura:1999:SLM

[AG99] Tuomas Aura and Dieter Gollmann. Software license management with Smart Cards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

http://www.usenix.org/publications/
library/proceedings/smartcard99/
aura.html.

Abrams:1990:PDA

[AGF90] M. D. Abrams, D. A. Gomberg, and C. W. Fink. Panel and discussion on access control. In USENIX [USE90i], page ?? LCCN QA 76.9 A25 U55 1990.

Auerbach:1994:CCL

[AGG^{+}94] Joshua S. Auerbach, Arthur P. Goldberg, Ajei S. Gopal, Mark T. Kennedy, and James R. Russell. Concert/C: a language for distributed programming. In USENIX [USE94l], pages 79–96. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/jrussell.html.

Allison:1998:FSS

[AHB^{+}98] Bridget Allison, Rob Hawley, Andrea Borr, Mark Muhlestein, and Dave Hitz. File system security: Secure network data sharing for NT and Unix. In USENIX [USE98e], pages 17–26. ISBN 1-880446-96-0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/allison.html;
http://www.usenix.org/publications/
library/proceedings/lisa-
nt98/allison_slides.

Angebranndt:1991:IAT

[AHL^{+}91] Susan Angebranndt, Richard L.

Hyde, Daphne Huetu Loung, Nagendra Siravara, and Chris Schmandt. Integrating audio and telephony in a distributed workstation environment. In USENIX [USE91g], pages 419–

436. LCCN QA 76.76 O63 U83 1991.

Ahlberg:1999:BRI

[Ahl99] Bo Ahlberg. Bring the “real”

Internet to Windows NT, 1999.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/ahlberg_slides.

Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Ahrendt:1993:BI

[Ahr93] Gunter Ahrendt. Big iron. ;login: the USENIX Association newsletter, 18(3):23–24, May/ June 1993. CODEN LOGNEM. ISSN 1044-6397.

Arnold:1995:MII

[AHS95] Ken Arnold, Kee Hinckley, and Eric Sheinbrood. Mediaindependent interfaces in a media-dependent world. In USENIX [USE95k], pages 195–

203. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/arnold.html.

Alexandrov:1997:EOS

[AISS97] Albert D. Alexandrov, Maximilian Ibel, Klaus E. Schauser, and Chris J. Scheiman. Extending the operating system at the user-level: the Ufo global file system. In USENIX [USE97e], pages 77–

90. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ana97/alexandrov.html;   http:
                                        //www.usenix.org/publications/
                                        library/proceedings/ana97/
                                        summaries/alexandrov.html.

Anderson:1994:SFS

[AK94] Karl A. Anderson and Brian H.

Kirouac. A simple and free system for automated network backups. In USENIX [USE94m], pages 63–68. ISBN ???? LCCN ????

Anderson:1996:TRC

[AK96] Ross Anderson and Markus Kuhn. Tamper resistance

— a cautionary note. In USENIX [USE96h], pages 1–

11. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/kuhn.html.

Anupam:1999:UPV

[AKM99] Vinod Anupam, David M. Kristol, and Alain Mayer. A user’s and programmer’s view of the new JavaScript security model. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/anupam.html.

Angelovich:1998:NGB

[AKS98] Stephen J. Angelovich, Kevin B.

Kenny, and Brion D. Sarachan. NBC’s GEnesis broadcast automation system: From prototype to production. In USENIX [USE98f], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998. URL

http://db.usenix.org/publications/
library/proceedings/tcl98/
angelovich.html.

Allison:1993:CCR

[AL93] Colin Allison and Mike Livesey. Coping with concurrency in real time groupware. In USE-NIX [USE93j], pages 289–

295. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sedms4/.

Alfieri:1994:EKB

[Alf94] Robert A. Alfieri. An efficient kernel-based implementation of POSIX threads. In USENIX [USE94j], pages 59–

72. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/alfieri.html.

Allman:1992:SFC

[All92] Eric Allman. San Francisco Conference wrapup. ;login: the USENIX Association newsletter, 17(2):17–22, 1992. CODEN LOGNEM. ISSN 1044-6397.

Allen:1995:FNV

[All95] Jeff R. Allen. Finding a needle in a virtual haystack: Whois++ and the Whois++ client library. In USENIX [USE95h], pages 25–34. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa95/allen.html.

Allen:1996:SAE

[All96] Jeff Allen. System administration employment from a student’s point of view. ;login: the USENIX Association newsletter, 21(1):17–18, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Allison:1997:SFP

[All97] Jeremy Allison. The Samba file and print server. ;login: the USENIX Association newsletter, 22(6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-11/samba.html.

Allen:1999:DRV

[All99] Jeff R. Allen. Driving by the rear-view mirror: Managing a network with Cricket. In USE-NIX [USE99g], page ?? ISBN 1-880446-36-7. LCCN TK5105.5 .C664 1999. URL http://www.

usenix.org/publications/library/
proceedings/neta99/allen.
html.

Abrams:1992:RFM

[ALLO92] Marshall Abrams, Leonard La-Padula, Manette Lazear, and Ingrid Olson. Reconciling a formal model and a prototype implementation: Lessons learned in implementing the ORGCON policy. In USENIX [USE92m], pages 313–327. ISBN 1-880446-46-4. LCCN ????

Almeida:1999:FFI

[Alm99] Danilo Almeida. FIFS: a framework for implementing usermode file systems in Windows NT. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt99/almeida.html.

Abrams:1990:BGA

[ALO90] Marshall D. Abrams, Leonard J. LaPadula, and Ingrid M. Olson. Building generalized access control on UNIX. In USENIX [USE90i], pages 65–70. LCCN QA 76.9 A25 U55 1990.

Alt:1993:RMS

[Alt93] Howard Alt. Removable media in Solaris. In USE-NIX [USE93l], pages 281–287. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Avery:1997:DDN

[AM97] Joel Avery and Andrew Macpherson. DNS —- doing nothing the same, 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/lisa97/invited_
                                        talks/avery/avery.pdf;   http:
                                        //www.usenix.org/publications/
                                        library/proceedings/lisa97/
                                        invited_talks/avery/avery.

ps. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Abbey:1998:GEC

[AM98a] Jonathan Abbey and Michael Mulvaney. Ganymede: An extensible and customizable directory management framework. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/abbey.html.

Anupam:1998:SWB

[AM98b] Vinod Anupam and Alain Mayer. Security of Web browser scripting languages: Vulnerabilities, attacks, and remedies. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/anupam.html.

Amaranth:1998:TBM

[Ama98]  Paul Amaranth.  A Tcl-

based multithreaded test harness. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/amaranth.
html.

Aahlad:1996:ANA

[AMML96] Yeturu Aahlad, Bruce E. Martin, Mod Marathe, and Chung Lee. Asynchronous notifications among distributed objects. In USENIX [USE96g], pages 83–95. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/aahlad.html.

Atkinson:1996:IIB

[AMP^{+}96] Randall J. Atkinson, Daniel L. McDonald, Bao G. Phan, Craig W. Metz, and Kenneth C. Chin. Implementation of IPv6 in 4.4 BSD. In USENIX [USE96k], pages 113–125 (or 113–126??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd96/atkinson.html.

Anderson:1998:ERL

[AMP98] Ross J. Anderson, Václav Matyásv Jr., and Fabien A. P. Petitcolas. The Eternal Resource Locator: An alternative means of establishing trust on the World Wide Web. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec98/anderson.html.

Anastacio:1999:ART

[Ana99] Rui Miguel Dias Anastacio. Agenda: a reminder tool. ;login: the USENIX Association newsletter, 24(1s): ??, January 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-1/

agenda.html. Special issue on

tools.

Anderson:1991:MPB

[And91] Paul Anderson. Managing program binaries in a heterogeneous UNIX network. In USENIX [USE91f], pages 1–10.

Anderson:1992:EUL

[And92] Paul Anderson. Effective use of local workstation disks in an NFS network. In USENIX [USE92e], pages 1–8. ISBN 1-880446-47-2. LCCN ????

Anderson:1993:RL

[And93] Paul Anderson. Report from LISA ’92. ;login: the USENIX Association newsletter, 18(2): 18–20, March/April 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Anderson:1994:THL

[And94]  Paul Anderson.  Towards

a high-level machine configuration system. In USE-NIX [USE94h], pages 19–26. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/anderson.html.

Anderson:1995:RSS

[And95] Eric Anderson. Results of the 1995 SANS survey. ;login: the USENIX Association newsletter, 20(5):13–18, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Anderson:1997:TPR

[And97] Paul Anderson. Teaching Perl to “real world” users. ;login: the USENIX Association newsletter, 22(2):27–??, April 1997. CODEN LOGNEM. ISSN 1044-

Announcement:1997:AUW

[Ann97] Announcement. Announcement: 3rd USENIX Workshop on Electronic Commerce. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

Annis:1998:BRI

[Ann98] William S. Annis. Book review: Intranet Security: Stories from the Trenches. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-4/
                                        intranet.html.

Anonymous:1992:SSH

[Ano92] Anonymous. System support for high performance multiprocessing. In USENIX [USE92l], pages 1–12. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Anonymous:1993:IV

[Ano93a] Anonymous. Index to volumes 1-5. In USENIX [USE93d], pages 37–50.

Anonymous:1993:SAC

[Ano93b] Anonymous. SAGE — AU conference. ;login: the USENIX Association newsletter, 18(3):14– ??, May/June 1993. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1993:TS

[Ano93c] Anonymous. The TeleRead solution. ;login: the USENIX Association newsletter, 18(5):45–??, September/October 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1994:UUB

[Ano94a] Anonymous. USL and UC Berkeley reach agreement. ;login: the USENIX Association newsletter, 19(2):12–??, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1994:WNa

[Ano94b] Anonymous. What’s new. ;login: the USENIX Association newsletter, 19(1):30–31, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1994:WNb

[Ano94c] Anonymous. What’s new. ;login:

the USENIX Association newsletter, 19(2):29–32, March/April 1994. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1994:WNd

[Ano94d] Anonymous. What’s new. ;login: the USENIX Association newsletter, 19(4):32–36, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1994:WNe

[Ano94e] Anonymous. What’s new. ;login: the USENIX Association newsletter, 19(5):35–42, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1994:WCR

[Ano94f] Anonymous. Winter ’94 conference reports. ;login: the USE-NIX Association newsletter, 19 (2):3–9, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1995:FS

[Ano95a]  Anonymous.  1994 financial

statements. ;login: the USE-NIX Association newsletter, 20 (3):7–??, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1995:LAA

[Ano95b]  Anonymous.  1995 lifetime

achievement award. ;login: the USENIX Association newsletter, 20(1):17–??, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1995:CNT

[Ano95c] Anonymous. Call for nominations for two awards. ;login: the USENIX Association newsletter, 20(5):9–??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1995:SAT

[Ano95d] Anonymous. Summary of the actions taken at the 1994 SAGE board of directors meetings. ;login: the USENIX Association newsletter, 20(2):32–34, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1996:DMM

[Ano96a] Anonymous. Dear miss management. ;login: the USE-NIX Association newsletter, 21 (4):27–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1996:HUC

[Ano96b] Anonymous. Highlights from 1995 USENIX conferences,

1996.  URL http://www.
usenix.org/publications/library/
proceedings/sd96/. Unpub-

lished invited talk at the USE-NIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Anonymous:1996:HSC

[Ano96c] Anonymous. Horror story contest winners. ;login: the USE-NIX Association newsletter, 21 (6):47–??, December 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1996:SCR

[Ano96d] Anonymous. SANS ’96 conference reports. ;login: the USE-NIX Association newsletter, 21 (4):24–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1996:ULS

[Ano96e] Anonymous. USACO: Letters from students. ;login: the USE-NIX Association newsletter, 21 (4):6–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:ASO

[Ano97a] Anonymous. Announcement:

3rd Symposium on Operating Systems Design and Implementation. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:AAT

[Ano97b] Anonymous. Announcement: 6th Annual Tcl/Tk Conference. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:AUS

[Ano97c] Anonymous. Announcement: 7th USENIX Security Symposium. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:BNW

[Ano97d]  Anonymous.  A brave new

world: UNIX developers in an NT land. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1997-12/
                                        brave.html.

Anonymous:1997:C

[Ano97e] Anonymous. The chase is on. ;login: the USENIX Association newsletter, 22(1):18–21, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:DSP

[Ano97f] Anonymous. Deadline for student proposals. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:LSS

[Ano97g] Anonymous. Large scale system administration of Windows NT Workshop ’97. ;login: the USE-NIX Association newsletter, 22 (6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/summaries.

Anonymous:1997:LP

[Ano97h] Anonymous. Letter from the President. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:LXC

[Ano97i] Anonymous. LISA XI conference on domain-specific languages. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:MLA

[Ano97j] Anonymous. More letters about ‘Whither USENIX?’. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:SBE

[Ano97k] Anonymous. SAGE board election results. ;login: the USE-NIX Association newsletter, 22 (1):8–??, February 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:SCE

[Ano97l] Anonymous. The SAGE code of ethics. ;login: the USENIX Association newsletter, 22(1):10– 11, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:SHM

[Ano97m] Anonymous. Seminar held for Maryland Virtual High School Program. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:SPU

[Ano97n] Anonymous. Stop the presses: Usenix/Uselinux in Anaheim. Linux Journal, 36:??, April 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:UWG

[Ano97o] Anonymous. USA wins gold at Central European Programming Contest. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1997:UWN

[Ano97p] Anonymous. USENIX Windows NT Workshop ’97. ;login: the USENIX Association newsletter, 22(6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/library/
proceedings/usenix-nt97/summaries.

Anonymous:1998:ACU

[Ano98a]  Anonymous.  Announcement

and call: 3rd USENIX Workshop on Electronic Commerce. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
reference.html.

Anonymous:1998:ACC

[Ano98b]  Anonymous.  Announcement

and call: COOTS’98. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-2/reference.html.

Anonymous:1998:ACL

[Ano98c]  Anonymous.  Announcement

and call: LISA ’98. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-2/reference.html.

Anonymous:1998:ACS

[Ano98d]  Anonymous.  Announcement

and call: SANE ’98. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-2/reference.html.

Anonymous:1998:ACN

[Ano98e] Anonymous. Announcement:

1st Conference on Network Administration. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AIS

[Ano98f] Anonymous. Announcement: 1st International SANE Conference. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AUWc

[Ano98g] Anonymous. Announcement: 1st USENIX Workshop on Intrusion Detection and Network Monitoring. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:ACD

[Ano98h] Anonymous. Announcement: 2nd Conference on Domain-Specific Languages. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AIW

[Ano98i] Anonymous. Announcement: 2nd IEEE Workshop on Mobile Computing Systems and Applications (WMCSA ’99). ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AUS

[Ano98j] Anonymous. Announcement: 2nd USENIX Symposium on Internet Technologies and Systems (USITS ’99). ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AUWa

[Ano98k] Anonymous. Announcement:

2nd USENIX Windows NT Symposium. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:ASO

[Ano98l] Anonymous. Announcement:

3rd Symposium on Operating Systems Design and Implementation. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AUWb

[Ano98m] Anonymous. Announcement:

3rd USENIX Workshop on Electronic Commerce. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:ACO

[Ano98n] Anonymous. Announcement:

5th Conference on Object-Oriented Technologies and Systems. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AAT

[Ano98o] Anonymous. Announcement:

6th Annual Tcl/Tk Conference. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AEU

[Ano98p] Anonymous. Announcement: Eighth USENIX Security Symposium. ;login: the USENIX Association newsletter, 23(3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397. Special issue on security.

Anonymous:1998:ALI

[Ano98q] Anonymous. Announcement: Large Installation System Administration of Windows NT Conference. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AL

[Ano98r] Anonymous. Announcement: LISA ’98. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AN

[Ano98s] Anonymous. Announcement: NordU99. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AO

[Ano98t] Anonymous. Announcement: OSDI ’99. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AFA

[Ano98u] Anonymous. Announcement: The Fifth Annual International Conference on Mobile Computing and Networking (MobiCom ’99). ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AUA

[Ano98v] Anonymous. Announcement:

USENIX 1999 Annual Technical Conference. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:AWN

[Ano98w] Anonymous. Announcement:

Windows NT Tutorials. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:CRC

[Ano98x] Anonymous. Conference report:

Conference on Domain-Specific Languages. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/summaries.

Anonymous:1998:CRE

[Ano98y] Anonymous. Conference report: Eleventh Systems Administration Conference (LISA ’97). ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

publications/library/proceedings/
lisa97/summaries.

Anonymous:1998:CRUd

[Ano98z] Anonymous. Conference report: 2nd USENIX 1998 Windows NT Symposium. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/summaries.html.

Anonymous:1998:CRUb

[Ano98-27] Anonymous. Conference report: 4th USENIX Conference on Object-Oriented Technologies and Systems (COOTS). ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
coots98/summaries.html.

Anonymous:1998:CRA

[Ano98-28] Anonymous. Conference report: 7th Annual System Administration, Networking and Security Conference (SANS 1998). ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
coots98/summaries.html.

Anonymous:1998:CRL

[Ano98-29] Anonymous. Conference report: Large Installation System Administration of Windows NT Conference. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/summaries.html.

Anonymous:1998:CRS

[Ano98-30] Anonymous. Conference report: Seventh USENIX Security Symposium. ;login: the USE-NIX Association newsletter, 23 (3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/summaries. Special issue

on security.

Anonymous:1998:CRT

[Ano98-31] Anonymous. Conference report: Third USENIX Workshop on Electronic Commerce. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
ec98/summaries.html.

Anonymous:1998:CRUc

[Ano98-32] Anonymous. Conference report:

USENIX 1998 Annual Technical Conference. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/summaries/index.
html.

Anonymous:1998:CRUa

[Ano98-33] Anonymous. Conference report: USENIX Symposium on Internet Technologies and Systems. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits97/summaries.html.

Anonymous:1998:ER

[Ano98-34] Anonymous. Election results. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:LEE

[Ano98-35] Anonymous. Letter to the editor: Erratum. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:MD

[Ano98-36] Anonymous. Member dues. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:NES

[Ano98-37] Anonymous. New editor for short topics series. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:NAM

[Ano98-38] Anonymous. Notice of annual meeting. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:RS

[Ano98-39] Anonymous. Riding in style.

;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:SPD

[Ano98-40] Anonymous. Student proposal deadlines. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:TM

[Ano98-41] Anonymous. Torn money. ;login:

the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:ULS

[Ano98-42] Anonymous. USACO — letters from students. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:UCC

[Ano98-43] Anonymous. USACO camp completed. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1998:UIP

[Ano98-44] Anonymous. USENIX international programs. ;login: the USENIX Association newsletter, 23(7):??, December 1998.

CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:OAA

[Ano99a] Anonymous. The 1998 outstanding achievement award. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
achievement_award.html.

Anonymous:1999:ASAb

[Ano99b] Anonymous. Announcement: 13th Systems Administration Conference (LISA ’99). ;login: the USENIX Association newsletter, 24(2):??, April 1999. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:AUAa

[Ano99c] Anonymous. Announcement: 1999 USENIX Annual Technical Conference. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:AUAb

[Ano99d] Anonymous. Announcement: 2000 USENIX Annual Technical Conference. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/usenix2000.

Anonymous:1999:AUAc

[Ano99e] Anonymous. Announcement: 2000 USENIX Annual Technical Conference. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/usenix2000/.

Anonymous:1999:ACD

[Ano99f] Anonymous. Announcement:

2nd Conference on Domain-Specific Languages. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/library/
proceedings/dsl99/.

Anonymous:1999:AUSa

[Ano99g] Anonymous. Announcement:

2nd USENIX Symposium on Internet Technologies and Systems. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/.

Anonymous:1999:AAA

[Ano99h] Anonymous. Announcement:

3rd Annual Atlanta Linux Showcase. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.linuxshowcase.org.

Anonymous:1999:AAL

[Ano99i] Anonymous. Announcement:

3rd Annual Linux Showcase. ;login: the USENIX Association newsletter, 24(5s):??, September 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

linuxshowcase.org. Special is-

sue on intrusion detection.

Anonymous:1999:ALIa

[Ano99j] Anonymous. Announcement: 3rd Large Installation System Administration of Windows NT / 2000 Conference. ;login: the USENIX Association newsletter, 24(5s):??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/lisa-nt2000. Special

issue on intrusion detection.

Anonymous:1999:ALIb

[Ano99k] Anonymous. Announcement: 3rd Large Installation System Administration of Windows NT/2000 Conference. ;login: the USENIX Association newsletter, 24(6s):??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/lisa-nt2000.

Anonymous:1999:AUWa

[Ano99l] Anonymous. Announcement: 3rd USENIX Windows Systems Symposium. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99.

Anonymous:1999:ASO

[Ano99m] Anonymous. Announcement:

4th Symposium on Operating Systems Design and Implementation (OSDI 2000). ;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/events/osdi2000.

Anonymous:1999:AUWb

[Ano99n] Anonymous. Announcement:

4th USENIX Windows Systems Symposium. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/usenix-win2000.

Anonymous:1999:AUT

[Ano99o] Anonymous. Announcement:

7th USENIX Tcl/Tk Conference. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/tcl2k.

Anonymous:1999:AUSb

[Ano99p] Anonymous. Announcement:

9th USENIX Security Symposium. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
events/sec2000. Special issue

on intrusion detection.

Anonymous:1999:AUSc

[Ano99q] Anonymous. Announcement: 9th USENIX Security Symposium. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        events/sec2000.

Anonymous:1999:AAM

[Ano99r] Anonymous. Announcement: ASA/MA ’99: Agent Systems and Applications/Mobile Agents. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:ACCa

[Ano99s] Anonymous. Announcement: COOTS ’99: 5th Conference on Object-Oriented Technologies and Systems. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:ACCb

[Ano99t] Anonymous. Announcement: COOTS ’99: 5th Conference on Object-Oriented Technologies and Systems. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:ALN

[Ano99u] Anonymous. Announcement: LISA-NT — 2nd Large Installation System Administration of Windows NT Conference. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt99.

Anonymous:1999:ASAa

[Ano99v] Anonymous. Announcement:

SAGE-AU’99: The 7th Annual Conference of The System Administrators Guild of Australia. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:ATS

[Ano99w] Anonymous. Announcement:

Tcl/2K: The Seventh USENIX Tcl/Tk Conference. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:AUN

[Ano99x] Anonymous. Announcement:

USENIX Networking: 1st Conference on Network Administration and 1st Workshop on Intrusion Detection and Network Monitoring. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:AWA

[Ano99y] Anonymous. Announcement:

Workshop on Applications of Embedded Systems. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/events/es2000.

Anonymous:1999:AWS

[Ano99z] Anonymous. Announcement: Workshop on Smartcard Technology. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:BRT

[Ano99-27] Anonymous. Book review: Thanks to reviewers. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:CRA

[Ano99-28] Anonymous. Conference report: 6th Annual Tcl/Tk Conference. ;login: the USENIX Association newsletter, 24(1s):??, January 1999. CODEN LOGNEM. ISSN 1044-6397. Special issue on tools.

Anonymous:1999:CRRf

[Ano99-29] Anonymous. Conference report: Report on the 5th USE-NIX Conference on Object-Oriented Technologies and Systems (COOTS). ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/coots99/summaries.

Anonymous:1999:CRRc

[Ano99-30] Anonymous. Conference report: Report on the Workshop on Embedded Systems. ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
es99/summaries.

Anonymous:1999:CRRb

[Ano99-31] Anonymous. Conference report: Reports on the 3rd Symposium on Operating Systems Design and Implementation (OSDI ’99). ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
osdi99/summaries/.

Anonymous:1999:CRRg

[Ano99-32] Anonymous. Conference report: Reports on the 8th USENIX Security Symposium. ;login: the USENIX Association newsletter, 24(6s):??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
sec99/summaries.

Anonymous:1999:CRRd

[Ano99-33] Anonymous. Conference report: Reports on the First Conference on Network Administration. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/

                                        neta99/summaries/. Special is-

sue on intrusion detection.

Anonymous:1999:CRRe

[Ano99-34] Anonymous. Conference report: Reports on the Workshop on Intrusion Detection and Network Monitoring. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        detection99/summaries/. Spe-

cial issue on intrusion detection.

Anonymous:1999:CRRa

[Ano99-35] Anonymous. Conference report: Reports on the Twelfth Systems Administration Conference (LISA ’98). ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa98/summaries/.

Anonymous:1999:MMW

[Ano99-36] Anonymous. In memoriam: Mark Weiser. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:IL

[Ano99-37] Anonymous. In this issue …. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-6/
                                        in_this_issue.html.

Anonymous:1999:Ia

[Ano99-38] Anonymous. In this issue. …;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
in_this_issue.html.

Anonymous:1999:Ib

[Ano99-39] Anonymous. In this issue. … ;login: the USENIX Association newsletter, 24(5s):??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-9/
inthisissue.html. Special is-

sue on intrusion detection.

Anonymous:1999:Ic

[Ano99-40] Anonymous. In this issue. …;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
inthisissue.html.

Anonymous:1999:Id

[Ano99-41] Anonymous. In this issue. … ;login: the USENIX Association newsletter, 24(6s):??, November 1999. CODEN LOGNEM. ISSN 1044-6397. Special issue on security.

Anonymous:1999:Ie

[Ano99-42] Anonymous. In this issue. …;login: the USENIX Association newsletter, 24(6):

??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
inthisissue.html.

Anonymous:1999:JCP

[Ano99-43] Anonymous. June crossword puzzle solution. ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
junesolution.pdf.

Anonymous:1999:LMC

[Ano99-44] Anonymous. Last month’s crossword puzzle solution. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-6/solution.pdf.

Anonymous:1999:LEF

[Ano99-45] Anonymous. Letter to the Editor: On free software. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
freesoftware.html.

Anonymous:1999:LEa

[Ano99-46] Anonymous. Letters to the Editor. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
letters.html.

Anonymous:1999:LEb

[Ano99-47] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
letters.html.

Anonymous:1999:LEd

[Ano99-48] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
letters.html.

Anonymous:1999:LEe

[Ano99-49] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
letters.html.

Anonymous:1999:LEC

[Ano99-50] Anonymous. Letters to the Editor: Corrections. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-4/letters.html.

Anonymous:1999:MUN

[Ano99-51] Anonymous. Meet USENIX new staff. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:NAM

[Ano99-52] Anonymous. Notice of annual meeting. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-4/
                                        annual_meeting.html.

Anonymous:1999:SCS

[Ano99-53] Anonymous. SAGE Certification Subcommittee briefs. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-4/
                                        sage_cert.html.

Anonymous:1999:TOC

[Ano99-54] Anonymous. Thanks to our 1998 contributors. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        thanks_contrib.html.

Anonymous:1999:UAA

[Ano99-55] Anonymous. USENIX annual awards. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:UCPb

[Ano99-56] Anonymous. The USENIX crossword puzzle. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/login/
1999-6/crossword.pdf.

Anonymous:1999:UCPc

[Ano99-57] Anonymous. The USENIX crossword puzzle. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:1999:UCPa

[Ano99-58] Anonymous. The USE-NIX crossword puzzle: April puzzle and February solution. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
crossword.html.

Anonymous:1999:UPS

[Ano99-59] Anonymous. The USENIX privacy statement. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
privacy.html.

Antonelli:1995:RML

[Ant95] Charles J. Antonelli. Report on the 2nd Mobile & Location-Independent Computing Symposium. ;login: the USE-NIX Association newsletter, 20 (3):11–16, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Anderson:1992:CMF

[AOG92] David P. Anderson, Yoshitomo Osawa, and Ramesh Govindan. The continuous media file system (real-time disk storage and retrieval of digital audio/video data). In USENIX [USE92g], pages 157–164. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Arjomandi:1996:SMOb

[AOW96a] Eshrat Arjomandi, William G. O’Farrell, and Gregory V. Wilson. Smart messages: An object-oriented communication mechanism. In USE-NIX [USE96g], pages 233–

240. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/arjomandi.html.

Arjomandi:1996:SMOa

[AOW96b] Eshrat Arjomandi, William G. O’Farrell, and Gregory V. Wilson. Smart Messages: An object-oriented communication mechanism for parallel systems. In USE-NIX [USE96a], pages 313–

329. URL http://www.usenix.
org/publications/library/

proceedings/coots96/arjomandi.
html.

Anderson:1997:ESM

[AP97] Eric Anderson and Dave Patterson. Extensible, scalable monitoring for clusters of computers. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/02.anderson.html.

Anderson:1999:RTY

[AP99] Eric Anderson and Dave Patterson. A retrospective on twelve years of LISA proceedings. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/anderson.
html.

Appelman:1996:WIM

[App96] Dan Appelman. What it may mean to you. ;login: the USE-NIX Association newsletter, 21 (2):23–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Arnold:1991:CCM

[AR91] Ed Arnold and Craig Ruff. Configuration control and management. In USENIX [USE91f], pages 195–198.

Avolio:1994:NPS

[AR94] Frederick M. Avolio and Marcus J. Ranum. A network perimeter with secure external access. In USENIX [USE94m], pages 1–14. ISBN ???? LCCN ????

Archer:1993:TPS

[Arc93]  Barrie Archer.  Towards a

POSIX standard for software administration. In USE-NIX [USE93f], pages 67–

79. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/archer.html.

Adamson:1995:JSR

[ARH95] William A. Adamson, Jim Rees, and Peter Honeyman. Joining security realms: a single login for NetWare and Kerberos. In USENIX [USE95f], pages 157–166. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/security95/
                                        adamson.html.

Arnold:1990:EOF

[Arn90] James Q. Arnold. A ELF: An object file to mitigate mischievous misoneism. In USENIX [USE90g], pages 1–10. LCCN QA76.8.U65 U81 1990.

Arnold:1991:IYS

[Arn91] Bob Arnold. If you’ve seen one UNIX, you’ve seen them all. In USENIX [USE91f], pages 11–20.

Arnott:1995:MUI

[Arn95a] Ken Arnott. Monitoring user’s idle time. ;login: the USE-NIX Association newsletter, 20 (5):13–??, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Arnow:1995:DLB

[Arn95b] David M. Arnow. DP: a library for building portable, reliable distributed applications. In USENIX [USE95e], pages 235–247. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/arnow.html.

Arnold:1998:AUC

[Arn98] Bob Arnold. Accountworks:

Users create accounts on SQL, Notes, NT, and UNIX. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
arnold.html.

Akyurek:1992:PRD

[AS92] Sedat Akyurek and Kenneth Salem. Placing replicated data to reduce seek delays. In USE-NIX [USE92h], pages 133–134. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Akyurek:1993:ABR

[AS93] Sedat Akyurek and Kenneth Salem. Adaptive block rearrangement under UNIX. In USENIX [USE93g], pages 307–

321. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/akyurek.html.

Arnold:1999:DDC

[ASB^{+}99] David Arnold, Bill Segall, Julian Boot, Andy Bond, Melfyn Lloyd, and Simon Kaplan. Discourse with disposable computers: How and why you will talk to your tomatoes. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        es99/arnold.html.

Adams:1993:INM

[ASG^{+}93] Norman Adams, Bill N. Schilit, Rich Gold, Michael Tso, and Roy Want. An infrared network for mobile computers. In USENIX [USE93i], pages 41–51. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mobile93/adams.html.

Abdel-Shafi:1999:EUL

[ASSB99] Hazim Abdel-Shafi, Evan Speight, and John K. Bennett. Efficient user-level thread migration and checkpointing on Windows NT clusters. In USE-NIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt99/abdel-shafi.html.

Allen:1991:PLO

[ASWW91] Larry W. Allen, Harminder G. Singh, Kevin G. Wallace, and Melanie B. Weaver. Program loading in OSF/1. In USENIX [USE91i], pages 145–160. LCCN QA 76.76 O63 U84 1992.

Alashqur:1992:RGS

[AT92] Abdullah Alashqur and Craig Thompson. O-R Gateway: a system for connecting C++ application programs and relational databases. In USENIX [USE92k], pages 151–169. ISBN 1-880446-45-6. LCCN ????

Atkinson:1996:IIS

[Atk96] Ran Atkinson. IPv6 & IPsec software available. ;login: the USENIX Association newsletter, 21(5):18–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Alexander:1994:EHS

[ATS94] D. Scott Alexander, C. Brendan S. Traw, and Jonathan M. Smith. Embedding high speed ATM in UNIX IP (abstract only). In USE-NIX [USE94f], pages 119–

121. ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
hispeed/index.html.

Antonelli:1999:PVS

[AUH99] C. J. Antonelli, M. Undy, and P. Honeyman. The Packet Vault: Secure storage of network data. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/antonelli.html.

Authors:1994:LL

[Aut94] Various Authors. Letters to ;login. ;login: the USENIX Association newsletter, 19(5):3–7, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Authors:1995:UTC

[Aut95a] Various Authors. 1995 USE-NIX technical conference reports. ;login: the USENIX Association newsletter, 20(2):8–25, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Authors:1995:LVC

[Aut95b] Various Authors. LISA VIII conference reports. ;login: the USENIX Association newsletter, 20(1):18–21, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Avolio:1996:GE

[Avo96] Frederick M. Avolio. Guest editorial. In USENIX [USE96d], pages 1–2.

Avolio:1998:PCP

[Avo98] Frederick M. Avolio. Practical cryptography — privacy for business and electronic commerce, 1998. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/lisa98/invited_
                                        talks/avolio_html/. Unpub-

lished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6– 11, 1998 Boston, Massachusetts, USA.

Asokan:1998:TFH

[AVS98] N. Asokan, Els Van Herreweghen, and Michael Steiner. Towards A framework for handling disputes in payment systems. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

URL http://www.usenix.org/
publications/library/proceedings/
ec98/asokan.html.

Abler:1999:BHV

[AW99a] Randal Abler and Gail Wells.

Supporting H.323 video and voice in an enterprise network. In USENIX [USE99g], page ?? ISBN 1-880446-36-

7. LCCN TK5105.5 .C664 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/abler.html.

Allen:1999:FSA

[AW99b] Jeff R. Allen and David Williamson. The four-star approach to network management,

1999. URL http://db.usenix.
org/publications/library/
proceedings/lisa99/invitedtalks/

allen.pdf. Unpublished invited talk at LISA XIII.

Blaze:1991:LTC

[BA91] Matt Blaze and Rafael Alonso.

Long-term caching strategies for very large distributed file systems. In USENIX [USE91g], pages 3–16. LCCN QA 76.76 O63 U83 1991.

Blaze:1992:IMS

[BA92] Matt Blaze and Rafael Alonso. Issues in massive scale distributed file systems. In USE-NIX [USE92h], pages 135–136. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Bertrand:1997:BLC

[BA97] Frédéric Bertrand and Michel Augeraud. BDL: a language to control the behavior of concurrent objects. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/dsl97/bertrand.
                                        html.

Babaoglu:1990:FTC

[Bab90] Ozalp Babaoglu. Fault-tolerant computing based on Mach. In USENIX [USE90e], pages 185– 199.

Bacon:1995:OTN

[Bac95] Jeff Bacon. Opinion: Thoughts on the net. ;login: the USE-NIX Association newsletter, 20 (4):4–5, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Bacon:19xx:OOR

[Bacxx] David F. Bacon. OptiMach: Optimistic recovery of Mach tasks. In ???? USENIX, Berkeley, CA, USA, 19xx.

Bahn:1990:BBL

[Bah90]  Pat Bahn.  Beyond Bell-

LaPadula: a security model for real applications. In USENIX [USE90i], pages 63–64. LCCN QA 76.9 A25 U55 1990.

Bahreman:1996:GEP

[Bah96] Alireza Bahreman. Generic electronic payment services: Framework and functional specification. In USENIX [USE96h], pages 87–103. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/bahreman.html.

Bailey:1997:CRR

[Bai97] Brian Bailey. Conference report:

Reports on the Fifth Tcl/Tk Workshop. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/summaries.

Bakker:1999:MAS

[Bak99]  Bastiaan Bakker.  Mutual

authentication with Smart Cards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/bakker.html.

Blewett:1992:XWB

[BAKW92] Doug Blewett, Scott Anderson, Meg Kilduff, and Mike Wish. X widget based software tools for UNIX. In USENIX [USE92j], pages 111–124.

Balan:1992:SIV

[Bal92] Ramesh Balan. A scalable implementation of virtual memory HAT layer for shared memory multiprocessor machines. In USENIX [USE92g], pages 107–

116. ISBN 1-880446-44-8. LCCN

QA 76.76 O63 U83 1992.

Ball:1996:SWB

[Bal96] Steve Ball. SurfIt! — A WWW browser. In USE-NIX [USE96e], pages 161–171. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/ball.html.

Ball:1998:XST

[Bal98] Steve Ball. XML support for Tcl. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl98/ball.html.

Barrera:1991:FMN

[Bar91] Joseph S. Barrera, III. A fast Mach network IPC implementation. In USENIX [USE91h], pages 1–12. LCCN QAX 27.

Barton:1992:BRC

[Bar92a] Billy Barton. Book review: Crossing the Internet Threshold. ;login: the USENIX Association newsletter, 17(6):41–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Barton:1992:BRI

[Bar92b] Billy Barton. Book review:

The Internet Companion. ;login: the USENIX Association newsletter, 17(6):42–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Barton:1992:BRW

[Bar92c] Billy Barton. Book review: The Whole Internet Guide and Catalog. ;login: the USENIX Association newsletter, 17(5):46–47, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Barton:1992:ZAI

[Bar92d] Billy Barton. Zen and the Art of Internet. ;login: the USENIX Association newsletter, 17(4): 50–??, July/August 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Barlow:1994:NSI

[Bar94a] John Perry Barlow. Notable Speeches of the Information Age. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1994. ISBN 1-56592-992-6. LCCN ???? US$9.95. USE-NIX Conference Keynote Address, January 17, 1994; San Francisco, CA. Audio tape: 90 minutes.

Barrett:1994:IIR

[Bar94b] Thomas Barrett. Internet information resources for the system administrator. In USENIX [USE94m], pages 39–48. ISBN ???? LCCN ????

Barron:1994:RUG

[Bar94c] Billy Barron. Review: a DOS user’s guide to the Internet — E-mail, netnews, and file transfer with UUCP. ;login: the USE-NIX Association newsletter, 19 (6):54–??, December 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Barron:1994:RZA

[Bar94d] Billy Barron. Review: Zen and the Art of the Internet, 3rd Edition. ;login: the USENIX Association newsletter, 19(6):53– 54, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Barron:1996:LE

[Bar96] Billy Barron. Letter to the editor. ;login: the USENIX Association newsletter, 21(3):31–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Barber:1997:ISA

[Bar97] Michael R. Barber. Increased server availability and flexibility through failover capability. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/13.barber.html.

Barish:1998:UTR

[Bar98] Greg Barish. Using Tcl to rapidly develop a scalable engine for processing dynamic application logic. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/barish.
html.

Baskett:1998:HCC

[Bas98] Forest Baskett. Here comes cc-NUMA: The revolution in computer architecture and its OS implications, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/invited_talks/
baskett_slides. Unpublished

keynote address at the 2nd USE-NIX Windows NT Symposium, 1998 August 3-4, 1998 Seattle, Washington, USA.

Balter:1991:AIG

[BB^{+}91] R. Balter, J. Bernadat, et al. Architecture and implementation of guide, an object-oriented distributed system. In USENIX [USE91d], pages 31–68.

Bernadat:1993:RMM

[BB93] Philippe Bernadat and David Black. Real memory Mach. In USENIX [USE93h], pages 235–251. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/bernadat.html.

Bakre:1995:HSS

[BB95a] Ajay Bakre and B. R. Badrinath. Handoff and systems support for indirect TCP/IP. In USENIX [USE95i], pages 11–24. ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995.

URL http://www.usenix.org/
publications/library/proceedings/
mob95/bakre.html.

Blaze:1995:SLE

[BB95b] Matt Blaze and Steven M. Bellovin. Session-layer encryption. In USENIX [USE95f], pages 85–94 (or 85–93??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5).

URL http://www.usenix.org/
publications/library/proceedings/
security95/index.html.

Baker:1996:SST

[BBD^{+}96] Mary Baker, Bob Bruce, William H. Davidow, Michael Tiemann, and Linus Torvalds. Selling stuff that’s free: the commercial side of free software, 1996. URL http://www.

usenix.org/publications/library/
proceedings/sd96/. Unpub-

lished invited talk at the USE-NIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Bruno:1999:PCB

[BBG^{+}99] John Bruno, José Brustoloni, Eran Gabber, Avi Silberschatz, and Christopher Small. Pebble: a component-based operating system for embedded applications. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

URL http://www.usenix.org/

publications/library/proceedings/
es99/bruno.html.

Buschkes:1999:TBA

[BBK99] Roland Büschkes, Mark Borning, and Dogan Kesdogan. Transaction-based anomaly detection. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/buschkes.html.

Brutch:1999:IUH

[BBP99] Paul C. Brutch, Tasneem G.

Brutch, and Udo Pooch. Indicators of UNIX host compromise. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-9/
features/compromise.html.

Special issue on intrusion detection.

Berman:1995:TPF

[BBS95] Andrew Berman, Virgil Bourassa, and Erik Selberg. TRON: Process-specific file protection for the UNIX operating system. In USENIX [USE95e], pages 165–175. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/bourassa.html.

Buck:1991:EID

[BC91] A. Lester Buck and Robert A.

Coyne, Jr. An experimental implementation of draft POSIX asynchronous I/O. In USENIX [USE91i], pages 289–306. LCCN QA 76.76 O63 U84 1992.

Black:1994:ERI

[BC94] Richard Black and Simon Crosby. Experience and results from implementation of an ATM socket family. In USENIX [USE94l], pages 143–

152. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/black.html.

Buchanan:1997:CTS

[BC97] Matt Buchanan and Andrew A. Chien. Coordinated thread scheduling for workstation clusters under Windows NT. In USENIX [USE97f], pages 47–

53. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt97/buchanan.html.

Boyer:1991:SOO

[BCC^{+}91] F. Boyer, J. Cayuela, P. Y. Chevalier, A. Freyssinet, and Daniel Hagimont. Supporting an object-oriented distributed system: Experience with UNIX, Mach and Chorus. In USENIX [USE91j], page ?? ISBN ???? LCCN ????

Blackwell:1994:SSC

[BCC^{+}94] Trevor Blackwell, Kee Chan, Koling Chang, Thomas Charuhas, James Gwertzman, Brad Karp, H. T. Kung, David Li, Dong Lin, Robert Morris, Rob Polansky, Diane Tang, Cliff Young, and John Zao. Secure short-cut routing for mobile IP. In USENIX [USE94j], pages 305–316. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/blackwell.html.

Butrico:1997:GRM

[BCC^{+}97] Maria A. Butrico, Henry Chang, Anthony Cocchi, Norman H. Cohen, Dennis G. Shea, and Stephen E. Smith. Gold Rush: Mobile transaction middleware with Java-object replication. In USENIX [USE97g], pages 91–101. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/butrico.html. See

also [LH97, Rie97, WW97].

Brock:1999:WNC

[BCC^{+}99] B. Brock, G. Carpenter, E. Chiprout, E. Elnozahy, M. Dean, D. Glasco, J. Peterson, R. Rajamony, F. Rawson, R. Rockhold, and A. Zimmerman. Windows NT in a ccNUMA system. In USE-NIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/brock.html.

Bryant:1991:SSS

[BCCR91] Ray Bryant, Paul Carini, Hung-Yang Chang, and Bryan Rosenburg. Supporting structured shared virtual memory under Mach. In USENIX [USE91h], pages 59–76. LCCN QAX 27.

Bergadano:1998:SWT

[BCE98] F. Bergadano, B. Crispo, and M. Eccettuato. Secure WWW transactions using Standard HTTP and Java applets. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec98/crispo.html.

Black:1991:OVM

[BCF^{+}91] David Black, Jeff Carter, George Feinberg, Rod MacDonald, Jim Van Sciver, Ping Wang, Shashi Mangalat, and Eric Shienbrood. OSF/1 virtual memory improvements. In USENIX [USE91h], pages 87–104. LCCN QAX 27.

Balter:1993:MTW

[BCF^{+}93] R. Balter, P. Y. Chevalier, A. Freyssinet, D. Hagimont, S. Lacourte, and X. Rousset de Pina. Is microkernel technology well suited for the support of object-oriented systems: The guide experience. In USENIX [USE93k], pages 1–11. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/balter.html.

Bumbulis:1991:IUW

[BCGS91] Peter Bumbulis, Donald Cowan, Eric Giguere, and Terry Stepien.

Integrating UNIX within a microcomputer-oriented development environment. In USE-NIX [USE91f], pages 29–36.

Baratloo:1998:FHR

[BCH^{+}98] Arash Baratloo, P. Emerald Chung, Yennun Huang, Sampath Rangarajan, and Shalini Yajnik. Filterfresh: Hot replication of Java RMI server objects. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/baratloo.html.

Bakerji:1994:DDM

[BCK94a] A. Bakerji, D. Cohn, and D. Kulkarni. Design, distribution, and management of object oriented software. In USENIX [USE94g], pages 109–

120. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/index.html.

Banerji:1994:DDM

[BCK94b] Arindam Banerji, David Cohn, and Dinesh Kulkarni. Design, distributions, and management of object oriented software. In USENIX [USE94g], pages 109–120. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/banerji.html.

Buonadonna:1999:MSC

[BCLC99] Philip Buonadonna, Joshua Coates, Spencer Low, and David E. Culler. Millennium Sort: a cluster-based application for Windows NT using DCOM, River primitives, and the Virtual Interface Architecture. In USE-NIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt99/buonadonna.html.

Brown:1990:PDN

[BCR90] D. S. Brown, J. Carlin, and J. Roth. Panel and discussion on network security. In USENIX [USE90i], page ?? LCCN QA 76.9 A25 U55 1990.

Bryant:1991:EDR

[BCR91a] Ray Bryant, Hung-Yang Chang, and Bryan Rosenburg. Experiences developing the RP3 operating system. In USENIX [USE91j], pages 1–18. ISBN ???? LCCN ????

Byrant:1991:EDR

[BCR91b] Ray Byrant, Hung-Yang Chang, and Bryan Rosenburg. Experience developing the RP3 operating system. In USENIX [USE91c], pages 183–216.

Burcowski:1991:PES

[BCVC91] Forbes J. Burcowski, Charles L. Clerke, Gordon J. Vreugdenhil, and Crispin Cowan. Performance evaluation of the Sylvan multiprocessor architecture. In USENIX [USE91j], pages 165–

184. ISBN ???? LCCN ????

Buhr:1992:ACP

[BD92] Peter A. Buhr and Glen Ditchfield. Adding concurrency to a programming language. In USENIX [USE92k], pages 207–

223. ISBN 1-880446-45-6. LCCN ????

Bowman:1993:EDW

[BD93] C. Mic Bowman and Chanda Dharap. The enterprise distributed white-pages service. In USENIX [USE93l], pages 349–359. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Bishop:1996:CRC

[BD96] Matt Bishop and Michael Dilger. Checking for race conditions in file accesses. In USENIX [USE96b], pages 131–152.

Banga:1997:MCW

[BD97] Gaurav Banga and Peter Druschel. Measuring the capacity of a Web server. In USE-NIX [USE97h], pages 61–71. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/banga.html.

Bender:1993:UNM

[BDD^{+}93] Michael Bender, Alexander Davidson, Clark Dong, Steven Drach, Anthony Glenning, Karl Jacob, Jack Jia, James Kempf, Nachiappan Periakaruppan, Gale Snow, and Becky Wong. UNIX for nomads: Making UNIX support mobile computing. In USE-NIX [USE93i], pages 53–67. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/bender.html.

Banino:1993:DIR

[BDKM93] Jean-Serge Banino, Jean Delcoigne, Claude Kaiser, and Gerard Morisset. The DUNE-iX real-time operating system. In USENIX [USE93a], pages 425– 480.

Banga:1999:RCN

[BDM99] Gaurav Banga, Peter Druschel, and Jeffrey C. Mogul. Resource containers: a new facility for resource management in server systems. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/banga.
html.

Barton-Davis:1993:ASA

[BDMVL93] Paul Barton-Davis, Dylan Mc-Namee, Raj Vaswani, and Edward Lazowska. Adding scheduler activations to Mach 3.0. In USENIX [USE93h], pages 119–136. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Banga:1997:ODW

[BDR97] Gaurav Banga, Fred Douglis, and Michael Rabinovich. Optimistic deltas for WWW latency reduction. In USE-NIX [USE97e], pages 289–

303. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/banga.html;   http:/
/www.usenix.org/publications/
library/proceedings/ana97/
summaries/banga.html.

Benecke:1998:SCI

[BE98] Carsten Benecke and Uwe Ellermann. Securing ‘Classical IP Over ATM Networks’. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/benecke.html.

Beattie:1994:TPT

[Bea94] Malcolm Beattie. TkPerl — A port of the Tk toolkit to Perl5. In USENIX [USE94k], pages 71–82. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/beattie.html.

Beazley:1996:SEU

[Bea96] David M. Beazley. SWIG:

An easy to use tool for integrating scripting languages with C and C++. In USE-NIX [USE96e], pages 129–139. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/beazley.html.

Beardsley:1997:BRR

[Bea97] Reginald Beardsley. Book review: Rapid development: Taming wild software schedules. ;login: the USENIX Association newsletter, 22(5):68–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Beardsley:1998:BRU

[Bea98] Reginald Beardsley. Book review: UNIX Power Tools, Second Edition. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Beck:1999:DPE

[Bec99] Robert Beck. Dealing with public Ethernet jacks — switches, gateways, and authentication. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa99/beck.html.

Bedichek:1990:SEA

[Bed90] Robert Bedichek. Some efficient architecture simulation techniques. In USENIX [USE90h], pages 53–64. LCCN QA76.8.U65 U82 1990.

Beecher:1992:DLD

[Bee92] Bryan Beecher. Dealing with lame delegations. In USENIX [USE92e], pages 127–133. ISBN 1-880446-47-2. LCCN ????

Bellovin:1990:PND

[Bel90] S. M. Bellovin. Pseudo-network drivers and virtual networks. In USENIX [USE90h], pages 229–

244. LCCN QA76.8.U65 U82 1990.

Bellovin:1992:TD

[Bel92] Steve Bellovin. There be dragons. In USENIX [USE92m], pages 1–16. ISBN 1-880446-46-

4. LCCN ????

Bellovin:1995:UDN

[Bel95]  Steven M. Bellovin.  Us-

ing the Domain Name System for system break-ins. In USENIX [USE95f], pages 199–208. ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
bellovin.html.

Bell:1996:SCF

[Bel96a] John D. Bell. A simple caching file system for application serving. In USENIX [USE96j], pages 171–179. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/jbell4.html.

Bellovin:1996:PAI

[Bel96b] Steven M. Bellovin. Problem areas for the IP security protocols.

In USENIX [USE96i], pages 205–214. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/bellovin.html.

Bellovin:1999:DF

[Bel99] Steven M. Bellovin. Distributed firewalls. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-11/
features/firewalls.html. Spe-

cial issue on security.

Bhide:1991:HAN

[BEM91] Anupam Bhide, Elmootazbellah N. Elnozahy, and Stephen P. Morgan. A highly available network file server. In USENIX [USE91i], pages 199–206. LCCN QA 76.76 O63 U84 1992.

Bent:1993:SAU

[Ben93] Wilson H. Bent, Jr. System administration as a user interface: An extended metaphor. In USENIX [USE93f], pages 209–212. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/bent.html.

Bennett:1994:CAM

[Ben94a] Kris K. Bennett. The change agent mind set for technology infusion. In USENIX [USE94m], pages 137–146. ISBN ???? LCCN ????

Benson:1994:LSC

[Ben94b] Brent W. Benson, Jr. libscheme:

Scheme as a C library. In USE-NIX [USE94k], pages 7–9 (or 7–20??). ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/benson.html.

Bent:1994:YV

[Ben94c] Win Bent. Your voice. ;login: the USENIX Association newsletter, 19(5):28–29, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Bent:1995:RL

[Ben95] Win Bent. Report from LISA ’95. ;login: the USENIX Association newsletter, 20(6):21– 22, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Berliner:1990:CIP

[Ber90] Brian Berliner. CVS II: Parallelizing software development. In USENIX [USE90h], pages 341–352. LCCN QA76.8.U65 U82 1990.

Bershad:1992:III

[Ber92] Brian Bershad. The increasing irrelevance of IPC performance for microkernel-based operating systems. In USENIX [USE92i], pages 205–212. ISBN 1-880446-42-1. LCCN QAX 32.

Berglund:1996:JAT

[Ber96] Eric Berglund. Jack of all trades.

;login: the USENIX Association newsletter, 21(5):21–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Bernhard:1998:IID

[Ber98] Lee F. Bernhard. Iclient/Iserver: Distributed objects using [incr Tcl]. In USENIX [USE98f], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/bernhard.
html.

Berkowitz:1999:ER

[Ber99a] Gale Berkowitz. Election results. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
election.html.

Berkowitz:1999:ICA

[Ber99b] Gale Berkowitz. Incident Cost Analysis and Modeling Project II (I-CAMP II). ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
icamp.html.

Berkowitz:1999:MDF

[Ber99c]  Gale Berkowitz.  Member

dues and 1998 financial statements. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

publications/login/1999-8/
statements.html.

Berkowitz:1999:SCU

[Ber99d] Gale Berkowitz. SAGE certification update. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
sagecert.html.

Berkowitz:1999:SR

[Ber99e] Gale Berkowitz. The SAGE report. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
sagereport.html.

Berkowitz:1999:SRG

[Ber99f] Gale Berkowitz. Student research grants. ;login: the USE-NIX Association newsletter, 24 (5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
research.html.

Balfanz:1999:HHC

[BF99] Dirk Balfanz and Edward W.

Felten. Hand-held computers can be better Smart Cards. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/balfanz.
html.

Bomberger:1992:KRN

[BFH^{+}92] Allen C. Bomberger, William S. Frantz, Ann C. Hardy, Norman Hardy, Charles R. Landau, and Jonathan S. Shapiro. The KeyKOS(R) nanokernel architecture. In USENIX [USE92i], pages 95–112. ISBN 1-880446-42-1. LCCN QAX 32.

Berni:1992:EIS

[BFM92] Alessandro Berni, Paolo Franchi, and Joy Marino. Experiences of Internet security in Italy. In USENIX [USE92m], pages 41–47. ISBN 1-880446-46-

4. LCCN ????

Bonachea:1999:HLP

[BFRS99] Dan Bonachea, Kathleen Fisher, Anne Rogers, and Frederick Smith. Hancock: a language for processing very largescale data. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        dsl99/bonachea.html.

Bac:1993:CCC

[BG93] Christian Bac and Edmond Garnier. Cohabitation and cooperation of Chorus and MacOS. In USENIX [USE93k], pages 61–71. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/bac.html.

Bergeman:1997:UTT

[BG97] Ellen R. Bergeman and Mark Graves. Using Tcl/Tk in biology research application development. In USENIX [USE97d], page 167. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_bergeman.html.

Bass:1999:GFI

[BG99] Tim Bass and Dave Gruber.

A glimpse into the future of ID. ;login: the USENIX Association newsletter, 24(5s):??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-9/
features/future.html. Spe-

cial issue on intrusion detection.

Bleichenbacher:1998:SPC

[BGG^{+}98] Daniel Bleichenbacher, Eran Gabber, Phillip B. Gibbons, Yossi Matias, and Alain Mayer. On secure and pseudonymous client-relationships with multiple servers. In USENIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http://www.

usenix.org/publications/library/
proceedings/ec98/bleichenbacher.
html.

Berkley:1990:TU

[BGH^{+}90] Robert Berkley, Skip Gilbrech, Timothy Hunt, Mark Luppi, and Richard Plevin. Tickerplants on UNIX. In USENIX [USE90h], pages 65–74. LCCN QA76.8.U65 U82 1990.

Batlivala:1992:ESC

[BGH^{+}92] Nariman Batlivala, Barry Gleeson, Jim Hamrick, Scott Lurndal, Darren Price, James Soddy, and Vadim Abrossimov. Experience with SVR4 over Chorus. In USENIX [USE92i], pages 223–

242. ISBN 1-880446-42-1. LCCN

QAX 32.

Bellare:1995:IFS

[BGH^{+}95] Mihir Bellare, Juan A. Garay, Ralf Hauser, Amir Herzberg, Hugo Krawczyk, Michael Steiner, Gene Tsudik, and Michael Waidner. iKP — A family of secure electronic payment protocols. In USE-NIX [USE95g], pages 89–

106. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec95/krawczyk.html.

Bellare:1998:VMP

[BGJY98] M. Bellare, J. Garay, C. Jutla, and M. Yung. VarietyCash: a multi-purpose electronic payment system. In USENIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/ec98/bellare.
                                        html.

Bettison:1991:SIU

[BGMR91] Andrew Bettison, Andrew Gollan, Chris Maltby, and Neil Russell. SHARE II — A user administration and resource control system for UNIX. In USE-NIX [USE91f], pages 51–60.

Bruno:1998:EOS

[BGÖS98] John Bruno, Eran Gabber, Banu Özden, and Abraham Silberschatz. The Eclipse operating system: Providing quality of service via reservation domains. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/bruno.html.

Bailey:1994:PPB

[BGP^{+}94] Mary L. Bailey, Burra Gopal, Michael A. Pagels, Larry L. Peterson, and Prasenjit Sarkar. PATHFINDER: a pattern-based packet classifier. In USE-NIX [USE94i], pages 115–

123. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/index.html.

Brace:1998:CNE

[BGT98] Don Brace, Andrew Gordon, and Scott Teel. Compaq’s new engineering compute farm environment: Moving forward with Windows NT. In USENIX [USE98e], pages 55–68. ISBN 1-880446-96-0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/brace.html.

Bond:1991:DDS

[BH91] Andy Bond and John H. Hine. DRUMS: a distributed statistical server for STARS. In USE-NIX [USE91i], pages 335–348. LCCN QA 76.76 O63 U84 1992.

Bederson:1995:APZ

[BH95] Benjamin B. Bederson and James D. Hollan. Advances in the Pad++ zoomable graphics widget. In USENIX [USE95j], pages 31–38 (or 31– 37??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/bederson.html.

Baker-Harvey:1999:ERD

[BH99] Miche Baker-Harvey. ETI resource distributor: Guaranteed resource allocation and scheduling in multimedia systems. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

                                        5. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi99/baker.
                                        html.

Baillie:1992:EKB

[BHJV92] Linda Baillie, Gary W. Hoglund, Lisa Jansen, and Eduardo M. Valcarce. ESSENSE: a knowledge-based security monitoring and control. In USENIX [USE92m], pages 155–170. ISBN 1-880446-46-4. LCCN ????

Bass:1990:SUI

[BHK^{+}90] Len Bass, Eric Hardy, Rick Kazman, Robert Seacord, and Brian Clapper. Serpent: a user interface management system. In USENIX [USE90h], pages 245–

258. LCCN QA76.8.U65 U82 1990.

Blackwell:1995:HCA

[BHS95] Trevor Blackwell, Jeffrey Harris, and Margo Seltzer. Heuristic cleaning algorithms in logstructured file systems. In USENIX [USE95e], pages 277–

288. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/blackwell.html.

Bianchi:1996:TRV

[Bia96a] David J. Bianchi. The terminal room: a volunteer’s view. ;login: the USENIX Association newsletter, 21(6):40–??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Bianco:1996:RBI

[Bia96b] David J. Bianco. Review: Bandits on the information superhighway. ;login: the USE-NIX Association newsletter, 21 (3):56–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Binde:1999:CHA

[Bin99]  Melissa Binde.  Call for

how-to authors and reviewers. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-8/
                                        callforhowto.html.

Birss:1990:BRS

[Bir90] Robert C. Birss. Book review: Software Engineering: Concept and Management. ;login: the USENIX Association newsletter, 15(6):16–??, November/December 1990. CODEN LOGNEM. ISSN 1044-6397.

Birss:1993:RIW

[Bir93] Bob Birss. Review: Inside Windows NT. ;login: the USE-NIX Association newsletter, 18 (3):37–38, May/June 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Birss:1994:RAD

[Bir94] Bob Birss. Review: Application Development for Distributed Environments. ;login: the USENIX Association newsletter, 19(6):51–??, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Birss:1995:BRC

[Bir95] Bob Birss. Book review: The Craft of Software Testing: Subsystem Testing Including Object-Based and Object Oriented Testing. ;login: the USENIX Association newsletter, 20(5):49– ??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Bishop:1990:EPC

[Bis90] Matt Bishop. An extensible password changing program. In USENIX [USE90i], pages 15–16. LCCN QA 76.9 A25 U55 1990.

Bishop:1991:AMU

[Bis91] Matt Bishop. An authentication mechanism for USENET. In USENIX [USE91i], pages 281– 288. LCCN QA 76.76 O63 U84 1992.

Bischofberger:1992:SPA

[Bis92a] Walter R. Bischofberger. Sniff:

a pragmatic approach to a C++ programming environment. In USENIX [USE92k], pages 67– 81. ISBN 1-880446-45-6. LCCN ????

Bishop:1992:APP

[Bis92b] Matt Bishop. Anatomy of a proactive password changer. In USENIX [USE92m], pages 171– 184. ISBN 1-880446-46-4. LCCN ????

Bishop:1992:PEE

[Bis92c] Matt Bishop. Privacy enhanced electronic mail. ;login: the USE-NIX Association newsletter, 17 (6):20–22, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Buzzard:1996:IHS

[BJM^{+}96] Greg Buzzard, David Jacobson, Milon Mackey, Scott Marovich, and John Wilkes. An implementation of the Hamlyn sendermanaged interface architecture. In USENIX [USE96f], pages 245–259. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/

publications/library/proceedings/
osdi96/buzzard.html.

Bentley:1991:SAA

[BK91] J. Bentley and B. Kernigan. A system for algorithm animation. In USENIX [USE91d], pages 5– 30.

Baldwin:1993:DSF

[BK93] Bob Baldwin and Jim Kubon. Dial-in security firewall software. In USENIX [USE93m], pages 59–62. ISBN 1-880446-55-3. LCCN QA 76.9 A25

U54 1993. URL http://www.
usenix.org/publications/library/
proceedings/sec4/.

Bailey:1997:NCB

[BK97] Brian Bailey and Joseph A. Konstan. Nsync — A constraint based toolkit for multimedia. In USENIX [USE97d], pages 107–114. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/bailey.html.

Banerji:1994:FBE

[BKC94] Arindam Banerji, Dinesh Kulkarni, and David Cohn. A framework for building extensible C++ class libraries. In USE-NIX [USE94e], pages 21–36. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/banerji.html.

Birk:1997:ICJ

[BKM97] Yitzhak Birk, Uri Kareev, and Mark Mokryn. IntelliJuke — a caching jukebox-based storage server. In USENIX [USE97f], page 150. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

Baran:1990:SBF

[BKS90] Fuat Baran, Howard Kaye, and Margarita Suarez. Security breaches: Five recent incidents at Columbia University. In USE-NIX [USE90i], pages 151–171. LCCN QA 76.9 A25 U55 1990.

Bala:1994:SPC

[BKW94] Kavita Bala, M. Frans Kaashoek, and William E. Weihl. Software prefetching and caching for translation lookaside buffers. In USENIX [USE94i], pages 243–253. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/bala.html.

Bacon:1990:PRT

[BL90a] David F. Bacon and Andy Lowry. A portable run-time system for the Hermes distributed programming language. In USE-NIX [USE90g], pages 39–50. LCCN QA76.8.U65 U81 1990.

Barkley:1990:DFS

[BL90b] Ronald E. Barkley and T. Paul Lee. A dynamic file system inode allocation and reclaim policy. In USENIX [USE90h], pages 1–10. LCCN QA76.8.U65 U82

Boykin:1990:MCA

[BL90c] Joseph Boykin and Alan Langerman. Mach/4.3BSD: a conservative approach to parallelization. In USENIX [USE90d], pages 69– 100.

Boneh:1996:RBS

[BL96] Dan Boneh and Richard J. Lipton. A revocable backup system. In USENIX [USE96i], pages 91–96. ISBN 1-880446-79-

0. LCCN QA76.9.A25 U83 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec96/boneh.html.

Blumofe:1997:ARP

[BL97] Robert D. Blumofe and Philip A. Lisiecki. Adaptive and reliable parallel computing on networks of workstations. In USENIX [USE97e], pages 133–

147. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ana97/blumofe.html;   http:
                                        //www.usenix.org/publications/
                                        library/proceedings/ana97/
                                        summaries/blumofe.html.

Black:1990:MTF

[Bla90] David L. Black. The Mach timing facility: An implementation of accurate low-overhead usage timing. In USENIX [USE90e], pages 53–72.

Black:1991:PPP

[Bla91] David L. Black. Processors, priority and policy: Mach scheduling for new environment. In USENIX [USE91i], pages 1–12. LCCN QA 76.76 O63 U84 1992.

Blaze:1992:NTP

[Bla92] Matt Blaze. NFS tracing by passive network monitoring. In USENIX [USE92j], pages 333– 344.

Black:1993:KAM

[Bla93]  John Black.  Keynote ad-

dress: a management view of systems administration. In USENIX [USE93f], page ?? ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993. URL

http://www.usenix.org/publications/
library/proceedings/lisa93/

.

Blaze:1994:KME

[Bla94] Matt Blaze. Key management in an encrypting file system. In USENIX [USE94j], pages 27–35. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/blaze.html.

Blackwood:1999:SRL

[Bla99] David Blackwood. Standard report: Is Linux the future of POSIX? ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
26.linux.html.

Blinne:1990:TPL

[Bli90] Richard Blinne. A type parameterization language for C++. In USENIX [USE90j], pages 137– 148.

Bender:1991:NMM

[BLO^{+}91] Walter Bender, Hakon Lie, Jonathan Orwant, Laura Teodosio, and Nathan Abramson. Newspace: Mass media and personal computing. In USE-NIX [USE91g], pages 329–348. LCCN QA 76.76 O63 U83 1991.

Bloch:1992:WS

[Blo92] Lisa A. Bloch. WIP summaries. ;login: the USENIX Association newsletter, 16(5):23–25, March/ April 1992. CODEN LOGNEM. ISSN 1044-6397.

Bellovin:1991:LKA

[BM91a] Steven M. Bellovin and Michael Merritt. Limitations of the Kerberos authentication system. In USENIX [USE91i], pages 253–267. LCCN QA 76.76 O63 U84 1992. URL ftp:

                                        //research.bell-labs.com/
                                        dist/kerblimit.usenix.ps.Z.

FTP - research.att.com:/dist/kerblimit.usenix.ps.Z; local - kerblimit.usenix.ps.

Bolinger:1991:PSH

[BM91b] Don Bolinger and Shashi Mangalat. Parallelizing signal handling and process management in OSF/1. In USE-NIX [USE91h], pages 105–122. LCCN QAX 27.

Bosch:1996:CFI

[BM96] Peter Bosch and Sape J. Mullender. Cut-and-paste filesystems: Integrating simulators and file-systems. In USE-NIX [USE96k], pages 307–318. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/bosch.html.

Baker:1998:DSJ

[BM98a] Brenda S. Baker and Udi Manber. Deducing similarities in Java sources from bytecodes. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/baker.html.

Banga:1998:SKP

[BM98b] Gaurav Banga and Jeffrey C.

Mogul. Scalable kernel performance for Internet servers under realistic loads. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix98/
banga.html.

Brooks:1998:VPW

[BMH98] Charles L. Brooks, Murray S. Mazer, and Frederick J. Hirsch. Visualizing personal Web caches with Caubview. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/brooks.
html.

Banatre:1991:HDR

[BMHR91] Michel Banatre, Gilles Muller, Pack Heng, and Bruno Rochat. How to design reliable servers using fault tolerant microkernel mechanisms. In USE-NIX [USE91h], pages 223–232. LCCN QAX 27.

Bernabeu:1996:ETU

[BMK96] Jose Bernabeu, Vlada Matena, and Yousef Khalidi. Extending a traditional OS using object-oriented techniques. In USENIX [USE96g], pages 53–63. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots96/khalidi.html.

Barnett:1997:WBB

[BMLH97] Tyler Barnett, Kyle McPeek, Larry S. Lile, and Ray Hyatt, Jr. A Web-based backup/ restore method for Intel-based PC’s. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/10.barnett.html.

Brachman:1992:TDL

[BN92] Barry Brachman and Gerald Neufeld. TDBM: a DBM library with atomic transactions. In USENIX [USE92g], pages 63–

80. ISBN 1-880446-44-8. LCCN

QA 76.76 O63 U83 1992.

Bahreman:1996:PMN

[BN96] Alireza Bahreman and Rajkuman Narayanaswamy. Payment method negotiation service: Framework and programming interface. In USENIX [USE96h], pages 299–314. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/bahreman2.html.

Bahn:1999:UFR

[BNMK99] Hyokyung Bahn, Sam H.

Noh, Sang Lyul Min, and Kern Koh. Using full reference history for efficient document replacement in Web caches. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/bahn.html.

Bobenrieth:1999:TS

[Bob99]  Guy Bobenrieth.  A Tcl

story. ;login: the USENIX Association newsletter, 24(1s): ??, January 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-1/

tcl.html. Special issue on tools.

Bogdanovich:1995:OTO

[Bog95] Pedja Bogdanovich. Objective-Tcl: An object-oriented Tcl environment. In USENIX [USE95j], pages 161–172 (or 161–171??). ISBN 1-880446-72-3. LCCN QA76.73.T44

                                        T44 1995. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/tcl95/bogdanovich.
                                        html.

Baquero:1995:ICC

[BOM95] Carlos Baquero, Rui Oliveira, and Francisco Moura. Integration of concurrency control in a language with subtyping and subclassing. In USENIX [USE95k], pages 173–184 (or 173–183??). ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/baquero.html.

Bonwick:1994:SAO

[Bon94] Jeff Bonwick. The slab allocator: An object-caching kernel memory allocator. In USE-NIX [USE94j], pages 87–98. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/bonwick.html.

Borenstein:1992:MMB

[Bor92] Nathaniel S. Borenstein. Multimedia mail from the bottom up -or- teaching dumb mailers to sing. In USENIX [USE92j], pages 79–92.

Borss:1993:RLK

[Bor93] Bob Borss. Review: Learning the Korn Shell. ;login: the USE-NIX Association newsletter, 18 (6):46–47, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Borr:1998:SSU

[Bor98] Andrea J. Borr. SecureShare:

Safe UNIX/Windows file sharing through multiprotocol locking. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/borr.html.

Bostic:1993:R

[Bos93] Keith Bostic. 4.4BSD release.

;login: the USENIX Association newsletter, 18(5):29–32, September/October 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Bossart:1996:HIV

[Bos96]  Pierre-Louis Bossart.  Hy-

pertools in image and volume visualization. In USE-NIX [USE96e], pages 221–230. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/bossart.html.

Bothner:1998:KCD

[Bot98] Per Bothner. Kawa — compiling dynamic languages to the Java VM. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/bothner.
slides.ps.

Bowers:1998:WJA

[Bow98] Neil Bowers. Weblint: Just another Perl hack. In USENIX [USE98d], page ?? ISBN ????

LCCN ???? URL http://www.
usenix.org/publications/library/
proceedings/usenix98/freenix/
bowers.ps.

Box:1996:PDC

[Box96] Don Box. Programming distributed components using network OLE and C++. In USE-NIX [USE96g], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

usenix.org/publications/library/
proceedings/coots96/.

Boyer:1991:CDS

[Boy91] Fabienne Boyer. A causal distributed shared memory based on external pagers. In USENIX [USE91h], pages 41–58. LCCN QAX 27.

Bhagwat:1993:MNS

[BP93] Pravin Bhagwat and Charles E. Perkins. A mobile networking system based on Internet protocol (IP). In USE-NIX [USE93i], pages 69–82. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/bhagwat.html.

Barnes:1999:PDC

[BP99] J. Fritz Barnes and Raju Pandey. Providing dynamic and customizable caching policies. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/barnes.html.

Budge:1992:HPS

[BPR92] K. G. Budge, J. S. Perry, and A. C. Robinson. Highperformance scientific computing using C++. In USENIX [USE92k], pages 131–150. ISBN 1-880446-45-6. LCCN ????

Belaid:1998:DMC

[BPT98] Djamel Belaid, Nicolas Provenzano, and Chantal Taconet. Dynamic management of CORBA Trader Federation. In USENIX [USE98c], page ?? ISBN ????

LCCN ???? URL http://www.
usenix.org/publications/library/
proceedings/coots98/belaid.
html.

Baumgartner:1994:ISC

[BR94] Gerald Baumgartner and Vince Russo. Implementing signatures for C++. In USE-NIX [USE94e], pages 37–56. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/baumgartner.html.

Bushnell:1998:AAW

[BR98] Thomas Bushnell and Karl Ramm. Anatomy of an Athena Workstation. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
bushnell.html.

Braunsdorf:1990:KMS

[Bra90] Kevin Braunsdorf. Keeping up with the manual system. In USENIX [USE90f], pages 71–80. LCCN QA76.76.O63 L37 1990.

Braunsdorf:1991:KMS

[Bra91] Kevin Braunsdorf. Keeping up with the manual system. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Bradley:1992:OBF

[Bra92a] Mitch Bradley. Open boot firmware. In USENIX [USE92j], pages 223–236.

Branagan:1992:BRG

[Bra92b] Linda Branagan. Book review: GNU Emacs. ;login: the USENIX Association newsletter, 17(2):34–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397.

Brassil:1999:EVE

[Bra99] John J. Brassil. The evolution of VLAN/ELAN Architecture at Vanderbilt University, 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neta99/ITs/brassil_html/.

Unpublished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Brecht:1993:IPA

[Bre93] Tim Brecht. On the importance of parallel application placement in NUMA multiprocessors. In USENIX [USE93j], pages 1–18. ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/brecht.html.

Bhoedjang:1993:PPP

[BRH^{+}93] Raoul Bhoedjang, Tim Ruhl, Rutger Hofman, Koen Langendoen, Henri Bal, and Frans Kaashoek. Panda: a portable platform to support parallel programming languages. In USENIX [USE93j], pages 213–

226. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/bhoedjang.html.

Brundrett:1997:MSW

[Bru97] Peter Brundrett. MS/Security in Windows NT 5.0. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/its/brundrett.

it. Invited talk.

Brundrett:1998:KAW

[Bru98] Peter Brundrett. Kerberos authentication in Windows NT 5.0 domains. ;login: the USE-NIX Association newsletter, 23 (3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-5/
brundrett.html. Special issue

on security.

Brumley:1999:IIR

[Bru99a] David Brumley. Invisible intruders: Rootkits in practice. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-9/
                                        features/rootkits.html. Spe-

cial issue on intrusion detection.

Brumley:1999:THI

[Bru99b] David Brumley. Tracking hackers on IRC. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-11/
                                        features/hackers.html. Spe-

cial issue on security.

Bentley:1999:SOE

[BRW99] Damian Bentley, Greg Rose, and Tara Whalen. ssmail: Opportunistic encryption in sendmail. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa99/bentley.
                                        html.

Brown:1990:CVP

[BS90] David S. Brown and E. Eugene Schultz, Jr. Communicating vulnerabilities: Perils and pitfalls. In USENIX [USE90i], pages 149–150. LCCN QA 76.9 A25 U55 1990.

Baker:1992:RBU

[BS92a] Mary Baker and Mark Sullivan.

The recovery box: Using fast recovery to provide high availability in the UNIX environment. In USENIX [USE92g], pages 31–

44. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Bhide:1992:HAL

[BS92b] Anupam Bhide and Spencer Shepler. A highly available lock manager for HA-NFS. In USE-NIX [USE92g], pages 177–184. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Burkhard:1992:SER

[BS92c] Walt Burkhard and Petar D.

Stojadinovic. Storage efficient reliable files. In USENIX [USE92j], pages 69–78.

Bolosky:1993:FSE

[BS93a] William J. Bolosky and Michael L.

Scott. False sharing and its effect on shared memory performance. In USENIX [USE93j], pages 57–71. ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/bolosky.html.

Breitbart:1993:PET

[BS93b] Y. Breitbart and A. Silberschatz. Performance evaluation of two multidatabase transaction management algorithms. In USENIX [USE93c], pages 245–

Brustoloni:1996:EBS

[BS96] Jose Carlos Brustoloni and Peter Steenkiste. Effects of buffering semantics on I/O performance. In USE-NIX [USE96f], pages 277–

291. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/brusto.html.

Brown:1997:RTS

[BS97a] David S. Brown and James M. Sharp. Restoring trust: The sysadmin’s Prime responsibility after a break-in. ;login: the USENIX Association newsletter, 22(1):17–18, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Burgess:1997:ALF

[BS97b] Mark Burgess and Demosthenes Skipitaris. Adaptive locks for frequently scheduled tasks with unpredictable runtimes. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/15.burgess.html.

Burgess:1998:MFA

[BS98] Mark Burgess and Demosthenes Skipitaris. Managing filesystem ACLs with GNU/ Cfengine. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
acls.html.

Ben-Shaul:1993:AMU

[BSKH93] Israel Z. Ben-Shaul, Gail E.

Kaiser, and George T. Heineman. An architecture for multi-user software development environments. In USENIX [USE93b], pages 65–103.

Badger:1995:DTE

[BSS^{+}95] Lee Badger, Daniel F. Sterne, David L. Sherman, Kenneth M. Walker, and Sheila A. Haghighat. A domain and type enforcement UNIX prototype. In USENIX [USE95f], pages 127–140. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
badger.html.

Black:1996:FLL

[BSSD96] David L. Black, Randall D.

Smith, Steven J. Sears, and Randall W. Dean. FLIPC: a low latency messaging system for distributed real time environments. In USENIX [USE96k], pages 229–238. ISBN 1-880446-76-6. LCCN QA 76.76 O63

U88 1996. URL http://www.
usenix.org/publications/library/
proceedings/sd96/black.html.

Badger:1996:DTE

[BSSW96] Lee Badger, Daniel F. Sterne, David L. Sherman, and Kenneth M. Walker. A domain and type enforcement USIX prototype. In USENIX [USE96d], pages 47–83.

Banerji:1997:PSL

[BTC97] Arindam Banerji, John M. Tracey, and David L. Cohn. Protected shared libraries — a new approach to modularity and sharing. In USE-NIX [USE97e], pages 59–

75. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/banerji.html. Best stu-

dent paper award.

Bickerstaff:1999:BJT

[BTS^{+}99] Cindy Bickerstaff, Ken True, Charles Smothers, Tod Oace, Jeff Sedayao, and Clinton Wong. Don’t just talk about the weather — manage it! A system for measuring, monitoring, and managing Internet performance and connectivity. In USENIX [USE99g], page ?? ISBN 1-880446-36-

7. LCCN TK5105.5 .C664 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/bickerstaff.html.

Bugaj:1996:PRH

[Bug96] Stephan Vladimir Bugaj. Passwords for real humans. ;login: the USENIX Association newsletter, 21(3):41–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Burgess:1995:SCE

[Bur95] Mark Burgess. A site configuration engine. In USENIX [USE95c], pages 309–337.

Burgess:1998:CI

[Bur98]  Mark Burgess.  Computer

immunology. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63

S97 1998. URL http://www.
usenix.org/publications/library/
proceedings/lisa98/burgess.
html.

Burgess:1999:MNSa

[Bur99a] Mark Burgess. Managing network security with Cfengine

— part 1. ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
features/cfengine1.html.

Burgess:1999:MNSb

[Bur99b] Mark Burgess. Managing network security with Cfengine, part 2. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/cfengine_p2.html.

Burgess:1999:MNSc

[Bur99c] Mark Burgess. Managing network security with Cfengine, part 3. ;login: the USENIX Association newsletter, 24(6):

??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/cfengine3.html.

Belani:1998:CWA

[BVAD98] Eshwar Belani, Amin Vahdat, Thomas Anderson, and Michael Dahlin. The CRI-SIS wide area security architecture. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/belani.html.

Birman:1997:MEC

[BVG^{+}97] K. Birman, W. Vogels, K. Guo, M. Hayden, T. Hickey, R. Friedman, R. van Renesse, A. Vaysburd, and S. Maffeis. Moving the ensemble communication system to NT and Wolfpack. In USENIX [USE97f], pages 107–112. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/birman.html.

Bulterman:1991:STD

[Bvv91] Dick C. A. Bulterman, Guido van Rossum, and Robert van Liere. A structure for transportable, dynamic multimedia documents. In USE-NIX [USE91g], pages 137–156. LCCN QA 76.76 O63 U83 1991.

Byrne:1993:EAP

[Byr93]  Sheri Byrne.  Establishing

and administering a public access Internet facility. In USENIX [USE93f], pages 227–

235. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/byrne.html.

Baker:1996:SMM

[BZCS96] Mary Baker, Xinhua Zhao, Stuart Cheshire, and Jonathan Stone. Supporting mobility in MosquitoNet. In USE-NIX [USE96k], pages 127–139. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/baker.html.

Bhargava:1991:ECS

[BZM91] Bharat Bhargava, Yongguang Zhang, and Enrique Mafla. Evolution of a communication system for distributed transaction processing in raid. In USENIX [USE91c], pages 277–314.

Bettati:1999:RTI

[BZT99] R. Bettati, W. Zhao, and D. Teodor. Real-time intrusion detection and suppression in ATM networks. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999. URL

http://www.usenix.org/publications/
library/proceedings/detection99/
bettati.html.

Clark:1996:BCT

[CA96] Dayton Clark and David M.

Arnow. Backtracking and constraints in Tcl-BC. In USE-NIX [USE96e], pages 49–59. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/clark.html.

Caetta:1997:DIS

[Cae97]  Jennifer Caetta.  Develop-

ing interim systems. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/lisa97/
                                        07.caetta.html.

Chutani:1992:EFS

[CAK^{+}92] Sailesh Chutani, Owen T. Anderson, Michael L. Kazar, Bruce W. Leverett, W. Anthony Mason, and Robert N. Sidebotham. The Episode file system. In USENIX [USE92j], pages 43–60.

Calabrese:1993:ATC

[Cal93] Christopher J. Calabrese. An advanced tape cataloging system for UNIX systems. In USENIX [USE93l], pages 289–

293. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Calabrese:1996:TBF

[Cal96] Christopher J. Calabrese. A tool for building firewall-router configurations. In USENIX [USE96c], pages 239–253.

Calabrese:1998:LEU

[Cal98] Christopher J. Calabrese. Letter to the Editor: Unique UIDs. ;login: the USENIX Association newsletter, 23(4):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Calabrese:1999:TB

[Cal99] Christopher Calabrese. The trouble with biometrics. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-8/features/biometrics.
html.

Campbell:1993:WHL

[Cam93]  Roy H. Campbell.  What-

ever happened to large packets or are tiny messages good? In USENIX [USE93j], pages 279–281. ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

Campbell:1994:CEE

[Cam94] Stephen Campbell. Campus email for everyone: Making it work in real life. In USENIX [USE94m], pages 23–38. ISBN ???? LCCN ????

Cargill:1990:DCR

[Car90a] T. A. Cargill. Does C++ really need multiple inheritance? In USENIX [USE90j], pages 315–

Carlin:1990:GSM

[Car90b] Jerry Carlin. Gateway security measures. In USENIX [USE90i], pages 145–148. LCCN QA 76.9 A25 U55 1990.

Cargill:1991:CCA

[Car91a] T. A. Cargill. Controversy: The case against multiple inheritance in C++. In USENIX [USE91d], pages 69–82.

Carroll:1991:PNI

[Car91b]  Martin Carroll.  Problems

with non-invasive inheritance in C++. In USENIX [USE91k], pages 13–28. LCCN QA76.73.C153 U27 1991.

Carlitz:1992:R

[Car92a] Robert D. Carlitz. Reflections. ;login: the USENIX Association newsletter, 17(3):16–??, May/ June 1992. CODEN LOGNEM. ISSN 1044-6397.

Carroll:1992:LE

[Car92b] Martin D. Carroll. Not a language extension. In USENIX [USE92k], pages 21–36. ISBN 1-880446-45-6. LCCN ????

Cargill:1993:CAM

[Car93a] T. A. Cargill. The case against multiple inheritance in C++. In Waldo [Wal93b], pages 101–109. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Carlin:1993:USU

[Car93b] Jerry M. Carlin. UNIX security update. In USE-NIX [USE93m], pages 119–

130. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/.

Carr:1993:PCI

[Car93c] Robert Carr. Pen-based computing and its impact. In USENIX [USE93l], page ?? ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/

sd93/. Keynote address.

Carson:1993:SS

[Car93d] Mark E. Carson. Sendmail without the superuser. In USE-NIX [USE93m], pages 139–

144. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/carson.html.

Carroll:1994:FPP

[Car94] Martin Carroll. Faster parsing via prefix analysis. In USE-NIX [USE94e], pages 73–84. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/index.html.

Cardelli:1995:LDS

[Car95] Luca Cardelli. A language with distributed scope. In USENIX [USE95d], pages 27–59.

Carberry:1998:LER

[Car98a] Mick Carberry. Letter to the editor: To Rik Farrow. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Carter:1998:PSA

[Car98b] Gerald Carter. Patch 32: a system for automated client OS updates. In USENIX [USE98e], pages 1–10. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/carter.html.

Carpenter:1999:WBC

[Car99a] Jeffrey J. Carpenter. Welcome to the Big City: Incident reporting helps the CERT Coordination Center keep pace with a rapidly expanding Internet. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-11/
features/cert.html. Special

issue on security.

Carter:1999:AWN

[Car99b] Gerald Carter. Administering Windows NT domains using a non-Windows NT PDC. In USENIX [USE99b], pages 43–

50. ISBN 1-880446-30-8. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa-nt99/carter.
html.

Casella:1995:SAO

[Cas95] Karen A. Casella. Security administration in an open networking environment. In USE-NIX [USE95h], pages 67–74 (or 67–73??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/casella.html.

Casella:1996:EIDb

[Cas96a] Karen Casella. Elementary intrusion detection, part 2. ;login: the USENIX Association newsletter, 21(3):13–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Casella:1996:EIDa

[Cas96b] Karen Casella. Elementary intrusion detection, part I. ;login: the USENIX Association newsletter, 21(2):27–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Cattey:1990:ETC

[Cat90] William Cattey. The evolution of Turnin: a classroom oriented file exchange service. In USE-NIX [USE90g], pages 171–182. LCCN QA76.8.U65 U81 1990.

Cate:1992:AGF

[Cat92] Vincent Cate. Alex — A global filesystem. In USE-NIX [USE92h], pages 1–12. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Chin-A-Young:1999:YTN

[CAY99] Albert Chin-A-Young. Ypmake: a tool for NIS. ;login: the USE-NIX Association newsletter, 24 (5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/ypmake.html.

Caswell:1990:IMD

[CB90] Deborah Caswell and David Black. Implementing a Mach debugger for multithreaded applications. In USENIX [USE90h], pages 25–40. LCCN QA76.8.U65 U82 1990.

Cheswick:1996:DFS

[CB96] Bill Cheswick and Steven M. Bellovin. A DNS filter and switch for packet-filtering gateways. In USENIX [USE96i], pages 15–19. ISBN 1-880446-79-

0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/cheswick.html.

Campbell:1991:PUS

[CBB^{+}91] Mark Campbell, Richard Barton, Jim Browning, Dennis Cervenka, Ben Curry, Tod Davis, Tracy Edmonds, Russ Holt, John Slice, Tucker Smith, and Rich Wescott. The parallelization of UNIX System V Release 4.0. In USENIX [USE91i], pages 307–324. LCCN QA 76.76 O63 U84 1992.

Christenson:1997:SNA

[CBB97a] Nick Christenson, David Beckemeyer, and Trent Baker. A scalable news architecture on a single spool. ;login: the USE-NIX Association newsletter, 22 (5):41–45, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Christenson:1997:HSE

[CBB97b] Nick Christenson, Tim Bosserman, and David Beckemeyer. A highly scalable electronic mail service using open systems. In USENIX [USE97h], pages 37–

47. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/christenson.html.

Chen:1993:ICM

[CBE93] Jiansan Chen, Omran A.

Bukhres, and Ahmed K. Elmagarmid. The implementation of cooperative mechanisms among system components in a heterogeneous multidatabase environment. In USENIX [USE93c], pages 207–243.

Chen:1990:IIV

[CBL90] Danny Chen, Ronald E. Barkley, and T. Paul Lee. Insuring improved VM performance: Some no-fault policies. In USENIX [USE90h], pages 11–24. LCCN QA76.8.U65 U82 1990.

Chiueh:1999:RSO

[cC99] Tzi cker Chiueh. RETHER:

a software-only real-time Ethernet for PLC networks. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999. URL

http://www.usenix.org/publications/
library/proceedings/es99/
chiueh.html.

Cohn:1992:UDP

[CCGS92] David L. Cohn, Michael R. Casey, Paul M. Greenawalt, and John E. Saldanha. A universal distributed programming paradigm for multiple operating systems. In USENIX [USE92l], pages 191–204. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Cheng:1991:OEE

[CCK^{+}91] Edward C. Cheng, Edward Chang, Johannes Klein, Dora Lee, Edward Lu, Alberto Lutgardo, and Ron Obermarck. An open and extensible event-based transaction manager. In USE-NIX [USE91g], pages 49–58. LCCN QA 76.76 O63 U83 1991.

Casas:1995:MMT

[CCK^{+}95] Jeremy Casas, Dan L. Clark, Ravi Konuru, Steve W. Otto, Robert M. Prouty, and Jonathan Walpole. MPVM: a migration transparent version of PVM. In USENIX [USE95b], pages 171– 216.

Cohen:1998:APT

[CCK98] Geoff A. Cohen, Jeffrey S. Chase, and David L. Kaminsky. Automatic program transformation with JOIE. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix98/
cohen.html.

Carson:1990:NUS

[CCM^{+}90] Mark Carson, Janet Cugini, Sohail Malik, Mythili Kannan, and Wen-Der Jiang. Networked UNIX without the superuser. In USENIX [USE90i], pages 137–

140. LCCN QA 76.9 A25 U55 1990.

Chan:1997:ILN

[CCW97] Andrea Chan, Freddie Chow, and Jonathan Wong. Implementing large NT networks in a heterogeneous computing environment at SLAC. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/chan.html.

Chen:1992:ICC

[CD92] Raymond C. Chen and Partha Dasgupta. Integrating consistency control and distributed shared memory: The travails of an implementation. In USE-NIX [USE92l], pages 279–296. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Castelluccia:1994:MCS

[CD94a] Claude Castelluccia and Walid Dabbous. Modular communication subsystem implementation using A synchronous approach. In USENIX [USE94f], pages 1–11. ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        hispeed/castelluccia.html.

Cheriton:1994:CMO

[CD94b] David R. Cheriton and Kenneth J. Duda. A caching model of operating system kernel functionality. In USE-NIX [USE94i], pages 179–

193. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/cheriton.html.

Cardelli:1997:SCW

[CD97] Luca Cardelli and Rowan Davies. Service combinators for Web computing. In USE-NIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/dsl97/cardelli.
                                        html.

Curtin:1998:BFS

[CD98] Matt Curtin and Justin Dolske. A brute force search of DES keyspace. ;login: the USE-NIX Association newsletter, 23 (3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-5/

curtin.html. Special issue on

security.

Cox:1999:NMT

[CD99] Phil Cox and Tina Darmohray. NT: (the manual is) ‘not there’! ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
nt.html.

Conran:1998:TCB

[CDGT98] David Conran, Barb Dijker, Tim Gassaway, and Kim Trudel. Taming the certification beast. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Chen:1999:PUL

[CDK^{+}99] Yuqun Chen, Stefanos N. Damianakis, Sanjeev Kumar, Xiang Yu, and Kai Li. Porting a user-level communication architecture to NT: Experiences and performance. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/chen.html.

Chi:1999:CPS

[CDL99] Chi-Hung Chi, Jing Deng, and Yan-Hong Lim. Compression proxy server: Design and implementation. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96. URL

http://www.usenix.org/publications/
library/proceedings/usits99/
chi.html.

Comer:1990:EIT

[CDM90] Douglas Comer, Ralph E.

Droms, and Thomas P. Murtagh.

An experimental implementation of the tilde naming system. In USENIX [USE90a], pages 487–516.

Chankhunthod:1996:HIO

[CDN^{+}96] Anawat Chankhunthod, Peter B. Danzig, Chuck Neerdaels, Michael F. Schwartz, and Kurt J. Worrell. A hierarchical Internet object cache. In USENIX [USE96k], pages 153–163 (or 153–164??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/danzig.html.

Cranor:1999:OSR

[CdR99] Charles D. Cranor and Theo de Raadt. Opening the source repository with anonymous CVS. In USENIX [USE99m], page ?? ISBN 1-880446-33-

2. LCCN ???? URL http:

//www.openbsd.org/papers/
anoncvs-paper.ps.

Chandra:1999:JCM

[CE99] Surendar Chandra and Carla Schlatter Ellis. JPEG compression metric as a qualityaware image transcoding. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96. URL

http://www.usenix.org/publications/
library/proceedings/usits99/
chandra.html.

Chinoy:1994:TIH

[CF94] Bilal Chinoy and Kevin Fall. TCP/IP and HIPPI performance in the CASA gigabit testbed. In USENIX [USE94f], pages 45–59. ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
hispeed/chinoy.html.

Coe:1995:DDC

[CF95] Diane E. Coe and Judith A.

Furlong. Developing and deploying corporate cryptographic systems. In USE-NIX [USE95g], pages 137–

146. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/coe.html.

Christenson:1998:TOI

[CF98] Nick Christenson and Dan Farmer. From the trenches: One ISP’s response to the problem of spam. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
trenches.html.

Chevalier:1993:ESO

[CFH^{+}93] P. Y. Chevalier, A. Freyssinet, D. Hagimont, S. Krakowiak, S. Lacourte, and X. Rousset de Pina. Experience with shared object support in the Guide system. In USENIX [USE93j], pages 157–

173. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/

                                        publications/library/proceedings/
                                        sedms4/chevalier.html.

Crovella:1999:CSW

[CFHB99] Mark E. Crovella, Robert Frangioso, and Mor Harchol-Balter. Connection scheduling in Web servers. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/crovella.html.

Cao:1994:ACF

[CFL94a] Pei Cao, Edward W. Felten, and Kai Li. Applicationcontrolled file caching policies. In USENIX [USE94j], pages 171–182. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/cao.html.

Cao:1994:IPA

[CFL94b] Pei Cao, Edward W. Felten, and Kai Li. Implementation and performance of applicationcontrolled file caching. In USENIX [USE94i], pages 165–

177. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/cao.html.

Cameron:1992:PIC

[CFLM92] Don Cameron, Paul Faust, Dmitry Lenkov, and Michey Mehta. A portable implementation of C++ exception handling. In USENIX [USE92k], pages 225–243. ISBN 1-880446-45-6. LCCN ????

Cordrey:1999:MLF

[CFSW99] Vincent Cordrey, Doug Freyburger, Jordan Schwartz, and Liza Weissler. Moving large filesystems on-line, including exiting HSM filesystems. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/cordrey.
html.

Comer:1990:NDD

[CG90] Douglas Comer and James Griffioen. A new design for distributed systems: The remote memory model. In USE-NIX [USE90g], pages 127–136. LCCN QA76.8.U65 U81 1990.

Comer:1991:VMX

[CG91] Douglas Comer and James Griffioen. Virtual memory Xinu. In USENIX [USE91j], pages 261– 270. ISBN ???? LCCN ????

Comer:1992:EOD

[CG92a] Douglas E. Comer and James N.

Griffioen. Efficient orderdependent communication in a distributed virtual memory environment. In USE-NIX [USE92l], pages 249–262. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Crowther:1992:TIO

[CG92b] Nancy Crowther and Joyce Graham. TCP/IP and OSI interoperability with the X Window System. In USENIX [USE92g], pages 243–253. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Chang:1999:AHG

[CG99a] Fay Chang and Garth A. Gibson. Automatic I/O hint generation through speculative execution. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/chang.
html.

Couch:1999:EDW

[CG99b] Alva L. Couch and Michael Gilfix. It’s elementary, dear Watson: Applying logic programming to convergent system management processes. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/couch.
html.

Cranor:2005:SUD

[CG05] Lorrie Faith Cranor and Simson Garfinkel. Security and usability: designing secure systems that people can use. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00827-9 (paperback). xviii + 714 pp. LCCN QA76.9.A25; QA76.9.A25 S3533 2005; QA76.9.A25 S43 2005eb; QA76.9.A25 S43 2005. US$44.95, CDN$62.95, UK£31.95. URL

http://www.oreilly.com/catalog/
securityusability/.

Cohn:1991:UKL

[CGCS91] David L. Cohn, Paul M.

Greenawalt, Michael R. Casey, and Matthew P. Stevenson. Using kernel-level support for distributed shared data. In USE-NIX [USE91j], pages 247–260. ISBN ???? LCCN ????

Cheng:1995:DIM

[CGHK95] Pau-Chen Cheng, Juan A.

Garay, Amir Herzberg, and Hugo Krawczyk. Design and implementation of modular key management protocol and IP secure tunnel on AIX. In USENIX [USE95f], pages 41–54. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
cheng.html.

Capobianchi:1992:LIA

[CGLR92] R. Capobianchi, R. Guerraoui, Agnes Lanusse, and Pierre Roux. Lessons from implementing active objects on a parallel machine. In USENIX [USE92l], pages 13–26. ISBN 1-880446-41-

3. LCCN QA76.9.D3 S954 1992.

Cohn:1997:SOA

[CGLR97] Robert Cohn, David Goodwin, P. Geoffrey Lowney, and Norman Rubin. Spike: An optimizer for Alpha/NT executables. In USENIX [USE97f], pages 17–

23. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/cohn.html.

Cargill:1991:HUM

[CGMV91] Tom Cargill, Keith Gorlen, Rob Murray, and Mike Vilot. How useful is multiple inheritance in C++? In USENIX [USE91k], pages 81–84. LCCN QA76.73.C153 U27 1991. Panel: Moderator — Andrew Koenig.

Costa:1996:LLL

[CGS^{+}96] Manuel Costa, Paulo Guedes, Manuel Sequeira, Nuno Neves, and Miguel Castro. Lightweight logging for lazy release consistent distributed shared memory. In USENIX [USE96f], pages 59–73. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/costa.html.

Chow:1995:NRP

[CH95] Chee-Seng Chow and Amir Herzberg. Network randomization protocol: a proactive pseudo-random generator. In USENIX [USE95f], pages 55–64 (or 55–63??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
chow.html.

Chernoff:1997:DFR

[CH97] Anton Chernoff and Ray Hookway. DIGITAL FX!32 — running 32-bit x86 applications on Alpha NT. In USENIX [USE97f], pages 9–

13. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/chernoff.html.

Chapman:1992:BMI

[Cha92a]  Brent Chapman.  Building

and managing the Interop ’91 Shownet. ;login: the USENIX Association newsletter, 17(2): 28–??, March/April 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Chapman:1992:MHM

[Cha92b] D. Brent Chapman. Majordomo: How I manage 17 mailing lists without answering “- request” mail. In USENIX [USE92e], pages 135–143. ISBN 1-880446-47-2. LCCN ????

Chapman:1992:NST

[Cha92c] D. Brent Chapman. Network (in)security through IP packet filtering. In USENIX [USE92m], pages 63–76. ISBN 1-880446-46-

4. LCCN ????

Chastain:1994:TAL

[Cha94] Michael Chastain. Taking analogies to the limit. ;login: the USENIX Association newsletter, 19(6):36–??, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Chapman:1996:CCW

[Cha96] Brent Chapman. The computer consultant’s workbookreview. ;login: the USENIX Association newsletter, 21(4):35– ??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Chapman:1998:ZLO

[Cha98] Brent Chapman. Zero to LISA in one year, 1998. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6–11, 1998 Boston, Massachusetts, USA.

Chapman:1999:BVL

[Cha99]  Brent Chapman.  Band-

width versus latency: Helping developers understand network performance limitations. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-4/
                                        bandwidth.html.

Cheswick:1990:DSI

[Che90] Bill Cheswick. The design of a secure Internet gateway. In USENIX [USE90g], pages 233– 237. LCCN QA76.8.U65 U81 1990.

Cheswick:1992:EBW

[Che92] Bill Cheswick. An evening with Berferd in which a cracker is lured, endured, and studied. In USENIX [USE92j], pages 163– 174.

Chen:1994:MBX

[Che94a]  J. Bradley Chen.  Mem-

ory behavior for an X11 Window System. In USE-NIX [USE94l], pages 189–200. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/chen.html.

Chen:1994:DTG

[Che94b] Yih-Farn (Robin) Chen. Dagger: a tool to generate program graphs. In USE-NIX [USE94g], pages 19–

35. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/chen.html.

Cheswick:1997:CCS

[Che97] Bill Cheswick. Cget, Cput, and stage-safe file transport tools for the Internet. In USENIX [USE97e], pages 185–

194. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/cheswick.html;   http:
//www.usenix.org/publications/
library/proceedings/ana97/
summaries/cheswick.html.

Chessin:1998:LEP

[Che98a] Steve Chessin. Letter to the Editor: Proportional representation? ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Cheswick:1998:KAS

[Che98b] Bill Cheswick. Keynote address: Security lessons from all over, 1998. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/sec98/technical.

html. Unpublished talk at the 7th USENIX Security Symposium, 1998, January 26–29, 1998, San Antonio, Texas, USA.

Chesson:19xx:DAB

[Chexx] Greg Chesson. Dogzilla: a battle computer for dogfight. In ????, page ?? USENIX, Berkeley, CA, USA, 19xx. Title listed in contents only, no paper, no abstract.

Chen:1999:CCS

[CHHL99] Michael Chen, Marti Hearst, Jason Hong, and James Lin. Cha-Cha: a system for organizing intranet search results. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/chen.html.

Chin:1997:BRW

[Chi97]  Eric Chin.  Book review:

Writing GNU Emacs Extensions. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1997-12/
                                        emacs.html.

Chalup:1998:DFW

[CHK^{+}98] Strata Rose Chalup, Christine Hogan, Greg Kulosa, Bryan McDonald, and Bryan Stansell. Drinking from the fire(walls) hose: Another approach to very large mailing lists. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/chalup.html.

Cho:1998:FAQ

[Cho98]  Kenjiro Cho.  A frame-

work for alternate queueing: Towards traffic management by PC-UNIX based routers. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/cho.html.

Choo:1999:VVC

[Cho99] Tse-Huong Choo. Vaulted VPN:

Compartmented virtual private networks on trusted operating systems. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/choo.html.

Christiansen:1990:AAM

[Chr90a] Tom Christiansen. The answer to all man’s problems. In USENIX [USE90f], pages 81–88. LCCN QA76.76.O63 L37 1990.

Christoph:1990:SCG

[Chr90b] Gary Christoph. Security considerations of going to a UNIX based supercomputer operating system. In USENIX [USE90i], pages 129–130. LCCN QA 76.9 A25 U55 1990.

Christiansen:1991:AAM

[Chr91] Tom Christiansen. The answer to all man’s problems. In USE-NIX [USE91e], page ?? ISBN ???? LCCN ????

Christiansen:1992:AP

[Chr92] Tom Christiansen. Autoloading in Perl. ;login: the USENIX Association newsletter, 17(5):25– 27, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Christiansen:1993:PO

[Chr93a] Tom Christiansen. Perl 5.0 overview. ;login: the USE-NIX Association newsletter, 18 (6):30–32, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Christiansen:1993:SIL

[Chr93b] Tom Christiansen. SANS II & LISA — what’s the difference? ;login: the USENIX Association newsletter, 18(2):20–??, March/ April 1993. CODEN LOGNEM. ISSN 1044-6397.

Christenson:1997:BRA

[Chr97a] Nick Christenson. Book review: Administering Usenet News Servers. ;login: the USE-NIX Association newsletter, 22 (7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
mcdermott.html.

Christoffel:1997:BTS

[Chr97b]  Jurgen Christoffel.  Bal —

A tool to synchronize document collections between computers. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/12.christoffel.html.

Christenson:1998:BRA

[Chr98a] Nick Christenson. Book review: Applying RCS and SCCS. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
rcs.html.

Christenson:1998:BRL

[Chr98b] Nick Christenson. Book review:

At Large: The Strange Case of the World’s Biggest Internet Invasion. ;login: the USE-NIX Association newsletter, 23 (4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
freedman.html.

Christenson:1998:BRCa

[Chr98c] Nick Christenson. Book review: Cisco TCP/IP Routing Professional Reference.

;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-2/
                                        reference.html.

Christenson:1998:BRCb

[Chr98d] Nick Christenson. Book review: Configuration and Capacity Planning for Solaris Servers. ;login: the USE-NIX Association newsletter, 23 (4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-6/
                                        wong.html.

Christenson:1998:BRMa

[Chr98e] Nick Christenson. Book review: Managing IP Networks with Cisco Routers. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Christenson:1998:BRMb

[Chr98f]  Nick Christenson.  Book

review: Managing Usenet. ;login: the USENIX Association newsletter, 23(4): ??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-6/
                                        spencer.html.

Christenson:1998:BRP

[Chr98g] Nick Christenson. Book review: Protecting Networks with SATAN. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
satan.html.

Campbell:1991:LGT

[CHS91a] Mark D. Campbell, Russ Holt, and John Slice. Lock granularity tuning mechanisms in SVR4/ MP. In USENIX [USE91j], pages 221–228. ISBN ???? LCCN ????

Cohen:1991:PII

[CHS91b] Bruce Cohen, Douglas Hahn, and Neil Soiffer. Pragmatic issues in the implementation of flexible libraries in C++. In USENIX [USE91k], pages 193–

202. LCCN QA76.73.C153 U27 1991. See [HS91a].

Corbridge:1991:PFI

[CHS91c] Bruce Corbridge, Robrt Henig, and Charles Slater. Packet filtering in an IP router. In USENIX [USE91f], pages 227–232.

Camp:1996:AAT

[CHTY96] Jean Camp, Michael Harkavy, J. D. Tygar, and Bennet Yee. Anonymous atomic transactions. In USENIX [USE96h], pages 123–133. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Chu:1996:ZCT

[Chu96] Hsiao-Keng Jerry Chu. Zerocopy TCP in Solaris. In USENIX [USE96k], pages 253–

264. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/chu.html.

Chubb:1999:JLA

[Chu99] Lucy Chubb. The John Lions Award. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
award.html.

Chung:1998:CMT

[CHW98] P. Emerald Chung, Yennun Huang, and Yi-Min Wang. COM on a multicast transport. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/full_papers/poster_
chung/chung.pdf;   http://
www.usenix.org/publications/
library/proceedings/usenix-
nt98/full_papers/poster_chung/
chung_html/chung.html.

Campbell:1993:TDF

[CI93] Roy H. Campbell and Nayeem Islam. A technique for documenting the framework of an object-oriented system. In USE-NIX [USE93a], pages 363–389.

Cao:1997:CAW

[CI97] Pei Cao and Sandy Irani. Costaware WWW proxy caching algorithms. In USENIX [USE97h], pages 193–206. ISBN 1-880446-91-X. LCCN TK5105.875.I57

U96 1997. URL http://www.
usenix.org/publications/library/
proceedings/usits97/cao.html.

Clear:1994:EHD

[CIC94] David Clear, Alan Ibbetson, and Peter Collinson. Exporting home directories on demand to PCs. In USE-NIX [USE94h], pages 139–

147. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/clear.html.

Campbell:1992:CFR

[CIM92] Roy H. Campbell, Nayeem Islam, and Peter Madany. Choices, frameworks and refinement. In USENIX [USE92c], pages 217–257.

Coq:1992:SUF

[CJ92] Philippe Coq and Sylvie Jean.

SysView: a user-friendly environment for administration of distributed UNIX systems. In USENIX [USE92e], pages 145–

151. ISBN 1-880446-47-2. LCCN ????

Crone:1995:TIA

[CJ95] Frank Crone and John Jarocki.

TCP/IP addressing strategies for switched networks. ;login: the USENIX Association newsletter, 20(1):27–37, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Candea:1998:VLS

[CJ98] George M. Candea and Michael B. Jones. Vassal: Loadable scheduler support for multipolicy scheduling. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

usenix.org/publications/library/
proceedings/usenix-nt98/candea.
html;   http://www.usenix.
org/publications/library/
proceedings/usenix-nt98/candea_
slides.

Cherepov:1999:HRT

[CJ99] Mike Cherepov and Chris Jones. Hard real-time with RTX on Windows NT. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/cherepov.html.

Collins:1992:AHP

[CJC^{+}92] Bill Collins, Lynn Jones, Granville Chorn, Ronald Christman, Danny Cook, Kathleen Kelly, D. Lynn Kluegel, Christina Mercier, and Cheryl Ramsey. The Los-Alamos highperformance data system. In USENIX [USE92h], pages 139–

140. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Colby:1998:OCT

[CJJ^{+}98] Christopher Colby, Lalita Jategaonkar Jagadeesan, Radha Jagadeesan, Konstantin Laufer, and Carlos Puchol. Objects and concurrency in Triveni: a telecommunication case study in Java. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/colby.html.

Cheriton:1996:OMB

[CK96a] David R. Cheriton and Robert A.

Kutter. Optimized memorybased messaging: Leveraging the memory system for highperformance communication. In USENIX [USE96c], pages 179– 215.

Cooperstock:1996:WUF

[CK96b] Jeremy R. Cooperstock and Steve Kotsopoulos. Why use a fishing line when you have a net? An adaptive multicast data distribution protocol. In USENIX [USE96k], pages 343–352. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/cooperstock.html.

Curry:1990:AAC

[CKDS90] David A. Curry, Samuel D.

Kimery, Kent C. De La Croix, and Jeffrey R. Schwab. AC-MAINT: An account creation and maintenance system for distributed UNIX systems. In USENIX [USE90f], pages 1–10. LCCN QA76.76.O63 L37 1990.

Crowley:1995:TTC

[CKM95] Charles Crowley, Joseph A. Konstan, and Michael J. McLennan. Tcl and Tk in the classroom: Lessons learned (panel). In USENIX [USE95j], pages 147–150 (or 147–149??). ISBN 1-880446-72-3. LCCN QA76.73.T44

                                        T44 1995. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/tcl95/index.html.

Cabrera:1991:SUD

[CL91a] Luis-Felipe Cabrera and Darell D. E. Long. Swift: Using distributed disk striping to provide high I/O data rates. In USENIX [USE91a], pages 405–436.

Cabrera:1991:EMS

[CL91b] Luis-Felipe Cabrera and Darrell D. E. Long. Exploiting multiple I/O streams to provide high data-rates. In USENIX [USE91g], pages 31–48. LCCN QA 76.76 O63 U83 1991.

Comer:1994:PTI

[CL94] Douglas E. Comer and John C. Lin. Probing TCP implementations. In USE-NIX [USE94j], pages 245–255. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/lin.html.

Chen:1997:IIL

[CL97] J. Bradley Chen and Bradley D. D. Leupen. Improving instruction locality with just-in-time code layout. In USENIX [USE97f], pages 25–

32. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/chen.html.

Carney:1998:CMI

[CL98] Michael Carney and Brian Loe.

A comparison of methods for implementing adaptive security policies. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/loe.html.

Castro:1999:PBF

[CL99] Miguel Castro and Barbara Liskov. Practical Byzantine fault tolerance. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/castro.
html.

Clark:1992:AC

[Cla92] Michael J. Clark. Automation in the 20th Century. ;login: the USENIX Association newsletter, 17(2):47–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397.

Clark:1996:MJM

[Cla96] Dave Clark. Making the jump:

Moving from permanent to contract employment. ;login: the USENIX Association newsletter, 21(4):33–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Clarke:1997:DKB

[Cla97]  De Clarke.  “dashboard”:

a knowledge-based real-time control panel. In USE-NIX [USE97d], pages 9–18. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/tcl97/
                                        declark.html.

Clark:1999:RYR

[Cla99] David Clark. Resume your resumé writing. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-4/
                                        resume.html.

Chandra:1997:ELW

[CLD^{+}97] Satish Chandra, James R. Larus, Michael Dahlin, Bradley Richards, Randolph Y. Wang, and Thomas E. Anderson. Experience with a language for writing coherence protocols. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        dsl97/chandra.html.

Carrer:1997:TTB

[CLL97] M. Carrer, L. Ligresti, and T. D. C. Little. A Tcl/Tkbased video annotation engine.

In USENIX [USE97d], pages 181–182. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_carrer.html.

Chaves:1991:KKC

[CLMS91] Eliseu M. Chaves, Jr., Thomas J.

LeBlanc, Brian D. Marsh, and Michael L. Scott. Kernel-kernel communication in a sharedmemory multiprocessor. In USENIX [USE91j], pages 105–

116. ISBN ???? LCCN ????

Cabrera:1992:FGA

[CLS92] Luis-Filipe Cabrera, Allen W.

Luniewski, and James W. Stamos. Fine-grained access control in a transactional objectoriented system. In USENIX [USE92c], pages 199–216.

Cohan:1994:IMR

[CM94] Scott Cohan and Steve Miano.

Installing and managing remote sites. In USENIX [USE94m], pages 103–106. ISBN ???? LCCN ????

Cha:1998:ECC

[CMB^{+}98a] Lloyd Cha, Chris Motta, Syed Babar, Mukul Agarwa, Jack Ma, Waseem Shaikh, and Istvan Marko. The evolution of the CMD computing environment: a case study in rapid growth. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/cha2.html.

Cha:1998:WDW

[CMB^{+}98b] Lloyd Cha, Chris Motta, Syed Babar, Mukul Agarwal, Jack Ma, Waseem Shaikh, and Istvan Marko. What to do when the lease expires: a moving experience. In USENIX [USE98g], page ?? ISBN 1-880446-40-

5. LCCN QA76.76.O63 S97

                                        1998. URL mailto:giray@
                                        coubros.com.

Chao:1990:MVA

[CMS90] Chia Chao, Milon Mackey, and Bart Sears. Mach on a virtually addressed cache architecture. In USENIX [USE90e], pages 31–??

Clemencon:1993:DSA

[CMS93] Christian Clemencon, Bodhisattwa Mukherjee, and Karsten Schwan. Distributed shared abstractions (DSA) on large-scale multiprocessors. In USENIX [USE93j], pages 227–

246. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sedms4/clemencon.html.

Czajkowski:1999:RCJ

[CMSvE99] Grzegorz Czajkowski, Tobias Mayr, Praveen Seshadri, and Thorsten von Eicken. Resource control for Java database extensions. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots99/czajkowski.html.

Chen:1990:BFT

[CN90] Rong Chen and Tony P. Ng.

Building a fault-tolerant system based on Mach. In USENIX [USE90e], pages 157–168.

Castro:1993:MDO

[CNTS93] Miguel Castro, Nuno Neves, Pedro Trancuso, and Pedro Sousa. MIKE: a distributed objectoriented programming platform on top of the Mach micro-kernel. In USENIX [USE93h], pages 253–272. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/castro.html.

Coffman:1997:BRW

[Cof97] Dewey Coffman. Book review:

The Windows NT device driver book: a guide for programmers. ;login: the USENIX Association newsletter, 22(4):69–??, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Coggins:1990:DCC

[Cog90] James M. Coggins. Design criteria for C++ libraries. In USE-NIX [USE90j], pages 25–35 (or 25–36??).

Collinson:1991:BRPa

[Col91a]  Peter Collinson.  Book re-

view: POSIX Programmer’s Guide. ;login: the USENIX Association newsletter, 16(4):39– ??, July/August 1991. CODEN LOGNEM. ISSN 1044-6397.

Collinson:1991:BRPb

[Col91b] Peter Collinson. Book review: The POSIX.1 standard. ;login: the USENIX Association newsletter, 16(4):39–??, July/August 1991. CODEN LOGNEM. ISSN 1044-6397.

Collinson:1992:BRA

[Col92a] Peter Collinson. Book review: Advanced Programming in the UNIX Environment. ;login: the USENIX Association newsletter, 17(5):44–45, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Collyer:1992:COS

[Col92b] Geoff Collyer. A comment on open systems. ;login: the USE-NIX Association newsletter, 17 (6):22–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Collinson:1993:RDB

[Col93] Peter Collinson. Review: DNS and BIND. ;login: the USE-NIX Association newsletter, 18 (6):44–??, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Collyer:1996:SIP

[Col96] Geoff Collyer. Setting interrupt priorities in software via interrupt queueing. In USENIX [USE96b], pages 119–131.

Cooper:1992:OR

[Coo92] Michael A. Cooper. Overhauling Rdist for the ’90s. In USENIX [USE92e], pages 175–188. ISBN 1-880446-47-2. LCCN ????

Cooper:1995:SSP

[Coo95] Michael A. Cooper. SPM: System for password management. In USENIX [USE95h], pages 149–170. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/mcooper.html.

Cook:1997:ATL

[Coo97a]  Jonathan Cook.  Assertions

for the Tcl language. In USENIX [USE97d], pages 73–

80. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/cook.html.

Cook:1997:OSL

[Coo97b] Robert P. Cook. Obtuse, a scripting language for migratory applications. In USE-NIX [USE97g], pages 67–

80. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/cook.html. See also

[LH97, Rie97, WW97].

Cooper:1999:IF

[Coo99] Simon Cooper. Inexpensive firewalls, 1999. URL http://db.

usenix.org/publications/library/
proceedings/lisa99/practicum/

cooper.pdf. Unpublished practicum talk at LISA XIII.

Coplien:1994:STO

[Cop94] James O. Coplien. Supporting truly object-oriented debugging of C++ programs. In USENIX [USE94e], pages 99–

108. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/coplien.html.

Courtney:1995:PIL

[Cou95] Antony Courtney. Phantom: An interpreted language for distributed programming. In USE-NIX [USE95k], pages 83–102 (or 83–101??). ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/courtney.html.

Couch:1996:SSE

[Cou96a] Alva L. Couch. SLINK: Simple, effective filesystem maintenance abstractions for communitybased administration. In USE-NIX [USE96j], pages 205–

212. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/couch.html.

Couch:1996:VHT

[Cou96b]  Alva L. Couch.  Visualiz-

ing huge tracefiles with Xscal. In USENIX [USE96j], pages 51–58. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/couch2.html.

Couch:1997:COS

[Cou97] Alva L. Couch. Chaos out of order: a simple, scalable file distribution facility for ‘intentionally heterogeneous’ networks. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/19.couch.html.

Cowan:1999:RUS

[Cow99] Crispin Cowan. Report from a USENIX student grant project: InDependence — inferring dependencies for system components. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
student.html.

Cox:1997:AXO

[Cox97a]  Phil Cox.  Active X: An

overview. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
cox.html.

Cox:1997:SSA

[Cox97b] Phil Cox. Securely setting ACL on NT 4.0: a starting point. ;login: the USENIX Association newsletter, 22(5):14–15, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Cox:1997:SNE

[Cox97c] Phil Cox. Spam: Nothing for everyone. ;login: the USE-NIX Association newsletter, 22 (4):20–21, August 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Cox:1997:WNI

[Cox97d]  Phil Cox.  Windows NT

5.0: Integration friendly? ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1997-12/
                                        windows.html.

Cox:1998:AUD

[Cox98a]  Phil Cox.  Auditing: The

ugly duckling of computers. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-2/
                                        auditing.html.

Cox:1998:LEC

[Cox98b] Phil Cox. Letter to the Editor: Clarification. ;login: the USE-NIX Association newsletter, 23 (4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Cox:1998:NAS

[Cox98c] Phil Cox. NT accounts: a starting point. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
ntaccounts.html.

Cox:1998:TSW

[Cox98d] Phil Cox. TCPDUMP: The spanner wrench of network monitoring. ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
tcpdump.html.

Christopher:1993:NIO

[CPA93] Wayne A. Christopher, Steven J.

Procter, and Thomas E. Anderson. The Nachos instructional operating system. In USENIX [USE93l], pages 481–

488. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/christopher.pdf. Best

paper winner.

Collingson:1994:SCR

[CPK94] Peter Collingson, Jerry Peek, and Lou Katz. Summer ’94 conference reports. ;login: the USE-NIX Association newsletter, 19 (5):16–23, October 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Cowan:1998:SAA

[CPM^{+}98] Crispan Cowan, Calton Pu, Dave Maier, Jonathan Walpole, Peat Bakke, Steve Beattie, Aaron Grier, Perry Wagle, Qian Zhang, and Heather Hinton. StackGuard: Automatic adaptive detection and prevention of buffer-overflow attacks. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec98/cowan.html.

CaraDonna:1993:MLP

[CPW93] Joseph P. CaraDonna, Noemi Paciorek, and Craig E. Wills. Measuring lock performance in multiprocessor operating system kernels. In USENIX [USE93j], pages 37–56. ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sedms4/.

Clark:1994:DFS

[CR94] Morgan Clark and Stephen Rago. The desktop file system. In USENIX [USE94j], pages 113–124. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/index.html.

Crabb:1990:PSL

[Cra90] Michele Crabb. Password security in a large distributed environment. In USENIX [USE90i], pages 17–30. LCCN QA 76.9 A25 U55 1990.

Crabb:1994:GFE

[Cra94a] Michele D. Crabb. Guarding the fortress: Efficient methods to monitor security on 300 systems. In USENIX [USE94m], pages 123–135. ISBN ???? LCCN ????

Crabb:1994:WTW

[Cra94b] Michele D. Crabb. Who’s trusting whom? How to audit and manage users’ .rhosts files. In USENIX [USE94m], pages 15–

22. ISBN ???? LCCN ????

Crabb:1996:SPL

[Cra96] Michele D. Crabb. Security policies — lead-weighted balloons or safety vests? ;login: the USE-NIX Association newsletter, 21 (3):38–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Crew:1997:ALE

[Cre97] Roger F. Crew. ASTLOG:

a language for examining abstract syntax trees. In USE-NIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/crew.html.

Cota-Robles:1999:CWD

[CRH99] Erik Cota-Robles and James P.

Held. A comparison of Windows driver model latency performance on Windows NT and Windows 98. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/cota.html.

Crowley:1995:TRR

[Cro95] Charles Crowley. TkReplay:

Record and replay for Tk. In USENIX [USE95j], pages 131–140. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/crowley.html.

Crosser:1998:IFT

[Cro98] Eugene G. Crosser. Ifmail: FidoNet — technology implementation of UNIX platform. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/crosser.
ps.

Chew:1993:KSR

[CRRS93] Kien-Mien Chew, Jyothy Reddy, Theodore H. Romer, and Abraham Silberschatz. Kernel support for recoverable-persistent virtual memory. In USE-NIX [USE93h], pages 215–234. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Cohen:1999:PTS

[CRS99] Ariel Cohen, Sampath Rangarajan, and Hamilton Slye. On the performance of TCP splicing for URL-aware redirection. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/cohen.html.

Carpenter:1997:RRT

[CRVZ97] Bill Carpenter, Mark Roman, Nick Vasilatos, and Myron Zimmerman. The RTX real-time subsystem for Windows NT. In USENIX [USE97f], pages 33–37. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/carpenter.html.

Curran:1990:CBC

[CS90] Stephan Curran and Michael Stumm. A comparison of basic CPU scheduling algorithms for multiprocessor UNIX. In USE-NIX [USE90a], pages 551–580.

Carson:1992:OWB

[CS92] Scott Carson and Sanjeev Setia. Optimal write batch size in log-structured file systems. In USENIX [USE92h], pages 79–

92. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Callaghan:1993:AA

[CS93] Brent Callaghan and Satinder Singh. The Autofs Automounter. In USENIX [USE93g], pages 59–68. ISBN 1-880446-50-2. LCCN QA 76.76 O63

U83 1993. URL http://www.
usenix.org/publications/library/
proceedings/cinci93/callaghan.
html.

Carson:1994:OWB

[CS94] Scott Carson and Sanjeev Setia.

Optimal write batch size in logstructured file systems. In USE-NIX [USE94b], pages 263–281.

Chevenet:1997:WWW

[CS97] François Chevenet and Guillaume Le Stum. A World Wide Web-to-database connectivity using Tcl/Tk: The hydrology-meteorology toolkit. In USENIX [USE97d], pages 165–166. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_chevenet.html.

Cost:1997:ADS

[CSL^{+}97] R. Scott Cost, Ian Soboroff, Jeegar Lakhani, Tim Finin, Ethan Miller, and Charles Nicholas. Agent development support for Tcl. In USE-NIX [USE97d], pages 177–

178. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_cost.html.

Cutler:1997:USH

[CSM97] Michal Cutler, Yungming Shih, and Weiyi Meng. Using the structure of HTML documents to improve retrieval. In USE-NIX [USE97h], pages 241–251. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/cutler.html.

Camp:1995:TNM

[CST95] L. Jean Camp, Marvin Sirbu, and J. D. Tygar. Token and notational money in electronic commerce. In USENIX [USE95g], pages 1–

12. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/camp.html.

Chin:1994:CDI

[CSYG94] Goodwin Chin, Dharini Sitaraman, Chung Yang, and Martin Giles. C++ design and implementation challenges in technology computer aided design frameworks. In USE-NIX [USE94e], pages 189–205. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/chin.html.

Cox:1995:NST

[CTS95] Benjamin Cox, J. D. Tygar, and Marvin Sirbu. NetBill security and transaction protocol. In USENIX [USE95g], pages 77–88. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/cox.html.

Cuccia:1990:BRU

[Cuc90] Nichlos H. Cuccia. Book review: UNIX System Administration Handbook. ;login: the USENIX Association newsletter, 15(4):13–??, July/August 1990. CODEN LOGNEM. ISSN 1044-6397.

Cuccia:1991:DIM

[Cuc91] Nichlos H. Cuccia. The design and implementation of a multihub electronic mail environment. In USENIX [USE91f], pages 37–44.

Cuccia:1995:SMB

[Cuc95] Nick Cuccia. SysAdmin at the movies: Braveheart. ;login: the USENIX Association newsletter, 20(6):33–??, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Cuccia:1996:SMS

[Cuc96] Nick Cuccia. SysAdmin at the movies strikes again. …;login: the USENIX Association newsletter, 21(2):36–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Cummings:1999:MID

[Cum99a] Curt Cummings. Microsoft’s internal deployment of Windows 2000, 1999. URL http://db.

                                        usenix.org/publications/library/
                                        proceedings/lisa99/invitedtalks.

html. Unpublished invited talk

at LISA XIII.

Cummings:1999:NMS

[Cum99b] Curt Cummings. NT 5.0 migration strategies at Microsoft,

                                        1999.  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/lisa-nt98/cummings_

html. Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Curry:1991:AAC

[Cur91] David Curry. ACMAINT: An account creation and maintenance system for distributed UNIX systems. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Curry:1994:PTD

[Cur94] Timothy W. Curry. Profiling and tracing dynamic library usage via interposition. In USENIX [USE94j], pages 267–278. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/curry.html.

Curry:1997:RUS

[Cur97] David Curry. Review: UNIX systems programming for SVR4. ;login: the USENIX Association newsletter, 22(1):58–59, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Curtin:1999:ESO

[Cur99] Matt Curtin. Electronic snake oil. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
snake_oil.html.

Cutler:1992:N

[Cut92] David Cutler. NT. In USE-NIX [USE92i], page ?? ISBN 1-880446-42-1. LCCN QAX 32. (paper not included in this proceedings).

Cotrozzi:1993:AAT

[CV93] Massimo Cotrozzi and David Vincenzetti. ATP — antitampering program. In USE-NIX [USE93m], pages 79–

89. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec4/.

Conn:1996:CCB

[CV96] Coco Conn and Zane Vella. CitySpace: Come build it yourself — A user-extensible virtual environment for real-time play, 1996. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/sd96/. Unpub-

lished invited talk at the USE-NIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Castagnoli:1992:DTU

[CW92a] Charisse Castagnoli and Charles Watt. Distributed trusted UNIX systems. In USENIX [USE92m], pages 347–?? ISBN 1-880446-46-4. LCCN ????

Colyer:1992:DTM

[CW92b] Wallace Colyer and Walter Wong. Depot: a tool for managing software environments. In USENIX [USE92e], pages 153–

162. ISBN 1-880446-47-2. LCCN

????

Clark:1996:HCD

[CW96] K. L. Clark and T. I. Wang. Highly concurrent distributed knowledge objects. In USE-NIX [USE96g], pages 251–

261. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/clark.html.

Cheriton:1990:BEU

[CWS90] David R. Cheriton, Gregory R.

Whitehead, and Edward W. Sznyter. Binary emulation of UNIX using the V kernel. In USENIX [USE90g], pages 73–

86. LCCN QA76.8.U65 U81 1990.

Carter:1997:RSS

[CWT^{+}97] Mark R. Carter, Jay Williamson, Paul Thomas, John Mastin, Jr., Brian Kirouac, Carolyn M. Hennings, Sandy Cash, and Idajean Fisher. Reports on SANS: The sixth annual system administration, networking and security conference. ;login: the USE-NIX Association newsletter, 22 (4):42–53, August 1997. CO-DEN LOGNEM. ISSN 1044-6397.

dAusbourg:1992:UOS

[d’A92a] Bruno d’Ausbourg. UNIX operating services on a multilevel secure machine. In USENIX [USE92m], pages 329–346. ISBN 1-880446-46-4. LCCN ????

Dean:1992:DMK

[DA92b] Randall W. Dean and François Armand. Data movement in kernelized systems. In USENIX [USE92i], pages 243–262. ISBN 1-880446-42-1. LCCN QAX 32.

Dasgupta:1991:LOS

[DAM^{+}91a] Partha Dasgupta, R. Ananthanarayanan, Sathis Menon, Ajay Mohindra, Mark Pearson, Raymond Chen, and Christopher Wilkenloh. Language and operating system support for distributed programming in Clouds. In USENIX [USE91j], pages 321–340. ISBN ???? LCCN ????

Dasgupte:1991:DPO

[DAM^{+}91b] Partha Dasgupte, R. Ananthanarayanan, Sathis Menon, Ajay Mohindra, and Raymond Chen. Distributed programming with objects and threads in the Clouds system. In USENIX [USE91c], pages 243–276.

Damon:1993:PAB

[Dam93] Lee Damon. Pragmatica: The alias builder. ;login: the USE-NIX Association newsletter, 18 (6):33–38, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Damon:1995:GHT

[Dam95] Lee Damon. On glass houses and thrown power bricks; or: a Solaris hater eats his hat. ;login: the USENIX Association newsletter, 20(3):34–35, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Damon:1997:IJ

[Dam97a] Lee Damon. IMHO: Jihad. ;login: the USENIX Association newsletter, 22(1):4–5, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Damon:1997:WII

[Dam97b] Lee Damon. WWWhither(ing)

Internet. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
imho.html.

Damon:1998:LER

[Dam98] Lee Damon. Letter to the editor:

In response. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

Damon:1999:IY

[Dam99] Lee Damon. IMHO: Y2K.

;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/imho.html.

Daneliuk:1994:CCC

[Dan94] Tim Daneliuk. Conquering corporate computing with messageoriented middleware. ;login: the USENIX Association newsletter, 19(4):28–31, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1991:SCS

[Dar91] Tine M. Darmohray. A sendmail.cf scheme for a large network. In USENIX [USE91f], pages 45–50.

Darmohray:1993:JDS

[Dar93a] Tina Darmohray, editor. Job Descriptions for System Administrators, October, 1993. USE-NIX for SAGE, the System Administrator’s Guild, Berkeley, CA, USA, October 1993.

Darmohray:1993:STS

[Dar93b] Tina Darmohray. Short topics in system administration #1. In USENIX [USE93e].

Darmohray:1993:SJD

[Dar93c] Tina M. Darmohray. SAGE job descriptions. ;login: the USE-NIX Association newsletter, 18 (5):18–22, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1995:E

[Dar95a] Tina Darmohray. From the editor. ;login: the USENIX Association newsletter, 20(6):27– ??, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1995:SEa

[Dar95b] Tina Darmohray. From the SAGE editor. ;login: the USE-NIX Association newsletter, 20 (3):10–11, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1995:SEb

[Dar95c] Tina M. Darmohray. From the SAGE editor. ;login: the USE-NIX Association newsletter, 20 (5):10–11, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Darmohray:1995:DY

[Dar95d] Tina M. Darmohray. I dare you! ;login: the USENIX Association newsletter, 20(2):28–29, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1995:SMN

[Dar95e] Tina M. Darmohray. Somewhere in the middle of nowhere. ;login: the USENIX Association newsletter, 20(4):12–13, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:E

[Dar96a] Tina Darmohray. From the editor. ;login: the USENIX Association newsletter, 21(1):9– ??, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:SER

[Dar96b] Tina Darmohray. SAGE election results. ;login: the USE-NIX Association newsletter, 21 (1):9–??, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:Y

[Dar96c] Tina Darmohray. You’goo. ;login: the USENIX Association newsletter, 21(1):19–??, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:BF

[Dar96d] Tina M. Darmohray. Back to the future. ;login: the USENIX Association newsletter, 21(6):6– ??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:DP

[Dar96e] Tina M. Darmohray. Disaster preparation. ;login: the USE-NIX Association newsletter, 21 (2):24–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:ESF

[Dar96f] Tina M. Darmohray. Editorial on SAGE focus: Being A consultant. ;login: the USE-NIX Association newsletter, 21 (4):30–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:FJF

[Dar96g] Tina M. Darmohray. Not fancy, just functional. ;login: the USE-NIX Association newsletter, 21 (5):20–??, October 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Darmohray:1996:SNE

[Dar96h] Tina M. Darmohray. SAGE news editorial. ;login: the USE-NIX Association newsletter, 21 (3):12–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1997:AL

[Dar97a]  Tina Darmohray.  [A]bused

leashes. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
leashes.html.

Darmohray:1997:BF

[Dar97b] Tina Darmohray. Back to the future. ;login: the USENIX Association newsletter, 22(2):8–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1997:IDA

[Dar97c] Tina Darmohray. Interview with Dan Appelman — the SAGE code of ethics. ;login: the USE-NIX Association newsletter, 22 (2):10–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1997:SB

[Dar97d] Tina Darmohray. Strange bedfellows. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
tmd.html.

Darmohray:1997:WYR

[Dar97e] Tina Darmohray. Would you repeat the question? ;login: the USENIX Association newsletter, 22(4):12–??, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1997:PBV

[Dar97f] Tina M. Darmohray. A peek at BIND version 8. ;login: the USENIX Association newsletter, 22(5):13–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1997:R

[Dar97g] Tina M. Darmohray. Reflections. ;login: the USENIX Association newsletter, 22(5):10–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1997:SYI

[Dar97h] Tina M. Darmohray. The sevenyear itch. ;login: the USE-NIX Association newsletter, 22 (1):7–8, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1998:AB

[Dar98a] Tina Darmohray. 80/20 across the board. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-6/
                                        editorial.html.

Darmohray:1998:CS

[Dar98b] Tina Darmohray. Coming up to speed. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-12/
                                        comingup.html.

Darmohray:1998:HSH

[Dar98c] Tina Darmohray. Home sweet home. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1998:MYC

[Dar98d]  Tina Darmohray.  Manag-

ing your career using the 80/ 20 rule. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/

publications/login/1998-4/
editorial.html.

Darmohray:1998:TS

[Dar98e] Tina Darmohray. A trip to SANS. ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Darmohray:1998:WR

[Dar98f] Tina Darmohray. Worth repeating. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
worth.html.

Darmohray:1999:BTB

[Dar99a] Tina Darmohray. Better than better. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
better.html.

Darmohray:1999:SW

[Dar99b] Tina Darmohray. Off to see the wizards. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
wizards.html.

Darmohray:1999:WH

[Dar99c] Tina Darmohray. Where’s the horse? ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
horse.html.

Darmohray:1999:WA

[Dar99d] Tina Darmohray. Wisdom of the aged. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
wisdom.html.

Darmohray:1999:WW

[Dar99e] Tina Darmohray. Wisdom of words. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
wisdom.html.

Dasgupta:1997:PPW

[Das97]  Partha Dasgupta.  Paral-

lel processing with Windows NT networks. In USE-NIX [USE97f], pages 113–

122. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/dasgupta.html.

Davenport:1991:ML

[Dav91] Glorianna Davenport. The MIT media laboratory. In USE-NIX [USE91g], pages 349–350. LCCN QA 76.76 O63 U83 1991.

Davey:1994:PMX

[Dav94a]  Paul Davey.  Porting and

maintaining with X and Motif: a retrospective view. In USENIX [USE94g], pages 171–

176. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/davey.html.

Davis:1994:RAC

[Dav94b] Donald T. Davis. Review:

Applied Cryptography. ;login: the USENIX Association newsletter, 19(2):46–??, March/April 1994. CODEN LOGNEM. ISSN 1044-6397.

Davis:1995:KPR

[Dav95] Don Davis. Kerberos plus RSA for world wide Web security. In USENIX [USE95g], pages 185–188. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/davis.html.

Davis:1996:CDP

[Dav96] Don Davis. Compliance defects in public key cryptography. In USENIX [USE96i], pages 171–178. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/davis.html.

Douglis:1996:TVC

[DB96a] Fred Douglis and Thomas Ball. Tracking and viewing changes on the Web. In USENIX [USE96k], pages 165–

176. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/douglis.html.

Druschel:1996:LRP

[DB96b] Peter Druschel and Gaurav Banga. Lazy receiver processing (LRP): a network subsystem architecture for server systems. In USENIX [USE96f], pages 261–275. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/druschel.html.

Daly:1999:NSO

[DBC^{+}99] Gregg Daly, Gary Buhrmaster, Matthew Campbell, Andrea Chan, Robert Cowles, Ernest Denys, Patrick Hancox, Bill Johnson, David Leung, and Jeff Lwin. NT security in an open academic environment. In USE-NIX [USE99b], pages 59–67. ISBN 1-880446-30-8. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa-nt99/daly.
html.

Davies:1995:NES

[DBCF95] Nigel Davies, Gordon S. Blair, Keith Cheverst, and Adrian Friday. A network emulator to support the development of adaptive applications. In USE-NIX [USE95i], pages 47–56 (or 47–55??). ISBN 1-880446-69-

3. LCCN TK5103.2 .U83 1995.

URL http://www.usenix.org/
publications/library/proceedings/
mob95/davies.html.

Denker:1999:MVP

[DBD^{+}99] John S. Denker, Steven M.

Bellovin, Hugh Daniel, Nancy L. Mintz, Tom Killian, and Mark A. Plotnick. Moat: a virtual private network appliance and services platform. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/denker.
html.

Dagenais:1993:LDS

[DBGL^{+}93] Michel Dagenais, Stéphane Boucher, Robert Grin-Lajoie, Pierre Laplante, Pierre Mailhot, and Centre de Recherche Informatique de Montréal. LUDE: a distributed software library. In USENIX [USE93f], pages 25–32. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/dagenais.html.

Daswani:1998:SSW

[DBGM^{+}98] Neil Daswani, Dan Boneh, Hector Garcia-Molina, Steven Ketchpel, and Andreas Paepcke. SWAPEROO: a simple wallet architecture for payments, exchanges, refunds, and other operations. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

URL http://www.usenix.org/

                                        publications/library/proceedings/
                                        ec98/daswani.html.

DeJong:1998:CME

[DBK98] Moses DeJong, Brian Bailey, and Joseph A. Konstan. Creating a multimedia extension for Tcl using the Java Media Framework. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl98/dejong.
                                        html.

dAusbourg:1993:USM

[dC93] Bruno d’Ausbourg and Christel Calas. UNIX services for multilevel storage and communications over a secure LAN. In USENIX [USE93l], pages 157–168. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Douglis:1994:SAM

[DCK^{+}94] Fred Douglis, Ramon Caceres, Frans Kaashoek, Kai Li, Brian Marsh, Joshua A. Tauber, and D. E. Shaw. Storage alternatives for mobile computers. In USENIX [USE94i], pages 25–37. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/douglis.html.

Dasgupta:1990:DIC

[DCM^{+}90] P. Dasgupta, R. C. Chen, S. Menon, M. P. Pearson, R. Ananthanarayanan, U. Ramachandran, M. Ahamad, R. J. LeBlanc, W. F. Appelbe, J. M. Bernabeu-Auban, P. W. Hutto, M. Y. A. Khalidi, and C. J. Wilkenloh. The design and implementation of the Clouds distributed operating system. In USENIX [USE90d], pages 11– 46.

Davies:1992:EHM

[DCWB92] Nigel Davies, Geoff Coulson, Neil Williams, and Gordon S. Blair. Experiences of handling multimedia in distributed open systems. In USENIX [USE92l], pages 173–190. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Daghi:1992:CFD

[DDD92] Afshin Daghi, Pierre Delisle, and Salil Deshpande. A communication facility for distributed object-oriented applications. In USENIX [USE92k], pages 263– 277. ISBN 1-880446-45-6. LCCN ????

DeAlvare:1990:HCC

[De 90] Ana Maria De Alvare. How crackers crack passwords. In USENIX [USE90i], pages 103– 112. LCCN QA 76.9 A25 U55 1990.

Dearle:1990:DPA

[Dea90]  Fergal Dearle.  Designing

portable application frameworks for C++. In USENIX [USE90j], pages 51–61 (or 51–62??).

Dean:1993:UCB

[Dea93] Randall W. Dean. Using continuations to build a userlevel threads library. In USE-NIX [USE93h], pages 136–151. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mach3/dean.html.

deLongpre:1999:HCD

[deL99]  Chris deLongpre.  How-to:

Configure DNS. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/sage/sysadmins/
                                        howto/3.dns.html.

Dennehy:1992:CLA

[Den92] Thomas G. Dennehy. Class libraries as an alternative to language extensions for distributed programming. In USE-NIX [USE92l], pages 313–326. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Dennehy:1995:DSA

[Den95]  Thomas G. Dennehy.  A

distributed software architecture for GPS-Driven mobile applications. In USE-NIX [USE95i], pages 99–

108. ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/dennehy.html.

Deno:1998:TGG

[Den98] Cynthia Deno. There’s gold in good works: a report on USENIX support of worthwhile projects. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
reference.html.

Deno:1999:NRB

[Den99a] Cynthia Deno. New releases of *BSD and Debian Linux OSes given away at USENIX Annual Conference. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
newreleases.html.

Deno:1999:UFH

[Den99b]  Cynthia Deno.  USENIX

funding helps improve the US patent process for software. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
funding.html.

Deno:1999:UTP

[Den99c]  Cynthia Deno.  USENIX

teams up to put on the 1999 Atlanta Linux Showcase. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-6/
                                        als.html.

Depp:1994:RBS

[Dep94] Gregory C. Depp. Review: Building a Successful Software Business. ;login: the USE-NIX Association newsletter, 19 (6):52–53, December 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Detke:1991:HAS

[Det91] John F. Detke. Host aliases and symbolic links -or- how to hide the servers’ real name. In USE-NIX [USE91f], pages 249–252.

Detlefs:1992:GCR

[Det92] David Detlefs. Garbage collection and run-time typing as a C++ library. In USENIX [USE92k], pages 37–56. ISBN 1-880446-45-6. LCCN ????

DeVaney:1993:UDL

[DeV93] Mike DeVaney. UNIX is dead; long live UNIX. ;login: the USE-NIX Association newsletter, 18 (1):23–24, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

DeWitt:1991:CPO

[DeW91] David DeWitt. C++ + persistence != an object-oriented DBMS. In USENIX [USE91k], page ?? LCCN QA76.73.C153 U27 1991.

Dewey:1997:HSM

[Dew97] Brian Dewey. How I spent my summer vacation, or, escape from the evil empire. ;login: the USENIX Association newsletter, 22(6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
summer.html.

DeHart:1993:USS

[DF93] Edward DeHart and Barbara Fraser. 1992 UNIX security symposium report. ;login: the USENIX Association newsletter, 18(1):4–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Don:1995:CCM

[DF95] Bruce Don and David Frelinger.

Can the conventional models apply? The microeconomics of the information revolution. In USENIX [USE95g], pages 23–31. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/don.html.

Defert:1996:MDA

[DFG^{+}96] Ph. Defert, E. Fernandez, M. Goossens, O. Le Moigne, A. Peyrat, and I. Reguero. Managing and distributing application software. In USE-NIX [USE96j], pages 213–

226. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/

publications/library/proceedings/
lisa96/reguero.html.

Defago:1997:RCE

[DFGG97] Xavier Défago, Pascal Felber, Benoît Garbinato, and Rachid Guerraoui. Reliability with CORBA event channels. In USENIX [USE97g], pages 237–

240. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/technical.html. See

also [LH97, Rie97, WW97].

Douglis:1997:RCO

[DFKM97] Fred Douglis, Anja Feldmann, Balachander Krishnamurthy, and Jeffrey Mogul. Rate of change and other metrics: a live study of the World Wide Web. In USE-NIX [USE97h], pages 147–158. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/douglis_rate.html.

Drapeau:1991:MDM

[DG91] George D. Drapeau and Howard Greenfield. MAEstro — A distributed multimedia authoring environment. In USE-NIX [USE91g], pages 315–

328. LCCN QA 76.76 O63

U83 1991. URL ftp://sioux.
stanford.edu/pub/MAEstro/
MAEstro.tar.Z.  FTP -

sioux.stanford.edu:/pub/MAEstro/MAEstro.tar.Z [./Documentation/UsenixPaper.ps]; local - maestro.ps.

Dooley:1994:CCC

[DG94] John F. Dooley and Vince Guarna. Creating a configurable compiler driver for System V Release 4. In USENIX [USE94g], pages 67–

77. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/dooley.html.

Davis:1995:KSC

[DG95] Don Davis and Daniel E. Geer.

Kerberos security with clocks adrift. In USENIX [USE95f], pages 35–40. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
davis.html.

DSouza:1996:AMD

[DG96] Desmond F. D’Souza and Petter Graff. Advanced modeling and design for Java systems. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

Davis:1996:KCA

[DGT96] Donald T. Davis, Daniel E.

Geer, and Theodore Ts’o. Kerberos with clocks adrift: History, protocols, and implementation. In USENIX [USE96d], pages 29–46.

Dunigan:1999:IDI

[DH99] Tom Dunigan and Greg Hinkel. Intrusion detection and intrusion prevention on a large network: a case study. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/detection99/
                                        dunigan.html.

Douglis:1997:HHM

[DHR97] Fred Douglis, Antonio Haro, and Michael Rabinovich. HPP: HTML macro-preprocessing to support dynamic document caching. In USE-NIX [USE97h], pages 83–94. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits97/douglis_hpp.html.

Dijker:1996:DKN

[Dij96a] Barbara L. Dijker. A different kind of networking. ;login: the USENIX Association newsletter, 21(4):28–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Dijker:1996:SC

[Dij96b] Barbara L. Dijker. Step away from the computer. ;login: the USENIX Association newsletter, 21(3):20–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Dijker:1997:AO

[Dij97] Barb Dijker. Apricots and oranges. ;login: the USENIX Association newsletter, 22(2): 11–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Dijker:1998:SCQ

[Dij98] Barbara L. Dijker. SAGE certification: Q&A. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

Dijker:1999:RRW

[Dij99a] Barbara L. Dijker. Round and round we go. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
round.html.

Dijker:1999:SCU

[Dij99b] Barbara L. Dijker. SAGE certification update. ;login: the USE-NIX Association newsletter, 24 (5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
sagecert.html.

Dilley:1995:OCF

[Dil95] John Dilley. OODCE: a C++ framework for the OSF distributed computing environment. In USENIX [USE95e], pages 107–118. ISBN 1-880446-67-7. LCCN QA 76.76 O63

U88 1995. URL http://www.
usenix.org/publications/library/

proceedings/neworl/dilley.
html.

deIcaza:1998:GDE

[dILM98] Miguel de Icaza, Elliot Lee, and Federica Mena. The GNOME desktop environment. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/icaza.ps.

Detlefs:1995:DSM

[DK95] Dave L. Detlefs and Bill Kaslow. Debugging storage management problems in garbagecollected environments. In USENIX [USE95k], pages 69–

82. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/detlefs.html.

Damon:1999:E

[DK99] Lee Damon and Rob Kolstad. Ethics, 1999. URL http://db.

usenix.org/publications/library/
proceedings/lisa99/practicum.

html. Unpublished practicum talk at LISA XIII.

Douglis:1995:ADSa

[DKB95a] Fred Douglis, P. Krishnan, and Brian Bershad. Adaptive disk spin-down policies for mobile computers. In USE-NIX [USE95a], pages 381–

413. URL http://www.usenix.
org/publications/library/
proceedings/mob95/douglis.
html.

Douglis:1995:ADSb

[DKB95b] Fred Douglis, P. Krishnan, and Brian Bershad. Adaptive disk spin-down policies for mobile computers. In USENIX [USE95i], pages 121–

137. ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995.

URL http://www.usenix.org/
publications/library/proceedings/
mob95/douglis.html.

Douglis:1994:TPH

[DKM94] Fred Douglis, P. Krishnan, and Brian Marsh. Thwarting the power-hungry disk. In USE-NIX [USE94l], pages 292–306. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/marsh.html.

DeSimone:1993:SDS

[DL93] Salvatore DeSimone and Christine Lombardi. Sysctl: a distributed system control package. In USENIX [USE93f], pages 131–143. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/deSimone.html.

deLeon:1995:TSA

[dL95] Arnold de Leon. From Thinnet to 10base-T, from Sys Admin to Network Manager. In USENIX [USE95h], pages 229–240. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/

                                        publications/library/proceedings/
                                        lisa95/deleon.html.

Doster:1990:UO

[DLM90] William A. Doster, Yew-Hong Leong, and Steven J. Mattson. Uniqname overview. In USENIX [USE90f], pages 27–36. LCCN QA76.76.O63 L37 1990.

Danzig:1992:DIA

[DLO92] Peter B. Danzig, Shih-Hao Li, and Katia Obraczka. Distributed indexing of autonomous Internet services. In USENIX [USE92a], pages 433–459.

deLeon:1993:OUR

[dLRT93] Laura de Leon, Mike Rodriquez, and Brent Thompson. Our users have root! In USENIX [USE93f], pages 17–24. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/deLeon.html.

DePauw:1998:EPO

[DLVW98] Wim De Pauw, David Lorenz, John Vlissides, and Mark Wegman. Execution patterns in object-oriented visualization. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots98/depauw.html.

Devarakonda:1992:IIC

[DM92] Murthy Devarakonda and Arup Mukherjee. Issues in implementation of cache-affinity scheduling. In USENIX [USE92j], pages 345–358.

Duska:1997:MAC

[DMF97] Bradley M. Duska, David Marwood, and Michael J. Feeley. The measured access characteristics of World-Wide-Web client proxy caches. In USE-NIX [USE97h], pages 23–35. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/duska.html.

Duby:1992:CMC

[DMR92] Carolyn K. Duby, Scott Meyers, and Steven P. Reiss. CCEL: a metalanguage for C++. In USENIX [USE92k], pages 99–

115. ISBN 1-880446-45-6. LCCN ????

Devarakonda:1995:EDA

[DMST95] Murthy Devarakonda, Ajay Mohindra, Jill Simoneaux, and William H. Tetzlaff. Evaluation of design alternatives for a cluster file system. In USE-NIX [USE95e], pages 35–46. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/devarakonda.html.

Dougan:1999:OIT

[DMY99] Cort Dougan, Paul Mackerras, and Victor Yodaiken. Optimizing the idle task and other MMU tricks. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

                                        5. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi99/dougan.
                                        html.

Draves:1998:IIW

[DMZ98] Richard P. Draves, Allison Mankin, and Brian D. Zill. Implementing IPv6 for Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt98/draves.html;
                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/draves_slides.pdf.

Douglis:1991:CTD

[DOKT91] Fred Douglis, John K. Ousterhout, M. Frans Kaashoek, and Andrew S. Tannenbaum. A comparison of two distributed systems: Amoeba and Sprite. In USENIX [USE91a], pages 353– 384.

Donahue:1999:HSA

[Don99]  Adam M. Donahue.  How-

to: Set up an Apache Web server. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-6/
                                        features/howto.html.

Doster:1991:UO

[Dos91]  Bill Doster.  UNIQNAME

overview. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Douglis:1993:CCU

[Dou93] Fred Douglis. The compression cache: Using on-line compression to extend physical memory. In USENIX [USE93l], pages 519–529. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Gauntt:1996:DCP

[dPG96] John du Pre Gauntt. Digital currency and public networks: So what if it is secure, is it money? In USENIX [USE96h], pages 77–86. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Gauntt:1998:ECM

[dPG98] John du Pré Gauntt. Electronic commerce on the move. In USE-NIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http://www.

usenix.org/publications/library/
proceedings/ec98/gauntt.html.

deRaadt:1998:O

[dR98] Theo de Raadt. OpenBSD.

In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/deraadt.
ps.

Daemen:1999:EBC

[DR99] Joan Daemen and Vincent Rijmen. Efficient block ciphers for Smartcards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/smartcard99/
                                        daemen.html.

Draves:1990:RII

[Dra90] Richard Draves. A revised IPC interface. In USENIX [USE90e], pages 101–122.

Draves:1991:PRR

[Dra91] Richard P. Draves. Page replacement and reference bit emulation in Mach. In USE-NIX [USE91h], pages 201–212. LCCN QAX 27.

Draves:1992:M

[Dra92] Richard Draves. Mach. In USENIX [USE92i], pages 11–

30. ISBN 1-880446-42-1. LCCN

QAX 32.

deRaadt:1999:COO

[dRHG^{+}99] Theo de Raadt, Niklas Hallqvist, Artur Grabowski, Angelos D. Keromytis, and Niels Provos. Cryptography in OpenBSD: An overview. In USENIX [USE99m], page ?? ISBN 1-880446-33-2. LCCN

                                        ????  URL http://www.
                                        openbsd.org/papers/crypt-
                                        paper.ps.

Davis:1992:NSP

[DS92] Don Davis and Ralph Swick.

Network security via privatekey certificates. In USENIX [USE92m], pages 239–242. ISBN 1-880446-46-4. LCCN ????

daSilva:1998:NCM

[dSdCF^{+}98] Fabio Q. B. da Silva, Juliana Silva da Cunha, Danielle M. Franklin, Luciana S. Varejão, and Rosalie Belian. An NFS configuration management system and its underlying object-oriented model. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
silva.html.

daSilveira:1998:CDS

[dSdS98] Glêdson Elias da Silveira and Fabio Q. B. da Silva. A configuration distribution system for heterogeneous networks. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/silveira.html.

Silva:1993:ANB

[dSG93] James da Silva and Olafur Gumundsson. The Amanda network backup manager. In USENIX [USE93f], pages 171–

182. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/

                                        publications/library/proceedings/
                                        lisa93/daSilva.html.

Silva:1992:PPN

[dSGM92] James da Silva, Olafur Gudmundsson, and Daniel Mosse. Performance of a parallel network backup manager. In USE-NIX [USE92g], pages 217–226. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

daSilva:1996:BNP

[dSM96] Fabio Queda Bueno da Silva and Silvio Lemos Meira. The Brazilian National Program of Open Systems Reference Centers. ;login: the USENIX Association newsletter, 21(1):28–31, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Dorward:1990:ANC

[DSS90] Sean M. Dorward, Ravi Sethi, and Jonathan E. Shopiro. Adding new code to a running C++ program. In USENIX [USE90j], pages 279–292.

Darmohray:1996:NLA

[DSS96] Tina Darmohray, John Schimmel, and John Sellens. Notes on the LISA ’96 Advanced Topics in System Administration Workshop. ;login: the USENIX Association newsletter, 21(6):41– 46, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Dinesh:1997:SBA

[DT97] T. B. Dinesh and Frank Tip. A slicing-based approach for locating type errors. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/dinesh.html.

Dubois:1994:IR

[Dub94] Paul Dubois. Imake response.

;login: the USENIX Association newsletter, 19(1):34–35, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Duchamp:1991:ETR

[Duc91] Dan Duchamp. Experience with threads and RPC in Mach. In USENIX [USE91j], pages 87–

104. ISBN ???? LCCN ????

Duchamp:1994:OLW

[Duc94]  Dan Duchamp.  Optimistic

lookup of whole NFS paths in a single operation. In USENIX [USE94j], pages 161–

169. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/duchamp.html.

Duchamp:1997:TAP

[Duc97]  Dan Duchamp.  A toolkit

approach to partially connected operation. In USE-NIX [USE97e], pages 305–

318. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/duchamp.html;   http:
//www.usenix.org/publications/
library/proceedings/ana97/
summaries/duchamp.html.

Duchamp:1999:PH

[Duc99] Dan Duchamp. Prefetching hyperlinks. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/duchamp.html.

Dunlop:1990:IS

[Dun90a]  Dominic Dunlop.  Interna-

tional standardization. ;login: the USENIX Association newsletter, 15(3):21–24, May/June 1990. CODEN LOGNEM. ISSN 1044-6397.

Dunlop:1990:RIIb

[Dun90b] Dominic Dunlop. Report on ISO/IEC JTC1/SC22/WG15 meeting. ;login: the USE-NIX Association newsletter, 15 (5):10–??, September/October 1990. CODEN LOGNEM. ISSN 1044-6397.

Dunlop:1990:RIIa

[Dun90c] Dominic Dunlop. Report on ISO/IEC JTC1/SC22/WG15 rapporteur group on internationalization meeting. ;login: the USENIX Association newsletter, 15(3):17–20, May/June 1990. CODEN LOGNEM. ISSN 1044-6397.

Dunlop:1990:VOS

[Dun90d] Dominic Dunlop. A view of the organisational structure of POSIX standards. EUUG Newsletter, 10(1):76–79, Spring 1990. CODEN EONLE8. ISSN 1011-4211.

Dunlop:1991:ESR

[Dun91a]  Dominic Dunlop.  EurOpen

standards report. ;login: the USENIX Association newsletter, 16(2):12–??, March/April 1991. CODEN LOGNEM. ISSN 1044-6397.

Dunlop:1991:RIP

[Dun91b] Dominic Dunlop. Report on ISO POSIX working group. ;login: the USENIX Association newsletter, 16(1):17–??, January/February 1991. CODEN LOGNEM. ISSN 1044-6397.

Dunlop:1991:ROP

[Dun91c] Dominic Dunlop. Report on OSI POSIX working group. ;login: the USENIX Association newsletter, 16(5):13–??, September/October 1991. CODEN LOGNEM. ISSN 1044-6397.

Dewan:1990:OMC

[DV90] Prasun Dewan and Eric Vasilic.

An object model for conventional operating systems. In USENIX [USE90a], pages 517– 550.

Dearle:1994:GOP

[DVdB^{+}94] Alan Dearle, Francis Vaughan, Rex di Bona, James Farrow, Frans Henskens, Anders Lindstrom, and John Rosenberg. Grasshopper: An orthogonally persistent operating system. In USENIX [USE94c], pages 289– 312.

DeDecker:1992:HID

[DVPL92] Bart De Decker, Els Van Herreweghen, Frank Piessens, and K. U. Leuven. Heterogeneous intra-domain authentication. In USENIX [USE92m], pages 285–

298. ISBN 1-880446-46-4. LCCN ????

Danielsson:1998:ARL

[DW98a] Johan Danielsson and Assar Westerlund. Arla — A really likeable AFS-client (slides). In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/arla1.ps.

Danielsson:1998:HKW

[DW98b] Johan Danielsson and Assar Westerlund. Heimdal — Kerberos 5 for the world (slides). In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/heimdal1.
ps.

Dahlin:1994:CCU

[DWAP94] Michael D. Dahlin, Randolph Y. Wang, Thomas E. Anderson, and David A. Patterson. Cooperative caching: Using remote client memory to improve file system performance. In USENIX [USE94i], pages 267–280. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/dahlin.html.

Dyker:1991:LVC

[Dyk91] Barbara J. Dyker. LISA V conference trip report. ;login: the USENIX Association newsletter, 16(6):16–??, November/December 1991. CODEN LOGNEM. ISSN 1044-6397.

Dyker:1992:TRN

[Dyk92] Barbara J. Dyker. Trip report on NeXTWORLD. ;login: the USENIX Association newsletter, 17(3):6–??, May/June 1992. CODEN LOGNEM. ISSN 1044-6397.

Deng:1997:OER

[DZL97] Z. Deng, L. Zhang, and J. Liu. An open environment for real-time applications, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/technical.html.

Unpublished poster session at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Epstein:1991:PSA

[E^{+}91] Jeremy Epstein et al. A prototype B3 Trusted X Window System. In Proceedings of the Seventh Annual Computer Security Applications Conference, San Antonio, TX, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 1991. The architecture for TRW’s high assurance multilevel secure X prototype. See also [ES90].

Epps:1999:NSD

[EBG99] Alan Epps, Glenn Bailey, and Douglas Glatz. NFS and SMB data sharing within a heterogeneous environment: a real world study. In USE-NIX [USE99b], pages 37–41. ISBN 1-880446-30-8. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa-nt99/epps.
                                        html.

Emtage:1992:AED

[ED92] Alan Emtage and Peter Deutsch. archie — an electronic directory service for the Internet. In USENIX [USE92j], pages 93– 110.

Ellis:1994:SEG

[ED94] John R. Ellis and David L. Detlefs. Safe, efficient garbage collection for C++. In USE-NIX [USE94e], pages 143–177. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        c++94/ellis.html.

Edelson:1992:SPT

[Ede92]  Daniel R. Edelson.  Smart

pointers: They’re smart, but they’re not pointers. In USENIX [USE92o], pages 1–19. ISBN 1-880446-45-6. LCCN ????

Edelson:1993:FIF

[Ede93] Daniel R. Edelson. Fault interpretation: Fine-grain monitoring of page accesses. In USENIX [USE93l], pages 395–

403. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/

publications/library/proceedings/
sd93/.

Eriksson:1996:IIM

[EFJ96] Joakim Eriksson, Niclas Finne, and Sverker Janson. Information and interaction in MarketSpace — towards an open agent-based market infrastructure. In USENIX [USE96h], pages 271–277. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/finne.html.

Evard:1997:SAS

[EG97] Rémy Evard and Michail Gomberg. Systems administration of scientific computing on NT. In USENIX [USE97a], page ?? ISBN ???? LCCN

???? URL http://www.mcs.
anl.gov/home/evard/talks/
lisa-nt-cp6;   http://www.
usenix.org/publications/library/
proceedings/nt-sysadmin97/
evard.html.

Eirich:1994:BTF

[Eir94] Thomas Eirich. Beam: a tool for flexible software update. In USENIX [USE94h], pages 75–82. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/eirich.html.

Eykholt:1992:BMM

[EKB^{+}92] J. R. Eykholt, S. R. Kleiman, S. Barton, R. Faulkner, D. Stein, M. Smith, A. Shivalingiah, J. Voll, M. Weeks, and D. Williams. Beyond multiprocessing: Multithreading the System V Release 4 kernel. In USENIX [USE92g], pages 11–

18. ISBN 1-880446-44-8. LCCN

QA 76.76 O63 U83 1992.

Elling:1992:USS

[EL92] Richard Elling and Matthew Long. user-setup: a system for custom configuration of user environments, or helping users help themselves. In USENIX [USE92e], pages 215–223. ISBN 1-880446-47-2. LCCN ????

Evard:1994:SSE

[EL94] Remy Evard and Robert Leslie. Soft: a software environment abstraction mechanism. In USE-NIX [USE94h], pages 65–74. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/leslie.html.

Ellison:1996:EIC

[Ell96] Carl Ellison. Establishing identity without certification authorities. In USENIX [USE96i], pages 67–76. ISBN 1-880446-79-

0. LCCN QA76.9.A25 U83 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec96/ellison.html.

Elliott:1997:MIM

[Ell97] Conal Elliott. Modeling interactive 3D and multimedia animation with an embedded language. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/elliott.html.

Elbaum:1999:IDT

[EM99] Sebastian Elbaum and John C.

Munson. Intrusion detection through dynamic software measurement. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/elbaum.html.

Ellson:1996:TTE

[EN96] John Ellson and Stephen North.

TclDG — A Tcl extension for dynamic graphs. In USE-NIX [USE96e], pages 37–47. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/ellson.html.

Engquist:1991:DUB

[Eng91] Jim Engquist. A database for UNIX backup. In USENIX [USE91f], pages 89–96.

English:1992:CIS

[Eng92] Robert M. English. A case for intelligent storage devices. In USENIX [USE92h], pages 147–148. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Engst:1993:SBR

[Eng93] Tonya Engst. SAGE book review: Help! The Art of Computer Technical Support. ;login: the USENIX Association newsletter, 18(1):16–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Engler:1997:IAS

[Eng97]  Dawson R. Engler.  Incor-

porating application semantics and control into compilation. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        dsl97/engler.html.

Edelson:1991:CCC

[EP91] Daniel Edelson and Ira Pohl. Copying collector for C++. In USENIX [USE91k], pages 85–

102. LCCN QA76.73.C153 U27

1991.

Eggert:1992:IFS

[EP92] Paul R. Eggert and D. Stott Parker. An intensional file system. In USENIX [USE92h], pages 145–146. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Eggert:1993:FSU

[EP93] Paul R. Eggert and D. Stott Parker. File systems in user space. In USENIX [USE93l], pages 229–240. ISBN 1-880446-48-0. LCCN QA 76.76 O63

                                        U84 1993. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/sd93/.

Epstein:1992:TCU

[Eps92] Jeremy Epstein. Is there a C2 UNIX system in the house? In USENIX [USE92m], pages 243–

252. ISBN 1-880446-46-4. LCCN ????

Epstein:1990:TXW

[ES90] Jeremy Epstein and Marvin Shugerman. A trusted X Window System server for trusted Mach. In USENIX [USE90e], pages 141–156. This paper describes the initial architecture of the Trusted X Window System prototype developed at TRW. This paper was superseded by the paper at the Seventh Annual Computer Security Applications Conference [E^{+}91].

English:1992:LSO

[ES92] Robert M. English and Alexander A. Stepanov. Loge: a selforganizing disk controller. In USENIX [USE92j], pages 237– 252.

Eustace:1995:AFI

[ES95] Alan Eustace and Amitabh Srivastava. ATOM: a flexible interface for building high performance program analysis tools. In USENIX [USE95e], pages 303–314. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/eustace.html.

Endo:1997:MWN

[ES97] Yasuhiro Endo and Margo I.

Seltzer. Measuring Windows NT

— possibilities and limitations. In USENIX [USE97f], pages 61–66. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt97/endo.html.

Esperanca:1997:TOW

[Esp97]  Claudio Esperança.  A Tk

OpenGL widget. In USE-NIX [USE97d], pages 89–

95. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/esperanca.html.

Essick:1990:EBF

[Ess90] Raymond B. Essick. An eventbased fair share scheduler. In USENIX [USE90h], pages 147–

162. LCCN QA76.8.U65 U82

1990.

Eisler:1996:SMI

[ESS96] Mike Eisler, Roland J. Schemers, and Raj Srinivasan. Security mechanism independence in ONC RPC. In USENIX [USE96i], pages 51–

65. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec96/eisler.html.

Evans:1993:OUR

[ESSB93] Steve Evans, Bart Smaalders, Dave Singleton, and Jeff Bonwick. Optimizing UNIX resource scheduling for user interaction. In USENIX [USE93g], pages 205–218. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/evans.html.

Evans:1993:HS

[Eva93a] Paul Evans. History of SAGE.

;login: the USENIX Association newsletter, 18(2):22–24, March/ April 1993. CODEN LOGNEM. ISSN 1044-6397.

Evard:1993:CNC

[Eva93b]  Rémy Evard.  Collabora-

tive networked communication: MUDs as systems tools. In USENIX [USE93f], pages 1–8. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/evard.html.

Evans:1994:HJS

[Eva94a] Paul Evans. Highlights from the June SAGE board. ;login: the USENIX Association newsletter, 19(4):18–??, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Evard:1994:MEG

[Eva94b] Remy Evard. Managing the ever-growing to do list. In USENIX [USE94h], pages 111–

116. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/evard2.html.

Evard:1994:TRE

[Eva94c] Remy Evard. Tenwen: The re-engineering of A computing environment. In USE-NIX [USE94h], pages 37–46. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/evard.html.

Evans:1995:AMS

[Eva95a] Paul Evans. Another maxim of system administration. ;login: the USENIX Association newsletter, 20(3):28–29, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Evans:1995:AP

[Eva95b] Paul Evans. Aspects of professionalism. ;login: the USE-NIX Association newsletter, 20 (6):32–??, December 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Evans:1995:CNN

[Eva95c] Paul Evans. Community news: a new form of professional recognition. ;login: the USENIX Association newsletter, 20(5):22– ??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Evans:1995:OP

[Eva95d] Paul Evans. Opinion: Professionalism. ;login: the USE-NIX Association newsletter, 20 (5):21–22, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Evans:1996:LPa

[Eva96a] Paul Evans. Letter from the President. ;login: the USE-NIX Association newsletter, 21 (4):18–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Evans:1996:LPb

[Eva96b] Paul Evans. Letter from the President. ;login: the USENIX Association newsletter, 21(6):7– ??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Evans:1996:LSP

[Eva96c] Paul Evans. Letter from the SAGE President. ;login: the USENIX Association newsletter, 21(3):13–??, June 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Evans:1996:YAC

[Eva96d] Paul Evans. Your 30-amp circuits. ;login: the USENIX Association newsletter, 21(2):26–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Evard:1997:AUS

[Eva97]  Remy Evard.  An analysis

of UNIX system configuration. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/20.evard.html.

Emer:1990:IIA

[EW90] Joel S. Emer and William E.

Weihl. Integrated interactive access to heterogeneous distributed services. In USE-NIX [USE90h], pages 101–116. LCCN QA76.8.U65 U82 1990.

Endo:1996:ULE

[EWCS96] Yasuhiro Endo, Zheng Wang, J. Bradley Chen, and Margo I.

Seltzer. Using latency to evaluate interactive system performance. In USENIX [USE96f], pages 185–199. ISBN 1-880446-82-0. LCCN QA76.76.O63

S96 1996. URL http://www.
usenix.org/publications/library/
proceedings/osdi96/endo.html.

Farmer:1992:SBR

[Far92] Dan Farmer. SAGE book reviews: Practical UNIX security. ;login: the USENIX Association newsletter, 17(6):13–14, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1994:RFI

[Far94] Rik Farrow. Review: Firewalls and Internet Security: Repelling the Wily Hacker. ;login: the USENIX Association newsletter, 19(3):32–??, May/ June 1994. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1995:Ma

[Far95a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 20(4):39–41, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1995:Mb

[Far95b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 20(5):34–35, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1995:MLI

[Far95c] Rik Farrow. Musings from LISA IX. ;login: the USENIX Association newsletter, 20(6):19– 21, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1995:RFU

[Far95d] Rik Farrow. Report on the Fifth USENIX UNIX security symposium. ;login: the USE-NIX Association newsletter, 20 (4):6–10, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1995:ROI

[Far95e] Rik Farrow. Review: O’Reilly’s Internet Talk Radio Tapes. ;login: the USENIX Association newsletter, 20(2):55–??, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:Ma

[Far96a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 21(1):26–27, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:Mb

[Far96b] Rik Farrow. Musings. ;login:

the USENIX Association newsletter, 21(2):43–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:Mc

[Far96c] Rik Farrow. Musings. ;login:

the USENIX Association newsletter, 21(3):32–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:Md

[Far96d] Rik Farrow. Musings. ;login:

the USENIX Association newsletter, 21(4):41–??, August 1996.

CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:ML

[Far96e] Rik Farrow. Musings from LISA 10. ;login: the USENIX Association newsletter, 21(6):27– 29, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:REJ

[Far96f] Rik Farrow. Review: Exploring Java. ;login: the USE-NIX Association newsletter, 21 (5):46–??, October 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Farrow:1996:UJa

[Far96g] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 21(5):39–42, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:UJb

[Far96h] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 21(6):24–26, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:UJN

[Far96i] Rik Farrow. Using Java: a new series of articles about Java. ;login: the USENIX Association newsletter, 21(5):39–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:BRW

[Far97a] Rik Farrow. Book review: Web Security Sourcebook; Web Security and Commerce. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/rubin.html.

Farrow:1997:Mb

[Far97b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 22(1):24–25, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:Mc

[Far97c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 22(4):25–27, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:Md

[Far97d] Rik Farrow. Musings. ;login:

the USENIX Association newsletter, 22(5):19–21, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:Me

[Far97e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/musings.html.

Farrow:1997:Mf

[Far97f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-12/musings.html.

Farrow:1997:MA

[Far97g] Rik Farrow. Musings from Anaheim. ;login: the USENIX Association newsletter, 22(2):35–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:SRS

[Far97h] Rik Farrow. Software review: SSH for Windows. ;login: the USENIX Association newsletter, 22(2):76–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:UJa

[Far97i]  Rik Farrow.  Using Java.

;login: the USENIX Association newsletter, 22(2):19– 23, April 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/java/javaindex.
                                        html.

Farrow:1997:UJb

[Far97j]  Rik Farrow.  Using Java.

;login: the USENIX Association newsletter, 22(4):33– 36, May 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/java/javaindex.
                                        html.

Farrow:1997:UJc

[Far97k] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 22(5):32–35, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1998:Mf

[Far98a]  Captain Farrow.  Musings.

;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
musings.html.

Farrow:1998:Ma

[Far98b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-2/musings.html.

Farrow:1998:Mb

[Far98c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-4/musings.html.

Farrow:1998:Mc

[Far98d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
musings.html.

Farrow:1998:Md

[Far98e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-8/musings.html.

Farrow:1998:Me

[Far98f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/musings.html.

Farrow:1998:UJ

[Far98g]  Rik Farrow.  Using Java.

;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/javaindex.
html;   http://www.usenix.
org/publications/java/usingjava9.
html.

Farrow:1999:Ma

[Far99a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-2/musings.html.

Farrow:1999:Mb

[Far99b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-4/musings.html.

Farrow:1999:Mc

[Far99c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-6/features/musings.html.

Farrow:1999:Md

[Far99d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-8/features/musings.html.

Farrow:1999:Me

[Far99e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-10/features/musings.
html.

Farrow:1999:Mf

[Far99f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-12/features/musings.
html.

Farrow:1999:TRS

[Far99g] Rik Farrow. A tribute to Rich Stevens. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
tribute.html.

Fiuczynski:1996:EPA

[FB96] Marc E. Fiuczynski and Brian N. Bershad. An extensible protocol architecture for applicationspecific networking. In USE-NIX [USE96k], pages 55–64. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/mef.html.

Feitelson:1992:IRT

[FBAE^{+}92] Dror G. Feitelson, Yosi Ben-Asher, Moshe Ben Ezra, Iaakov Exman, Lior Picherski, Larry Rudolph, and Dror Zernik. Issues in run-time support for tightly-coupled parallel processing. In USENIX [USE92l], pages 27–42. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Feng:1996:APR

[FC96] Li-Chi Feng and Ruei-Chuan Chang. Achieving predictable response time with an intelligent file system updater. In USENIX [USE96c], pages 217–237.

Feeley:1994:ICR

[FCNL94] Michael J. Feeley, Jeffrey S. Chase, Vivek R. Narasayya, and Henry M. Levy. Integrating coherency and recovery in distributed systems. In USENIX [USE94i], pages 215–227. ISBN 1-880446-66-9.

LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/feeley.html.

Flick:1998:UTT

[FD98] C. Allen Flick and James S.

Dixson. Using TCL/TK for an automatic test engine. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/flick.html.

Fels:1997:IIG

[FEBM97] Sidney Fels, Silvio Esser, Armin Bruderlin, and Kenji Mase. InvenTcl: Interpretive 3D graphics using Open Inventor and Tcl/Tk/[incr Tcl]. In USE-NIX [USE97d], pages 185–

186. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_esser.html.

Fennessy:1996:LS

[Fen96] Quentin Fennessy. Linkage and symbolism. ;login: the USE-NIX Association newsletter, 21 (4):43–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Ferguson:1995:AOE

[Fer95] John Ferguson. Avoiding openended questions. ;login: the USENIX Association newsletter, 20(6):32–33, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Feldman:1990:CPN

[FG90] Stuart I. Feldman and W. Morven Gentleman. Controversy: Portability — A no longer solved problem. In USENIX [USE90b], pages 359–380.

Faulkner:1991:PFS

[FG91] Roger Faulkner and Ron Gomes. The process file system and process model in UNIX System V. In USENIX [USE91i], pages 243–252. LCCN QA 76.76 O63

                                        U84 1992. URL http://obits.
                                        mlive.com/obituaries/grandrapids/
                                        obituary.aspx?pid=180588279;
                                         http://thenewstack.io/remembering-
                                        roger-faulkner/;   https:/
                                        /www.usenix.org/memoriam-
                                        roger-faulkner;   https://
                                        www.usenix.org/sites/default/
                                        files/usenix_winter91_faulkner.
                                        pdf.

Forin:1991:SM

[FGB91] Alessandro Forin, David Golub, and Brian Bershad. An I/O system for Mach 3.0. In USE-NIX [USE91h], pages 163–176. LCCN QAX 27.

Finlayson:1993:VVE

[FHG93] Ross Finlayson, Mark D. Honnecke, and Steven Goldberg. From V to Vanguard: The evolution of a distributed objectoriented microkernel interface. In USENIX [USE93k], pages 27–30. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/finlayson.html.

Fowler:1993:ULR

[FHKR93] Glenn S. Fowler, Yennun Huang, David Korn, and Herman Rao. A user-level replicated file system. In USE-NIX [USE93g], pages 279–290. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/.

Ford:1996:MMR

[FHL^{+}96] Bryan Ford, Mike Hibler, Jay Lepreau, Patrick Tullman, Godmar Back, and Steven Clawson. Microkernels meet recursive virtual machines. In USENIX [USE96f], pages 137–

151. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/hibler.html.

Ford:1999:IEM

[FHL^{+}99] Bryan Ford, Mike Hibler, Jay Lepreau, Roland McGrath, and Patrick Tullmann. Interface and execution models in the Fluke kernel. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/lepreau.
html.

Forin:1999:HPD

[FHLW99] Alessandro Forin, Galen Hunt, Li Li, and Yi-Min Wang. Highperformance distributed objects over system area networks. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usenix-nt99/forin.
                                        html.

Fiander:1994:RS

[Fia94] David J. Fiander. Review: Sendmail. ;login: the USENIX Association newsletter, 19(2): 48–??, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Fiander:1995:CND

[Fia95] David J. Fiander. Caldera network desktop. ;login: the USE-NIX Association newsletter, 20 (6):51–52, December 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Finkel:1990:DBM

[Fin90] David Finkel. Developing benchmarks to measure the performance of the Mach operating system. In USENIX [USE90e], pages 83–100.

Fine:1991:CS

[Fin91] Thomas Fine. A console server. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Finke:1994:APC

[Fin94a] Jon Finke. Automating printing configuration. In USE-NIX [USE94h], pages 175–

183. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/finke2.html.

Finke:1994:MUW

[Fin94b] Jon Finke. Monitoring usage of workstations with a relational database. In USE-NIX [USE94h], pages 149–

157. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/finke.html.

Finke:1995:SAG

[Fin95a] Jon Finke. SQL 2 HTML: Automatic generation of HTML database schemas. In USE-NIX [USE95h], pages 133–

138. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/finke.html.

Finke:1995:WWM

[Fin95b] Jon Finke. What we mean by trust. ;login: the USENIX Association newsletter, 20(5):29– 31, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Finke:1996:IWP

[Fin96] Jon Finke. Institute white pages as a system administration problem. In USE-NIX [USE96j], pages 233–

240. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/finkej.html.

Fink:1997:ID

[Fin97a]  Bob Fink.  IPv6 deploy-

ment on the 6bone, 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/invited_talks/fink/

fink.ps. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Finke:1997:ASC

[Fin97b]  Jon Finke.  Automation

of site configuration management. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/18.finke.html.

Finke:1997:MAU

[Fin97c] Jon Finke. Monitoring application use with license server logs. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/03.finke.html.

Finke:1999:SMN

[Fin99a] Jon Finke. Servicetrak meets NLOG/NMAP. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

                                        URL http://db.usenix.org/
                                        publications/library/proceedings/
                                        lisa99/finke.html.

Finke:1999:UIP

[Fin99b] Jon Finke. University issues panel, 1999. URL http://db.

usenix.org/publications/library/
proceedings/lisa99/practicum/

finke.ppt. Unpublished practicum talk at LISA XIII.

Fisher:1996:CUV

[Fis96a] Idajean M. Fisher. Confessions of a USENIX volunteer. ;login: the USENIX Association newsletter, 21(6):40–??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Fisher:1996:SJD

[Fis96b] Idajean M. Fisher. The SAGE job descriptions for system administrators as a tool. ;login: the USENIX Association newsletter, 21(3):24–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Fisk:1996:AAH

[Fis96c]  Michael Fisk.  Automating

the administration of heterogeneous LANs. In USE-NIX [USE96j], pages 181–

186. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/mfisk.html.

Fisher:1997:UAM

[Fis97] Idajean M. Fisher. USENIX annual meeting. ;login: the USE-NIX Association newsletter, 22 (2):4–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Fowler:1992:SUL

[FK92] Robert J. Fowler and Leonidas I. Kontothanassis. Supporting user-level exception handling on a multiprocessor microkernel: Experiences with platinum. In USENIX [USE92l], pages 217–232. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Frolund:1998:QSS

[FK98] Svend Frolund and Jari Koistinen. Quality of service specification in distributed object systems design. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/frolund.html.

Frolund:1999:QSA

[FK99] Svend Frølund and Jari Koistinen. Quality of service– aware distributed object systems. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/frolund.html.

Fowler:1994:FBP

[FKSV94] Glenn S. Fowler, David G. Korn, John J. Snyder, and Kiem-Phong Vo. Feature-based portability. In USENIX [USE94k], pages 197–207. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/fowler.feature.html.

Fujimura:1999:DTC

[FKT^{+}99] Ko Fujimura, Hiroshi Kuno, Masayuki Terada, Kazuo Matsuyama, Yasunao Mizuno, and Jun Sekine. Digital-ticket-controlled digital ticket circulation. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/fujimura.
html.

Frantz:1993:OOT

[FL93] William S. Frantz and Charles R.

Landau. Object-oriented transaction processing in the KeyKOS microkernel. In USENIX [USE93k], pages 13–26. ISBN 1-880446-52-9. LCCN QA76.9.A73

U83 1993. URL http://www.
usenix.org/publications/library/
proceedings/micro93/frantz.
html.

Ford:1994:EMM

[FL94] Bryan Ford and Jay Lepreau.

Evolving Mach 3.0 to a migrating thread model. In USENIX [USE94l], pages 97–

114. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/ford.html.

Fried:1996:BHA

[FL96] Benjamin Fried and Andrew Lowry. BigDog: Hierarchical authentication, session control, and authorization for the Web. In USENIX [USE96h], pages 165–172. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec96/index.html.

Fulmer:1998:ANC

[FL98] Robert Fulmer and Alex Levine. AutoInstall for NT: Complete NT installation over the network. In USENIX [USE98e], pages 27–35. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa-nt98/fulmer.html.

Freeh:1994:DFE

[FLA94] Vincent W. Freeh, David K. Lowenthal, and Gregory R. Andrews. Distributed filaments: Efficient fine-grain parallelism on a cluster of workstations. In USENIX [USE94i], pages 201–213. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/freeh.html.

Flanagan:1997:RJD

[Fla97]  David Flanagan.  Review:

JavaScript, the definitive guide (beta Edition). ;login: the USE-NIX Association newsletter, 22 (1):60–61, February 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Fiuczynski:1998:DII

[FLB98] Marc E. Fiuczynski, Vincent K. Lam, and Brian N. Bershad. The design and implementation of an IPv6/IPv4 network address and protocol translator. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/fiuczynski.html.

Fletcher:1992:DNS

[Fle92a] Mark Fletcher. doit: a network software management tool. In USENIX [USE92e], pages 189–

196. ISBN 1-880446-47-2. LCCN ????

Fletcher:1992:NNP

[Fle92b] Mark Fletcher. nlp: a network printing tool. In USENIX [USE92e], pages 245–256. ISBN 1-880446-47-2. LCCN ????

Fleisch:1993:RDS

[Fle93] Brent D. Fleisch. The role of distributed shared memory in future experimental distributed systems. In USE-NIX [USE93j], pages 273–

278. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

Fletcher:1996:L

[Fle96] Mark Fletcher. An LPD for the 90s. In USENIX [USE96j], pages 73–79. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/mark.html.

Flynt:1997:CTR

[Fly97] Clifton Flynt. Coding techniques for reducing code maintenance. In USENIX [USE97d], pages 183–184. ISBN 1-880446-87-1. LCCN QA76.73.T44

T35 1997. URL http://www.
usenix.org/publications/library/
proceedings/tcl97/poster_
flynt.html.

Flynt:1998:TTP

[Fly98a]  Clif Flynt.  Tcl/Tk pro-

gram development tools. In USENIX [USE98f], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998. URL

http://db.usenix.org/publications/
library/proceedings/tcl98/
full_papers/ranganathan_poster/
ranganathan_html/ranganathan.
html.

Flynt:1998:BRT

[Fly98b] Clifton Flynt. Book review: Tcl/Tk Tools. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
tcltk.html.

Flynt:1999:BRC

[Fly99a] Clif Flynt. Book review: C/ C++ Treasure Chest. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-2/treasure_chest.html.

Flynt:1999:TSb

[Fly99b] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
tclsh.html.

Flynt:1999:TSc

[Fly99c] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
features/tclsh.html.

Flynt:1999:TSd

[Fly99d] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
features/tclsh.html.

Flynt:1999:TSe

[Fly99e] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/tclsh.html.

Flynt:1999:TSA

[Fly99f] Clif Flynt. The Tclsh spot:

Adding hypertext links and image display to htmlview.tcl. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/tclsh.html.

Forin:1994:MFU

[FM94] Alessandro Forin and Gerald Malan. An MS-DOS filesystem for UNIX. In USE-NIX [USE94l], pages 337–354. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/forin.html.

Fineberg:1999:RBT

[FM99] Samuel A. Fineberg and Pankaj Mehra. The record-breaking terabyte sort on a Compaq cluster. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/fineberg.html.

Fontana:1991:EDC

[FN91] Mary Fontana and Martin Neath. Experiences in the design of a C++ class library. In USENIX [USE91k], pages 179–

192. LCCN QA76.73.C153 U27 1991.

Fujimura:1998:GPD

[FN98] Ko Fujimura and Yoshiaki Nakajima. General-purpose digital ticket framework. In USENIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http://www.

usenix.org/publications/library/
proceedings/ec98/fujimura.
html.

Faith:1997:KSR

[FNP97] Rickard E. Faith, Lars S. Nyland, and Jan F. Prins. KHEP-ERA: a system for rapid implementation of domain specific languages. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/faith.html.

Furlani:1996:AYM

[FO96] John L. Furlani and Peter W. Osel. Abstract yourself with modules. In USE-NIX [USE96j], pages 193–

203. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/pwo.html.

Forte:1998:SMJ

[For98] Dario Forte. The security model of JDK 1.2. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
jdk.html.

Forte:1999:BUT

[For99a] Dario Forte. Biometrics: Untruths and the truth. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1999-4/biometrics.html.

Forte:1999:IDS

[For99b] Dario Forte. Intrusion-detection systems: Guaranteeing the safety of a network beyond using a firewall. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        intrusion.html.

Fowler:1993:SS

[Fow93] Glenn S. Fowler. The shell as a service. In USE-NIX [USE93g], pages 267–277. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/fowler.html.

Fowler:1994:CFF

[Fow94] Glenn Fowler. cql — A flat file database query language. In USENIX [USE94l], pages 11–21. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/fowler.html.

Francisco:1990:MSS

[FP90] Dodi Francisco and Lois C. Price. Multiple site source reconciliation. In USENIX [USE90h], pages 331–340. LCCN QA76.8.U65 U82 1990.

Fall:1993:EKD

[FP93] Kevin Fall and Joseph Pasquale. Exploiting in-kernel data paths to improve I/O throughput and CPU availability. In USENIX [USE93l], pages 327–333. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Farmer:1998:T

[FPA98] Dan Farmer, Brad Powell, and Matthew Archibald. TI-TAN. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/farmer.html.

Fortier:1991:EDD

[FPSW91] Paul J. Fortier, David V.

Pitts, John C. Sieg, Jr., and C. Thomas Wilkes. Experiences with distributed data management in real-time C3 systems. In USENIX [USE91j], pages 19–34. ISBN ???? LCCN ????

Fine:1990:CS

[FR90] Thomas A. Fine and Steven M.

Romig. A console server. In USENIX [USE90f], pages 97–

100. LCCN QA76.76.O63 L37 1990.

Franklin:1992:TVM

[Fra92] James W. Franklin. Tiled virtual memory for UNIX. In USE-NIX [USE92g], pages 99–106.

ISBN 1-880446-44-8. LCCN QA

76.76 O63 U83 1992.

Franklin:1993:P

[Fra93] Wm. Randolph Franklin. PEM. ;login: the USENIX Association newsletter, 18(6):33–??, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Freedman:1991:EBP

[Fre91] Dan Freedman. Experience building a process migration subsystem for UNIX. In USE-NIX [USE91i], pages 349–356. LCCN QA 76.76 O63 U84 1992.

Fredrick:1997:ULL

[Fre97] Michael Fredrick. Utilizing lowlevel NTFS disk structures for rapid disk searching. In USE-NIX [USE97a], page ?? ISBN ???? LCCN ????

Fritchie:1997:CNF

[Fri97] Scott Lystig Fritchie. The Cyclic News Filesystem: Getting INN to do more with less. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/14.fritchie.html.

Farmer:1990:CSC

[FS90] Dan Farmer and Eugene H. Spafford. The Cops security checker system. In USE-NIX [USE90g], pages 165–170. LCCN QA76.8.U65 U81 1990.

Ferreira:1994:GCD

[FS94] Paulo Ferreira and Marc Shapiro.

Garbage collection and DSM consistency. In USENIX [USE94i], pages 229–241. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994. URL http://www.

usenix.org/publications/library/
proceedings/osdi/ferr.html.

Ford:1996:CIS

[FS96] Bryan Ford and Sai R. Susarla.

CPU inheritance scheduling. In USENIX [USE96f], pages 91–105. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/ford.html.

Fujikawa:1991:MPS

[FSM^{+}91] Kazutoshi Fujikawa, Shinji Shimojo, Toshio Matsuura, Shojiro Nishio, and Hideo Miyahara. Multimedia presentation system “harmony” with temporal and active media. In USENIX [USE91g], pages 75–

94. LCCN QA 76.76 O63 U83 1991.

Fyfe:1991:CSS

[FST91] Alastair Fyfe, Ivan Soleimanipour, and Vijay Tatkar. Compiling from saved state: Fast incremental compilation with traditional UNIX compilers. In USENIX [USE91i], pages 161–

172. LCCN QA 76.76 O63 U84

Fernandez:1999:DSD

[FST99] Mary Fernández, Dan Suciu, and Igor Tatarinov. Declarative specification of dataintensive Web sites. In USE-NIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/fernandez.html.

Fowler:1994:EUS

[FSV94] Glenn S. Fowler, John J. Snyder, and Kiem-Phong Vo. Enduser systems, reusability and high-level design. In USE-NIX [USE94k], pages 101–

118. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/fowler.enduser.html.

Finn:1999:PNP

[FT99] Gregory G. Finn and Joe Touch. The Personal Node (PN). In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
es99/finn.html.

Fuchs:1997:DSL

[Fuc97] Matthew Fuchs. Domain specific languages for ad hoc distributed applications. In USE-NIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/fuchs.html.

Fulton:1996:IPP

[Ful96] Jim Fulton. Introduction to the Python programming language. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

Furlani:1991:MPF

[Fur91] John L. Furlani. Modules: Providing a flexible user environment. In USENIX [USE91f], pages 141–152.

Futakata:1995:PCM

[Fut95] Atsushi Futakata. Patch control mechanism for large scale software. In USENIX [USE95h], pages 213–220 (or 213–219??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/futakata.html.

Forslund:1990:EWD

[FWF^{+}90] David W. Forslund, Charles Wingate, Peter Ford, J. Stephen Junkins, Jeffrey Jackson, and Stephen C. Pope. Experiences in writing a distributed particle simulation code in C++. In USENIX [USE90j], pages 177– 190.

Frankel:1998:RMU

[FY98] Yair Frankel and Moti Yung.

Risk management using threshold RSA cryptosystems. ;login: the USENIX Association newsletter, 23(3):??, May 1998.

CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-5/frankel.html. Special

issue on security.

Griffioen:1994:RFS

[GA94] Jim Griffioen and Randy Appleton. Reducing file system latency using a predictive approach. In USE-NIX [USE94j], pages 197–207. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/griffioen.html.

Gould:1996:ORL

[GADG96] Ed Gould, Dan Appelman, Mitch Dembin, and Mike Godwin. Opinions on recent legal decisions, 1996.

URL http://www.usenix.org/
publications/library/proceedings/

sd96/. Unpublished invited talk panel discussion at the USENIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Gaines:1994:DLI

[Gai94] R. Stockton Gaines. Dixie languages and interpreter issues. In USENIX [USE94k], pages 191–196. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/gaines.html.

Gallmeister:1993:DH

[Gal93] Bill Gallmeister. DOS digs in its heels. ;login: the USENIX Association newsletter, 18(1):26–??, January/February 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Galvin:1996:PKD

[Gal96] James M. Galvin. Public key distribution with secure DNS. In USENIX [USE96i], pages 161–170. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/galvin.html.

Gamble:1993:IEC

[Gam93]  Todd Gamble.  Implement-

ing execution controls in Unix. In USENIX [USE93f], pages 237–242. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/gamble.html.

Garg:1990:PSM

[Gar90a] Arun Garg. Parallel STREAMS:

a multi-processor implementation. In USENIX [USE90h], pages 163–176. LCCN QA76.8.U65 U82 1990.

Gart:1990:ABU

[Gar90b] Mitchell Gart. Ada and binary UNIX standards. In USE-NIX [USE90h], pages 353–362. LCCN QA76.8.U65 U82 1990.

Garberson:1998:LEP

[Gar98] John D. Garberson. Letter to the Editor: Programming Python. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Gassaway:1996:SS

[Gas96] Timothy Gassaway. From the SAGE secretary. ;login: the USENIX Association newsletter, 21(4):23–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Gassaway:1998:YPM

[Gas98] Tim Gassaway. Are you preparing for the millenium? ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

Gautron:1991:PEC

[Gau91]  Philippe Gautron.  Porting

and extending the C++ task system with the support of lightweight processes. In USE-NIX [USE91k], pages 135–146. LCCN QA76.73.C153 U27 1991.

Gautron:1992:AMB

[Gau92] Philippe Gautron. An assertion mechanism based on exceptions. In USENIX [USE92k], pages 245–262. ISBN 1-880446-45-6. LCCN ????

Galvin:1992:SAU

[GB92] James M. Galvin and David M. Balenson. Security aspects of a UNIX PEM implementation. In USENIX [USE92m], pages 119–

131. ISBN 1-880446-46-4. LCCN ????

Gribble:1997:SDI

[GB97] Steven D. Gribble and Eric A.

Brewer. System design issues for Internet middleware services: Deductions from a large client trace. In USE-NIX [USE97h], pages 207–218. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/gribble.html.

Girardin:1998:VAM

[GB98] Luc Girardin and Dominique Brodbeck. A visual approach for monitoring logs. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
girardin.html.

Ginsberg:1993:UMC

[GBB93] Michael Ginsberg, Robert V.

Baron, and Brian N. Bershad. Using the Mach communication primitives in X11. In USENIX [USE93h], pages 103–110. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/ginsberg.html.

Goldick:1995:MAA

[GBK^{+}95] Jonathan S. Goldick, Kathy Benninger, Christopher Kirby, Christopher Maher, and Bill Zumach. Multi-resident AFS: An adventure in mass storage. In USENIX [USE95e], pages 47–58. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neworl/zumach.html.

Golding:1995:IS

[GBS^{+}95] Richard Golding, Peter Bosch, Carl Staelin, Tim Sullivan, and John Wilkes. Idleness is not sloth. In USE-NIX [USE95e], pages 201–212. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neworl/wilkes.html.

Grass:1990:CIA

[GC90] Judith E. Grass and Yih-Farn Chen. The C++ information abstractor. In USENIX [USE90j], pages 265–277 (or 265–278??).

Greenwald:1996:SBN

[GC96] Michael Greenwald and David Cheriton. The synergy between non-blocking synchronization and operating system structure. In USENIX [USE96f], pages 123–136. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi96/greenwald.html.

Grubb:1998:SSS

[GC98] Michael Fleming Grubb and Rob Carter. Single sign-on and the system administrator. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/grubb.html.

Golub:1991:MDM

[GD91] David B. Golub and Richard P.

Draves. Moving the default memory manager out of the Mach kernel. In USE-NIX [USE91h], pages 177–188. LCCN QAX 27.

Geer:1996:TMC

[GD96a] Daniel E. Geer and Donald T.

Davis. Token-mediated certification and electronic commerce. In USENIX [USE96h], pages 13–22. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/geer.html.

Goel:1996:LDD

[GD96b] Shantanu Goel and Dan Duchamp. Linux device driver emulation in Mach. In USE-NIX [USE96k], pages 65–73. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/goel.html.

Golub:1990:UAP

[GDFR90] David Golub, Randall Dean, Alessandro Forin, and Richard Rashid. UNIX as an application program. In USENIX [USE90g], pages 87–96. LCCN QA76.8.U65 U81 1990.

Geer:1990:UU

[Gee90] Daniel E. Geer, Jr. An Ultrix 4.0 uniserver. In USENIX [USE90e], page ??

Geer:1994:BBS

[Gee94] Dan Geer. Breaking into banks: Security lessons learned from financial services. In USENIX [USE94f], page ?? ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/lisa94/index.
                                        html.

Geer:1998:DFW

[Gee98] Daniel E. Geer, Jr. Driving the future: Why the financial community rules. ;login: the USE-NIX Association newsletter, 23 (3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-5/

geer.html. Special issue on se-

curity.

Geer:1999:PRW

[Gee99] Dan Geer. Privacy in the real world. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-10/
                                        features/privacy.html.

Gomberg:1998:CLS

[GES98] Michail Gomberg, Rémy Evard, and Craig Stacey. A comparison of large-scale software installation methods on NT and UNIX. In USENIX [USE98e], pages 37–47. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/gomberg.html.

Galvin:1995:MOS

[GF95] James M. Galvin and Mark S.

Feldman. MIME object security services: Issues in a multi-user environment. In USENIX [USE95f], pages 217–

223. ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
galvin.html.

Gopinath:1991:HAL

[GG91] Prabha Gopinath and Rajiv Gupta. A hybrid approach to load balancing in distributed systems. In USENIX [USE91j], pages 133–148. ISBN ???? LCCN ????

Garbinato:1997:USD

[GG97] Benoît Garbinato and Rachid Guerraoui. Using the strategy design pattern to compose reliable distributed protocols. In USENIX [USE97g], pages 221–232. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/garbinato.html. See

also [LH97, Rie97, WW97].

Gil:1998:CTS

[GG98] Joseph Gil and Zvi Gutterman. Compile time symbolic derivation with C++ templates. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots98/gil.html.

Gerrity:1991:CWS

[GGI^{+}91] G. W. Gerrity, A. Goscinski, J. Indulska, W. Toomey, and W. Zhu. Can we study design issues of distributed operating systems in a generalized way? In USENIX [USE91j], pages 301–

320. ISBN ???? LCCN ????

Goyal:1996:HCS

[GGV96] Pawan Goyal, Xingang Guo, and Harrick M. Vin. A hierarchical CPU scheduler for multimedia operating system. In USENIX [USE96f], pages 107–121. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi96/vin.html.

Goldberg:1999:NJ

[GGWB99] Ian Goldberg, Steven D. Gribble, David Wagner, and Eric A. Brewer. The Ninja Jukebox. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/goldberg.html.

Golan:1993:DVH

[GH93] Michael Golan and David R.

Hanson. DUEL — A very high-level debugging language. In USENIX [USE93l], pages 107–117. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Gassaway:1998:SB

[GH98] Tim Gassaway and Helen Harrison. Speakers bureau. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Guy:1990:IFR

[GHM^{+}90] Richard G. Guy, John S. Heidemann, Wai Mak, Thomas W. Page, Jr., Gerald J. Popek, and Dieter Rothmeir. Implementation of the Ficus Replicated File System. In USE-NIX [USE90g], pages 63–

72. LCCN QA76.8.U65 U81

1990. URL ftp://shemp.cs.
ucla.edu/pub/ficus/usenix\
_summer\_90.ps.Z; local -
--   ficus.ps.  FTP -

shemp.cs.ucla.edu:/pub/ficus/usenix summer 90.ps.Z; local - ficus.ps.

Galyean:1990:TMW

[GHN90] Tinsley Galyean, Trent Hein, and Evi Nemeth. Trouble-MH: a work-queue management package for a ¿3 ring circus. In USENIX [USE90f], pages 93–96. LCCN QA76.76.O63 L37 1990.

Galyean:1991:TMW

[GHN91] Tinsley Galyean, Trent Hein, and Evi Nemeth. Trouble-MH: a work-queue management system for A ¿3-ring circus. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Gill:1990:RUP

[Gil90] David Gill. Roles for users and privileges for system processes: High trust mechanisms for low trust systems. In USENIX [USE90i], pages 61–62. LCCN QA 76.9 A25 U55 1990.

Gillies:1993:RLU

[Gil93] Betsy Gillies. Review: Learning the UNIX Operating System. ;login: the USENIX Association newsletter, 18(6):48–??, November/December 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Gilmore:1996:YDN

[Gil96] John Gilmore. Is your domain name system secure? ;login: the USENIX Association newsletter, 21(5):31–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Girard:1999:WSP

[Gir99a]  Pierre Girard.  Which se-

curity policy for multiapplication Smart Cards? In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/smartcard99/
                                        girard.html.

Girardin:1999:ENI

[Gir99b]  Luc Girardin.  An eye on

network intruder-administrator shootouts. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/girardin.html.

Ganger:1997:EIE

[GK97] Gregory R. Ganger and M. Frans Kaashoek. Embedded inodes and explicit grouping: Exploiting disk bandwidth for small files. In USENIX [USE97e], pages 1–17. ISBN 1-880446-84-7. LCCN QA76.76.O63

U88 1997. URL http://www.
usenix.org/publications/library/
proceedings/ana97/ganger.
html;   http://www.usenix.
org/publications/library/
proceedings/ana97/summaries/
ganger.html.  Best paper

award.

Goldschlag:1999:BCC

[GK99] David M. Goldschlag and David W. Kravitz. Beyond cryptographic conditional access. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/goldschlag.html.

Gamsa:1999:TML

[GKAS99] Ben Gamsa, Orran Krieger, Jonathan Appavoo, and Michael Stumm. Tornado: Maximizing locality and concurrency in a shared memory multiprocessor operating system. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/gamsa.
html.

Grass:1990:OOR

[GKS90] Judith E. Grass, Chandra Kintala, and Ravi Sethi. Objectoriented redesign using C++: Experience with parser generators. In USENIX [USE90j], pages 75–85 (or 75–86??).

Gunaseelan:1993:DOT

[GL93] L. Gunaseelan and Richard J. LeBlanc, Jr. Debugging objects and threads in a shared memory system. In USE-NIX [USE93j], pages 175–

194. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/gunaseelan.html.

Goodwin:1997:SOA

[GL97] David Goodwin and P. Geoffrey Lowney. Spike: An optimizer for Alpha/NT executables, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/technical.html.

Unpublished poster session at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Guyer:1999:ALO

[GL99] Samuel Z. Guyer and Calvin Lin. An annotation language for optimizing software libraries. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/guyer.html.

Glass:1996:SA

[Gla96] Graham Glass. STL in action. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

Guillemont:1991:SGM

[GLOR91] Marc Guillemont, Jim Lipkis, Doug Orr, and Marc Rosier. A second-generation micro-kernel based UNIX: Lessons in performance and compatibility. In USENIX [USE91i], pages 13–22. LCCN QA 76.76 O63 U84 1992.

Gettys:1993:ANT

[GLP^{+}93] James Gettys, Thomas M. Levergood, Andrew C. Payne, Lawrence C. Stewart, and G. Winfield Treese. AudioFile: a network-transparent audio server. In USE-NIX [USE93g], pages 219–236. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/gettys.html.

Golding:1994:RDB

[GLW94] Richard A. Golding, Darrell D. E. Long, and John Wilkes. The refdbms distributed bibliographic database system. In USENIX [USE94l], pages 47–62. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/golding.html.

Graham:1995:ABP

[GLW95] Susan L. Graham, Steven Lucco, and Robert Wahbe. Adaptable binary programs. In USENIX [USE95e], pages 315–

325. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neworl/wahbe.html.

Gokhale:1993:ISD

[GM93] Maya B. Gokhale and Ronald G. Minnich. An implementation of the shared data formats standard for distributed shared memories. In USE-NIX [USE93j], pages 109–

121. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sedms4/gokhale.html.

Gopal:1999:ICB

[GM99] Burra Gopal and Udi Manber. Integrating content-based access mechanisms with hierarchical file systems. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/gopal.
html.

Green:1999:ATD

[GMNR99] John Green, David Marchette, Stephen Northcutt, and Bill Ralph. Analysis techniques for detecting coordinated attacks and probes. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25

W67 1999. URL http://www.
usenix.org/publications/library/
proceedings/detection99/green.
html.

Gong:1997:GBS

[GMPS97] Li Gong, Marianne Mueller, Hemma Prafullchandra, and Roland Schemers. Going beyond the sandbox: An overview of the new security architecture in the Java Development Kit 1.2. In USE-NIX [USE97h], pages 103–112. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/gong.html.

Golub:1993:MER

[GMR93] David Golub, Ravi Manikundalam, and Freeman L. Rawson, III. MVM — an environment for running multiple DOS, Windows and DPMI programs on the microkernel. In USENIX [USE93h], pages 173–190. ISBN 1-880446-49-9.

LCCN QA 76.76 O63 U85 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mach3/.

Gittler:1995:MSA

[GMR95] Xev Gittler, W. Phillip Moore, and J. Rambhaskar. Morgan Stanley’s Aurora system: a next generation global production Unix environment. In USENIX [USE95h], pages 47–

58. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa95/gittler.html.

Grady:1996:QLS

[GMS96] Steven Grady, G. S. Madhusudan, and Marc Sugiyama. QuaSR: a large-scale automated, distributed testing environment. In USENIX [USE96e], pages 61–68. ISBN 1-880446-78-2. LCCN QA 76.73 T44

                                        T35 1996. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/tcl96/grady.html.

Gani-Naor:1997:GAS

[GNMR97] Ilana Gani-Naor, Udi Margolin, and Raz Rafaeli. GeN-MSim — the agent simulator. In USENIX [USE97d], pages 139–147. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/margolin.html.

Goldberg:1990:TRM

[Gol90a] Arthur Goldberg. Transparent recovery of Mach applications. In USENIX [USE90e], pages 169–184.

Goldberg:1990:MUA

[Gol90b] David Goldberg. The MITRE user authentication system. In USENIX [USE90i], pages 1–4. LCCN QA 76.9 A25 U55 1990.

Golding:1992:WCA

[Gol92] Richard A. Golding. A weakconsistency architecture for distributed information systems. In USENIX [USE92a], pages 379– 405.

Golding:1995:ARU

[Gol95]  Richard Golding.  About

Refdbms and USENIX references. ;login: the USENIX Association newsletter, 20(2):55–??, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Goldenhar:1999:NHR

[Gol99] Josh Goldenhar. NetMapper:

Hostname resolution based on client network location. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/goldenhar.
html.

Gong:1996:EES

[Gon96]  Li Gong.  Enclaves: En-

abling secure collaboration over the Internet. In USE-NIX [USE96i], pages 149– 159. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/gong.html.

Gordon:1990:BRD

[Gor90] Edware Gordon. Book review: Designing Object-Oriented Software. ;login: the USENIX Association newsletter, 15(5):15–??, November/December 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Ganger:1994:MUP

[GP94] Gregory R. Ganger and Yale N. Patt. Metadata update performance in file systems. In USENIX [USE94i], pages 49–

60. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/ganger.html.

Ghormley:1998:SES

[GPRA98] Douglas P. Ghormley, David Petrou, Steven H. Rodrigues, and Thomas E. Anderson. SLIC: An extensibility system for commodity operating systems. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/ghormley.html.

Granston:1991:SBP

[GR91] Elana Granston and Vincent Russo. Signature-based polymorphism for C++. In USENIX [USE91k], pages 65–80. LCCN QA76.73.C153 U27 1991.

Granston:1993:SBP

[GR93] Elana D. Granston and Vincent F. Russo. Signature-based polymorphism for C++. In Waldo [Wal93b], pages 121–133. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Grass:1991:RSM

[Gra91]  Judith E. Grass.  Report

from SUUG meeting. ;login: the USENIX Association newsletter, 16(5):20–??, November/December 1991. CODEN LOGNEM. ISSN 1044-6397.

Grass:1992:CSD

[Gra92a] Judith E. Grass. Cdiff: a syntax directed diff for C++ programs. In USENIX [USE92k], pages 181–193. ISBN 1-880446-45-6. LCCN ????

Grass:1992:OOD

[Gra92b]  Judith E. Grass.  Object-

oriented design archaeology with CIA++. In USENIX [USE92d], pages 5–69.

Grass:1993:TS

[Gra93] Translated by Judith E. Grass.

True science. ;login: the USE-NIX Association newsletter, 18 (6):23–??, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Gray:1996:ATF

[Gra96] Robert S. Gray. Agent Tcl:

a flexible and secure mobileagent system. In USE-NIX [USE96e], pages 9–23.

ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/gray.html.

Gray:1998:LSC

[Gra98a] Bob Gray. Loading source code UNIX on the PC. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-8/source.html.

Gray:1998:PHS

[Gra98b] Bob Gray. PC hardware for source code UNIX. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
source.html.

Gray:1998:SCUc

[Gra98c]  Bob Gray.  Source code

UNIX: Help a friend get online, cheap. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
sourcecode.html.

Gray:1998:SCUb

[Gra98d] Bob Gray. Source code UNIX on the PC. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
sourcecode.html.

Gray:1998:SCUa

[Gra98e] Robert Gray. Source code UNIX for PCs. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
source.html.

Gray:1999:SMR

[Gra99a]  Bob Gray.  Software mini-

review: Red Hat Linux 6.0. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
features/redhat.html.

Gray:1999:SCUd

[Gra99b]  Bob Gray.  Source code

UNIX. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/sourcecode.html.

Gray:1999:SCUa

[Gra99c] Bob Gray. Source code UNIX:

Embedding source code UNIX in the product. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

                                        publications/login/1999-4/
                                        sourcecode.html.

Gray:1999:SCUb

[Gra99d] Bob Gray. Source code UNIX: Security on a source code UNIX system. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-6/
                                        features/sourcecode.html.

Gray:1999:SCUc

[Gra99e]  Bob Gray.  Source code

UNIX: What’s your data worth? ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-8/
                                        features/sourcecode.html.

Greely:1991:FFS

[Gre91] J. Greely. A flexible file system cleanup utility. In USENIX [USE91f], pages 105–108.

Gregoire:1993:DUH

[Gré93] Jean-Charles Grégoire. Delegation: Uniformity in heterogeneous distributed administration. In USENIX [USE93f], pages 113–117. ISBN 1-880446-56-1. LCCN QA76.76.O63

                                        L37 1993. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/lisa93/gregoire.
                                        html.

Greenberg:1999:IVP

[Gre99]  Eric Greenberg.  Interop-

erable virtual private networks (VPNs), directory services, and security, 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/ITs/greenberg.pdf.

Unpublished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Grgoire:1993:DUH

[Grg93] Jean-Charles Grgoire. Delegation: Uniformity in heterogeneous distributed administration. In USENIX [USE93f], pages 113–117. ISBN 1-880446-56-1. LCCN QA76.76.O63

L37 1993. URL http://www.
usenix.org/publications/library/
proceedings/lisa93/.

Griffin:1997:TAF

[Gri97] David Griffin. Tcl in AltaVista forum. In USENIX [USE97d], pages 1–8. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/griffin.html.

Griffin:1998:CTS

[Gri98] Dave Griffin. Charity telethon supported by Tcl/Tk. In USENIX [USE98f], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998. URL

http://db.usenix.org/publications/
library/proceedings/tcl98/
griffin.html.

Grubb:1996:HGT

[Gru96]  Michael Grubb.  How to

get there from here: Scaling the enterprise-wide mail infrastructure. In USE-NIX [USE96j], pages 131–

138. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/mg.html.

Gourhant:1990:FCF

[GS90] Yvon Gourhant and Marc Shapiro. FOG/C++: a fragmented object generator. In USENIX [USE90j], pages 63–74.

Gold:1994:WDO

[GS94a] Israel Gold and Uri Shani. Wrapping DCE/OSF client/ server applications. In USE-NIX [USE94g], pages 1–17. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/appdev94/
                                        gold.html.

Goldstein:1994:OBI

[GS94b] Theodore Goldstein and Alan Sloane. The object binary interface: C++ objects for evolvable shared class libraries. In USENIX [USE94e], pages 1–19. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        c++94/goldstein.html.

Guyton:1994:EST

[GS94c] James D. Guyton and Michael F.

Schwartz. Experiences with a survey tool for discovering network time protocol servers. In USENIX [USE94j], pages 257–265. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/index.html.

Gabber:1996:AMD

[GS96a] Eran Gabber and Abraham Silberschatz. Agora: a minimal distributed protocol for electronic commerce. In USE-NIX [USE96h], pages 223–

232. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/gabber.html.

Gwertzman:1996:WWW

[GS96b] James Gwertzman and Margo Seltzer. World Wide Web cache consistency. In USENIX [USE96k], pages 141–151 (or 141–152??). ISBN 1-880446-76-6. LCCN QA 76.76 O63

U88 1996. URL http://www.
usenix.org/publications/library/
proceedings/sd96/seltzer.
html.

Ganapathy:1998:GPO

[GS98] Narayanan Ganapathy and Curt Schimmel. General purpose operating system support for multiple page sizes. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix98/
                                        ganapathy.html.

Ghosh:1999:SUN

[GS99a] Anup K. Ghosh and Aaron Schwartzbard. A study in using neural networks for anomaly and misuse detection. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

                                        1999. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/sec99/ghosh.html.

Gupta:1999:PAN

[GS99b] Minaxi Gupta and Mani Subramanian. Preprocessor algorithm for network management codebook. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        detection99/gupta.html.

Garrett:1993:LSS

[GSB^{+}93] W. E. Garrett, M. L. Scott, R. Bianchini, L. I. Kontothanassis, R. A. McCallum, J. A. Thomas, R. Wisniewski, and S. Luk. Linking shared segments. In USE-NIX [USE93l], pages 13–27. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Gschwind:1995:PGIb

[Gsc95a] Michael Gschwind. A practical guide to internationalization, part 2: Using national characters in hardware. ;login: the USENIX Association newsletter, 20(5):23–29, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Gschwind:1995:PGIa

[Gsc95b]  Michael Gschwind.  Practi-

cal guide to internationalization, part I. ;login: the USE-NIX Association newsletter, 20 (4):29–36, August 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Gifford:1995:PSO

[GSPT95] David K. Gifford, Lawrence C.

Stewart, Andrew C. Payne, and G. Winfield Treese. Payment switches for open networks. In USENIX [USE95g], pages 69–75. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/gifford.html.

Goel:1998:SFC

[GSPW98] Ashvin Goel, David Steere, Calton Pu, and Jonathan Walpole. SWiFT: a feedback control and dynamic reconfiguration toolkit. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/full_papers/poster_
goel/goel.pdf;   http://www.

                                        usenix.org/publications/library/
                                        proceedings/usenix-nt98/full_
                                        papers/poster_goel/goel_html/
                                        goel.html.

Golub:1993:ADD

[GSR93] D. B. Golub, G. G. Sotomayor, and F. L. Rawson. An architecture for device drivers executing as user-level tasks. In USENIX [USE93h], pages 153–172. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

Ghosh:1999:LPB

[GSS99a] Anup K. Ghosh, Aaron Schwartzbard, and Michael Schatz. Learning program behavior profiles for intrusion detection. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/detection99/
                                        ghosh.html.

Gomberg:1999:SRA

[GSS99b] Michail Gomberg, Craig Stacey, and Janet Sayre. Scalable, remote administration of Windows NT. In USE-NIX [USE99b], pages 1–9. ISBN 1-880446-30-8. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa-nt99/gomberg.
                                        html.

Gobioff:1996:SCH

[GSTY96] Howard Gobioff, Sean Smith, J. D. Tygar, and Bennet Yee. Smart Cards in hostile environments. In USENIX [USE96h], pages 23–28. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/gobioff.html.

Guedes:1992:WCS

[Gue92] Paulo Guedes. Writing a clientserver application in C++. In USENIX [USE92k], pages 279–

293. ISBN 1-880446-45-6. LCCN ????

Gunther:1997:NML

[Gun97] Neil Gunther. NT to the max … (NoT). ;login: the USE-NIX Association newsletter, 22 (6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
max.html.

Gunther:1998:ATN

[Gun98a]  Neil Gunther.  The ABCs

of TPCs and NT scalability, II. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
tpc.html.

Gunther:1998:NSI

[Gun98b] Neil Gunther. NT scalability III. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
nt.html.

Gutmann:1996:SDD

[Gut96] Peter Gutmann. Secure deletion of data from magnetic and solid-state memory. In USENIX [USE96i], pages 77–

89. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/gutmann.html.

Guthery:1998:SC

[Gut98a] Scott Guthery. Smart cards. ;login: the USENIX Association newsletter, 23(3): ??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-5/

guthery.html. Special issue on security.

Gutmann:1998:SGP

[Gut98b] Peter Gutmann. Software generation of practically strong random numbers. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998. URL http:

//www.cs.auckland.ac.nz/~
pgut001/;   http://www.cs.
auckland.ac.nz/~pgut001/pubs/
usenix98.pdf;   http://www.
usenix.org/publications/library/
proceedings/sec98/gutmann.
html.

Gutmann:1999:DCS

[Gut99] Peter Gutmann. The design of a cryptographic security architecture. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/gutmann.
html.

Grimshaw:1991:FDS

[GV91] Andrew S. Grimshaw and Vigrilio E. Vivas. FALCON: a distributed scheduler for MIMD architectures. In USENIX [USE91j], pages 149–164. ISBN ???? LCCN ????

Gronvall:1999:DMB

[GWP99] Björn Grönvall, Assar Westerlund, and Stephen Pink. The design of a multicast-based distributed file system. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/gronvall.
html.

Goldberg:1996:SEU

[GWTB96] Ian Goldberg, David Wagner, Randi Thomas, and Eric A. Brewer. A secure environment for untrusted helper applications. In USENIX [USE96i], pages 1–13. ISBN 1-880446-79-

0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/goldberg.html.

Gwyn:1990:BRL

[Gwy90] Douglas A. Gwyn. Book review:

Life with UNIX — A Guide For Everyone. ;login: the USE-NIX Association newsletter, 15 (3):13–??, May/June 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Gottschick:1994:BIE

[GZ94] Jan Gottschick and Malte Zimmermann. Building an integrated and enterprise-specific configuration management. In USENIX [USE94m], pages 69–

79. ISBN ???? LCCN ????

Hansen:1992:CSM

[HA92] Stephen E. Hansen and E. Todd Atkins. Centralized system monitoring with swatch. In USENIX [USE92m], pages 105–

117. ISBN 1-880446-46-4. LCCN

????

Hansen:1993:ASM

[HA93] Stephen E. Hansen and E. Todd Atkins. Automated system monitoring and notification with Swatch. In USE-NIX [USE93f], pages 145–

155. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/hansen.html.

Haahr:1990:MBW

[Haa90] Paul Haahr. Montage: Breaking windows into small pieces. In USENIX [USE90g], pages 289–

298. LCCN QA76.8.U65 U81

1990.

Hitz:1998:MNU

[HAB^{+}98] Dave Hitz, Bridget Allison, Andrea Borr, Rob Hawley, and Mark Muhlestein. Merging NT and UNIX filesystem permissions. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/hitz.html.

HackVan:1998:IPD

[Hac98a] Stig HackVän. Interview with L.

Peter Deutsch. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
interview.html.

Honeyman:1998:SV

[HAC^{+}98b] Peter Honeyman, Andy Adamson, Kevin Coffman, Janani Janakiraman, Rob Jerdonek, and Jim Rees. Secure videoconferencing. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/honeyman.html.

Haemer:1990:UUCa

[Hae90a] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(1):34–43, January/February 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Haemer:1990:UUCb

[Hae90b] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(3):25– 54, May/June 1990. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1990:UUCc

[Hae90c] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(4):21– 58, July/August 1990. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1990:UUCd

[Hae90d] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(5):16–24, September/October 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Haemer:1990:UUR

[Hae90e] Jeffrey S. Haemer. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 15(6):17–39, November/December 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Haemer:1991:BRO

[Hae91a] Jeff Haemer. Book review: Open Systems: A Business Strategy for the 1990’s. ;login: the USE-NIX Association newsletter, 16 (5):10–??, September/October 1991. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1991:UURa

[Hae91b] Jeffrey S. Haemer. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(1):24–34, January/February 1991. CO-DEN LOGNEM. ISSN 1044-6397.

Haemer:1991:UURb

[Hae91c] Jeffrey S. Haemer. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(2):22– 25, March/April 1991. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1992:CUD

[Hae92] Jeff Haemer. Can UNIX designers learn anything from PCs? ;login: the USENIX Association newsletter, 17(3):17–18, May/ June 1992. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1994:IRM

[Hae94a]  Jeffery S. Haemer.  Imake

rhymes with mistake. ;login: the USENIX Association newsletter, 19(1):32–33, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1994:NOO

[Hae94b] Jeffrey S. Haemer. A new object-oriented programming language: sh. In USE-NIX [USE94j], pages 1–13. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/haemer.html.

Haemer:1995:RJI

[Hae95a] Jeff Haemer. Review: Japanese information processing. ;login: the USENIX Association newsletter, 20(1):45–47, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Haemer:1995:VHL

[Hae95b] Jeff Haemer. Very high level languages symposium report. ;login: the USENIX Association newsletter, 20(1):5–10, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Haines:1997:DLT

[Hai97]  Matthew Haines.  On de-

signing lightweight threads for substrate software. In USE-NIX [USE97e], pages 243–

255. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/haines.html.

Hall:1994:FGP

[Hal94] J. Storrs Hall. Fornax: a general-purpose programming language. In USENIX [USE94k], pages 147–156. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/hall.html.

Hall:1998:EPPa

[Hal98a]  Joseph N. Hall.  Effective

Perl programming. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
perl.html.

Hall:1998:EPPb

[Hal98b] Joseph N. Hall. Effective Perl programming. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
perl.html.

Hall:1998:EPPc

[Hal98c] Joseph N. Hall. Effective Perl programming: Simplicity is a good object. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
perl.html.

Hall:1998:KOP

[Hal98d] Joseph N. Hall. Knit one, Perl 5.005. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
knitone.html.

Hall:1999:DC

[Hal99a] Joseph N. Hall. Don’t CPANic.

;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/cpanic.html.

Hall:1999:EPPb

[Hal99b]  Joseph N. Hall.  Effective

Perl programming: a Perl CPANorama. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-6/
                                        features/perl.html.

Hall:1999:EPPa

[Hal99c] Joseph N. Hall. Effective Perl programming: Analysis without paralysis. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-4/
                                        perl.html.

Halprin:1999:SAB

[Hal99d] Geoff Halprin. The system administrator’s body of knowledge, 1999. URL http://db.

                                        usenix.org/publications/library/
                                        proceedings/lisa99/invitedtalks/

halprin.pdf. Unpublished in-

vited talk at LISA XIII.

Hamilton:1996:IOS

[Ham96] Jennifer Hamilton. Interlanguage object sharing with SOM. In USENIX [USE96g], pages 35–51. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots96/hamilton.html.

Hamilton:1997:MSM

[Ham97a] David Hamilton. Microsoft and systems management. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/its/hamilton.

it. Invited talk.

Hamilton:1997:MSP

[Ham97b] Jennifer Hamilton. Montana smart pointers: They’re smart, and they’re pointers. In USENIX [USE97g], pages 21–39. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/hamilton.html. See

also [LH97, Rie97, WW97].

Hanks:1992:CMU

[Han92] Stanley P. Hanks. Creating MANs using LAN technology: Sometimes you gotta break the rules. In USENIX [USE92j], pages 439–451.

Hanson:1997:BRS

[Han97] Steve Hanson. Book review:

sendmail, 2d ed. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/costales.html.

Hand:1999:SPN

[Han99a] Steven M. Hand. Self-paging in the Nemesis operating system. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

                                        5. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi99/hand.html.

Hanson:1999:BRN

[Han99b]  Steve Hanson.  Book re-

view: NT Backup and Restore. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        nt_backup.html.

Harrison:1990:DMS

[Har90] Hellen E. Harrison. A domain mail system on dissimilar computers: Trials and tribulations of SMTP. In USENIX [USE90f], pages 133–136. LCCN QA76.76.O63 L37 1990.

Harrison:1991:DMS

[Har91] Helen E. Harrison. A domain mail system on dissimilar computers — trials and tribulations of SMTP. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Harrison:1992:MWL

[Har92] Helen E. Harrison. So many workstations, so little time. In USENIX [USE92e], pages 79– 87. ISBN 1-880446-47-2. LCCN ????

Harlander:1994:CSA

[Har94]  Magnus Harlander.  Cen-

tral system administration in a heterogeneous Unix environment: GeNUAdmin. In USENIX [USE94h], pages 1–

8. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/harlander.html.

Harris:1995:UID

[Har95] Cindy A. Harris. User interface development tools. ;login: the USENIX Association newsletter, 20(1):22–24, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Harrison:1996:DRW

[Har96a] Helen E. Harrison. A dozen reasons to write a paper. ;login: the USENIX Association newsletter, 21(2):35–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Harrison:1996:RC

[Har96b] Tim H. Harrison. Report on COOTS ’96. ;login: the USE-NIX Association newsletter, 21 (4):13–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Harker:1997:SRS

[Har97a]  Robert Harker.  Selectively

rejecting SPAM using Sendmail. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/22.harker.html.

Harold:1997:JNP

[Har97b] Elliotte Rusty Harold. Java Network Programming Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-227-1. xvii + 422 pp. LCCN QA76.73.J38 H37 1997. US$29.95. URL http://www.

ora.com/catalog/{javanetwk}
/noframes.html;   http://
www.oreilly.com/catalog/javanetwk.

Harris:1999:LMN

[Har99a] Trey Harris. Look, Ma, no hands! Coping with RSI,

1999. URL http://db.usenix.
org/publications/library/
proceedings/lisa99/practicum/
harris/harris.html. Unpub-

lished practicum talk at LISA XIII.

Harrison:1999:MOS

[Har99b]  Mark Harrison.  Managing

open source software. ;login: the USENIX Association newsletter, 24(1s):??, January 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-1/
opensource.html. Special issue

on tools.

Haugh:1992:ISP

[Hau92] John F. Haugh, II. Introduction to the shadow password suite. In USENIX [USE92m], pages 133–

144. ISBN 1-880446-46-4. LCCN ????

Hauser:1994:SUL

[Hau94]  Carl Hauser.  Speeding up

UNIX login by caching the initial environment. In USE-NIX [USE94h], pages 117–

124. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/hauser.html.

Hawley:1990:POA

[Haw90a] Michael Hawley. The personal orchestra, or audio data compression by 10,000:1. In USE-NIX [USE90b], pages 289–330.

Hawley:1990:BRE

[Haw90b] Mike Hawley. Book review:

Elements of Computer Music. ;login: the USENIX Association newsletter, 15(1):16–??, January/February 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Haxby:1996:LGC

[Hax96] Andy Haxby. The last game of cricket. ;login: the USE-NIX Association newsletter, 21 (6):47–48, December 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Hunt:1999:DBI

[HB99] Galen Hunt and Doug Brubacher.

Detours: Binary interception of Win32 functions. In USE-NIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/hunt.html.

Hagen:1998:SSN

[HBvdRA98] Teus Hagen, Frances Brazier, Wytze van der Raay, and Jos Alsters. A second start for the NLnet Foundation. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Hipp:1997:PUT

[HC97] D. Richard Hipp and Mike Cruse. PtTcl: Using Tcl with Pthreads. In USE-NIX [USE97d], pages 125–

130. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/hipp.html.

Hawblitzel:1998:IMP

[HCC^{+}98] Chris Hawblitzel, Chi-Chao Chang, Grzegorz Czajkowski, Deyu Hu, and Thorsten von Eicken. Implementing multiple protection domains in Java. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/hawblitzel.html.

Hogan:1995:DDS

[HCH95] Christine Hogan, Aoife Cox, and Tim Hunter. Decentralising distributed systems administration. In USENIX [USE95h], pages 139–148 (or 139–147??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/index.html.

Huang:1998:NSS

[HCK^{+}98] Yennun Huang, P. Emerald Chung, Chandra Kintala, Chung-Yih Wang, and De-Ron Liang. NT-SwiFT: Software implemented fault tolerance on Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/huang.html;   http://
www.usenix.org/publications/
library/proceedings/usenix-
nt98/huang_slides.

Ho:1995:OPD

[HCL95] W. Wilson Ho, Wei-Chau Chang, and Lilian H. Leung. Optimizing the performance of dynamically-linked programs. In USENIX [USE95e], pages 225– 234 (or 225–233??). ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995. URL http://

www.usenix.org/publications/
library/proceedings/neworl/
ho.html.

Hewett:1997:BDA

[HdJB97] Carl Hewett, Peter de Jong, and Nat Brown. Building distributed applications — CORBA and DCOM, 1997.

URL http://www.usenix.org/
publications/library/proceedings/

                                        usenix-nt97/technical.html.

Unpublished invted talk and panel at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Hearn:1997:LSN

[Hea97] Steven Hearn. Large scale NT operations management. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

Hecker:1994:RSL

[Hec94] Frank Hecker. Review: a Second Look. ;login: the USENIX Association newsletter, 19(2):45– 46, March/April 1994. CODEN LOGNEM. ISSN 1044-6397.

Heckman:1996:FMP

[Hec96] Carey Eugene Heckman. Forming a more perfect net governance, 1996. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/sd96/. Unpub-

lished invited talk at the USE-NIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Hedstrom:1998:GLD

[Hed98] Leif Hedstrom. Got LDAP? Deploying and using the Lightweight Directory Access Protocol, 1998. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6–11, 1998 Boston, Massachusetts, USA.

Heimlich:1990:TCN

[Hei90] Steven A. Heimlich. Traffic characterization of the NSFNET national backbone. In USE-NIX [USE90h], pages 207–228. LCCN QA76.8.U65 U82 1990.

Heintze:1996:SDF

[Hei96] Nevin Heintze. Scalable document fingerprinting. In USE-NIX [USE96h], pages 191–

200. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Heiss:1999:ERJ

[Hei99] Jason Heiss. Enterprise rollouts with JumpStart. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/heiss.
html.

Hellerstein:1992:CCC

[Hel92] Joseph L. Hellerstein. Control considerations for CPU scheduling in UNIX systems. In USE-NIX [USE92j], pages 359–374.

Hendricks:1990:FSD

[Hen90] David Hendricks. A filesystem for software development. In USENIX [USE90g], pages 333–

340. LCCN QA76.8.U65 U81 1990.

Herlocker:1996:NN

[Her96] Jonathan L. Herlocker. The NR newsreader. In USE-NIX [USE96e], pages 75–81. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/herlocker.html.

Hertel:1997:OSF

[Her97] Christopher R. Hertel. One size fits all: Stretching Net-BIOS to fit the enterprise. ;login: the USENIX Association newsletter, 22(6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1997-11/
                                        size.html.

Heuser:1990:IRT

[Heu90]  Mark Heuser.  An imple-

mentation of real-time thread synchronization. In USENIX [USE90g], pages 97–106. LCCN QA76.8.U65 U81 1990.

Herrin:1991:ADF

[HF91] Erin H. Herrin and Raphael Finkel. An ASCII database for fast queries of relatively stable data. In USENIX [USE91b], pages 127–156.

Herrin:1996:STT

[HF96] Eric H. Herrin, II and Raphael A. Finkel. Schema and tuple trees: An intuitive structure for representing relational data. In USE-NIX [USE96b], pages 93–118.

Hall:1993:CPP

[HG93] Robert J. Hall and Aaron J. Goldberg. Call path profiling of monotonic program resources in UNIX. In USE-NIX [USE93g], pages 1–13.

ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/hall.html.

Harris:1996:DIN

[HG96] J. Archer Harris and Gregory Gingerich. The design and implementation of a network account management system. In USENIX [USE96j], pages 33–41. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/arch.html.

Hjalmtysson:1998:DCC

[HG98] Gísli Hjálmtysson and Robert Gray. Dynamic C++ classes: a lightweight mechanism to update code in a running program. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/hjalmtysson.html.

Huston:1993:DOA

[HH93] Larry B. Huston and Peter Honeyman. Disconnected operation for AFS. In USE-NIX [USE93i], pages 1–10. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/huston.html.

Huston:1995:PCOa

[HH95a] L. B. Huston and P. Honeyman.

Partially connected operation.

In USENIX [USE95a], pages

                                        365–379. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/mob95/huston.
                                        html.

Huston:1995:PCOb

[HH95b] L. B. Huston and P. Honeyman. Partially connected operation. In USENIX [USE95i], pages 91–98 (or 91–97??). ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/mob95/huston.
                                        html.

Hohndel:1999:AIL

[HH99] Dirk Hohndel and Fabian Herschel. Automated installation of Linux systems using YaST. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa99/hohndel.
                                        html.

Hitz:1990:UUO

[HHLS90] David Hitz, Guy Harris, James K. Lau, and Allan M. Schwartz. Using UNIX as one component of a lightweight distributed kernel for microprocessor file servers. In USENIX [USE90h], pages 285–296. LCCN QA76.8.U65 U82 1990.

Honeyman:1992:HA

[HHS92] P. Honeyman, L. B. Huston, and M. T. Stolarchuk. Hijacking AFS. In USENIX [USE92j], pages 175–182.

Hughes:1997:JNP

[HHSW97] Merlin Hughes, Conrad Hughes, Michael Shoffner, and Maria Winslow, editors. Java Network Programming. Prentice-Hall and Manning Publications, Englewood Cliffs, NJ 07632, USA and Greenwich, CT, USA, January 1997. ISBN 1-884777-34-1 (Manning), 0-13-841206-5 (Prentice-Hall). xxiv + 519 pp. LCCN QA76.73.J38J377 1997. US$44.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0138412065.html.  Includes

CD-ROM.

Hollinden:1992:EIM

[HHW92] David Y. Hollinden, Debra A.

Hensgen, and Phillip A. Wilsey. Experiences implementing the Mintabs System on a MasPar MP-1. In USENIX [USE92l], pages 43–58. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Hickstein:1999:MPP

[Hic99] Jim Hickstein. A management perspective on privileged access to computer systems. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
mgmt_persp.html.

Hideyo:1994:OMU

[Hid94] Imazu Hideyo. OMNICONF — making OS upgrades and disk crash recovery easier. In USENIX [USE94h], pages 27–

31. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/hideyo.html.

Hietaniemi:1992:IPP

[Hie92] Jarkko Hietaniemi. ipasswd — proactive password security. In USENIX [USE92e], pages 105–

125. ISBN 1-880446-47-2. LCCN

????

Hildebrand:1992:AOQ

[Hil92] Dan Hildebrand. An architectural overview of QNX. In USENIX [USE92i], pages 113–

126. ISBN 1-880446-42-1. LCCN

QAX 32.

Hill:1996:PSF

[Hil96] Brian C. Hill. Priv: Secure and flexible privileged access dissemination. In USENIX [USE96j], pages 1–8. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/bchill.html.

Hill:1997:MTY

[Hil97] Paul Hill. More than you ever wanted to know about NT login authentication. ;login: the USE-NIX Association newsletter, 22 (4):18–19, August 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Hirsch:1997:BGW

[Hir97] Frederick J. Hirsch. Building a graphical Web history using Tcl/Tk. In USE-NIX [USE97d], pages 159–

160. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_hirsch.html.

Hastings:1992:PTD

[HJ92] Reed Hastings and Bob Joyce.

Purify: a tool for detecting memory leaks and access errors in C and C++ programs. In USENIX [USE92j], pages 125– 138.

Hansen:1997:SSN

[HJ97] Jørgen Sværke Hansen and Eric Jul. A scheduling scheme for network saturated NT multiprocessors. In USE-NIX [USE97f], pages 39–

46. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/hansen.html.

Hemmer:1992:MSG

[HJK^{+}92] F. Hemmer, E. Jagel, A. Kumar, L. Robertson, B. Segal, and A. Trannoy. Mainframe services from gigabit-networked workstations. In USENIX [USE92g], pages 165–176. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Haynes:1992:SSN

[HK92] Rena A. Haynes and Suzanne M.

Kelly. Software security for a network storage service. In USENIX [USE92m], pages 253– 264. ISBN 1-880446-46-4. LCCN ????

Hamilton:1993:SNM

[HK93] Graham Hamilton and Panos Kougiouris. The Spring nucleus: a microkernel for objects. In USENIX [USE93g], pages 147–159. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/hamilton.html.

Hannum:1995:NPH

[HK95a] Charles Hannum and John Kohl. The NetBSD project: a highly portable UNIX-like system. ;login: the USENIX Association newsletter, 20(6):39– 41, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Herlocker:1995:TCM

[HK95b] Jonathan L. Herlocker and Joseph A. Konstan. Tcl commands as media in a distributed multimedia toolkit. In USENIX [USE95j], pages 205–212. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/herlocker.html.

Hanson:1997:SEG

[HK97] David R. Hanson and Jeffrey L. Korn. A simple and extensible graphical debugger. In USENIX [USE97e], pages 173–184. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/

ana97/hanson.html;   http:
//www.usenix.org/publications/
library/proceedings/ana97/
summaries/hanson.html.

Hodes:1999:DBF

[HK99] Todd D. Hodes and Randy H.

Katz. A document-based framework for Internet application control. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/hodes.html.

Hoogenboom:1993:CSP

[HL93] Peter Hoogenboom and Jay Lepreau. Computer system performance problem detection using time series models. In USENIX [USE93g], pages 15–32. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/hoogen.html.

Lin:1998:SRT

[hLhCN98] Chih han Lin, Hao hua Chu, and Klara Nahrstedt. A soft real-time scheduling server on the Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/lin.html.

Hitz:1994:FSD

[HLM94] Dave Hitz, James Lau, and Michael Malcolm. File system design for an NFS file server appliance. In USE-NIX [USE94l], pages 235–246. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/hitz.html.

Hylands:1997:TUI

[HLR97] Christopher Hylands, Edward A. Lee, and H. John Reekie. The Tycho user interface system. In USENIX [USE97d], pages 149–157. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/hylands.html.

Hardy:1992:BAS

[HM92] Darren R. Hardy and Herb M. Morreale. buzzerd: Automated systems monitoring with notification in a network environment. In USENIX [USE92e], pages 203–210. ISBN 1-880446-47-2. LCCN ????

Hutchinson:1999:LVP

[HMF^{+}99] Norman C. Hutchinson, Stephen Manley, Mike Federwisch, Guy Harris, Dave Hitz, Steven Kleiman, and Sean O’Malley. Logical vs. physical file system backup. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/hutchinson.
html.

Harrison:1994:PFN

[HMS94] Helen E. Harrison, Mike C.

Mitchell, and Michael E. Shaddock. Pong: a flexible network services monitoring system. In USENIX [USE94h], pages 167–

173. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/harrison.html.

Harkavy:1998:IPP

[HMT^{+}98] Michael Harkavy, Andrew Myers, J. D. Tygar, Alma Whitten, and H. Chi Wong. Invited presentations on Public Key Infrastructure. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

URL http://www.usenix.org/
publications/library/proceedings/
ec98/pki.html.

Howden:1997:HCM

[HN97] Sally J. Howden and Frank B.

Northrup. How to control and manage change in a commercial data center without losing your mind. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/06.howden.html.

Hagiya:1990:POO

[HO90a] Masami Hagiya and Kouji Ohtani. Parallel objectoriented UIMS with macro and micro stubs. In USE-NIX [USE90h], pages 259–274. LCCN QA76.8.U65 U82 1990.

Hartmann:1990:PMM

[HO90b] John H. Hartmann and John L. Ousterhout. Performance measurements of a multiprocessor Sprite kernel. In USE-NIX [USE90g], pages 279–288. LCCN QA76.8.U65 U81 1990.

Hartman:1992:ZSN

[HO92] John H. Hartman and John K. Ousterhout. Zebra: a striped network file system. In USE-NIX [USE92h], pages 71–78. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Hayes:1994:XIG

[HO94] Jordan M. Hayes and Charles A. Ocheret. The Xt intrinsics as a general purpose application development platform or A user interface toolkit with optional users. In USENIX [USE94g], pages 147–157. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994. URL http://

                                        www.thinkbank.com/jordan/
                                        papers/USENIX/AppDev.html.

Hoff:1999:FTT

[Hof99]  Timothy Hoff.  Fundamen-

tals of troubleshooting TCP/ IP. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-10/
                                        features/tcp.html.

Hogan:1995:MM

[Hog95]  Christine Hogan.  Metrics

for management. In USE-NIX [USE95h], pages 125–

132. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/hogan.html.

Holland:1997:BMB

[Hol97a] Richard J. Holland. A better E-mail bouncer. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997. URL

http://www.usenix.org/publications/
library/proceedings/lisa97/
23.holland.html.

Holmwood:1997:IIF

[Hol97b] John Holmwood. Implementing an infrastructure framework. ;login: the USENIX Association newsletter, 22(6): ??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
implementing.html.

Honeyman:1994:GE

[Hon94] Peter Honeyman. Guest editorial. In USENIX [USE94b], pages 173–??

Honeyman:1995:DFUb

[Hon95a] P. Honeyman. Digest of the first USENIX workshop on electronic commerce. In USENIX [USE95g], pages 209–225 (or 209–223??). ISBN 1-880446-74-X. LCCN HF5548.33. U84

1995(1). URL http://www.
usenix.org/publications/library/
proceedings/ec95/index.html.

Honeyman:1995:DUWa

[Hon95b] P. Honeyman. Digest of the USENIX workshop on electronic commerce. ;login: the USENIX Association newsletter, 20(6):6– 19, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Hopkins:1992:CNC

[Hop92] Sharon Hopkins. Camels and needles: Computer poetry meets the Perl programming language. In USENIX [USE92j], pages 391–404.

Hopkirk:1996:VTB

[Hop96] Mike Hopkirk. Visual Tcl: Building a distributed Multi-Personality GUI toolkit for Tcl. In USENIX [USE96e], pages 91–102. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/hopkirk.html.

Horton:1992:UG

[Hor92] Mark Horton. Using Graph-Tab. ;login: the USENIX Association newsletter, 17(5):28–31, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Houle:1996:MWI

[Hou96] Bill Houle. MajorCool: a Web interface to Majordomo. In USENIX [USE96j], pages 145–153. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/bhoule.html.

Hoven:1991:MIS

[Hov91] Rand Hoven. Mach interfaces to support guest O.S. debugging. In USENIX [USE91h], pages 131–148. LCCN QAX 27.

Howes:1990:IXD

[How90]  Timothy Howes.  Integrat-

ing X.500 directory service into a large campus computing environment. In USENIX [USE90f], pages 125–132. LCCN QA76.76.O63 L37 1990.

Howes:1991:IXD

[How91] Timothy Howes. Integrating X.500 directory service into a large campus computing environment. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Howlett:1995:TBL

[How95] George A. Howlett. A tablebased layout editor. In USE-NIX [USE95j], pages 39–42 (or 39–41??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/howlett.html.

Howell:1998:FM

[How98a] Jon Howell. Faq-O-Matic. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-6/faq.html.

Howlett:1998:DO

[How98b] George A. Howlett. Data objects. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/howlett.
html.

Hu:1997:HPW

[HPS97] James C. Hu, Irfan Pyarali, and Douglas C. Schmidt. High performance Web servers on Windows NT: Design and performance. In USENIX [USE97f], page 149. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/single_hu.html.

Heller:1991:KSN

[HR91] Jeffrey Heller and Franklin Reynolds. Kernel support for network protocol servers. In USENIX [USE91h], pages 149–

162. LCCN QAX 27.

Hillyer:1992:IBB

[HR92] Bruce K. Hillyer and Bethany S. Robinson. Issues in BBFS, a broadband filesystem. In USE-NIX [USE92h], pages 129–130. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Haahr:1993:SHO

[HR93] Paul Haahr and Byron Rakitzis. Es: a shell with higherorder functions. In USE-NIX [USE93l], pages 51–60. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Hollander:1996:KWS

[HRT96] Isaac Hollander, P. Rajaram, and Constantin Tanno. Kerberos on Wall Street. In USENIX [USE96i], pages 105–

112. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/hollander.html.

Hahn:1991:PII

[HS91a] Douglas Hahn and Neil Soiffer. Pragmatic issues in the implementation of flexible libraries for C++. In USE-NIX [USE91k], pages 193–202. LCCN QA76.73.C153 U27 1991. See [CHS91b].

Hodges:1991:PEM

[HS91b] Matthew Hodges and Russel Sasnett. Plastic editors for multimedia documents. In USE-NIX [USE91g], pages 463–473. LCCN QA 76.76 O63 U83 1991.

Howell:1991:WME

[HS91c] Bud Howell and Bjorn Satdeva.

We have met the enemy, an informal survey of policy practices in the internetworked community. In USENIX [USE91f], pages 159–170.

Hume:1991:FSS

[HS91d] Andrew Hume and Daniel Sunday. Fast string searching. In USENIX [USE91g], pages 221–

234. LCCN QA 76.76 O63 U83

Hambridge:1993:HBD

[HS93a] Sally Hambridge and Jeffrey C. Sedayao. Horses and barn doors: Evolution of corporate guidelines for Internet usage. In USENIX [USE93f], pages 9–16. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/hambridge.html.

Hardy:1993:ERD

[HS93b] Darren R. Hardy and Michael F. Schwartz. Essence: a resource discovery system based on semantic file indexing. In USENIX [USE93l], pages 361–

373. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Hagsand:1994:WSR

[HS94] Olof Hagsand and Peter Sjodin. Workstation support for realtime multimedia communication. In USENIX [USE94l], pages 133–142. ISBN 1-880446-58-8. LCCN QA 76.76 O63

U84 1994. URL http://www.
usenix.org/publications/library/
proceedings/sf94/hagsand.
html.

Hauser:1995:GEW

[HS95a] Ralf Hauser and Michael Steiner. Generic extensions of WWW browsers. In USENIX [USE95g], pages 147–154. ISBN 1-880446-74-X. LCCN HF5548.33. U84

1995(1). URL http://www.
usenix.org/publications/library/
proceedings/ec95/hauser.html.

Heidrich:1995:AGT

[HS95b] Wolfgang Heidrich and Philipp Slusallek. Automatic generation of Tcl bindings for C and C++ libraries. In USE-NIX [USE95j], pages 85–94 (or 85–93??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/heidrich.html.

Hardjono:1996:SEC

[HS96] Thomas Hardjono and Jennifer Seberry. Strongboxes for electronic commerce. In USENIX [USE96h], pages 135–

145. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/hardjono.html.

Honeyman:1997:LMH

[HS97] Peter Honeyman and Joe Saul. Logging and monitoring: How, why, and when, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/invited_talks/honey_
saul/honeyman.pdf;   http:
//www.usenix.org/publications/
library/proceedings/lisa97/
invited_talks/honey_saul/
honeyman.ps;   http://www.
usenix.org/publications/library/
proceedings/lisa97/invited_
talks/honey_saul/saul.pdf;
 http://www.usenix.org/publications/

                                        library/proceedings/lisa97/
                                        invited_talks/honey_saul/

saul.ps. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Howard:1999:BIA

[HS99a] Michael Howard and Christopher S. Sontag. Bringing the Internet to all electronic devices. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        es99/howard.html.

Hunt:1999:CAD

[HS99b] Galen C. Hunt and Michael L. Scott. The Coign automatic distributed partitioning system. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

                                        5. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi99/hunt.html.

Hunt:1999:IIC

[HS99c] Galen C. Hunt and Michael L. Scott. Intercepting and instrumenting COM applications. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots99/hunt.html.

Heybey:1996:CDS

[HSE96] Andrew Heybey, Mark Sullivan, and Paul England. Calliope: a distributed, scalable multimedia server. In USE-NIX [USE96k], pages 75–86. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/heybey.html.

Helme:1999:OD

[HSK99] Arne Helme and Tage Stabell-Kulø. Offline delegation. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

http://db.usenix.org/publications/
library/proceedings/sec99/
helme.html.

Hambridge:1999:BTM

[HSOS99] Sally Hambridge, Charles Smothers, Tod Oace, and Jeff Sedayao. Just type make! Managing Internet firewalls using make and other publicly available utilities. In USENIX [USE99g], page ?? ISBN 1-880446-36-

7. LCCN TK5105.5 .C664 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/hambridge.html.

Heydon:1992:SCU

[HT92] Allan Heydon and J. D. Tygar.

Specifying and checking UNIX security constraints. In USENIX [USE92m], pages 211–226. ISBN 1-880446-46-4. LCCN ????

Heydon:1994:SCU

[HT94] Allan Heydon and J. D. Tygar. Specifying and checking UNIX security constraints. In USENIX [USE94d], pages 91–112.

Hauser:1996:SI

[HT96] Ralf Hauser and Gene Tsudik. On shopping incognito. In USENIX [USE96h], pages 251–

257. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec96/tsudik.html.

Harkavy:1998:EAP

[HTK98] Michael Harkavy, J. D. Tygar, and Hiroaki Kikuchi. Electronic auctions with private bids. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec98/harkavy.html.

Heintze:1996:MCE

[HTWW96] Nevin Heintze, J. D. Tygar, Jeannette Wing, and H. Chi Wong. Model checking electronic commerce protocols. In USENIX [USE96h], pages 146–

164. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec96/wong.html.

Huang:1998:SEP

[Hua98] Ping Huang. Some emerging peripherals interconnect technologies. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
peripherals.html.

Hubbard:1998:IPO

[Hub98] Jordan K. Hubbard. IMHO:

Pulling on one end of the rope. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Hudak:1997:PDS

[Hud97] Paul Hudak. The promise of domain-specific languages, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/technical.html. Un-

published keynote address at the Conference on Domain-Specific Languages (DSL), 1997 October 15–17, 1997 Santa Barbara, California, USA.

Hughes:1996:UVS

[Hug96] Doug Hughes. Using visualization in system and network administration. In USENIX [USE96j], pages 59–

66. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/doug.html.

Hume:1996:DCS

[Hum96a] Andrew Hume. The demise of computing systems. ;login: the USENIX Association newsletter, 21(5):17–??, October 1996.

CODEN LOGNEM. ISSN 1044-6397.

Hume:1996:LP

[Hum96b] Andrew Hume. Letter from the President. ;login: the USE-NIX Association newsletter, 21 (5):7–??, October 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Hume:1997:BRM

[Hum97a] Andrew Hume. Book review: Mastering regular expressions. ;login: the USENIX Association newsletter, 22(5):67–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Hume:1997:WNU

[Hum97b] Andrew Hume. Whither now, USENIX? ;login: the USE-NIX Association newsletter, 22 (5):2–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Hume:1998:BRS

[Hum98a] Andrew Hume. Book review: Sun Performance and Tuning: Java and the Internet. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-10/
                                        sun.html.

Hume:1998:LP

[Hum98b] Andrew Hume. Letter from the President. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Hume:1998:LPR

[Hum98c] Andrew Hume. Letter from the President: Rainless in Seattle. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
mailing.html.

Hume:1999:C

[Hum99a] Andrew Hume. Changes.

;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
changes.html.

Hume:1999:DSR

[Hum99b] Andrew Hume. The dark side of regular expressions. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
reg_exp.html.

Hungerford:1995:ME

[Hun95] Debby Hungerford. A moving experience. ;login: the USE-NIX Association newsletter, 20 (2):30–32, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Hunt:1997:CUM

[Hun97]  Galen C. Hunt.  Creat-

ing user-mode device drivers with a proxy. In USE-NIX [USE97f], pages 55–

59. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/hunt.html.

vonEicken:1997:DYN

[hvESV^{+}97] horsten von Eicken, Margo I. Seltzer, Werner Vogels, P. Geoffrey Lowney, Nick Vasilatos, and Brian Bershad. Do you need source?, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/technical.html.

Unpublished panel session at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Hunter:1993:GSA

[HW93] Tim Hunter and Scott Watanabe. Guerrilla system administration: Scaling small group systems administration to a larger installed base. In USENIX [USE93f], pages 99–

105. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/hunter.html.

Ioannidis:1993:AIN

[IB93] John Ioannidis and Matt Blaze. The architecture and implementation of network layer security in UNIX. In USE-NIX [USE93m], pages 29–

39. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/.

Islam:1992:DCS

[IC92] Nayeem Islam and Roy H.

Campbell. Design considerations for shared memory multiprocessor message systems. In USENIX [USE92l], pages 149–

172. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Iyengar:1997:IWS

[IC97] Arun Iyengar and Jim Challenger. Improving Web server performance by caching dynamic data. In USENIX [USE97h], pages 49–60. ISBN 1-880446-91-X. LCCN TK5105.875.I57

U96 1997. URL http://www.
usenix.org/publications/library/
proceedings/usits97/iyengar.
html.

IEEE:1990:PIW

[IEE90] Proceedings: (Second) IEEE Workshop on Experimental Distributed Systems, October 11– 12, 1990, Huntsville, Alabama. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2092-7 (paperback), 0-8186-9092-5 (case), 0-8186-6092-9 (microfiche). LCCN QA76.9.D5 I33 1990. Sponsored by IEEE Computer Society, TC on Distributed Processing, in cooperation with USENIX Association.

Israel:1991:EPN

[IFT^{+}91] Robert K. Israel, Antony W. Foster, Arun Taylor, Tracy M. Taylor, and Neil Webber. Evolutionary path to network storage management. In USENIX [USE91i], pages 185–198. LCCN QA 76.76 O63 U84 1992.

Burge:1999:JJB

[IG99] Legand L. Burge III and K. M. George. JMAS: a Java-based mobile actor system for distributed parallel computation. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots99/burge.html.

Imielinski:1995:EED

[IGP95] Tomasz Imielinski, Monish Gupta, and Sarma Peyyeti. Energy efficient data filtering and communication in mobile wireless computing. In USENIX [USE95i], pages 109–120 (or 109–119??). ISBN 1-880446-69-

3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/imielinski.html.

Itoi:1998:PAM

[IH98] Naomaru Itoi and Peter Honeyman. Pluggable authentication modules for Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/

usenix-nt98/itoi.html;   http:
//www.usenix.org/publications/
library/proceedings/usenix-
nt98/itoi_slides.

Itoi:1999:SIK

[IH99] Naomaru Itoi and Peter Honeyman. Smartcard integration with Kerberos V5. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

http://www.usenix.org/publications/
library/proceedings/smartcard99/
itoiKerberos.html.

Itoi:1999:SUF

[IHR99] Naomaru Itoi, Peter Honeyman, and Jim Rees. SCFS: a UNIX filesystem for Smartcards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/itoiSCFS.html.

Itoh:1995:UMO

[iIYL95] Jun ichiro Itoh, Yashuhiko Yokote, and Rodger Lea. Using meta-objects to support optimisation in the Apertos operating system. In USENIX [USE95k], pages 147–158 (or 147–157??). ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/itoh.html.

Ingham:1997:TEM

[IJ97] Jim Ingham and Raymond Johnson. TclOSAScript — exec for Mac Tcl. In USE-NIX [USE97d], pages 43–

48. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/ingham.html.

Iyengar:1995:TDP

[IK95] Sunanda Iyengar and Joseph A. Konstan. TclProp: a datapropagation formula manager for Tcl and Tk. In USE-NIX [USE95j], pages 25–

30. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/iyengar.html.

Interrante:1990:RAT

[IL90] John A. Interrante and Mark A. Linton. Runtime access to type information in C++. In USE-NIX [USE90j], pages 233–240.

Interrante:1993:RAT

[IL93] John A. Interrante and Mark A. Linton. Runtime access to type information in C++. In Waldo [Wal93b], pages 189–196. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Ilacqua:1999:FI

[Ila99]  Spike Ilacqua.  The first

ISP. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
isp.html.

Ioannidis:1993:DIM

[IM93] John Ioannidis and Gerald Q.

Maguire, Jr. The design and implementation of a mobile Internetworking architecture. In USENIX [USE93l], pages 489–

502. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Islam:1993:PDA

[IMC93] Nayeem Islam, Robert E. Mc-Grath, and Roy H. Campbell. Parallel distributed application performance and message passing: a case study. In USENIX [USE93j], pages 73–88. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

Ingham:1996:TTO

[Ing96] Jim Ingham. Tcl/Tk as an OpenDoc scripting part. In USENIX [USE96e], pages 115–

123. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/ingham.html.

Instenes:1995:GKS

[Ins95a] Shawn Instenes. Getting to know SATAN. ;login: the USE-NIX Association newsletter, 20 (3):23–26, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1995:MAA

[Ins95b] Shawn Instenes. Musings about authentication. ;login: the USE-NIX Association newsletter, 20 (4):22–23, August 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Instenes:1995:SRR

[Ins95c] Shawn Instenes. A secure rsh replacement. ;login: the USE-NIX Association newsletter, 20 (5):19–??, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Instenes:1995:SP

[Ins95d] Shawn Instenes. Security potpourri. ;login: the USENIX Association newsletter, 20(6):28– ??, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1996:FLE

[Ins96a] Shawn Instenes. Food for the line eater. ;login: the USE-NIX Association newsletter, 21 (3):20–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1996:FIS

[Ins96b] Shawn Instenes. The future of IP security. ;login: the USE-NIX Association newsletter, 21 (2):30–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1996:MOB

[Ins96c] Shawn Instenes. My own boss. ;login: the USENIX Association newsletter, 21(4):39–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1996:RBI

[Ins96d] Shawn Instenes. Review: Building Internet firewalls. ;login: the USENIX Association newsletter, 21(1):47–??, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1996:WBT

[Ins96e]  Shawn Instenes.  Watching

browser technologies. ;login: the USENIX Association newsletter, 21(5):22–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1996:WAD

[Ins96f] Shawn Instenes. Where’s the ark? Dealing with SYN flooding. ;login: the USENIX Association newsletter, 21(6):8– ??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Instenes:1997:SEU

[Ins97a] Shawn Instenes. Spam, email, Usenet, and spam. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/instenes.html.

Instenes:1997:SSW

[Ins97b] Shawn Instenes. Stack smashing: What to do? ;login: the USENIX Association newsletter, 22(2):18–??, April 1997. CODEN LOGNEM. ISSN 1044-

Ingham:1994:DWA

[IP94] David B. Ingham and Graham D. Parrington. Delayline: a wide-area network emulation tool. In USENIX [USE94c], pages 313–332.

Inoue:1991:STE

[IS91] Seiki Inoue and Masahiro Shibata. Spacio-temporal editing for HDTV program production. In USENIX [USE91g], pages 95– 104. LCCN QA 76.76 O63 U83 1991.

Idaho:1993:O

[IS93] Duncan Idaho and Nick Stoughton. Opinion. ;login: the USE-NIX Association newsletter, 18 (4):29–33, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Itzkovitz:1999:MMF

[IS99] Ayal Itzkovitz and Assaf Schuster. MultiView and Millipage — fine-grain sharing in pagebased DSMs. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/itzkovitz.
html.

Isaak:1995:II

[Isa95] Jim Isaak. Information infrastructure. ;login: the USE-NIX Association newsletter, 20 (5):42–43, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Islam:1995:RUC

[Isl95] Nayeem Islam. Report on the USENIX conference on objectoriented technologies. ;login: the USENIX Association newsletter, 20(5):6–8, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Ismert:1998:MCP

[Ism98] Michael Ismert. Making commodity PCs fit for signal processing. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/ismert.html.

Itzkovitz:1997:MUL

[ISS97] Ayal Itzkovitz, Assaf Schuster, and Lea Shalev. Millipede: a user-level NT-based distributed shared memory system with thread migration and dynamic run-time optimization of memory references. In USENIX [USE97f], page

148. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/single_schuster.

html. Extended abstract only.

Wu:1991:LIL

[iW91] Shanun inn Wu. LogiC++:

An integrated logic and objectoriented programming language. In USENIX [USE91k], pages 235–243. LCCN QA76.73.C153 U27 1991.

Juul:1998:ECD

[JABA98] Niels Christian Juul, Kim Viborg Andersen, and Niels Bjørn-Andersen. Electronic commerce in Denmark: The spread of EDI in business-to-business transactions. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec98/juul.html.

Jackson:1993:ISA

[Jac93] Janet Jackson. Impressions of SAGE-AU ’93. ;login: the USE-NIX Association newsletter, 18 (6):13–15, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Jacobson:1994:LPT

[Jac94] Van Jacobson. Learning from the present — things that IP got right and ATM got wrong. In USENIX [USE94f], page ?? ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        hispeed/index.html. Unpub-

lished keynote address.

Jaeger:1995:MWT

[Jae95]  Shannon Jaeger.  Mega-

widgets in Tcl/Tk: Evaluation and analysis. In USENIX [USE95j], pages 43–

52. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/jaeger.html.

Jaspan:1996:DWE

[Jas96] Barry Jaspan. Dual-workfactor encrypted key exchange: Efficiently preventing password chaining and dictionary attacks. In USENIX [USE96i], pages 43–50. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/jaspan.html.

Jennings:1999:VVE

[JB99] James Jennings and Eric Beuscher. Verischemelog: Verilog embedded in Scheme. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663

1999.  URL http://www.
usenix.org/publications/library/
proceedings/dsl99/jennings.
html.

Julin:1991:GES

[JCS^{+}91] Daniel P. Julin, Jonathan J.

Chew, J. Mark Stevenson, Paulo Guedes, Paul Neves, and Paul Roy. Generalized emulation services for Mach 3.0 — overview, experiences and current status. In USENIX [USE91h], pages 13–

26. LCCN QAX 27.

Jagannathan:1994:AEI

[JD94] R. Jagannathan and Chris Dodd. Application experience with an implicitly parallel composition language. In USENIX [USE94k], pages 209–

226. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        vhll/jagganathan.html.

Jeffery:1994:GPI

[Jef94] Clinton L. Jeffery. Graphics programming in Icon version

9. In USENIX [USE94k], pages 157–168. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        vhll/jeffery.html.

Jenson:1997:BRP

[Jen97] Rob Jenson. Book review: Programming with GNU software. ;login: the USENIX Association newsletter, 22(4):66–67, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Jenson:1998:BRT

[Jen98] Rob Jenson. Book review: TCP/IP Network Administration, 2nd ed. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-8/
                                        tcpip.html.

Jain:1994:FNM

[JHC94] Parag K. Jain, Norman C. Hutchinson, and Samuel T. Chanson. A framework for the non-monolithic implementation of protocols in the Xkernel. In USENIX [USE94f], pages 13–30. ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
hispeed/jain.html.

Jensen:1990:WGC

[JJ90] Preben Fisker Jensen and Peter Juhl. Writing a gateway in C++. In USENIX [USE90j], pages 205–214.

Jung:1993:CIS

[JKS^{+}93] Hyun Tae Jung, Hae Lyong Kim, Yang Min Seo, Ghun Choe, Sang Min, Chong Sang Kim, and Kern Koh. Caller ID system in the Internet environment. In USENIX [USE93m], pages 69–78. ISBN 1-880446-55-3. LCCN QA 76.9 A25

U54 1993. URL http://www.
usenix.org/publications/library/
proceedings/sec4/.

Jaeger:1998:OSP

[JLI98] Trent Jaeger, Jochen Liedtke, and Nayeem Islam. Operating system protection for fine-grained programs. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998. URL

http://www.usenix.org/publications/
library/proceedings/sec98/
jaeger.html.

Johnson:1994:CM

[JM94] Stephen C. Johnson and Cleve Moler. Compiling Matlab. In USENIX [USE94k], pages 119–127. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        vhll/index.html.

Jermyn:1999:DAG

[JMM^{+}99] Ian Jermyn, Alain Mayer, Fabian Monrose, Michael K. Reiter, and Aviel D. Rubin. The design and analysis of graphical passwords. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

                                        http://db.usenix.org/publications/
                                        library/proceedings/sec99/
                                        jermyn.html.

Johnson:1990:PFP

[Joh90] S. C. Johnson. Postloading for fun and profit. In USE-NIX [USE90h], pages 325–330. LCCN QA76.8.U65 U82 1990.

Johnson:1992:PL

[Joh92] Stephen C. Johnson. President’s letter. ;login: the USENIX Association newsletter, 17(6):6–??, November/December 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Johnson:1993:PLSa

[Joh93a] Stephen C. Johnson. President’s letter: Some of your best friends are marketeers. … ;login: the USENIX Association newsletter, 18(1):3–4, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1993:PLSb

[Joh93b] Stephen C. Johnson. President’s letter: State of the year. ;login: the USENIX Association newsletter, 18(2):3–4, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1993:PL

[Joh93c] Steve Johnson. President’s letter. ;login: the USENIX Association newsletter, 18(4):3–4, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1993:UWF

[Joh93d] Steve Johnson. United we fall, or killing the goose that laid the golden egg. In USENIX [USE93d], pages 29–36.

Johnson:1994:OO

[Joh94a] Stephen C. Johnson. Objecting to objects. In USENIX [USE94l], page ?? ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994. URL http://

www.usenix.org/publications/
library/proceedings/sf94/
johnson.html.

Johnson:1994:PLa

[Joh94b] Steve Johnson. President’s letter. ;login: the USENIX Association newsletter, 19(5):4–5, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1994:PLb

[Joh94c]  Steve Johnson.  President’s

letter. ;login: the USENIX Association newsletter, 19(6):4– 5, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1994:PLT

[Joh94d] Steve Johnson. President’s letter — A tale of two companies. ;login: the USENIX Association newsletter, 19(1):3–4, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

John:1995:DVD

[Joh95a] Aju John. Dynamic vnodes — design and implementation. In USENIX [USE95e], pages 11–24 (or 11–23??). ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/john.html.

Johnson:1995:PLa

[Joh95b]  Steve Johnson.  President’s

letter. ;login: the USENIX Association newsletter, 20(1):4– ??, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1995:PLb

[Joh95c] Steve Johnson. President’s letter. ;login: the USENIX Association newsletter, 20(3):6–??, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1995:PLc

[Joh95d] Steve Johnson. President’s letter. ;login: the USENIX Association newsletter, 20(5):6–??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1996:F

[Joh96a] Steve Johnson. Farewell. ;login: the USENIX Association newsletter, 21(3):5–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1996:PL

[Joh96b]  Steve Johnson.  President’s

letter. ;login: the USENIX Association newsletter, 21(1):7– 8, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1997:BRH

[Joh97a] Steve Johnson. Book review:

The hackers dictionary, third ed. ;login: the USENIX Association newsletter, 22(5):68–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Johnson:1997:TLM

[Joh97b] Steve Johnson. A technologist looks at management, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/invited_talks/johnson/
johnson.pdf;   http://www.
usenix.org/publications/library/
proceedings/lisa97/invited_
talks/johnson/johnson.ps.

Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26– 31, 1997 San Diego, California, USA.

Johnson:1998:CAU

[Joh98]  Stephen C. Johnson.  C

and the AT&T Unix port

— A personal history, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/invited_talks/johnson.

ps. Unpublished invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Jones:1991:BCL

[Jon91] Michael B. Jones. Bringing the C libraries with us into a multithreaded future. In USENIX [USE91i], pages 81–92. LCCN QA 76.76 O63 U84 1992.

Jones:1994:CWC

[Jon94] Vicki E. Jones. Celebrating women in computing. ;login: the USENIX Association newsletter, 19(4):23–24, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Joncheray:1995:SAA

[Jon95] Laurent Joncheray. A simple active attack against TCP. In USENIX [USE95f], pages 7–20 (or 7–19??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/security95/
                                        joncheray.html.

Jones:1998:CRN

[Jon98] George M. Jones. Conference report: NTnix … You Are There. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt98/ntnix.html.

Joshi:1999:WMR

[JP99] Nikhil Joshi and Tom Phillips.

Windows(R) management roadmap,

1999.  URL http://www.
usenix.org/publications/library/
proceedings/lisa-nt98/phillips_

html. Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Jaeger:1996:BSF

[JPR96] Trent Jaeger, Atul Prakash, and Avi Rubin. Building systems that flexibly download executable content. In USENIX [USE96i], pages 131–

148. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/.

Jones:1991:CCH

[JR91] George M. Jones and Steven M.

Romig. Cloning customized hosts(or customizing cloned hosts). In USENIX [USE91f], pages 233–242.

Jones:1999:CRT

[JR99] Michael B. Jones and John Regehr. CPU reservations and time constraints: Implementation experience on Windows NT. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/jones.html.

Jacob:1993:MBT

[JS93] Betty Jacob and Nancy Shoemaker. The Myers-Briggs type indicator: An interpersonal tool for system administrators. In USENIX [USE93f], pages 7–15. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/.

Johnson:1995:CPS

[JS95] Ray Johnson and Scott Stanton. Cross platform support in Tk. In USENIX [USE95j], pages 77–84 (or 77– 83??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/johnson.html.

Jain:1997:SCP

[JS97] Prashant Jain and Douglas C. Schmidt. Service configurator: a pattern for dynamic configuration of services. In USENIX [USE97g], pages 209–

219. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots97/jain.html. See also

[LH97, Rie97, WW97].

Jossman:1990:CCL

[JSS90] Paul R. Jossman, Edward N. Schiebel, and Jere C. Shank. Climbing the C++ learning tree. In USENIX [USE90j], pages 11– 23 (or 11–24??).

Juszczak:1994:IWP

[Jus94]  Chet Juszczak.  Improv-

ing the write performance of an NFS server. In USE-NIX [USE94l], pages 247–259. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/juszczak.html.

Jordan:1995:SCM

[JV95a] M. Jordan and M. L. Van De Vanter. Software configuration management in an object-oriented database. In USENIX [USE95k], pages 55–

68. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/jordan.html.

Jordan:1995:CMO

[JV95b] Mick Jordan and Michael Van De Vanter. Configuration management in an object-oriented database. In USENIX [USE95k], pages 55–67. ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/index.html.

Johnson:1999:HN

[JW99] Steve Johnson and Dusty White.

How now. ;login: the USE-NIX Association newsletter, 24 (6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/hownow.html.

Jain:1998:DPM

[JWS98] Prashant Jain, Seth Widoff, and Douglas C. Schmidt. The design and performance of Med-Java. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots98/jain.html.

Jutla:1996:PAS

[JY96] Charanjit Jutla and Moti Yung. PayTree: “amortized-signature” for flexible MicroPayments. In USENIX [USE96h], pages 213–221. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec96/index.html.

Kendall:1994:SBT

[KA94] Samuel Kendall and Glenn Allinn. Sharing between translation units in C++ program databases. In USE-NIX [USE94e], pages 247–263. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        c++94/kendall.html.

Kordale:1996:OCCa

[KAD96a] R. Kordale, M. Ahamad, and M. Devarakonda. Object caching in a CORBA compliant system. In USE-NIX [USE96a], pages 377–

                                        404. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/coots96/kordale.
                                        html.

Kordale:1996:OCCb

[KAD96b] R. Kordale, M. Ahamad, and M. Devarakonda. Object caching in a CORBA compliant system. In USENIX [USE96g], pages 65–81. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/kordale.html.

Kahn:1995:SUX

[Kah95] Brian L. Kahn. Safe use of X Window System protocol across a firewall. In USENIX [USE95f], pages 105–116. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
kahn.html.

Kamens:1993:RNS

[Kam93] Jonathan I. Kamens. Retrofitting network security to third-party applications — the SecureBase experience. In USENIX [USE93m], pages 41–

57. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/kamens.html.

Kamath:1998:MUH

[Kam98a] Sean Kamath. Making users happy. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
comingup.html.

Kamp:1998:MR

[Kam98b] Poul-Henning Kamp. Malloc(3) revisited. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix98/freenix/kamp.ps.

Kandori:1991:WBM

[Kan91] Keishi Kandori. A workstationbased multi-media environment for broadcast television. In USE-NIX [USE91g], pages 455–461. LCCN QA 76.76 O63 U83 1991.

Kaplan:1992:HCO

[Kap92a] Alan E. Kaplan. A history of the COSNIX operating system: Assembly language UNIX 1970 to July 1991. In USENIX [USE92j], pages 429–438.

Kapor:1992:BOR

[Kap92b] Mitch Kapor. Building the open road: The Internet as a testbed for the National Public Network. In USENIX [USE92j], pages 1–2.

Kaplan:1994:HAL

[Kap94]  Harry Kaplan.  Highly au-

tomated low personnel system administration in a Wall Street environment. In USE-NIX [USE94h], pages 185–

189. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/kaplan.html.

Kaplan:1997:DJ

[Kap97] Philippe Kaplan. Debugging in Java. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/javaindex.
html;   http://www.usenix.
org/publications/java/usingjava8.
html.

Kardel:1990:FF

[Kar90] Frank Kardel. Frozen files. In USENIX [USE90i], pages 83–86. LCCN QA 76.9 A25 U55 1990.

Karn:1993:QCD

[Kar93] Phil Karn. The Qualcomm CDMA digital cellular system. In USENIX [USE93i], pages 35–39. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/karn.html.

Karp:1996:HFI

[Kar96] Yonah D. Karp. Highlights from the fourth international WWW conference. ;login: the USE-NIX Association newsletter, 21 (2):47–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Katz:1994:RB

[Kat94] Lou Katz. Review of BSD/ 386. ;login: the USENIX Association newsletter, 19(1):29–30, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Kawin:1996:WIR

[Kaw96] Rick Kawin. When it rains. … ;login: the USENIX Association newsletter, 21(2):25–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Khalidi:1996:SMM

[KBM^{+}96] Yousef A. Khalidi, Jose M. Bernabeu, Vlada Matena, Ken Shirriff, and Moti Thadani. Solaris MC: a multi computer OS. In USENIX [USE96k], pages 191–203 (or 191–204??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/khalidi.html.

Kling:1993:RC

[KC93] Rob Kling and Lisa Covi. Review: Connections. ;login: the USENIX Association newsletter, 18(4):41–42, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Kotz:1996:ELS

[KC96] David Kotz and Preston Crow. The expected lifetime of singleaddress-space operating systems. In USENIX [USE96c], pages 155–178.

claffy:1999:SAI

[kc99a]  kc claffy.  State of the

art in Internet measurement and data analysis: Topology, workload, performance and routing statistics, 1999.

URL http://db.usenix.org/
publications/library/proceedings/
lisa99/invitedtalks.html.

Unpublished invited talk at LISA XIII.

Kon:1999:SAC

[KC99b] Fabio Kon and Roy H. Campbell. Supporting automatic configuration of component-based distributed systems. In USE-NIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/kon.html.

Krajewski:1994:ASC

[KCCT94] Marjan Krajewski, Jr., John C.

Chipchak, David A. Chodorow, and Jonathan T. Trostle. Application of smart cards to network user authentication. In USENIX [USE94d], pages 75–89.

Keleher:1994:TMD

[KCDZ94] Peter Keleher, Alan L. Cox, Sandhya Dwarkadas, and Willy Zwaenepoel. Tread marks: Distributed shared memory on standard workstations and operating system. In USE-NIX [USE94l], pages 115–132. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/keleher.html.

Karoly:1998:WTI

[KCG98] David Karoly, Todd Copeland, and David Gardner. WinACIF: a telecom IC support tool using Tcl/Tk. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/

                                        proceedings/tcl98/karoly.
                                        html.

Krishna:1995:CBA

[KCVP95] P. Krishna, M. Chatterjee, N. H. Vaidya, and D. K. Pradhan. A cluster-based approach for routing in ad-hoc networks. In USENIX [USE95i], pages 1–10. ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/krishna.html.

Kantarjiev:1993:EXW

[KDF^{+}93] Christopher A. Kantarjiev, Alan Demers, Ron Frederick, Robert T. Krivacic, and Mark Weiser. Experiences with X in a wireless environment. In USENIX [USE93i], pages 117–

128. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mobile93/kantarjiev.html.

Keith:1990:PNF

[Kei90]  Bruce E. Keith.  Perspec-

tives on NFS file server performance characterization. In USE-NIX [USE90g], pages 267–278. LCCN QA76.8.U65 U81 1990.

Kellum:1992:WTa

[Kel92a] Jeff Kellum. What’s out there? ;login: the USENIX Association newsletter, 17(3):3–??, May/ June 1992. CODEN LOGNEM. ISSN 1044-6397.

Kellum:1992:WTb

[Kel92b] Jeff Kellum. What’s out there?

;login: the USENIX Association newsletter, 17(5):20–22, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Kellem:1993:WT

[Kel93] Jeff Kellem. What’s out there?

;login: the USENIX Association newsletter, 18(3):15–18, May/ June 1993. CODEN LOGNEM. ISSN 1044-6397.

Keller:1995:SCV

[Kel95] Arthur M. Keller. Smart catalogs and virtual catalogs. In USENIX [USE95g], pages 125– 131 (or 125–132??). ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1). URL http://www.

usenix.org/publications/library/
proceedings/ec95/keller.html.

Keleher:1999:THL

[Kel99] Peter J. Keleher. Tapeworm:

High-level abstractions of shared accesses. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/keleher.
html.

Kempf:1993:IHR

[Kem93] James Kempf. Integrating handwriting recognition into UNIX. In USENIX [USE93g], pages 187–204. ISBN 1-880446-50-2.

LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/kempf.html.

Kenyon:1990:IIA

[Ken90] Simon Kenyon. IAW — the Intelligence Analyst Workbench. In USENIX [USE90g], pages 193–198. LCCN QA76.8.U65 U81 1990.

Kennedy:1991:FOO

[Ken91] Brian Kennedy. The features of the object-oriented abstract type hierarchy (OATH). In USENIX [USE91k], pages 41–

50. LCCN QA76.73.C153 U27

1991.

Kenny:1996:TAC

[Ken96] Kevin B. Kenny. TclSolver: An algebraic constraint manager for Tcl. In USE-NIX [USE96e], pages 69–74. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/kenny.html.

Kernighan:1982:PLT

[Ker82] Brian W. Kernighan. PIC: a language for typesetting graphics. Software—Practice and Experience, 12(1):1–22, January 1982.

Kernighan:1991:ACT

[Ker91] Brian Kernighan. An AWK to C++ translator. In USE-NIX [USE91k], pages 217–228. LCCN QA76.73.C153 U27 1991.

Kernighan:1995:ETT

[Ker95] Brian W. Kernighan. Experience with Tcl/Tk for scientific and engineering visualization. In USENIX [USE95j], pages 269–278. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/kernighan.html.

Keves:1993:OSF

[Kev93] Brian William Keves. Open systems formal evaluation process. In USENIX [USE93f], pages 87–91. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/keves.html.

Kumar:1998:IA

[KF98] Manoj Kumar and Stuart I.

Feldman. Internet auctions. In USENIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http:/

/www.usenix.org/publications/
library/proceedings/ec98/
kumar_auctions.html.

Koch:1994:MDR

[KFJ94] Povl T. Koch, Robert J. Fowler, and Eric Jul. Message-driven relaxed consistency in a software distributed shared memory. In USENIX [USE94i], pages 75–85. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/koch.html.

Kolstad:1993:ORA

[KG93] Rob Kolstad and Dan Geer. Opening remarks and announcements. In USENIX [USE93l], page ?? ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Katseff:1991:ELN

[KGL^{+}91] Howard P. Katseff, Robert D. Gaglianello, Thomas B. London, Bethany S. Robinson, and Donald B. Swicker. Experiences with the Liaison network multimedia workstation. In USENIX [USE91j], pages 341–347. ISBN ???? LCCN ????

Ketchpel:1996:UPU

[KGMP^{+}96] Steven P. Ketchpel, Hector Garcia-Molina, Andreas Paepcke, Scott Hassan, and Steve Cousins. U-PAI: a universal payment application interface. In USENIX [USE96h], pages 105–121. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Kamin:1997:SPL

[KH97a] Samuel N. Kamin and David Hyatt. A special-purpose language for picture-drawing. In USENIX [USE97b], pages 297–

310. ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/kamin.html. This paper

describes FPIC, an implementation in ML of a programmable extension to a PIC-like [Ker82] language. The authors raise interesting points about the difficulty of doing this correctly in ML, and conclude that a more powerful base language (e.g., Haskell) might be more suitable. FPIC includes support for both PostScript and L^{A}TEX output.

Kouvelas:1997:OWS

[KH97b] Isidor Kouvelas and Vicky Hardman. Overcoming workstation scheduling problems in a real-time audio tool. In USENIX [USE97e], pages 235–

242. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/kouvelas.html.

Khalili:1998:SFD

[Kha98]  Aram Khalili.  Samba fu-

ture development plans. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/allison.
ps.

Kedem:1999:BFA

[KI99] Gershon Kedem and Yuriko Ishihara. Brute force attack on UNIX passwords with SIMD computer. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

http://db.usenix.org/publications/

                                        library/proceedings/sec99/
                                        kedem.html.

King:1990:APC

[Kin90] Maria King. An automatic policy checker for controlling undesirable program behaviors. In USENIX [USE90i], pages 31–32. LCCN QA 76.9 A25 U55 1990.

Kint:1991:SSC

[Kin91] Richard W. Kint. SCRAPE (system configuration, resource and process exception) monitor. In USENIX [USE91f], pages 217–226.

Kislitzin:1990:NMS

[Kis90] Katy Kislitzin. Network monitoring by scripts. In USENIX [USE90f], pages 101–106. LCCN QA76.76.O63 L37 1990.

Kislitzin:1991:NMS

[Kis91] Katy Kislitzin. Network monitoring by scripts. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Koblas:1992:S

[KK92a] David Koblas and Michelle R. Koblas. SOCKS. In USENIX [USE92m], pages 77–83. ISBN 1-880446-46-4. LCCN ????

Krell:1992:CCO

[KK92b] Eduardo Krell and Balachander Krishnamurthy. COLA: Customized overlaying. In USENIX [USE92j], pages 3–8.

Kommerling:1999:DPT

[KK99] Oliver Kömmerling and Markus G. Kuhn. Design principles for tamper-resistant Smartcard processors. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/kommerling.html.

Katseff:1993:FDB

[KL93] Howard P. Katseff and Thomas B.

London. The Ferret document browser. In USE-NIX [USE93g], pages 101–110. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/katseff.html.

Kroeger:1996:PFS

[KL96] Thomas M. Kroeger and Darrell D. E. Long. Predicting file-system actions from prior events. In USENIX [USE96k], pages 319–328. ISBN 1-880446-76-6. LCCN QA 76.76 O63

U88 1996. URL http://www.
usenix.org/publications/library/
proceedings/sd96/kroeger.
html.

Kroeger:1997:EBW

[KL97] Thomas M. Kroeger and Darrell D. E. Long. Exploring the bounds of Web latency reduction from caching and prefetching. In USE-NIX [USE97h], pages 13–22. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/kroeger.html.

Kazar:1990:DFS

[KLA^{+}90] Michael L. Kazar, Bruce W. Leverett, Owen T. Anderson, Vasilis Apostolides, Ben A. Bottos, Sailesh Chutani, Craig F. Everhart, W. Antony Mason, Shu-Tsui Tu, and Edward R. Zayas. DEcorum file system architectural overview. In USE-NIX [USE90g], pages 151–164. LCCN QA76.8.U65 U81 1990.

Khetawat:1998:IBD

[KLB98] Amit Khetawat, Hemang Lavana, and Franc Brglez. Internetbased desktops in Tcl/Tk: Collaborative and recordable. In USENIX [USE98f], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998. URL

                                        http://db.usenix.org/publications/
                                        library/proceedings/tcl98/
                                        khetawat.html.

Klein:1990:FCS

[Kle90a]  Daniel Klein.  Foiling the

cracker: a survey of, and improvements to, password security. In USENIX [USE90i], pages 5–14. LCCN QA 76.9 A25 U55 1990.

Klein:1990:SIP

[Kle90b] Daniel Klein. A survey of, and improvements to, password security. In USENIX [USE90i], pages 5–14. LCCN QA 76.9 A25 U55 1990.

Klein:1994:ADU

[Kle94]  Daniel Klein.  Applications

development with a UIMS. In USENIX [USE94g], pages 37–56. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/klein.html.

Klein:1997:CTP

[Kle97] Dan Klein. Call for tutorial proposals. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Klein:1998:SDSa

[Kle98a]  Daniel Klein.  Succumb-

ing to the dark side of the force: The Internet as seen from an adult Website, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/invited_talks/klein.

ps. Unpublished invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Klein:1998:SDSb

[Kle98b]  Daniel Klein.  Succumb-

ing to the dark side of the force: The Internet as seen from an adult Website, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/invited_talks/klein.
pdf;   http://www.usenix.
org/publications/library/
proceedings/lisa98/invited_
talks/klein_html/. Unpub-

lished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6– 11, 1998 Boston, Massachusetts, USA.

Klein:1999:CTP

[Kle99a] Daniel V. Klein. Call for tutorial proposals. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-6/
                                        tutorial.html.

Klein:1999:DAW

[Kle99b]  Daniel V. Klein.  Defend-

ing against the wily surfer

— Web-based attacks and defenses. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        detection99/klein.html.

Klein:1999:HTB

[Kle99c] Daniel V. Klein. Hot topics/buzzword bingo, 1999.

                                        URL http://db.usenix.org/
                                        publications/library/proceedings/
                                        lisa99/practicum.html. Un-

published practicum talk at

LISA XIII.

Kluth:1990:SSS

[Klu90] Bernd T. Kluth. SIC — A system for stochastic simulation in C++. In USENIX [USE90j], pages 125–136.

Koblas:1992:PRM

[KM92] David Koblas and Paul M. Moriarty. PITS: a request management system. In USENIX [USE92e], pages 197–202. ISBN 1-880446-47-2. LCCN ????

Kaashoek:1992:TFT

[KMBT92] M. Frans Kaashoek, Raymond Michiels, Henri E. Bal, and A. Tanenbaum. Transparent fault-tolerance in parallel Orca programs. In USENIX [USE92l], pages 297–312. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Koivisto:1991:OOO

[KMR91] Juha Koivisto, Juhani Malka, and James Reilly. OTSO — an object-oriented approach to distributed computation. In USE-NIX [USE91k], pages 163–178. LCCN QA76.73.C153 U27 1991.

Keeton:1993:PCO

[KMS^{+}93] Kimberly Keeton, Bruce A.

Mah, Srinivasan Seshan, Randy H. Katz, and Domenico Ferrari. Providing connection-oriented network services to mobile hosts. In USENIX [USE93i], pages 83–102. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/keeton.html.

Khalidi:1993:IUO

[KN93a] Yousef A. Khalidi and Michael N.

Nelson. An implementation of UNIX on an object-oriented operating system. In USE-NIX [USE93l], pages 469–479. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Khalidi:1993:FEP

[KN93b] Yousef A. Khalidi and Mike N. Nelson. A flexible external paging interface. In USE-NIX [USE93k], pages 127–

140. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/khalidi.html.

Knight:1996:SSA

[Kni96] Zanna Knight. SAGE 1995 system administrator profile and salary survey. ;login: the USE-NIX Association newsletter, 21 (1):10–??, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Kawaguchi:1995:FBF

[KNM95] Atsuo Kawaguchi, Shingo Nishioka, and Hiroshi Motoda. A flash-memory based file system. In USENIX [USE95e], pages 155–164. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neworl/kawaguchi.html.

Knowles:1997:BRS

[Kno97] Brad Knowles. Book review: sed & awk. 2nd ed. ;login: the USE-NIX Association newsletter, 22 (4):68–??, August 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Koenig:1992:SET

[Koe92] Andrew Koenig. Space-efficient trees in C++. In USENIX [USE92k], pages 117–129. ISBN 1-880446-45-6. LCCN ????

Koenig:1994:AAM

[Koe94] Andrew Koenig. An anecdote about ML type inference. In USENIX [USE94k], pages 1–5. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/koenig.html.

Kohl:1999:WTA

[Koh99]  Neil Kohl.  Webmaster’s

toolbox: Apache modules. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
features/webmasters.html.

Kolstad:1990:SUR

[Kol90]  Rob Kolstad.  A summer

USENIX retrospective. UNIX review, 8(9):74–??, September 1990. CODEN UNRED5. ISSN 0742-3136.

Kolstad:1991:NSB

[Kol91] Rob Kolstad. A new step in backup and restore technology. In USENIX [USE91f], pages 73– 80.

Kolstad:1992:SVC

[Kol92] Rob Kolstad. SAGE views:

Counterpoint. ;login: the USE-NIX Association newsletter, 17 (5):17–??, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1993:MAA

[Kol93a] Rob Kolstad. Middle aged addiction. ;login: the USENIX Association newsletter, 18(2):33– ??, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1993:RUP

[Kol93b] Rob Kolstad. Review: UNIX power tools. ;login: the USE-NIX Association newsletter, 18 (3):38–39, May/June 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1993:UCO

[Kol93c] Rob Kolstad. USA computing Olympiad. ;login: the USENIX Association newsletter, 18(4): 26–27, July/August 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:LTE

[Kol94a] Rob Kolstad. 1992 LISA time expenditure survey. ;login: the USENIX Association newsletter, 19(3):14–17, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:ETM

[Kol94b] Rob Kolstad. From the editor: Technology marches on. ;login: the USENIX Association newsletter, 19(4):3–??, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:EWJ

[Kol94c] Rob Kolstad. From the editor: The world just keeps getting more complicated. ;login: the USENIX Association newsletter, 19(6):3–??, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:EWI

[Kol94d] Rob Kolstad. From the editor:

Woodstock II. ;login: the USE-NIX Association newsletter, 19 (5):3–??, October 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:RQC

[Kol94e] Rob Kolstad. Review: a Quarter Century of UNIX. ;login: the USENIX Association newsletter, 19(3):35–??, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:RCE

[Kol94f] Rob Kolstad. Review: Computer Ethics: Cautionary Tales and Ethical Dilemmas in Computing. ;login: the USENIX Association newsletter, 19(3):34– ??, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:REC

[Kol94g] Rob Kolstad. Review: Ethical Conflicts in Information and Computer Science, Technology, and Business. ;login: the USE-NIX Association newsletter, 19 (4):42–43, August 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:RIS

[Kol94h] Rob Kolstad. Review: Information Security Policies Made Easy. ;login: the USENIX Association newsletter, 19(3):35– ??, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:RLP

[Kol94i] Rob Kolstad. Review: Learning Perl. ;login: the USENIX Association newsletter, 19(1):49–??, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1994:YWH

[Kol94j] Rob Kolstad. So you want to hire a cheap system administrator. ;login: the USENIX Association newsletter, 19(1):28–??, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:BRE

[Kol95a] Rob Kolstad. Book review: Expert C Programming: Deep C Secrets. ;login: the USENIX Association newsletter, 20(5):49– ??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:EDF

[Kol95b] Rob Kolstad. From the editor: Dreams fulfilled. ;login: the USENIX Association newsletter, 20(4):3–??, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:EGI

[Kol95c] Rob Kolstad. From the editor: Good ideas. ;login: the USE-NIX Association newsletter, 20 (3):3–??, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:ESP

[Kol95d] Rob Kolstad. From the editor:

Sustaining performance. ;login: the USENIX Association newsletter, 20(2):3–??, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:EWR

[Kol95e] Rob Kolstad. From the editor:

Wherein RBK’s eyes are opened. ;login: the USENIX Association newsletter, 20(5):3–4, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:EW

[Kol95f] Rob Kolstad. From the editor: Words. ;login: the USE-NIX Association newsletter, 20 (1):3–??, February 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:IJO

[Kol95g] Rob Kolstad. Interview with John Ousterhout. ;login: the USENIX Association newsletter, 20(6):34–36, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:LN

[Kol95h] Rob Kolstad. Love those numbers. ;login: the USENIX Association newsletter, 20(6):3– ??, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:RCR

[Kol95i] Rob Kolstad. Review: Computer-Related Risks. ;login: the USE-NIX Association newsletter, 20 (1):47–??, February 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1995:TB

[Kol95j] Rob Kolstad. Thanks to Bryan. ;login: the USENIX Association newsletter, 20(2):28–??, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:EAR

[Kol96a] Rob Kolstad. From the editor: Another revelation. ;login: the USENIX Association newsletter, 21(3):3–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:EJ

[Kol96b] Rob Kolstad. From the editor: Journalism. ;login: the USE-NIX Association newsletter, 21 (5):3–??, October 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:H

[Kol96c] Rob Kolstad. Hype. ;login: the USENIX Association newsletter, 21(2):3–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:IAH

[Kol96d] Rob Kolstad. Interview with Andrew Hume. ;login: the USE-NIX Association newsletter, 21 (4):40–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:IDA

[Kol96e] Rob Kolstad. Interview with Dan Appelman. ;login: the USENIX Association newsletter, 21(2):37–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:IJW

[Kol96f] Rob Kolstad. Interview with Jim Waldo. ;login: the USE-NIX Association newsletter, 21 (3):27–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:IMO

[Kol96g] Rob Kolstad. Interview with Mike O’Dell. ;login: the USE-NIX Association newsletter, 21 (6):12–13, December 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:ISJ

[Kol96h] Rob Kolstad. Interview with Steve Johnson. ;login: the USE-NIX Association newsletter, 21 (1):20–22, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:LAF

[Kol96i] Rob Kolstad. Life is always full of surprises. ;login: the USE-NIX Association newsletter, 21 (4):3–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:UTF

[Kol96j] Rob Kolstad. USACO taps four for international Olympiad. ;login: the USENIX Association newsletter, 21(4):5–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:UPC

[Kol96k] Rob Kolstad. The user/product conundrum. ;login: the USENIX Association newsletter, 21(6):1– ??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1996:W

[Kol96l] Rob Kolstad. Working. ;login: the USENIX Association newsletter, 21(1):3–??, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:BRD

[Kol97a] Rob Kolstad. Book review: Death March. ;login: the USE-NIX Association newsletter, 22 (7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
yourdon.html.

Kolstad:1997:DSI

[Kol97b] Rob Kolstad. Discussing security: An interview with Char Sample and Mark Teicher. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
interview.html.

Kolstad:1997:GO

[Kol97c] Rob Kolstad. Growing old. ;login: the USENIX Association newsletter, 22(1):1–??, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:IRI

[Kol97d] Rob Kolstad. Is it real or is it Memorex? ;login: the USE-NIX Association newsletter, 22 (5):1–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:MPS

[Kol97e] Rob Kolstad. Marching to Pretoria (or somewhere). ;login: the USENIX Association newsletter, 22(4):1–??, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:Ma

[Kol97f] Rob Kolstad. motd. ;login:

the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:Mb

[Kol97g] Rob Kolstad. motd. ;login:

the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:TAF

[Kol97h] Rob Kolstad. Thanks for all the fish. ;login: the USENIX Association newsletter, 22(2):1–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1997:TSL

[Kol97i] Rob Kolstad. Tuning sendmail for large mailing lists. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997. URL

http://www.usenix.org/publications/
library/proceedings/lisa97/
21.kolstad.html.

Kolstad:1997:UM

[Kol97j] Rob Kolstad. USACO marches on. ;login: the USENIX Association newsletter, 22(2):6–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1998:UOU

[Kol98a] Rob Kolstad. 1998 US Open USACO programming championships. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
tcpip.html.

Kolstad:1998:IBC

[Kol98b] Rob Kolstad. Interview with Bill Cheswick. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
interview.html.

Kolstad:1998:IDH

[Kol98c] Rob Kolstad. Interview with Dan Hildebrand. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
interview.html.

Kolstad:1998:IDC

[Kol98d]  Rob Kolstad.  Interview

with Dr. Clair W. Goldsmith. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
interview.html.

Kolstad:1998:IJS

[Kol98e] Rob Kolstad. Interview with John Stewart. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
interview.html.

Kolstad:1998:Ma

[Kol98f] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-4/mvhs.html.

Kolstad:1998:Mb

[Kol98g] Rob Kolstad. motd. ;login:

the USENIX Association newsletter, 23(4), June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1998:Mc

[Kol98h] Rob Kolstad. motd. ;login:

the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-8/tcpip.html.

Kolstad:1998:Md

[Kol98i] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/mailing.html.

Kolstad:1998:Me

[Kol98j] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-12/satan.html.

Kolstad:1998:UBG

[Kol98k] Rob Kolstad. USA bags gold at IOI. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
satan.html.

Kolstad:1998:UTS

[Kol98l] Rob Kolstad. USA team scores gold. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
reference.html.

Kolstad:1998:UST

[Kol98m] Rob Kolstad. USACO spring tournament results. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Kolstad:1999:GCD

[Kol99a] Rob Kolstad. The great certification debate: Rob Kolstad interviews Barb Dijker. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:1999:Ma

[Kol99b] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-2/motd.html.

Kolstad:1999:Mb

[Kol99c] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-4/motd.html.

Kolstad:1999:Mc

[Kol99d] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-6/motd.html.

Kolstad:1999:Md

[Kol99e] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1999-8/motd.html.

Kolstad:1999:Me

[Kol99f] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1999-10/motd.html.

Kolstad:1999:Mf

[Kol99g] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1999-12/motd.html.

Kolstad:1999:UWB

[Kol99h] Rob Kolstad. USACO wins Baltic Olympiad. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/login/
                                        1999-6/usaco.html.

Kolstad:19xx:ERS

[Kolxx] Rob Kolstad. The evolving role of the system administrator. In ????, page ?? USENIX, Berkeley, CA, USA, 19xx. Title listed in contents only, no paper, no abstract.

Koontz:1993:PBM

[Koo93] Kenneth W. Koontz. Port buffers: a Mach IPC optimization for handling large volumes of small messages. In USENIX [USE93h], pages 89–

102. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Kim:1995:ISR

[KOO95] Gene H. Kim, Hilarie Orman, and Sean O’Malley. Implementing a secure rlogin environment: a case study of using a secure network layer protocol. In USENIX [USE95f], pages 65–74 (or 65–73??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
kim.html.

Korn:1994:KEH

[Kor94] David G. Korn. ksh: An extensible high level language. In USENIX [USE94k], pages 129–146. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/korn.ksh.html.

Korn:1996:TTL

[Kor96] Jeffrey Korn. Tksh: a Tcl library for KornShell. In USENIX [USE96e], pages 149–

159. ISBN 1-880446-78-2.

LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/korn.html.

Korn:1997:UWN

[Kor97a]  David Korn.  From UNIX

to Windows NT — A long day’s journey into night, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/invited_talks/scarr/

sld028.htm. Unpublished keynote address at the Large Scale System Administration of Windows NT Workshop, 1997, August 14–16, 1997, Seattle, Washington, USA.

Korn:1997:WWG

[Kor97b] David Korn. Working with Win32: The good, the bad, and the ugly. ;login: the USENIX Association newsletter, 22(6): ??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
win32.html.

Korn:1997:PUW

[Kor97c]  David G. Korn.  Porting

UNIX to Windows NT. In USENIX [USE97e], pages 43–

57. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/korn.html;   http://
www.usenix.org/publications/
library/proceedings/ana97/
summaries/korn.html.

Korn:1997:UUWa

[Kor97d]  David G. Korn.  UWIN

— UNIX for Windows. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/korn.html.

Korn:1997:UUWb

[Kor97e] David G. Korn. UWIN — UNIX for Windows. In USE-NIX [USE97f], pages 133–

145. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/korn.html.

Kotz:1992:MFS

[Kot92] David Kotz. Multiprocessor file system interfaces. In USE-NIX [USE92h], pages 149–150. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Kotz:1994:DDM

[Kot94]  David Kotz.  Disk-directed

I/O for MIMD multiprocessors. In USENIX [USE94i], pages 61–74. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/kotz.html.

Kottaridis:1998:BRP

[Kot98] Chris Kottaridis. Book review: PPP Design and Debugging. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
ppp.html.

Kay:1993:MAI

[KP93a] Jonathan Kay and Joseph Pasquale. Measurement, analysis, and improvement of UDP/IP throughput for the DECstation 5000. In USE-NIX [USE93l], pages 249–258. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Knister:1993:IDT

[KP93b] Michael Knister and Atul Prakash. Issues in the design of toolkit for supporting multiple group editors. In USENIX [USE93b], pages 135–166.

Kim:1997:FOO

[KP97] Ted H. Kim and Gerald J. Popek. Frigate: An objectoriented file system for ordinary users. In USE-NIX [USE97g], pages 115–

129. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/kim.html. See also

[LH97, Rie97, WW97].

Keromytis:1998:OII

[KPI98] Angelos D. Keromytis, Niels Provos, and John Ioannidis. OpenBSD IPsec implementation. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/keromytis.
ps.

Kuenning:1994:ATD

[KPR94] Geoffrey H. Kuenning, Gerald J. Popek, and Peter L. Reiher. An analysis of trace data for predictive file caching in mobile computing. In USE-NIX [USE94j], pages 291–303. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/kuenning.html.

Kangas:1999:UMC

[KR99] Kari Kangas and Juha Röning.

Using mobile code interfaces to control ubiquitous embedded systems. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
es99/kangas.html.

Kranenburg:1998:MUN

[Kra98a] Paul Kranenburg. Monitoring utilization in an NT workstation lab. In USENIX [USE98e], pages 11–16. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/kranenburg.html.

Krannig:1998:S

[Kra98b]  Annette Krannig.  Towards

Web security using PLASMA. In USENIX [USE98i], page ??

ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/sec98/
                                        krannig.html.

Kramp:1999:SNB

[Kra99]  William D. Kramp.  Se-

curing an NT-based DNS server. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-10/
                                        features/securing.html.

Kreiling:1994:IID

[Kre94] Amy K. Kreiling. An introduction to Internet discovery & retrieval tools (invited). In USE-NIX [USE94m], pages 49–61. ISBN ???? LCCN ????

Kreiling:1997:IFT

[Kre97] Amy K. Kreiling. Impressions of a first time SAGE board member. ;login: the USENIX Association newsletter, 22(5):12–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Kumar:1998:SPI

[KRJM98] Manoj Kumar, Anand Rangachari, Anant Jhingran, and Rakesh Mohan. Sales promotions on the Internet. In USE-NIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/ec98/kumar_promotions.
                                        html.

Koenig:1990:EHC

[KS90] Andrew Koenig and Bjarne Stroustrup. Exception handling for C++ (revised). In USENIX [USE90j], pages 149–176.

Koenig:1993:EHC

[KS93] Andrew Koenig and Bjarne Stroustrup. Exception of handling for C++. In Waldo [Wal93b], pages 137–171. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Kim:1994:ETU

[KS94a] Gene H. Kim and Eugene H.

Spafford. Experiences with Tripwire: Using integrity checkers for intrusion detection. In USENIX [USE94m], pages 89–

101. ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/kim.html.

Kim:1994:WSE

[KS94b] Gene H. Kim and Eugene H.

Spafford. Writing, supporting, and evaluating Tripwire: a publicly available security tool. In USENIX [USE94g], pages 89–107. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/index.html.

Kofler:1994:PMA

[KS94c] Thomas Kofler and Bruno Schaeffer. A poor man’s approach to string-based interfacing of C++ objects. In USE-NIX [USE94e], pages 237–246.

ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/kofler.html.

Kumar:1995:FSR

[KS95] Puneet Kumar and M. Satyanarayanan. Flexible and safe resolution of file conflicts. In USE-NIX [USE95e], pages 95–106 (or 95–105??). ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/kumar.html.

Klarlund:1997:DSL

[KS97] Nils Klarlund and Michael I. Schwartzbach. A domainspecific language for regular sets of strings and trees. In USE-NIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/klarlund.
html.

Kelsey:1998:ECS

[KS98] John Kelsey and Bruce Schneier. Electronic commerce and the Street Performer Protocol. In USENIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http:/

/www.usenix.org/publications/
library/proceedings/ec98/
schneier.html.

Kelsey:1999:AST

[KS99] John Kelsey and Bruce Schneier. Authenticating secure tokens using slow memory access. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

http://www.usenix.org/publications/
library/proceedings/smartcard99/
kelsey.html.

Kittur:1996:FTD

[KSAL96] Sunil Kittur, Douglas Steel, François Armand, and Jim Lipkis. Fault tolerance in a distributed CHORUS/MiX system. In USENIX [USE96k], pages 219–228. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/kittur.html.

Kimball:1999:ACS

[KSKD99] Conrad E. Kimball, Vincent D.

Skahan, Jr., David J. Kasik, and Roger. L. Droz. Automated client-side integration of distributed application servers. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/kimball.
html.

Kohl:1993:HUL

[KSS93] John T. Kohl, Carl Staelin, and Michael Stonebraker. High-Light: Using a log-structured file system for tertiary storage management. In USE-NIX [USE93l], pages 435–447. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Krieger:1992:EAM

[KSU92] Orran Krieger, Michael Stumm, and Ron Unrau. Exploiting the advantages of mapped files for stream I/O. In USENIX [USE92j], pages 27–42.

Khanna:1992:RSS

[KSZ92] Sandeep Khanna, Michael Sebree, and John Zolnowsky. Realtime scheduling in SunOS 5.0. In USENIX [USE92j], pages 375– 390.

Karsenty:1993:GSE

[KTBL93] Alain Karsenty, Christophe Tronche, and Michel Beaudouin-Lafon. GroupDesign: Shared editing in a heterogeneous environment. In USENIX [USE93b], pages 167–195.

Kitayama:1993:RII

[KTN93] Takuro Kitayama, Hideyuki Tokuda, and Tatsuo Nakajima. RT-IPC: An IPC extension for real-time Mach. In USENIX [USE93k], pages 91–104. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/kitayama.html.

Kimbrel:1996:TDC

[KTP^{+}96] Tracy Kimbrel, Andrew Tomkins, R. Hugo Patterson, Brian Bershad, Pei Cao, Edward W. Felten, Garth Gibson, Anna R. Karlin, and Kai Li. A tracedriven comparison of algorithms for parallel prefetching and caching. In USE-NIX [USE96f], pages 19–

34. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/kimbrel.html.

Kubicki:1992:CSM

[Kub92] Carol Kubicki. Customer satisfaction metrics and measurement. In USENIX [USE92e], pages 63–68. ISBN 1-880446-47-

2. LCCN ????

Kubicki:1993:SAM

[Kub93]  Carol Kubicki.  The Sys-

tem Administration Maturity Model — SAMM. In USE-NIX [USE93f], pages 213–

225. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/kubicki.html.

Kupfer:1993:SM

[Kup93] Michael D. Kupfer. Sprite on Mach. In USENIX [USE93h], pages 307–322. ISBN 1-880446-49-9. LCCN QA 76.76 O63

U85 1993. URL http://www.
usenix.org/publications/library/
proceedings/mach3/.

Kuras:1990:ESA

[Kur90] Joseph Kuras. An expert systems approach to security inspection of UNIX. In USENIX [USE90i], pages 55–56. LCCN QA 76.9 A25 U55 1990.

Korn:1991:SSF

[KV91] David G. Korn and K. Phong Vo. SFIO: Safe/fast string/file I/O. In USENIX [USE91g], pages 235–256. LCCN QA 76.76 O63 U83 1991.

Keith:1993:LNG

[KW93] Bruce E. Keith and Mark Wittle. LADDIS: The next generation in NFS file server benchmarking. In USE-NIX [USE93g], pages 111–128. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/wittle.html.

Kindred:1996:FAC

[KW96] Darrell Kindred and Jeannette Wing. Fast, automatic checking of security protocols. In USENIX [USE96h], pages 41–52. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/kindred.html.

Krishnamurthy:1997:SPC

[KW97] Balachander Krishnamurthy and Craig E. Wills. Study of piggyback cache validation for proxy caches in the World Wide Web. In USE-NIX [USE97h], pages 1–12. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/krishnamurthy.html.

Kakkad:1999:ATS

[KW99] Sheetal V. Kakkad and Paul R.

Wilson. Address translation strategies in the Texas Persistent Store. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/kakkad.html.

Krishnaswamy:1998:EIJ

[KWB^{+}98] Vijaykumar Krishnaswamy, Dan Walther, Sumeer Bhola, Ethendranath Bommaiah, George Riley, Brad Topol, and Mustaque Ahamad. Efficient implementation of Java Remote Method Invocation (RMI). In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/krishnaswamy.html.

Konstantinou:1999:MSD

[KYBR99] Alexander V. Konstantinou, Yechiam Yemini, Sandeep Bhatt, and S. Rajagopalan. Managing security in dynamic networks. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/konstantinou.
html.

Levien:1998:ART

[LA98] Raph Levien and Alex Aiken.

Attack-resistant trust metrics for public key certification. In USENIX [USE98i], page ?? ISBN 1-880446-92-8.

LCCN QA76.9.A25 U83 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec98/levien.html.

Labiaga:1999:EAA

[Lab99] Ricardo Labiaga. Enhancements to the Autofs automounter. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa99/labiaga.
                                        html.

Lacy:1993:CCS

[Lac93]  John B. Lacy.  CryptoLib:

Cryptography in software. In USENIX [USE93m], pages 1–

17. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec4/.

Laeufer:1995:FHF

[Lae95] K. Laeufer. A framework for higher-order functions in C++. In USENIX [USE95k], pages 103–116. ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/index.html.

Lepreau:1996:WIW

[LAG^{+}96] Jay Lepreau, Marc Auslander, Michel Gien, Rob Gingell, Larry McVoy, Jeff Mogul, Rob Short, and John Sontag. What the OS industry wants from OS research, 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/

osdi96/panel/panel.html. Un-

published panel at the USENIX 2nd Symposium on OS Design and Implementation (OSDI ’96), October 28–31, 1996, Seattle, Washington.

Lajoie:1996:NAC

[Laj96] Josee Lajoie. New ANSI C++ features. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

Lalonde:1998:SRN

[Lal98] Chris Lalonde. Software review: Network Flight Recorder. ;login: the USENIX Association newsletter, 23(3): ??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-5/

lalonde.html. Special issue on security.

Lammert:1990:ABS

[Lam90] Steve Lammert. The AFS 3.0 backup system. In USENIX [USE90f], pages 143–147. LCCN QA76.76.O63 L37 1990.

Lammert:1991:ABS

[Lam91a] Steve Lammert. The AFS 3.0 backup system. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Lampen:1991:AFA

[Lam91b] Andreas Lampen. Advancing files to attributed software objects. In USENIX [USE91i], pages 219–230. LCCN QA 76.76 O63 U84 1992.

Lam:1995:DMW

[Lam95] Ioi K. Lam. Designing mega widgets in the Tix library. In USENIX [USE95j], pages 53–60 (or 53–59??). ISBN 1-880446-72-3. LCCN QA76.73.T44

                                        T44 1995. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/tcl95/lam.html.

Lampson:1997:OSS

[Lam97] Butler Lampson. Operating system security meets the Internet, 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/usenix-nt97/technical.

html. Unpublished keynote address at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Langston:1990:LLM

[Lan90] Peter S. Langston. Little languages for music. In USENIX [USE90b], pages 193–288.

Lansky:1991:MDM

[Lan91] Paul Lansky. Musical dreams and musical reality. In USENIX [USE91g], pages 1–2. LCCN QA 76.76 O63 U83 1991.

Langston:1992:RWM

[Lan92] Peter S. Langston. Report on 1992 workshop on microkernels and other kernel architectures. ;login: the USENIX Association newsletter, 17(4):4– 8, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Lance:1994:SAC

[Lan94a] Kate Lance. SAGE-AU code of ethical conduct. ;login: the USENIX Association newsletter, 19(4):19–20, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Lance:1994:SCE

[Lan94b] Kate Lance. SAGE-AU code of ethical conduct. ;login: the USENIX Association newsletter, 19(4):19–20, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Langeveld:1996:SSP

[Lan96] Hank Langeveld. Shell script programming. ;login: the USE-NIX Association newsletter, 21 (4):4–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Langston:19xx:AGE

[Lanxx] Peter S. Langston. Auditals: the graftals of the ear. In ???? USE-NIX, Berkeley, CA, USA, 19xx. Title listed in contents only, no paper, no abstract.

LaPadula:1994:RSA

[LaP94] Leonard J. LaPadula. A rule-set approach to formal modeling of a trusted computer system. In USENIX [USE94d], pages 113– 167.

Lary:1991:DDI

[Lar91] Ellen Lary. DIDDLY: Digital’s Integrated Distributed Database LaboratorY. In USE-NIX [USE91g], pages 107–106. LCCN QA 76.76 O63 U83 1991.

Lark:1994:BTG

[Lar94] Jay S. Lark. Bridging the technology generation gap: Upgrading a network management application to a new technology base. In USE-NIX [USE94g], pages 159–

169. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/lark.html.

Larkins:1999:IRD

[Lar99] D. Brian Larkins. Internet routing and DNS voodoo in the enterprise. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

URL http://db.usenix.org/
publications/library/proceedings/
lisa99/larkins.html.

LaStrange:1990:SXW

[LaS90] Thomas E. LaStrange. swm: An X window manager shell. In USENIX [USE90g], pages 299–

306. LCCN QA76.8.U65 U81 1990.

Laufer:1995:FHO

[Lau95] Konstantin Laufer. A framework for higher-order functions in C++. In USENIX [USE95k], pages 103–116. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/laufer.html.

Law:1994:TBD

[Law94] Ted Law. Template base delegation. In USENIX [USE94e], pages 179–187. ISBN 1-880446-60-X. LCCN QA 76.73 C153

C22 1994. URL http://www.
usenix.org/publications/library/
proceedings/c++94/law.html.

Lazowska:1992:SSH

[Laz92]  Edward Lazowska.  System

support for high performance multiprocessing. In USENIX [USE92l], pages 1–12. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Lazenby:1997:BRN

[Laz97] Daniel Lazenby. Book review:

Netizens: On the History and Impact of Usenet and the Internet. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
hauben.html.

Lazenby:1998:BRD

[Laz98]  Daniel Lazenby.  Book re-

view: Developing Imaging Applications with XIELib. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Lazenby:1999:BRS

[Laz99] Daniel Lazenby. Book review:

SNA and TCP/IP Enterprise Networking. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

                                        publications/login/1999-4/
                                        sna.html.

Laumann:1994:EEL

[LB94] Oliver Laumann and Carsten Bormann. Elk: The Extension Language Kit. In USENIX [USE94a], pages 419–449.

Lai:1996:PCU

[LB96] Kevin Lai and Mary Baker. A performance comparison of UNIX operating systems on the Pentium. In USENIX [USE96k], pages 265–277 (or 265–278??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd96/lai.html.

Loon:1997:ALB

[LB97] Tong Sau Loon and Vaduvur Bharghavan. Alleviating the latency and bandwidth problems in WWW browsing. In USE-NIX [USE97h], pages 219–230. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits97/tong.html.

Lavana:1998:WST

[LB98] Hemang Lavana and Franc Brglez. WebWiseTclTk: a Safe-Tcl/Tk-based toolkit enhanced for the World Wide Web. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/

proceedings/tcl98/lavana.
html.

Lee:1999:RSL

[LBBA99] Dennis Lee, Jean-Loup Baer, Brian Bershad, and Tom Anderson. Reducing startup latency in Web and desktop applications. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/lee.html.

Langerman:1990:HPM

[LBLM90] Alan Langerman, Joseph Boykin, Susan LoVerso, and Shashi Mangalat. A highly-parallelized Mach-based Vnode filesystem. In USENIX [USE90h], pages 297–312. LCCN QA76.8.U65 U82 1990.

Luerkens:1997:IWN

[LBSL97] Cameron D. Luerkens, Mark J.

Bartelme, Kevin K. Sizer, and James ‘Bud’ Lande. Integrating Windows 95/NT into an existing UNIX Enterprise Network. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.collins.
rockwell.com/presentation/
ntpaper/index.htm;   http:
//www.usenix.org/publications/
library/proceedings/nt-sysadmin97/
luerkens.html.

Linton:1993:DII

[LC93] Mark A. Linton and Paul R.

Calder. The design and implementation of InterViews. In Waldo [Wal93b], pages 75–86.

ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Li:1999:SWC

[LC99] Dan Li and David R. Cheriton. Scalable Web caching of frequently updated objects using reliable multicast. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/li.html.

Lee:1994:HHP

[LCC94] Chao-Hsien Lee, Meng Chang Chen, and Ruei-Chuan Chang. HiPEC: High performance external virtual memory caching. In USENIX [USE94i], pages 153–164. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/lee.html.

Lehman:1992:CNM

[LCH92] R. Lehman, G. Carpenter, and N. Hien. Concurrent network management with a distributed management tool. In USENIX [USE92e], pages 235–244. ISBN 1-880446-47-2. LCCN ????

Luerkens:1998:SDP

[LCL98] Cameron D. Luerkens, H. John Cole, and Danielle R. Legg. Software distribution to PC clients in an enterprise network. In USENIX [USE98e], pages 49–54. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/luerkens.html.

Lei:1997:AAF

[LD97] Hui Lei and Dan Duchamp.

An analytical approach to file prefetching. In USE-NIX [USE97e], pages 275–

288. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/lei.html;   http://
www.usenix.org/publications/
library/proceedings/ana97/
summaries/lei.html.

Levy:1998:STS

[LDOW98] Jacob Y. Levy, Laurent Demailly, John K. Ousterhout, and Brent B. Welch. The Safe-Tcl security model. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix98/
levy.html.

Lea:1990:CC

[Lea90a] Douglas Lea. Customization in C++. In USENIX [USE90j], pages 301–314 (or 301–315??).

Leach:1990:BRU

[Lea90b] George W. Leach. Book review:

UNIX System Software Readings. ;login: the USENIX Association newsletter, 15(1):15–??, January/February 1990. CO-DEN LOGNEM. ISSN 1044-

Lea:1992:RTT

[Lea92a] Doug Lea. Run time type information and class design. In USENIX [USE92k], pages 341–

347. ISBN 1-880446-45-6. LCCN

????

Leach:1992:BRP

[Lea92b] George Leach. Book review: Practical UNIX Security. ;login: the USENIX Association newsletter, 17(3):29–??, May/ June 1992. CODEN LOGNEM. ISSN 1044-6397.

Leach:1993:RCP

[Lea93] George W. Leach. Review: C++ programming style. ;login: the USENIX Association newsletter, 18(6):45–46, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Leach:1994:PRO

[Lea94] Paul Leach. Panel: Radical operating systems structures for extensibility. In USENIX [USE94i], pages 195–

200. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/index.html.

Leach:1995:BRO

[Lea95a] George W. Leach. Book review: The Object Advantage: Business Process Reengineering with Object Technology. ;login: the USENIX Association newsletter, 20(5):46–47, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Leach:1995:RCC

[Lea95b] George W. Leach. Review:

Creating cool Web pages with HTML. ;login: the USENIX Association newsletter, 20(6):50– ??, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Leach:1996:HNT

[Lea96a] George W. Leach. Hands-on Netscape: a tutorial for Windows users. ;login: the USE-NIX Association newsletter, 21 (4):56–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Leach:1996:RCT

[Lea96b] George W. Leach. Review: Cultural treasures of the Internet. ;login: the USENIX Association newsletter, 21(5):48–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Leach:1996:RHJ

[Lea96c] George W. Leach. Review:

Hooked on Java. ;login: the USENIX Association newsletter, 21(3):55–??, June 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Leach:1996:RHD

[Lea96d] George W. Leach. Review:

HTML: The definitive guide. ;login: the USENIX Association newsletter, 21(5):47–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Leach:1996:RJP

[Lea96e] George W. Leach. Review:

The Java programming language. ;login: the USENIX Association newsletter, 21(5):46– ??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Leach:1996:REH

[Lea96f] George W. Leach. Review: To engineer is human: The role of failure in successful design. ;login: the USENIX Association newsletter, 21(1):47–48, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Leach:1997:BRJ

[Lea97] George W. Leach. Book review: Java Security: Hostile Applets, Holes, and Antidotes. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
mcgraw.html.

Leach:1998:BRL

[Lea98] George W. Leach. Book review: LDAP: Programming Directory-Enabled Applications with Lightweight Directory Access Protocol. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
ldap.html.

Lear:1999:MYN

[Lea99] Eliot Lear. Managing your network(s): Corporate mergers & acquisitions, or, you got your chocolate in my peanut butter,

1999. URL http://db.usenix.
org/publications/library/
proceedings/lisa99/invitedtalks.

html. Unpublished invited talk at LISA XIII.

Leber:1997:CBE

[Leb97] Jody Schivley Leber. Choosing a backup environment. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/schivley.html.

LeBlanc:1999:SWN

[LeB99]  David LeBlanc.  Securing

Windows NT services, 1999.

URL http://db.usenix.org/
publications/library/proceedings/
lisa-nt99/invitedtalks.html.

Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Leedom:1994:HSN

[Lee94] Casey Leedom. High speed networking symposium report. ;login: the USENIX Association newsletter, 19(5):8–16, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Lee:1998:RAC

[Lee98]  Jong-Hyeon Lee.  A re-

silient access control scheme for secure electronic transactions. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec98/lee.html.

LeFebvre:1992:RNA

[LeF92] William LeFebvre. Restricting network access to system daemons under SunOS. In USENIX [USE92m], pages 93–103. ISBN 1-880446-46-4. LCCN ????

LeFebvre:1994:KMT

[LeF94]  William LeFebvre.  Kernel

mucking in Top. In USE-NIX [USE94h], pages 47–55. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/leFebvre.html.

LeFebvre:1999:RRD

[LeF99] William LeFebvre. Rapid reverse DNS lookups for Web servers. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/lefebvre.html.

Lehenbauer:1998:NEW

[Leh98] Karl Lehenbauer. NeoWeb-Script: Enabling Webpages with active content using Tcl. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl98/lehenbauer.
                                        html.

Leighton:1998:NDU

[Lei98] Luke Kenneth Casson Leighton.

NT 3.5 / 4.0 domains for UNIX. In USENIX [USE98e], pages 85–88. ISBN 1-880446-96-0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/leighton.html.

Leir:1999:FTS

[Lei99] Rick Leir. FreeBSD: Tracking stable. ;login: the USE-NIX Association newsletter, 24 (5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/freebsd.html.

Leonard:1991:KAS

[Leo91] Anne G. Leonard. Keynote address: Structural revelation: Towards a mythology of the system. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Levy:1996:TNP

[Lev96] Jacob Levy. A Tk Netscape plugin. In USENIX [USE96e], pages 31–35. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/levy.html.

Levy:1998:TCA

[Lev98] Elias Levy. Trends in computer attacks. ;login: the USE-NIX Association newsletter, 23 (3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-5/

levy.html. Special issue on security.

Lewis:1996:FBC

[Lew96] Brian T. Lewis. An on-the-fly bytecode compiler for Tcl. In USENIX [USE96e], pages 103–114. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/lewis.html.

Long:1992:VRQ

[LFB92] Kevin Brook Long, Jerry Fowler, and Stan Barber. VNS retriever: Querying MEDLINE over the Internet (developing a network version of grateful med). In USENIX [USE92g], pages 81–92. ISBN 1-880446-44-

8. LCCN QA 76.76 O63 U83 1992.

Limoncelli:1998:PRN

[LFR^{+}98] Thomas A. Limoncelli, Robert Fulmer, Thomas Reingold, Alex Levine, and Ralph Loura. Providing reliable NT desktop services by avoiding NT Server. In USENIX [USE98e], pages 75–83. ISBN 1-880446-96-0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/limoncelli.html.

Langston:1990:CSU

[LH90] Peter S. Langston and Michael Hawley. Computing Systems: USENIX music. University of California Press, Berkeley, CA, USA, 1990. 1 sound disc.

Lea:1997:DPC

[LH97] Doug Lea and David Holmes.

Design patterns for concurrent programming: a Java example. Tutorial session at [USE97g].,

1997. URL http://dxsting.
cern.ch/sting/COOTS97.html.

Lichota:1996:VCP

[LHB96] Randall W. Lichota, Grace L.

Hammonds, and Stephen H. Brackin. Verifying cryptographic protocols for electronic commerce. In USE-NIX [USE96h], pages 53–

65. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/hammond.html.

Lepreau:1993:KSM

[LHF^{+}93] Jay Lepreau, Mike Hibler, Bryan Ford, Jeffrey Law, and Douglas Orr. In-kernel servers on Mach 3.0: Implementation and performance. In USE-NIX [USE93h], pages 39–55. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/lepreau.html.

Libes:1990:ECU

[Lib90a]  Don Libes.  expect: Cur-

ing those uncontrollable fits of interaction. In USE-NIX [USE90g], pages 183–192. LCCN QA76.8.U65 U81 1990.

                                        URL ftp://durer.cme.nist.
                                        gov/pub/expect.ps.Z. FTP -

durer.cme.nist.gov:/pub/expect.ps.Z; local - expect.ps.

Libes:1990:UEA

[Lib90b]  Don Libes.  Using expect

to automate system administration tasks. In USENIX [USE90f], pages 107–114. LCCN QA76.76.O63 L37 1990.

Libes:1991:ESC

[Lib91a] Don Libes. expect: Scripts for controlling interactive processes. In USENIX [USE91b], pages 99– 126.

Libes:1991:UEA

[Lib91b] Don Libes. Using expect to automate system administration tasks. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Libes:1992:RTC

[Lib92] Don Libes. Regression testing and conformance testing interactive programs. In USENIX [USE92g], pages 135–144. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Libes:1994:HPS

[Lib94] Don Libes. Handling passwords with security and reliability in background processes. In USENIX [USE94h], pages 57–64. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/libes.html.

Libes:1995:REE

[Lib95] Don Libes. Review: Exploring Expect. ;login: the USE-NIX Association newsletter, 20 (2):53–54, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Libes:1996:HAL

[Lib96a] Don Libes. How to avoid learning Expect -or- automating interactive programs. In USENIX [USE96j], pages 67–

71. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/libes.html.

Libes:1996:WCS

[Lib96b]  Don Libes.  Writing CGI

scripts in Tcl. In USE-NIX [USE96e], pages 189–201. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/libes.html.

Libes:1997:MTN

[Lib97a] Don Libes. Managing Tcl’s namespaces collaboratively. In USENIX [USE97d], pages 115–

123. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/libes_managing.html.

Libes:1997:WTE

[Lib97b] Don Libes. Writing a Tcl extension in only … 7 years. In USENIX [USE97d], pages 55–65. ISBN 1-880446-87-1.

LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/libes_writing.html.

Best paper award.

Liedtke:1992:FTM

[Lie92] Jochen Liedtke. Fast thread management and communication without continuations. In USENIX [USE92i], pages 213–

222. ISBN 1-880446-42-1. LCCN QAX 32.

Lieberman:1999:WSE

[Lie99]  Larry Lieberman.  Win-

dows 2000 support engineer for premier enterprise, 1999.

URL http://db.usenix.org/
publications/library/proceedings/
lisa-nt99/invitedtalks.html.

Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Limoncelli:1994:P

[Lim94] Tom Limoncelli. On professionalism. ;login: the USE-NIX Association newsletter, 19 (5):26–17, October 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Limoncelli:1997:TCU

[Lim97] Tom Limoncelli. Turning the corner: Upgrading yourself from “system clerk” to “system advocate”. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/

publications/library/proceedings/
lisa97/05.limoncelli.html.

Limoncelli:1998:PQ

[Lim98] Tom Limoncelli. Please quit.

;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Limoncelli:1999:SIN

[Lim99a] Thomas Limoncelli. Splitting IP networks: The 1999 update, 1999. URL http://www.

usenix.org/publications/library/
proceedings/neta99/invitedtalks.

html. Unpublished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Limoncelli:1999:DUR

[Lim99b] Thomas A. Limoncelli. Deconstructing user requests and the nine step model. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/limoncelli.
html.

Limoncelli:1999:PRN

[Lim99c] Thomas A. Limoncelli. Providing reliable NT desktop services by avoiding NT Server,

1999. URL http://db.usenix.
org/publications/library/
proceedings/lisa99/invitedtalks.

html. Unpublished invited talk at LISA XIII.

Limoncelli:1999:TYC

[Lim99d] Thomas A. Limoncelli. Tricks you can do if your firewall is a bridge. In USENIX [USE99g], page ?? ISBN 1-880446-36-

7. LCCN TK5105.5 .C664 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neta99/limoncelli.html.

Linn:1990:GSS

[Lin90a] John Linn. Generic security service application program interface. In USENIX [USE90i], pages 33–54. LCCN QA 76.9 A25 U55 1990.

Linton:1990:ED

[Lin90b] Mark A. Linton. The evolution of Dbx. In USE-NIX [USE90g], pages 211–

220. LCCN QA76.8.U65 U81 1990. URL ftp:/

                                        /interviews.stanford.edu/
                                        pub/papers/dbx.ps.Z. FTP -

interviews.stanford.edu:/pub/papers/dbx.ps.Z; local - dbx.ps.

Linton:1992:ECL

[Lin92] Mark Linton. Encapsulating a C++ library. In USENIX [USE92k], pages 57–66. ISBN 1-880446-45-6. LCCN ????

LoVerso:1993:SPF

[LIN^{+}93] Sue J. LoVerso, Marshall Isman, Andy Nanopoulos, William Nesheim, Eric L. Rowe, and Richard Wheeler. sfs: a parallel file system for the CM-

5. In USENIX [USE93g], pages 291–305. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/loverso.html.

Lindblad:1994:UTC

[Lin94] Christopher J. Lindblad. Using Tcl to control a computerparticipative multimedia programming environment. In USENIX [USE94k], pages 57–

69. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/lindblad.html.

Lipson:1995:CMT

[Lip95] Sam Lipson. Capital markets trading floors, current practice. In USENIX [USE95h], pages 35–46 (or 35–45??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/lipson.html.

Littman:1992:AT

[Lit92] Jay Littman. Applying threads.

In USENIX [USE92j], pages 209–222.

Lear:1996:RTM

[LKCT96] Eliot Lear, Jennifer Katinsky, Jeff Coffin, and Diane Tharp. Renumbering: Threat or menace? In USENIX [USE96j], pages 91–

96. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/ariana.html.

Li:1994:QAD

[LKHA94] Kester Li, Roger Kumpf, Paul Horton, and Thomas Anderson. A quantitative analysis of disk drive power management in portable computers. In USENIX [USE94l], pages 279–291. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/anderson.html.

Lakshman:1993:DIM

[LKY93] K. Lakshman, G. Kuthethoor, and R. Yavatkar. Design and implementation of a multimedia protocol suite in a BSD UNIX kernel. In USE-NIX [USE93g], pages 129–146. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/.

Lockard:1998:SSP

[LL98] John Lockard and Jason Larke. Synctree for single point installation, upgrades, and OS patches. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/lockard.html.

Laffra:1994:HVD

[LM94] Chris Laffra and Ashok Malhotra. HotWire — A visual debugger for C++. In USE-NIX [USE94e], pages 109–122. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/laffra.html.

LoVerso:1997:CDW

[LM97] John Robert LoVerso and Murray S. Mazer. Caubweb: Detaching the Web with Tcl. In USENIX [USE97d], pages 19–29. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/loverso.html.

Leijen:1999:DSE

[LM99a] Daan Leijen and Erik Meijer. Domain-specific embedded compilers. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/leijen.html.

Li:1999:CSL

[LM99b] Du Li and Richard R. Muntz. A collaboration specification language. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999. URL http://www.

usenix.org/publications/library/
proceedings/dsl99/li.html.

Long:1994:SRD

[LMC94] Darrell D. E. Long, Bruce R.

Montague, and Luis-Felipe Cabrera. Swift/RAID: a distributed RAID system. In USE-NIX [USE94c], pages 333–359.

Lenkov:1993:TIC

[LMU93] Dmitry Lenkov, Michey Mehta, and Shankar Unni. Type identification in C++. In Waldo [Wal93b], pages 197–213. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Lange:1995:PEP

[LN95] Danny B. Lange and Yuichi Nakamura. Program explorer: a program visualizer for C++. In USENIX [USE95k], pages 39–54. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/lange.html.

Leinwand:1990:TNM

[LO90] Allan Leinwand and Jeff Okamoto. Two network management tools(how many packets would a packet router route if a packet router could route packets). In USENIX [USE90h], pages 195–

206. LCCN QA76.8.U65 U82

1990.

Levy:1995:STT

[LO95] Jacob Y. Levy and John K. Ousterhout. A safe Tcl toolkit for electronic meeting places. In USENIX [USE95g], pages 133– 135 (or 133–136??). ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1). URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/ec95/levy.html.

Lodin:1993:CSB

[Lod93] Steven W. Lodin. The corporate software bank. In USENIX [USE93f], pages 33–

42. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/lodin.html.

Loepere:1996:CSM

[Loe96]  Keith Loepere.  Composing

special memory allocators in C++. In USENIX [USE96g], pages 15–24. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/loepere.html.

Lidl:1994:DFM

[LOM94] Kurt J. Lidl, Josh Osborne, and Joe Malcolm. Drinking from the firehose: Multicast USENET news. In USE-NIX [USE94l], pages 33–45. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/lidl.html.

Lonczewski:1995:PUT

[Lon95]  Frank Lonczewski.  PLUG-

IN: Using Tcl/Tk for planbased user guidance. In USE-NIX [USE95j], pages 141–

144. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/longczews.html.

Lord:1995:AGI

[Lor95] Thomas Lord. An anatomy of Guile: The interface to Tcl/ Tk. In USENIX [USE95j], pages 95–114. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/lord.html.

Linton:1994:ITI

[LP94] Mark Linton and Douglas Pan. Interface translation and implementation filtering. In USE-NIX [USE94e], pages 227–236. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        c++94/linton.html.

Liedtke:1998:HPC

[LPJI98] Jochen Liedtke, Vsevolod Panteleenko, Trent Jaeger, and Nayeem Islam. High-performance caching with the Lava hitserver. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix98/liedtke.html.

LoVerso:1991:OUF

[LPLF91] Susan LoVerso, Noemi Paciorek, Alan Langerman, and George Feinberg. The OSF/1 UNIX filesystem (UFS). In USENIX [USE91i], pages 207–218. LCCN QA 76.76 O63 U84 1992.

Lee:1999:AID

[LPS99] Wenke Lee, Christopher T.

Park, and Salvatore J. Stolfo. Automated intrusion detection using NFR: Methods and experiences. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/lee.html.

Lindemann:1991:AIM

[LPV^{+}91] Eric Lindemann, Miller Puckette, Eric Viara, Maurizio De Cecco, François Dechelle, and Bennett Smith. The architecture of the IRCAM musical workstation. In USE-NIX [USE91g], pages 209–220. LCCN QA 76.76 O63 U83 1991.

Ladd:1994:TAL

[LR94] David A. Ladd and J. Christopher Ramming. Two application languages in software production. In USENIX [USE94k], pages 169–177. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/ladd.html.

Lowe:1998:IDE

[LR98a] James S. Lowe and Luigi Rizzo. ISC DHCP EDistribution. In USENIX [USE98d], page ?? ISBN ???? LCCN

???? URL http://www.isc.
org/dhcp-freenix.html.

Lowe:1998:MDS

[LR98b] James S. Lowe and Luigi Rizzo. Multimedia driver support in the FreeBSD operating system. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/lowe.ps.

Limoncelli:1997:CNL

[LRNL97] Tom Limoncelli, Tom Reingold, Ravi Narayan, and Ralph Loura. Creating a network for Lucent Bell Labs Research South. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/16.limoncelli.html.

Litzkow:1992:SCP

[LS92] Michael Litzkow and Marvin Solomon. Supporting checkpointing and process migration outside the UNIX kernel. In USENIX [USE92j], pages 283– 290.

Li:1993:NDP

[LS93] Hui Li and Kenneth C. Sevcik. NUMACROS: Data parallel programming on NUMA multiprocessors. In USE-NIX [USE93j], pages 247–

263. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/li.html.

Lam:1997:JTI

[LS97] Ioi K. Lam and Brian C.

Smith. Jacl: a Tcl implementation in Java. In USENIX [USE97d], pages 31–

36. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.cs.cornell.
edu/home/ioi/Jacl;   http:
//www.usenix.org/publications/
library/proceedings/tcl97/
lam.html.

Lee:1998:DMA

[LS98a] Wenke Lee and Salvatore J.

Stolfo. Data mining approaches for intrusion detection. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/lee.html.

Little:1998:JTI

[LS98b] M. C. Little and S. K. Shrivastava. Java transactions for the Internet. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/little.html.

Lee:1999:LIE

[LS99] Daniel D. Lee and H. Sebastian Seung. Learning in intelligent embedded systems. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
es99/lee.html.

Liu:1995:TDN

[LSR95] Peter T. Liu, Brian Smith, and Lawrence Rowe. Tcl-DP name server. In USE-NIX [USE95j], pages 1–14 (or 1–13??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/liu.html.

Leong:1991:UPE

[LT91] Philip Leong and Chris Tham. UNIX password encryption considered unsecure. In USENIX [USE91i], pages 269–280. LCCN QA 76.76 O63 U84 1992.

Leffler:1997:MLA

[LT97] Sam Leffler and Melange Tortuba. Mailing list archive tools. In USENIX [USE97e], pages 207–217. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ana97/leffler.html.

Lu:1997:IOD

[Lu97] Paul Lu. Implementing optimized distributed data sharing using scoped behaviour and a class library. In USE-NIX [USE97g], pages 145–

158. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots97/lu.html. See also

[LH97, Rie97, WW97].

Lenkov:1991:TIC

[LUM91] Dmitry Lenkov, Shankar Unni, and Michey Mehta. Type identification in C++. In USE-NIX [USE91k], pages 103–118. LCCN QA76.73.C153 U27 1991.

Lunt:1990:EK

[Lun90] Steven Lunt. Experiences with Kerberos. In USENIX [USE90i], pages 113–120. LCCN QA 76.9 A25 U55 1990.

Lutz:1999:UP

[Lut99] Mark Lutz. Using Python.

;login: the USENIX Association newsletter, 24(1s): ??, January 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-1/

python.html. Special issue on tools.

Liang:1999:CPS

[LV99] Sheng Liang and Deepa Viswanathan.

Comprehensive profiling support in the Java Virtual Machine. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/liang.html.

Lucks:1999:RKE

[LW99] Stefan Lucks and Rüdiger Weis. Remotely keyed encryption using non-encrypting Smart Cards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/lucks.html.

Lindblad:1994:VAI

[LWS^{+}94] Christopher J. Lindblad, David J. Wetherall, William F. Stasior, Joel F. Adam, Henry H. Houh, Mike Ismert, David R. Bacher, Brent M. Phillips, and David L. Tennenhouse. View-Station applications: Intelligent video processing over a broadband local area network. In USENIX [USE94f], pages 103–117. ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
hispeed/lindblad.html.

Wong:1999:EWN

[lWS99] Alexander Ya li Wong and Margo I. Seltzer. Evaluating Windows NT Terminal Server performance. In USE-NIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/wong.html.

Lycklama:1994:BWD

[Lyc94] Ed Lycklama. Better widget design: a practitioner’s approach. In USENIX [USE94g], pages 121–132. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/lycklama.html.

Lycklama:1996:OSP

[Lyc96] Heinz Lycklama. Open systems, POSIX, and Windows NT

— another point of view. ;login: the USENIX Association newsletter, 21(1):40–44, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Lee:1997:BTI

[LZ97] Han Bok Lee and Benjamin G.

Zorn. BIT: a tool for instrumenting Java bytecodes. In USENIX [USE97h], pages 73–

82. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/lee.html.

McNamee:1990:EME

[MA90] Dylan McNamee and Katherine Armstrong. Extending the Mach external pager interface to accommodate user-level page replacement policies. In USENIX [USE90e], pages 17–30.

Macklem:1991:LLT

[Mac91] Rick Macklem. Lessons learned tuning the 4.3BSD reno implementation of the NFS protocol. In USENIX [USE91i], pages 53–

64. LCCN QA 76.76 O63 U84 1992.

Macklem:1994:QNS

[Mac94] Rick Macklem. Not quite NFS, soft cache consistency for NFS. In USENIX [USE94l], pages 261–278. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/macklem.html.

Madsen:1992:GCD

[Mad92] Jørgen Bo Madsen. The greatest cracker-case in Denmark: The detecting, tracing, and arresting of two international crackers. In USENIX [USE92m], pages 17–

40. ISBN 1-880446-46-4. LCCN ????

Maffeis:1993:EMD

[Maf93] Silvano Maffeis. Electra — making distributed programs object-oriented. In USE-NIX [USE93j], pages 143–

156. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/maffeis.html.

Maffeis:1995:AGC

[Maf95]  Silvano Maffeis.  Adding

group communication and faulttolerance to CORBA. In USENIX [USE95k], pages 135–

146. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/maffeis.html.

Maffeis:1996:OGD

[Maf96]  Silvano Maffeis.  The ob-

ject group design pattern. In USENIX [USE96g], pages 155–

163. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/maffeis.html.

Mahler:1991:OTU

[Mah91] Alex Mahler. Organizing tools in a uniform environment framework. In USENIX [USE91i], pages 231–242. LCCN QA 76.76 O63 U84 1992.

Mahmoud:1996:TCS

[Mah96] Qusay H. Mahmoud. Tcl, CGI, and security. ;login: the USE-NIX Association newsletter, 21 (3):42–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Miller:1998:UCD

[MAH98] Ethan L. Miller, Kennedy Akala, and Jeffrey K. Hollinsworth. Using content-derived names for package management in Tcl. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/akala.html.

McDonald:1995:OTP

[MAM95] Daniel L. McDonald, Randall J.

Atkinson, and Craig Metz. One-Time Passwords in Everything (OPIE): Experiences with building and using stronger authentication. In USENIX [USE95f], pages 177–186. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
mcdonald.html.

Manandhar:1991:ASY

[Man91]  Sanjay Manandhar.  Activ-

ity server: You can run but you can’t hide. In USE-NIX [USE91g], pages 299–312. LCCN QA 76.76 O63 U83 1991.

Manber:1994:FSF

[Man94] Udi Manber. Finding similar files in a large file system. In USENIX [USE94l], pages 1–10. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/manber.1.html.

Manasse:1995:MPE

[Man95]  Mark S. Manasse.  The

Millicent protocols for electronic commerce. In USENIX [USE95g], pages 117–123 (or 117–124??). ISBN 1-880446-74-X. LCCN HF5548.33. U84

1995(1). URL http://www.
usenix.org/publications/library/
proceedings/ec95/manasse.
html.

Manber:1997:CPW

[Man97] Udi Manber. Creating a personal Web notebook. In USE-NIX [USE97h], pages 183–191. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/manber_creating.html.

Mann:1998:JCS

[Man98] Steve Mann. Joint closing session: Reconfigured self as basis for humanistic intelligence, 1998. URL http://www.

usenix.org/publications/library/
proceedings/usenix98/invited_
talks/mann.ps. Unpublished

invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Maples:1996:P

[Map96] Greg Maples. Pentagon. ;login: the USENIX Association newsletter, 21(6):48–??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Markey:1991:EHS

[Mar91a] Brian D. Markey. Emerging hypermedia standards — hypermedia marketplace prepares for HyTime and MHEG. In USENIX [USE91g], pages 59– 74. LCCN QA 76.76 O63 U83 1991.

Martin:1991:SII

[Mar91b] Bruce Martin. The separation of interface and implementation in C++. In USENIX [USE91k], pages 51–64. LCCN QA76.73.C153 U27 1991.

Martin:1993:SII

[Mar93] Bruce Martin. The separation of interface and implementation in C++. In Waldo [Wal93b], pages 249–265. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Martin:1996:ICC

[Mar96a] Bruce Martin. Introduction to CORBA and CORBA services. In USENIX [USE96g], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/

publications/library/proceedings/
coots96/.

Martin:1996:AWC

[Mar96b]  Ken Martin.  Automated

wrapping of a C++ class library into Tcl. In USE-NIX [USE96e], pages 141–148. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/martin.html.

Martin:1996:DPD

[Mar96c] Robert Martin. Design patterns for dealing with dual inheritance hierarchies in C++. In USENIX [USE96g], pages 143–153. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/martin.html.

Markey:1997:SAV

[Mar97] Bruce Markey. A system administrator’s view of LDAP. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
markey.html.

Martin:1998:IAT

[Mar98]  David M. Martin.  Inter-

net anonymizing techniques. ;login: the USENIX Association newsletter, 23(3): ??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-5/

martin.html. Special issue on security.

Marchette:1999:SMP

[Mar99] David Marchette. A statistical method for profiling network traffic. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/marchette.html.

Masui:1991:UIC

[Mas91] Toshiyuki Masui. User interface construction based on parallel and sequential execution specification. In USENIX [USE91i], pages 117–126. LCCN QA 76.76 O63 U84 1992.

Mata:1997:TOR

[Mat97] Cristian S. Mata. TxRx: An ONC RPC interpreter. In USENIX [USE97d], pages 179–

180. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_mata.html.

Matheson:1998:UFC

[Mat98]  Branson Matheson.  Using

FreeBSD for a console server. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/matheson.
ps.

Matus:1999:SLF

[Mat99] Juan Matus. Setting up a Linux firewall. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-11/
                                        features/linux.html. Special

issue on security.

Mauck:1998:SPU

[Mau98] Aaron Mauck. Stop the presses: USENIX 1998. Linux Journal, 53:??, September 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Maybee:1990:PNO

[May90] Paul Maybee. pdb: a network oriented symbolic debugger. In USENIX [USE90h], pages 41– 52. LCCN QA76.8.U65 U82 1990.

Maybee:1992:NNE

[May92] Paul Maybee. NED: The network extensible debugger. In USENIX [USE92g], pages 145– 156. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Mayer:1995:TTL

[May95] Ken Mayer. Top 10 tools for lonely sys admins. ;login: the USENIX Association newsletter, 20(5):11–12, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Mazumdar:1996:IDM

[Maz96]  Subrata Mazumdar.  Inter-

domain management: CORBA, OSI, SNMP. In USENIX [USE96g], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

usenix.org/publications/library/
proceedings/coots96/.

Mafla:1991:IPC

[MB91] Enrique Mafla and Bharat Bhargava. Implementation and performance of a communication facility for the RAID distributed transaction processing system. In USENIX [USE91j], pages 69–

86. ISBN ???? LCCN ????

Maetzel:1996:FPA

[MB96] Kai-Uwe Maetzel and Walter Bischofberger. The Any framework:A pragmatic approach to flexibility. In USE-NIX [USE96g], pages 179–

190. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/maetzel.html.

Manber:1997:SB

[MB97] Udi Manber and Peter A. Bigot.

The search broker. In USE-NIX [USE97h], pages 231–239. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/manber_search.html.

Ma:1998:CCB

[MB98] Chaoying Ma and Jean Bacon. COBEA: a CORBAbased event architecture. In USENIX [USE98c], page ??

ISBN ???? LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots98/ma.html.

Minnich:1994:GSM

[MBH94] Ron Minnich, Dan Burns, and Frank Hady. A 1.2 Gbit/sec, 1 microsecond latency ATM interface. In USENIX [USE94f], pages 73–90. ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        hispeed/index.html.

McIlroy:1993:ERS

[MBM93] Peter M. McIlroy, Keith Bostic, and M. Douglas McIlroy. Engineering radix sort. In USENIX [USE93d], pages 5–27.

Maffeis:1995:GMT

[MBM95] Silvano Maffeis, Walter Bischofberger, and Kai-Uwe Maetzel. A generic multicast transport service to support disconnected operation. In USE-NIX [USE95i], pages 79–90 (or 79–89??). ISBN 1-880446-69-

3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/maffeis.html.

Mogul:1995:PIM

[MBMS95] Jeffrey C. Mogul, Joel F. Bartlett, Robert N. Mayo, and Amitabh Srivastava. Performance implications of multiple pointer sizes. In USE-NIX [USE95e], pages 187–200. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/mogul.html.

Macaire:1999:PND

[MC99] Alain Macaire and David Carlier. A personal naming and directory service for mobile Internet users. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/macaire.html.

McCluskey:1997:UCBa

[McC97a] Glen McCluskey. Using C++ as a better C. ;login: the USE-NIX Association newsletter, 22 (4):30–32, August 1997. CO-DEN LOGNEM. ISSN 1044-6397.

McCluskey:1997:UCBb

[McC97b] Glen McCluskey. Using C++ as a better C. ;login: the USE-NIX Association newsletter, 22 (5):49–53, June 1997. CODEN LOGNEM. ISSN 1044-6397.

McCluskey:1997:UCBc

[McC97c] Glen McCluskey. Using C++ as a better C. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
cpluplus.html.

McCluskey:1997:UCBd

[McC97d] Glen McCluskey. Using C++ as a better C. ;login: the USE-NIX Association newsletter, 22 (7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
cplusplus.html.

McCluskey:1998:JPd

[McC98a] Glen McCluskey. Java I/O performance. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
javaio.html.

McCluskey:1998:JPa

[McC98b] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
jperf.html.

McCluskey:1998:JPb

[McC98c] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
java_perf.html.

McCluskey:1998:JPc

[McC98d] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
java_performance.html.

McCluskey:1998:UCBa

[McC98e] Glen McCluskey. Using C++ as a better C. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
c++.html.

McCluskey:1998:UCBb

[McC98f] Glen McCluskey. Using C++ as a better C. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
c%2B%2B.html.

McCriskard:1998:EIA

[McC98g] D. Scott McCriskard. Enhancing interfaces with animated widgets. In USENIX [USE98f], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/full_papers/
declarke_poster/declarke_
html/declarke.html.

McCluskey:1999:JPa

[McC99a] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

                                        publications/login/1999-4/
                                        java_perf.html.

McCluskey:1999:JPb

[McC99b] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-8/
                                        features/performance.html.

McCluskey:1999:JPc

[McC99c] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-12/
                                        features/performance.html.

McCluskey:1999:JPM

[McC99d] Glen McCluskey. Java performance: Memory fragmentation. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        java_perf.html.

McDonald:1992:SRW

[McD92] Bryan McDonald. SAGE report on: World Conference on System Administration & Security. ;login: the USENIX Association newsletter, 17(5):19–??, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

McDonald:1993:SEa

[McD93a] Bryan McDonald. From the SAGE editor. ;login: the USE-NIX Association newsletter, 18 (4):20–??, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

McDonald:1993:SEb

[McD93b] Bryan McDonald. From the SAGE editor. ;login: the USE-NIX Association newsletter, 18 (5):23–??, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

McDonald:1994:E

[McD94] Bryan McDonald. From the editor. ;login: the USENIX Association newsletter, 19(4):17– ??, August 1994. CODEN LOGNEM. ISSN 1044-6397.

McDonald:1996:B

[McD96] Bryan McDonald. BayLISA 1996. ;login: the USENIX Association newsletter, 21(6):9– ??, December 1996. CODEN LOGNEM. ISSN 1044-6397.

McDonald:1999:PCC

[McD99]  Bryan McDonald.  Prepar-

ing to consider certification. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
preparing.html.

McKay:1993:USS

[McK93a] Calum D. McKay. UNIX security symposium report. ;login: the USENIX Association newsletter, 18(6):4–8, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

McKusick:1993:CG

[McK93b] Kirk McKusick. A changing of the guard. ;login: the USENIX Association newsletter, 18(1): 5–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

McKusick:1995:VFI

[McK95] Marshall Kirk McKusick. The virtual filesystem interface in 4.4BSD. In USENIX [USE95d], pages 3–25.

McKusick:1998:CSC

[McK98] Kirk McKusick. Complete set of CSRG BSD releases available. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
mailing.html.

McLennan:1995:NIT

[McL95] Michael J. McLennan. The new [incr Tcl]: Objects, mega-widgets, namespaces and more. In USENIX [USE95j], pages 151–160 (or 151–159??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995. URL

http://www.usenix.org/publications/
library/proceedings/tcl95/
mclennan.html.

Mecklenburg:1994:DDP

[MCLY94] Robert Mecklenburg, Charles Clark, Gary Lindstrom, and Benny Yih. A dossier driven persistent objects facility. In USENIX [USE94e], pages 265–

281. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/mecklenburg.html.

McNutt:1993:RBS

[McN93]  Dinah McNutt.  Role-based

system administration or who, what, where, and how. In USENIX [USE93f], pages 107–

112. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/mcnutt.html.

McRae:1993:HPK

[McR93]  Andrew McRae.  Hardware

profiling of kernels. In USE-NIX [USE93l], pages 375–386. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

McRobert:1995:TCL

[McR95] Stuart McRobert. From twisting country lanes to multilane Ethernet superhighways. In USENIX [USE95h], pages 221–228. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/mcrobert.html.

Modzelewski:1997:IGO

[MCU97] Brent E. Modzelewski, David Cyganski, and Marian V. Underwood. Interactive-group object-replication fault tolerance for CORBA. In USE-NIX [USE97g], pages 241–

244. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots97/panel.modzelewski.

html. See also [LH97, Rie97,

WW97].

McKenney:1992:EDI

[MD92] Paul E. McKenney and Ken F. Dove. Efficient demultiplexing of incoming TCP packets. In USENIX [USE92b], pages 141– 157.

Montenegro:1995:SIN

[MD95] Gabriel Montenegro and Steve Drach. System isolation and network fast fail capability in Solaris. In USENIX [USE95i], pages 67–78. ISBN 1-880446-69-

3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/montenegro.html.

Markatos:1996:IRR

[MD96] Evangelos P. Markatos and George Dramitinos. Implementation of a reliable remote memory pager. In USENIX [USE96k], pages 177–189 (or 177–190??). ISBN 1-880446-76-6. LCCN QA 76.76 O63

                                        U88 1996. URL http://www.
                                        usenix.org/publications/library/

proceedings/sd96/markatos.
html.

McLaughlin:1998:DPS

[MD98] Donald McLaughlin and Partha Dasgupta. Distributed preemptive scheduling on Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/full_papers/poster_
mclaughlin/mclaughlin.pdf;
 http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/full_papers/poster_mclaughlin/
mclaughlin_html/mclaughlin.
html.

Mohindra:1995:DIO

[MDC95] Ajay Mohindra, Murthy Devarakonda, and George Copeland. Dynamic insertion of object services. In USE-NIX [USE95k], pages 13–

20. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/mohindra.html.

Mowry:1996:ACI

[MDK96] Todd C. Mowry, Angela K.

Demke, and Orran Krieger. Automatic compiler-inserted I/O prefetching for out-of-core applications. In USENIX [USE96f], pages 3–17. ISBN 1-880446-82-0. LCCN QA76.76.O63

S96 1996. URL http://www.
usenix.org/publications/library/
proceedings/osdi96/mowry.

html. Award paper.

Miller:1999:SSC

[MdR99] Todd C. Miller and Theo

                                        de Raadt. strlcpy and strlcat

— consistent, safe, string copy and concatenation. In USE-NIX [USE99m], page ?? ISBN 1-880446-33-2. LCCN ????

                                        URL http://www.openbsd.
                                        org/papers/strlcpy-paper.
                                        ps.

Messerges:1999:IPA

[MDS99] Thomas S. Messerges, Ezzy A. Dabbish, and Robert H. Sloan. Investigations of power analysis attacks on Smartcards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/smartcard99/
                                        messerges.html.

May:1999:RDC

[ME99] Kenneth O. May and P. Eng. Radio dial-in connectivity to NT networks. In USE-NIX [USE99b], pages 51–57. ISBN 1-880446-30-8. LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa-nt99/may.
                                        html.

Meeks:1997:BNI

[Mee97] W. Scott Meeks. Building a notification infrastructure using Tcl, Zephyr and Linda. In USENIX [USE97d], pages 161–162. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/

publications/library/proceedings/
tcl97/poster_meeks.html.

Meek:1998:AAM

[Mee98] Jon Meek. Apache authentication module. ;login: the USE-NIX Association newsletter, 23 (3):??, May 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-5/

meek.html. Special issue on security.

Mellis:1996:FOW

[Mel96a] Mark K. Mellis. Fifty-one weeks.

;login: the USENIX Association newsletter, 21(4):38–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Mellis:1996:MYU

[Mel96b] Mark K. Mellis. Managing your UPS in an emergency, or (don’t) wait until dark. ;login: the USE-NIX Association newsletter, 21 (2):25–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Mellis:1997:BP

[Mel97a] Mark Mellis. The big picture. ;login: the USENIX Association newsletter, 22(2):12–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Mellis:1997:CN

[Mel97b] Mark K. Mellis. Computational nostalgia. ;login: the USE-NIX Association newsletter, 22 (5):16–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Mellis:1997:SD

[Mel97c] Mark K. Mellis. Surprises in the DMZ. ;login: the USE-NIX Association newsletter, 22 (1):23–??, February 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Mellis:1998:RE

[Mel98] Mark K. Mellis. Riding the escalator. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
escalator.html.

Mellis:1999:RWI

[Mel99a] Mark K. Mellis. Real world intrusion detection, 1999. URL

http://db.usenix.org/publications/
library/proceedings/lisa99/
invitedtalks.html. Unpub-

lished invited talk at LISA XIII.

Melski:1999:BBR

[Mel99b]  Eric Melski.  Burt: The

Backup and Recovery Tool. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/melski.
html.

MPC:1993:PDF

[Mem93] Members of the Program Committee. Panel discussion: The future of experimental distributed systems research. In USENIX [USE93j], page ?? ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

Menter:1993:MMC

[Men93] E. Scott Menter. Managing the mission critical environment. In USENIX [USE93f], pages 81–86. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/menter.html.

Mentor:1994:GBF

[Men94] E. Scott Mentor. On the glory of being first. ;login: the USENIX Association newsletter, 19(3):9– 10, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Menezes:1998:ECR

[Men98]  Alfred Menezes.  Elliptic

curves — ready for prime time. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/invited_talks/menezes/
ecdsa.ps;   http://www.usenix.
org/publications/library/
proceedings/sec98/invited_
talks/menezes/slides-4up.
ps.

Metz:1995:SSC

[Met95] Melissa Metz. Some SANS conference highlights. ;login: the USENIX Association newsletter, 20(4):14–19, August 1995. CODEN LOGNEM. ISSN 1044-

Meek:1998:WAN

[MET98a] Jon T. Meek, Edwin S. Eichert, and Kim Takayama. Wide area network ecology. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/lisa98/
                                        meek.html.

Metz:1998:NI

[Met98b] Craig Metz. NRL IPv6+IPsec. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix98/freenix/metz.ps.

Meyers:1991:IPM

[Mey91] Randall Meyers. The interaction of pointers to members and virtual base classes in C++. In USENIX [USE91k], pages 1–12. LCCN QA76.73.C153 U27 1991.

Mosedale:1995:AVH

[MFM95] Dan Mosedale, William Foss, and Rob McCool. Administering very high volume Internet services. In USE-NIX [USE95h], pages 95–102. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa95/mosedale.html.

Minnich:1998:FCC

[MGM^{+}98] Ron Minnich, Maya Gokhale, Aaron Marks, Jim Kaba, and John Degood. Freeware for cluster computing. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/minnich.
ps.

Milojicic:1993:ELD

[MGZ93] Dejan S. Milojicic, Peter Giese, and Wolfgang Zint. Experiences with load distribution on top of the Mach microkernel. In USENIX [USE93j], pages 19–36. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/milojicic.html.

Muntz:1992:MLC

[MH92] D. Muntz and P. Honeyman. Multi-level caching in distributed file systems -or- your cache ain’t nuthin’ but trash. In USENIX [USE92j], pages 305– 314.

Hall:1996:UC

[mH96] John “maddog” Hall. USELINUX conference. ;login: the USE-NIX Association newsletter, 21 (3):26–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Moskowitz:1999:BS

[MH99] Adam Moskowitz and Gregory H. Hamm. Budgeting for sysadmins, 1999. URL http://

db.usenix.org/publications/
library/proceedings/lisa99/
practicum.html. Unpublished

practicum talk at LISA XIII.

Manning:1993:UWS

[MI93] Craig Manning and Tim Irvin. Upgrading 150 workstations in a single sitting. In USE-NIX [USE93f], pages 17–

27. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/.

Milazzo:1991:SVU

[Mil91]  Paul G. Milazzo.  Shared

video under UNIX. In USE-NIX [USE91g], pages 369–384. LCCN QA 76.76 O63 U83 1991.

Miller:1993:MEC

[Mil93]  John E. Miller.  Managing

an ever-changing user base. In USENIX [USE93f], pages 1–5. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/.

Miller:1995:SSB

[Mil95] Hal Miller. Summary of SAGE board activities, 1995. ;login: the USENIX Association newsletter, 20(6):27–28, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Miller:1996:R

[Mil96a] Hal Miller. Reflection. ;login: the USENIX Association newsletter, 21(2):27–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Miller:1996:CRU

[Mil96b] James R. Miller. Class relationships and user extensibility in solid geometric modeling. In USENIX [USE96g], pages 209–218. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/miller.html.

Miller:1997:SP

[Mil97a] Hal Miller. From the SAGE President. ;login: the USE-NIX Association newsletter, 22 (5):11–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Miller:1997:ISC

[Mil97b] Hal Miller. An introduction to the SAGE code of ethics. ;login: the USENIX Association newsletter, 22(1):9–10, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Miller:1997:RDA

[Mil97c] Hal Miller. A response to Dan Appelman. ;login: the USE-NIX Association newsletter, 22 (4):24–??, August 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Miller:1997:SBE

[Mil97d] Hal Miller. SAGE board election changes. ;login: the USE-NIX Association newsletter, 22 (4):13–14, August 1997. CO-DEN LOGNEM. ISSN 1044-

Miller:1997:SPL

[Mil97e] Hal Miller. SAGE President’s letter. ;login: the USENIX Association newsletter, 22(2):9–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Miller:1997:YC

[Mil97f] Hal Miller. A year of change. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Miller:1998:CE

[Mil98a] Hal Miller. The case for expansion. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Miller:1998:EYR

[Mil98b] Hal Miller. End of year report. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

Miller:1998:FC

[Mil98c] Hal Miller. Flattening the curve. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Miller:1998:MSR

[Mil98d] Hal Miller. Member survey results. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-4/
                                        survey.html.

Miller:1998:PLa

[Mil98e] Hal Miller. President’s letter. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
worth.html.

Miller:1998:PLb

[Mil98f] Hal Miller. President’s letter. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
editorial.html.

Miller:1998:TY

[Mil98g] Hal Miller. Thanks to you. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Miller:1998:FUP

[Mil98h] Richard Miller. The first Unix port, 1998. URL http://www.

usenix.org/publications/library/
proceedings/usenix98/invited_
talks/miller.ps. Unpublished

invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Milburn:1999:EA

[Mil99a]  Shane B. Milburn.  Email

archiving. ;login: the USE-NIX Association newsletter, 24 (5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/archiving.html.

Miller:1999:AP

[Mil99b] Hal Miller. Approaching a petabyte, 1999. URL http://

db.usenix.org/publications/
library/proceedings/lisa99/
invitedtalks/miller.pdf. Un-

published invited talk at LISA XIII.

Miller:1999:BP

[Mil99c] Hal Miller. Broken paradigm. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
paradigm.html.

Miller:1999:EOG

[Mil99d] Hal Miller. Extending our goal set. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
goal_set.html.

Miller:1999:HST

[Mil99e] Hal Miller. How-to: Set up a time server. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/sage/sysadmins/
howto/4.time.html.

Miller:1999:MA

[Mil99f]  Hal Miller.  Moving along.

;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
moving.html.

Miller:1999:NP

[Mil99g] Hal Miller. Note from the President. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
president.html.

Miller:1999:SA

[Mil99h]  Hal Miller.  SAGE AU

’99. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
sageau.html.

Miller:1999:TS

[Mil99i] Hal Miller. The third SAGE!

;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
thirdsage.html.

Miller:1999:TJS

[Mil99j]  Hal Miller.  Truth, jus-

tice, and the sysadmin way.

;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
sysadminway.html.

Mills:1999:ANS

[Mil99k] Kevin L. Mills. AirJava: Networking for smart spaces. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999. URL

http://www.usenix.org/publications/
library/proceedings/es99/
mills.html.

Minnich:1993:AFC

[Min93a] Ronald G. Minnich. The AutoCacher: a file cache which operates at the NFS level. In USENIX [USE93l], pages 77–83. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Minnich:1993:MNM

[Min93b] Ronald G. Minnich. Mether-NFS: a modified NFS which supports virtual shared memory. In USENIX [USE93j], pages 89–107. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/minnich.html.

Minear:1995:PPC

[Min95] Spencer E. Minear. Providing policy control over object operations in a Mach-Based system. In USENIX [USE95f], pages 141–156 (or 141–155??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
minear.html.

Mitchell:1991:MRC

[Mit91] David W. Mitchell. Mach resource control in OSF/1. In USENIX [USE91h], pages 123–

130. LCCN QAX 27.

Mitchell:1993:MSR

[Mit93] Dave Mitchell. Mach symposium report. ;login: the USE-NIX Association newsletter, 18 (3):3–7, May/June 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Mitchell:1995:RFS

[Mit95] Dave Mitchell. Report on the first symposium on operating systems design and implementation. ;login: the USENIX Association newsletter, 20(1):11– 16, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Mitchell:1996:JBF

[Mit96]  Jim Mitchell.  Java OS:

Back to the future. In USENIX [USE96f], pages 1– ?? ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/.

Mitchell:1997:JBF

[Mit97]  Jim Mitchell.  Java OS:

Back to the future, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/invited_talks/mitchell/
abstract.html. Unpublished

invited talk at the USENIX 2nd Symposium on OS Design and Implementation (OSDI ’96), October 28–31, 1996, Seattle, Washington.

Moran:1992:CME

[MJ92] William Moran, Jr. and Farnam Jahanian. Cheap mutual exclusion. In USENIX [USE92g], pages 55–62. ISBN 1-880446-44-

8. LCCN QA 76.76 O63 U83 1992.

McCanne:1993:BPF

[MJ93] Steven McCanne and Van Jacobson. The BSD packet filter: a new architecture for user-level packet capture. In USENIX [USE93l], pages 259–

269. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/mccanne.pdf. Best stu-

dent paper winner.

Malan:1997:SPB

[MJS97] G. Robert Malan, Farnam Jahanian, and Sushila Subramanian. Salamander: a pushbased distribution substrate for Internet applications. In USE-NIX [USE97h], pages 171–181. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/malan.html.

McVoy:1991:ELP

[MK91] L. W. McVoy and S. R. Kleiman.

Extent-like performance from a UNIX file system. In USENIX [USE91l], pages 33–43. FTP - ? ucbarpa.berkeley.edu:/pub/mcvoy.clust.usenix.ps.Z.

Metz:1992:DDJ

[MK92] Melissa Metz and Howie Kaye.

DeeJay — the dump jockey: a heterogeneous network backup system. In USENIX [USE92e], pages 127–133. ISBN 1-880446-47-2. LCCN ????

Miller:1993:AFM

[MK93] Ethan L. Miller and Randy H.

Katz. An analysis of file migration in a UNIX supercomputing environment. In USENIX [USE93l], pages 421–

433. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Miller:1995:REA

[MK95] Ethan L. Miller and Randy H.

Katz. RAMA: Easy access to a high-bandwidth massively parallel file system. In USENIX [USE95e], pages 59–70 (or 59– 69??). ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/miller.html.

Miller:1998:UTS

[MK98] Michael L. Miller and Srikumar Kareti. Using Tcl to script CORBA interactions in a distributed system. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl98/miller.
                                        html.

Montgomery:1999:SOS

[MK99] Michael Montgomery and Ksheerabdhi Krishna. Secure object sharing in Java Card. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/smartcard99/
                                        montgomery.html.

McKusick:1990:PMB

[MKB90] Marshall Kirk McKusick, Michael J. Karels, and Keith Bostic. A pageable memory based file system. In USENIX [USE90g], pages 137–144. LCCN QA76.8.U65 U81 1990.

Mock:1992:IAO

[MKC92] Michael Mock, Reinhold Kroeger, and Vinny Cahill. Implementing atomic objects with the RelaX transaction facility. In USENIX [USE92c], pages 259–304.

Markatos:1999:SSM

[MKPF99] Evangelos P. Markatos, Manolis G. H. Katevenis, Dionisis Pnevmatikatos, and Michail Flouris. Secondary storage management for Web proxies. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/markatos.html.

Miller:1990:ROO

[ML90] David J. Miller and R. Charleene Lennox. RIPE: An object oriented robot independent programming environment. In USENIX [USE90j], pages 115– 124.

Meyers:1991:ADC

[ML91a] Scott Meyers and Moises Lejter.

Automatic detection of C++ programming errors: Initial thoughts on a lint++. In USE-NIX [USE91k], pages 29–40. LCCN QA76.73.C153 U27 1991.

Milliken:1991:NT

[ML91b] Walter Milliken and Gregory Lauer. A network toolkit. In USENIX [USE91k], pages 203–

216. LCCN QA76.73.C153 U27 1991.

Morrison:1992:TTT

[ML92] Brad Morrison and Karl Lehenbauer. Tcl and Tk: Tools for the system administrator. In USENIX [USE92e], pages 225–

234. ISBN 1-880446-47-2. LCCN ????

Moran:1993:RMF

[ML93] Joe Moran and Bob Lyon.

The Restore-o-Mounter: The file motel revisited. In USE-NIX [USE93g], pages 45–58. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/moran.html.

Milojicicc:1998:MOA

[MLC98] Dejan S. Milojicicc, William LaForge, and Deepika Chauhan. Mobile objects and agents (MOA). In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/milojicic.html.

Moon:1995:UTT

[MLL95] George C. Moon, Alex Lee, and Stephen Lindsey. Using Tcl/Tk to program a full functional geographic information system. In USENIX [USE95j], pages 123–129 (or 123–130??). ISBN 1-880446-72-3. LCCN QA76.73.T44

T44 1995. URL http://www.
usenix.org/publications/library/
proceedings/tcl95/moon.html.

Mullender:1994:OSS

[MLM94] Sape J. Mullender, Ian M. Leslie, and Derek McAuley. Operating system support for distributed multimedia. In USENIX [USE94j], pages 209–

219. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/mullender.html.

Magnusson:1998:SSV

[MLM^{+}98] Peter S. Magnusson, Fredrik Larsson, Andreas Moestedt, Bengt Werner, Fredrik Dahlgren, Magnus Karlsson, Fredrik Lundholm, Jim Nilsson, Per Stenström, and Håkan Grahn. Sim-ICS/sun4m: a virtual workstation. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/magnusson.html.

Mutka:1992:OAT

[MM92] Matt W. Mutka and Philip K.

McKinley. The OpenSim approach — tools for management and analysis of simulation jobs. In USENIX [USE92j], pages 291–304.

McCormack:1993:SSF

[MM93] Joel McCormack and Bob McNamara. Sketch of the smart frame buffer. In USE-NIX [USE93l], pages 169–179. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Miller:1996:CAD

[MM96a] Mark Miller and Joe Morris. Centralized administration of distributed firewalls. In USENIX [USE96j], pages 19–23. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/

publications/library/proceedings/
lisa96/marcus.html.

Muenchaisri:1996:SCE

[MM96b] Pornsiri Muenchaisri and Toshimi Minoura. Software composition with extended entityrelationship diagrams. In USE-NIX [USE96g], pages 113–

127. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/muench.html.

Muller:1997:HFE

[MMBC97] Gilles Muller, Bárbara Moura, Fabrice Bellard, and Charles Consel. Harissa: a flexible and efficient Java environment mixing bytecode and compiled code. In USENIX [USE97g], pages 1–20. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

URL http://www.irisa.fr/
compose/harissa/coots_web.
ps.gz;   http://www.irisa.
fr/compose/harissa/harissa.
html;   http://www.oasis.
leo.org/java/development/
java-to-c/Harissa.dsc.html;
 http://www.usenix.org/publications/
library/proceedings/coots97/

muller.html. See also [LH97, Rie97, WW97].

Minshall:1994:ONO

[MMP94] Greg Minshall, Drew Major, and Kyle Powell. An overview of the NetWare operating system. In USENIX [USE94l], pages 355–372. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/minshall.html.

Minnear:1992:ISM

[MMR92] Robert E. Minnear, Patrick A.

Muckelbauer, and Vincent F. Russo. Integrating the Sun microsystems XDR/RPC protocols into the C++ stream model. In USENIX [USE92k], pages 295–311. ISBN 1-880446-45-6. LCCN ????

Matias:1997:LSP

[MMS97] Yossi Matias, Alain Mayer, and Avi Silberschatz. Lightweight security primitives for Ecommerce. In USENIX [USE97h], pages 95–102. ISBN 1-880446-91-X. LCCN TK5105.875.I57

U96 1997. URL http://www.
usenix.org/publications/library/
proceedings/usits97/matias.
html.

Mason:1999:AOO

[MO99] Jeff Mason and Emil S. Ochotta.

The application of objectoriented design techniques to the evolution of the architecture of a large legacy software system. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/mason.html.

Moen:1992:DEH

[Moe92] Doug Moen. A discipline of error handling. In USENIX [USE92g], pages 123–134. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Moffat:1992:EIL

[Mof92] Alistair Moffat. Economical inversion of large text files. In USENIX [USE92b], pages 125– 139.

Mogg:1990:ATC

[Mog90] Gill Mogg. AT&T column: C++ and object oriented programming. EUUG Newsletter, 10(1):66–68, Spring 1990. CO-DEN EONLE8. ISSN 1011-4211.

Mogul:1992:RPS

[Mog92] Jeffrey C. Mogul. A recovery protocol for Spritely NFS. In USENIX [USE92h], pages 93–110. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Mogul:1994:BUP

[Mog94a]  Jeffrey C. Mogul.  A bet-

ter update policy. In USE-NIX [USE94j], pages 99–111. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/mogul.html.

Mogul:1994:RSN

[Mog94b] Jeffrey C. Mogul. Recovery in Spritely NFS. In USENIX [USE94b], pages 201–262.

Mogul:1999:SSC

[Mog99] Jeffrey Mogul. Speedier squid: a case study of an Internet server performance problem. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
squid.html.

Mohler:1999:SPSa

[Moh99a] Bruce W. Mohler. System profiles with syssumm. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
syssumm.html.

Mohler:1999:SPSb

[Moh99b] Bruce W. Mohler. System profiles with syssumm — A follow-up. ;login: the USE-NIX Association newsletter, 24 (5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/syssumm.html.

Molitor:1995:AAP

[Mol95] Andrew Molitor. An architecture for advanced packet filtering and access policy. In USENIX [USE95f], pages 117–126. ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
molitor.html.

Monroe:1993:STR

[Mon93] Mark A. Monroe. Security tool review: TCP wrappers. ;login: the USENIX Association newsletter, 18(6):15–16, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Moriarty:1992:SVW

[Mor92] Paul Moriarty. SAGE views: What one quality do you value most in a system administrator? ;login: the USENIX Association newsletter, 17(6):12–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Morris:1993:KA

[Mor93] Robert H. Morris, Sr. Keynote address. In USENIX [USE93m], page ?? ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec4/.

Moskowitz:1994:RCI

[Mos94a] Adam Moskowitz. Review: Connecting to the Internet. ;login: the USENIX Association newsletter, 19(4):44–??, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Moskowitz:1994:RLK

[Mos94b] Adam S. Moskowitz. Review: Learning the Korn Shell. ;login: the USENIX Association newsletter, 19(1):48–49, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Moskowitz:1996:RTI

[Mos96] Adam S. Moskowitz. Review: TCP/IP illustrated, volume 3. ;login: the USENIX Association newsletter, 21(4):54–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Mott:1991:LGA

[Mot91] Arch Mott. Link globally, act locally: a centrally maintained database of symlinks. In USE-NIX [USE91f], pages 127–128.

Mott:1995:BMH

[Mot95] Archibald C. R. Mott. Bringing the MBONE home: Experiences with internal use of multicast-based conferencing tools. In USENIX [USE95h], pages 103–109 (or 103–110??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/mott.html.

Massalin:1990:FGA

[MP90] Henry Massalin and Calton Pu.

Fine-grain adaptive scheduling using feedback. In USENIX [USE90d], pages 139–174.

Massalin:1992:RSK

[MP92] Henry Massalin and Calton Pu.

Reimplementing the synthesis kernel. In USENIX [USE92i], pages 177–186. ISBN 1-880446-42-1. LCCN QAX 32.

McNutt:1993:WDA

[MP93] Dinah McNutt and Michael Pearlman. Where did all the bytes go? In USE-NIX [USE93f], pages 157–

163. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/pearlman.html.

Mosberger:1996:MPE

[MP96] David Mosberger and Larry L. Peterson. Making paths explicit in the Scout operating system. In USENIX [USE96f], pages 153–167. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/mosberger.html.

Menon:1999:CSL

[MP99] Vijay Menon and Keshav Pingali. A case for sourcelevel transformations in MAT-LAB. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999. URL http://www.

usenix.org/publications/library/
proceedings/dsl99/menon.html.

McDaniel:1999:AFF

[MPH99] Patrick McDaniel, Atul Prakash, and Peter Honeyman. Antigone: a flexible framework for secure group communication. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

http://db.usenix.org/publications/
library/proceedings/sec99/
mcdaniel.html.

Madukkarumukumana:1998:HUL

[MPS98] Rajesh S. Madukkarumukumana, Calton Pu, and Hemal V. Shah. Harnessing user-level networking architectures for distributed object computing over high-speed networks. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/madukkarum.html.

Mohindra:1998:PNC

[MPZD98] Ajay Mohindra, Apratim Purakayastha, Deborra Zukowski, and Murthy Devarakonda. Programming network components using NetPebbles: An early report. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/mohindra.html.

Moreau:1997:DSQ

[MQ97] Luc Moreau and Christian Queinnec. Design and semantics of Quantum: a language to control resource consumption in distributed computing. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/moreau.html.

Moruzzi:1991:WSS

[MR91] Carla Moruzzi and Greg Rose.

Watson share scheduler. In USENIX [USE91f], pages 129– 134.

Muckelbauer:1995:LFI

[MR95] Patrick A. Muckelbauer and Vincent F. Russo. Lingua Franca: An IDL for structural subtyping distributed object systems. In USENIX [USE95k], pages 117–133 (or 117–134??). ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/muckelbauer.html.

Mogul:1996:ERL

[MR96] Jeffrey C. Mogul and K. K. Ramakrishnan. Eliminating receive livelock in an interruptdriven kernel. In USENIX [USE96k], pages 99–111 (or 99– 112??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/mogul.html.

Medvidovic:1997:DCS

[MR97a] Nenad Medvidovic and David S. Rosenblum. Domains of concern in software architectures and architecture description languages. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/medvidovic.html.

Murray:1997:EUI

[MR97b] Mark Murray and Troy Roberts. Effective use of individual user profiles with software distribution. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/murray.html.

McLeod:1998:TMP

[MR98] Jeanette McLeod and Greg Rose. Torn money and the PGP web of trust. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
tornmoney.html.

Miller:1997:EGM

[MRK97] Bradley N. Miller, John T.

Riedl, and Joseph A. Konstan. Experience with GroupLens: Making Usenet useful again. In USENIX [USE97e], pages 219–233. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/miller.html.

McClennan:1992:IMS

[MS92] Michael McClennan and Stuart Sechrest. Introducing multistructured file naming into UNIX. In USENIX [USE92h], pages 151–152. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

McKenney:1993:EKM

[MS93a] Paul E. McKenney and Jack Slingwine. Efficient kernel memory allocation on sharedmemory multiprocessor. In USENIX [USE93l], pages 295–

305. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Mukherjee:1993:ERM

[MS93b] Bodhisattwa Mukherjee and Karsten Schwan. Experimentation with a reconfigurable microkernel. In USENIX [USE93k], pages 45–60. ISBN 1-880446-52-

9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/mukherjee.html.

Mummert:1994:LGC

[MS94] Lily B. Mummert and M. Satyanarayanan. Large granularity cache coherence for intermittent connectivity. In USENIX [USE94j], pages 279–

289. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/mummert.html.

Mathaisel:1995:AEK

[MS95] Bernard F. Mathaisel and Timothy S. Simcoe. Alliance ecology: a key to multimedia strategic success. In USE-NIX [USE95g], pages 189–

208. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec95/index.html.

McGregor:1996:MTE

[MS96a] John McGregor and S. Srinivas. A measure of testing effort. In USENIX [USE96g], pages 129–141. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

                                        URL http://www.usenix.org/

publications/library/proceedings/
coots96/mcgregor.html.

McVoy:1996:LPT

[MS96b] Larry McVoy and Carl Staelin.

lmbench: Portable tools for performance analysis. In USENIX [USE96k], pages 279–294. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996. URL ftp://

ftp.bitmover.com/lmbench/;
http://bitmover.com/lmbench/
;   http://www.usenix.org/
publications/library/proceedings/
sd96/mcvoy.html.

Manley:1997:WFF

[MS97] Stephen Manley and Margo Seltzer. Web facts and fantasy. In USENIX [USE97h], pages 125–133. ISBN 1-880446-91-X. LCCN TK5105.875.I57

U96 1997. URL http://www.
usenix.org/publications/library/
proceedings/usits97/manley.
html.

Mehta:1998:EES

[MS98] Nimisha V. Mehta and Karen R.

Sollins. Expanding and extending the security features of Java. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/mehta.html.

Manber:1997:WCB

[MSG97] Udi Manber, Michael Smith, and Burra Gopal. WebGlimpse — combining browsing and searching. In USE-NIX [USE97e], pages 195–

206. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/manber.html.

Machines:1994:DES

[MSL94] Memory Machines, Daniel J. Scales, and Monica S. Lam. The design and evaluation of a shared object system for distributed. In USE-NIX [USE94i], pages 101–

114. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/index.html.

Miller:1993:SSA

[MSN93] Todd Miller, Christopher Stirlen, and Evi Nemeth. satool — A system administrator’s cockpit, an implementation. In USENIX [USE93f], pages 119–

129. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/miller.html.

Mitchell:1998:FSA

[MSS98] John C. Mitchell, Vitaly Shmatikov, and Ulrich Stern. Finitestate analysis of SSL 3.0. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998. URL

http://www.usenix.org/publications/
library/proceedings/sec98/
mitchell.html.

MacDonald:1999:OOP

[MSS99] Steve MacDonald, Duane Szafron, and Jonathan Schaeffer. Objectoriented pattern-based parallel programming with automatically generated frameworks. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/macdonald.html.

McCanne:1993:RSC

[MT93] Steven McCanne and Chris Torek. A randomized sampling clock for CPU utilization estimation and code profiling. In USENIX [USE93l], pages 387–394. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Minsky:1998:USH

[MU98] Naftaly H. Minsky and Victoria Ungureanu. Unified support for heterogeneous security policies in distributed systems. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

URL http://www.usenix.org/
publications/library/proceedings/
sec98/ungureanu.html.

Mudge:1999:HAI

[Mud99] Mudge. A hacker’s approach to ID. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-9/
                                        features/hackers.html. Spe-

cial issue on intrusion detection.

Mueller:1993:LIP

[Mue93] Frank Mueller. A library implementation of POSIX threads under UNIX. In USE-NIX [USE93l], pages 29–41. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Mueller:1994:BSS

[Mue94a] Scott Hazen Mueller. Business survival skills 101: The interview. ;login: the USENIX Association newsletter, 19(6):33– 36, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1994:OYS

[Mue94b] Scott Hazen Mueller. Optimizing your shell scripts. ;login: the USENIX Association newsletter, 19(4):25–28, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1994:UUW

[Mue94c] Scott Hazen Mueller. Understanding UNIX workstation performance. ;login: the USE-NIX Association newsletter, 19 (5):32–34, October 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Mueller:1995:I

[Mue95a] Scott Hazen Mueller. Impact. ;login: the USENIX Association newsletter, 20(5):35–37, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1995:OO

[Mue95b] Scott Hazen Mueller. Object oriented? ;login: the USE-NIX Association newsletter, 20 (3):32–33, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1995:PFU

[Mue95c] Scott Hazen Mueller. Possible future user interfaces. ;login: the USENIX Association newsletter, 20(4):38–39, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1995:STE

[Mue95d] Scott Hazen Mueller. Some thoughts on the economics of information. ;login: the USE-NIX Association newsletter, 20 (1):25–27, February 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Mueller:1995:WGG

[Mue95e] Scott Hazen Mueller. What good is a gig? ;login: the USE-NIX Association newsletter, 20 (2):45–46, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1996:PLH

[Mue96a] Harald Mueller. Pattern languages for handing C++ resources in an exception-safe way. In USENIX [USE96g], pages 165–178. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/mueller.html.

Mueller:1996:CUS

[Mue96b] Scott Hazen Mueller. Can USIX survive? ;login: the USE-NIX Association newsletter, 21 (1):5–6, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1996:GSL

[Mue96c] Scott Hazen Mueller. Good software, lousy installation. ;login: the USENIX Association newsletter, 21(3):36–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1996:ID

[Mue96d] Scott Hazen Mueller. The importance of documentation. ;login: the USENIX Association newsletter, 21(5):35–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Mueller:1998:SL

[Mue98] Scott Hazen Mueller. Spam and the law. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-4/
                                        spam.html.

Muffet:1995:WHA

[Muf95] Alec Muffet. WAN-hacking with AutoHack: Auditing security behind the firewall. In USENIX [USE95f], pages 21–34. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

                                        www.usenix.org/publications/

library/proceedings/security95/
muffett.html.

Muffett:1999:SBD

[Muf99] Alec Muffett. SENSS Bruce:

Developing a tool to aid intranet security. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-11/
features/senss.html. Special

issue on security.

Mullerworth:1997:FGD

[Mul97] S. D. Mullerworth. A flexible GUI design system. In USENIX [USE97d], pages 163–

164. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_mullerworth.
html.

Murray:1992:STA

[Mur92] Robert B. Murray. A statically typed abstract representation for C++ programs. In USENIX [USE92k], pages 83–

97. ISBN 1-880446-45-6. LCCN ????

Murray:1993:BWB

[Mur93] R. B. Murray. Building wellbehaved type relationships in C++. In Waldo [Wal93b], pages 87–97. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Murphy:1999:SEC

[Mur99a] Finnbarr P. Murphy. The single European currency. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
standards/23.euro.html.

Murphy:1999:SRW

[Mur99b] Finnbarr P. Murphy. Standard report: Whither the X Window System? ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
29.xwindow.html.

Manber:1994:GTS

[MW94] Udi Manber and Sun Wu. GLIMPSE: a tool to search through entire file systems. In USENIX [USE94l], pages 23–32. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/manber.2.html.

Mullen:1996:MBN

[MW96] Tracy Mullen and Michael P. Wellman. Market-based negotiation for Digital library services. In USENIX [USE96h], pages 259–269. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Maples:1998:LER

[MW98a] Greg Maples and Jon Williams.

Letter to the editor: In response to Lee Damon. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Mullen:1998:AMM

[MW98b] Tracy Mullen and Michael P.

Wellman. The Auction Manager: Market middleware for large-scale electronic commerce. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

URL http://www.usenix.org/
publications/library/proceedings/
ec98/mullen.html.

Manheimer:1990:DFS

[MWCR90] Kenneth Manheimer, Barry A.

Warsaw, Stephen N. Clark, and Walter Rowe. The depot: a framework for sharing software installation across organizational and UNIX platform boundaries. In USENIX [USE90f], pages 37–46. LCCN QA76.76.O63 L37 1990.

Manheimer:1991:DFS

[MWCR91] Ken Manheimer, Barry Warsaw, Steve Clark, and Walter Rowe. The depot: a framework for sharing software installation across organizational and UNIX platform boundaries. In USE-NIX [USE91e], page ?? ISBN ???? LCCN ????

Murray:1993:DIO

[MWO^{+}93] Kevin Murray, Tim Wilkinson, Peter Osmon, Ashley Saulsbury, Tom Stiemerling, and Paul Kelly. Design and implementation of an object-oriented 64bit single address space microkernel. In USENIX [USE93k], pages 31–43. ISBN 1-880446-52-

9. LCCN QA76.9.A73 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
micro93/murray.html.

Murphy:1996:APM

[MZA96] B. J. Murphy, S. Zeadally, and C. J. Adams. An analysis of process and memory models to support high-speed networking in a UNIX environment. In USENIX [USE96k], pages 239–251 (or 239–252??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/murphy.html.

Milojicic:1993:TMT

[MZDG93] Dejan Milojicic, Wolfgang Zint, Andreas Dangel, and Peter Giese. Task migration on the top of the Mach microkernel. In USENIX [USE93h], pages 273–289. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Nadkami:1992:PFS

[Nad92] Ashok V. Nadkami. The “processor” file system in UNIX SVR4.2. In USENIX [USE92h], pages 131–132. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Nagaratnam:1996:JAA

[Nag96] Nataraj Nagaratnam. Java applets and the AWT. In USE-NIX [USE96g], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

usenix.org/publications/library/
proceedings/coots96/.

Nakatani:1999:JDE

[NAOP99] Lloyd H. Nakatani, Mark A.

Ardis, Robert G. Olsen, and Paul M. Pontrelli. Jargons for domain engineering. In USE-NIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/nakatani.html.

Nassif:1996:DN

[Nas96] Nicholas S. Nassif. Defamation and the net. ;login: the USE-NIX Association newsletter, 21 (5):27–??, October 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Nather:1992:SVW

[Nat92] Wendy Nather. SAGE views:

Whither the customer, part two. ;login: the USENIX Association newsletter, 17(6):10–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Natter:1993:SV

[Nat93] Wendy Natter. SAGE views. ;login: the USENIX Association newsletter, 18(4):22–23, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Nather:1996:WDP

[Nat96] Wendy Nather. What do politics have to do with system administration? ;login: the USE-NIX Association newsletter, 21 (1):16–17, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Neuman:1993:UPS

[NAU93] B. Clifford Neuman, Steven Seger Augart, and Shantaprasad Upasani. Using Prospero to support integrated locationindependent computing. In USENIX [USE93i], pages 29–

34. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mobile93/neuman.html.

Nackman:1994:BCC

[NB94] Lee Nackman and John Barton. Base class composition with multiple derivation and virtual bases. In USE-NIX [USE94e], pages 57–71. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        c++94/nackman.html.

Nagaratnam:1997:RAC

[NB97] Nataraj Nagaratnam and Steven B.

Byrne. Resource access control for an Internet user agent. In USENIX [USE97g], pages 199–208. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/nagaratnam.html. See

also [LH97, Rie97, WW97].

Nelson:1992:HWS

[NC92] Bruce Nelson and Yu-Ping Cheng. How and why SCSI is better than IPI for NFS. In USENIX [USE92j], pages 253– 270.

Neuman:1994:OSM

[NC94] Michael Neuman and Gary Christoph. The operator shell: a means of privilege distribution under Unix. In USENIX [USE94m], pages 107–115. ISBN ???? LCCN ????

Neely:1990:SDV

[Nee90] Richard Neely. System design and verification for secure applications under UNIX. In USE-NIX [USE90i], pages 125–128. LCCN QA 76.9 A25 U55 1990.

Nemeth:1999:UNC

[Nem99a] Evi Nemeth. 2000 USE-NIX nominating committee. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

                                        publications/login/1999-10/
                                        nominating.html.

Nemeth:1999:STY

[Nem99b] Mike Nemeth. Sysadmin tools your vendor never told you about — revisited. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1999-8/tools.html.

Nessett:1999:MDS

[Nes99] Dan Nessett. Massively distributed systems: Design issues and challenges. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/es99/
                                        nessett.html.

Neuman:1992:PFSb

[Neu92a] B. Clifford Neuman. The Prospero file system. In USE-NIX [USE92h], pages 13–28. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Neuman:1992:PFSa

[Neu92b] B. Clifford Neuman. The Prospero File System: a global file system based on the virtual system model. In USENIX [USE92a], pages 407–432.

New:1995:IIC

[New95]  Darren New.  Internet in-

formation commerce: The First Virtual(R) approach. In USENIX [USE95g], pages 33–

68. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/new.html.

Neve:1999:FSC

[NFQ99] Amaury Nève, Denis Flandre, and Jean-Jacques Quisquater. Feasibility of Smart Cards in silicon-on-insulator (SOI) technology. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/neve.html.

Nelson:1993:HPD

[NH93] Michael N. Nelson and Graham Hamilton. High performance dynamic linking through caching. In USENIX [USE93g], pages 253–265. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/nelson.html.

Nilsen:1994:RRT

[Nil94] Kelvin Nilsen. Reliable real-time garbage collection of C++. In USENIX [USE94a], pages 467– 504.

Nikander:1998:JBC

[NK98] Pekka Nikander and Arto Karila. A Java beans component architecture for cryptographic protocols. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec98/nikander.html.

Nemeth:1999:IM

[Nkc99] Evi Nemeth and k. claffy. Internet measurements, 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neta99/invitedtalks.html.

Unpublished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Nelson:1993:EBF

[NKM93] Michael N. Nelson, Yousef A. Khalidi, and Peter W. Madany. Experience building a file system on a highly modular operating system. In USE-NIX [USE93j], pages 123–

141. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sedms4/nelson.html.

Nishioka:1996:PLK

[NKM96] Shingo Nishioka, Atsuo Kawaguchi, and Hiroshi Motoda. Process labeled kernel profiling: a new facility to profile system activities. In USENIX [USE96k], pages 295–306. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd96/nishioka.html.

Nakajima:1993:ERT

[NKT93] Tatsuo Nakajima, Takuro Kitayama, and Hideyuki Tokuda.

Experiments with real-time servers in real-time Mach. In USENIX [USE93h], pages 1–

19. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Necula:1996:SKE

[NL96] George C. Necula and Peter Lee. Safe kernel extensions without run-time checking. In USENIX [USE96f], pages 229–243. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/necula.html. Award

paper.

Nagaratnam:1998:SDD

[NL98] Nataraj Nagaratnam and Doug Lea. Secure delegation for distributed object environments. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/nagaratnam.html.

Nimmagadda:1999:PPA

[NLG^{+}99] S. Nimmagadda, C. Liyanaarachchi, A. Gopinath, D. Niehaus, and A. Kaushal. Performance patterns: Automated scenario-based ORB performance evaluation. In USE-NIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/nimmagadda.html.

Nimmagadda:1999:HEW

[NLZ99] Srinivas Nimmagadda, Joshua LeVasseur, and Rumi Zahir. High-end workstation compute farms using Windows NT. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt99/nimmagadda.html.

Narasimhan:1997:EII

[NMMS97a] P. Narasimhan, L. E. Moser, and P. M. Melliar-Smith. Exploiting the Internet Inter-ORB Protocol Interface to provide CORBA with fault tolerance. In USENIX [USE97g], pages 81–90. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots97/narasimhan.html. See

also [LH97, Rie97, WW97].

Narasimhan:1997:IAR

[NMMS97b] Priya Narasimhan, L. E. Moser, and P. M. Melliar-Smith. The interception approach to reliable distributed CORBA objects. In USENIX [USE97g], pages 245–248. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots97/panel.narasimhan.

html. See also [LH97, Rie97,

WW97].

Neville-Neil:1992:BRD

[NN92a] George Neville-Neil. Book review: The Design and Implementation of the 4.3 BSD UNIX OS Answer Book. ;login: the USENIX Association newsletter, 17(2):35–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neill:1992:FSG

[NN92b] George Neville-Neill. In favor of student grants. ;login: the USENIX Association newsletter, 17(5):24–??, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neill:1992:DMM

[NN92c] George V. Neville-Neill. Distributed multi-media: The bandwidth bandit. ;login: the USENIX Association newsletter, 17(6):28–29, November/ December 1992. CODEN LOGNEM. ISSN 1044-6397.

Neumann:1993:WXT

[NN93a] Gustaf Neumann and Stefan Nusser. Wafe — an X toolkit based frontend for application programs in various programming languages. In USE-NIX [USE93l], pages 181–192. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Neville-Neil:1993:RXR

[NN93b] George Neville-Neil. Review:

The X Resource. ;login: the USENIX Association newsletter, 18(4):42–??, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1994:RAT

[NN94a] George Neville-Neil. Review: Asynchronous Transfer Mode Solution for Broadband ISDN. ;login: the USENIX Association newsletter, 19(6):51–52, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1994:RAU

[NN94b] George Neville-Neil. Review: ATM User-Network Interface Specification Version 3.0. ;login: the USENIX Association newsletter, 19(6):51–52, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1994:RIT

[NN94c] George Neville-Neil. Review: Internetworking with TCP/IP. ;login: the USENIX Association newsletter, 19(4):43–44, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1995:RFH

[NN95a] George Neville-Neil. Review: The FDDI Handbook. ;login: the USENIX Association newsletter, 20(1):44–45, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1995:REH

[NN95b] George V. Neville-Neil. Review: The Early History of Data Networks. ;login: the USE-NIX Association newsletter, 20 (3):46–47, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1996:RII

[NN96a] George Neville-Neil. Review:

IPng: Internet protocol next generation. ;login: the USE-NIX Association newsletter, 21 (4):56–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Neville-Neil:1996:RTI

[NN96b] George V. Neville-Neil. Review: TCP/IP illustrated Vol.I (the protocols), TCP/IP illustrated Vol.II (the implementation). ;login: the USENIX Association newsletter, 21(2):58–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Naor:1998:CRC

[NN98] Moni Naor and Kobbi Nissim. Certificate revocation and certificate update. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998. URL

http://www.usenix.org/publications/
library/proceedings/sec98/
nissim.html.

Noer:1998:CFW

[Noe98]  Geoffrey J. Noer.  Cyg-

win32: a free Win32 porting layer for UNIX(R). applications. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/noer.html;   http:
//www.usenix.org/publications/
library/proceedings/usenix-
nt98/noer_slides.

Norwood:1991:NUO

[Nor91] Peter Norwood. A NonStop UNIX operating system. In USENIX [USE91i], pages 325–

334. LCCN QA 76.76 O63 U84

1992.

Neumann:1999:EED

[NP99] Peter G. Neumann and Phillip A. Porras. Experience with EMERALD to date. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/detection99/
                                        neumann.html.

Noble:1995:PIAa

[NPS95a] Brian B. Noble, Morgan Price, and M. Satyanarayanan. A programming interface for applicationaware adaptation in mobile computing. In USENIX [USE95a], pages 345–363. URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/mob95/
                                        noble.html.

Noble:1995:PIAb

[NPS95b] Brian B. Noble, Morgan Price, and M. Satyanarayanan. A programming interface for applicationaware adaptation in mobile computing. In USENIX [USE95i], pages 57–66. ISBN 1-880446-69-

3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/noble.html.

Nichol:1993:CTT

[NPW93] John Nichol, David V. Pitts, and C. Thomas Wilkes. Convergence: a triple threat or how ATM will change the world. In USENIX [USE93j], pages 283–287. ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

Nelson:1994:UNS

[NR94] Michael N. Nelson and Sanjay R.

Radia. A uniform name service for Spring’s UNIX environment. In USENIX [USE94l], pages 201–209. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/nelson.html.

Namprempre:1999:ICL

[NSM99] Chanathip Namprempre, Jeremy Sussman, and Keith Marzullo. Implementing causal logging using OrbixWeb interception. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999. URL http://www.

usenix.org/publications/library/
proceedings/coots99/namprempre.
html.

Reporters:1994:LCR

[Num94a] Numerous Reporters. LISA ’93 conference reports. ;login: the USENIX Association newsletter, 19(1):6–15, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

NumerousReporters:1994:WCR

[Num94b] Numerous Reporters. Winter ’94 conference reports. ;login: the USENIX Association newsletter, 19(2):3–9, March/April 1994. CODEN LOGNEM. ISSN 1044-6397.

Nagamatsu:1997:EPT

[NUO^{+}97] M. Nagamatsu, H. Uno, A. Obana, H. Yoshida, M. Asai, Y. Oohata, and R. Hamatsu. Experience of prototyping Tcl/Tkbased GUI for Geant4: an object-oriented toolkit for simulation in high energy physics. In USENIX [USE97d], page

187. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_nagamatsu.html.

North:1993:DGL

[NV93] Stephen C. North and Kiem-Phong Vo. Dictionary and graph libraries. In USE-NIX [USE93l], pages 1–11. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Nicol:1992:EAH

[NWEF92] John R. Nicol, C. Thomas Wilkes, Richard D. Edmiston, and J. C. Fitzgerald. Experiences with accommodating heterogeneity in a large scale telecommunications infrastructure. In USENIX [USE92l], pages 233–248. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Newman-Wolfe:1991:BOD

[NWRP^{+}91] R. E. Newman-Wolfe, C. L.

Ramirez, H. Pelimuhandiram, M. Montes, M. Webb, and D. L. Wilson. A brief overview of the DCS distributed conferencing system. In USENIX [USE91g], pages 437–452. LCCN QA 76.76 O63 U83 1991.

Nahum:1994:PIP

[NYKT94] Erich M. Nahum, David J.

Yates, James F. Kurose, and Don Towsley. Performance issues in parallelized network protocols. In USE-NIX [USE94i], pages 125–

137. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/nahum.html.

Nakajima:1991:MRE

[NYM91] Jun Nakajima, Masatomo Yazaki, and Hitoshi Matsumoto. Multimedia/realtime extensions for the Mach operating system. In USENIX [USE91g], pages 183–

198. LCCN QA 76.76 O63 U83 1991.

Nakajima:1992:MRE

[NYM92] Jun Nakajima, Masatomo Yazaki, and Hitoshi Matsumoto. Multimedia/realtime extensions for Mach 3.0. In USENIX [USE92i], pages 161–176. ISBN 1-880446-42-1. LCCN QAX 32.

Nystrom:1999:PCT

[Nys99]  Magnus Nyström.  PKCS

#15 — A cryptographictoken information format standard. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/nystrom.html.

Nater:1993:SH

[NZ93a] Wendy Nater and Elizabeth Zwicky. SAGE humor. ;login: the USENIX Association newsletter, 18(4):24–??, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Nather:1993:SV

[NZ93b] Wendy Nather and Elizabeth D. Zwicky. SAGE views. ;login: the USENIX Association newsletter, 18(6):17–19, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Neumann:1999:FLS

[NZ99] Gustaf Neumann and Uwe Zdun. Filters as a language support for design patterns in object-oriented scripting languages. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/neumann.html.

OBrien:1999:WNL

[O’B99a]  Robert O’Brien.  Windows

NT lights out operation guide,

1999.  URL ftp://ftp.
msftlabs.com/loop;   http:
//www.usenix.org/publications/
library/proceedings/lisa-
nt98/obrien_slides. Unpub-

lished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Oliva:1999:DIG

[OB99b] Alexandre Oliva and Luiz Eduardo Buzato. The design and implementation of Guaraná. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/oliva.html.

Olsson:1990:DGI

[OCH90] Ronald A. Olsson, Richard H.

Crawford, and W. Wilson Ho. Dalek: a GNU, improved programmable debugger. In USE-NIX [USE90g], pages 221–232. LCCN QA76.8.U65 U81 1990.

ODell:1990:G

[O’D90a] Michael D. O’Dell. Greetings. In USENIX [USE90b], pages 191– 192.

ODell:1990:PUV

[O’D90b] Michael D. O’Dell. Putting UNIX on very fast computers. In USENIX [USE90g], pages 239–

246. LCCN QA76.8.U65 U81 1990.

ODell:1994:Ga

[O’D94a] Michael D. O’Dell. Greetings. In USENIX [USE94c], pages 285–

ODell:1994:Gb

[O’D94b] Michael D. O’Dell. Greetings. In USENIX [USE94a], pages 415– 417.

Oetiker:1998:SSI

[Oet98a] Tobia Oetiker. SEPP — software installation and sharing system. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/oetiker2.html.

Oetiker:1998:MMR

[Oet98b] Tobias Oetiker. MRTG — the Multi Router Traffic Grapher. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/oetiker.html.

Osel:1995:OIS

[OG95] Peter W. Osel and Wilfried Gansheimer. OpenDist — incremental software distribution. In USENIX [USE95h], pages 181–194 (or 181–193??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/osel.html.

Ott:1995:PPW

[OH95] Maximilian Ott and John Hearn. Plug-and-play with wires. In USENIX [USE95j], pages 229–234 (or 229–233??). ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995. URL

http://www.usenix.org/publications/
library/proceedings/tcl95/
ott.html.

Oldroyd:1993:DC

[OK93] J. R. Oldroyd and Rob Kolstad.

Debate: To certify of not? ;login: the USENIX Association newsletter, 18(6):24–29, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Oestreicher:1999:OLJ

[OK99] Marcus Oestreicher and Ksheerabdhi Krishna. Object lifetimes in Java Card. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

http://www.usenix.org/publications/
library/proceedings/smartcard99/
oestreicher.html.

Okamoto:1992:NHH

[Oka92] Jeff Okamoto. “Nightly”: How to handle multiple scripts on multiple machines with one configuration file. In USENIX [USE92e], pages 171–173. ISBN 1-880446-47-2. LCCN ????

Okamoto:1998:RSIa

[Oka98a]  Jeff Okamoto.  Repetitive

strain injury: Causes, treatment, and prevention, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/invited_talks/okamoto_

html. Unpublished invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Okamoto:1998:RSIb

[Oka98b] Jeff Okamoto. Repetitive Strain Injury (RSI): Causes, treatment, and prevention, 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa98/invited_talks/okamoto_

html/. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6–11, 1998 Boston, Massachusetts, USA.

Orr:1993:FFS

[OLB93] Douglas B. Orr, Jay Lepreau, and John Bonn. Fast and flexible shared libraries. In USENIX [USE93g], pages 237–251. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/.

Oldroyd:1994:LVC

[Old94a] J. R. Oldroyd. LISA VIII conference reports. ;login: the USE-NIX Association newsletter, 19 (6):5–11, December 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Oldroyd:1994:SCR

[Old94b] J. R. Oldroyd. Summer ’94 conference reports. ;login: the USE-NIX Association newsletter, 19 (6):11–12, December 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Olson:1993:DII

[Ols93] Michael A. Olson. The design and implementation of the inversion file system. In USENIX [USE93l], pages 205–

217. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

ONeel:1997:BRJa

[O’N97a] Bruce O’Neel. Book review:

Java Network Programming. ;login: the USENIX Association newsletter, 22(4):33– 36, August 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/javaindex.

html. Reviews of [Har97b, HHSW97].

ONeel:1997:BRJb

[O’N97b]  Bruce O’Neel.  Book re-

view: Java AWT Reference. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
zukowski.html.

ONeel:1998:BRD

[O’N98] Bruce O’Neel. Book review:

Designing with JavaScript. ;login: the USENIX Association newsletter, 23(4): ??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
heinle.html.

Orman:1993:FGI

[OOM^{+}93] Hilarie Orman, Sean O’Malley, Edwin Menze, III, Larry Peterson, and Richard Schroeppel. A fast and general implementation of Mach IPC in a network. In USENIX [USE93h], pages 75–88. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mach3/.

Oppenheimer:1996:ISP

[Opp96] David L. Oppenheimer. Information security policies made easy. ;login: the USENIX Association newsletter, 21(4):58– ??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Orji:1992:DFS

[OS92] Cyril U. Orji and Jon A. Solworth. The Delta file system. In USENIX [USE92h], pages 143–144. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

OToole:1994:OLE

[OS94] James O’Toole and Liuba Shrira. Opportunistic log: Efficient installation reads in a reliable storage server. In USENIX [USE94i], pages 39–

48. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/otoole.html.

Okamoto:1992:MKA

[OSS^{+}92] Toshio Okamoto, Hideo Segawa, Sung Ho Shin, Hiroshi Nozue, Ken ichi Maeda, and Mitsuo Saito. A micro kernel architecture for next generation processor. In USENIX [USE92i], pages 83–94. ISBN 1-880446-42-1. LCCN QAX 32.

Ohta:1990:FTF

[OT90] Masataka Ohta and Hiroshi Tezuka. A fast /tmp file system by delay mount option. In USE-NIX [USE90g], pages 145–150. LCCN QA76.8.U65 U81 1990.

Ousterhout:1990:WAO

[Ous90a] John Ousterhout. Why aren’t operating systems getting faster as fast as hardware? In USE-NIX [USE90g], pages 247–256. LCCN QA76.8.U65 U81 1990.

Ousterhout:1990:TEC

[Ous90b] John K. Ousterhout. Tcl: An embeddable command language. In USENIX [USE90h], pages 133–146. LCCN QA76.8.U65 U82 1990.

Ousterhout:1991:OSR

[Ous91a] John Ousterhout. Operating systems research: Fads and true value. In USENIX [USE91h], page ?? LCCN QAX 27.

Ousterhout:1991:XTB

[Ous91b]  John K. Ousterhout.  An

X11 toolkit based on the Tcl language. In USE-NIX [USE91i], pages 105–115. LCCN QA 76.76 O63 U84

1992. URL ftp://allspice.
berkeley.edu/tkUsenix91.ps;
ftp://ftp.uu.net/languages/

tcl/tkUsenix91.ps. FTP - all-

spice.berkeley.edu:/tkUsenix91.ps; local - tkUsenix91.ps.

Ousterhout:1996:WTB

[Ous96] John Ousterhout. Why threads are A bad idea (for most purposes), 1996. Unpublished invited talk at the USENIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Overholt:1991:IRT

[Ove91] Jennifer Overholt. Integrating real-time video with Sun workstation. In USENIX [USE91g], pages 351–352. LCCN QA 76.76 O63 U83 1991.

Oyang:1994:ODM

[OW94] Yen-Jen Oyang and Le-Chun Wu. Optimal design of megabyte second-level caches for minimizing bus traffic in sharedmemory shared-bus multiprocessors. In USENIX [USE94c], pages 393–408.

Owen:1997:IMA

[Owe97] Charles B. Owen. The ImageTcl multimedia algorithm development system. In USE-NIX [USE97d], pages 97–

105. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/owen.html.

Pucci:1990:UHD

[PA90] Marc F. Pucci and J. L. Alberi. Using hints in DUNE remote procedure calls. In USE-NIX [USE90d], pages 47–68.

Pack:1993:RWG

[Pac93]  Jeff Pack.  Real-world gi-

gabit networking. In USE-NIX [USE93f], pages 29–

32. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/.

Page:1998:CDS

[Pag98] Christopher R. Page. Configuring database systems. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
page.html.

Palay:1992:CCE

[Pal92] Andrew J. Palay. C++ in a changing environment. In USE-NIX [USE92k], pages 195–206. ISBN 1-880446-45-6. LCCN ????

Phelan:1993:PM

[PAO93] James M. Phelan, James Arendt, and Gary R. Ormsby. An OS/2 personality on Mach. In USENIX [USE93h], pages 191–201. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Parrington:1990:RDP

[Par90] Graham D. Parrington. Reliable distributed programming in C++: The Arjuna approach. In USENIX [USE90j], pages 37–50.

Parseghian:1992:PMS

[Par92] P. Parseghian. Panel: Models of system administration. In USE-NIX [USE92e], page ?? ISBN 1-880446-47-2. LCCN ????

Parrington:1993:RDP

[Par93a] Graham D. Parrington. Reliable distributed programming in C++: The Arjuna approach. In Waldo [Wal93b], pages 235–248. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Parseghian:1993:SBMa

[Par93b] Pat Parseghian. SAGE board meeting highlights. ;login: the USENIX Association newsletter, 18(2):21–22, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Parseghian:1993:SBMb

[Par93c] Pat Parseghian. SAGE board meeting highlights. ;login: the USENIX Association newsletter, 18(5):16–17, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Partridge:1994:HPS

[Par94] Craig Partridge. High performance systems and protocols — myths and realities. In USENIX [USE94f], page ?? ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/

hispeed/index.html. Unpub-

lished keynote address.

Parrington:1995:SGS

[Par95] Graham D. Parrington. A stub generation system for C++. In USENIX [USE95b], pages 135– 169.

Parson:1999:UJR

[Par99a] Dale Parson. Using Java reflection to automate extension language parsing. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/parson.html.

Parter:1999:L

[Par99b]  David Parter.  LISA ’99.

;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
lisa.html.

Patience:1993:RSC

[Pat93] Simon Patience. Redirecting system calls in Mach 3.0, an alternative to the emulator. In USENIX [USE93h], pages 57–73. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Paterson:1997:TPR

[Pat97] Jessie Paterson. Teaching Perl to “real world” users. ;login: the USENIX Association newsletter, 22(2):27–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Paxson:1991:BRL

[Pax91] Vern Paxson. Book review: lex & yacc. ;login: the USENIX Association newsletter, 16(1): 23–??, January/February 1991. CODEN LOGNEM. ISSN 1044-6397.

Paxson:1993:RHP

[Pax93] Vern Paxson. Review: High Performance Computing. ;login: the USENIX Association newsletter, 18(6):47–48, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Paxson:1994:RTI

[Pax94] Vern Paxson. Review: TCP/IP illustrated. ;login: the USENIX Association newsletter, 19(2): 47–??, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Paxson:1998:BSD

[Pax98] Vern Paxson. Bro: a system for detecting network intruders in real-time. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec98/paxson.html.

Pardyak:1996:DBE

[PB96] Przemysław Pardyak and Brian Bershad. Dynamic binding for an extensible system. In USENIX [USE96f], pages 201–

212. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/pardyak.html.

Poger:1997:SPI

[PB97] Elliot Poger and Mary G. Baker.

Secure Public Internet Access Handler (SPINACH). In USE-NIX [USE97h], pages 113–123. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/poger.html.

Parter:1997:PDM

[PBEDT97] David Parter, David N. Blank-Edelman, Jim Duncan, and Christopher Thorpe. Panel discussion: Managing computers for academics, 1997.

URL http://www.ccs.neu.
edu/home/dnb/presentations/
97-10-LISA97/html/lisa97academia/

index.htm. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Porat:1996:COC

[PBF^{+}96] Sara Porat, David Bernstein, Yaroslav Fedorov, Joseph Rodrigue, and Eran Yahav. Compiler optimization of C++ virtual function calls. In USENIX [USE96g], pages 3–

14. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/

                                        publications/library/proceedings/
                                        coots96/porat.html.

Plank:1995:LTC

[PBKL95] James S. Plank, Micah Beck, Gerry Kingsley, and Kai Li. Libckpt: Transparent checkpointing under UNIX. In USENIX [USE95e], pages 213–

224. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neworl/plank.html.

Puaut:1991:EEB

[PBR91] Isabelle Puaut, Michel Banatre, and Jean-Paul Routeau. Early experience with building and using the Gothic distributed operating system. In USENIX [USE91j], pages 271–282. ISBN ???? LCCN ????

Parson:1997:TBS

[PBS97] Dale Parson, Paul Beatty, and Bryan Schlieder. A Tclbased self-configuring embedded system debugger. In USE-NIX [USE97d], pages 131–

138. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/parson.html.

Paas:1998:WAE

[PBS98] Sven M. Paas, Thomas Bemmerl, and Karsten Scholtyssik. Win32 API emulation on UNIX for software DSM. In USENIX [USE98a], page ?? ISBN 1-880446-95-2.

LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/paas.html;   http:
//www.usenix.org/publications/
library/proceedings/usenix-
nt98/paas_slides.pdf.

Pasieka:1991:DMH

[PCMI91] Michael Pasieka, Paul Crumley, Ann Marks, and Ann Infortuna. Distributed multimedia: How can the necessary data rates be supported? In USE-NIX [USE91g], pages 169–182. LCCN QA 76.76 O63 U83 1991.

Pai:1999:ILU

[PDZ99] Vivek S. Pai, Peter Druschel, and Willy Zwaenepoel. IO-Lite: a unified I/O buffering and caching system. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/pai.html.

Pascale:1992:VWS

[PE92] Rita Pascale and Jeremy Epstein. Virtual window systems: a new approach to supporting concurrent heterogeneous windowing systems. In USENIX [USE92g], pages 117–122. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992. A generalization of TRW’s prototype to nonsecurity applications.

Peacock:1992:FSM

[Pea92] J. Kent Peacock. File system multithreading in System V Release 4 MP. In USENIX [USE92g], pages 19–30. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Peak:1993:SDZ

[Pea93] Chris Peak. The San Diego “Zoo”: a multicomputer stress test suite. In USENIX [USE93l], pages 119–130. ISBN 1-880446-48-0. LCCN QA 76.76 O63

                                        U84 1993. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/sd93/.

Pease:1999:PWW

[Pea99] Holly Brackett Pease. Problems with world-wide networking, 1999. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/neta99/ITs/pease_

html/. Unpublished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Pedneault:1990:HLU

[Ped90] Michael Pedneault. A high level user interface toolkit for the X Window System and character terminals. In USE-NIX [USE90g], pages 307–314. LCCN QA76.8.U65 U81 1990.

Peek:1996:LEI

[Pee96] Jerry Peek. Letters to the editor: Improvements for the webmaster. ;login: the USENIX Association newsletter, 21(2):3–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Perkins:1992:FSW

[Per92] Drew Perkins. 1992 File Systems Workshop Report. ;login: the USENIX Association newsletter, 17(5):4–7, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Perrine:1998:SI

[Per98a] Tom Perrine. Security as infrastructure, 1998. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6–11, 1998 Boston, Massachusetts, USA.

Perry:1998:LEE

[Per98b] Alan Perry. Letter to the editor: Erratum. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Peterson:1993:TLA

[Pet93] Larry Peterson. Is there life after microkernels? In USENIX [USE93j], page ?? ISBN 1-880446-54-5. LCCN QA76.9.D5

U84 1993. URL http://www.
usenix.org/publications/library/
proceedings/sedms4/.

Petersen:1994:TTP

[Pet94] Karin Petersen. Tcl/Tk for a Personal Digital Assistant. In USENIX [USE94k], pages 41–55. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/index.html.

Petersen:1996:ORA

[Pet96] Karin Petersen. Opening remarks and awards presentation, 1996. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/osdi96/opening_
                                        remarks/proceedings.html.

Unpublished foreword at the USENIX 2nd Symposium on OS Design and Implementation (OSDI ’96), October 28–31, 1996, Seattle, Washington.

Pomeranz:1993:SV

[PEZ93] Hal Pomeranz, Paul Evans, and Elizabeth D. Zwicky. SAGE views. ;login: the USENIX Association newsletter, 18(3):8– 12, May/June 1993. CODEN LOGNEM. ISSN 1044-6397.

Pham:1992:MDA

[PG92] Thuan Q. Pham and Pankaj K. Garg. On migrating a distributed application to a multithreaded environment. In USE-NIX [USE92g], pages 45–54. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Page:1991:MRV

[PGH^{+}91] Thomas W. Page, Jr., Richard G. Guy, John S. Heidemann, Gerald J. Popek, Wai Mak, and Dieter Rothmeier. Management of replicated volume location data in the Ficus Replicated File System. In USENIX [USE91g], pages 17–30. LCCN QA 76.76 O63 U83 1991.

Peterson:1999:MR

[PH99] John Peterson and Greg Hager.

Monadic robotics. In USE-NIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/peterson.html.

Phelps:1995:TYT

[Phe95]  Thomas A. Phelps.  Two

years with TkMan: Lessons and innovations. In USE-NIX [USE95j], pages 243–

250. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/phelps.html.

Phillips:1991:MEM

[Phi91] Richard L. Phillips. MediaView:

An editable multimedia publishing system developed with an object-oriented toolkit. In USE-NIX [USE91g], pages 125–136. LCCN QA 76.76 O63 U83 1991.

Phillips:1995:SNB

[Phi95] Gretchen Phillips. From something to nothing (and back). In USENIX [USE95h], pages 117–124. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/phillips.html.

Pyarali:1996:DPO

[PHS96a] I. Pyarali, T. Harrison, and D. Schmidt. Design and performance of an object-oriented framework for high-performance electronic medical imaging. In USENIX [USE96g], pages 191–

208. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/pyarali.html.

Pyrali:1996:DPO

[PHS96b] Irfan Pyrali, Timothy H. Harrison, and Douglas C. Schmidt. Design and performance of an object-oriented framework for high-speed electronic medical imaging. In USENIX [USE96a], pages 331–375.

Poutanen:1998:NLP

[PHS98] Tomi Poutanen, Heather Hinton, and Michael Stumm. Net-Cents: a lightweight protocol for secure micropayments. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

URL http://www.usenix.org/
publications/library/proceedings/
ec98/poutanen.html.

Piele:1994:RIC

[Pie94a] Don Piele. Results of the International Computer Olympiad. ;login: the USENIX Association newsletter, 19(6):37–??, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Piele:1994:FIO

[Pie94b] Donald T. Piele. The Fifth International Olympiad in Informatics. ;login: the USENIX Association newsletter, 19(1):

36–38, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Piele:1995:UWM

[Pie95] Don Piele. USA wins medals at computer Olympics in the Netherlands. ;login: the USE-NIX Association newsletter, 20 (5):37–??, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Pierce:1996:ISA

[Pie96] Clinton Pierce. The Igor system administration tool. In USENIX [USE96j], pages 9–

18. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/cpierce.html.

Piele:1999:IRT

[Pie99a] Don Piele. IOI report from Turkey. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
turkey.html.

Piele:1999:UCO

[Pie99b]  Don Piele.  USA Com-

puting Olympiad picks final four. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
olympiad.html.

Pike:1991:PWS

[Pik91a] Rob Pike. 8-1/2, the Plan 9 window system. In USE-NIX [USE91g], pages 257–266. LCCN QA 76.76 O63 U83 1991.

                                        URL ftp://research.bell-
                                        labs.com/dist/plan9doc/4.

FTP - research.att.com:/dist/plan9doc/4; local - 8half.ps.

Pike:1991:MGU

[Pik91b]  Rob Pike.  A minimalist

global user interface. In USE-NIX [USE91g], pages 267–280. LCCN QA 76.76 O63 U83 1991.

Pike:1994:AUI

[Pik94] Rob Pike. Acme: a user interface for programmers. In USENIX [USE94l], pages 223–

234. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/pike.html.

Pintar:1992:PEV

[Pin92] Douglas L. Pintar. Practice and experience: Voices and FAXes and UNIX (oh, my!). ;login: the USENIX Association newsletter, 17(2):30–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397.

Pawlowski:1994:NVD

[PJS^{+}94] Brian Pawlowski, Chet Juszczak, Peter Staubach, Carl Smith, Diane Lebel, and Dave Hitz. NFS Version 3: Design and implementation. In USENIX [USE94j], pages 137–

151. ISBN 1-880446-62-6.

LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/index.html.

Polk:1991:ECB

[PK91] Jeff Polk and Rob Kolstad. Engineering a commercial backup program. In USENIX [USE91f], pages 173–180.

Patel:1993:UTD

[PK93] Dorab Patel and Scott D.

Kalter. A UNIX toolkit for distributed synchronous collaborative applications. In USENIX [USE93b], pages 105–133.

Perkovic:1996:ODR

[PK96] Dejan Perkovic and Pete Keleher. Online data-race detection via coherency guarantees. In USENIX [USE96f], pages 47–57. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/perkovic.html.

Peacock:1998:FCC

[PKA98] J. Kent Peacock, Ashvin Kamaraju, and Sanjay Agrawal. Fast consistency checking for the Solaris file system. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix98/
peacock.html.

Powell:1991:SMT

[PKB^{+}91] M. L. Powell, S. R. Kleiman, S. Barton, D. Shah, D. Stein, and M. Weeks. SunOS multithread architecture. In USENIX [USE91i], pages 65–80. LCCN QA 76.76 O63 U84 1992.

Plank:1993:JFP

[Pla93] James S. Plank. Jgraph — A filter for plotting graphs in PostScript. In USE-NIX [USE93l], pages 61–66. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Plastina:1997:ITM

[Pla97] Dan Plastina. Invited talk: MS/zero administration Windows. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/its/plastina.
it.

Plastina:1999:WNA

[Pla99] Dan Plastina. Windows NT administration: Change and configuration management, 1999.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/plastina_slides.

Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Pendry:1995:UM

[PM95a] Jan-Simon Pendry and Marshall Kirk McKusick. Union mounts in 4.4BSD-Lite. In USENIX [USE95e], pages 25–

33. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/mckusick.html.

Powell:1995:LEP

[PM95b] Patrick Powell and Justin Mason. LPRng — an enhanced printer spooler system. In USENIX [USE95h], pages 13–24. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/papowell.html.

Purakayastha:1998:SMN

[PM98] Apratim Purakayastha and Ajay Mohindra. System management with NetScript. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
purakayastha.html.

Provos:1999:FAP

[PM99] Niels Provos and David Mazieres.

A future-adaptable password scheme. In USENIX [USE99m], page ?? ISBN 1-880446-33-2.

LCCN ???? URL http://www.
openbsd.org/papers/bcrypt-
paper.ps.

Periakaruppan:1999:GGT

[PN99] Ram Periakaruppan and Evi Nemeth. GTrace — A graphical traceroute tool. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/periakaruppan.
html.

Polk:1993:ACW

[Pol93] W. Timothy Polk. Approximating Clark-Wilson access triples with basic UNIX commands. In USENIX [USE93m], pages 145–154. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/.

Pollice:1994:UVH

[Pol94]  Gary F. Pollice.  Using a

very high level language to build families of high quality reusable components. In USENIX [USE94k], pages 179–

190. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/pollice.html.

Polichar:1996:OMF

[Pol96] Valerie E. Polichar. An office MUD for fun and profit? Or maybe just better communication. ;login: the USENIX Association newsletter, 21(3):21–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Pollack:1997:LSD

[Pol97] Dan Pollack. A large scale data warehouse application case study. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/08.pollack.html.

Pomeranz:1993:PPa

[Pom93a] Hal Pomeranz. Perl practicum.

;login: the USENIX Association newsletter, 18(4):20–22, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1993:PPb

[Pom93b] Hal Pomeranz. Perl practicum.

;login: the USENIX Association newsletter, 18(5):24–25, September/October 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1993:PPc

[Pom93c] Hal Pomeranz. Perl practicum.

;login: the USENIX Association newsletter, 18(6):20–22, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1993:PKT

[Pom93d] Hal Pomeranz. PLOD: Keep track of what you’re doing. In USENIX [USE93f], pages 183–187. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/pomeranz.html.

Pomeranz:1994:LCM

[Pom94a] Hal Pomeranz. Living with CC:mail. ;login: the USENIX Association newsletter, 19(2): 15–16, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1994:NNC

[Pom94b] Hal Pomeranz. A new network for the cost of one SCSI cable: a simple caching strategy for third-party applications. In USENIX [USE94m], pages 117– 122. ISBN ???? LCCN ????

Pomeranz:1994:PPa

[Pom94c] Hal Pomeranz. Perl practicum. ;login: the USENIX Association newsletter, 19(1):25–26, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1994:PPb

[Pom94d] Hal Pomeranz. Perl practicum. ;login: the USENIX Association newsletter, 19(2):18–19, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1994:PPP

[Pom94e] Hal Pomeranz. Perl practicum: a plea for clarity. ;login: the USENIX Association newsletter, 19(6):20–21, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1994:SSX

[Pom94f] Hal Pomeranz. Surviving Solaris 2.X. ;login: the USE-NIX Association newsletter, 19 (5):27–28, October 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1994:SAT

[Pom94g] Hal Pomeranz. System administration tools your vendor never told you about: Painkiller. ;login: the USENIX Association newsletter, 19(6):24–??, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1995:PPD

[Pom95a] Hal Pomeranz. Perl practicum:

DB or not DB? ;login: the USE-NIX Association newsletter, 20 (6):30–32, December 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1995:PPF

[Pom95b] Hal Pomeranz. Perl practicum:

Fun with formats. ;login: the USENIX Association newsletter, 20(3):19–23, June 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1995:PPC

[Pom95c] Hal Pomeranz. Perl practicum:

The camel spins a Web. ;login: the USENIX Association newsletter, 20(4):19–22, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1996:MMD

[Pom96a] Hal Pomeranz. Many mail domains, one machine: The forwarding mailer. In USE-NIX [USE96j], pages 123– 130. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/hal.html.

Pomeranz:1996:PP

[Pom96b] Hal Pomeranz. Perl practicum. ;login: the USENIX Association newsletter, 21(2):31–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1996:PPF

[Pom96c] Hal Pomeranz. Perl practicum: Failed to understand the reference. ;login: the USENIX Association newsletter, 21(1):10– 13, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1996:PPNa

[Pom96d] Hal Pomeranz. Perl practicum: Network wiles (part I). ;login: the USENIX Association newsletter, 21(4):21–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1996:PPNb

[Pom96e] Hal Pomeranz. Perl practicum: Network wiles (part II). ;login: the USENIX Association newsletter, 21(5):23–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1996:PPD

[Pom96f] Hal Pomeranz. Perl practicum: The devil in the details. ;login: the USENIX Association newsletter, 21(3):15–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1997:BRS

[Pom97a] Hal Pomeranz. Book review:

System security: a management perspective. ;login: the USE-NIX Association newsletter, 22 (5):69–70, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1997:PPN

[Pom97b] Hal Pomeranz. Perl practicum:

Network wiles. ;login: the USE-NIX Association newsletter, 22 (1):12–16, February 1997. CO-DEN LOGNEM. ISSN 1044-6397.

Pomeranz:1997:PPE

[Pom97c] Hal Pomeranz. Perl practicum:

The email of the species. ;login: the USENIX Association newsletter, 22(2):14–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Pomeranz:1999:HIA

[Pom99] Hal Pomeranz. How-to: Install anonymous FTP. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-2/anonymous_ftp.html.

Pooquet:1994:BWK

[Poo94] Jean Pooquet. Beneath who knows how many Freudian and Calvinistic strata. ;login: the USENIX Association newsletter, 19(4):25–??, August 1994. CODEN LOGNEM. ISSN 1044-

Pyarali:1999:AOP

[POS^{+}99] Irfan Pyarali, Carlos O’Ryan, Douglas Schmidt, Nanbor Wang, Vishal Kachroo, and Aniruddha Gokhale. Applying optimization principle patterns to design real-time ORBs. In USE-NIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/pyarali.html.

Potter:1997:BRL

[Pot97]  Stephen Potter.  Book re-

view: Learning Perl, 2nd ed. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
schwartz.html.

Povey:1999:EWF

[Pov99] Dean Povey. Enforcing wellformed and partially formed transactions for UNIX. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

http://db.usenix.org/publications/
library/proceedings/sec99/
povey.html.

Pitkow:1999:MLR

[PP99] Jim Pitkow and Peter Pirolli. Mining longest repeating subsequences to predict World Wide Web surfing. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57

U96.  URL http://www.
usenix.org/publications/library/
proceedings/usits99/pitkow.
html.

Paul:1997:RWS

[PPC97] Santanu Paul, Edwin Park, and Jarir Chaar. Rain-Man: a workflow system for the Internet. In USE-NIX [USE97h], pages 159–170. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/paul.html.

Pike:1995:PBL

[PPD^{+}95] Rob Pike, Dave Presotto, Sean Dorward, Bob Flandrena, Ken Thompson, Howard Trickey, and Phil Winterbottom. Plan 9 from Bell labs. In USENIX [USE95c], pages 221–254.

Pande:1994:STD

[PR94] Hemant Pande and Barbara Ryder. Static type determination for C++. In USE-NIX [USE94e], pages 85–97. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/pande.html.

Presotto:1992:PDS

[Pre92] David Presotto. Plan 9, A distributed system. In USENIX [USE92i], pages 31–38. ISBN 1-880446-42-1. LCCN QAX 32.

Presotto:1995:E

[Pre95] David L. Presotto. Editorial. In USENIX [USE95d], pages 1–??

Preston:1997:EBR

[Pre97] W. Curtis Preston. Enterprise backup and recovery — do you need a commercial utility?, 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/lisa97/invited_
                                        talks/preston/preston.pdf;
                                         http://www.usenix.org/publications/
                                        library/proceedings/lisa97/
                                        invited_talks/preston/preston.

ps. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Preston:1998:UGE

[Pre98] W. Curtis Preston. Using Gigabit Ethernet to backup six terabytes. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa98/preston.html.

Prevelakis:1999:SSN

[Pre99] Vassilis Prevelakis. A secure station for network monitoring and control. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

                                        1999. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/sec99/prevelakis.
                                        html.

Pukatzki:1992:ANM

[PS92] Dieter Pukatzki and Johann Schumann. AUTOLOAD: The network management system. In USENIX [USE92e], pages 97–

104. ISBN 1-880446-47-2. LCCN ????

Paxson:1993:GUL

[PS93] Vern Paxson and Chris Saltmarsh. Glish: a user-level software bus for loosely-coupled distributed systems. In USE-NIX [USE93l], pages 141–155. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Perl:1996:SWN

[PS96] Sharon E. Perl and Richard L.

Sites. Studies of Windows NT performance using dynamic execution traces. In USENIX [USE96f], pages 169–

183. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/perl.html.

Pyarali:1997:RTC

[PS97] Irfan Pyarali and Peter H. Salus.

Reports on the third conference on object-oriented technologies and systems (COOTS). ;login: the USENIX Association newsletter, 22(4):53–60, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Perham:1997:RTD

[PSJL97] Mike Perham, Brian C. Smith, Tibor Janosi, and Ioi K. Lam. Redesigning Tcl-DP. In USENIX [USE97d], pages 49–53. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/perham.html.

Peacock:1992:EMS

[PST^{+}92] J. Kent Peacock, Sunil Saxena, Dean Thomas, Fred Yang, and Wilfred Yu. Experiences from multithreading system V release 4. In USENIX [USE92l], pages 77–92. ISBN 1-880446-41-

3. LCCN QA76.9.D3 S954 1992.

Parrington:1995:DIA

[PSWL95] Graham D. Parrington, Santosh K. Shrivastava, Stuart M. Wheater, and Mark C. Little. The design and implementation of Arjuna. In USENIX [USE95c], pages 255–308.

Paciorek:1992:OOF

[PT92] Noemi Paciorek and Marc Teller. An object oriented, file system independent, distributed file server. In USE-NIX [USE92h], pages 45–62. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Paciorek:1993:RFS

[PT93a] Noemi Paciorek and Marc J. Teller. Response to file system workshop report. ;login: the USENIX Association newsletter, 18(1):25–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Pike:1993:HWD

[PT93b] Rob Pike and Ken Thompson. Hello world (distributed OS text format). In USE-NIX [USE93l], pages 43–50. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Proebsting:1997:TJA

[PTB^{+}97] Todd A. Proebsting, Gregg Townsend, Patrick Bridges, John H. Hartman, Tim Newsham, and Scott A. Watterson. Toba: Java for applications — A Way Ahead of Time (WAT) compiler. In USE-NIX [USE97g], pages 41–53. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997. URL

ftp://ftp.cs.arizona.edu/
reports/1997/TR97-01.pdf;
 ftp://ftp.cs.arizona.edu/
reports/1997/TR97-01.ps;
http://www.usenix.org/publications/
library/proceedings/coots97/
proebsting.html. See also

[LH97, Rie97, WW97].

Pucci:1991:CDM

[Puc91a] Marc F. Pucci. Configurable data manipulation in an attached multiprocessor. In USE-NIX [USE91c], pages 217–242.

Pucci:1991:IDE

[Puc91b] Marc F. Pucci. The ION data engine. In USENIX [USE91j], pages 35–50. ISBN ???? LCCN ????

Pucella:1999:HOC

[Puc99] Riccardo Pucella. Higher-order concurrent Win32 programming. In USENIX [USE99d], page ?? ISBN 1-880446-29-4.

                                        LCCN ???? URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usenix-nt99/pucella.
                                        html.

Pultar:1998:ASC

[Pul98]  Giray Pultar.  Automati-

cally selecting a close mirror based on network topology. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa98/pultar.html.

Pultar:1999:DIF

[Pul99]  Giray Pultar.  Design and

implementation of a failsafe print system. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

                                        URL http://db.usenix.org/
                                        publications/library/proceedings/
                                        lisa99/pultar.html.

Pozzetti:1995:ITE

[PV95] Ennio Pozzetti and Vidar Vetland. Improving the trustworthiness of evidence derived from security trace files. In USE-NIX [USE95f], pages 187–198 (or 187–197??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
pozzetti.html.

Paciorek:1991:DMO

[PVL91] Noemi Paciorek, Susan Lo Verso, and Alan Langerman. Debugging multiprocessor operating system kernels. In USE-NIX [USE91j], pages 185–202. ISBN ???? LCCN ????

Presotto:1993:ONP

[PW93] Dave Presotto and Phil Winterbottom. The organization of networks in Plan 9. In USENIX [USE93l], pages 271–

280. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Proebsting:1997:KDJ

[PW97] Todd A. Proebsting and Scott A.

Watterson. Krakatoa: Decompilation in Java (does bytecode reveal source?). In USE-NIX [USE97g], pages 185–

197. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/proebsting2.html.

The authors demonstrate that, because of the strongly-type nature of the Java Virtual Machine, it is possible to reverseengineer Java byte codes back to quite readable correct Java source code, making it difficult to conceal Java source code by distributing only compiled byte codes. This has significant ramifications for commercial software development in Java.

Quarterman:1998:FI

[Qua98] John S. Quarterman. The future of the Internet, 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix98/invited_talks/quarterman.
                                        ps;   http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix98/invited_talks/quarterman_

html. Unpublished invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Quin:1994:TRP

[Qui94] Liam R. E. Quin. A text retrieval package for the Unix operating system. In USE-NIX [USE94j], pages 231–243. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/quin.html.

Quong:1991:BRC

[Quo91] Russell Quong. Book review: A C++ Toolkit and C++ Primer, 2nd Edition. ;login: the USE-NIX Association newsletter, 16 (6):15–??, November/December 1991. CODEN LOGNEM. ISSN 1044-6397.

Ranum:1994:TMI

[RA94] Marcus K. Ranum and Frederick M. Avolio. A toolkit and methods for Internet firewalls.

In USENIX [USE94j], pages 37–44. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/index.html.

Reed:1998:OOF

[RA98] John A. Reed and Abdollah A. Afjeh. An objectoriented framework for distributed computational simulation of aerospace propulsion systems. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/reed.html.

Rodrigues:1997:HPL

[RAC97] Steven H. Rodrigues, Thomas E.

Anderson, and David E. Culler. High-performance local area communication with fast sockets. In USENIX [USE97e], pages 257–274. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/rodrigues.html.

Radatti:1997:MH

[Rad97] Peter V. Radatti. MrMean the hacker. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
mrmean.html.

Rosenthal:1992:PCC

[RAF92] Douglas Rosenthal, Wayne Allen, and Kenneth Fiduk. Process control and communication in distributed CAD environments. In USENIX [USE92j], pages 271–282.

Ramey:1994:BGS

[Ram94] Chet Ramey. Bash — the GNU shell. ;login: the USENIX Association newsletter, 19(6):25– 33, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Reiter:1998:DHS

[RAM98] Michael K. Reiter, Vinod Anupam, and Alain Mayer. Detecting hit shaving in click-through payment schemes. In USENIX [USE98b], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74 1998. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/ec98/reiter.html.

Rankin:1997:FAN

[Ran97] Brian C. Rankin. Filesharing with Appletalk and NetBios. ;login: the USENIX Association newsletter, 22(2):23–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Ranum:1999:SSF

[Ran99] Marcus J. Ranum. Selling security: Fear leads to … the dark side. ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-11/
                                        features/darkside.html. Spe-

cial issue on security.

Rao:1998:UJa

[Rao98a] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava11.html.

Rao:1998:UJb

[Rao98b]  Prithvi Rao.  Using Java.

;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava12.
html.

Rao:1998:UJc

[Rao98c]  Prithvi Rao.  Using Java.

;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava13.
html.

Rao:1998:UJS

[Rao98d]  Prithvi Rao.  Using Java:

Security outside the JVM. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava14.
html.

Rao:1999:UJ

[Rao99a]  Prithvi Rao.  Using Java.

;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava19.
html.

Rao:1999:UJA

[Rao99b] Prithvi Rao. Using Java: Applets and animation. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava16.html.

Rao:1999:UJI

[Rao99c]  Prithvi Rao.  Using Java:

Input and output streams. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava18.
html.

Rao:1999:UJR

[Rao99d]  Prithvi Rao.  Using Java:

Remote Method Invocation. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava17.
html.

Rao:1999:UJJ

[Rao99e]  Prithvi Rao.  Using Java:

The Java Native Interface. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava15.
html.

Ranganathan:1998:MS

[RASA98] M. Ranganathan, Laurent Andrey, Virginie Schall, and Anurag Acharya. Mobile streams. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/full_papers/
mccrickard_poster/mccrickard_
html/mccrickard.html.

Ranganathan:1997:NAM

[RASS97] Mudumbai Ranganathan, Anurag Acharya, Shamik Sharma, and Joel Saltz. Network-aware mobile programs. In USE-NIX [USE97e], pages 91–

103. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/ranganathan.html;
http://www.usenix.org/publications/
library/proceedings/ana97/
summaries/ranganathan.html.

Rath:1993:AP

[Rat93] Christopher Rath. About PEM.

;login: the USENIX Association newsletter, 18(2):32–??, March/ April 1993. CODEN LOGNEM. ISSN 1044-6397.

Rath:1994:BSL

[Rat94] Christopher Rath. The BNR standard login (A login configuration manager). In USE-NIX [USE94h], pages 125–

138. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/rath.html.

Rattner:1998:NPB

[Rat98] Justin Rattner. The new power behind Windows NT, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/invited_talks/

rattner. Unpublished keynote address at the 2nd USENIX Windows NT Symposium, 1998 August 3-4, 1998 Seattle, Washington, USA.

Rauch:1999:FDF

[Rau99] Jeremy Rauch. Full disclosure: The future of vulnerability disclosure? ;login: the USENIX Association newsletter, 24(6s): ??, November 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-11/
features/disclosure.html.

Special issue on security.

Rajagopalan:1995:MIM

[RAV95] Ramki Rajagopalan, Sridhar Alagar, and S. Venkatesan. MCE: An integrated mobile computing environment and simulation testbed. In USENIX [USE95i], pages 33–

46. ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995.

URL http://www.usenix.org/
publications/library/proceedings/
mob95/raja.html.

Rawson:1993:ADD

[Raw93] Freeman L. Rawson, III. An architecture for device drivers executing as user-level tasks. In USENIX [USE93h], pages 153–171. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Rangan:1991:TMD

[RBB^{+}91] P. Venkat Rangan, Walter A.

Burkhard, Robert W. Bowdidge, Harrick M. Vin, John W. Lindwall, Kashun Chan, Ingvar A. Aaberg, Linda M. Yamamoto, and Ian G. Harris. A testbed for managing digital video and audio storage. In USENIX [USE91g], pages 199–

208. LCCN QA 76.76 O63 U83 1991.

Reiher:1991:DTW

[RBJ91] Peter L. Reiher, Steven Bellenot, and David Jefferson. Debugging the Time Warp operating system and its application programs. In USENIX [USE91j], pages 203–220. ISBN ???? LCCN ????

Reiher:1993:OR

[RC93] Peter Reiher and David Cohn.

Opening remarks. In USENIX [USE93j], page ?? ISBN 1-880446-54-5. LCCN QA76.9.D5

U84 1993. URL http://www.
usenix.org/publications/library/
proceedings/sedms4/.

Rouse:1995:TCC

[RC95] Forest R. Rouse and Wayne Christopher. A Tcl to C compiler. In USENIX [USE95j], pages 115–122. ISBN 1-880446-72-3. LCCN QA76.73.T44

T44 1995. URL http://www.
usenix.org/publications/library/
proceedings/tcl95/rouse.html.

Rouse:1997:TSO

[RC97] Forest Rouse and Wayne Christopher. A typing system for an optimizing multiplebackend Tcl compiler. In USENIX [USE97d], pages 37–

41. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL ftp.dnai.com/users/i/
icemcfd/tcl;   http://www.
icemcfd.com/tcl/ice.html;
 http://www.usenix.org/publications/
library/proceedings/tcl97/

rouse.html. The backends include C and the Java Virtual Machine. Demonstration licenses for the Tcl Lint system are also available.

Riddle:1995:AAM

[RDM95] Paul Riddle, Paul Danckaert, and Matt Metaferia. AGUS: An automatic multi-platform account generation system. In USENIX [USE95h], pages 171–

180. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/riddle.html.

Reddy:1998:OLN

[Red98] Ian Reddy. Overview of the LISA/NT conference, 1998. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6–11, 1998 Boston, Massachusetts, USA.

Rees:1995:GE

[Ree95] Jim Rees. Guest editorial. In USENIX [USE95a], pages 343– 344.

Reed:1997:RGH

[Ree97]  Nancy E. Reed.  Report

from a Grace Hopper celebrant. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
imho.html.

Reiher:1992:EOS

[Rei92a] Peter L. Reiher. Experiences with optimistic synchronization for distributed operating systems. In USENIX [USE92l], pages 59–76. ISBN 1-880446-41-

3. LCCN QA76.9.D3 S954 1992.

Reiser:1992:SIR

[Rei92b] John F. Reiser. Static initializers: Reducing the value added tax on programs. In USENIX [USE92k], pages 171–180. ISBN 1-880446-45-6. LCCN ????

Reid:1997:WUM

[Rei97]  Jim Reid.  When UNIX

met air traffic control, 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/invited_talks/reid/

reid.pdf;   http://www.usenix.
org/publications/library/
proceedings/lisa97/invited_
talks/reid/reid.ps. Unpub-

lished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Reid:1998:LEC

[Rei98a] Jim Reid. Letter to the Editor: Clarification. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

Reinfelds:1998:TSP

[Rei98b] Juris Reinfelds. The transition from single-platform to multiplatform software development,

1998.  URL http://www.
usenix.org/publications/library/
proceedings/usenix98/invited_
talks/reinfelds.ps. Unpub-

lished invited talk at the USE-NIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Reid:1999:NCD

[Rei99] Jim Reid. New challenges and dangers for the DNS, 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/ITs/reid.pdf. Unpub-

lished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Reynolds:1992:AOA

[Rey92] Franklin Reynolds. An architectural overview of alpha: a real-time, distributed kernel. In USENIX [USE92i], pages 127–

146. ISBN 1-880446-42-1. LCCN QAX 32.

Ramsey:1995:NJM

[RF95] Norman Ramsey and Mary F.

Fernandez. The New Jersey machine-code toolkit. In USENIX [USE95e], pages 289–

302. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/ramsey.html.

Reguero:1998:LSP

[RFD98] Ignacio Reguero, David Foster, and Ivan Deloose. Large scale print spool service. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
reguero.html.

Romig:1999:CFL

[RFR99] Steve Romig, Mark Fullmer, and Suresh Ramachandran. Cisco flow logs and intrusion detection at the Ohio State University. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-9/

osu.html. Special issue on intrusion detection.

Reinhardt:1993:KSW

[RFW93] Steven K. Reinhardt, Babak Falsafi, and David A. Wood.

Kernel support for the Wisconsin Wind Tunnel. In USENIX [USE93k], pages 73–

89. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
micro93/reinhardt.html.

Ramm:1995:ESS

[RG95] Karl Ramm and Michael Grubb. Exu — A system for secure delegation of authority on an insecure network. In USENIX [USE95h], pages 89–94 (or 89– 93??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/ramm.html.

Rubin:1993:LRJ

[RH93] A. D. Rubin and P. Honeyman. Long running jobs in an authenticated environment. In USENIX [USE93m], pages 19–28. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/.

Reid:1998:AFS

[RH98a] Jim Reid and Anton Holleman. ADAPT: a flexible solution for managing the DNS, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/invited_talks/reid.
ps;   http://www.usenix.org/
publications/library/proceedings/
usenix98/invited_talks/reid_

html. Unpublished invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Reid:1998:DNS

[RH98b] Jim Reid and Anton Holleman. Domain Name System: The origin solution, 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/invited_talks/reid.
whitepaper.ps. Unpublished

invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Rhett:1998:RVM

[Rhe98] Joe Rhett. Request v3: a modular, extensible task tracking tool. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/rhett.html.

Record:1998:LES

[RHG98] Ronald Joe Record, Michael Hopkirk, and Steve Ginzburg. Linux emulation for SCO. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/record.ps.

Radia:1995:SOM

[RHKP95] Sanjay Radia, Graham Hamilton, Peter Kessler, and Michael L. Powell. The Spring object model. In USENIX [USE95k], pages 159–172. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/radia.html.

Reiher:1994:RFC

[RHR^{+}94] Peter Reiher, John Heidemann, David Ratner, Gregory Skinner, and Gerald J. Popek. Resolving file conflicts in the Ficus file system. In USE-NIX [USE94j], pages 183–195. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/heidemann.html.

Richter:1997:ADU

[Ric97] Thomas Richter. Another disk usage tool. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Riddle:1994:AUL

[Rid94] Paul Riddle. Automated upgrades in a lab environment. In USENIX [USE94h], pages 33–36. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa94/riddle.html.

Rieken:1997:IJB

[Rie97]  Bill Rieken.  An introduc-

tion to Java Beans. Tutorial session at [USE97g]., 1997.

                                        URL http://dxsting.cern.
                                        ch/sting/COOTS97.html.

Ritchie:1990:KA

[Rit90] Dennis Ritchie. Keynote address. In USENIX [USE90g], page ?? LCCN QA76.8.U65 U81 1990.

Rizzo:1998:DFE

[Riz98] Luigi Rizzo. Dummynet and forward error correction. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/rizzo.ps.

Ranum:1997:CAE

[RK97] Marcus Ranum and Rob Kolstad. A conversation about entrepreneurship between Marcus Ranum and Rob Kolstad. ;login: the USENIX Association newsletter, 22(5):21–24, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Rich:1991:HFD

[RL91] Kenneth Rich and Scott Leadley.

hobgoblin: a file and directory auditor. In USENIX [USE91f], pages 199–208.

Reekie:1998:TSC

[RL98] H. John Reekie and Edward A.

Lee. The Tycho Slate: Complex drawing and editing in Tcl/Tk. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/reekie.
html.

Ringel:1999:ATP

[RL99] Matthew F. Ringel and Thomas A. Limoncelli. Adverse termination procedures -or- “How To Fire A System Administrator”. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/ringel.
html.

Romer:1994:DPM

[RLBC94] Theodore H. Romer, Dennis Lee, Brian N. Bershad, and J. Bradley Chen. Dynamic page mapping policies for cache conflict resolution on standard hardware. In USENIX [USE94i], pages 255–

266. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/index.html.

Ranum:1997:IGT

[RLS^{+}97] Marcus J. Ranum, Kent Landfield, Mike Stolarchuk, Mark Sienkiewicz, Andrew Lambeth, and Eric Wall. Implementing a generalized tool for network monitoring. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997. URL

http://www.usenix.org/publications/
library/proceedings/lisa97/
01.ranum.html.

Reiss:1990:FSC

[RM90] Steven P. Reiss and Scott Meyers. FIELD support for C++. In USENIX [USE90j], pages 293– 299 (or 293–300??).

Rouillard:1994:CMI

[RM94a] John P. Rouillard and Richard B.

Martin. Config: a mechanism for installing and tracking system configurations. In USENIX [USE94h], pages 9–

17. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/rouillard.html.

Rouillard:1994:DLM

[RM94b] John P. Rouillard and Richard B.

Martin. Depot-Lite: a mechanism for managing software. In USENIX [USE94h], pages 83–91. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/martin.html.

Reverbel:1997:MCO

[RM97] Francisco C. R. Reverbel and Arthur B. Maccabe. Making CORBA objects persistent: the object database adapter approach. In USENIX [USE97g], pages 55–65. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/reverbel.html. See

also [LH97, Rie97, WW97].

Russo:1990:COS

[RMC90] Vincent F. Russo, Peter W.

Madany, and Roy H. Campbell. C++ and operating systems performance: a case study. In USENIX [USE90j], pages 103– 114.

Rashid:1991:MA

[RMGB91] Richard Rashid, Gerald Malan, David Golub, and Robert Baron. DOS as a Mach 3.0 application. In USENIX [USE91h], pages 27–40. LCCN QAX 27.

Roussopoulos:1999:PLR

[RMS^{+}99] Mema Roussopoulos, Petros Maniatis, Edward Swierk, Kevin Lai, Guido Appenzeller, and Mary Baker. Person-level routing in the Mobile People Architecture. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/roussopoulos.html.

Ritchie:1993:ULI

[RN93] D. Stuart Ritchie and Gerald W. Neufeld. User level IPC and device management in the Raven kernel. In USE-NIX [USE93k], pages 111–

125. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/ritchie.html.

Rosenblum:1990:LSM

[RO90] Mendel Rosenblum and John K. Ousterhout. The LFS storage manager. In USE-NIX [USE90g], pages 315–324. LCCN QA76.8.U65 U81 1990.

Roberts:1990:WC

[Rob90] William Roberts. Windows column. EUUG Newsletter, 10(1): 72–74, Spring 1990. CODEN EONLE8. ISSN 1011-4211.

Robertson:1992:USC

[Rob92] Bruce Robertson. Using SCSI to control almost anything. In USENIX [USE92h], pages 127–

128. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Robinson:1999:ANB

[Rob99]  David Robinson.  The ad-

vancement of NFS Benchmarking: SFS 2.0. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/robinson.
html.

Rochlis:1997:HVD

[Roc97] Jon A. Rochlis. Happy valentine’s day via the Web. ;login: the USENIX Association newsletter, 22(1):34–35, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Rodriguez:1993:PHI

[Rod93] Eduardo Rodriguez. The persistent hacker: An intruder attacks A new Internet host. In USENIX [USE93m], pages 131–137. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/rodriquez.html.

Roesch:1999:SLI

[Roe99]  Martin Roesch.  Snort —

lightweight intrusion detection for networks. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

                                        URL http://db.usenix.org/
                                        publications/library/proceedings/
                                        lisa99/roesch.html.

Rohnert:1996:POS

[Roh96]  Hans Rohnert.  Pattern-

oriented software architecture. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots96/.

Romig:1990:BOS

[Rom90] Steven M. Romig. Backup at Ohio State, take 2. In USENIX [USE90f], pages 137–142. LCCN QA76.76.O63 L37 1990.

Romig:1991:BOS

[Rom91a] Steve Romig. Backup at Ohio-State, take 2. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Romig:1991:SUC

[Rom91b] Steven M. Romig. Some useful changes for boot RC files. In USENIX [USE91f], pages 245– 248.

Roome:1992:TOF

[Roo92] William D. Roome. 3DFS: a time-oriented file server. In USENIX [USE92j], pages 405– 418.

Rooker:1997:BRI

[Roo97]  Terry Rooker.  Book re-

view: Internet Economics. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
mcknight.html.

Rooker:1998:BRJ

[Roo98a]  Terry Rooker.  Book re-

view: Java Network Security. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-10/
java_sec.html.

Rooker:1998:BRP

[Roo98b]  Terry Rooker.  Book re-

view: Programming Python. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
python.html.

Rooker:1999:BRI

[Roo99] Terry Rooker. Book review: Intranet Security. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
intranet.html.

Rosenthal:1990:EVI

[Ros90] David S. H. Rosenthal. Evolving the Vnode interface. In USE-NIX [USE90g], pages 107–118. LCCN QA76.8.U65 U81 1990.

Rosenberg:1992:AOS

[Ros92] John Rosenberg. Architectural and operating system support for orthogonal persistence. In USENIX [USE92c], pages 305– 335.

Rose:1993:SCR

[Ros93] Greg Rose. Summer ’93 conference report. ;login: the USE-NIX Association newsletter, 18 (5):10–13, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Rose:1994:GB

[Ros94] Greg Rose. The god of backup. ;login: the USENIX Association newsletter, 19(2):28–29, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Rose:1995:PPZ

[Ros95a] Greg Rose. PGP, Phil Zimmerman, Life, the universe, and so on. ;login: the USENIX Association newsletter, 20(2):4–7, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Rose:1995:SUP

[Ros95b] Greg Rose. Status update on PGP and legal actions. ;login: the USENIX Association newsletter, 20(5):9–??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Rose:1995:UTC

[Ros95c] Greg Rose. UNIX tip: I can never remember those pesky options. ;login: the USENIX Association newsletter, 20(4):36– 38, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Roseman:1995:WOO

[Ros95d] Mark Roseman. When is an object not an object? In USENIX [USE95j], pages 197–

204. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/roseman.html.

Rose:1996:IUP

[Ros96a] Greg Rose. Instructions for the USENIX PGP key-signing. ;login: the USENIX Association newsletter, 21(4):10–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Rose:1996:MC

[Ros96b] Greg Rose. A major contributor. ;login: the USENIX Association newsletter, 21(2):46–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Rose:1996:PMI

[Ros96c] Greg Rose. The PGP Moose — implementation and experience. In USENIX [USE96j], pages 155–160. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/ggr.html.

Rose:1996:PZH

[Ros96d] Greg Rose. Phil Zimmermann is off the hook. ;login: the USE-NIX Association newsletter, 21 (2):6–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Rose:1996:UCI

[Ros96e] Greg Rose. Update on cryptography issues. ;login: the USE-NIX Association newsletter, 21 (1):4–??, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Rose:1996:UPKa

[Ros96f] Greg Rose. USENIX PGP keysigning service. ;login: the USE-NIX Association newsletter, 21 (3):10–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Rose:1996:UPKb

[Ros96g] Greg Rose. USENIX PGP keysigning service. ;login: the USE-NIX Association newsletter, 21 (4):8–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Roseman:1996:MCT

[Ros96h] Mark Roseman. Managing complexity in TeamRooms, a Tclbased Internet groupware application. In USENIX [USE96e], pages 1–8. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/roseman.html.

Roseman:1996:RFA

[Ros96i] Mark Roseman. Reports on the Fourth Annual Tcl/Tk Workshop. ;login: the USENIX Association newsletter, 21(5):15– ??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Rose:1997:UUP

[Ros97] Greg Rose. Update on the USE-NIX PGP key signing service. ;login: the USENIX Association newsletter, 22(1):6–??, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Rose:1998:NUP

[Ros98a] Greg Rose. News from the USENIX PGP key signing service. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
pgp.html.

Rose:1998:UPK

[Ros98b] Greg Rose. USENIX PGP key signing service to be discontinued. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
tcpip.html.

Roth:1990:HAF

[Rot90] Jeffrey Roth. Hardening anonymous FTP. In USENIX [USE90i], pages 141–144. LCCN QA 76.9 A25 U55 1990.

Rotaru:1995:RRO

[Rot95] Alexandru Rotaru. ROSE ’94 — the Romanian open systems conference & exhibition report. ;login: the USENIX Association newsletter, 20(1):16–17, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Roth:1999:NMM

[Rot99a] Dave Roth. A networked machine management system. In USENIX [USE99b], pages 11– 26. ISBN 1-880446-30-8. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa-nt99/roth.
html.

Roth:1999:PSP

[Rot99b] Helmut Roth. POSIX.1h SRASS and POSIX.1m checkpoint/ restart. ;login: the USENIX Association newsletter, 24(1): ??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
24.posix1h.html.

Roth:1999:SRP

[Rot99c] Helmut Roth. Standard report: POSIX.1h SRASS and POSIX.1m Checkpoint Restart. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
28.srass.html.

Rouland:1997:DEM

[Rou97a] Chris Rouland. Domain engineering for mission critical environment. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/rouland.html.

Rouland:1997:WLI

[Rou97b] Chris Rouland. WINS in the Lehman infrastructure. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/wips/rouland.
html.

Roy:1993:UFA

[Roy93] Paul J. Roy. UNIX file access and caching in a multicomputer environment. In USE-NIX [USE93h], pages 21–37. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/roy.html.

Rozier:1992:COC

[Roz92] Marc Rozier. Chorus (overview of the Chorus distributed operating system). In USENIX [USE92i], pages 39–70. ISBN 1-880446-42-1. LCCN QAX 32.

Rosenstein:1992:MWC

[RP92] Mark Rosenstein and Ezra Peisach. Mkserv — workstation customization and privatization. In USENIX [USE92e], pages 89–

95. ISBN 1-880446-47-2. LCCN ????

Reichwein:1999:SSI

[RRB99] James Reichwein, Gregg Rothermel, and Margaret Burnett.

Slicing spreadsheets: An integrated methodology for spreadsheet testing and debugging. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663

1999.  URL http://www.
usenix.org/publications/library/
proceedings/dsl99/reichwein.
html.

Raizen:1991:BSL

[RS91] Helen S. Raizen and Stephen C. Shwarm. Building a semi-loosely coupled multiprocessor system based on network process extension. In USENIX [USE91j], pages 51–68. ISBN ???? LCCN ????

Ruane:1990:PSU

[Rua90] Lawrence M. Ruane. Process synchronization in the UTS kernel. In USENIX [USE90c], pages 387–422.

Rubine:1991:IGR

[Rub91]  Dean Rubine.  Integrating

gesture recognition and direct manipulation. In USE-NIX [USE91g], pages 281–299. LCCN QA 76.76 O63 U83 1991.

Rubin:1995:IOT

[Rub95]  Aviel D. Rubin.  Indepen-

dent one-time passwords. In USENIX [USE95f], pages 167–

176. ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
rubin.html.

Rubin:1996:RSU

[Rub96a] Avi Rubin. Reports on the sixth USENIX security symposium. ;login: the USENIX Association newsletter, 21(5):8–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Rubin:1996:IOT

[Rub96b] Aviel D. Rubin. Independent one-time passwords. In USENIX [USE96d], pages 15–27.

Rubin:1998:HUS

[Rub98] Avi Rubin. Highlights from the 7th USENIX security symposium, 1998. URL http://www.

usenix.org/publications/library/
proceedings/usenix98/invited_
talks/security_highlights.

Unpublished invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Rudorfer:1997:MPO

[Rud97]  Gottfried Rudorfer.  Man-

aging PC operating systems with a Revision Control System. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/11.rudorfer.html.

Ruefenacht:1996:RMP

[Rue96]  Craig Ruefenacht.  RUST:

Managing problem reports and to-do lists. In USE-NIX [USE96j], pages 81–

89. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/ruefenac.html.

Russo:1991:PSS

[Rus91] Vincent F. Russo. Process scheduling and synchronization in the renaissance objectoriented multiprocessor operating system. In USENIX [USE91j], pages 117–132. ISBN ???? LCCN ????

Russinovich:1999:WNN

[Rus99a] Mark Russinovich. Windows NT — A new O/S that architecturally isn’t so new, 1999. Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Russo:1999:MAS

[Rus99b] Christopher M. Russo. The magic, art, and science of customer support. ;login: the USE-NIX Association newsletter, 24 (6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/customer.html.

Riedel:1998:PSS

[RvIG98] Erik Riedel, Catharine van Ingen, and Jim Gray. A performance study of sequential I/O on Windows NT^{T M}

4. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/

publications/library/proceedings/
usenix-nt98/riedel.html.

Romer:1997:IOW

[RVL^{+}97] Ted Romer, Geoff Voelker, Dennis Lee, Alec Wolman, Wayne Wong, Hank Levy, Brian Bershad, and J. Bradley Chen. Instrumentation and optimization of Win32/Intel executables using etch. In USENIX [USE97f], pages 1–7. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/romer.html.

Rosenblum:1991:RSA

[RW91] David Rosenblum and Alexander Wolf. Representing semantically analyzed C++ code with Reprise. In USE-NIX [USE91k], pages 119–135. LCCN QA76.73.C153 U27 1991.

Ruemmler:1993:UDA

[RW93] Chris Ruemmler and John Wilkes. UNIX disk access patterns. In USE-NIX [USE93l], pages 405–420. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Ruddock:1997:TIW

[RW97] Rodney Ruddock and Stephen

Walli. telnetd and inetd on

Windows NT with OpenNT. ;login: the USENIX Association newsletter, 22(6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
opennt.html.

Riggs:1996:PSJa

[RWWB96a] Roger Riggs, Jim Waldo, Ann Wollrath, and Krishna Bharat. Pickling state in the Java system. In USE-NIX [USE96a], pages 291–

312. URL http://www.usenix.
org/publications/library/
proceedings/coots96/riggs.
html.

Riggs:1996:PSJb

[RWWB96b] Roger Riggs, Jim Waldo, Ann Wollrath, and Krishna Bharat. Pickling state in the Java system. In USE-NIX [USE96g], pages 261–??, 241–250. ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/riggs.html.

Schiller:1995:SWT

[SA95a] Jeffrey I. Schiller and Derek Atkins. Scaling the web of trust: Combining Kerberos and PGP to provide large scale authentication. In USE-NIX [USE95e], pages 83–94. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/index.html.

Sedayao:1995:LTB

[SA95b] Jeff Sedayao and Kotaro Akita. LACHESIS: a tool for benchmarking Internet service providers. In USENIX [USE95h], pages 111–115. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/sedayao.html.

Smith:1998:TTH

[SA98] Sean W. Smith and Vernon Austel. Trusting trusted hardware: Towards a formal model for programmable secure coprocessors. In USENIX [USE98b], page ?? ISBN 1-880446-97-

9. LCCN HF5004 .U74 1998.

URL http://www.usenix.org/
publications/library/proceedings/
ec98/smith.html.

Sabetella:1990:ISL

[Sab90] Marc Sabetella. Issues in shared libraries design. In USENIX [USE90g], pages 11–24. LCCN QA76.8.U65 U81 1990.

Saell:1998:NES

[Sae98] Jan Saell. News from EurOpen.SE. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
tcpip.html.

Safdar:1992:GCT

[Saf92]  Shabbir J. Safdar.  Giving

customers the tools to protect themselves. In USENIX [USE92m], pages 145–154. ISBN 1-880446-46-4. LCCN ????

Sah:1995:MTC

[Sah95] Adam Sah. Multiple trace composition and its uses. In USENIX [USE95j], pages 15–

24. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/sah.html.

Sakkinen:1992:CCI

[Sak92a] Markku Sakkinen. Corrigendum to “A critique of the inheritance principles of C++”. In USENIX [USE92c], pages 361–363.

Sakkinen:1992:CIP

[Sak92b] Markku Sakkinen. A critique of the inheritance principles of C++. In USENIX [USE92d], pages 69–110.

Salus:1992:Bd

[Sal92a] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 17(5):42–43, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Salus:1992:Be

[Sal92b] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 17(6):40–41, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Salus:1992:Ba

[Sal92c] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 17(2):33–??, March/ April 1992. CODEN LOGNEM. ISSN 1044-6397.

Salus:1992:Bb

[Sal92d] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 17(3):28–??, May/ June 1992. CODEN LOGNEM. ISSN 1044-6397.

Salus:1992:Bc

[Sal92e] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 17(4):48– 49, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Salz:1992:USC

[Sal92f]  Rich Salz.  1992 USENIX

Summer conference BOF report. ;login: the USENIX Association newsletter, 17(4):11– 13, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Salz:1992:IUT

[Sal92g]  Rich Salz.  InterNetNews:

Usenet transport for Internet sites. In USENIX [USE92g], pages 93–98. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Salz:1992:PP

[Sal92h] Rich Salz. A perspective on pathalias. ;login: the USE-NIX Association newsletter, 17 (6):23–25, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:Ba

[Sal93a] Peter Salus. The bookworm.

;login: the USENIX Association newsletter, 18(1):47–48, January/February 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Salus:1993:Bb

[Sal93b] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 18(2):46–47, March/ April 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:Bc

[Sal93c] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 18(3):34–35, May/ June 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:Bd

[Sal93d] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 18(4):38–39, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:Bf

[Sal93e] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 18(6):41–42, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:RWW

[Sal93f] Peter Salus. Review: Words, Words, Words. ;login: the USE-NIX Association newsletter, 18 (6):42–43, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:BRM

[Sal93g] Peter H. Salus. Book review: The Multimedia Maven. ;login: the USENIX Association newsletter, 18(2):48–??, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1993:Be

[Sal93h] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 18(5):46– 47, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Salus:1994:Ba

[Sal94a] Peter Salus. The bookworm.

;login: the USENIX Association newsletter, 19(1):46–47, January/February 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Salus:1994:Bb

[Sal94b] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 19(2):46–47, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Salus:1994:Bc

[Sal94c] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 19(3):31–32, May/ June 1994. CODEN LOGNEM. ISSN 1044-6397.

Salus:1994:A

[Sal94d] Peter H. Salus. Announcements. In USENIX [USE94c], pages 287–??

Salus:1994:Bd

[Sal94e] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 19(4):41–42, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Salus:1994:Be

[Sal94f] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 19(5):47–48, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Salus:1994:Bf

[Sal94g] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 19(6):49–51, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Salus:1995:Be

[Sal95a] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 20(5):44–46, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Salus:1995:Ba

[Sal95b] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 20(1):43–44, February 1995. CODEN LOGNEM. ISSN 1044-6397.

Salus:1995:Bb

[Sal95c] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 20(2):52–53, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Salus:1995:Bc

[Sal95d] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 20(3):45–46, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:CIE

[Sal96a] Emily W. Salus. The cybermeds’ invasion of earth: Patients and physicians online. ;login: the USENIX Association newsletter, 21(4):45–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:LC

[Sal96b] Peter Salus. Language and communication. ;login: the USE-NIX Association newsletter, 21 (5):37–??, October 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Salus:1996:BRB

[Sal96c] Peter H. Salus. Book reviews:

The bookworm. ;login: the USENIX Association newsletter, 21(5):45–48, October 1996. CODEN LOGNEM. ISSN 1044-6397. Review of eight books on Java and JavaScript.

Salus:1996:Ba

[Sal96d] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 21(1):45–46, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:Bb

[Sal96e] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 21(2):56–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:Bc

[Sal96f] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 21(3):53–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:Bd

[Sal96g] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 21(4):52–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:Bf

[Sal96h] Peter H. Salus. The bookworm. ;login: the USENIX Association newsletter, 21(6):49–50, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1996:EI

[Sal96i] Peter H. Salus. Early insecurity. ;login: the USENIX Association newsletter, 21(3):33–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Salus:1997:Bb

[Sal97a] Peter Salus. The bookworm. ;login: the USENIX Association newsletter, 22(2):74–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Salus:1997:RA

[Sal97b] Peter Salus. Report on ACM97. ;login: the USENIX Association newsletter, 22(2):60–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Salus:1997:BRBa

[Sal97c]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/bookworm.html.

Salus:1997:BRBb

[Sal97d]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-12/bookworm.html.

Salus:1997:Ba

[Sal97e] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 22(1):57–58, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Salus:1997:Bd

[Sal97f] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 22(4):65–66, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Salus:1997:Be

[Sal97g] Peter H. Salus. The bookworm.

;login: the USENIX Association newsletter, 22(5):64–65, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Salus:1997:CRY

[Sal97h]  Peter H. Salus.  Confer-

ence report: Are you Tclish? ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/summaries/.

Salus:1997:TYA

[Sal97i] Peter H. Salus. Twenty years ago in UNIX NEWS. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Salus:1998:LEU

[Sal98a] Peter Salus. Letter to the Editor: USENIX events calendar. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Salus:1998:BRBa

[Sal98b]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1998-2/bookworm.html.

Salus:1998:BRBb

[Sal98c]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1998-4/bookworm.html.

Salus:1998:BRBc

[Sal98d] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
bookworm.html.

Salus:1998:BRBd

[Sal98e]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-8/bookworm.html.

Salus:1998:BRBe

[Sal98f]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/bookworm.html.

Salus:1998:BRBf

[Sal98g]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-12/bookworm.html.

Salus:1998:TYAa

[Sal98h] Peter H. Salus. Twenty years ago in ;login:. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-2/
                                        reference.html.

Salus:1998:TYAb

[Sal98i] Peter H. Salus. Twenty years ago in ;login:. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-4/
                                        pgp.html.

Salus:1998:TYAc

[Sal98j] Peter H. Salus. Twenty years ago in ;login:. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-6/
                                        diffie.html.

Salus:1998:TYAd

[Sal98k] Peter H. Salus. Twenty years ago in ;login:. ;login: the USE-NIX Association newsletter, 23 (5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-8/
                                        tcpip.html.

Salus:1998:TYAe

[Sal98l] Peter H. Salus. Twenty years ago in ;login:. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1998-10/
                                        mailing.html.

Salus:1998:TYAf

[Sal98m] Peter H. Salus. Twenty years ago in USENIX. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-12/satan.html.

Salus:1999:YAUa

[Sal99a] Peter H. Salus. 20 years ago in UNIX. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
20yearsago.html.

Salus:1999:YAUb

[Sal99b] Peter H. Salus. 20 years ago in U[SE]NIX. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
20years.html.

Salus:1999:YAUc

[Sal99c] Peter H. Salus. 20 years ago in USENIX. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
20yearsago.html.

Salus:1999:YAUd

[Sal99d] Peter H. Salus. 20 years ago in USENIX. ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
20years.html.

Salus:1999:YAUe

[Sal99e] Peter H. Salus. 20 years ago in USENIX. ;login: the USE-NIX Association newsletter, 24 (6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
20yrs.html.

Salus:1999:YS

[Sal99f] Peter H. Salus. 30 years since 1969. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
30years.html.

Salus:1999:BRBa

[Sal99g]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-2/bookworm.html.

Salus:1999:BRBb

[Sal99h]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-4/bookworm.html.

Salus:1999:BRBc

[Sal99i]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-6/bookworm.html.

Salus:1999:BRBd

[Sal99j]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-8/bookworm.html.

Salus:1999:BRBe

[Sal99k]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 24(5):??, October 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-10/bookworm.html.

Salus:1999:BRBf

[Sal99l]  Peter H. Salus.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-12/bookworm.html.

Salus:1999:MJL

[Sal99m] Peter H. Salus. In memoriam: John Lions. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        lions.html.

Sammons:1995:MPI

[Sam95] Nathaniel Sammons. Multiplatform interrogation and reporting with Rscan. In USE-NIX [USE95h], pages 75–88 (or 75–87??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa95/sammons.html.

Sanger:1990:CDS

[San90]  Colston Sanger.  Call Doc

Strange. EUUG Newsletter, 10 (1):51–60, Spring 1990. CODEN EONLE8. ISSN 1011-4211.

Sanders:1993:IHH

[San93] Tony Sanders. International hypertext is here. ;login: the USE-NIX Association newsletter, 18 (5):33–34, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Sarwono:1999:IMN

[Sar99]  Edhi Sarwono.  Inside the

Microsoft Network (MSN),

                                        1999. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa-nt99/invitedtalks.

html. Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Satdeva:1990:PSI

[Sat90] Bjorn Satdeva. Perspectives and solutions for increasing security in UNIX system administration. In USENIX [USE90i], pages 131–136. LCCN QA 76.9 A25 U55 1990.

Satdeva:1992:APD

[Sat92] Bjorn Satdeva. Audit: a policy driven security checker for a heterogeneous environment. In USENIX [USE92m], pages 185–

202. ISBN 1-880446-46-4. LCCN ????

Satdeva:1993:AF

[Sat93a] Bjorn Satdeva. After the fact.

;login: the USENIX Association newsletter, 18(5):26–28, September/October 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Satdeva:1993:MMO

[Sat93b] Bjorn Satdeva. Methods for maintaining one source tree in a heterogeneous environment. In USENIX [USE93f], pages 57–65. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/satdeva.html.

Satdeva:1993:ORA

[Sat93c] Bjorn Satdeva. Opening remarks and announcements. In USENIX [USE93f], page ??

ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/lisa93/

.

Satdeva:1993:ON

[Sat93d] Bjorn Satdeva. Opinion on NT. ;login: the USENIX Association newsletter, 18(4):25– ??, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Satdeva:1993:SVC

[Sat93e] Bjorn Satdeva. SAGE views: Communication: An important aspect of UNIX system administration. ;login: the USENIX Association newsletter, 18(1): 12–15, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Satdeva:1994:MSA

[Sat94] Bjorn Satdeva. “make” as a system administration tool. In USENIX [USE94m], pages 81–

87. ISBN ???? LCCN ????

Satyanarayanan:1995:RIM

[Sat95] M. Satyanarayanan. Report on the IEEE Mobile Computing Systems and Application Workshop. ;login: the USENIX Association newsletter, 20(2):35–44, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Savage:1999:STB

[Sav99] Stefan Savage. Sting: a TCPbased network measurement tool. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/savage.html.

Smith:1999:ERE

[SAYZ99] Ben Smith, Anurag Acharya, Tao Yang, and Huican Zhu. Exploiting result equivalence in caching dynamic Web content. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

URL http://www.usenix.org/
publications/library/proceedings/
usits99/smith.html.

Schaps:1993:PAN

[SB93] Gary L. Schaps and Peter Bishop. A practical approach to NFS response time monitoring. In USENIX [USE93f], pages 165–169. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/.

Sah:1994:NAI

[SB94] Adam Sah and Jon Blow. A new architecture for the implementation of scripting languages. In USENIX [USE94k], pages 21–38. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/sah.html.

Smaragdakis:1997:DTL

[SB97a] Yannis Smaragdakis and Don Batory. DiSTiL: a transformation library for data structures. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/smaragdakis.html.

Speight:1997:BTG

[SB97b] Evan Speight and John K. Bennett. Brazos: a third generation DSM system. In USENIX [USE97f], pages 95–

106. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/speight.html.

Saito:1998:TMS

[SB98a] Yasushi Saito and Brian Bershad. A transactional memory service in an extensible operating system. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix98/
saito.html.

Sudame:1998:TTF

[SB98b] Pradeep Sudame and B. R. Badrinath. Transformer Tunnels: a framework for providing route-specific adaptations. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/sudame.html.

Sirer:1999:UPG

[SB99] Emin Gün Sirer and Brian N. Bershad. Using production grammars in software testing. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999. URL http://www.

usenix.org/publications/library/
proceedings/dsl99/sirer.html.

Sah:1996:PIS

[SBB96] Adam Sah, Kevin Brown, and Eric Brewer. Programming the Internet from the serverside with Tcl and audience. In USENIX [USE96e], pages 183–188. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/sah.html.

Siegel:1990:DFD

[SBM90] Alex Siegel, Kenneth Birman, and Keith Marzullo. Deceit: a flexible distributed file system. In USENIX [USE90g], pages 51–

62. LCCN QA76.8.U65 U81 1990.

Seltzer:1993:ILS

[SBMS93] Margo Seltzer, Keith Bostic, Marshall Kirk McKusick, and Carl Staelin. An implementation of a log-structured file system for UNIX. In USE-NIX [USE93l], pages 307–326. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/seltzer.pdf. Best pre-

sentation winner.

Solana:1996:ARE

[SBRH96] E. Solana, V. Baggiolini, M. Ramluckun, and J. Harms. Automatic and reliable elimination of E-mail loops based on statistical analysis. In USENIX [USE96j], pages 139–

144. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/solana.html.

Sekar:1999:PIP

[SBS99] R. Sekar, T. Bowen, and M. Segal. On preventing intrusions by process behavior monitoring. In USENIX [USE99l], page ?? ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        detection99/sekar.html.

Sibert:1995:DSP

[SBV95] Olin Sibert, David Bernstein, and David Van Wie. The Digi-Box: a self-protecting container for information commerce. In USENIX [USE95g], pages 171– 183 (or 171–184??). ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1). URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/ec95/sibert.html.

Spencer:1992:ICH

[SC92] Henry Spencer and Geoff Collyer. #ifdef considered harmful or portability experience with C News. In USE-NIX [USE92g], pages 185–198. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

URL https://www.usenix.
org/legacy/publications/library/
proceedings/sa92/spencer.
pdf.

Song:1993:PPB

[SC93] Inshik Song and Yookun Cho.

Page prefetching based on fault history. In USE-NIX [USE93h], pages 203–213. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mach3/.

Stanton:1996:TTP

[SC96] Scott Stanton and Ken Corey.

TclJava: Toward portable extensions. In USENIX [USE96e], pages 25–29. ISBN 1-880446-78-2. LCCN QA 76.73 T44

T35 1996. URL http://www.
usenix.org/publications/library/
proceedings/tcl96/corey.html.

Sorenson:1999:RSS

[SC99] Eric Sorenson and Rose Chalup.

RedAlert: a scalable system for application monitoring. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/sorenson.
html.

Scarr:1997:OLS

[Sca97a] Phil Scarr. Overview of the Large Scale System Administration of Windows NT Workshop, 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/lisa97/invited_

talks/scarr. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Scarr:1997:WWC

[Sca97b] Phil Scarr. When worlds collide. ;login: the USENIX Association newsletter, 22(4):22–23, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Scarr:1998:PDT

[Sca98] Phil Scarr. Panel discussion: TCP/IP futures, 1998. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6– 11, 1998 Boston, Massachusetts, USA.

Stodolsky:1993:FIP

[SCB93] Daniel Stodolsky, J. Brad Chen, and Brian Bershad. Fast interrupt priority management in operating system kernels. In USENIX [USE93k], pages 105–110. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        micro93/stodolsky.html.

Santhanam:1990:EUC

[SCD90] Vatsa Santhanam, Paul Chan, and Manoj Dadoo. Evolution of the Ucode compiler intermediate language. In USE-NIX [USE90g], pages 199–210. LCCN QA76.8.U65 U81 1990.

Swartz:1994:AEH

[SCD94] Karl L. Swartz, Les Cottrell, and Marty Dart. Adventures in the evolution of a high-bandwidth network for central servers. In USE-NIX [USE94h], pages 159–

166. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/swartz.html.

Schmidt:1990:GPH

[Sch90] Douglas C. Schmidt. GPERF: a perfect hash function generator. In USENIX [USE90j], pages 87– 101 (or 87–102??).

Schaefer:1991:NMU

[Sch91] Stephen P. Schaefer. newu:

Multihost user setup. In USE-NIX [USE91e], page ?? ISBN ???? LCCN ????

Schafer:1992:BPC

[Sch92a] Peg Schafer. bbn-public — contributions from the user community. In USENIX [USE92e], pages 211–213. ISBN 1-880446-47-2. LCCN ????

Schafer:1992:CSA

[Sch92b] Peg Schafer. Is centralized system administration the answer? In USENIX [USE92e], pages 55–

61. ISBN 1-880446-47-2. LCCN ????

Schaufler:1992:RWP

[Sch92c] Robin Schaufler. Realtime workstation performance for MIDI.

In USENIX [USE92j], pages 139–152.

Schafer:1993:SB

[Sch93a] Peg Schafer. From the SAGE board. ;login: the USENIX Association newsletter, 18(6):12– ??, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Schafer:1993:SPN

[Sch93b] Peg Schafer. SAGE: a proposed new model. ;login: the USENIX Association newsletter, 18(1): 18–22, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Schimmel:1993:CSM

[Sch93c]  John Schimmel.  A case

study on moves and mergers. In USENIX [USE93f], pages 93–98. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/schimmel.html.

Scherr:1994:RPG

[Sch94a] Andy Scherr. Rapid programming with graph rewrite rules. In USENIX [USE94k], pages 83–100. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/index.html.

Schmidt:1994:AOO

[Sch94b] Douglas Schmidt. ASX: An object-oriented framework for developing distributed applications. In USENIX [USE94e], pages 207–226. ISBN 1-880446-60-X. LCCN QA 76.73 C153

C22 1994. URL http://www.
usenix.org/publications/library/
proceedings/c++94/schmidt.
html.

Schuerr:1994:RPG

[Sch94c] Andy Schürr. Rapid programming with graph rewrite rules. In USENIX [USE94k], pages 83–100. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

URL http://www.usenix.org/
publications/library/proceedings/
vhll/index.html.

Schemers:1995:LLB

[Sch95a] Roland J. Schemers, III. lbnamed: a load balancing name server in Perl. In USENIX [USE95h], pages 1–11 (or 1– 12??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/schemers.html.

Schimmel:1995:RUS

[Sch95b] Curt Schimmel. Review: UNIX systems for modern architectures, symmetric multiprocessing, and caching for kernel programmers. ;login: the USE-NIX Association newsletter, 20 (2):54–55, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Schimmel:1995:WTF

[Sch95c] John Schimmel. What’s up on the Tcl front? ;login: the USE-NIX Association newsletter, 20 (5):20–21, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Schimmel:1995:FNM

[Sch95d] John E. Schimmel. Free network management using Tcl. ;login: the USENIX Association newsletter, 20(4):23–28, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Schimmel:1995:FDR

[Sch95e] John E. Schimmel. Future directions: a report form the LISA ’95 advanced topics workshop. ;login: the USENIX Association newsletter, 20(6):29–30, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Schimmel:1995:WTS

[Sch95f] John E. Schimmel. So what’s this tickle stuff all about? ;login: the USENIX Association newsletter, 20(3):26–27, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Schultzer:1995:EBE

[Sch95g] Daniel Schultzer. Electronic banking and electronic commerce on the SuperInformation highway. In USE-NIX [USE95g], pages 225–

263. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec95/index.html.

Schimmel:1996:CPD

[Sch96a] John Schimmel. Commercial product design in Tcl. ;login: the USENIX Association newsletter, 21(2):34–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Schimmel:1996:MNS

[Sch96b] John Schimmel. Modern name service directions. ;login: the USENIX Association newsletter, 21(3):17–??, June 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Schimmel:1996:MBD

[Sch96c] John Schimmel. Moving big data fast. ;login: the USE-NIX Association newsletter, 21 (4):19–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Schimmel:1996:RSB

[Sch96d] John E. Schimmel. Review: In search of a better Internet language. ;login: the USENIX Association newsletter, 21(5):17– ??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Schimmel:1996:TOA

[Sch96e] John E. Schimmel. Tcl, the other applet language. ;login: the USENIX Association newsletter, 21(1):13–14, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Schmidt:1996:GE

[Sch96f] Douglas C. Schmidt. Guest editorial. In USENIX [USE96a], pages 261–263.

Schmidt:1996:ODP

[Sch96g] Douglas C. Schmidt. OO design patterns for concurrent, parallel, and distributed systems. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

Schmidt:1996:OR

[Sch96h] Douglas C. Schmidt. Opening remarks. In USENIX [USE96g], page ?? ISBN 1-880446-77-

4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

Schneier:1996:CC

[Sch96i] Bruce Schneier. Cryptography in the 21st Century. In USENIX [USE96k], page 352. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/

sd96/. Unpublished invited talk.

Schimmel:1997:HLF

[Sch97a] John Schimmel. A historical look at firewall technologies. ;login: the USENIX Association newsletter, 22(1):21–23, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Schubert:1997:UTR

[Sch97b] Volker Schubert. Using Tk as remote GUI frontend for 4GL-database applications. In USENIX [USE97d], pages 171–

172. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_schubert.html.

Schmidt:1999:AOA

[Sch99] Douglas C. Schmidt. An architectural overview of the ACE framework. ;login: the USE-NIX Association newsletter, 24 (1s):??, January 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-1/

ace.html. Special issue on tools.

Seltzer:1990:DSR

[SCO90] Margo Seltzer, Peter Chen, and John Ousterhout. Disk scheduling revisited. In USE-NIX [USE90h], pages 313–324. LCCN QA76.8.U65 U82 1990.

Scott:1993:PPP

[Sco93] Michael L. Scott. The prospects for parallel programs on distributed systems. In USE-NIX [USE93j], pages 265–

272. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

Scott:1997:ASR

[Sco97] Peter Scott. Automating 24 × 7 support response to telephone requests. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/04.scott.html.

Salus:1993:SCR

[SCP^{+}93] Peter H. Salus, Peter Collinson, Jerry Peek, Henry Spencer, Peg Schafer, and O’Reilly and Inc. Associates. Summer ’93 conference reports. ;login: the USENIX Association newsletter, 18(4):6–15, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Sampson:1999:DMW

[SCW^{+}99] Jason Sampson, Elwood Coslett, Gary Washington, Bob Paauwe, Russ Craft, and Kevin Wheeler. Deployment of Microsoft Windows NT in a design engineering environment. In USE-NIX [USE99b], pages 69–75. ISBN 1-880446-30-8. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa-nt99/sampson.
html.

Simmons:1992:SRL

[SD92] Steve Simmons and Barbara Dyker. SAGE report: 1992 LISA VI conference. ;login: the USENIX Association newsletter, 17(6):8–9, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Sardesai:1998:CWN

[SD98a] Shantanu Sardesai and Partha Dasgupta. Chime: a Windows NT based parallel processing system. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/

usenix-nt98/full_papers/poster_
sardesai/sardesai.pdf;   http:
//www.usenix.org/publications/
library/proceedings/usenix-
nt98/full_papers/poster_sardesai/
sardesai_html/sardesai.html.

Shah:1998:CHP

[SD98b] Alex Shah and Tony Darugar.

Creating high performance Web applications using Tcl, display templates, XML, and database content. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/shah.html.

Sweeney:1996:SXF

[SDH^{+}96] Adam Sweeney, Don Doucette, Wei Hu, Curtis Anderson, Mike Nishimoto, and Geoff Peck. Scalability in the XFS file system. In USE-NIX [USE96k], pages 1–14. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/sweeney.html.

Sheard:1999:DIU

[SeaBP99] Tim Sheard, Zine el-abidine Benaissa, and Emir Pasalic. DSL implementation using staging and monads. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/sheard.html.

Sebes:1991:DTM

[Seb91] Edward John Sebes. Distributed trusted Mach architecture. In USENIX [USE91h], pages 251–

262. LCCN QAX 27.

Sechrest:1995:OGS

[Sec95] John Sechrest. Opinion: Gnu and Solaris. ;login: the USE-NIX Association newsletter, 20 (4):5–??, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Seeley:1990:SLO

[See90]  Donn Seeley.  Shared li-

braries as objects. In USENIX [USE90g], pages 25–38. LCCN QA76.8.U65 U81 1990.

Seiwald:1994:JMR

[Sei94]  Christopher Seiwald.  Jam

— Make(1) redux. In USE-NIX [USE94g], pages 79–

88. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        appdev94/seiwald.html.

Schwartz:1992:CIR

[SEKN92] Michael F. Schwartz, Alan Emtage, Brewster Kahle, and B. Clifford Neuman. A comparison of Internet resource discovery approaches. In USENIX [USE92a], pages 461–493.

Seliger:1990:ECS

[Sel90]  Robert Seliger.  Extending

C++ to support remote procedure call, concurrency, exception handling, and garbage collection. In USENIX [USE90j], pages 241–264.

Sellens:1991:SMC

[Sel91] John Sellens. Software maintenance in a campus environment: The Xhier approach. In USE-NIX [USE91f], pages 21–28.

Sellens:1994:SLT

[Sel94] John Sellens. SPP — low tech, practical, UNIX software portability. In USE-NIX [USE94g], pages 57–

66. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/sellens.html.

Sellens:1995:FFT

[Sel95a] John Sellens. filetsf: a file transfer system based on lpr/lpd. In USENIX [USE95h], pages 195–212. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/sellen.html.

Sellens:1995:S

[Sel95b] John Sellens. The seminar. ;login: the USENIX Association newsletter, 20(5):12–13, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Sellens:1997:AAR

[Sel97a] John Sellens. Adding ‘areas’ to Req. ;login: the USENIX Association newsletter, 22(1):26– 28, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Sellens:1997:R

[Sel97b] John Sellens. On reliability. ;login: the USENIX Association newsletter, 22(5):17–18, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Sellens:1997:RCH

[Sel97c] John Sellens. On reliability — computing hardware. ;login: the USENIX Association newsletter, 22(4):15–17, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Sellens:1997:RNS

[Sel97d]  John Sellens.  On relia-

bility — networks and services. ;login: the USENIX Association newsletter, 22(5): ??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
sellens.html.

Sellens:1997:RSA

[Sel97e]  John Sellens.  On reliabil-

ity — system administration. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
reliability.html.

Sellens:1998:RBR

[Sel98a] John Sellens. On reliability — backups, restores, and recovery? ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
reliability.html.

Sellens:1998:RRR

[Sel98b]  John Sellens.  On reliabil-

ity — restores and recovery. ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
reliability.html.

Sellens:1998:RWA

[Sel98c]  John Sellens.  On relia-

bility — what about yourself? ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
reliability.html.

Sellens:1999:R

[Sel99a]  John Sellens.  On reliabil-

ity. ;login: the USENIX Association newsletter, 24(5s):??, September 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-9/
features/reliability.html.

Special issue on intrusion detection.

Sellens:1999:RSR

[Sel99b]  John Sellens.  On reliabil-

ity: Security and reliability. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-9/
                                        features/reliability.html.

Sellens:1999:RYY

[Sel99c] John Sellens. On reliability: You and your users. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        reliability.html.

Seow:1993:BRG

[Seo93a] Choong Huei Seow. Book review: Guide to Writing DCE Applications. ;login: the USE-NIX Association newsletter, 18 (2):49–51, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Seow:1993:BRU

[Seo93b] Choong Huei Seow. Book review: Understanding DCE. ;login: the USENIX Association newsletter, 18(2):48–49, March/ April 1993. CODEN LOGNEM. ISSN 1044-6397.

Seltzer:1996:DDS

[SESS96] Margo I. Seltzer, Yasuhiro Endo, Christopher Small, and Keith A. Smith. Dealing with disaster: Surviving misbehaved kernel extensions. In USENIX [USE96f], pages 213– 227. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/seltzer.html.

Steves:1996:PST

[SEYG96] Douglas H. Steves, Chris Edmondson-Yurkanan, and Mohamed Gouda. A protocol for secure transactions. In USENIX [USE96h], pages 201–212. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/steves.html.

Stevenson:1997:PLS

[SF97a] Daniel E. Stevenson and Margaret M. Fleck. Programming language support for digitized images or, the monsters in the closet. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

URL http://www.usenix.org/
publications/library/proceedings/
dsl97/stevenson.html.

Sunkara:1997:CSD

[SF97b] Ramu Sunkara and Steve Fanshier. Case studies: Deep ports, 1997. URL http://www.

usenix.org/publications/library/
proceedings/usenix-nt97/technical.

html. Unpublished panel session at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Spinellis:1997:LLS

[SG97a] Diomidis Spinellis and V. Guruprasad. Lightweight languages as software engineering tools. In USENIX [USE97b], page ?? ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        dsl97/spinellis.html.

Stichnoth:1997:CCI

[SG97b] James M. Stichnoth and Thomas Gross. Code composition as an implementation language for compilers. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/dsl97/stichnoth.
                                        html.

Steere:1999:FDP

[SGG^{+}99] David C. Steere, Ashvin Goel, Joshua Gruenberg, Dylan Mc-Namee, Calton Pu, and Jonathan Walpole. A feedback-driven proportion allocator for realrate scheduling. In USENIX [USE99h], page ?? CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

                                        5. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi99/steere.
                                        html.

Staelin:1991:SF

[SGM91] Carl Staelin and Hector Garcia-Molina. Smart filesystems. In USENIX [USE91i], pages 45–52. LCCN QA 76.76 O63 U84 1992.

Sanders:1994:RC

[SH94] Tony Sanders and Bill Hunter. Response: On certification. ;login: the USENIX Association newsletter, 19(1):27–??, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Sarkar:1996:ECC

[SH96] Prasenjit Sarkar and John Hartman. Efficient cooperative caching using hints. In USENIX [USE96f], pages 35–46. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/sarkar.html.

Schulz:1998:ENV

[SH98a] Martin Schulz and Hermann Hellwagner. Extending NT virtual memory by SCI-based hardware DSM. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

usenix.org/publications/library/
proceedings/usenix-nt98/full_
papers/poster_schulz/schulz_
html/schulz.html.

Sullivan:1998:TSM

[SH98b] Mark Sullivan and Andrew Heybey. Tribeca: a system for managing large databases of network traffic. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/sullivan.html.

Shannon:1992:NUM

[Sha92a] Karen Shannon. Nostalgia: the USENET map. ;login: the USE-NIX Association newsletter, 17 (6):26–27, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Sharp:1992:RTT

[Sha92b] James M. Sharp. Request: a tool for training new sys admins and managing old ones. In USENIX [USE92e], pages 69–72. ISBN 1-880446-47-2. LCCN ????

Shao:1997:TCI

[Sha97] Zhong Shao. Typed common intermediate format. In USE-NIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/dsl97/shao.html.

Schmidt:1995:OOC

[SHAS95] Douglas C. Schmidt, Tim Harrison, and Ehab Al-Shaer. Objectoriented components for highspeed network programming. In USENIX [USE95k], pages 21–38. ISBN 1-880446-71-5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/schmidt.html.

Sang:1993:DIS

[ShCR93] Janche Sang, Ke hsiung Chung, and Vernon Rego. Design and implementation of a simulation library using lightweight processes. In USE-NIX [USE93g], pages 33–44. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/sang.html.

Shein:1992:BRB

[She92a] Barry Shein. Book review: 3:16 Bible Texts Illuminated. ;login: the USENIX Association newsletter, 17(1):31–??, January/February 1992. CODEN LOGNEM. ISSN 1044-6397.

Shein:1992:BRI

[She92b] Barry Shein. Book review: Illustrating Computer Documentation: The Art of Presenting Information Graphically on Paper and Online. ;login: the USENIX Association newsletter, 17(1): 32–??, January/February 1992. CODEN LOGNEM. ISSN 1044-6397.

Shein:1993:LYAa

[She93a] Barry Shein. ;login: 50 and 100 years ago. ;login: the USENIX Association newsletter, 18(4): 33–??, July/August 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Shein:1993:LYAb

[She93b] Barry Shein. ;login: 50 and 100 years ago. ;login: the USE-NIX Association newsletter, 18 (5):32–??, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Shein:1993:LYAc

[She93c] Barry Shein. ;login: 50 and 100 years ago. ;login: the USE-NIX Association newsletter, 18 (6):29–??, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Shein:1993:RUS

[She93d] Barry Shein. Review: UNIX System V network programming. ;login: the USENIX Association newsletter, 18(4):40– ??, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Shein:1994:LYAa

[She94a] Barry Shein. ;login: 50 and 100 years ago. ;login: the USENIX Association newsletter, 19(2): 49–??, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Shein:1994:LYAb

[She94b] Barry Shein. ;login: 50 and 100 years ago. ;login: the USE-NIX Association newsletter, 19 (4):45–??, August 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Sheehan:1995:ICO

[She95]  Dean Sheehan.  Interpreted

C++, object oriented Tcl, what next? In USENIX [USE95j], pages 183–196 (or 183–195??). ISBN 1-880446-72-3. LCCN QA76.73.T44

                                        T44 1995. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/tcl95/sheehan.
                                        html.

Shipley:1991:AAD

[Shi91] Carl Shipley. Adding additional database features to the man system. In USENIX [USE91f], pages 135–140.

Shirriff:1997:BDP

[Shi97] Ken Shirriff. Building distributed process management on an object-oriented framework. In USENIX [USE97e], pages 119–131. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/shirriff.html;   http:
//www.usenix.org/publications/
library/proceedings/ana97/
summaries/shirriff.html.

Shibla:1998:KOM

[Shi98] Matt Shibla. K–12 outreach:

MVHS and the student network. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
mvhs.html.

Schuller:1992:PBO

[SHK92] P. Schuller, H. Hartig, and W. E.

Kunhauser. Performance of the BirLix operating system. In USENIX [USE92i], pages 147–

160. ISBN 1-880446-42-1. LCCN QAX 32.

Sims:1993:PCS

[SHM93] David L. Sims, Debra A. Hensgen, and Lantz Moore. Performance of concurrent servers generated automatically from sequential servers. In USE-NIX [USE93j], pages 195–

211. ISBN 1-880446-54-5. LCCN QA76.9.D5 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sedms4/sims.html.

Short:1997:WTM

[Sho97] Rob Short. What a tangled mess! Untangling user-visible complexity in windows systems, 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/usenix-nt97/technical.

html. Unpublished keynote address at the USENIX Windows NT Workshop 1997, August 11-13, 1997, Seattle, Washington, USA.

Shumway:1991:DOI

[Shu91a] Steve Shumway. A distributed operator interaction system. In USENIX [USE91f], pages 97– 104.

Shumway:1991:ILB

[Shu91b] Steve Shumway. Issues in on-line backup. In USENIX [USE91f], pages 81–88.

Sienkiewicz:1996:HLG

[Sie96] Carolyn J. Sienkiewicz. How local groups can make a difference. ;login: the USENIX Association newsletter, 21(4):29–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Sienkiewicz:1997:BRW

[Sie97a] Carolyn Sienkiewicz. Book review: Web Client Programming with Perl. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
wong.html.

Sienkiewicz:1997:BRN

[Sie97b] Carolyn J. Sienkiewicz. Book review: Network planning, procurement, & management. ;login: the USENIX Association newsletter, 22(4):67–68, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Sienkiewicz:1998:BRW

[Sie98a] Carolyn J. Sienkiewicz. Book review: Windows NT in a Nutshell. ;login: the USE-NIX Association newsletter, 23 (4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
pearce.html.

Sienkiewicz:1998:UJ

[Sie98b]  Mark Sienkiewicz.  Using

Java. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/java/usingjava10.
html.

Simmons:1990:LR

[Sim90] Steve Simmons. Life without root. In USENIX [USE90f], pages 89–92. LCCN QA76.76.O63 L37 1990.

Simmons:1991:LR

[Sim91a] Steven C. Simmons. Life without root. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Simmons:1991:RPC

[Sim91b] Steven C. Simmons. Redundant printer configuration. In USE-NIX [USE91f], pages 253–254.

Simonson:1991:SRA

[Sim91c] John Simonson. System resource accounting on UNIX. In USE-NIX [USE91f], pages 61–72.

Simmons:1992:SBRb

[Sim92a] Steve Simmons. SAGE book review: Managing NFS and NIS. ;login: the USENIX Association newsletter, 17(5):18–??, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1992:SBRa

[Sim92b] Steve Simmons. SAGE book reviews: Keeping the link, telecommunication wiring. ;login: the USENIX Association newsletter, 17(4):22–??, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1992:SLS

[Sim92c] Steve Simmons. Simmons’ laws of system administration. ;login: the USENIX Association newsletter, 17(4):24–25, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1993:SBR

[Sim93a] Steve Simmons. SAGE book review: UNIX System Performance Tuning. ;login: the USE-NIX Association newsletter, 18 (1):17–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1993:STS

[Sim93b] Steve Simmons. SAGE: Tools (system administration tools your vendor never told you about: The logbook). ;login: the USENIX Association newsletter, 18(1):11–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1994:RCU

[Sim94a] Steve Simmons. Review: a C user’s guide to ANSI C. ;login: the USENIX Association newsletter, 19(3):33–??, May/ June 1994. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1994:SDI

[Sim94b] Steve Simmons. Software design for installability. In USE-NIX [USE94g], pages 177–

189. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/simmons.html.

Simmons:1996:OMB

[Sim96a] Steve Simmons. 36.8% overhead or money, the bottom line on consulting. ;login: the USE-NIX Association newsletter, 21 (4):31–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Simmons:1996:RII

[Sim96b] Steve Simmons. Review: IPng:

Internet protocol next generation. ;login: the USENIX Association newsletter, 21(4):56– ??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Simmons:1996:SR

[Sim96c] Steve Simmons. Stress relief. ;login: the USENIX Association newsletter, 21(1):16–??, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Simpson:1996:MRC

[Sim96d] Ian Simpson. Modeling the risks and costs of digitally signed certificates in electronic commerce. In USENIX [USE96h], pages 287–297. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Simon:1997:UC

[Sim97] Josh Simon. USENIX conferences. ;login: the USENIX Association newsletter, 22(2):5–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Simpson:1999:IIC

[Sim99] William Allen Simpson. IKE/ ISAKMP considered harmful. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
features/harmful.html.

Singer:1997:TEP

[Sin97a]  Daniel E. Singer.  Tool-

Man edits pipes — interactively! ;login: the USENIX Association newsletter, 22(5):

??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
toolman.html.

Singer:1997:TMP

[Sin97b] Daniel E. Singer. ToolMan meets PatchReport. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-12/toolman.html.

Singer:1997:TVD

[Sin97c] Daniel E. Singer. ToolMan visualized disk usage. ;login: the USENIX Association newsletter, 22(4):27–29, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Singer:1997:TAD

[Sin97d] Daniel E. Singer. ToolMan’s approach to documenting UNIX directories. ;login: the USE-NIX Association newsletter, 22 (5):45–48, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Singer:1998:TFS

[Sin98a]  Daniel E. Singer.  Tool-

man features Steve Kinzler. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
toolman.html.

Singer:1998:TGW

[Sin98b] Daniel E. Singer. Toolman: Generating Web pages with sh and make. ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
toolman.html.

Singer:1998:TSA

[Sin98c]  Daniel E. Singer.  Tool-

man: Sorting and archiving email. ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
toolman.html.

Singer:1998:TUT

[Sin98d] Daniel E. Singer. ToolMan: Upcoming tools; analyzing paths. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
toolman.html.

Singer:1999:TGW

[Sin99a]  Daniel E. Singer.  Tool-

man: Generating web Pages with sh and make, part 2. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/

publications/login/1999-2/
toolman.html.

Singer:1999:TRC

[Sin99b]  Daniel E. Singer.  Tool-

man: Revision control revisited. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/toolman.html.

Stevenson:1995:MUU

[SJ95] J. Mark Stevenson and Daniel P.

Julin. Mach-US: UNIX on generic OS object servers. In USENIX [USE95e], pages 119–

130. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/stevenson.html.

Sjlin:1999:SDS

[Sjö99]  Martin Sjölin.  State-driven

software installation for Windows NT. In USENIX [USE99b], pages 27–36. ISBN 1-880446-30-8. LCCN ????

URL http://db.usenix.org/
publications/library/proceedings/
lisa-nt99/sjolin.html.

Schneier:1996:PPS

[SK96] Bruce Schneier and John Kelsey.

A peer-to-peer software metering system. In USE-NIX [USE96h], pages 279–

286. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/

                                        publications/library/proceedings/
                                        ec96/index.html.

Schneier:1998:CSS

[SK98] Bruce Schneier and John Kelsey. Cryptographic support for secure logs on untrusted machines. In USENIX [USE98i], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec98/schneier.html.

Skahan:1999:MMB

[SK99] Vincent D. Skahan, Jr. and Robert Katz. MJDLM: Majordomo based Distribution List Management. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

                                        URL http://db.usenix.org/
                                        publications/library/proceedings/
                                        lisa99/skahan.html.

Skarra:1994:ULS

[Ska94] Andrea H. Skarra. Using OS locking services to implement a DBMS: An experience report. In USENIX [USE94j], pages 73–86. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/index.html.

Stabell-Kulo:1999:PAM

[SKAM99] Tage Stabell-Kulø, Ronny Arild, and Per Harald Myrvang. Providing authentication to messages signed with a Smart Card in hostile environments. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/stabell.html.

Safonov:1997:ETO

[SKCB97] Alex Safonov, Joseph A. Konstan, John V. Carlis, and Brian Bailey. Extending traces with OAT: An object attri bute trace package for Tcl/Tk. In USENIX [USE97d], pages 81–88. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/safonov.html.

Skeen:1992:IBA

[Ske92] Dale Skeen. An information bus architecture for large-scale, decision-support environments. In USENIX [USE92j], pages 183–196.

Sklower:1991:TBP

[Skl91] Keith Sklower. A tree-based packet routing table for Berkeley UNIX. In USENIX [USE91i], pages 93–104. LCCN QA 76.76 O63 U84 1992.

Satyanarayanan:1993:EDO

[SKM^{+}93] M. Satyanarayanan, James J.

Kistler, Lily B. Mummert, Maria R. Ebling, Puneet Kumar, and Qi Lu. Experience with disconnected operation in a mobile environment. In USENIX [USE93i], pages 11–28. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mobile93/satyanarayanan.html.

Speer:1994:IUK

[SKP94] Steven E. Speer, Rajiv Kumar, and Craig Partridge. Improving UNIX kernel and networking performance using profile based optimization. In USENIX [USE94l], pages 181–

188. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/partridge.html.

Steere:1990:EUL

[SKS90] David C. Steere, James J. Kistler, and M. Satyanarayanan. Efficient user-level file cache management on the Sun vnode interface. In USE-NIX [USE90g], pages 325–332. LCCN QA76.8.U65 U81 1990.

Staelin:1992:CFS

[SKS92] Carl Staelin, John Kohl, and Mike Stonebraker. The Coconut file system — utilizing tapebased robotic storage. In USE-NIX [USE92h], pages 141–142. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Stroustrup:1992:RTT

[SL92] Bjarne Stroustrup and Dmitry Lenkov. Run time type identification for C++. In USENIX [USE92k], pages 313–339. ISBN 1-880446-45-6. LCCN ????

Schonwalder:1993:HKT

[SL93] J. Schönwälder and H. Langendörfer. How to keep track of your network configuration. In USENIX [USE93f], pages 189–193 (or 189–194??). ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993. URL

http://www.usenix.org/publications/
library/proceedings/lisa93/
schonwalder.html.

Scales:1994:DES

[SL94] D. J. Scales and M. S. Lam.

The design and evaluation of a shared object system for distributed memory machines. In USENIX [USE94i], pages 101–114. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/scales.html.

Shonwalder:1995:TEN

[SL95] J. Shönwalder and H. Langendörfer. Tcl extensions for network management applications. In USENIX [USE95j], pages 279–288. ISBN 1-880446-72-3. LCCN QA76.73.T44

T44 1995. URL http://www.
usenix.org/publications/library/
proceedings/tcl95/schoenwalder.
html.

Scales:1996:TFT

[SL96] Daniel J. Scales and Monica S. Lam. Transparent fault tolerance for parallel applications on networks of workstations. In USENIX [USE96k], pages 329–341 (or 329–342??).

ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd96/scales.html.

Slattery:1997:UJ

[Sla97] Terry Slattery. Using Java. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/java/javaindex.
                                        html;   http://www.usenix.
                                        org/publications/java/usingjava7.
                                        html.

Strayer:1994:XTP

[SLC94] W. T. Strayer, M. Lewis, and R. E. Cline, Jr. XTP as a transport protocol for distributed parallel processing. In USENIX [USE94f], pages 91–101. ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        hispeed/strayer.html.

Scott:1990:IIP

[SLM^{+}90] Michael L. Scott, Thomas J. LeBlanc, Brian D. Marsh, Timothy G. Becker, Cezary Dubnicki, Evangelos P. Marcatos, and Neil G. Smithline. Implementation issues for the psyche multiprocessor operating system. In USENIX [USE90d], pages 101– 138.

Salemi:1992:PMS

[SLS92] Charles Salemi, Eric Lund, and Suryakanta Shah. A privilege mechanism for System V Release 4 operating systems. In USENIX [USE92g], pages 235–

242. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Satdeva:1991:FDB

[SM91] Bjorn Satdeva and Paul M. Moriarty. Fdist: a domain based file distribution system for heterogeneous environment. In USENIX [USE91f], pages 109–126.

Shapiro:1992:DAL

[SM92] Marc Shapiro and Mesaac Makpangou. Distributed abstractions, lightweight references. In USENIX [USE92i], pages 263–

268. ISBN 1-880446-42-1. LCCN QAX 32.

Small:1997:RKS

[SM97] Christopher Small and Stephen Manley. A revisitation of kernel synchronization schemes. In USENIX [USE97e], pages 31–41. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/small.html;   http:/
/www.usenix.org/publications/
library/proceedings/ana97/
summaries/small.html.

Skousen:1998:SDS

[SM98a] Alan Skousen and Donald Miller. The Sombrero distributed single address space operating system project. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/full_papers/poster_skousen/
skousen.pdf;   http://www.
usenix.org/publications/library/
proceedings/usenix-nt98/full_
papers/poster_skousen/skousen_
html/skousen.html.

Staelin:1998:MAM

[SM98b] Carl Staelin and Larry McVoy. mhz: Anatomy of a microbenchmark. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63

U88 1998. URL http://www.
usenix.org/publications/library/
proceedings/usenix98/staelin.
html.

Schreder:1999:RSA

[SM99] Willem A. (Vlakkies) Schreüder and Maria Murillo. rat: a secure archiving program with fast retrieval. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

URL http://db.usenix.org/
publications/library/proceedings/
lisa99/schreuder.html.

Smallwood:1990:GTS

[Sma90] Kevin C. Smallwood. Guidelines and tools for software maintenance. In USENIX [USE90f], pages 47–70. LCCN QA76.76.O63 L37 1990.

Smallwood:1991:GTS

[Sma91] Kevin C. Smallwood. Guidelines and tools for software maintenance ina productional environment. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Smallwood:1992:BBR

[Sma92a] Kevin Smallwood. BSDI BOF report. ;login: the USENIX Association newsletter, 17(3):9– ??, May/June 1992. CODEN LOGNEM. ISSN 1044-6397.

Smallwood:1992:SVW

[Sma92b] Kevin Smallwood. SAGE views:

Whither the customer? ;login: the USENIX Association newsletter, 17(5):15–16, September/October 1992. CODEN LOGNEM. ISSN 1044-6397.

Small:1997:TCS

[Sma97] Christopher Small. A tool for constructing safe extensible C++ systems. In USE-NIX [USE97g], pages 175–

184. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/small.html. See also

[LH97, Rie97, WW97].

Stein:1994:SUG

[SMD^{+}94] Lincoln Stein, Andre Marquis, Robert Dredge, Mary Pat Reeve, Mark Daly, Steve Rozen, and Nathan Goodman. Splicing UNIX into a genome mapping laboratory. In USE-NIX [USE94j], pages 221–229. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/stein.html.

Simao:1997:SSG

[SMDP97] Jorge Paulo F. Simão, José A. Legatheaux Martins, Henrique João L. Domingos, and Nuno Manuel R. Preguiça. Supporting synchronous groupware with peer object-groups. In USENIX [USE97g], pages 233–236. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/panel.simao.html.

See also [LH97, Rie97, WW97].

Shaddock:1995:HUW

[SMH95] Michael E. Shaddock, Michael C. Mitchell, and Helen E. Harrison. How to upgrade 1500 workstations on Saturday, and still have time to mow the yard on Sunday. In USENIX [USE95h], pages 59–66 (or 59– 65??). ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/shaddock.html.

Smith:1991:SDL

[Smi91] Jonathan M. Smith. The software design laboratory. In USE-NIX [USE91a], pages 385–404.

Subramanian:1998:IMP

[SMPR98] Indira Subramanian, Cliff Mather, Kurt Peterson, and Balakrishna Raghunath. Implementation of multiple pagesize support in HP-UX. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/subramanian.html.

Santos:1998:ERG

[SMSM98] Michael D. Santos, P. M.

Melliar-Smith, and L. E. Moser. An extensible remote graphical interface for an ATM network simulator. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/santos.
html.

Snoopy:1996:NFP

[Sno96] Snoopy. New fangled phone systems pose new challenges for system administrators. In USENIX [USE96j], pages 241–

249. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/snoopy.html.

Seltzer:1992:LPM

[SO92a] Margo Seltzer and Michael Olson. LIBTP: Portable, modular transactions for UNIX. In USE-NIX [USE92j], pages 9–26.

Shirriff:1992:TDA

[SO92b] Ken Shirriff and John K.

Ousterhout. A trace-driven analysis of name and attribute caching in a distributed system. In USENIX [USE92j], pages 315–332.

Shirriff:1994:SHB

[SO94] Ken Shirriff and John Ousterhout. Sawmill: a high bandwidth logging file system. In USENIX [USE94j], pages 125–

136. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/shirriff.html.

Seltzer:1999:CED

[SO99] Margo I. Seltzer and Michael A. Olson. Challenges in embedded database system administration. In USENIX [USE99k], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        es99/seltzer.html.

Sommer:1997:DN

[Som97]  Steven Sommer.  Dreams

in a nutshell. In USE-NIX [USE97f], pages 81–

86. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt97/sommer.html.

Soques:1997:CUI

[Soq97] Kimbik Soques. Cloning and unattended installation. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

Sosic:1995:DDS

[Sos95] Rok Sosic. The dynascope directing server: Design and implementation. In USENIX [USE95b], pages 107–134.

Sechrest:1991:ULP

[SP91a] Stuart Sechrest and Yoonho Park. User-level physical memory management for Mach. In USENIX [USE91h], pages 189–

200. LCCN QAX 27.

Smith:1991:PMC

[SP91b] Thomas G. Aguierre Smith and Natalio C. Pincever. Parsing movies in context. In USE-NIX [USE91g], pages 157–168. LCCN QA 76.76 O63 U83 1991.

Spellvin:1994:PC

[SP94] George Spellvin and Hal Pomeranz. Point/counterpoint. ;login: the USENIX Association newsletter, 19(6):21–22, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Stevens:1995:P

[SP95] W. Richard Stevens and Jan-Simon Pendry. Portals in 4.4BSD. In USENIX [USE95e], pages 1–10. ISBN 1-880446-67-7. LCCN QA 76.76 O63

U88 1995. URL http://www.
usenix.org/publications/library/
proceedings/neworl/stevens.
html.

Smith:1996:OES

[SP96] Sean Smith and Paul Pedersen. Organizing electronic services into security taxonomies. In USENIX [USE96h], pages 233–241. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/smith.html.

Salemme:1999:CWS

[SP99a] Anne Salemme and Jag Patel. A couple of Web servers, a small staff, thousands of users, and millions of Web pages… how we manage (sort of),

                                        1999. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa99/practicum/

salemme.pdf. Unpublished practicum talk at LISA XIII.

Spatscheck:1999:DAD

[SP99b] Oliver Spatscheck and Larry L. Peterson. Defending against denial of service attacks in Scout. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

                                        5. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi99/spatscheck.
                                        html.

Spafford:1992:ORP

[Spa92] Eugene Spafford. Observations on reusable password choices. In USENIX [USE92m], pages 299– 312. ISBN 1-880446-46-4. LCCN ????

Spencer:1991:AMS

[Spe91] Henry Spencer. Awk as a major systems programming language. In USENIX [USE91i], pages 137–144. LCCN QA 76.76 O63 U84 1992.

Spencer:1992:FSF

[Spe92] Henry Spencer. Faster string functions. In USENIX [USE92j], pages 419–428.

Spencer:1993:RNR

[Spe93a] Henry Spencer. Review: Numerical Recipes in FORTRAN. ;login: the USENIX Association newsletter, 18(3):36–37, May/ June 1993. CODEN LOGNEM. ISSN 1044-6397.

Spencer:1993:TCC

[Spe93b] Henry Spencer. The Ten Commandments for C programmers (annotated edition). ;login: the USENIX Association newsletter, 18(2):29–31, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Spencer:1995:BRT

[Spe95a] Henry Spencer. Book review:

Tcl and the Tk Toolkit. ;login: the USENIX Association newsletter, 20(5):47–48, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Spencer:1995:BRC

[Spe95b] Henry Spencer. Book review:

The Cartoon Guide to Statistics. ;login: the USENIX Association newsletter, 20(5):48–??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Spencer:1996:SMH

[Spe96] Henry Spencer. Shuse: Multihost account administration. In USENIX [USE96j], pages 25–32. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/henry.html.

Spencer:1997:BRS

[Spe97a] Henry Spencer. Book review: Software solutions in C. ;login: the USENIX Association newsletter, 22(5):69–??, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Spencer:1997:STM

[Spe97b] Henry Spencer. Shuse at two: Multi-host account administration. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/09.spencer.html.

Spencer:1997:SMS

[Spe97c] Henry Spencer. Simple multilingual support for Tcl. In USENIX [USE97d], pages 67–

72. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/spencer.html.

Spector:19xx:CTT

[Spexx] Alfred Spector. Critical technologies for transaction systems. In ???? USENIX, Berkeley, CA, USA, 19xx. Listed in contents only, no paper, no abstract.

Spinellis:1994:WUT

[Spi94] Diomidis Spinellis. Wux: UNIX tools under Windows. In USENIX [USE94l], pages 325–

336. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/

publications/library/proceedings/
sf94/spin.html.

Spinellis:1997:CAP

[Spi97] Diomidis D. Spinellis. A conversation about Perl and the shell: Choosing the implementation vehicle. ;login: the USE-NIX Association newsletter, 22 (5):25–31, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Spinellis:1998:STA

[Spi98]  Diomidis Spinellis.  Small

tools for automatic text generation. ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-8/
tools.html.

Spinellis:1999:OPA

[Spi99]  Diomidis Spinellis.  Opti-

mal peripheral access using pipe-based double-buffering. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
features/peripheral.html.

Safonov:1996:LNV

[SPK^{+}96] Alex Safonov, Douglas Perrin, Joseph Konstan, John Carlis, and Robert Elde. Lessons from the neighborhood viewer: Building innovative collaborative applications in Tcl and Tk. In USENIX [USE96e], pages 203–213. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/safonov.html.

Segev:1996:FEI

[SPR96a] Arie Segev, Jaana Porra, and Malu Roldan. Financial EDI over the Internet: Case study II. In USENIX [USE96h], pages 173–190. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/roldan.html.

Spruce:1996:HPG

[Spr96b] Darren Spruce. High performance graphic display with Tcl/ Tk. In USENIX [USE96e], pages 215–220. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/spruce.html.

Saxena:1993:PMS

[SPY^{+}93] Sunil Saxena, J. Kent Peacock, Fred Yang, Vijaya Verma, and Mohan Krishnan. Pitfalls in multithreading SVR4 STREAMS and other weightless processes. In USE-NIX [USE93l], pages 85–96. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Srinidhi:1993:MDR

[Sri93] Hassan N. Srinidhi. Managing data redundancy in interoperable heterogeneous environments. In USENIX [USE93c], pages 285–317.

Stein:1992:ILT

[SS92] D. Stein and D. Shah. Implementing lightweight threads. In USENIX [USE92g], pages 1–10. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Shopiro:1994:UCC

[SS94a] Susan E. Waggoner Shopiro and Jonathan Shopiro. USENIX C++ conference report. ;login: the USENIX Association newsletter, 19(5):23–25, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Spasojevic:1994:UPE

[SS94b] Mirjana Spasojevic and M. Satyanarayanan. A usage profile and evaluation of a widearea distributed file system. In USENIX [USE94l], pages 307–323. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/spasojevic.html.

Swartz:1995:RRI

[SS95] Jonathan Swartz and Brian C.

Smith. RIVL: a resolution independent video language. In USENIX [USE95j], pages 235–242. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/swartz.html.

Small:1996:CET

[SS96a] Christopher Small and Margo Seltzer. A comparison of OS extension technologies. In USE-NIX [USE96k], pages 41–54. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/small.html.

Smith:1996:CFD

[SS96b] Keith A. Smith and Margo Seltzer. A comparison of FFS disk allocation policies. In USE-NIX [USE96k], pages 15–25 (or 15–26??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/smith.html.

Souto:1997:DSM

[SS97] Pedro A. Souto and Eugene W. Stark. A distributed shared memory facility for FreeBSD. In USENIX [USE97e], pages 149–162. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/souto.html;   http:/
/www.usenix.org/publications/
library/proceedings/ana97/
summaries/souto.html.

Schneier:1999:BHD

[SS99] Bruce Schneier and Adam Shostack. Breaking up is hard to do: Modeling security threats for Smart Cards. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

http://www.usenix.org/publications/
library/proceedings/smartcard99/
schneier.html.

Shaver:1992:ILF

[SSB92] Dave Shaver, Eric Schnoebelen, and George Bier. An implementation of large files for BSD UNIX. In USENIX [USE92j], pages 61–68.

Seltzer:1995:FSL

[SSB^{+}95] Margo Seltzer, Keith A. Smith, Hari Balakrishnan, Jacqueline Chang, Sara McMains, and Venkata Padmanabhan. File system logging versus clustering: a performance comparison. In USENIX [USE95e], pages 249–264. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/seltzer.html.

Srouji:1998:TCF

[SSBK98] Johny Srouji, Paul Schuster, Maury Bach, and Yulik Kuzmin. A transparent checkpoint facility on NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/srouji.html;
http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/srouji_slides.

Swanson:1993:DSW

[SSCK93] Mark Swanson, Leigh Stoller, Terence Critchlow, and Robert Kessler. The design of the schizophrenic workstation system. In USENIX [USE93h], pages 291–306. ISBN 1-880446-49-9. LCCN QA 76.76 O63

U85 1993. URL http://www.
usenix.org/publications/library/
proceedings/mach3/.

Safford:1993:TSP

[SSH93a] Dave Safford, Douglas Lee Schales, and David K. Hess. The TAMU security package: An ongoing response to Internet intruders in an academic environment. In USE-NIX [USE93m], pages 91–

118. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/tamu.saf.html.

Safford:1993:SRA

[SSH93b] David R. Safford, Douglas Lee Schales, and David K. Hess. Secure RPC authentication (SRA) for TELNET and FTP. In USENIX [USE93m], pages 63–67. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sec4/rpc.saf.html.

Safford:1996:TUA

[SSH96] David Safford, Douglas Schales, and David Hess. Texas A&M University anarchistic key authorization (AKA). In USE-NIX [USE96i], pages 179–

185. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/safford.html.

Seshan:1997:SSP

[SSK97] Srinivasan Seshan, Mark Stemm, and Randy H. Katz. SPAND: Shared Passive Network Performance Discovery. In USE-NIX [USE97h], pages 135–146. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/seshan.html.

Skinner:1996:TTD

[SSL96] John M. Skinner, Robert M.

Sweet, and Richard S. LaBarca. Tcl/Tk in the development of user-extensible graphical user interfaces. In USE-NIX [USE96e], pages 83–89. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/skinner.html.

Spencer:1999:FSA

[SSL^{+}99] Ray Spencer, Stephen Smalley, Peter Loscocco, Mike Hibler, Dave Andersen, and Jay Lepreau. The Flask Security Architecture: System support for diverse security policies. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/spencer.
html.

Silverstein:1990:USM

[SST90] Joel D. Silverstein, Steven E. Sommars, and Yio-Chian Tao. The UNIX system math library, a status report. In USENIX [USE90h], pages 117–132 (or 117–131??). LCCN QA76.8.U65 U82 1990.

Snapp:1992:DDI

[SST92] Steven R. Snapp, Stephen E. Smaha, and Daniel M. Teal. The DIDS (distributed intrusion detection system) prototype. In USENIX [USE92g], pages 227–

234. ISBN 1-880446-44-8. LCCN

QA 76.76 O63 U83 1992.

Sousa:1993:DPI

[SSZ^{+}93] Pedro Sousa, Manuel Sequeira, Andre Zuquete, Paulo Ferreira, Cristina Lopes, Jose Pereira, Paulo Guedes, and Jose Alves Marques. Distribution and persistence in the IK platform: Overview and evaluation. In USENIX [USE93a], pages 391– 424.

Sarachan:1995:PNG

[SSZ95] Brion D. Sarachan, Alexandra J. Schmidt, and Steven A. Zahner. Prototyping NBC’s GEnesis broadcast automation system using Tcl/Tk. In USENIX [USE95j], pages 251–

260. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/sarachan.html.

Savage:1993:RTM

[ST93] Stefan Savage and Hideyuki Tokuda. Real time — Mach timers: Exporting time to the user. In USENIX [USE93h], pages 111–118. ISBN 1-880446-49-9. LCCN QA 76.76 O63

U85 1993. URL http://www.
usenix.org/publications/library/
proceedings/mach3/.

Su:1996:BBA

[ST96] Jiawen Su and J. D. Tygar. Building blocks for atomicity in electronic commerce. In USENIX [USE96i], pages 97–103. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/tygar.html.

Staelin:1998:MSP

[Sta98]  Carl Staelin.  mkpkg: a

software packaging tool. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
staelin.html.

Stewart:1992:SCK

[Ste92] John N. Stewart. SunOS, C2 and Kerberos — A comparative review. In USENIX [USE92m], pages 265–284. ISBN 1-880446-46-4. LCCN ????

Sterling:1993:WSD

[Ste93a]  Nicholas Sterling.  WAR-

LOCK — A static data race analysis tool. In USE-NIX [USE93l], pages 97–106. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Stewart:1993:ABA

[Ste93b] John N. Stewart. AMD — the Berkeley automounter, part 1. ;login: the USENIX Association newsletter, 18(3):19–??, May/ June 1993. CODEN LOGNEM. ISSN 1044-6397.

Stevens:1995:RCN

[Ste95]  W. Richard Stevens.  Re-

view: Casting the Net; From ARPANET to Internet and Beyond. ;login: the USENIX Association newsletter, 20(3):47–??, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Sterne:1999:SAC

[STM^{+}99] Daniel F. Sterne, Gregg W. Tally, C. Durward McDonell, David L. Sherman, David L. Sames, Pierre X. Pasturel, and E. John Sebes. Scalable access control for distributed object systems. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/sterne.
html.

Stolarchuk:1992:FAS

[Sto92] Michael Stolarchuk. Faster AFS: a stacked Vnode implementation. In USENIX [USE92h], pages 137–138. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Stolarchuk:1993:FA

[Sto93a] Michael T. Stolarchuk. Faster AFS. In USENIX [USE93l], pages 67–75. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/library/
proceedings/sd93/.

Stolfa:1993:SSA

[Sto93b]  Roland J. Stolfa.  Simpli-

fying system administration tasks: The UAMS approach. In USENIX [USE93f], pages 203–208. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/stolfa.html.

Stoughton:1993:UURd

[Sto93c] Nicholas Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(4):34– 37, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1993:UURe

[Sto93d] Nicholas Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(5):36– 45, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1993:UURc

[Sto93e] Nicholas M. Stoughton. An update on UNIX-related standards activities. ;login: the USE-NIX Association newsletter, 18 (3):25–33, May/June 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1993:UURf

[Sto93f] Nicholas M. Stoughton. An update on UNIX-related standards activities. ;login: the USE-NIX Association newsletter, 18 (6):39–40, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1994:UURb

[Sto94a] Nicholas Stoughton. An update on UNIX-related activities. ;login: the USENIX Association newsletter, 19(2):34–43, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1994:UURa

[Sto94b] Nicholas Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 19(1): 39–45, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1994:UURc

[Sto94c] Nick Stoughton. An update in UNIX-related standards activities. ;login: the USENIX Association newsletter, 19(3):27– 30, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1994:USRa

[Sto94d] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 19 (4):37–40, August 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1994:USRb

[Sto94e] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 19 (5):43–46, October 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1994:USRc

[Sto94f] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 19 (6):44–48, December 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1995:USRe

[Sto95a] Nicholas Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 20 (5):38–42, October 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1995:CN

[Sto95b] Nick Stoughton. Community news. ;login: the USENIX Association newsletter, 20(6):25– ??, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1995:USRa

[Sto95c] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 20 (1):38–42, February 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1995:USRb

[Sto95d] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 20 (2):47–51, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1995:USRc

[Sto95e] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 20 (3):40–44, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Stokely:1996:BVK

[Sto96a] Celeste Stokely. Breadth of vision — A key to successful consulting. ;login: the USE-NIX Association newsletter, 21 (4):36–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Storm:1996:IIS

[Sto96b] Kaye Storm. Industry initiatives for science and math education (IISME). ;login: the USE-NIX Association newsletter, 21 (1):31–32, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1996:USRc

[Sto96c] Nicholas M. Stoughton. An update on standards relevant to USENIX. ;login: the USE-NIX Association newsletter, 21 (3):45–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1996:USRd

[Sto96d] Nicholas M. Stoughton. An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 21(4):50–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1996:SWE

[Sto96e] Nick Stoughton. Are standards worth the effort? ;login: the USENIX Association newsletter, 21(5):29–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1996:USRa

[Sto96f] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 21 (1):33–40, February 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Stoughton:1996:USRb

[Sto96g] Nick Stoughton. An update on standards relevant to USE-NIX members. ;login: the USE-NIX Association newsletter, 21 (2):50–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1997:SRUa

[Sto97a] Nicholas M. Stoughton. Standards report: An update on standards relevant to USENIX members. ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/

.

Stoughton:1997:SRUb

[Sto97b] Nicholas M. Stoughton. Standards report: An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/

.

Stoughton:1997:USRa

[Sto97c] Nicholas M. Stoughton. An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 22(1):49–56, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1997:USRb

[Sto97d] Nicholas M. Stoughton. An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 22(2):64–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1997:USRdd

[Sto97e] Nicholas M. Stoughton. An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 22(4):61–64, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1997:USRe

[Sto97f] Nicholas M. Stoughton. An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 22(5):54–63, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Stoughton:1998:SRH

[Sto98a] Nicholas M. Stoughton. Standard report: Happy birthday, POSIX. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
16.editorial.html.

Stoughton:1998:SRUa

[Sto98b] Nicholas M. Stoughton. Standards report: An update on standards relevant to USENIX members. ;login: the USENIX Association newsletter, 23(1): ??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards.

Stoughton:1998:SRUb

[Sto98c] Nicholas M. Stoughton. Standards report: An update on standards relevant to USENIX members. ;login: the USE-NIX Association newsletter, 23 (2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards.

Stoughton:1998:SRUc

[Sto98d] Nicholas M. Stoughton. Standards report: An update on standards relevant to USENIX members. ;login: the USE-NIX Association newsletter, 23 (6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
21.posix.html.

Stoughton:1998:SRUd

[Sto98e] Nicholas M. Stoughton. Standards report: An update on standards relevant to USE-NIX members (A programmer’s overview of the POSIX.1d Draft Standard: Additional Realtime Extensions). ;login: the USE-NIX Association newsletter, 23 (7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
22.posix.html.

Stoughton:1999:SRP

[Sto99a] Nicholas M. Stoughton. Standard report: POSIX Update. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
27.austin.html.

Stoughton:1999:SRO

[Sto99b] Nicholas M. Stoughton. Standard report: Open source — A standards success story?

;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/standards/
25.opensource.html.

Strack:1990:EAC

[Str90] Hermann Strack. Extended access controls in UNIX System V: ACLs and context. In USENIX [USE90i], pages 87–102. LCCN QA 76.9 A25 U55 1990.

Stroustrup:1993:EC

[Str93a] Bjarne Stroustrup. The evolution of C++: 1985 to 1989. In Waldo [Wal93b], pages 13–

52. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Stroustrup:1993:PDC

[Str93b]  Bjarne Stroustrup.  Possi-

ble directions for C++. In Waldo [Wal93b], pages 53–73. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Stradtman:1999:NDC

[Str99] Chris Stradtman. NT without domain controllers, 1999.

URL http://db.usenix.org/
publications/library/proceedings/
lisa99/invitedtalks.html.

Unpublished invited talk at LISA XIII.

Spezzano:1990:CPS

[STV90] G. Spezzano, D. Talia, and M. Vanneschi. A concurrent programming support for distributed systems. In USENIX [USE90c], pages 423–448.

Schilit:1993:CMA

[STW93a] Bill N. Schilit, Marvin M. Theimer, and Brent B. Welch. Customizing mobile applications. In USENIX [USE93i], pages 129–138. ISBN 1-880446-51-0. LCCN QA 76.76 O63

U86 1993. URL http://www.
usenix.org/publications/library/
proceedings/mobile93/schilit.
html.

Schmidtmann:1993:DIM

[STW93b] Carl Schmidtmann, Michael Tao, and Steven Watt. Design and implementation of a multi-threaded Xlib. In USE-NIX [USE93l], pages 193–203. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Sekar:1999:SFI

[SU99] R. Sekar and P. Uppuluri. Synthesizing fast intrusion prevention/detection systems from high-level specifications. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

http://db.usenix.org/publications/
library/proceedings/sec99/
sekar.html.

Subramanian:1991:MDP

[Sub91] Indira Subramanian. Managing discardable pages with an external pager. In USENIX [USE91h], pages 77–86. LCCN QAX 27.

Subbarao:1998:BRF

[Sub98] Kartik Subbarao. Book review: First Things First. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
first.html.

Suggs:1992:SSA

[Sug92] Darrell Suggs. Secure superuser access via the Internet. In USENIX [USE92m], pages 203–

210. ISBN 1-880446-46-4. LCCN ????

Stumm:1992:DSO

[SUK92] Michael Stumm, Ronald Unrau, and Orran Krieger. Designing a scalable operating system for shared memory multiprocessor. In USENIX [USE92i], pages 285–303. ISBN 1-880446-42-1. LCCN QAX 32.

Summit:1994:FDU

[Sum94] Steve Summit. Filesystem daemons as unifying mechanism for network information access. In USENIX [USE94l], pages 63–77. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/summit.html.

Sundaresan:1997:EST

[Sun97] Neelakantan Sundaresan. Extending the standard template library for parallelism in Coir<Futures>. In USE-NIX [USE97g], pages 159–

173. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/sundaresan.html. See

also [LH97, Rie97, WW97].

Szymaszek:1998:BSE

[SUZ98] Jakub Szymaszek, Andrzej Uszok, and Krzysztof Zielinski. Building a scalable and efficient component oriented system using CORBA — active badge system case study. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/szymaszek.html.

Shaefer:1990:NMH

[SV90a] Stephen P. Shaefer and Satyanarayana R. Vemulakonda. newu: Multi-host user setup. In USENIX [USE90f], pages 23–26. LCCN QA76.76.O63 L37 1990.

Stolfa:1990:UUD

[SV90b] Roland J. Stolfa and Mark J. Vasoll. UDB — user data base system. In USENIX [USE90f], pages 11–16. LCCN QA76.76.O63 L37 1990.

Stolfa:1991:UUD

[SV91] Roland J. Stolfa and Mark J. Vasoll. UDB — user data base system. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Scharf:1994:STM

[SV94] Jerry Scharf and Paul Vixie.

SENDS: a tool for managing domain naming and electronic mail in a large organization. In USENIX [USE94h], pages 93–103. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/scharf.html.

Szczur:1990:TPT

[SW90] Martha Szczur and Karl R.

Wolf. TAE plus: Transportable application environment plus a user interface development tool for building graphic oriented applications. In USE-NIX [USE90h], pages 363–373. LCCN QA76.8.U65 U82 1990.

Shipley:1991:MAL

[SW91a] Carl Shipley and Chingyow Wang. Monitoring activity on a large UNIX network with Perl and Syslogd. In USENIX [USE91f], pages 209–216.

Szabo:1991:DCJ

[SW91b] Bernard I. Szabo and Gregory K. Wallace. Design considerations for JPEG video and synchronized audio in a UNIX workstation environment. In USENIX [USE91g], pages 353–

368. LCCN QA 76.76 O63 U83 1991.

Schauer:1992:IG

[SW92a] Herve Schauer and Christophe Wolfhugel. An Internet gatekeeper. In USENIX [USE92m], pages 49–61. ISBN 1-880446-46-

4. LCCN ????

Stern:1992:NPN

[SW92b] Hal L. Stern and Brian L. Wong. NFS performance and network loading. In USENIX [USE92e], pages 33–38. ISBN 1-880446-47-

2. LCCN ????

Skinner:1993:SVP

[SW93] Glenn C. Skinner and Thomas K. Wong. “stacking” vnodes: a progress report. In USE-NIX [USE93g], pages 161–174. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/skinner.html.

Stajano:1995:TCD

[SW95] Frank Stajano and Rob Walker. Taming the complexity of distributed multimedia applications. In USENIX [USE95j], pages 213–228. ISBN 1-880446-72-3. LCCN QA76.73.T44

T44 1995. URL http://www.
usenix.org/publications/library/
proceedings/tcl95/stajano.
html.

Savage:1996:AFR

[SW96] Stefan Savage and John Wilkes. AFRAID — A frequently redundant array of independent disks. In USENIX [USE96k], pages 27–39. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/wilkes.html.

Santos:1998:IEL

[SW98] Jonathan Santos and David Wetherall. Increasing effective link bandwidth by suppressing replicated data. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998. URL

http://www.usenix.org/publications/
library/proceedings/usenix98/
santos.html.

Swartz:1992:ORI

[Swa92] Karl L. Swartz. Optimal routing of IP packets to multi-homed servers. In USENIX [USE92e], pages 9–16. ISBN 1-880446-47-

2. LCCN ????

Swartz:1993:FDR

[Swa93] Karl L. Swartz. Forecasting disk resource requirements for a Usenet server. In USE-NIX [USE93f], pages 195–

202. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/swartz.html.

Swartz:1996:BLT

[Swa96] Karl L. Swartz. The brave little toaster meets Usenet. In USENIX [USE96j], pages 161–170. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/kls.html.

Swartz:1997:ARTa

[Swa97a] Karl L. Swartz. Adding response time measurement of CIFS file server performance to Net-Bench. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/swartz.html.

Swartz:1997:ARTb

[Swa97b]  Karl L. Swartz.  Adding

response time measurement of CIFS file server performance to NetBench. In USENIX [USE97f], pages 87–

94. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/swartz.html;
 http://www.usenix.org/publications/
library/proceedings/nt-sysadmin97/
swartz.ppt.

Stiemerling:1992:IDT

[SWS92] T. Stiemerling, T. Wilkinson, and A. Saulsbury. Implementing DVSM on the TOPSY multicomputer. In USENIX [USE92l], pages 263–278. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Seltzer:1991:NHP

[SY91] Margo Seltzer and Ozan Yigit. A new hashing package for UNIX. In USENIX [USE91i], pages 173–184. LCCN QA 76.76 O63 U84 1992.

Szeto:1999:ASA

[Sze99] Bailey Szeto. Applying sendmail anti-spam checks at the user level. ;login: the USENIX Association newsletter, 24(3): ??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
features/antispam.html.

Thompson:1992:SPN

[TA92] J. David Thompson and Kate Arndt. A secure public network access mechanism. In USENIX [USE92m], pages 227–238. ISBN 1-880446-46-4. LCCN ????

Taber:1994:CCL

[Tab94] Dave Taber. Computers could be like autos. ;login: the USE-NIX Association newsletter, 19 (1):45–??, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Talbot:1995:TAO

[Tal95] Jacques Talbot. Turning the AIX operating system into an MP-capable OS. In USENIX [USE95e], pages 143–154 (or 143–153??). ISBN 1-880446-67-7. LCCN QA 76.76 O63

U88 1995. URL http://www.
usenix.org/publications/library/
proceedings/neworl/talbot.
html.

Tang:1995:SPM

[Tan95] Lei Tang. A set of protocols for micropayments in distributed systems. In USENIX [USE95g], pages 107–116 (or 107–115??). ISBN 1-880446-74-X. LCCN HF5548.33. U84

1995(1). URL http://www.
usenix.org/publications/library/
proceedings/ec95/tang.html.

Tang:1996:FBE

[Tan96] Lei Tang. A framework for building an electronic currency system. In USE-NIX [USE96i], pages 113–

122. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec96/.

Tardo:1990:PKB

[TAP90] Joe Tardo, Kannan Alagappan, and Richard Pitkin. Public-key-based authentication using Internet certificates. In USENIX [USE90i], pages 121–124. LCCN QA 76.9 A25 U55 1990.

Taplin:1999:CMP

[Tap99] Jeremy Bargenand Seth Taplin. Cro-Magnon: a patch hunter-gatherer. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN ????

                                        URL http://db.usenix.org/
                                        publications/library/proceedings/
                                        lisa99/bargen.html.

Taunton:1991:CEE

[Tau91] Mark Taunton. Compressed executables: An exercise in thinking small. In USENIX [USE91g], pages 385–404. LCCN QA 76.76 O63 U83 1991.

Taylor:1995:WDN

[Tay95a] Dave Taylor. Web designer notes. ;login: the USENIX Association newsletter, 20(5):31– 34, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1995:WMC

[Tay95b] Dave Taylor. The Web master:

Context-sensitive Web pages with CGI. ;login: the USE-NIX Association newsletter, 20 (6):36–39, December 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Taylor:1996:WM

[Tay96a] Dave Taylor. The Web master.

;login: the USENIX Association newsletter, 21(1):22–26, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1996:W

[Tay96b] Dave Taylor. The webmaster. ;login: the USENIX Association newsletter, 21(3):30–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1996:WLY

[Tay96c] Dave Taylor. The Webmaster logging in to your Web pages. ;login: the USENIX Association newsletter, 21(4):47–??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1996:WTJ

[Tay96d] Dave Taylor. The Webmaster:

Traffic jam on iway-80. ;login: the USENIX Association newsletter, 21(2):39–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1996:WWS

[Tay96e] Dave Taylor. The webmaster:

Web site memories with cookies. ;login: the USENIX Association newsletter, 21(5):32–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1997:WMSa

[Tay97a] Dave Taylor. The Web master: a slick frame trick to ease navigation. ;login: the USENIX Association newsletter, 22(5):36–40, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1997:WMR

[Tay97b] Dave Taylor. The Web master: Roll your own search with grep. ;login: the USENIX Association newsletter, 22(4):37–41, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1997:WMSb

[Tay97c] Dave Taylor. The Web master: Surveys online! ;login: the USE-NIX Association newsletter, 22 (5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-10/
webmaster.html.

Taylor:1997:WMV

[Tay97d]  Dave Taylor.  The Web

master: The virtual poet. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
webmaster.html.

Taylor:1997:WMU

[Tay97e] Dave Taylor. The Web master: Using frames to enhance your Web site. ;login: the USE-NIX Association newsletter, 22 (2):32–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1998:Wa

[Tay98a] Dave Taylor. The Webmaster. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-4/
webmaster.html.

Taylor:1998:Wb

[Tay98b]  Dave Taylor.  The Web-

master. ;login: the USE-NIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
webmaster.html.

Taylor:1999:WPP

[Tay99] Dave Taylor. The Webmaster:

Password-protecting areas of your Web site. ;login: the USE-NIX Association newsletter, 24 (2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-4/
webmaster.html.

Tracey:1994:DDI

[TB94] John Michael Tracey and Arindam Banerji. Device driver issues in high-performance networking. In USENIX [USE94f], pages 31–43. ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        hispeed/banerji.html.

Tokuda:1999:ATM

[TB99] Lance Tokuda and Don Batory. Automating three modes of evolution for object-oriented software architectures. In USE-NIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots99/tokuda.html.

Tilbrook:1992:LSP

[TC92] David Tilbrook and Russell Crook. Large scale porting through parameterization. In USENIX [USE92g], pages 209–

216. ISBN 1-880446-44-8. LCCN

QA 76.76 O63 U83 1992.

Trotter:1995:WAA

[TC95] John Trotter and Mark Cravatt. A wireless adapter architecture for mobile computing. In USE-NIX [USE95i], pages 25–32 (or 25–31??). ISBN 1-880446-69-

3. LCCN TK5103.2 .U83 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        mob95/trotter.html.

Thornley:1998:SSH

[TCI98] John Thornley, K. Mani Chandy, and Hiroshi Ishii. A system for structured high-performance multithreaded programming in Windows NT. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

usenix.org/publications/library/
proceedings/usenix-nt98/thornley.
html;   http://www.usenix.
org/publications/library/
proceedings/usenix-nt98/thornley_
slides.

Tait:1992:EVC

[TD92] Carl D. Tait and Dan Duchamp. An efficient, variableconsistency, replicated file service. In USENIX [USE92h], pages 111–126. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Tennenhouse:1996:AN

[Ten96]  David Tennenhouse.  Ac-

tive networks. In USE-NIX [USE96f], pages 89– ?? ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/invited_talks/tennenhouse/
abstract.html.

Traugott:1998:BI

[TH98] Steve Traugott and Joel Huddleston. Bootstrapping an infrastructure. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/traugott.html.

Foundation:1998:SSN

[the98] the Board of the NLnet Foundation. A second start for the NLnet Foundation. ;login: the USENIX Association newsletter, 23(4), June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Thomas:1990:UOS

[Tho90a] Cedric Thomas. From Unix to open systems. EUUG Newsletter, 10(1):2–8, Spring 1990. CODEN EONLE8. ISSN 1011-4211.

Thompson:1990:KLE

[Tho90b] Tim Thompson. Keynote — A language and extensible graphic editor for music. In USENIX [USE90b], pages 331–358.

Thomas:1996:ERO

[Tho96] Dave Thomas. Experiences on the road to object utopia: An industrial research and development perspective. In USENIX [USE96g], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/coots96/. Keynote

address.

Thorpe:1998:SES

[Tho98a] Christopher Thorpe. SSU: Extending SSH for secure root administration. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63

                                        S97 1998. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/lisa98/thorpe.
                                        html.

Thorpe:1998:MID

[Tho98b] Jason R. Thorpe. A machineindependent DMA framework for NetBSD. In USENIX [USE98d], page ?? ISBN ????

LCCN ???? URL http://www.
usenix.org/publications/library/
proceedings/usenix98/freenix/
thorpe.dma.ps.

Thorpe:1998:NOS

[Tho98c]  Jason R. Thorpe.  The

NetBSD operating system. In USENIX [USE98d], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/thorpe.netbsd.
ps.

Tiemann:1990:EHI

[Tie90] Michael D. Tiemann. An exception handling implementation for C++. In USENIX [USE90j], pages 215–232.

Tiemann:1993:EHI

[Tie93a] Michael D. Tiemann. An exception of handling implementation for C++. In Waldo [Wal93b], pages 173–185. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Tiemann:1993:SRP

[Tie93b] Michael D. Tiemann. Solving the RPC problem in GNU C++. In Waldo [Wal93b], pages 217–233. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Tsuda:1997:ISM

[TIF^{+}97] Yoshiyuki Tsuda, Masahiro Ishiyama, Atsushi Fukumoto, Atsushi Inoue, and Ken ichi Yokoyama. Implementing security and mobility functions in kernel drivers. In USENIX [USE97f], page 147. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997. URL

http://www.usenix.org/publications/
library/proceedings/usenix-
nt97/single_tsuda.html.

Tierney:1994:SII

[TJH^{+}94] Brian Tierney, Bill Johnston, Hanan Herzog, Gary Hoo, Guojun Jin, and Jason Lee. System issues in implementing high speed distributed parallel storage systems. In USENIX [USE94f], pages 61–72. ISBN 1-880446-63-

4. LCCN TK5105.5 .U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
hispeed/tierney.html.

Thompson:1999:CBA

[TJM^{+}99] Mary Thompson, William Johnston, Sr., Srilekha Mudumbai, Gary Hoo, Keith Jackson, and Abdelilah Essiari. Certificate-based access control for widely distributed resources. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/thompson.
html.

Tang:1996:CHT

[TL96] Lei Tang and Steve Low. Chrghttp: a tool for micropayments on the World Wide Web. In USENIX [USE96i], pages 123–129. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/tang.chrg-http.html.

Tang:1999:MBM

[TLK99] Ty Tang, Vipul Lal, and Shesha Krishnapura. MTEX

— A bridge for migrating CAD design environment from UNIX to NT. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/tang.html.

Tumblin:1997:XDT

[TM97] Henry R. Tumblin and Charles McElwain. xmb — dancing in the tar pits: Using Tcl/Tk to build a CASE environment. In USENIX [USE97d], pages 173–175. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_tumblin.html.

Tamches:1999:FGD

[TM99] Ariel Tamches and Barton P.

Miller. Fine-grained dynamic instrumentation of commodity operating system kernels. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/tamches.
html.

Thibault:1997:DSL

[TMC97] Scott Thibault, Renaud Marlet, and Charles Consel. A domain specific language for video device drivers: From design to implementation. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/dsl97/thibault.
                                        html.

Tevanian:1991:STN

[TMJY91] Avadis Tevanian, Trey Matteson, David Jaffe, and Bryan Yamamoto. Software technology at NeXT computer. In USE-NIX [USE91g], pages 313–328. LCCN QA 76.76 O63 U83 1991.

Tokuda:1991:ERT

[TN91] Hideyuki Tokuda and Tatsuo Nakajima. Evaluation of real-time synchronization in real-time Mach. In USE-NIX [USE91h], pages 213–222. LCCN QAX 27.

Tezuka:1996:SCM

[TN96] Hiroshi Tezuka and Tatsuo Nakajima. Simple continuous media storage server on real-time Mach. In USE-NIX [USE96k], pages 87–98. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd96/tekuza.html.

Tokuda:1990:RTM

[TNR90] Hideyuki Tokuda, Tatsuo Nakajima, and Prithvi Rao. Realtime Mach: Towards a predictable real-time system. In USENIX [USE90e], pages 73–82.

Tompkins:1996:NTT

[Tom96] Raven Tompkins. A new twist on teaching system administration. In USENIX [USE96j], pages 227–231. ISBN 1-880446-81-2. LCCN QA76.76.O63

S97 1996. URL http://www.
usenix.org/publications/library/
proceedings/lisa96/rtompkin.
html.

Toma:1998:BGM

[Tom98] Rory Toma. Branchstart — A generic, multi-OS installation server, 1998. Unpublished invited talk at the 12th Systems Administration Conference (LISA ’98) December 6–11, 1998 Boston, Massachusetts, USA.

Toomey:1998:USL

[Too98] Warren Toomey. Use the source, Luke! Again. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
luke.html.

Torvalds:1996:LAE

[Tor96] Linus Torvalds. Linux: Architecture, experiences, and future. In USENIX [USE96k], page ?? ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/

sd96/. Unpublished invited

talk.

Terek:1991:EAC

[TP91] Robert Terek and Joseph Pasquale. Experiences with audio conferencing using the X Window System, UNIX and TCP/ IP. In USENIX [USE91g], pages 405–418. LCCN QA 76.76 O63 U83 1991.

Trammell:1990:BPV

[Tra90] R. D. Trammell. The big picture: Visualizing system behavior in real time. In USE-NIX [USE90g], pages 257–266. LCCN QA76.8.U65 U81 1990.

Trocki:1996:PAT

[Tro96] Jim Trocki. PC administration tools: Using Linux to manage personal computers. In USENIX [USE96j], pages 187–191. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/trockij.html.

Trudel:1998:DHD

[Tru98] Kim Trudel. Do help desks help? ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

Tso:1997:MEE

[Ts’97] Theodore Ts’o. Microsoft ‘embraces and extends’ Kerberos V5. ;login: the USENIX Association newsletter, 22(6):

??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
embraces.html.

Thompson:1990:TIA

[TSE90] Bruce Thompson, Daryl Stolte, and David Ellis. A transparent integration approach for rewritable optical autochangers. In USENIX [USE90g], pages 119–126. LCCN QA76.8.U65 U81 1990.

Takahashi:1995:FBN

[TSM95] Toshinari Takahashi, Atsushi Shimbo, and Masao Murota. File-based network collaboration system. In USE-NIX [USE95f], pages 95–

104. ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
takahashi.html.

Thomas:1996:UHA

[TSO^{+}96] Gregory S. Thomas, James O.

Schroeder, Merrilee E. Orcutt, Desiree C. Johnson, Jeffrey T. Simmelink, and John P. Moore. UNIX host administration in a heterogeneous distributed computing environment. In USENIX [USE96j], pages 43–

50. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL http://www.usenix.org/
publications/library/proceedings/
lisa96/gthomas.html.

Treese:1993:XTF

[TW93] G. Winfield Treese and Alec Wolman. X through the firewall, and other application relays. In USENIX [USE93g], pages 87–99. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/treese.html.

Tygar:1996:WEC

[TW96] J. D. Tygar and Alma Whitten. WWW electronic commerce and Java Trojan horses. In USENIX [USE96h], pages 243–250. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec96/whitten.html.

Uhler:1990:MSP

[Uhl90] Stephen A. Uhler. MTX — A shell that permits dynamic rearrangement of process connections and Windows. In USE-NIX [USE90h], pages 275–285. LCCN QA76.8.U65 U82 1990.

Uhler:1991:HMT

[Uhl91] Stephen A. Uhler. $HOME MOVIE — tools for building demos on a Sparcstation. In USE-NIX [USE91i], pages 127–136. LCCN QA 76.76 O63 U84 1992.

Uhler:1993:PMT

[Uhl93] Stephen A. Uhler. PhoneStation, moving the telephone onto the virtual desktop. In USENIX [USE93l], pages 131–

140. ISBN 1-880446-48-0.

LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/uhler.pdf. Best presen-

tation winner.

Uhler:1995:GUI

[Uhl95] Stephen A. Uhler. A graphical user interface builder for Tk. In USENIX [USE95j], pages 145–146. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/uhler.html.

Uhler:1996:SPM

[Uhl96] Stephen A. Uhler. In search of the perfect mega-widget. In USENIX [USE96e], pages 125–128. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

URL http://www.usenix.org/
publications/library/proceedings/
tcl96/uhler.html.

Unrau:1994:ELN

[UKGS94] Ronald C. Unrau, Orran Krieger, Benjamin Gamsa, and Michael Stumm. Experiences with locking in a NUMA multiprocessor operating system kernel. In USENIX [USE94i], pages 139–152. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

URL http://www.usenix.org/
publications/library/proceedings/
osdi/unrau.html.

Ulferts:1995:IWO

[Ulf95] Mark L. Ulferts. [incr Widgets] an object-oriented mega-widget set. In USENIX [USE95j], pages 61–76. ISBN 1-880446-72-3.

LCCN QA76.73.T44 T44 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl95/ulferts.html.

Ullmer:1997:DIT

[Ull97] Brygg Ullmer. 3wish: Distributed [incr Tcl] extensions for physical-world interfaces. In USENIX [USE97d], pages 169–170. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl97/poster_ullmer.html.

Umali:1995:MLT

[Uma95] Rick Umali. Making a living in technical sales support. ;login: the USENIX Association newsletter, 20(3):33–34, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Umali:1996:RUH

[Uma96a] Rick Umali. Review: The USENET handbook: a user’s guide to netnews. ;login: the USENIX Association newsletter, 21(1):48–49, February 1996. CODEN LOGNEM. ISSN 1044-6397.

Umali:1996:RUT

[Uma96b] Rick Umali. Review: UNIX tamed. ;login: the USENIX Association newsletter, 21(4):57– ??, August 1996. CODEN LOGNEM. ISSN 1044-6397.

Umali:1996:RZA

[Uma96c] Rick Umali. Review: Zen and the Art of the Internet, 4th ed. ;login: the USENIX Association newsletter, 21(3):56–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Umali:1997:BRM

[Uma97a] Rick Umali. Book review: Mastering regular expressions. ;login: the USENIX Association newsletter, 22(5):66–67, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Umali:1997:RPP

[Uma97b] Rick Umali. Review: Programming Perl, 2nd Edition. ;login: the USENIX Association newsletter, 22(1):59–60, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Umali:1998:BRM

[Uma98a] Rick Umali. Book review:

Managing Mailing Lists. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/mailing.html.

Umali:1998:BRP

[Uma98b] Rick Umali. Book review: Privacy on the Line. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-6/
diffie.html.

Urban:1990:GRG

[Urb90]  Michael Urban.  GAUD:

RAND’s group and user database.

In USENIX [USE90f], pages 17–

22. LCCN QA76.76.O63 L37

1990.

Urban:1991:GRG

[Urb91] Michael Urban. Gaud: RAND’s group and user database. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Urner:1997:PSP

[Urn97] Douglas L. Urner. Pinpointing system performance issues. In USENIX [USE97c], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa97/17.urner.html.

USENIX:1990:CSF

[USE90a] USENIX, editor. Computing Systems, Fall, 1990. USENIX, Berkeley, CA, USA, Fall 1990.

USENIX:1990:CSSa

[USE90b] USENIX, editor. Computing Systems, Spring, 1990. USE-NIX, Berkeley, CA, USA, Spring 1990.

USENIX:1990:CSSb

[USE90c] USENIX, editor. Computing Systems, Summer, 1990. USE-NIX, Berkeley, CA, USA, Summer 1990.

USENIX:1990:CSW

[USE90d] USENIX, editor. Computing Systems, Winter, 1990. USE-NIX, Berkeley, CA, USA, Winter 1990.

USENIX:1990:MWC

[USE90e] USENIX, editor. Mach Workshop Conference Proceedings, October 4–5, 1990. Burlington, VT. USENIX, Berkeley, CA, USA, 1990.

USENIX:1990:PFL

[USE90f] USENIX, editor. Proceedings of the fourth Large Installation System Administrator’s Conference: October 17–19, 1990, Colorado Springs, Colorado, USA. USENIX, Berkeley, CA, USA, 1990. LCCN QA76.76.O63 L37 1990.

USENIX:1990:PSU

[USE90g] USENIX, editor. Proceedings of the Summer 1990 USENIX Conference: June 11–15, 1990, Anaheim, California, USA. USE-NIX, Berkeley, CA, USA, 1990. LCCN QA76.8.U65 U81 1990.

USENIX:1990:PWU

[USE90h] USENIX, editor. Proceedings of the Winter 1990 USENIX Conference, January 22–26, 1990, Washington, DC, USA. USE-NIX, Berkeley, CA, USA, 1990. LCCN QA76.8.U65 U82 1990.

USENIX:1990:USI

[USE90i] USENIX, editor. UNIX Security II: USENIX workshop proceedings, August 27–28, 1990, Portland, Oregon. USENIX, Berkeley, CA, USA, 1990. LCCN QA 76.9 A25 U55 1990.

USENIX:1990:UCC

[USE90j] USENIX, editor. USENIX C++ conference proceedings: C++ Conference, San Francisco, California, April 9–11, 1990. USE-NIX, Berkeley, CA, USA, Spring 1990.

USENIX:1991:CSF

[USE91a] USENIX, editor. Computing Systems, Fall, 1991. USENIX, Berkeley, CA, USA, Fall 1991.

USENIX:1991:CSSa

[USE91b] USENIX, editor. Computing Systems, Spring, 1991. USE-NIX, Berkeley, CA, USA, Spring 1991.

USENIX:1991:CSSb

[USE91c] USENIX, editor. Computing Systems, Summer, 1991, volume 4(3). USENIX, Berkeley, CA, USA, Summer 1991.

USENIX:1991:CSW

[USE91d] USENIX, editor. Computing Systems, Winter, 1991. USE-NIX, Berkeley, CA, USA, Winter 1991.

USENIX:1991:LIC

[USE91e] USENIX, editor. LISA IV Conference Proceedings, October 24–25, 1991. Colorado Springs, CO. USENIX, Berkeley, CA, USA, October 24–25, 1991. ISBN ???? LCCN ????

USENIX:1991:PFL

[USE91f] USENIX, editor. Proceedings of the fifth Large Installation Systems Administration Conference: September 30–October 3, 1991, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1991.

USENIX:1991:PSU

[USE91g] USENIX, editor. Proceedings of the Summer 1991 USE-NIX Conference: June 10– 14, 1991, Nashville, Tennessee, USA. USENIX, Berkeley, CA, USA, 1991. LCCN QA 76.76 O63 U83 1991.

USENIX:1991:PUM

[USE91h] USENIX, editor. Proceedings of the USENIX Mach Symposium: November 20–22, 1991, Monterey, California, USA. USE-NIX, Berkeley, CA, USA, 1991. LCCN QAX 27.

USENIX:1991:PWU

[USE91i] USENIX, editor. Proceedings of the Winter 1991 USENIX Conference: January 21–January 25, 1991, Dallas, TX, USA. USENIX, Berkeley, CA, USA, 1991. LCCN QA 76.76 O63 U84 1992.

USENIX:1991:SED

[USE91j]  USENIX, editor.  Sympo-

sium on Experiences with Distributed and Multiprocessor Systems (SEDMS II), March 21– 22, 1991. Atlanta GA. USENIX, Berkeley, CA, USA, March 21– 22, 1991. ISBN ???? LCCN ????

USENIX:1991:UCC

[USE91k] USENIX, editor. USENIX C++ conference proceedings: Washington, DC, April 22–25. 1991. USENIX, Berkeley, CA, USA, 1991. LCCN QA76.73.C153 U27 1991.

USENIX:1991:UCPc

[USE91l]  USENIX, editor.  USENIX

Conference Proceedings, Winter, 1991. Dallas, TX. USE-NIX, Berkeley, CA, USA, Winter 1991.

USENIX:1992:CSF

[USE92a] USENIX, editor. Computing Systems, Fall, 1992, volume 5(4). USENIX, Berkeley, CA, USA, Fall 1992.

USENIX:1992:CSSa

[USE92b] USENIX, editor. Computing Systems, Spring, 1992. USE-NIX, Berkeley, CA, USA, Spring 1992.

USENIX:1992:CSSb

[USE92c] USENIX, editor. Computing Systems, Summer, 1992. USE-NIX, Berkeley, CA, USA, Summer 1992.

USENIX:1992:CSW

[USE92d] USENIX, editor. Computing Systems, Winter, 1992. USE-NIX, Berkeley, CA, USA, Winter 1992.

USENIX:1992:PSS

[USE92e] USENIX, editor. Proceedings of the sixth Systems Administration Conference (LISA VI):

October 19–23, 1992, Long Beach, California, USA. USE-NIX, Berkeley, CA, USA, 1992. ISBN 1-880446-47-2. LCCN ????

USENIX:1992:PSE

[USE92f] USENIX, editor. Proceedings of the Spring 1992 EurOpen/ USENIX Workshop, April 6–9, 1992, Jersey, Channel Islands. EurOpen, Buntingford, Herts, UK, 1992. ISBN 1-873611-03-X. LCCN ????

USENIX:1992:PSU

[USE92g] USENIX, editor. Proceedings of the Summer 1992 USENIX Conference: June 8–12, 1992, San Antonio, Texas, USA. USE-NIX, Berkeley, CA, USA, Summer 1992. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

USENIX:1992:PUF

[USE92h] USENIX, editor. Proceedings of the USENIX File Systems Workshop, 21–22 May, 1992, Ann Arbor, MI, USA. USE-NIX, Berkeley, CA, USA, 1992. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

USENIX:1992:PUW

[USE92i] USENIX, editor. Proceedings of the USENIX Workshop on Micro-Kernels and Other Kernel Architectures: 27–28 April, 1992, Seattle, WA, USA. USE-NIX, Berkeley, CA, USA, 1992. ISBN 1-880446-42-1. LCCN QAX 32.

USENIX:1992:PWU

[USE92j] USENIX, editor. Proceedings of the Winter 1992 USENIX Conference: January 20 — January 24, 1992, San Francisco, California. USENIX, Berkeley, CA, USA, 1992.

USENIX:1992:PUC

[USE92k] USENIX, editor. Proceedings: USENIX C++ Technical Conference, August 10–13, 1992, Portland, OR. USENIX, Berkeley, CA, USA, 1992. ISBN 1-880446-45-6. LCCN ????

USENIX:1992:SED

[USE92l]  USENIX, editor.  Sympo-

sium on Experiences with Distributed and Multiprocessor Systems (SEDMS III), March 26– 27, 1992. Newport Beach, CA. USENIX, Berkeley, CA, USA, March 26–27, 1992. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

USENIX:1992:USI

[USE92m] USENIX, editor. UNIX Security III Symposium, September 14–17, 1992. Baltimore, MD. USENIX, Berkeley, CA, USA, September 14–17, 1992. ISBN 1-880446-46-4. LCCN ????

USENIX:1992:UCC

[USE92n] USENIX, editor. USENIX C++ Conference Proceedings, August 10–13, 1992. Portland, OR. USENIX, Berkeley, CA, USA, August 10–13, 1992.

USENIX:1992:UCT

[USE92o] USENIX, editor. USENIX C++ Technical Conference: proceedings: August 10–13, 1992, Portland, Oregon. USENIX, Berkeley, CA, USA, 1992. ISBN 1-880446-45-6. LCCN ????

USENIX:1993:CSF

[USE93a] USENIX, editor. Computing Systems, Fall, 1993. USENIX, Berkeley, CA, USA, Fall 1993.

USENIX:1993:CSSa

[USE93b] USENIX, editor. Computing Systems, Spring, 1993. USE-NIX, Berkeley, CA, USA, Spring 1993.

USENIX:1993:CSSb

[USE93c] USENIX, editor. Computing Systems, Summer, 1993. USE-NIX, Berkeley, CA, USA, Summer 1993.

USENIX:1993:CSW

[USE93d] USENIX, editor. Computing Systems, Winter, 1993. USE-NIX, Berkeley, CA, USA, Winter 1993.

USENIX:1993:JDS

[USE93e] USENIX, editor. Job Descriptions for System Administrators, October, 1993. USENIX for SAGE, the System Administrator’s Guild, Berkeley, CA, USA, October 1993.

USENIX:1993:PSS

[USE93f] USENIX, editor. Proceedings of the Seventh Systems Administration Conference (LISA VII): November 1–5, 1993, Monterey, CA, USA. USE-NIX, Berkeley, CA, USA, 1993. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/.

USENIX:1993:PSU

[USE93g] USENIX, editor. Proceedings of the Summer 1993 USE-NIX Conference: June 21–25, 1993, Cincinnati, Ohio, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/.

USENIX:1993:PUMa

[USE93h] USENIX, editor. Proceedings of the USENIX Mach III Symposium: April 19–21, 1993, Santa Fe, New Mexico, USA. USE-NIX, Berkeley, CA, USA, 1993. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

USENIX:1993:PUMb

[USE93i]  USENIX, editor.  Proceed-

ings of the USENIX Mobile and Location-Independent Computing Symposium: August 2–3, 1993, Cambridge, Massachusetts, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/.

USENIX:1993:PUSb

[USE93j] USENIX, editor. Proceedings of the USENIX Symposium on Experiences with Distributed and Multiprocessor Systems (SEDMS IV): September 22– 23, 1993, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-54-

5. LCCN QA76.9.D5 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sedms4/.

USENIX:1993:PUSa

[USE93k] USENIX, editor. Proceedings of the USENIX Symposium on Microkernels and Other Kernel Architectures: September 20– 21, 1993, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-52-9. LCCN QA76.9.A73 U83 1993.

USENIX:1993:PWU

[USE93l] USENIX, editor. Proceedings of the Winter 1993 USENIX Conference: January 25–29, 1993, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

USENIX:1993:USI

[USE93m] USENIX, editor. UNIX Security IV Symposium proceedings: October 4–6, 1993, Santa Clara, CA, USA. USENIX, Berkeley, CA, USA, October 4– 6, 1993. ISBN 1-880446-55-3.

LCCN QA 76.9 A25 U54 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec4/.

USENIX:1993:L

[USE93n] ;Login:, 1993. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA. The USENIX Association newsletter. Current frequency: Bimonthly.

USENIX:1994:CSF

[USE94a] USENIX, editor. Computing Systems, Fall, 1994. USENIX, Berkeley, CA, USA, Fall 1994.

USENIX:1994:CSSa

[USE94b] USENIX, editor. Computing Systems, Spring, 1994. USE-NIX, Berkeley, CA, USA, Spring 1994.

USENIX:1994:CSSb

[USE94c] USENIX, editor. Computing Systems, Summer, 1994. USE-NIX, Berkeley, CA, USA, Summer 1994.

USENIX:1994:CSW

[USE94d] USENIX, editor. Computing Systems, Winter, 1994. USE-NIX, Berkeley, CA, USA, Winter 1994.

USENIX:1994:PUC

[USE94e]  USENIX, editor.  Proceed-

ings of the 1994 USENIX C++ Conference: April 11–14, 1994, Cambridge, MA. USE-NIX, Berkeley, CA, USA, 1994. ISBN 1-880446-60-X. LCCN QA 76.73 C153 C22 1994.

URL http://www.usenix.org/
publications/library/proceedings/
c++94/index.html.

USENIX:1994:PUSa

[USE94f] USENIX, editor. Proceedings of the 1994 USENIX Symposium on High-Speed Networking, August 1–3, 1994, Oakland, California, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994.

USENIX:1994:PUU

[USE94g] USENIX, editor. Proceedings of the 1994 USENIX UNIX Applications Development Symposium: April 25–28, 1994, Toronto, Ontario, Canada. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL http://www.usenix.org/
publications/library/proceedings/
appdev94/index.html.

USENIX:1994:PES

[USE94h] USENIX, editor. Proceedings of the Eighth Systems Administration Conference (LISA VIII): September 19–23, 1994, San Diego, CA, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

USENIX:1994:PFU

[USE94i]  USENIX, editor.  Proceed-

ings of the First USENIX Symposium on Operating Systems Design and Implementation (OSDI): November 14– 17, 1994, Monterey, California, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/index.html.

USENIX:1994:PSU

[USE94j] USENIX, editor. Proceedings of the Summer 1994 USENIX Conference: June 6–10, 1994, Boston, Massachusetts, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

USENIX:1994:PUSb

[USE94k] USENIX, editor. Proceedings of the USENIX Symposium on Very High Level Languages (VHLL): October 26–28, 1994, Santa Fe, New Mexico, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-65-

0. LCCN QA76.7 .U74 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        vhll/index.html.

USENIX:1994:PWU

[USE94l] USENIX, editor. Proceedings of the Winter 1994 USENIX Conference: January 17–21, 1994, San Francisco, California, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/index.html.

USENIX:1994:TAS

[USE94m] USENIX, editor. The Third Annual System Administration, Networking and Security Conference (SANS III), April 4–8, 1994. Washington, DC. USE-NIX, Berkeley, CA, USA, April 4–8, 1994. ISBN ???? LCCN ????

USENIX:1995:CSF

[USE95a] USENIX, editor. Computing Systems, Fall, 1995. USENIX, Berkeley, CA, USA, Fall 1995.

USENIX:1995:CSSa

[USE95b] USENIX, editor. Computing Systems, Spring, 1995. USE-NIX, Berkeley, CA, USA, Spring 1995.

USENIX:1995:CSSc

[USE95c] USENIX, editor. Computing Systems, Summer, 1995. USE-NIX, Berkeley, CA, USA, Summer 1995.

USENIX:1995:CSW

[USE95d] USENIX, editor. Computing Systems, Winter, 1995. USE-NIX, Berkeley, CA, USA, Winter 1995.

USENIX:1995:PUT

[USE95e] USENIX, editor. Proceedings of the 1995 USENIX Technical Conference: January 16–20, 1995, New Orleans, Louisiana, USA. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/index.html.

USENIX:1995:PFUa

[USE95f] USENIX, editor. Proceedings of the fifth USENIX UNIX Security Symposium: June 5–7, 1995, Salt Lake City, Utah, USA. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
index.html.

USENIX:1995:PFUb

[USE95g] USENIX, editor. Proceedings of the first USENIX Workshop of Electronic Commerce: July 11–12, 1995, New York, New York, USA. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/index.html.

USENIX:1995:PNS

[USE95h] USENIX, editor. Proceedings of the Ninth Systems Administration Conference (LISA IX): September 17–22, 1995, Monterey, CA, USA. USE-NIX, Berkeley, CA, USA, 1995. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/index.html.

USENIX:1995:PSU

[USE95i]  USENIX, editor.  Proceed-

ings of the second USENIX Symposium on Mobile and Location-Independent Computing: April 10–11, 1995, Ann Arbor, Michigan, USA. USE-NIX, Berkeley, CA, USA, 1995. ISBN 1-880446-69-3. LCCN TK5103.2 .U83 1995.

URL http://www.usenix.org/
publications/library/proceedings/
mob95/index.html.

USENIX:1995:PTT

[USE95j] USENIX, editor. Proceedings of the Tcl/Tk Workshop, July 6–8, 1995, Toronto, Ontario, Canada, Proceedings of the Tcl Tk Workshop — USE-NIX Association 1995; 3rd. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-72-3. LCCN QA76.73.T44 T44 1995.

URL http://www.usenix.org/
publications/library/proceedings/
tcl95/index.html.

USENIX:1995:PUC

[USE95k] USENIX, editor. Proceedings of the USENIX Conference on Object-Oriented Technologies (COOTS): June 26– 29, 1995, Monterey, California, USA. USENIX, Berkeley, CA, USA, 1995. ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/index.html.

USENIX:1995:UCP

[USE95l]  USENIX, editor.  USENIX

Conference Proceedings, January 16–20, 1995. New Orleans, LA. USENIX, Berkeley, CA, USA, January 16–20, 1995.

USENIX:1996:CSF

[USE96a] USENIX, editor. Computing Systems, Fall, 1996. USENIX, Berkeley, CA, USA, Fall 1996.

USENIX:1996:CSSa

[USE96b] USENIX, editor. Computing Systems, Spring, 1996. USE-NIX, Berkeley, CA, USA, Spring 1996.

USENIX:1996:CSSb

[USE96c] USENIX, editor. Computing Systems, Summer, 1996. USE-NIX, Berkeley, CA, USA, Summer 1996.

USENIX:1996:CSW

[USE96d] USENIX, editor. Computing Systems, Winter, 1996. USE-NIX, Berkeley, CA, USA, Winter 1996.

USENIX:1996:PFA

[USE96e]  USENIX, editor.  Proceed-

ings of the fourth annual Tcl/ Tk Workshop, July 10–13, 1996, Monterey, California. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/.

USENIX:1996:PSS

[USE96f] USENIX, editor. Proceedings of the Second Symposium on Operating Systems Design and Implementation (OSDI ’96): October 28–31, 1996, Seattle, Washington. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-82-0.

LCCN QA76.76.O63 S96 1996.

URL http://www.usenix.org/
publications/library/proceedings/
osdi96/.

USENIX:1996:PSUb

[USE96g] USENIX, editor. Proceedings of the Second USENIX Conference on Object-Oriented Technologies and Systems (COOTS), June 17–21, 1996, Toronto, Canada. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/.

USENIX:1996:PSUa

[USE96h] USENIX, editor. Proceedings of the second USE-NIX Workshop on Electronic Commerce: November 18– 21, 1996, Oakland, California. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-83-

9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

USENIX:1996:PSA

[USE96i]  USENIX, editor.  Proceed-

ings of the sixth annual USE-NIX Security Symposium, focusing on applications of cryptography, July 22–25, 1996, San Jose, California. USENIX, Berkeley, CA, USA, July 22– 25, 1996. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/.

USENIX:1996:PTS

[USE96j]  USENIX, editor.  Proceed-

ings of the Tenth Systems Administration Conference (LISA X), September 29–October 4, 1996, Chicago, IL, USA. USE-NIX, Berkeley, CA, USA, 1996. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

USENIX:1996:PUA

[USE96k] USENIX, editor. Proceedings of the USENIX 1996 annual technical conference: January 22– 26, 1996, San Diego, California, USA, USENIX Conference Proceedings 1996. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

USENIX:1997:DLS

[USE97a] USENIX, editor. Digest of the Large Scale System Administration of Windows NT Workshop, August 14–16, 1997, Seattle, Washington, USA. USE-NIX, Berkeley, CA, USA, 1997. ISBN ???? LCCN ????

USENIX:1997:PCD

[USE97b] USENIX, editor. Proceedings of the Conference on Domain-Specific Languages, October 15– 17, 1997, Santa Barbara, California. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-89-

8. LCCN QA76.7 .C663 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        dsl97/.

USENIX:1997:PES

[USE97c]  USENIX, editor.  Proceed-

ings of the Eleventh Systems Administration Conference (LISA XI), October 26– 31, 1997, San Diego, CA, USA. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL http://www.usenix.org/
publications/library/proceedings/
lisa97/.

USENIX:1997:PFA

[USE97d] USENIX, editor. Proceedings of the fifth annual Tcl/ Tk Workshop, July 14–17, 1997, Boston, Massachusetts. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/.

USENIX:1997:PUA

[USE97e] USENIX, editor. Proceedings of the USENIX 1997 annual Technical Conference, January 6– 10, 1997, Anaheim, California, USA. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

USENIX:1997:PUW

[USE97f]  USENIX, editor.  Proceed-

ings of the USENIX Windows NT Workshop: August 11– 13, 1997, Seattle, Washington. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

USENIX:1997:TUC

[USE97g] USENIX, editor. The Third USENIX Conference on Object-Oriented Technologies and Systems (COOTS), June 16–19, 1997. Portland, Oregon. USE-NIX, Berkeley, CA, USA, June 16–20, 1997. ISBN 1-880446-86-3. LCCN QA76.64 .U845

1997. URL http://dxsting.
cern.ch/sting/COOTS97.html.

See also [LH97, Rie97, WW97].

USENIX:1997:USI

[USE97h] USENIX, editor. USENIX Symposium on Internet Technologies and Systems Proceedings, Monterey, California, December 8–11, 1997. USE-NIX, Berkeley, CA, USA, 1997. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/.

USENIX:1998:PUWa

[USE98a] USENIX, editor. Proceedings of the 2nd USENIX Windows NT Symposium: August 3–5, 1998, Seattle, Washington. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-95-

2. LCCN QA76.76.O63 U885

1998. URL http://db.usenix.
org/publications/library/
proceedings/usenix-nt98.

USENIX:1998:PUWb

[USE98b] USENIX, editor. Proceedings of the 3rd USENIX Workshop on Electronic Commerce: August 31–September 3, 1998, Boston, Mass. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-97-9. LCCN HF5004 .U74

1998. URL http://db.usenix.
org/publications/library/
proceedings/ec98.

USENIX:1998:PFU

[USE98c] USENIX, editor. Proceedings of the fourth USENIX Conference on Object-Oriented Technologies and Systems (COOTS): April 27–30, 1998, Santa Fe, NM. USENIX, Berkeley, CA, USA, 1998. ISBN ???? LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/coots98.

USENIX:1998:PFT

[USE98d] USENIX, editor. Proceedings of the FreeNIX Track: USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USE-NIX, Berkeley, CA, USA, 1998. ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/freenix/.

USENIX:1998:PLI

[USE98e] USENIX, editor. Proceedings of the Large Installation System Administration of Windows NT Conference, August 5–8, 1998, Seattle, Washington. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/.

USENIX:1998:PSA

[USE98f] USENIX, editor. Proceedings of the sixth annual Tcl/Tk Conference, September 18–24 [i.e. 14–18], 1998, San Diego, California. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-98-7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl98/.

USENIX:1998:PTS

[USE98g] USENIX, editor. Proceedings of the Twelfth Systems Administration Conference (LISA XII). USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-40-5. LCCN QA76.76.O63 S97

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa98.

USENIX:1998:PUA

[USE98h] USENIX, editor. Proceedings of the USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-94-4. LCCN QA76.76.O63 U88

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/usenix98.

USENIX:1998:SUS

[USE98i] USENIX, editor. Seventh USE-NIX Security Symposium proceedings: conference proceedings: San Antonio, Texas, January 26–29, 1998. USE-NIX, Berkeley, CA, USA, 1998. ISBN 1-880446-92-

8. LCCN QA76.9.A25 U83

1998. URL http://db.usenix.
org/publications/library/
proceedings/sec98.

USENIX:1999:PCD

[USE99a] USENIX, editor. Proceedings of the 2nd Conference on Domain-Specific Languages (DSL ’99), October 3–5, 1999, Austin, Texas, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663 1999.

URL http://www.usenix.org/
publications/library/proceedings/
dsl99/.

USENIX:1999:PLI

[USE99b] USENIX, editor. Proceedings of the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999. USE-NIX, Berkeley, CA, USA, 1999. ISBN 1-880446-30-8. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa-nt99/.

USENIX:1999:PUS

[USE99c] USENIX, editor. Proceedings of the 2nd USENIX Symposium on Internet Technologies and Systems (USITS ’99), October 11– 14, 1999, Boulder, Colorado, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-26-X. LCCN TK5105.875.I57

U96. URL http://db.usenix.

                                        org/publications/library/
                                        proceedings/usits99.

USENIX:1999:PUWb

[USE99d] USENIX, editor. Proceedings of the 3rd USENIX Windows NT Symposium: July 12–15, 1999, Seattle, Washington, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-29-4. LCCN ????

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt99/.

USENIX:1999:PEU

[USE99e]  USENIX, editor.  Proceed-

ings of the eighth USENIX Security Symposium (Security ’99), August 23–26, 1999, Washington, DC, USA. USE-NIX, Berkeley, CA, USA, 1999. ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec99/.

USENIX:1999:PFU

[USE99f] USENIX, editor. Proceedings of the fifth USENIX Conference on Object-Oriented Technologies and Systems (COOTS ’99): May 3–7, 1999, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1999. ISBN ???? LCCN QA76.64 .U84 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots99/.

USENIX:1999:PFC

[USE99g] USENIX, editor. Proceedings of the first Conference on Network Administration (NETA ’99): April 7–10, 1999, Santa Clara, California. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-36-7. LCCN TK5105.5 .C664 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/.

USENIX:1999:PTSb

[USE99h] USENIX, editor. Proceedings of the Third Symposium on Operating Systems Design and Implementation: (OSDI ’99) February 22–25, 1999, New Orleans, Louisiana, volume 32(5) of Operating Systems Review. USENIX, Berkeley, CA, USA, 1999. CODEN OS-RED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/.

USENIX:1999:PTSa

[USE99i] USENIX, editor. Proceedings of the Thirteenth Systems Administration Conference (LISA XIII): November 7—12, 1999, Seattle, WA, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-25-1. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
lisa99/.

USENIX:1999:PUWa

[USE99j]  USENIX, editor.  Proceed-

ings of the USENIX Workshop on Smartcard Technology (Smartcard ’99): May 10–11, 1999, Chicago, Illinois, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/.

USENIX:1999:PWE

[USE99k] USENIX, editor. Proceedings of the Workshop on Embedded Systems, March 29–31, 1999, Cambridge, Massachusetts, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
es99/.

USENIX:1999:PWI

[USE99l] USENIX, editor. Proceedings of the Workshop on Intrusion Detection and Network Monitoring (ID ’99): April 9-12, 1999, Santa Clara, California. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-37-5. LCCN QA76.9.A25 W67 1999.

URL http://www.usenix.org/
publications/library/proceedings/
detection99/.

USENIX:1999:UAT

[USE99m] USENIX, editor. Usenix Annual Technical Conference. June 6–11, 1999. Monterey, California, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-33-2. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/usenix99.

Vahdat:1998:TRC

[VA98] Amin Vahdat and Thomas Anderson. Transparent result caching. In USENIX [USE98h], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix98/vahdat.html.

VanSciver:1990:ZGC

[Van90] Jim Van Sciver. Zone garbage collection. In USENIX [USE90e], pages 1–16.

VanWyk:1991:CLS

[Van91]  Christopher Van Wyk.  A

class library for solving simultaneous equations. In USE-NIX [USE91k], pages 229–234. LCCN QA76.73.C153 U27 1991.

vanRenesse:1992:A

[van92] Robbert van Renesse. Amoeba.

In USENIX [USE92i], pages 1–

10. ISBN 1-880446-42-1. LCCN QAX 32.

VanMeter:1997:OEM

[Van97] Rodney Van Meter. Observing the effects of multi-zone disks. In USENIX [USE97e], pages 19–30. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/summaries/vanmeter.
html;   http://www.usenix.
org/publications/library/
proceedings/ana97/vanmeter.
html.

VanDinter:1998:LEC

[Van98] Theo Van Dinter. Letter to the Editor: Computer attacks. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397.

Vanderburg:1999:BRH

[Van99] Glenn Vanderburg. Book review: Handbook of Programming Languages. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-8/
                                        bookreview.html.

VanBaak:19xx:FDD

[Vanxx] Thomas Van Baak. The file disk drive. In ???? USENIX, Berkeley, CA, USA, 19xx.

Various:1992:SWGa

[Var92a] Various. SAGE working groups. ;login: the USENIX Association newsletter, 17(5):15–16, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Various:1992:SWGb

[Var92b] Various. SAGE working groups. ;login: the USENIX Association newsletter, 17(6):15–18, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Various:1994:WN

[Var94] Various. What’s new? ;login: the USENIX Association newsletter, 19(3):18–26, May/June 1994. CODEN LOGNEM. ISSN 1044-6397.

Varian:1995:EMD

[Var95a]  Hal R. Varian.  Economic

mechanism design for computerized agents. In USENIX [USE95g], pages 13–21 (or 13– 22??). ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

URL http://www.usenix.org/
publications/library/proceedings/
ec95/varian.html.

Various:1995:TTA

[Var95b] Various. Top ten anagrams for information superhighway. ;login: the USENIX Association newsletter, 20(4):41–??, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Contributors:1996:RTS

[Var96] Various Contributors. Reports on the tenth systems administration conference (LISA ’96). ;login: the USENIX Association newsletter, 21(6):29–40, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Contributors:1997:RSS

[Var97] Various Contributors. Reports on the second symposium on operating design and implementation (OSDI ’96). ;login: the USENIX Association newsletter, 22(1):36–58, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Vasilatos:1991:PMC

[Vas91] Nick Vasilatos. Partitioned multiprocessors and the coexistence of heterogeneous operating systems. In USENIX [USE91i], pages 23–32. LCCN QA 76.76 O63 U84 1992.

Vaughan:1990:PDA

[Vau90]  Francis Vaughan.  A per-

sistent distributed architecture supported by the Mach operating system. In USENIX [USE90e], pages 123–140.

VanEpp:1992:DMC

[VB92] Peter Van Epp and Bill Baines. Dropping the mainframe without crushing the users: Mainframe to distributed UNIX in nine months. In USENIX [USE92e], pages 39–53 (or 39– 54??). ISBN 1-880446-47-2. LCCN ????

vanRenesse:1992:RMB

[vBC^{+}92] Robbert van Renesse, Ken Birman, Robert Cooper, Bradford Glade, and Patrick Stephenson. Reliable multicast between micro-kernels. In USE-NIX [USE92i], pages 269–284. ISBN 1-880446-42-1. LCCN QAX 32.

Vaughan:1992:CCA

[VBD^{+}92] Francis Vaughan, Tracy Lo Basso, Alan Dearle, Chris Marlin, and Chris Barter. Casper: a cached architecture supporting persistence. In USENIX [USE92c], pages 337–359.

Vo:1992:ITA

[VC92] Kiem-Phong Vo and Yih-Farn Chen. Incl: a tool to analyze include files. In USENIX [USE92g], pages 199–208. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Varadarajan:1997:SIC

[VcC97] Srinidhi Varadarajan and Tzi cker Chiueh. SASE: Implementation of a compressed text search engine. In USE-NIX [USE97h], pages 253–262. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usits97/varadarajan.html.

Vckovski:1998:WDC

[Vck98] Andrej Vckovski. wshdbg — A debugger for CGI applications. In USENIX [USE98f], page ?? ISBN 1-880446-98-

7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/vckovski.
html.

Vangala:1992:SDM

[VCV92] Ram R. Vangala, Michael J.

Cripps, and Raj G. Varadarajan. Software distribution and management in a networked environment. In USENIX [USE92e], pages 163–170. ISBN 1-880446-47-2. LCCN ????

Vogels:1998:SMC

[VDA^{+}98] Werner Vogels, Dan Dumitriu, Ashutosh Agrawal, Teck Chia, and Katherine Guo. Scalability of the Microsoft Cluster Service. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt98/vogels.html;
                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/vogels_slides.pdf.

Vahdat:1999:ANF

[VDAA99] Amin Vahdat, Michael Dahlin, Thomas Anderson, and Amit Aggarwal. Active names: Flexible location and transport of wide-area resources. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usits99/
                                        vahdat.html.

Vella:1995:UCP

[Vel95] Zana Vella. USENIX and the CitySpace project. ;login: the USENIX Association newsletter, 20(4):10–??, August 1995. CODEN LOGNEM. ISSN 1044-6397.

Venema:1992:TWN

[Ven92] Wietse Venema. TCP WRAP-PER: Network monitoring, access control and booby traps. In USENIX [USE92m], pages 85–

92. ISBN 1-880446-46-4. LCCN

????

Venema:1996:MLC

[Ven96]  Wietse Venema.  Murphy’s

law and computer security. In USENIX [USE96i], pages 187–193. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

                                        URL http://www.usenix.org/

publications/library/proceedings/
sec96/venema.html.

Verdoes:1997:SCS

[Ver97] J. A. Verdoes. In search of a complete and scalable systems administration suite. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
nt-sysadmin97/verdoes.html.

Viles:1995:ALW

[VF95] Charles L. Viles and James C.

French. Availability and latency of world wide Web information servers. In USENIX [USE95d], pages 61–91.

Vahalia:1995:MLN

[VGT95] Uresh Vahalia, Cary G. Gray, and Dennis Ting. Metadata logging in an NFS server. In USENIX [USE95e], pages 265–

276. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/vahalia.html.

Vinoski:1996:BDA

[Vin96]  Steve Vinoski.  Building

distributed applications with CORBA and C++. In USE-NIX [USE96g], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

usenix.org/publications/library/
proceedings/coots96/.

Viswanadha:1996:PDA

[Vis96] Sreenivasa Rao Viswanadha.

Preliminary design of ADL/C++ — A specification language for C++. In USENIX [USE96g], pages 97–111. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

URL http://www.usenix.org/
publications/library/proceedings/
coots96/viswa.html.

Vixie:1992:M

[Vix92] Paul Vixie. Modula-3. ;login: the USENIX Association newsletter, 17(6):30–34, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

URL https://archive.org/
details/login_nov92/page/
n29/mode/2up?view=theater.

Vixie:1995:DBS

[Vix95]  Paul A. Vixie.  DNS and

BIND security issues. In USENIX [USE95f], pages 209–

216. ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
vixie.html.

Vixie:1998:LEW

[Vix98] Paul Vixie. Letter to the Editor: What potential abuses? ;login: the USENIX Association newsletter, 23(4):??, June 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Vixie:1999:DSB

[Vix99] Paul Vixie. Deep space BIND,

1999. URL http://db.usenix.
org/publications/library/
proceedings/lisa99/invitedtalks.

html. Unpublished invited talk at LISA XIII.

Viswanadha:1998:ILI

[VK98] Sreenivasa Viswanadha and Deepak Kapur. IBDL: a language for interface behavior specification and testing. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/viswanadha.html.

vonLwis:1998:PND

[vL98] Martin von Löwis. A portable NTFS driver. In USENIX [USE98d], page ?? ISBN ????

LCCN ???? URL http://www.
usenix.org/publications/library/
proceedings/usenix98/freenix/
loewis.ps.

Vellon:1998:ADV

[VMMD98] Manny Vellon, Kirk Marple, Don Mitchell, and Steven Drucker. The architecture of a distributed virtual worlds system. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/vellon.html.

Vo:1996:VSE

[Vo96]  Kiem-Phong Vo.  Vmal-

loc: The search ends?, 1996.

URL http://www.usenix.org/
publications/library/proceedings/

sd96/. Unpublished invited talk at the USENIX 1996 Annual Technical Conference, January 22–26, 1996, San Diego, CA.

Vo:1997:CGE

[Vo97]  Kiem-Phong Vo.  Cdt: a

general and efficient container data type library. In USE-NIX [USE97e], pages 163–

171. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/summaries/vo.html;
http://www.usenix.org/publications/
library/proceedings/ana97/
vo.html.

vanDriel:1995:XSH

[vR95] Ronald van Driel and Daan Reuhman. The X station: a high-end workstation in disguise. ;login: the USENIX Association newsletter, 20(3):30–31, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Vijaykrishnan:1999:TBP

[VR99] N. Vijaykrishnan and N. Ranganathan. Tuning branch predictors to support virtual method invocation in Java. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
coots99/vijaykrishnan.html.

Vessey:1990:IBS

[VS90] Ian Vessey and Glenn Skinner. Implementing Berkeley sockets in System V Release 4. In USE-NIX [USE90h], pages 177–194. LCCN QA76.8.U65 U82 1990.

Vincenzetti:1995:SST

[VTB95] David Vincenzetti, Stefano Taino, and Fabio Bolognesi. STEL: Secure TELnet. In USE-NIX [USE95f], pages 75–84 (or 75–83??). ISBN 1-880446-70-

7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/
vincenzetti.html.

vanWinkel:1998:SIC

[vW98] Jan Christiaan van Winkel.

SANE ’98 — an international conference. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397.

Valian:1999:NAD

[VW99a] Peter Valian and Todd K. Watson. NetReg: An automated DHCP registration system. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/valian.
html.

Herreweghen:1999:RPU

[VW99b] Els Van Herreweghen and Uta Wille. Risks and potentials of using EMV for Internet payments. In USENIX [USE99j], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL http://www.usenix.org/
publications/library/proceedings/
smartcard99/herreweghen.html.

Viega:1998:MGM

[VWM98] John Viega, Barry Warsaw, and Ken Manheimer. Mailman: The GNU Mailing List Manager. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa98/viega.html.

Veber:1991:CUD

[VZ91] Mark A. Veber and Elizabeth D. Zwicky. Configurable user documentation -or- how I came to write a language with a future conditional. In USENIX [USE91f], pages 153–158.

Wagner:1990:DSH

[Wag90] Bernhard Wagner. Distributed spooling in a heterogeneous environment. In USENIX [USE90c], pages 449–478.

Waggoner:1992:CCR

[Wag92]  Susan E. Waggoner.  1992

C++ conference report. ;login: the USENIX Association newsletter, 17(6):3–5, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Wang:1997:ZAS

[WAKS97] Daniel C. Wang, Andrew W. Appel, Jeff L. Korn, and Christopher S. Serra. The Zephyr abstract syntax description language. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/wang.html.

Waldo:1991:CCM

[Wal91a] Jim Waldo. Controversy: The case for multiple inheritance in C++. In USENIX [USE91b], pages 157–172.

Walli:1991:RJI

[Wal91b]  Stephen Walli.  Report on

July IEEE POSIX meeting. ;login: the USENIX Association newsletter, 16(5):17–20, September/October 1991. CO-DEN LOGNEM. ISSN 1044-6397.

Walli:1991:UURa

[Wal91c] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(4):18– 38, July/August 1991. CODEN LOGNEM. ISSN 1044-6397.

Walli:1991:UURb

[Wal91d] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(5):27–42, September/October 1991. CO-DEN LOGNEM. ISSN 1044-6397.

Walli:1991:UURc

[Wal91e] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(6):18–??, November/December 1991. CO-DEN LOGNEM. ISSN 1044-

Waldon:1992:BRI

[Wal92a] Tom Waldon. Book review: Inside small talk: Vols I and II. ;login: the USENIX Association newsletter, 17(2):35–??, March/ April 1992. CODEN LOGNEM. ISSN 1044-6397.

Walli:1992:IMR

[Wal92b] Stephen Walli. ISO monitor report. ;login: the USENIX Association newsletter, 17(4):45– 47, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Walli:1992:RE

[Wal92c] Stephen Walli. Report from EurOpen. ;login: the USENIX Association newsletter, 17(1): 13–??, January/February 1992. CODEN LOGNEM. ISSN 1044-6397.

Walli:1992:UURc

[Wal92d] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(3):19– 27, May/June 1992. CODEN LOGNEM. ISSN 1044-6397.

Walli:1992:UURd

[Wal92e] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(4):30– 44, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Walli:1992:UURe

[Wal92f] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(5):32–43, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Walli:1992:UURf

[Wal92g] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(6):35–39, November/December 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Walli:1992:UURa

[Wal92h] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 17(1):17–??, January/February 1992. CO-DEN LOGNEM. ISSN 1044-6397.

Walli:1992:UURb

[Wal92i] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 17(2):38– 46, March/April 1992. CODEN LOGNEM. ISSN 1044-6397.

Waldo:1993:CAM

[Wal93a] Jim Waldo. The case against multiple inheritance in C++. In The Evolution of C++: Language Design in the Marketplace of Ideas [Wal93b], pages 111–120. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993. Editor: Jim Waldo.

Waldo:1993:ECL

[Wal93b] Jim Waldo, editor. The Evolution of C++: Language Design in the Marketplace of Ideas.

USENIX and MIT Press, Berkeley, CA, USA and Cambridge, MA, USA, 1993. ISBN 0-262-73107-X. LCCN QA76.73.C15 E96 1993.

Walli:1993:UURa

[Wal93c] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(1):27–46, January/February 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Walli:1993:UURb

[Wal93d] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(2):36– 45, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Waldo:1994:URC

[Wal94a] Jim Waldo. USENIX retires C++ conference. UNIX review, 12(8):59–??, August 1994. CO-DEN UNRED5. ISSN 0742-3136.

Walli:1994:UAD

[Wal94b] Stephen Walli. UNIX applications development symposium report. ;login: the USENIX Association newsletter, 19(4): 5–8, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Walker:1995:WOS

[Wal95a] Stephen T. Walker. Why are our systems insecure? Must they always be? In USENIX [USE95f], page ?? ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/
library/proceedings/security95/

index.html. Keynote address.

Walli:1995:OSP

[Wal95b]  Stephe Walli.  Open sys-

tems, POSIX, and NT. ;login: the USENIX Association newsletter, 20(6):43–47, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Walli:1995:EYE

[Wal95c] Stephen R. Walli. Everything you’ve ever wanted to know about X/Open UNIX, and the single UNIX specification. ;login: the USENIX Association newsletter, 20(3):35–39, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Walters:1995:THC

[Wal95d] Rex Walters. Tracking hardware configurations in a heterogeneous network with syslogd. In USENIX [USE95h], pages 241–246. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

URL http://www.usenix.org/
publications/library/proceedings/
lisa95/walters.html.

Walli:1997:WTT

[Wal97a] Stephen Walli. Win32 Tcl/Tk GUIs on UNIX apps on Windows NT. ;login: the USE-NIX Association newsletter, 22 (6):??, November 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-11/
guis.html.

Walli:1997:OTU

[Wal97b] Stephen R. Walli. OPENNT^{T M}: UNIX application portability to Windows NT^{T M} via an alternative environment subsystem. In USENIX [USE97f], pages 123–132. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt97/walli.html.

Walli:1998:OUA

[Wal98] Stephen R. Walli. OPENNT?: UNIX(R) application portability to Windows NT via an alternative environment subsystem, 1998. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/usenix98/invited_
                                        talks/walli.ps. Unpublished

invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Wang:1991:SSL

[Wan91] Edward Wang. Staying small in a large installation: Autonomy and reliability (and a cute hack). In USENIX [USE91f], pages 243–244.

Wang:1999:VLB

[WAP99] Randolph Y. Wang, Thomas E. Anderson, and David A. Patterson. Virtual log based file systems for a programmable disk. In USENIX [USE99h], page ?? CODEN OSRED8. ISBN ???? ISSN 0163-5980. LCCN QA76.76.O63 U832 1999; QA1 .A85472 v.32, no.

5. URL http://www.usenix.
org/publications/library/
proceedings/osdi99/wang.html.

Ware:1994:PCD

[War94] Willis H. Ware. Policy consideration for data networks. In USE-NIX [USE94d], pages 1–44.

Wargaski:1999:LNC

[War99a] Christopher J. Wargaski. The little NIC that could, 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/ITs/wargaski.pdf. Un-

published invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Wargaski:1999:NMC

[War99b] Rob Wargaski. Network management on the cheap, 1999.

URL http://www.usenix.org/
publications/library/proceedings/
neta99/invitedtalks.html.

Unpublished invited talk at the First Conference on Network Administration April 7-10, 1999 Santa Clara, California, USA.

Watzman:1996:CAC

[Wat96] Anne Watzman. CMU Andrew Consortium releases C++ version of Andrew UI and WP system. ;login: the USENIX Association newsletter, 21(5):18– ??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Watson:1993:LAM

[WB93] Terri Watson and Brian Bershad. Local area mobile computing on stock hardware and mostly stock software. In USENIX [USE93i], pages 109–

115. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL http://www.usenix.org/
publications/library/proceedings/
mobile93/watson.html.

Woo:1994:SIS

[WBSL94] Thomas Y. C. Woo, Raghuram Bindignavle, Shaowen Su, and Simon S. Lam. SNP: An interface for secure network programming. In USE-NIX [USE94j], pages 45–58. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL http://www.usenix.org/
publications/library/proceedings/
bos94/woo.html.

Wills:1993:CUC

[WCM93] Craig E. Wills, Kirstin Cadwell, and William Marrs. Customization in a UNIX computing environment. In USE-NIX [USE93f], pages 43–

49. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/wills.html.

Whang:1993:RSI

[WCN93] Whan-Kyu Whang, Sharma Chakravarthy, and Shamkant B. Navathe. Relational schema integration: Dealing with interrelational correspondences and querying over component relations. In USENIX [USE93c], pages 319–352.

Wichers:1990:ACL

[WCO^{+}90] David Wichers, Douglas Cook, Ronald Olsson, John Crossley, Paul Kerchen, Karl Levitt, and Raymond Lo. An access control list approach to anti-viral security. In USENIX [USE90i], pages 71–83. LCCN QA 76.9 A25 U55 1990.

Williamson:1999:NLN

[WCR99] David Williamson, Gerald W.

Carter, and Greg Rose. NETA/ LISA-NT/security highlights,

1999. URL http://db.usenix.
org/publications/library/
proceedings/lisa99/practicum/

rose.ppt. Unpublished practicum talk at LISA XIII.

Wood:1993:FSD

[WCS93] David C. M. Wood, Sean S.

Coleman, and Michael F. Schwartz. Fremont: a system for discovering network characteristics and problems. In USENIX [USE93l], pages 335–347. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Winkler:1995:IST

[WD95] Ira S. Winkler and Brian Dealy.

Information security technology? Don’t rely on it. A case study in social engineering. In USENIX [USE95f], pages 1–6 (or 1–5??). ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-1995(5). URL http://

www.usenix.org/publications/

                                        library/proceedings/security95/
                                        winkler.html.

Webber:1993:OSS

[Web93] Neil Webber. Operating system support for portable filesystem extensions. In USE-NIX [USE93l], pages 219–225. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Weissler:1991:BT

[Wei91] Lisa Y. Weissler. Backup without tapes. In USENIX [USE91f], pages 191–194.

Weich:1995:GCD

[Wei95a] Carsten Weich. Generic containers for a distributed object store. In USENIX [USE95k], pages 185–194 (or 185–193??). ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots95/weich.html.

Weiser:1995:UC

[Wei95b] Mark Weiser. Ubiquitous computing, 1995. URL http:

                                        //www.ubiq.com/hypertext/
                                        weiser/Usenix95.html. Un-

published keynote address at the USENIX 1995 Technical Conference Proceedings, January 16–20, 1995, New Orleans, Louisiana.

Welch:1991:FSB

[Wel91a] Brent Welch. The file system belongs in the kernel. In USE-NIX [USE91h], pages 233–250. LCCN QAX 27.

Welch:1991:MPCa

[Wel91b] Brent B. Welch. Measured performance of caching in the Sprite network file system. In USENIX [USE91c], pages 315–342.

Welch:1991:MPCb

[Wel91c] Brent B. Welch. Measured performance of caching in the Sprite network file system. In USENIX [USE91j], pages 229–247. ISBN ???? LCCN ????

Welch:1992:CVS

[Wel92] Brent Welch. A comparison of the Vnode and Sprite File System architectures. In USE-NIX [USE92h], pages 29–44. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Welch:1994:CTD

[Wel94] Brent Welch. A comparison of three distributed file system architectures: Vnode, Sprite, and Plan 9. In USENIX [USE94b], pages 175–199.

Welch:1995:CFE

[Wel95] Brent Welch. Customization and flexibility in the exmh mail user interface. In USENIX [USE95j], pages 261–268 (or 261–267??). ISBN 1-880446-72-3. LCCN QA76.73.T44

T44 1995. URL http://www.
usenix.org/publications/library/
proceedings/tcl95/welch.html.

Wakeman:1995:IRT

[WGC^{+}95] Ian Wakeman, Atanu Ghosh, Jon Crowcroft, Van Jacobson, and Sally Floyd. Implementing real time packet forwarding policies using streams. In USENIX [USE95e], pages 71–82. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/wakeman.html.

Wagner:1992:ASP

[WH92] Bernard Wagner and Bruce K. Haddon. Application software: Project management and privileges. In USENIX [USE92j], pages 197–208.

Wansbrough:1997:MMA

[WH97] Keith Wansbrough and John Hamer. A modular monadic action semantics. In USENIX [USE97b], page ?? ISBN 1-880446-89-8. LCCN QA76.7 .C663 1997. URL http://www.

usenix.org/publications/library/
proceedings/dsl97/wansbrough.
html.

Whitted:19xx:ISP

[Whixx] Turner Whitted. Image synthesis on personal computers. In ???? USENIX, Berkeley, CA, USA, 19xx. Title listed in contents only, no paper, no abstract.

Wiecek:1992:MPV

[Wie92] Cheryl A. Wiecek. A model and prototype of VMS using the Mach 3.0 kernel. In USENIX [USE92i], pages 187–204. ISBN 1-880446-42-1. LCCN QAX 32.

Walpole:1992:MIM

[WIK92] Jonathan Walpole, Jon Inouye, and Ravindranath Konuru. Modularity and interfaces in microkernel design and implementation: a case study of Chorus on the HP PA-Risc. In USENIX [USE92i], pages 71–82. ISBN 1-880446-42-1. LCCN QAX 32.

Williams:1991:REa

[Wil91a] Alain Williams. Report from EurOpen. ;login: the USENIX Association newsletter, 16(1): 12–??, January/February 1991. CODEN LOGNEM. ISSN 1044-6397.

Williams:1991:REb

[Wil91b] Alain Williams. Report from EurOpen. ;login: the USE-NIX Association newsletter, 16 (5):24–??, September/October 1991. CODEN LOGNEM. ISSN 1044-6397.

Wilkes:1992:DRP

[Wil92a] John Wilkes. DataMesh research project, phase 1. In USENIX [USE92h], pages 63– 70. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Williams:1992:RE

[Wil92b] Alain Williams. Report from EurOpen. ;login: the USE-NIX Association newsletter, 17 (3):15–??, May/June 1992. CO-DEN LOGNEM. ISSN 1044-

Willison:1993:TIB

[Wil93] Pace Willison. The TCP/IP bakeoff. ;login: the USENIX Association newsletter, 18(3):20– ??, May/June 1993. CODEN LOGNEM. ISSN 1044-6397.

Wilson:1994:FFT

[Wil94] Pat Wilson. Factoids: Fun things to tell and know. ;login: the USENIX Association newsletter, 19(1):21–??, January/February 1994. CODEN LOGNEM. ISSN 1044-6397.

Wilson:1995:QS

[Wil95a] Pat Wilson. 10 questions for sysadmins. ;login: the USE-NIX Association newsletter, 20 (3):27–28, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Wilson:1995:CNEb

[Wil95b] Pat Wilson. Call for nominations for election to SAGE. ;login: the USENIX Association newsletter, 20(5):22–??, October 1995. CODEN LOGNEM. ISSN 1044-6397.

Wilson:1995:CNEa

[Wil95c] Pat Wilson. Call for nominations for election to SAGE board of directors. ;login: the USE-NIX Association newsletter, 20 (4):13–??, August 1995. CO-DEN LOGNEM. ISSN 1044-6397.

Williams:1997:EPC

[Wil97a]  Stephen Williams.  Embed-

ded programming with C++. In USENIX [USE97g], pages 131–143. ISBN 1-880446-86-

3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/williams.html. See

also [LH97, Rie97, WW97].

Wilson:1997:TPL

[Wil97b] Pat Wilson. A thousand pints of lite. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Wilson:1998:SAR

[Wil98] Pat Wilson. San Antonio revisited: Five years of SAGE. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CO-DEN LOGNEM. ISSN 1044-6397.

Willoughby:1999:OCM

[Wil99] Steve Willoughby. Organizing the chaos: Managing request tickets in a large environment. In USENIX [USE99i], page ?? ISBN 1-880446-25-1. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa99/willoughby.
html.

Winckler:1994:DLA

[Win94a]  Andreas Winckler.  A dis-

tributed look-ahead workload assignment algorithm for interdependent tasks. In USENIX [USE94c], pages 361–391.

Winterbottom:1994:ADB

[Win94b] Phil Winterbottom. ACID: a debugger built from a language.

In USENIX [USE94l], pages 211–222. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/index.html.

Winkler:1996:NTT

[Win96] Ira S. Winkler. The nontechnical threat to computing systems. In USENIX [USE96d], pages 3–14.

Wiz:1996:RAD

[Wiz96] Marc Wiz. Response about DCE. ;login: the USENIX Association newsletter, 21(2):4–??, April 1996. CODEN LOGNEM. ISSN 1044-6397.

Wittle:1993:LNG

[WK93] M. Wittle and B. E. Keith. LADDIS: The next generation in NFS file server benchmarking. In USENIX [USE93g], pages 111–128. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/.

Wilson:1998:RNC

[WKR98] Pat Wilson, Amy Kreiling, and Greg Rose. Report of the nominating committee. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
comingup.html.

Wetherall:1995:ETD

[WL95] David Wetherall and Christopher J. Lindblad. Extending Tcl for dynamic objectoriented programming. In USE-NIX [USE95j], pages 173–182 (or 173–181??). ISBN 1-880446-72-3. LCCN QA76.73.T44

T44 1995. URL http://www.
usenix.org/publications/library/
proceedings/tcl95/wetherall.
html.

Wang:1998:CCE

[WL98] Yi-Min Wang and Woei-Jyh Lee. COMERA: COM extensible remoting architecture. In USENIX [USE98c], page ?? ISBN ???? LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
coots98/wang.html.

Wu:1992:AFA

[WM92] Sun Wu and Udi Manber. agrep — A fast approximate patternmatching tool. In USENIX [USE92j], pages 153–162.

Wu:1999:BIT

[WMB99] Thomas Wu, Michael Malkin, and Dan Boneh. Building intrusion-tolerant applications. In USENIX [USE99e], page ?? ISBN 1-880446-28-

6. LCCN QA76.9.A25 U83

1999. URL http://db.usenix.
org/publications/library/
proceedings/sec99/wu.html.

Watson:1992:LMV

[WN92] Andy Watson and Bruce Nelson. LADDIS: a multi-vendor and vendor-neutral SPEC NFS benchmark. In USENIX [USE92e], pages 17–32. ISBN 1-880446-47-2. LCCN ????

Woodhull:1998:MAC

[WN98] Gordon Woodhull and Stephen C. North. Montage — an ActiveX container for dynamic interfaces. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt98/woodhull.html.

Wong:1990:SSU

[Won90] Raymond Wong. A survey of secure UNIX operating systems. In USENIX [USE90i], pages 57–

60. LCCN QA 76.9 A25 U55

1990.

Wong:1993:LDD

[Won93]  Walter C. Wong.  Local

disk depot — customizing the software environment. In USENIX [USE93f], pages 51–

55. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa93/wong.html.

Woolley:1994:PEL

[Woo94] David R. Woolley. PLATO: The emergence of on-line community. ;login: the USENIX Association newsletter, 19(2):20–27, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Woodard:1998:BEP

[Woo98]  Ben Woodard.  Building

an enterprise printing system. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa98/woodard.html.

Willison:1994:SCR

[WP94] Pace Willison and Jerry Peek.

Summer ’94 conference reports. ;login: the USENIX Association newsletter, 19(4):8–14, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Wang:1998:EIU

[WR98] Zheng Wang and Norm Rubin. Evaluating the importance of user-specific profiling. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/wang.html;   http:
//www.usenix.org/publications/
library/proceedings/usenix-
nt98/wang_slides.pdf.

Wright:1996:LSP

[Wri96] Benjamin Wright. Legal signatures and proof in electronic commerce. In USE-NIX [USE96h], pages 67–

75. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996.

URL http://www.usenix.org/
publications/library/proceedings/
ec96/index.html.

Wieland:1992:EPP

[WRJ92] Frederick Wieland, Peter Reiher, and David Jefferson. Experience in parallel performance measurement: The speedup bias. In USENIX [USE92l], page ?? ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Wollrath:1996:DOMa

[WRW96a] Ann Wollrath, Roger Riggs, and Jim Waldo. A distributed object model for the Java system. In USENIX [USE96a], pages 265–

                                        289. URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/coots96/wollrath.
                                        html.

Wollrath:1996:DOMb

[WRW96b] Ann Wollrath, Roger Riggs, and Jim Waldo. A distributed object model for the Java system. In USENIX [USE96g], pages 219–231. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots96/wollrath.html.

Willison:1993:MCS

[WS93] Pace Willison and Peter H. Salus. Mobile computing symposium reports. ;login: the USE-NIX Association newsletter, 18(5):3–8, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Wagner:1996:ASP

[WS96] David Wagner and Bruce Schneier. Analysis of the SSL 3.0 protocol. In USE-NIX [USE96h], pages 29–40. ISBN 1-880446-83-9. LCCN HF5004 .U74 1996. URL http:

//www.counterpane.com/ssl.
html;   http://www.usenix.
org/publications/library/
proceedings/ec96/wagner.html.

Warner:1999:NDW

[WS99] Wade Warner and Rajshekhar Sunderraman. Network documentation: a Web-based relational database approach. In USENIX [USE99g], page ?? ISBN 1-880446-36-7. LCCN TK5105.5 .C664 1999. URL

http://www.usenix.org/publications/
library/proceedings/neta99/
warner.html.

Walker:1996:CRP

[WSB^{+}96] Kenneth M. Walker, Daniel F.

Sterne, M. Lee Badger, Michael J. Petkac, David L. Sherman, and Karen A. Oostendorp. Confining root programs with domain and type enforcement. In USENIX [USE96i], pages 21–36. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/walker.html.

Wallace:1993:WCR

[WSR93] John J. Wallace, Peg Schafer, and Christopher Rath. Winter 1993 Conference Reports. ;login: the USENIX Association newsletter, 18(2):9–17, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Wilkinson:1992:IDT

[WSS92] T. Wilkinson, T. Stiemerling, and A. Saulsbury. Implementing DVSM on the TOPSY multicomputer. In USENIX [USE92l], pages 263–278. ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Whitten:1999:WJC

[WT99] Alma Whitten and J. D. Tygar. Why Johnny can’t encrypt: a usability evaluation of PGP 5.0. In USENIX [USE99e], pages 169–184. ISBN 1-880446-28-6. LCCN QA76.9.A25 U83

                                        1999. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/sec99/whitten.
                                        html;   http://www.cs.berkeley.
                                        edu/~tygar/papers/Why_Johnny_
                                        Cant_Encrypt/OReilly.pdf;
                                         http://www2.cs.cmu.edu/~
                                        alma/johnny.pdf. Reprinted in

[CG05, pp. 679–702].

Welch:1996:TTH

[WU96] Brent Welch and Steve Uhler. Tcl/Tk HTML tools. In USENIX [USE96e], pages 173–182. ISBN 1-880446-78-2. LCCN QA 76.73 T44 T35 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        tcl96/welch.html.

Welch:1997:WEA

[WU97] Brent Welch and Stephen Uhler. Web enabling applications. In USENIX [USE97d], pages 189–190. ISBN 1-880446-87-1. LCCN QA76.73.T44 T35 1997.

URL http://www.usenix.org/
publications/library/proceedings/
tcl97/poster_welch.html.

Wilson:1990:GXG

[WUB^{+}90] Matt Wilson, John Uhley, Upinder Bhalla, David Bilitch, Mark Nelson, and James Bower. GENESIS and XODUS, general purpose neural network simulation tool. In USENIX [USE90h], pages 75–88. LCCN QA76.8.U65 U82 1990.

Williams:1992:USC

[WUS^{+}92] David N. Williams, Stephen Uhler, Vivek Singhal, Sheetal V. Kakkad, Vladimir Bacvanski, J. G. Posthuma, J. Scholten, and J. G. Wijnstra. 1992 USE-NIX summer conference worksin-progress report. ;login: the USENIX Association newsletter, 17(4):9–11, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Wolman:1999:OBA

[WVS^{+}99] Alec Wolman, Geoff Voelker, Nitin Sharma, Neal Cardwell, Molly Brown, Tashana Landray, Denise Pinnel, Anna Karlin, and Henry Levy. Organizationbased analysis of Web-object sharing and caching. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96. URL

http://www.usenix.org/publications/
library/proceedings/usits99/
wolman.html.

Wolman:1994:LAT

[WVT94] Alec Wolman, Geoff Voelker, and Chandramohan A. Thekkath. Latency analysis of TCP on an ATM network. In USENIX [USE94l], pages 167–179. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sf94/wolman.html.

Witten:1991:NOC

[WW91] Matthew Witten and Robert E. Wyatt. Neural orchestration: From cortical simulation to cortical symphony. In USE-NIX [USE91g], pages 107–124. LCCN QA 76.76 O63 U83 1991.

Waldspurger:1994:LSF

[WW94] Carl A. Waldspurger and William E. Weihl. Lottery scheduling: Flexible proportionalshare resource management. In USENIX [USE94i], pages 1–11. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/waldspurger.html.

Waldo:1997:DCJ

[WW97] Jim Waldo and Ann Wollrath. Distributed computing with Java remote method invocation. Tutorial session at [USE97g].,

                                        1997. URL http://dxsting.
                                        cern.ch/sting/COOTS97.html.

Workman:1998:DOE

[WWDF98] Kevin Workman, Earl Waud, Steven Downs, and Mikel Featherstone. Designing an optimized enterprise Windows NT/ 95/98 client backup solution with intelligent data management. In USENIX [USE98e], pages 69–73. ISBN 1-880446-96-

0. LCCN TK5105.5 .L37 1998.

URL http://www.usenix.org/
publications/library/proceedings/
lisa-nt98/workman.html.

Weiser:1994:SRC

[WWDS94] Mark Weiser, Brent Welch, Alan Demers, and Scott Shenker. Scheduling for reduced CPU energy. In USENIX [USE94i], pages 13–23. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994. URL http://www.

usenix.org/publications/library/
proceedings/osdi/welch.html.

Wollrath:1995:SAD

[WWW95] Ann Wollrath, Geoff Wyant, and Jim Waldo. Simple activation for distributed objects. In USENIX [USE95k], pages 1–12 (or 1–11??). ISBN 1-880446-71-

5. LCCN QA76.64 .U84 1995.

URL http://www.usenix.org/
publications/library/proceedings/
coots95/wollrath.html.

Waldo:1995:ERB

[WWWK95] Jim Waldo, Ann Wollrath, Geoff Wyant, and Samuel C. Kendall. Events in an RPC based distributed system. In USENIX [USE95e], pages 131–

142. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL http://www.usenix.org/
publications/library/proceedings/
neworl/waldo.html.

Wybolt:1990:ECO

[Wyb90] Nicholas Wybolt. Experiences with C++ and object-oriented software development. In USE-NIX [USE90j], pages 1–9.

Wada:1993:MCE

[WYOT93] Hiromi Wada, Takashi Yozawa, Tatsuya Ohnishi, and Yasunori Tanaka. Mobile computing environment based on Internet packet forwarding. In USENIX [USE93l], pages 503–

517. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sd93/.

Yin:1999:HCC

[YADL99] Jian Yin, Lorenzo Alvisi, Mike Dahlin, and Calvin Lin. Hierarchical cache consistency in a WAN. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usits99/yin.html.

Yahalom:1994:STC

[Yah94] Raphael Yahalom. Secure timeliness: On the cost of nonsynchronized clocks. In USENIX [USE94a], pages 451–465.

Yarvin:1993:ARL

[YBA93] Curtis Yarvin, Richard Bukowski, and Thomas Anderson. Anonymous RPC: Lowlatency protection in a 64-Bit address space. In USE-NIX [USE93g], pages 175–186.

ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/yarvin.html.

Yuhara:1994:EPD

[YBMM94] Masanobu Yuhara, Brian N.

Bershad, Chris Maeda, and J. Eliot B. Moss. Efficient packet demultiplexing for multiple endpoints and large messages. In USENIX [USE94l], pages 153–

165. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

URL http://www.usenix.org/
publications/library/proceedings/
sf94/yuhara.html.

Yoshikawa:1997:USC

[YCE^{+}97] Chad Yoshikawa, Brent Chun, Paul Eastham, Amin Vahdat, Thomas Anderson, and David Culler. Using smart clients to build scalable services. In USENIX [USE97e], pages 105–

117. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

URL http://www.usenix.org/
publications/library/proceedings/
ana97/summaries/yoshikawa.
html;   http://www.usenix.
org/publications/library/
proceedings/ana97/yoshikawa.
html.

Yang:1994:CTD

[YCO94] Cheng-Zen Yang, Chih-Chung Chen, and Yen-Jen Oyang. Clue tables: a distributed, dynamicbinding naming mechanism. In USENIX [USE94j], pages 153–160. ISBN 1-880446-62-6.

LCCN QA 76.76 O63 U83 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        bos94/yang.html.

Yang:1994:IGD

[YD94] Cui-Qing Yang and Shrinand Desai. Implementing a generalized drag-and-drop in X. In USENIX [USE94g], pages 133–145. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        appdev94/yang.html.

Yan:1995:STW

[YGM95] Tak W. Yan and Hector Garcia-Molina. SIFT — a tool for widearea information dissemination. In USENIX [USE95e], pages 177–186. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        neworl/yan.html.

Yeom:1996:IMT

[YHK96] Heon Y. Yeom, Jungsoo Ha, and Ilhwan Kim. IP multiplexing by transparent portaddress translator. In USE-NIX [USE96j], pages 113–

121. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        lisa96/yeom.html.

Yigit:1997:REA

[Yig97] Ozan S. Yigit. Review: Effective awk programming. ;login: the USENIX Association newsletter, 22(2):75–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Yahalom:1994:TBN

[YKB94] Raphael Yahalom, Birgit Klein, and Thomas Beth. Trust-based navigation in distributed systems. In USENIX [USE94d], pages 45–73.

Yang:1998:DIA

[YL98] C. S. Yang and M. Y. Luo.

Design and implementation of an administration system for distributed Web server. In USENIX [USE98g], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
yang.html.

Yang:1999:ESC

[YL99] Chu-Sing Yang and Mon-Yen Luo. Efficient support for content-based routing in Web server clusters. In USENIX [USE99c], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57 U96. URL

http://www.usenix.org/publications/
library/proceedings/usits99/
yang.html.

Yavatkar:1993:DIM

[YLK93] Raj Yavatkar, K. Lakshman, and Giri Kuthethoor. Design and implementation of a multimedia protocol suite in a BSD UNIX kernel. In USE-NIX [USE93g], pages 129–146.

ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        cinci93/lakshman.html.

Ylonen:1996:SSL

[Ylo96] Tatu Ylonen. SSH — secure login connections over the Internet. In USENIX [USE96i], pages 37–42. ISBN 1-880446-79-

0. LCCN QA76.9.A25 U83 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        sec96/ylonen.html.

Yu:1999:BID

[YM99] Jay Yu and Bryan McDonald. Building Internet datacenters,

                                        1999. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa99/invitedtalks.

html. Unpublished invited talk

at LISA XIII.

York:1999:BLC

[Yor99] Dan York. Building a Linux certification program. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1999-6/features/linuxcert.
                                        html.

Young:1993:BMSa

[You93a] Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 18(1):6– 9, January/February 1993. CO-DEN LOGNEM. ISSN 1044-6397.

Young:1993:BMSb

[You93b] Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 18(2):4– 5, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Young:1993:BMSe

[You93c] Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 18(5): 8–9, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Young:1993:BMSf

[You93d] Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 18(6):8– 9, November/December 1993. CODEN LOGNEM. ISSN 1044-6397.

Young:1994:BMSb

[You94a] Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 19(2): 10–11, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Young:1994:BMSc

[You94b] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 19 (3):10–11, May/June 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Young:1994:BMSd

[You94c] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 19 (4):14–15, August 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Young:1994:UBMf

[You94d] Ellie Young. USENIX board meeting summary. ;login: the USENIX Association newsletter, 19(6):13–15, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Young:1995:MD

[You95a] Ellie Young. Member dues. ;login: the USENIX Association newsletter, 20(3):8–10, June 1995. CODEN LOGNEM. ISSN 1044-6397.

Young:1995:UBMa

[You95b] Ellie Young. USENIX board meeting summary. ;login: the USENIX Association newsletter, 20(2):25–26, April 1995. CODEN LOGNEM. ISSN 1044-6397.

Young:1995:UBMb

[You95c] Ellie Young. USENIX board meeting summary. ;login: the USENIX Association newsletter, 20(6):23–25, December 1995. CODEN LOGNEM. ISSN 1044-6397.

Young:1996:BMSa

[You96a] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 21 (4):12–??, August 1996. CO-DEN LOGNEM. ISSN 1044-6397.

Young:1996:BMSb

[You96b] Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 21(6):3– 4, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Young:1996:MD

[You96c] Ellie Young. Member dues.

;login: the USENIX Association newsletter, 21(3):7–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Young:1997:EBD

[You97a]  Ellie Young.  1998 elec-

tion for Board of Directors. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
imho.html.

Young:1997:UNC

[You97b] Ellie Young. 1998 USENIX Nominating Committee. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397.

Young:1997:BMSb

[You97c] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 22 (2):2–??, April 1997. CODEN LOGNEM. ISSN 1044-6397.

Young:1997:BMSd

[You97d] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 22 (4):4–??, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Young:1997:MD

[You97e] Ellie Young. Member dues. ;login: the USENIX Association newsletter, 22(5):3–6, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Young:1998:EBD

[You98a] Ellie Young. 1998 election for Board of Directors. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

Young:1998:BMS

[You98b] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
reference.html.

Young:1998:SUB

[You98c] Ellie Young. Summary of USE-NIX Board meetings. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/mailing.html.

Young:1998:TOV

[You98d] Ellie Young. Thanks to our volunteers. ;login: the USE-NIX Association newsletter, 23 (1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-2/
reference.html.

Young:1999:EBD

[You99a]  Ellie Young.  2000 elec-

tion for Board of Directors. ;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-12/
election.html.

Young:1999:BMSa

[You99b] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-2/
meeting.html.

Young:1999:BMSb

[You99c] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 24 (3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-6/
boardmtg.html.

Young:1999:BMSc

[You99d] Ellie Young. Board meeting summary. ;login: the USE-NIX Association newsletter, 24 (4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
boardmtg.html.

Young:1999:TOV

[You99e] Ellie Young. Thanks to our volunteers. ;login: the USE-NIX Association newsletter, 24 (1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.org/
                                        publications/login/1999-2/
                                        thanks_volunteers.html.

Yoo:1993:UKS

[YR93] Hyuck Yoo and Tom Rogers. UNIX kernel support for OLTP performance. In USENIX [USE93l], pages 241–247. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993. URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/sd93/.

Yee:1995:SCE

[YT95] Bennett Yee and J. D. Tygar. Secure coprocessors in electronic commerce applications. In USENIX [USE95g], pages 155–170. ISBN 1-880446-74-X. LCCN HF5548.33. U84 1995(1).

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        ec95/yee.html.

Young:1991:MAD

[YTJ91] Michael Wayne Young, Dean S. Thompson, and Elliot Jaffe. A modular architecture for distributed transaction processing. In USENIX [USE91i], pages 357–363. LCCN QA 76.76 O63 U84 1992.

Zhang:1999:HAC

[ZBCS99] Xiaolan Zhang, Michael Barrientos, J. Bradley Chen, and Margo Seltzer. HACC: An architecture for cluster-based Web servers. In USENIX [USE99d], page ?? ISBN 1-880446-29-4. LCCN ????

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt99/zhang.html.

Zadok:1993:DHR

[ZD93a] Erez Zadok and Dan Duchamp. Discovery and hot replacement of replicated readonly file systems, with application to mobile computing. In USENIX [USE93g], pages 69–85. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL http://www.usenix.org/
publications/library/proceedings/
cinci93/zadok.html.

Zadok:1993:HDE

[ZD93b] Erez Zadok and Alexander Dupuy. HLFSD: Delivering email to your $HOME. In USENIX [USE93f], pages 243–

254. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL http://www.usenix.org/
publications/library/proceedings/
lisa93/zadok.html.

Zeadally:1997:DHQ

[Zea97] Sherali Zeadally. Delivery of high quality uncompressed video over ATM to Windows NT desktop. In USENIX [USE97f], pages 67–80. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt97/zeadally.html.

Zielinski:1993:SVK

[Zie93] Stephan Zielinski. SAGE views: Know your UNIX system administrator — A field guide. ;login: the USENIX Association newsletter, 18(2):25–27, March/ April 1993. CODEN LOGNEM. ISSN 1044-6397.

Zhou:1996:PET

[ZIL96] Yuanyuan Zhou, Liviu Iftode, and Kai Li. Performance evaluation of two home-based lazy release consistency protocols for shared virtual memory systems. In USENIX [USE96f], pages 75–88. ISBN 1-880446-82-0. LCCN QA76.76.O63 S96 1996.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi96/zhou.html.

Zweig:1990:CCA

[ZJ90] Jonathan M. Zweig and Ralph E. Johnson. The conduit: a communication abstraction in C++. In USENIX [USE90j], pages 191–203 (or 191–204??).

Zwicky:1994:SSA

[ZK94] Elizabeth Zwicky and Zanna Knight. SAGE 1994 system administrator profile and salary survey. ;login: the USENIX Association newsletter, 19(6):16– 19, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Zerkle:1996:NMH

[ZL96] Dan Zerkle and Karl Levitt. NetKuang — A multi-host configuration vulnerability checker. In USENIX [USE96i], pages 195–201. ISBN 1-880446-79-0. LCCN QA76.9.A25 U83 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sec96/zerkle.html.

Zukowski:1997:MTC

[ZPMD97] Deborra J. Zukowski, Apratim Purakayastha, Ajay Mohindra, and Murthy Devarakonda. Metis: a thin-client application framework. In USE-NIX [USE97g], pages 103–

114. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.usenix.org/
publications/library/proceedings/
coots97/zukowski.html. See

also [LH97, Rie97, WW97].

Zajcew:1993:OUM

[ZRB^{+}93] Roman Zajcew, Paul Roy, David Black, Chris Peak, Paulo Guedes, Bradford Kemp, John LoVerso, Michael Leibensperger, Michael Barnett, Faramarz Rabii, and Durriya Netterwala. An OSF/1 UNIX for massively parallel multicomputers. In USENIX [USE93l], pages 449–468. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

URL http://www.usenix.org/
publications/library/proceedings/
sd93/.

Zhou:1999:ASH

[ZS99] Dong Zhou and Karsten Schwan.

Adaptation and specialization for high performance mobile agents. In USENIX [USE99f], page ?? ISBN ???? LCCN QA76.64 .U84 1999.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        coots99/zhou.html.

Zekauskas:1994:SWD

[ZSB94] Matthew J. Zekauskas, Wayne A. Sawdon, and Brian N. Bershad. Software write detection for distributed shared memory. In USENIX [USE94i], pages 87–100. ISBN 1-880446-66-9. LCCN QA76.76.O63 U832 1994.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        osdi/zekauskas.html.

Zhang:1998:RAC

[ZSBC98] Xiaolan Zhang, Ravi Shanmugan, Michael Barreintos, and J. Bradley Chen. RACC: An approach to cluster-based Web servers. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

                                        URL http://www.usenix.org/
                                        publications/library/proceedings/
                                        usenix-nt98/full_papers/poster_
                                        zhang/zhang.pdf;   http://
                                        www.usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/full_papers/poster_zhang/
                                        zhang_html/zhang.html.

Zwicky:1990:PSA

[ZSD90] Elizabeth D. Zwicky, Steve Simmons, and Ron Dalton. Policy as a system administration tool. In USENIX [USE90f], pages 115–

124. LCCN QA76.76.O63 L37

1990.

Zhou:1992:DSR

[ZSG92] H. Zhou, K. Schwan, and A. Gheith. Dynamic synchronization of real-time threads for multiprocessor systems. In USE-NIX [USE92l], page ?? ISBN 1-880446-41-3. LCCN QA76.9.D3 S954 1992.

Zeisset:1996:NAD

[ZTM96] Stephan Zeisset, Stefan Tritscher, and Martin Mairandres. A new approach to distributed memory management in the Mach microkernel. In USENIX [USE96k], pages 205–217 (or 205–218??). ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

URL http://www.usenix.org/
publications/library/proceedings/
sd96/zeisset.html.

Zwicky:1991:EYA

[Zwi91a] Elizabeth D. Zwicky. Enhancing your apparent psychic abilities through software. In USE-NIX [USE91f], pages 171–172.

Zwicky:1991:TTB

[Zwi91b] Elizabeth D. Zwicky. Torturetesting backup and archive programs: Things you ought to know but probably would rather not. In USENIX [USE91f], pages 181–190.

Zwicky:1992:NFU

[Zwi92a] Elizabeth Zwicky. Name frequency on Usenet. ;login: the USENIX Association newsletter, 17(4):26–29, July/August 1992. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1992:SBR

[Zwi92b] Elizabeth Zwicky. SAGE book reviews: UNIX for super-users.

;login: the USENIX Association newsletter, 17(6):13–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1992:SVC

[Zwi92c] Elizabeth Zwicky. SAGE views: The customer isn’t always right; the customer isn’t always even a customer. ;login: the USE-NIX Association newsletter, 17 (6):11–??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1992:TBC

[Zwi92d] Elizabeth D. Zwicky. Typecast: Beyond cloned hosts. In USENIX [USE92e], pages 73– 78. ISBN 1-880446-47-2. LCCN ????

Zwicky:1993:SR

[Zwi93a] Elizabeth Zwicky. SAGE review. ;login: the USENIX Association newsletter, 18(4):23– ??, July/August 1993. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1993:SVA

[Zwi93b] Elizabeth Zwicky. SAGE views: Analogies to live without: You don’t need a manual to drive a car. ;login: the USENIX Association newsletter, 18(2):27, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1993:SVS

[Zwi93c] Elizabeth Zwicky. SAGE views: System administration tools your vendor never told you about: The Laminator. ;login: the USENIX Association newsletter, 18(2):28–??, March/April 1993. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1993:SSD

[Zwi93d] Elizabeth D. Zwicky. SAGE:

Solaris developers conference report. ;login: the USENIX Association newsletter, 18(3):12– 13, May/June 1993. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1993:SAT

[Zwi93e]  Elizabeth D. Zwicky.  Sys-

tem administration tools. ;login: the USENIX Association newsletter, 18(5):23–??, September/October 1993. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:AP

[Zwi94a] Elizabeth Zwicky. Attitude problems. ;login: the USE-NIX Association newsletter, 19 (3):10–??, May/June 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:LSP

[Zwi94b] Elizabeth Zwicky. Letter from the SAGE president. ;login: the USENIX Association newsletter, 19(4):17–??, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:SVa

[Zwi94c] Elizabeth Zwicky. SAGE views.

;login: the USENIX Association newsletter, 19(1):20–21, January/February 1994. CO-DEN LOGNEM. ISSN 1044-

Zwicky:1994:SVb

[Zwi94d] Elizabeth Zwicky. SAGE views. ;login: the USENIX Association newsletter, 19(2):20–21, March/ April 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:SAM

[Zwi94e] Elizabeth Zwicky. System administration models. ;login: the USENIX Association newsletter, 19(5):29–31, October 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:SATb

[Zwi94f] Elizabeth Zwicky. System administration tools your vendor never told you about: The beeper. ;login: the USENIX Association newsletter, 19(4):21– ??, August 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:SATa

[Zwi94g] Elizabeth Zwicky. System administration tools your vendor never told you about: the door. ;login: the USENIX Association newsletter, 19(3):13–??, May/ June 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:SATc

[Zwi94h] Elizabeth Zwicky. System administration tools your vendor never told you about: The toy. ;login: the USENIX Association newsletter, 19(6):24–??, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:WYG

[Zwi94i] Elizabeth Zwicky. Wherever you go, there you are. ;login: the USENIX Association newsletter, 19(6):23–24, December 1994. CODEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:WMP

[Zwi94j] Elizabeth Zwicky. Who are my peers? ;login: the USE-NIX Association newsletter, 19 (4):22–??, August 1994. CO-DEN LOGNEM. ISSN 1044-6397.

Zwicky:1994:GMW

[Zwi94k] Elizabeth D. Zwicky. Getting more work out of work tracking systems. In USE-NIX [USE94h], pages 105–

110. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

URL http://www.usenix.org/
publications/library/proceedings/
lisa94/zwicky.html.

Zwicky:1997:WPS

[Zwi97] Elizabeth Zwicky. Why I am not A professional system administrator. ;login: the USE-NIX Association newsletter, 22 (7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1997-12/
leashes.html.

Zwicky:1998:ESD

[Zwi98]  Elizabeth Zwicky.  Enough

SNMP to be dangerous, part

1. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1998-12/
snmp.html.

Zwicky:1999:ESDa

[Zwi99a]  Elizabeth Zwicky.  Enough

SNMP to be dangerous — part 2. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-8/
features/snmp2.html.

Zwicky:1999:ESDb

[Zwi99b]  Elizabeth Zwicky.  Enough

SNMP to be dangerous, part

3. ;login: the USENIX Association newsletter, 24(5): ??, October 1999. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.org/
publications/login/1999-10/
features/snmp_p3.html.

Zwicky:1999:NFE

[Zwi99c]  Elizabeth Zwicky.  NT in

the firewall environment, 1999.

URL http://db.usenix.org/
publications/library/proceedings/
lisa-nt99/invitedtalks.html.

Unpublished invited talk at the 2nd Large Installation System Administration of Windows NT Conference, Seattle, Washington, USA, July 14–17, 1999.

Zurko:1999:JEI

[ZWM^{+}99] Mary Ellen Zurko, John Wray, Ian Morrison, Mike Shanzer, Mike Crane, Pat Booth, Ellen McDermott, Warren Macek, Ann Graham, Jim Wade, and Tom Sandlin. Jonah: Experience implementing PKIX reference freeware. In USENIX [USE99e], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL http://db.

usenix.org/publications/library/
proceedings/sec99/zurko.html.

Zabatta:1998:TPC

[ZY98] Fabian Zabatta and Kevin Ying.

A thread performance comparison: Windows NT and Solaris on a symmetric multiprocessor. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998.

URL http://www.usenix.org/
publications/library/proceedings/
usenix-nt98/zabatta.html;
 http://www.usenix.org/publications/
library/proceedings/usenix-
nt98/zabatta_slides.pdf.